A Complete Bibliography of Publications in Computing in Science and Engineering

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

16 March 2024 Version 1.93

Title word cross-reference

+ [Luo12]. 2 [CFA04, Hun07, OMKdSB11, YWC02]. 3 [Ama00, Ano15-38, Ano15-39, BB07, CCSS08, CY00, CHC^{+}11, CS14, CS15, DiP18b, GWA^{+}07, HMA00, HL00, LWF10, MJAK09, RV11, SDS00, SYP08, Sul02c, Weg00, XHL^{+}13, YWC02, YCK03, YCKK03, YaL10, ZZPC06, ZDW^{+}07, ZCXM99]. 4 [WNZ^{+}17]. A = B [BS00a]. D [UZC^{+}12]. k [ALM19]. B NEW ⇐ B_{0} ⊕ (B_{1} ∨ B_{2}) [Sul02b]. N [YB12]. ≠ [Sul04c]. QR [Par00a].

-body [YB12]. -Vector [ALM19].

1 [Bak10]. 11th [HPKS04]. 19 [AM20, BPMKC21, BCH^{+}22, Com20, FGHW20, HP20, Jef21, KKO^{+}20, PHW^{+}21, PQQ20, TGU21, VSB^{+}21, WHG21, Wes21].

2 [CW05b, FGG^{+}22, XDK^{+}20]. 20-Bit [TJ14]. 20-Year [BLT^{+}23]. 2000 [MFD^{+}09]. 2009 [SEPC10]. 2010 [SSW11]. 2011 [Ano11c]. 2013 [Thi13a]. 2014 [Ano13j, Ano14-31]. 2016 [Ano16b, CSW17, DE17]. 2017 [Ano17a]. 2019 [Ano18b, Ano19d, Ano19e, Ano19z, WBP^{+}19]. 2020 [Ano20x, Ano20y, Ano20-37, Ano20-54, Ano20-56]. 2021 [HCK22]. 2023 [Ano22w, Ano22a, Ano22b, Ano22-68, RHC^{+}23]. 20th [SCBT18]. 21st [Ano22-67, Ano23-47, Ano23-48, Ano23-49, Ano14-49, Ano22-66, Ric99]. 2Guest [BLMR21]. 2nd [Seg99].

3 [KBLD08, SGA^{+}22]. 3-D [SGA^{+}22]. 39th [Ano15b]. 3D [Lew99b]. 3M [CW05d]. 3Ms [CW05b, CW05c, CW05a].

43 [UZC^{+}12]. 4th [CHC17].

6000 [Tou02b]. 6i [Shi01a].

7 [Sim13].

8 [Bur99].

9 [MKM^{+}14]. 95 [GRE99].

Abinit [PBD^{+}11]. Abridged [FB04]. Academic [Bot16, LTD11]. Accelerate [WCH12]. Accelerated [Ben04, FPRK16, NGGS22, PHL^{+}10, SWDR22]. Accelerating [DdS23a, DdS23b, EVA^{+}21, EKCS12, KSB07, LFK^{+}19, Men18, MSM13, SRKS22, Thi12a, TCD^{+}14, UZC^{+}12, VMK20, WZS^{+}10, WOAEAG10]. Acceleration [AAAH^{+}16, FKB^{+}13, NLGNJ13, SKP^{+}10, STH22, SMSC22]. Accelerator [AMS14, DLB^{+}07, SG10]. Accelerators [Eis17, HGV^{+}08, KWB^{+}10]. Acceptance [PS17]. Access [ACKW01, AGC^{+}16, DJ02, Gal11, GDDR16, LPB13, PKST08b, PLW17, QRP22, MMG^{+}05]. Accessible [Oli13]. Accidental [BPH^{+}13, HMS^{+}00, PA22]. Accuracy [LPCY19, PSS20]. Accurate [TM14]. ACE2 [XDK^{+}20]. AceDB [STM99]. ACES [YLCZ05]. ACES-iSERVO [YLCZ05]. Achieving [OS04]. Acid [Maj03]. Acknowledged [KFMG20]. ACM [BTL19]. Acquisition [Azo06, Cas16, CN03, CC99, LGW^{+}17, MM16, NC03, PAN^{+}16a, PNL^{+}16, PS17, Shi01a]. Acquisitions [DCWH07]. across [PS17]. Action [Lat16]. Active [GHKZ17, SD11, YYG^{+}19]. Activities [Par12]. Activity [Har23, Muz19, NWP19]. Actuator [YMHQ19]. Ad [Ano18q, Ano18-39, Ano18-35, Ano18-36, Ano18-40, Ano18-46, Ano18-44, Ano20w, Ano18b, Ano18-53, Ano18-54]. Adaptation [GKG^{+}15]. Adaptations [DBB^{+}21]. Adapting [STWK15]. Adaptive [Beh05, BLT^{+}23, BMP^{+}06, Bry99, DBB^{+}21, GPZ^{+}04, KRR^{+}12, MCAA05, NSLD99, VSG^{+}02, WB03]. Adaptively [BW06]. Adder [GBPR11]. Adding [Mal00]. Additional [AAGH17a, AAGH17b]. Addressable [Hin20a]. Addressing [Bot16, FAFX20, HG02]. Adds [Got02a, Shi01a]. Adjoint [CL14, DAEJ18]. Adjoint-Based [CL14]. Adjustable [Par00b]. Adjusting [KS13]. Administrative [Ano18z, Ano19y]. Adobe [Tof08, Tof09a, Tof09b]. Adopting [Sch23a, SWPB00, Tan21]. Adoption [PS17]. Adsorption [KM99, WWJH20]. Adsorption-Desorption [KM99]. Advance [Ano15-37, TMC^{+}13]. Advanced [Akl18, DGG^{+}21, DRR^{+}04, Got15, Got16, Har23, MPR18, Men18, MHC^{+}18, MSD10, STG11, SNTL13, TMMB18, Tow18, Wat21, WZZ11, dJM18]. Advancement [Che15]. Advances [AMS14, CK09, ECK^{+}15, GL99, GK18, HP14a]. Advancing [CJL^{+}18, Gor13, KMB^{+}08, LG10, PV00, SYM^{+}21]. Advantage [Mil10]. Adventures [Nob00b]. Advertisement [Ano13g, Ano13h, Ano13p, Ano13n, Ano14d, Ano14e, Ano14g, Ano14h, Ano14i, Ano14l, Ano14j, Ano14m, Ano14t, Ano14-32, Ano14v, Ano14z, Ano14-29, Ano14w, Ano14-27, Ano14-28, Ano14-31, Ano14-30, Ano14-39, Ano14-40, Ano14-41, Ano14-42, Ano14-45, Ano14-46, Ano14-47, Ano14-48, Ano14-49, Ano15b, Ano15c, Ano15e, Ano15f, Ano15i, Ano15g, Ano15k, Ano15l, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-35, Ano15-36, Ano15-37, Ano15-43, Ano15-40, Ano15-41, Ano15-42, Ano15-39, Ano15-45, Ano15-49, Ano16c, Ano16d, Ano16e, Ano16l, Ano16t, Ano16q, Ano16u, Ano16r, Ano16z, Ano16v, Ano16n, Ano16s, Ano16-27, Ano16w, Ano16o, Ano16x, Ano16p, Ano16-28, Ano16-29, Ano16-30, Ano16-38, Ano16-39, Ano16-40, Ano16-47, Ano16-48, Ano18-37, Ano13b]. advertisement [Ano13c, Ano13d, Ano13e, Ano13k, Ano13o, Ano14f, Ano14k, Ano14x, Ano14-44, Ano14-56, Ano15h, Ano15j, Ano15-28, Ano15-38, Ano15-44, Ano15-48, Ano17-28, Ano17-29, Ano17-30]. Advertising [Day10b]. Advice [Day09a]. Aerodynamic [LQZL19]. Affects [KRH^{+}99]. Affordable [Weg00]. Africa [Amo18]. African [Sca16]. African-American [Sca16]. After [AM20, DF21, Key05]. Again [Cho07b, Sul07b]. Against [FGHW20, HP20, VSB^{+}21, WHG21, Wes21]. Age [Gar06, PQQ20, Thi02]. Agencies [MBB^{+}22]. Agency [Smi16]. Agent [AM05, DYY^{+}17, FGRS17, HXMC05, PI16, SU22, SMSC22]. Agent-Based [AM05, HXMC05, PI16, SU22, SMSC22]. Agent-Oriented [FGRS17]. Agents [BO04]. Aggregates [KLS01]. Aggregating [DMXR^{+}14]. Aggregation [LLQ18]. Agile [ABC^{+}14, PK18, SHPL12, VB08, Var08, FKB^{+}13]. Agilent [Tou01]. Aging [Dub15a, Fen06, dOMdO^{+}04]. Agreement [Smi99c]. AI [Ano22c, Guo23]. Aided [Ass00, Day17a, Gig00, JS99, Lew02b]. Aims [For01, TMC^{+}13]. Ain’t [Dub08c, Sul04f]. AIP [Ano18c]. Air [EDJ^{+}10, PAN^{+}16b]. Aircraft [KNKP14, MM16]. Airspace [Don03]. al [Lud13]. Alan [Lov04, Lov04]. Alert [SSG16]. Algebra [FGP99, HH22, LPV00, Los03, Sch14, ZZPC06]. Algebraic [Fal06, RLRML04a]. Algorithm [ALH15, ATG05, ACF18, BST^{+}13, Bea00, BS00c, BS00d, DCC10, Ebr10, FAFX20, Lui06, MBS^{+}00, OOB17, PL07, PSS20, Par00a, Pre09, Roc00, Rus03, SSP06, SR12, Thi13c, YYL^{+}18, ZZYNH06, ZLTX19]. Algorithmic [Fra07]. Algorithms [BFL^{+}22, DS00, Eng09, LGG^{+}22, MBS^{+}00, MSS09, NKV99, Rag07, Sul00c, Wep15]. Alignment [XHL^{+}13]. All-Digital [Thi15a]. All-Optical [GBPR11]. Alliance [CMN00]. Almost [Shi00a]. AlphaGo [Che16]. Alternative [ALH15, CF03]. Alternatives [EHG01]. Am [Sul07a]. Amazon [FPRK16, JRD^{+}13]. Ambient [Dac16]. America [MHC^{+}18, Smi00b]. American [Ano18d, Ano18e, Ano18f, Sca16]. among [Cho12]. AMP [Zhu16]. AMR [ZSM^{+}22]. Analog [Azo06]. Analogous [CB02]. Analyses [Cor07, HE05, PLH^{+}18]. Analysis [ALH15, AAH^{+}08, Ano15e, BAD^{+}21, BT01, BKK15, BCC^{+}99, CR15, DVP^{+}17, Dal99, DH12, DLLM04, DMR^{+}09, DKWL17, EHG01, GARS^{+}20, GM02, GYJL20, GNB^{+}09, HAB17, HBB08, HB08, JSNR11, JXY^{+}19, JRP^{+}17, JON^{+}21, KHE13, KL07, LFC01, MRU^{+}15, MAFM21, MB07, O’L06f, OMKdSB11, Ome06, PI16, PAN^{+}16b, RD05a, RD05b, SSCN11, SCW^{+}17, Shi99, SPB^{+}20, Sul06b, TGP^{+}06, TKM^{+}18, TGEA09, VN99, VGM^{+}09, WT12, XYC^{+}09, YLZ^{+}19, ZLW^{+}19, SOV^{+}13]. Analytic [NGGS22]. Analytical [DPBS16]. Analytics [AAB^{+}13, Ano14-44, Ano16-39, ERF21, GP15, KHE13, RSZ^{+}20, RSZ^{+}21, SAK^{+}13, SBZB13, Wan18]. Analyzing [ABK^{+}02, CS01a, DL00, O’L07a, O’L07b, SPW^{+}13, Vla12, WGJ16]. ANARI [SGA^{+}22]. Anatomic [CZ07, LYC07]. Anatomical [GWA^{+}07]. Anatomy [LCY08, YCL05]. Anderson [SS11]. Anecdotes [Got14a]. Aneurysm [WNZ^{+}17]. Angle [Nob02b, O’L12]. Animal [DL00]. Animating [Sil00, YCZ07]. Animation [LJWC06, WLCJ12]. Anisotropic [FL05]. Annals [Ano21i, Ano21j, Ano21k, Ano22i, Ano22j, Ano23c]. Anniversary [SCBT18]. Annotated [Wep15]. Annotating [WGJ16]. Announcing [Leu17]. Annual [Ano99, Ano00a, Ano01, Ano02a, Ano03, Ano04a, Ano05a, Ano06, Ano08, Ano09a, Ano11c, Ano15b]. Anomaly [Smi99d]. Answer [Sul05a]. Antennas [PAN^{+}16b]. Anthology [Nai15]. Anticocaine [Gor06a]. Antiproton [SSK02]. Antiproton-Hydrogen [SSK02]. Any [Pre09]. Anymore [Cho05b]. Anything [Sul99b]. apeNEXT [BST^{+}06]. API [RR21, SGA^{+}22]. App [Day14a]. App-Scale [Day14a]. Apparatus [ZDW^{+}07]. Appliance [SDCV10]. Appliances [How12]. Application [ACQ^{+}20, Bas02, BAD^{+}21, BHC^{+}15, DGK16, DMA^{+}21, DD07, HRRS09, Lau05, Läu06, LTG07, MM18, NdSS17, NdS17, SSCN11, SCW^{+}17, SKP^{+}10, SB00, WY12, ZZS^{+}19]. Applications [Ano15b, Ano20z, ACF18, Ara99, BFF12, BC05a, BC05b, Bry11, CJL^{+}18, CF99a, Che19, DBH^{+}02, Di 11, DSK15, DYY^{+}17, DG12, EVA^{+}21, Fox03d, GZC14, HPMJ12, JSNR11, JPE20, KNG10, KVP^{+}16, KVP^{+}17, KSM^{+}17, KSB07, KTG08, KHC^{+}07, LZZ17, LGG^{+}22, MDW^{+}22, Mem16, MV20, MHC^{+}18, NGGS22, OKS10, PF04, PMFM14, PAN^{+}16b, Rag07, Ran06, RLRML04b, SBZ^{+}08, SSP06, SKC02, SL03, Shi07, SBZB13, SF11, SPB^{+}20, Tec18, WCAL14, WD06, ZLW^{+}19, Wep15, Ano20a, Ano20f, Ano21a, Ano21b, Ano21l, Ano22k, Ano22l, Ano22m, Ano22n, Ano23d, Ano23e]. Applied [Coo14, MT00, Rei21a, Rei21b, RHC^{+}23]. Apply [Eng15]. Applying [ST05, SFSK01, TX08, TMC^{+}13, YLR02]. Approach [ABNZ09, Ama00, AM05, Bas14, Bet99, BZL^{+}07, Che99, Che03, DPG^{+}12, EGFL12, FGR^{+}07, FGRS17, Gan02, JMEL08, Kal19, KC19, KPA^{+}16, Kyr08, O’L05c, RRN20, RGD13, Roo06, SMF^{+}23, SR12, Ste00, SKC05, VMK20, VWP12, Wri16, WCH12, XKG05, YMLJ06, YM14, ZS07, ZWYL20, GGD^{+}05]. Approaches [Cos22, MVUSK14, MMR22, NG20, Pat02, TGP^{+}06]. Approximate [WQLZ18].

Approximating [GBDW04].

Approximation [Hin15a, Kus06a, Sul10a]. Arbitrary [GLTZ10]. Arboreal [Shi03]. Archaeology [Day16a]. Architecture [DADY15, DYY^{+}17, EWN^{+}13, EVA^{+}21, GHKR11, GHKZ17, HKB12, HDB^{+}04, HRRS09, Jer13, KS13, Läu06, LTG07, PKST08a, PKST08b, PKST08c, SES^{+}11, Lau05]. Architectures [Dec15, DGG^{+}21, KTG08, Kin09, Kog09, MSD10, Sch15, SBB^{+}15]. Archivable [AISR^{+}21]. Archive [Nei08, TSKG03, TSFG08, Tha08a]. Archives [BDS13]. Archiving [Gor06d]. Arctic [EKLY07, MKJ07, THGS07]. Argentina [BPMKC21]. Argonne [CJ16]. Argument [Fox18]. Arithmetic [Bai05, O’L06b, PPE00, Smi03]. ARL’s [War18]. Arm [DiP18b, Els21]. Array [YBBP15, vdWCV11, CBF^{+}22]. Array-Structured [YBBP15]. Arrays [Chi21]. Arrival [BBWW22, OM03]. Art [Bei12d, Gor05b, Sul10a]. Arterial [VSE01]. Arteriovenous [DKCL14]. ARTICLES [Ano21v, Ano20q, Ano20r, Ano20o, Ano20p, Ano21-29, Ano21-30, Ano21-31, Ano21-27, Ano22v, Ano22u, Ano21-28]. Artifacts [CGZ20, HBG^{+}20]. Artificial [Ano22w, Ano22a, BB20, Gor06b, PI16]. Arts [SL18, Tan21]. ASE [LVLA14]. Ask [Sul03b]. Aspects [GC00, SGA03]. Assays [WZS^{+}10]. Assessing [HPC20, LMC20, XHL^{+}13]. Assessment [FCT^{+}10, KL07, SPB^{+}20]. Assignment [O’L06e, O’L07b, XXK^{+}02]. Assimilation [WZZ11]. Associated [TAM^{+}14, Van12]. Assurance [HLS^{+}16, RPEB14, SSW21]. Asteroid [GWMG04]. Astronomers [Day14f, Gor08c, RTSS14b]. Astronomical [Day15a, WAS^{+}12]. Astronomy [BDS13, HJLH03, LL13, MSL02, PBSS14, WBB^{+}20]. Astrophysical [CDF^{+}04, JON^{+}21, Owe01, RCD^{+}00]. Astrophysics [ABNZ09, CK09].

Asymmetric [GBPR11]. Asynchronous [DAEJ18, DMA^{+}21]. ATI [BCH^{+}09]. ATLAS [Cas16]. Atmosphere [FSD02]. Atmospheric [Kel10, PD02, Beh05]. Atom [SSK02, WLL^{+}14]. Atomic [Che03, HBD^{+}23, WPZ00]. Atomistic [Adl03, BMCC21, GL99, GIF^{+}12, KL10, TL08b, VKN99]. Atomistic-Continuum [GIF^{+}12]. Atoms [Vla12]. Atos [Els21]. Atque [Cho08a]. Atresia [yFZDY13]. atrial [SOV^{+}13]. Attempt [Far99]. Attempts [Kul07]. Attitude [HAB17]. Attitudes [CHH^{+}13, Sca16]. Attributing [ACG^{+}20]. Atypical [SOH13]. Audience [SNCM16]. Audience-Targeted [SNCM16]. Audio [BBN03]. Auditory [XYC^{+}09]. Augmented [ACQ^{+}20, DBJ^{+}20, MV20, SPB^{+}20]. Aural [yFZDY13]. Aurora [Lum07]. Australia [MMG^{+}05]. Authentic [GARS^{+}20]. Authoring [CBB06, Sil02]. automan [Ram18]. Automata [MT00, Sch20, Sch23c]. Automated [Dav12, Edd09, JPE20, KC09a, KC09b, KC09c, MSB^{+}14, Ter11, XXK^{+}02]. Automatic [GMPR11, MAC08, RGD13]. Automatically [XHL^{+}13]. Automation [OSM^{+}19, Ram18]. Autonomous [LL19, VVNV18]. Autoregressive [RD05b]. Autotuner [Vud22]. Autotuning [YB12]. AV [HWPS16, HLS^{+}16]. Available [CE18]. Avalanches [Boe00, KPD^{+}99]. Avatars [YWC02]. Ave [Cho08a]. Average [Smi00b]. Award [Ano14y, Ano15g, Ano16n, Ano16o, Ano16p, Ano17p, Ano20-36, Ano21-48, Ano14y, Ano14-43, Ano16b, Ano17a, Ano18p, Ano18b, Ano18-38, Ano20-36, Ano21-47, Ano21-48, Ano21-46, Ano22b, Ano22-75]. Awards [Ano16c, Ano16d, Ano17b, Ano18h, Ano18q, Ano20n, Ano21y, Ano21z, Ano23-37, Ano23-38, Ano23-39, Ano20-51, Ano20-52, Ano20-53]. Aware [GHKZ17, TFF05]. Awareness [MM12]. Away [ERS^{+}03]. awesome [Day12f].

B [Ano17a]. Baaack [Dub06a]. Babbage [Ano16b, Ano18b, Ano22b]. Babel [Cho09b]. Back [Cho08b, Her21, Sch17b, Thi11a]. Backbone [XXK^{+}02]. Background [BCJK99, BKK15, Tho99a]. Bacterial [Hin17a]. Bad [Sul04b]. Bags [Day10a]. Balance [Kul07, WSC^{+}04, YYL^{+}18]. Balanced [CS01a]. Balances [BZL^{+}07]. Bang [BKK15]. Bangs [MKM^{+}14]. Bar [SSW21]. Barrier [JC02]. Based [ALH^{+}20, AAGH17a, AAGH17b, Ama00, Ano16-40, ACF18, AM05, AGC^{+}16, Aya14, BCC^{+}09, CL14, CF99b, CYW01, DVP^{+}17, DCWH07, DLLZ19, DLLZ20, DAKM16, DMA^{+}21, FOdLVF^{+}11, GGJD23, GW15, GPMSC20, GYJL20, GNB^{+}09, HLS^{+}16, HWPS16, HLYQ19, HGV^{+}08, HSJ^{+}19, HXMC05, JJ15, JLYL19, Jq19, KSW^{+}12, KMSH10, KPA^{+}16, KVP^{+}16, KBPW15, LFC01, Les16, LGJ^{+}19, Luo12, Mil10, MWC^{+}16, MMR22, NLGNJ13, NW13, NWP19, Osk07, PL07, PTML11, PCY14, Peo20, PQQ20, PI16, Pos09, PMW20, QSEQJFH20, Raf16, Ram18, SÜP^{+}11, SSCN11, SU22, SSG16, SBW^{+}19, SDCV10, SWDR22, SAK^{+}13, Sny13, SR12, SMSC22, SGRK^{+}18, TMMB18, Tec18, TKM^{+}18, TNV^{+}02, VMK20, VWP12, WCHRM21, WLCJ12, WCC^{+}02, WPM^{+}12, Wan18, WCH12, WCC^{+}19, YYL^{+}18, ZDW^{+}07, ZZC^{+}19, ZWYL20, ZZYNH06, vGDS18, GGD^{+}05]. Baseline [SBW^{+}19]. Bases [RLRML04a, RLRML04b]. Basic [HW15, Rus01a, Rus01b, Rus02, Rus03, Tof08, Shi00b]. Basin [WSC^{+}04]. Basis [Lau02, LWG19, Ste12]. Bat [Don06b]. Batch [LT08]. Batteries [Dub07e]. Battle [FGHW20, WHG21]. Battlespace [CAP^{+}10]. Bayes [PL07]. Bayes-Based [PL07]. Bayesian [BT01]. Be [Alt10, Ano16q, Ano16r, Ano18s, Ano19u, BS02, Cho07e, Dub08a, Got01, Sul04f, Sul06c, Roa04]. Beacon [BHC^{+}15]. Beam [SG00]. Bean [Dub05a]. Bean-Counted [Dub05a]. Bearing [ZZC^{+}19]. Beat [DC04]. Become [Lew00c]. Becoming [Day11b, Got14b]. Been [Day07b, Fla17, Smi99a]. Beetles [O’L07a, O’L07b]. Before [DC04, Day15a]. Begin [Bei09a]. Beginning [Bei09a, TW17]. Behavior [CF99b, HS12, LL11, Roh10]. Behavioral [Ano15e]. Behaviors [MSR15]. Being [AH07, Fox02c]. Belonging [Wri16]. Benefits [CMN00, Tou02b, YCD^{+}21]. Bent [KCPFT02]. Bent-Double [KCPFT02]. Beowulf [Gob05, VGM^{+}09]. Best [ACG^{+}20, BBC^{+}11, CJL^{+}18, DE17, For16a, NRG^{+}17, Tou00, Beh05, Ano19-31, Ano19-32]. Better [Ano17x, Ano17y, Ano17z, Ano17-27, GAB^{+}22, Gor07a, Lan06, Wil06, Ano17-28]. Between [BHKW03, MGD^{+}08, Rei21a, Rei21b, GRE99, OW01]. Beyond [Cus13, Lof03, Mei03, Pie04, SFC07, Sin18, The03, Thi11a, dSRT16, Ano18j]. Big [AAB^{+}21, AHS11, Ano14-44, Ano16z, Ano16-27, Ano16-28, Ano16-29, Ano16-38, Ano16-37, Ano18-27, Ano20j, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21x, Ano22q, BKK15, Bre17, Cus13, Day13d, HP15, Kus07, Lew01b, LL18, MKM^{+}14, Pos16, SAK^{+}13, WCAL14, WLL^{+}14]. Bilevel [JG12]. Billiards [Bäc07b]. Bills [SW10]. Binary [Bez08, BVT^{+}21, GBPR11, SETK05]. Binary-Coded [GBPR11]. Bind [Cot03]. Binding [TL04a, XDK^{+}20]. Bindings [JPE20, LFN^{+}11]. Biochemistry [PP20]. Biocomputation [KS02]. Biodiversity [DBCN03]. Bioengineering [CL01, Eth01]. Bioinformatics [GL08, SF11]. Biological [Bor02, GW15, Gor06a, MHK^{+}06, SMSC22, TLR10, dOMdO^{+}04]. Biologically [For00]. Biologists [Kin16]. Biology [Cho06c, DV99, LA18, MPRU23, MS99, MGS07, Wep15]. Biomechanical [ZCXM99]. Biomechanics [VSE01]. Biomedical [FKB^{+}13, Joh12, WZS^{+}10].

Biomolecular [ABK^{+}02, AM20, CR15]. Bionic [LQZL19]. Biophysics [Eth01]. Bipartite [XXK^{+}02]. Birds [Gor06d]. Bisexual [Wri16]. Bit [TJ14]. Bite [Day07d]. Bits [Day07d]. Black [BkCP22, BVT^{+}21, CHM^{+}20b, CHM^{+}20a, Day16d, PRCL^{+}22, WMB20, Hig04]. Blended [LPB13]. Blind [O’L05e, O’L05d]. Blockchain [Di 17]. Blocks [TBM^{+}19]. Blogging [Day13a]. Blood [GIF^{+}12, Luo13]. Bloom [Sul05b]. Blue [Hsu06]. Blur [Day17c]. Board [Ano13k, Ano18-29, Ano19v, Ano20-50, Ano21-64, Ano21-65, Ano21-66, Ano21-67, Ano22-55, Ano22-56, Ano22-57, Ano22-58]. Boards [HRWS06]. Bodies [EVL^{+}17, QSEQJFH20]. Body [MW14, Ron14, YB12]. Boltzmann [Bog03, SFSK01, WNZ^{+}17]. Bombs [Day11d]. Bond [Day09b, JC02]. Bone [DZW^{+}05]. Book [Ara99, BC02, Bar12, Bas14, Bil00, Cyb01, Fel00, Kwa17, Lov04, Lud13, McK00, Nai15, Seg99, Sha14, Sny13, Sul05a, Tur14b, Vog13]. Books [Ano10a, Ano11a, Ano11b, Ano12a, Bac10, CG08, Hai10, LC12, TCCC13]. Boost [Moo21, PK18]. Boosting [MHC^{+}18]. Borders [HH99, Day11c]. Boring [Sul06b]. Born [Jon15, Sul06a]. Bose [KF03, STTV05]. Bot [Sul06c]. Both [BBC^{+}11, O’L05f, Sma12]. Bottleneck [DPBS16]. bottom [LC12]. Bottoms [Che99]. Bottoms-Up [Che99]. Bound [BS02]. Boundaries [KSSF11]. Boundary [BS21, Moy06]. Bounding [Iac21]. Boys [WMB20]. Brain [Dub08a, Gor08b, GIF^{+}12]. Breakthrough [McM09, Wol16]. Breast [Gig00]. Bridge [DLW^{+}19, MGD^{+}08, ZLTX19]. Bridges [MTG^{+}12]. Bridging [Gyu99]. Brief [QL19]. Bringing [MM13, Mei10, SSK13, VGD^{+}11, YCD^{+}21]. Brittle [RcK99]. Broad [Mes15]. Broadband [GHK^{+}08, SKP^{+}10].

Broadened [Cho06b]. Broadening [Slo16]. Browser [SSC18]. Browsing [Day11c]. Bubbles [Smi01a]. Bugs [Mil21]. Build [DEK03]. Builder [ASM^{+}14]. Building [BBM^{+}21, BDCT05, Den16, FIG^{+}23, Gor07a, Kni05, Maj03, NCM^{+}14, Roa04, TBM^{+}19]. Bundle [Kra03]. Burning [HBB08]. Bus [SG10]. Bushing [Jq19]. Business [WY12]. Butt [JXY^{+}19]. Buying [Day11a].

C [JPE20, CNC10, Cot03, EGFL12, GRE99, HC99, PMM^{+}08, RMX12, UZC^{+}12, Zak18, Zhu16]. C-to-CUDA [UZC^{+}12]. Cached [PE09]. Caching [DB21, XLLJ04]. CAD [SWDR22]. CAD-Based [SWDR22]. Cafe [Dub04]. Calculating [RLHGA^{+}13]. Calculations [Eis17, Kyr08, WOAEAG10, dKCAY00]. Calculators [Got02a]. CALL [Ano21v, Ano13a, Ano14d, Ano14a, Ano14b, Ano14c, Ano14y, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15w, Ano16c, Ano16d, Ano16l, Ano16n, Ano16o, Ano16p, Ano17a, Ano17b, Ano17p, Ano18g, Ano18b, Ano20q, Ano20r, Ano20o, Ano20p, Ano20n, Ano20s, Ano20t, Ano20u, Ano20v, Ano20-36, Ano20-44, Ano20-45, Ano20-54, Ano21-29, Ano21-30, Ano21-31, Ano21-27, Ano21-28, Ano21y, Ano21z, Ano21-32, Ano21-33, Ano21-48, Ano21-53, Ano21-54, Ano21-55, Ano21-56, Ano21-57, Ano21-58, Ano22c, Ano22v, Ano22x, Ano22y, Ano22w, Ano22z, Ano22-41, Ano22-42, Ano22u, Ano23w, Ano23v, Ano23x, Ano23y, Ano23-41, Lat16]. Calls [Ano14u]. Can [Alt10, Bar11, Eng15, FLV^{+}09, HGA23, Kus07, NO03, Roc00, Ste14, Sul99b, TM15]. Canal [yFZDY13]. Cancer [Bre17, Dub08a, Hei20]. Candle [OR12]. Cannibalism [O’L07a, O’L07b]. Canonical [GF04]. Can’t [Ben09, DEK03]. Capabilities [Bas02, Tou01]. Capability [EDJ^{+}10]. Capstone [DAKM16]. Capture [Dav12]. Car [Bei12c]. Carbon [CJ16, JXY^{+}19, SMC01, TAM^{+}14]. Cards [Azo06]. Care [Bei09b, Jef22]. Career [Ano19-28, Ano19-27, Ano20-57, Ano22-43, Ano23o, Ano23z, Ano23-27, Ano23-28, CMK22, CHM^{+}20b, CHM^{+}20a, Dub07a, Ano16y]. Careers [Bot16, DVS22, GN08]. Caring [Hin12a]. Carlo [Ama06, BS06a, BS06b, BOS07, DCC10, Day13e, Eng09, EKCS12, Fel00, GH00, LSDP^{+}04, Lui06, OASFLAB09, Ork09, PPE00, SWDR22, SG00, Sul17, VWP12]. Carlo-Based [VWP12]. Carpentry [Wil06]. Cars [Day16b]. Case [AHL^{+}11, BBW^{+}20, BAD^{+}21, BDCT05, CDL^{+}23, COS^{+}15, DPBS16, DM12, DL23, DDV^{+}08, FAFX20, GHM^{+}16, Joh12, KMSH10, KPM10, Mas06, MB11, MR13, MM14, Mem16, MSS09, NCM^{+}14, NC18, PSSP15, PHW^{+}21, PBD^{+}11, RPEB14, SDA20, SM17, Sch18, Sch20, Sch21, Sch23b, SLK^{+}20, Sim13, SV14, Ste02, Tan21, WHH22]. Cases [SSW21]. CasJobs [LT08]. CASP [GHKZ17]. CASTEP [STH22]. Cat [Bei12c]. Catalog [Tha08a, TSFG08]. Catalonia [FGG^{+}22]. Catalyst [Rad21]. Catalysts [WPZ00]. Catalytic [VWP12]. Catalyzing [Par22]. Cathode [FMB^{+}07]. Causal [SSP06]. Cause [LG10]. Causes [Dal99]. Cautionary [Ste12]. Caveat [DC04]. Cavities [LPV00]. CDAT [SPW^{+}13]. CDE [Guo12]. Celebration [BTL19]. Celebrities [Day10e]. Celestial [Ono01]. Cell [CAS^{+}07, CS15, DV99, Dec15, LWT^{+}13, SÜP^{+}11, YZC^{+}13, CS14, GHK^{+}08, SKP^{+}10]. Cell-Based [SÜP^{+}11]. Cells [BMCC21, CFA04, Has08]. Cellular [CS15, FL05, MT00, Sch20, Sch23c]. Cellular-Automata [MT00]. Center [Ano16q, Ano16r, Ano18s, Ano19u, Ano22-43, Ano23z, Ano23-27, Ano23-28, NRG^{+}17, NdS17, RF00, Sul99a, MOBD^{+}22, WG15]. Centered [KPA^{+}16, NSP12, SGRK^{+}18]. Centric [DAKM16]. Century [Ano22-67, Ano23-47, Ano23-48, Ano23-49, Ano14-49, Ano22-66, Ric99]. Cerebral [Luo13]. Certainty [Sul01c]. Certification [KFMG20]. CFD [CCD^{+}22, TGEA09]. CG&A [Ano14u]. CGA [Ano20w]. Ch [CNC10]. Chain [She07]. Chains [Ran06]. Challenge [Cyb00c, Dal99, Day10d, Day11c, Kal19, SRKS22]. Challenges [AHL^{+}11, BLO^{+}22, BCJK99, BB01, Das00, DPG^{+}12, EUD15, GK22, GPL09, HL01, Joh12, KL15, Kog09, Kus07, MJM^{+}06, Muh23, Pos10, SBH^{+}00, TGP13, WPW11, YCD^{+}21]. Challenging [Leu20]. Chance [Thi15b]. Chancellor [Day12c]. Change [Cho06b, DM12, EJ09, HGA23, Joh06, MKJ07, O’L06e, O’L06g, SDA20, WPW11, YCD^{+}21]. Changed [Cho06d, DTL^{+}17]. Changes [Bei11a, EKLY07, Got17, WM00]. Changing [CHH^{+}13, KMB^{+}19, Pos13, SLM12]. Channel [FGR^{+}07, KLMS99]. Chaos [Bäc07b, Hin16, KB09, O’L07a, O’L07b]. Chaotic [Mar02]. Characteristics [Azo06, Biz16, OHK23, Rao16]. Characterizing [MB11]. Charge [YYL^{+}18]. Charles [Ano16b, Ano18b, Ano22b, Day09a]. Charts [Bei11c]. ChatGPT [Guo23]. Chatter [Kus06a]. Checkpointing [DAEJ18]. Chemical [Bor02, FSD02, MM04, MDK16, Par16, Tof08, Tof09a, Tof09b]. Chemistry [BW14, Das00, GS03, HS03, IBPV03, PP20, Sch17b, SDD^{+}08, TM00, UM08, WOAEAG10]. Chemora [SBB^{+}15]. Chess [Hsu06]. Chicago [Ree16]. Chief [Ano22-59, Ano22-68, Ano23-50, Cyb99a]. Chihuahua [Day10a]. Chile [BB20]. China [JJZC10, QL19, YLCZ05]. Chip [For01]. Chiral [Mal07b]. Chirp [Don06b]. Choice [Bea00]. Choose [Cou22, Ano16y]. Chore [Toh07]. Christof [Lov04]. Chromodynamics [Ale15, GHK^{+}08]. Cilk [Rob13]. Circuit [For01, MAFM21, PGC21, WCC^{+}19].

Circulation [YLR02]. CiSE

[Ano05b, Ano09b, Ano10c, Ano11c, Ano11d, Ano16e, Bei10d, Che17a, Cho06f, Day12c, Don99, Gyu99, SCBT18, Thi11a, Thi13a, Thi14, Thi15a]. CiSE-Reading [Day12c]. CiSEiest [Day10c]. Citation [CGZ20, Dru20, HBG^{+}20, HPC20, KHC^{+}20, WBB^{+}20]. Citizen [Chr15, COS^{+}15, JJ15, Mem15, MSR15]. Citizen-Based [JJ15]. CitScale [LMC20]. City [LMC20]. Clarify [BMS99]. Class [Ron14]. Classical [RD05a]. Classification [LPY18, LPCY19, PSS20, ZLW^{+}19]. Classified [MO03]. Classifier [DKCL14]. Classifying [KCPFT02]. Classroom [Mei10, SAC15]. Classrooms [NG20]. Clearance [YWMM04]. Clearly [NO03]. Clerman [Mol12]. Cleve [Mar99a]. Click [ERS^{+}03]. CLiiME [NCM^{+}14]. Climate [Bak21, Bal15, DPBS16, EJ09, EUD15, EAF^{+}23, FKS15, GKG^{+}15, Liu15, MSM13, Pie04, PD02, RRH^{+}02, SPW^{+}13, SBW^{+}19, SNCT13, SS02, SKA^{+}02, TM15, WPW11, WBS^{+}22, WC15, YCD^{+}21]. Clojure [KM12]. Closer [MM13]. Cloth [ZJW08]. Cloud [AAB^{+}21, Ano13g, Ano14v, Ano14-47, Ano14-48, Ano15w, Ano16l, FPRK16, GB20, Gor08a, How12, JCPS14, KKO^{+}20, KFS18, KILZ13, LHZ21, MB20b, MWC^{+}16, PARD13, RVG^{+}10, SSG16, SAK^{+}13, Sul08b, Sul09a, Thi10, TP13, Tsa14, VPL18, WZZ11, EAF^{+}23]. Cloud-Based [MWC^{+}16, SSG16, SAK^{+}13]. Cloud-Enabled [Thi10]. Cloud-Native [AAB^{+}21]. Cloud-Resolving [WZZ11]. CloudCast [KILZ13]. Clouds [Joh09, LTD11]. Cloudy [SS09, Thi15b]. Cluster [BUS21, BDCT05, DD05, DD07, Gob05, HRAB05, HKW03, Lui06, SES^{+}11, TFF05, Thi05]. Clustered [SSP06]. Clustering [HH99, KNV03, MO03, YCK03, Zhu02]. Clusters [BDCT05, DJS13, DSSS05, EKCS12]. CMD [BTL19]. CMD-IT [BTL19]. CNNs [LPY18]. Co [NRG^{+}17]. Co-Design [NRG^{+}17]. Coal [MRKK17]. Coalition [TGU21]. Coarse [BMCC21]. Coarse-Grained [BMCC21]. Coat [JC02]. Coatings [JC02]. Code [BJ02, Bar20b, BFH21, BW14, CGZ20, Dub21, GP21, Hin17c, Hin18d, KRR^{+}12, KMSH10, KPM10, LVLA14, PKST08c, RGD13, RCD^{+}00, Ter11, Van12, Wil06]. Coded [GBPR11]. Codes [Ben04, CDF^{+}04, Dec15, ECK^{+}15, KRR^{+}12, Roa04, VMK20, WJLY08]. Coding [MAFM21]. Coefficient [XBK10]. Coevolutionary [Boe00]. Coexistence [Hus22]. CoFlaVis [MRKK17]. Cognitive [Fox18, Thi13b]. Coherence [Kus06a]. Cohesive [GGJD23]. Cold [FAFX20]. Cold-Start [FAFX20]. Collaboration [BLT^{+}23, FL21, KE05, Les16, MK10, MHC^{+}18, PRCL^{+}22, SMS15]. Collaborative [CJL^{+}18, GDDR16, HHR^{+}13, PSR^{+}20, STWK15, Tur14a, Weg00]. Collaboratory [MKH^{+}23, PSR^{+}20]. Collage [KRR^{+}12]. Collapse [Hin19, MSS09, Ott16]. Colleague [Ano19-38]. Collection [JJ15]. College [CHM^{+}20b, CHM^{+}20a]. Colleges [SL18]. Collegeville [Her21, HCK22]. Collision [SSK02]. Colloids [Day13e]. Colony [CS14, CS15]. Color [Oli13, Smi16]. Color-Vision-Deficient [Oli13]. Colormaps [Thy20]. combination [Ano16e]. Combinatorial [GRS08]. Combinatorics [BS08, Gut01]. Combine [Gia22]. Combustion [AMCH07]. Come [Day18b, Gar06, Sul10b]. Comes [DS23, GM06, Sul02b]. Coming [CW05e, Cho08c, Gor08a, Kra15, OM03]. Commends [Ano12b]. Commentary [CHM^{+}20b, CHM^{+}20a]. Committee [Ano18z, Ano19y]. Common [OS04]. Commons [GHM^{+}16]. Communication [BC99, GDH^{+}23]. Communications [HOÖ99]. Communities [FIG^{+}23, GZC14, PS17, TTT15]. Community [ABM^{+}22, BBM^{+}21, Chu21, Den16, ESO08, Fom15, GSK^{+}23, GPMSC20, HPC20, KMB^{+}19, PMP21, PK18, RBC^{+}19, SGRK^{+}18, TS10, VGD^{+}11, WDC18, WBB^{+}20]. Community-Cyberinfrastructure-Enabled [ESO08]. Company [Ano15h]. Comparative [AAH^{+}08, DLLM04]. Comparing [HBG^{+}20, JRD^{+}13, Mis02, Sil02]. Comparison [Mus20]. Comparisons [Eng09]. Compartmental [PHW^{+}21]. Compatible [Bri23]. Compendium [Ome06]. Competition [SS06]. Competitive [Mil10]. Compilation [AE22]. Compiled [Dub00]. Compiler [Pad00]. Complete [Sch07]. Complex [BNNM04, BMS99, Don03, Ebr10, GF04, HHR02, LSDP^{+}04, LL11, Naj08, Nob07, Par00b, PSA14, SZM^{+}13, SBZB13, SKC05, TB11, TX07, Var08]. Complexity [Mer02, Sul00a]. Component [OMKdSB11]. Components [Dub02, Lan19, SL99]. Composable [Rob13]. Compose [Day11e]. Composite [JXY^{+}19, SGW02]. Composition [YEC^{+}19, ZL09]. Comprehensive [PGF^{+}15, Tes15, TBVP^{+}21]. Compressing [Ama00]. Compression [SSP06]. Compressors [GvdWT07]. COMPSAC [Ano19d, Ano19e, Ano20x, Ano20y]. Computability [Lau02]. Computation [Bai05, BC99, BSD07, Cho09a, Cho12, CS01a, Day07a, HLS^{+}16, Hin20b, Hu07, Kar99, Kin16, Kir03, KRH^{+}99, LG10, Liu11, Mei10, Ott16, PBSS14, PS02, Raf16, Rei02, SSC18, SJDV09, Ste00, Sul99b, TFF05, TB99, Win06, YMK11, vdWCV11]. Computation-Based [HLS^{+}16]. Computational [AM18, Adl20, Aya07, Aya14, Bäc07a, Bar19, Bar20a, Bar21b, Bar21a, BLMR21, BW14, BHV21, BERT09, CK09, CL14, Car09a, Car12, CE14, CHC17, CB02, Chu21, CG09, Cyb99a, DSPY05, Dal99, Das00, Dav12, Day06b, Day11b, Day12a, DV99, DL00, DD23, DM12, DPG^{+}12, DMR^{+}09, DG12, Ebr10, EI11, FKD21, FGHW20, FG01, Fox02a, FKSS08, FWGB07, GC00, GPL09, Gor10, GCV08, Gor13, Got14b, GL20, HS03, HL01, Hei22, Hei20, Hin15a, Hin15b, Hin17c, Hin23, HG00, HPMJ12, HMB^{+}14, JHJ01, JH18, KLS01, KHS09, KSP12, KSM^{+}17, KLMS99, Lan04, Lan06, LC09, LPB13, LPB15, LHN^{+}12, LA18, LNI^{+}19, LWSK07, LM07b, MM12, MPR18, Mar17, Mas06, MK10, MB11, MR13, MM14, Mem16, Mer02, MS99, MB17, MB20b, Mes15, MCGA22, Mil17, MHC^{+}18, MS07, Naj08, NMCM22]. Computational [NL99, O’L06a, Oug03, PARD13, Par12, PLH^{+}18, Pat02, PP20, PQQ20, PGH^{+}05, Pos07, Pos09, Pos10, PG17, Ric99, Roo06, RHC^{+}23, Run00, RF12, SH10, SBH^{+}00, Sch15, Sch17b, SM17, SGA03, SDD^{+}08, ST08, Sim13, SV14, SK01, SR13, SMC01, STB03, SMS15, Sul03a, Sul04a, TGP^{+}06, TB11, TK06, Tes15, TAHP23, Tha14, Thi02, TB04, Thi09a, THLK10, Thi13a, Thi15c, TX08, TS10, TGU21, THGS07, TP04, TM00, Tur14a, VGD^{+}11, VB22, Vir16, VSE01, WHG21, WPW11, WM00, Wep08, WL09, WCH12, YRT^{+}00, YMLJ06, YM14, Yas17a, Day12e, PAN^{+}16a, PNL^{+}16]. Computationally [LCG^{+}20, Sch07, SYM^{+}21]. Computations [Bar20b, DM04, DKK05, FS12, Ful06, GBDW04, Mus20, PE09, SKC00, VCvdG^{+}09, Wat21]. Compute [Day09b, HRWS06, Has08, HRRS09, Sul06a]. Computer [AAGH17a, AAGH17b, Ano13i, Ano13h, Ano14z, Ano14-29, Ano14w, Ano14-27, Ano14y, Ano14x, Ano14-28, Ano15x, Ano15y, Ano15z, Ano15-27, Ano16t, Ano16q, Ano16u, Ano16r, Ano16z, Ano16v, Ano16n, Ano16s, Ano16-27, Ano16w, Ano16o, Ano16x, Ano16p, Ano16m, Ano16y, Ano17o, Ano17j, Ano17k, Ano17l, Ano17m, Ano17p, Ano17n, Ano17-31, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano18s, Ano18r, Ano19t, Ano19u, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19v, Ano19s, Ano20z, Ano20-33, Ano20-34, Ano20-35, Ano20f, Ano20-46, Ano20-38, Ano20-39, Ano20-47, Ano20-51, Ano20-40, Ano20-48, Ano20-52, Ano20-41, Ano20-49, Ano20-44, Ano20-42, Ano20-43, Ano20-45, Ano20-53, Ano20-50, Ano21-42, Ano21-43, Ano21-44, Ano21-40, Ano21-41, Ano21-45, Ano21l, Ano21-49, Ano21-64]. Computer [Ano21-53, Ano21-59, Ano21-54, Ano21-65, Ano21-60, Ano21-55, Ano21-56, Ano21-61, Ano21-50, Ano21-57, Ano21-62, Ano21-51, Ano21-63, Ano21-66, Ano21-52, Ano21-58, Ano21-67, Ano22x, Ano22y, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22k, Ano22l, Ano22m, Ano22n, Ano22-39, Ano22-44, Ano22-48, Ano22-55, Ano22-41, Ano22-45, Ano22-40, Ano22-56, Ano22-46, Ano22-52, Ano22-57, Ano22-49, Ano22-53, Ano22-59, Ano22-58, Ano22-47, Ano22-50, Ano22-42, Ano22-54, Ano22-51, Ano22-43, Ano23t, Ano23d, Ano23e, Ano23-31, Ano23u, Ano23w, Ano23-37, Ano23-29, Ano23v, Ano23-34, Ano23-38, Ano23z, Ano23-30, Ano23-32, Ano23-35, Ano23-39, Ano23x, Ano23-27, Ano23-33, Ano23-36, Ano23y, Ano23-28, Ass00, BCE^{+}22, BMP^{+}06, BT10b, Boe00, BkCP22, CDL^{+}23, CBF^{+}22, CF99a, CSS00, Day06c, Day17a, Dec15, FL99, FM02]. Computer [FGP99, Gig00, GPMSC20, Gor08b, GH00, HHR02, HT99, HGA23, Jav12, JS99, Kad04, KS00, KS13, KPD^{+}99, KBPW15, LDAS19, LSBC22, Les16, Lew02b, LWT^{+}13, LPV00, Los03, MMTD^{+}17, MB99, New00, OW01, O’L06b, PKST08a, PKST08b, PKST08c, PA21, PR01, Ran06, Ree16, Saa09, SDA20, SW10, Sca16, Sch14, Sch15, SS06, Shi01b, Slo16, TS02, TNR^{+}23, Tre99, Var08, Vla12, WHH22, WMB20, WCP17, dKCAY00, Mat05, Ano18y, Ano19-29, Ano19-30, Ano22-27, Wil01, Ano20a, Ano20-72, Ano21a, Ano21b, Sny13]. Computer-Aided [Ass00, Day17a, Gig00, JS99, Lew02b]. Computer-Based [KBPW15]. Computer-Guided [BT10b]. Computer-Simulated [Tre99]. Computers [Ano15b, Bal17, CL01, Cra03, Cre99, Day12d, Day16b, Dun09, FHM99, GS13b, JT01, McC21, PSA14, SDA^{+}14, WSY^{+}22, Ano20u, Ano20v, Ano21-32, Ano21-33, Ano21r, Ano23j]. Computing [Akl18, AE22, Ale18, Amo18, AMS14, Ano13b, Ano17c, Ano18-41, Ano19x, Ano23l, Ano23m, Ano23n, Ano23f, Bak10, BFF12, Bal15, Bal99, BT17, BKRT21, BPW^{+}20, BTL19, Bar21c, BCH^{+}09, BMC99, BS99b, BS99a, BS00a, BS00b, BS08, BST^{+}06, BCC^{+}09, Biz16, BT01, Bog05, BC05a, BC05b, BCJK99, BB01, BCH^{+}22, BHC^{+}15, BBM^{+}21, Bry11, CDL^{+}23, CHB19, CR15, CF99b, CN03, Che15, Che16, Che17b, CHJC05, Cho06d, Cho06g, CNC10, CG09, CW20, CT00, CHM^{+}20b, CHM^{+}20a, Cyb02, DD05, Dau99, DCWH07, Day06a, Day07c, Day13b, Day15e, Day17d, Den16, DMXR^{+}14, DW01, DSSS05, DTL^{+}17, DRR^{+}04, DKWL17, DMT^{+}21, Dub22, ES18, EDJ^{+}10, EHG01, FKS15, FLV^{+}09, FM13, For00, FG01, Fox01, Fox02b, Fox03b, Fox03c, Fra02, FPRK16, GHT^{+}10, GRS08, GB20, GHK^{+}08, Gor05a, Gor06a, Gor07b, Gor07d, Got06]. Computing [GS13a, Got14a, Got15, Got16, Got17, Gro09, HP14a, HP14b, HP20, HLRW17, HC99, Har23, HRAB05, HJLH03, HB08, HGV^{+}08, Hig04, Hin17b, Hin18c, HG02, HP04, HPMJ12, How12, JH16, Jef22, JR10, JLNR19, JCC^{+}10, Joh12, JPMG08, JON^{+}21, KKO^{+}20, KM99, KT08, KFS18, Kel10, KSB07, KTG08, Kin09, KWB^{+}10, KT11, Kin12, KS13, KLQ19, KILZ13, Kup03, KBLD08, Lan19, LMPV13, LM08, Lat16, LUMM14, LA18, LSBC22, Lew02c, Lew02a, LZZ17, LAY04, Lun01, MP09, MS22, MB20a, MWE08, MR06, MMTD^{+}17, Men18, Mes17, MMG08, MKM^{+}14, Muc09, MSD10, NG20, NC03, Nob00a, Oli07, Osk07, PGF^{+}15, Pap16, Par22, PA12, PLW17, PG07, PGH11, PMP21, Pos11, Pos13, QL19, Rag07, Ram18, RVG^{+}10, SA08a, SA08b, SBZ^{+}08, SBB^{+}15, SKC02, SBW^{+}19, SRKS22, Sch18]. Computing [Sha14, SES^{+}11, SL99, SOH13, Smi00a, Smi16, SKA^{+}02, SS09, SLM12, Sto12, SGS10, Str10, SJF^{+}23, ST99, Sul09a, Sza11, Ter11, Thi05, Thi09b, TP13, Tho99a, Tho99b, Tho00, Tho01, Tow18, TBVP^{+}21, Tsa14, VB08, VPL18, VGD^{+}11, VM15, WCGB05, WCAL14, WBP^{+}19, WG15, WBS^{+}22, WR00, Wes21, Wri16, YLZ17, ZFS12, ZGR^{+}17, ZAF^{+}01, Beh05, Ano03, Ano05a, Ano12b, Ano13g, Ano14v, Ano15w, Ano16l, Ano16-30, Ano18-28, Ano19a, Ano19b, Ano20q, Ano20r, Ano20b, Ano20c, Ano20g, Ano20m, Ano20k, Ano20l, Ano21-29, Ano21-30, Ano21-31, Ano21i, Ano21j, Ano21k, Ano21s, Ano21t, Ano21u, Ano22v, Ano22d, Ano22e, Ano22f, Ano22i, Ano22j, Ano22r, Ano22s, Ano22t, Ano23a, Ano23b, Ano23c, Ano23g, Ano23k]. ComputingEdge [Ano20d, Ano20e, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22g, Ano22h]. Concepts [BFS04, DR05c, HW15, PL02, Sch23a]. Conceptual [Ikk16]. Conceptualization [CGK^{+}18]. Concern [CGZ20]. Concurrency [DS12, Vin12]. Concurrent [ZL09]. Condensates [KF03, STTV05]. Condensed [IBPV03]. Condensed-Phase [IBPV03]. Condition [GYJL20]. Conditions [Dar18, Moy06]. Condor [KMSH10]. Conference [Ano13c, Ano22w, Ano22a, OW01, RHC^{+}23, Ano15b]. Conferences [Ano15i, BTL19, Dau99, HGA23]. Confidence [DTA21]. Confidence-Guided [DTA21]. Configuration [Gob05, JS99, MWE08]. Conformational [BH02]. Congenital [yFZDY13]. Congress [Ano19z]. Conjecture [Rei21a, Rei21b]. Connected [Ano19f, Ano19-40, Ano15-45, Ano18-45, Ano19-39, Ano20-64]. Connecting [MBB^{+}22]. Connection [Com99]. Conquer [O’L04e]. Consciences [Day18a]. Consciousness [KNKP14]. Consensus [SETK05, YYL^{+}18]. Conservation [AM05]. Considerations [SNCM16]. Considered [LGJ^{+}19, TLG06]. Consilience [Kal99]. Consortium [BCH^{+}22, HP20]. Constellation [Lo99]. Constellations [DSSS05]. Constituents [FSD02]. Constrained [XXK^{+}02]. Constructive [FL21]. Consultants [Sch23a]. Consumer [DC04]. Consumption [SPJ^{+}14]. Contact [BW01]. Containers [HLRW17]. Containment [FGG^{+}22, ZSM^{+}22]. Contend [Sul05b]. Content [Hin20a, TL04b, XLLJ04]. Content-Addressable [Hin20a]. Contents [Ano14-50, Ano14-51, Ano14-53, Ano15-47, Ano16-41, Ano16-42, Ano16-43, Ano16-44, Ano16-45, Ano16-46, Ano17-37, Ano17-32, Ano17-33, Ano17-34, Ano17-36, Ano18-47, Ano18-48, Ano18-49, Ano18-50, Ano18-51, Ano18-52, Ano19-41, Ano19-42, Ano19-44, Ano19-45, Ano19-46, Ano20-70, Ano20-66, Ano20-67, Ano20-68, Ano20-69, Ano21-75, Ano21-76, Ano21-77, Ano21-78, Ano21-79, Ano21-80, Ano22-69, Ano22-70, Ano22-71, Ano22-72, Ano22-73, Ano22-74, Ano23-51, Ano23-52, Ano23-53, Ano23-54, Ano13q, Ano13r, Ano14-52, Ano14-54, Ano15-46, Ano17-35, Ano19-43, Ano20-65]. Contest [Don99, LSV^{+}07, MHDM99]. Context [Dav12, GHKZ17, Luo12]. Context-Aware [GHKZ17]. Contingency [Bez08]. Continuing [ERF21]. Continuous [GK22]. Continuum [GIF^{+}12]. Contour [GMPR11, ZDW^{+}07]. Contours [QPCJ07]. Contradictions [Dub07a]. Contribution [HH22, MMG^{+}05]. Control [BHL99, Bet99, CXC^{+}20, Cho08g, Day14b, DDV^{+}08, EHG01, HAB17, HLT09, KB07, LCC^{+}19, OS03, PLW17, RSC^{+}14, SZM^{+}13, Var08, YMHQ19]. Controlling [RcK99, SGW02]. Convection [MGZ00]. Convective [CFA04]. Convenient [CC99]. Converge [DKWL17]. Convergence [HF04]. Convergent [ERF21, XBK10]. Conversations [Cho12]. Conversions [CY00]. Convex [Muc09]. Convey [Bak10]. Convolutional [ZZC^{+}19]. Convolutions [DR05a]. Cooperating [PGH^{+}05]. Cooperation [Day13c]. Coordinates [HW15, Vor01a]. Coordination [YYL^{+}18]. Cope [HHR02]. Coprocessors [BHC^{+}15]. Copyright [Sto09]. Core [Ano16-47, Ano16-48, BFL^{+}22, CWOL11, GHKR11, HKB12, MSS09, Ott16, Pes03, HKB12]. Core-Collapse [Ott16]. Cores [ZSM^{+}22]. Corneal [VSG^{+}02]. Corner [CF99a, CF99b, Che99, CY00, CYW01, DADY15, LRRK00, Weg00, ZCXM99]. Corps [Den16]. Correction [Nan11]. Corrections [CHM^{+}20b, DLLZ20, RSZ^{+}21]. Correctness [CRDO16, Dub05b, Fos17]. Correlated [WOAEAG10]. Correlation [FM19]. Cosmic [BCJK99, BKK15]. Cosmological [MRU^{+}15]. Cosmology [AAH^{+}08, Bry99, CDKF15, Jon15, TB99, Wei21]. Cost [CJTH^{+}13, JPK01, SW10, TS02]. Cost-Effective [TS02]. Costa [Men18]. Costs [BHL99, RLHGA^{+}13]. Could [Gor07d, Peg12, Smi01a, WJ04]. Counted [Dub05a]. Counting [BOS07, Bei12a, Fen06, SSCN11, Cho08d]. Countries [AM18]. Coupled [CBS14, DLW^{+}19, GIF^{+}12, JSNR11]. Coupling [CFCD04, STG11, ZLTX19]. Course [Ass00, Aya07, Bog05, GDDR16]. Courses [BB20, Cho06d, Ful06, GL08, Pes03, Win06]. Courseware [Thi12a]. CoV [FGG^{+}22, XDK^{+}20]. Cover [Ano14n, Ano14o, Ano14q, Ano14s, Ano15r, Ano15n, Ano15o, Ano15p, Ano15q, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano17i, Ano17e, Ano17f, Ano17g, Ano17h, Ano18o, Ano18k, Ano18l, Ano18m, Ano18n, Ano19g, Ano19h, Ano19i, Ano19k, Ano19l, Ano19m, Ano20-32, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano21-34, Ano21-35, Ano21-37, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano23p, Ano23q, Ano23r, Ano23s, Ano13f, Ano14p, Ano14r, Ano15m, Ano19j, Ano20-27, Ano21-36, Ano21-38, Ano21-39, Ano22-28]. Covered [Ano20-46, Ano20-47, Ano20-48, Ano20-49, Ano21-59, Ano21-60, Ano21-61, Ano21-62, Ano21-63, Ano22-48, Ano22-49, Ano22-50, Ano22-51, Ano23-31, Ano23-32, Ano23-33]. COVID [AM20, BPMKC21, BCH^{+}22, Com20, FGHW20, HP20, Jef21, KKO^{+}20, PHW^{+}21, PQQ20, TGU21, VSB^{+}21, WHG21, Wes21]. COVID-19 [AM20, BPMKC21, BCH^{+}22, Com20, FGHW20, HP20, Jef21, KKO^{+}20, PHW^{+}21, PQQ20, TGU21, VSB^{+}21, WHG21, Wes21]. CPU [EH22]. CPUG [Lan04]. CPUs [Alt10, AAAH^{+}16, WJLY08]. Cracks [Mar99b]. Craigslist [Day08b]. Crash [Bog05, YAA^{+}00]. CREATE [HWPS16, PAN^{+}16a, PNL^{+}16, Deb18, Hin16, HLS^{+}16, KPA^{+}16, KVP^{+}16, KVP^{+}17, LGW^{+}17, PAN^{+}16b, PS17, WQT^{+}16]. CREATE-AV [HWPS16, HLS^{+}16]. CREATE-GV [LGW^{+}17]. CREATE-SH [WQT^{+}16]. Creating [Guo12, OASFLAB09, SL18, Tof09a]. Creation [PI16]. Creativity [Day14b, Sch99]. Crevasse [GGJD23]. Crisis [Day17b]. Critical [Lan99, Run03, Ste99]. Cross [DKK05]. Cross-Site [DKK05]. Crossing [Gue18]. Crosstalk [Cho12]. Crosswell [CL14]. Crowdsourcing [PSSP15]. Crucial [Smi01b]. Cruising [Lew99a]. Crunched [Peg12]. Crunching [KM12, Peg12]. Cryoelectron [FL05]. Cryospheric [DM12]. Crystals [CCPS12, Mal07a, Mal07b]. CS [Ano15-36, Ano16b, Ano18b, Ano22b, Ano14-55, Ano14-56, Ano15-35, Ano15-48, Ano15-49, Ano18-30, Bei12d, MDK16]. CSE [Bea00, BB06, Bet99, Car09b, Haa99, WLCD01, YAA^{+}00]. CSGF [Rad21]. CSIFT [LPCY19]. cT [Shi00c]. Cuban [CE18]. Cubes [DSC^{+}09, WY12]. Cuckoo [FAFX20]. CUDA [DLLZ20, DLLZ19, UZC^{+}12, Vog13]. CUDA-Based [DLLZ20, DLLZ19]. Cultivating [vGDS18]. Culturally [NG20]. Culture [BBM^{+}21, Cyb99a, KMB^{+}19, SLM12]. Curb [Sul08a]. Curettage [LWT^{+}13]. Current [CCJ04, LCC^{+}19, MBB^{+}22]. Curricular [BCB07, CBB06, Joh06]. Curriculum [AAGH17a, AAGH17b, Got17, HC17, LG10, TK06]. Curve [LWG19, LL19]. Curvelets [HH06, MP09]. Customer [Dub07b]. Customized [TGEA09]. Cut [BS09]. Cutting [Bea00]. Cyber [SKL10]. Cyber-Enabled [SKL10]. Cyberinfrastructure [ESO08, HEB^{+}11, OHK23, PSR^{+}20, RP20, SMS15]. Cyberplatform [CHP^{+}18]. Cybersecurity [Ano14-45, Ano14-46, Ano15-43, Ano15-40, Ano15-41, Ano15-42]. Cycles [Ome06]. Cycling [OSM^{+}19]. Cyclone [DRA11, SNTL13, SNCT13]. Cython [BBC^{+}11].

D [Ano14y, Ano20-36, Ano21-47, Ano21-48, Sny13, Ama00, Ano15-38, Ano15-39, BB07, CCSS08, CFA04, CY00, CHC^{+}11, CS14, CS15, DiP18b, GWA^{+}07, HMA00, HL00, Hun07, LWF10, MJAK09, OMKdSB11, RV11, SDS00, SYP08, SGA^{+}22, Sul02c, WNZ^{+}17, Weg00, XHL^{+}13, YWC02, YCK03, YCKK03, YaL10, ZZPC06, ZDW^{+}07, ZCXM99]. D&I [Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano23-29, Ano23-30]. D4Science [CCP23]. Damage [BP99, MSR^{+}16]. Dantzig [Nas00, O’L05b]. Darnedest [Dub07b]. Data [AAB^{+}21, AHL^{+}11, AMCH07, ALH^{+}20, AHS11, Ama00, Ano14-44, Ano16z, Ano16-27, Ano16-38, Ano16-37, Azo06, Bak21, Ben09, Ber99, BAD^{+}21, BT01, BCG^{+}99, BKK15, Bre17, BBM^{+}21, Bry11, BCC^{+}99, Bur18, CCSS08, CDL^{+}23, CF03, Cas16, CHP^{+}18, CHB19, CN03, CHC^{+}11, Chi21, CS18, Chr15, CC99, Col18, Com20, CSR^{+}23, Cus13, CF13, Cus14, DVP^{+}17, DPBS16, DBH^{+}02, DM12, DH12, DDC04, Dra00, EUD15, EAF^{+}23, EWN^{+}13, EPHY18, Fei05, FAFX20, Fox03a, FB04, Gal11, GP15, Gor06d, GP21, GM02, GHM^{+}16, GNB^{+}09, HP15, Har04a, HW15, HKW03, Hin12a, HE05, HPMJ12, HSJ^{+}19, IKMK13, JJ15, JMEL08, JON^{+}21, KWT99, KL15, Kar02, KHC^{+}20, KS20, LL18, Ma03, MRU^{+}15, MAFM21, MM13, MO03, MR13, MGD^{+}08, MSR15, NKV99, NC03, Nei08, NCB^{+}05, NSLD99, PARD13, PT14]. Data [PSR^{+}20, PCY14, PLW17, PFS21, PGH^{+}05, Poi10, PTH13, Pos16, QRP22, RV11, RBK02, RP20, SSP06, SYM^{+}21, SDCV10, SKNV03, SRM^{+}07, SCW^{+}17, Shi01a, Shn06, SAK^{+}13, SR13, SPB^{+}20, STG08, Sza11, TAF^{+}18, TR08, TL04a, Thi12b, TNV^{+}02, VGM^{+}09, Vil08, VCGS11, Vor01a, WY12, WCAL14, WLL^{+}14, Wan18, WJ04, Weg00, WZZ11, WGJ16, YKD^{+}03, YBBP15, ZMM03, dSRT16, IK05, Ano16-28, Ano16-29, Ano18-27, Ano20j, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21x, Ano22q]. Data-Compression [SSP06]. Data-Driven [Col18, PSR^{+}20, PGH^{+}05]. Data-Enabled [PARD13]. Data-Intensive [AHL^{+}11, Bry11, HPMJ12, NCB^{+}05, Sza11]. Data-Loading [STG08]. Data-Management [NKV99]. Data-Proximate [JON^{+}21]. Data-Scarce [RBK02]. Database [BO03, Gaa03, IKMK13, Sch01, Smi99f, STM99, SBZB13, TSFG08]. Databases [Cho03, CC03, HBG^{+}20, Lew00a, TDT^{+}22, TSKG03, GGD^{+}05, KBLE15]. DataCenterHub [CHP^{+}18]. Dataflow [EVA^{+}21, PA12]. DataPort [Ano19w]. Datasets [BBW^{+}20, DLLZ19, DLLZ20, Luo12, RRN20, SPW^{+}13]. DataSpace [GM02]. Datoviz [RR21]. Dawn [LA18]. Day [Wil16]. Daydreaming [Hin13a]. Days [Day15b, Sha14, CED^{+}21]. dbX [SDCV10]. Dealing [Bei02, Hin19]. Death [Cyb00b, Day07b]. Deblurring [CO15, NO03]. Debris [JJ15]. Debt [Hin15b]. Debugging [HMB^{+}14]. Decimal [GBPR11]. Decimation [PCY14]. Decision [LLQ18, RSZ^{+}20, RSZ^{+}21, TGU21]. Decision-Making [LLQ18, RSZ^{+}20, RSZ^{+}21]. Decisions [Asr04, MBH14]. Decoding [WW17]. Decomposition [FSED10, SCW^{+}17]. Decompositional [Ste00]. Decompositions [GW15]. Deconvolution [O’L05e, O’L05d]. Deep [NWP19, NSLD99, Hsu06]. Defect [LWT^{+}13]. Defense [HG02, OKS10, Pos07, PS17, PK18, Bar21c]. Deficient [Oli13]. Define [HH99]. Defined [Smi99c]. Defining [COS^{+}15]. Definition [SB00]. Deformations [YCKK03]. Deformed [CCPS12]. Degeneracy [Bei02]. Degradation [LGJ^{+}19]. Degree [Lan04, LGJ^{+}19, Peo20]. Delays [BPMKC21]. Deleted [Smi99f]. Delivering [MWC^{+}16, Wil16]. Delivery [Goo17, QRP22, Sil02]. Dell [Kra03]. Demise [Day11c]. Democratizing [QRP22]. Demultiplexer [GBPR11]. Demystified [Thi13a]. Denoising [HH06, Tas00]. Dense [VCvdG^{+}09]. Density [ZMM03]. Department [LTD11, Leu17, Pos07, Bar21c]. Dependencies [Dru20]. Dependent [BMCC21]. Deploying [JWLG14]. Deployment [BHC^{+}15]. Depth [Wep08].

Derivative [Rei21a, Rei21b]. Derived [PMM^{+}08]. Derminant [BS00b]. Description [PMFM14]. Descriptions [Eng09]. Deserve [Ano16s]. Design [BKB20, Cho08g, DAKM16, Don03, Dub21, For01, Fra07, Gor06a, Ikk16, Jer13, JS99, KFS18, KB07, Kwa17, Lo99, Lun01, MSS09, NRG^{+}17, PAN^{+}16b, QSEQJFH20, SNCM16, SLK^{+}20, She07, SGRK^{+}18, TJ14, WW17, WQT^{+}16, XZL^{+}19]. Designing [BS21, DD07, Dub02, GW15, KWBB22, SGW02, WZS^{+}10, ZFS12]. Designs [FMB^{+}07, SW10]. Desktop [PR01, TS02]. Desorption [KM99]. Destination [WBS^{+}22]. Detect [KSSF11]. Detection [Bai00, DM12, HEH^{+}10, HSJ^{+}19, LM07a, LL19, TMMB18, WCC^{+}19]. Detectives [Gor05b]. Determining [BS00b]. Deterministic [CL12]. Detonations [BPH^{+}13]. Develop [ARAG19, SGRK^{+}18]. Developed [KMB^{+}19]. Developing [AM18, JWLG14, KB09, KB04, MPR18, RRAB06, WD06, YBD10]. Development [ABC^{+}14, BB20, BW14, CCP23, CAS^{+}07, DF21, DGK16, Fox04b, GPMSC20, Gyu99, Hin13c, KMSH10, KVP^{+}16, Läu08, MOBD^{+}22, MBB^{+}09, NCM^{+}14, NC18, Peo20, Pos14, PK18, PBD^{+}11, PMW20, QL19, Saa22, STWK15, SHPL12, SPJ^{+}14, MMG^{+}05]. Developmental [SLK^{+}20]. Developments [SS06]. Device [HRRS09]. Devices [KL10, YLZ17, Zhu02]. DEVS [Zei17]. DFT [Lew10]. Diagnosing [DRA11]. Diagnosis [Gig00]. Diagnostic [WZS^{+}10, WCC^{+}19]. Dialog [FL21]. Diamond [CJ16]. Diamond-Like [CJ16]. Dictionary [GDH^{+}23, WLL^{+}14]. Did [Day18b, Hin20b, TL08a]. Diego [LC09]. Difference [Bar11, HLYQ19, Smi01b, UZC^{+}12]. Differences [O’L05g, SRM^{+}07, WCP17]. Different [AK04, SL18, Wep15]. Differential [GWW09, JWEK06, JHJ01, Lud13, MSL^{+}07, MW14]. Difficult [Hin15a].

Diffraction [Tre99]. Diffuse [SGA03]. Diffusion [Mal07b, OHK23, WLCD01]. Digging [Thi12b]. Digital [Ano13d, Ano14e, Ano14f, Ano14g, Ano14h, BLT^{+}23, Cho08c, CSR^{+}23, Gar06, Gor05b, HBG^{+}20, LVWK02, Lew99a, ML02, Mas06, Nei08, Sza99, Tha08b, Thi15a, Toh08]. Dimension [ARO^{+}11, GYL^{+}17, Nob00b, SL03, Vor01a, dSRT16]. Dimensional [BBW^{+}20, CN03, CS18, GWMG04, HKW03, Maj03, MB99, NC03, Shn06, SR13, dSRT16]. Dimensionality [JG12, Vil08]. Dinosaurs [Lew02b]. Direct [Bet99, CCSS08, CXC^{+}20, HSG03, O’L05c]. Directed [PL02]. Directed-Energy [PL02]. Direction [OM03]. Direction-of-Arrival [OM03]. Directions [DSSS05, GB20]. Directives [BBG^{+}01]. Disabilities [SLK^{+}20]. Disaster [WPM^{+}12]. Disasters [FCT^{+}10]. Disbelief [Sul01d]. Disciplinary [Biz16, PV00]. Disciplines [Cyb00b]. Discontinuous [YLZ^{+}19]. Discourage [Bot16]. Discovering [Chi21, LSPN21, MSM13]. Discovery [CJ16, ESO08, JWLG14, LFK^{+}19, Lun01, Men18, QRP22, SSP06, TWE14, TCD^{+}14, WBB^{+}20]. Discrete [BPMKC21, Cas16, Gra07, JCC^{+}10, JG13, NW15, Sch21, Sch23b, Wai16]. Discrete-Event [NW15]. Discrete-Time [BPMKC21]. Discretization [YLR02]. Discriminant [FOdLVF^{+}11]. Discussions [Can99]. Diseases [Dal99]. Disk [ZMM03]. Dislocation [CCPS12]. Disparate [BHC^{+}08]. Dispersion [PAF08]. Display [Toh08]. Displays [WJ04]. Dissemination [MK10]. Dissimilar [Has12]. Dissipation [PAF08]. Distance [Day07b, Luo12, SSCN11, Thi12a, WLCJ12]. Distance-Based [Luo12, WLCJ12]. Distinguished [Ano14-43]. Distribute [Reb99]. Distributed [DPP^{+}01, DPG^{+}12, GHT^{+}10, MPP14, MKM^{+}14, MSL02, PE09, PBD^{+}11, SKA^{+}02, YKD^{+}03].

Distributed-Memory [DPP^{+}01]. Distribution [BBM^{+}15, SM17, XLLJ04]. Distributions [Tho01]. District [SDA20]. District-Level [SDA20]. Diverse [BLO^{+}22, Deb18]. Diversion [LWF10]. Diversity [Akl18, Ano23m, Ano23n, BTL19, Cor07, HGA23, Leu17, Leu20, SL18, WBP^{+}19]. Dividing [SB00]. Diving [NSLD99]. Djange [Dub07c]. DLR [ERS^{+}03]. DNA [Lew02c, Mye99, Rei02, SBH^{+}00, WCH12]. DO [Day12f, Bar11, Day10b, Deb18, Has08, HGA23, Lew00a, Rag06, SHPL12, Sul02c, Thi13d]. Docking [NLGNJ13, WCH12]. Docs [Day09d]. Document [KHE13]. Documentation [PSSP15]. Documents [LVWK02, O’L14, SD11]. DoD [MM16]. DOE [AHL^{+}11, BLMR21, BHV21]. Does [Deb18, MMTD^{+}17]. Doing [Dar18, Thi09a]. Domain [CPdlF^{+}12, DPBS16, Hin18c, JWLG14, Läu08, Oug03]. Domain-Driven [Läu08]. Domain-Specific [Hin18c, JWLG14]. Domains [RBK02]. Donald [Day10b]. Done [Alt10, Dub15b]. Dongarra [HH22, Par22]. Don’t [Tho99a]. Doors [DiD03]. Dorrit [Day09c]. Dot.com [Smi00d]. Double [DMA^{+}21, KCPFT02, OOB17]. Down [Sul02b]. Downdating [O’L06e, O’L06g]. Downward [Cho04]. Drag [LQZL19]. Draper [Day10b]. Drawing [Tof08, Tof09a, Tof09b]. DRC [Rei21a, Rei21b]. Dream [DG12, Hin17b, Sic09]. Drinking [Tha08b]. Drive [Ano23m, Ano23n]. Driven [APC^{+}19, Cas16, Col18, Läu08, NCM^{+}14, NC18, PSR^{+}20, PCY14, PGH^{+}05]. Driving [LL19]. Droplets [YZZ04]. Drops [BW01, MPR18]. Drug [KFS18, Lun01, Sch99]. Drugs [WZS^{+}10]. DSP [TJ14]. Dual [KEF07]. Dual-Lookup [KEF07]. Dual-Modality [KEF07]. Dubois [Dub04]. During [Leu20, RSZ^{+}20, RSZ^{+}21]. Dust [CF99b]. DVDs [Kra03]. Dwarf [DMA^{+}21]. Dye [KSW^{+}12]. Dye-Based [KSW^{+}12]. Dynamic [AGM^{+}00, BWC01, Bec15, DLW^{+}19, DBH^{+}02, Mar99b, VN99, VKN99, YLZ^{+}19]. Dynamical [O’L07a, O’L07b, WB03]. Dynamically [DGR^{+}05]. Dynamics [BST^{+}13, BH02, CJTH^{+}13, CFCD04, CS14, CS15, GLS07, HSG03, Hym05, IBPV03, KSP12, KSM^{+}17, LNI^{+}19, MB17, MB20b, Muz19, NKV99, PAF08, PZJS10, SDA^{+}14, Sch20, Sch23c, TL08b, VB22, Vor01b, ZZS^{+}19]. Dystopia [Bei10a].

e-Infrastructure [MTG^{+}12].

E-infrastructures [SPS15]. E-Reader [Day14c]. E-Science [Fox02a, KE05, RTSS14a, RTSS14b]. E3SM [WSY^{+}22]. Ear [yFZDY13, ZDW^{+}07]. Early [BHV21, ERF21, PHW^{+}21, Pos16, Sha14, Wil18]. Earth [GVB15, HSP^{+}23, KL15, Run05, Asr04, Bet17, DRR^{+}04, GPZ^{+}04, HDB^{+}04, KB04, LM07b, MGZ00, Run00, SCW^{+}17, WBS^{+}22, Zeb00]. Earthquake [CHC^{+}11, DPG^{+}12, EI11, FCT^{+}10, Gor07b, HEB^{+}11, HS12, JRP^{+}17, KMM^{+}11, McK11, PSR^{+}00, RF12, TB11, UZC^{+}12, WPM^{+}12, YMK11]. Earthquakes [Day11d, STHR12]. Easier [Dav12]. Eastern [Ale18]. Easy [Dub08c, Sul02a, Tho99c, Vor01b]. Eau [Smi99b]. EC2 [JRD^{+}13]. ECG [WGJ16]. Eclipse [WD06]. EcoG [SES^{+}11]. Ecological [GYF^{+}10]. Economist [Fla17]. Economy [Wes03]. Ecoregion [HH99]. Ecosystem [Bar21b, DBCN03, PGH11, Tow18, WCGB05, TBVP^{+}21]. Eddy [DJS13, YWMM04]. Eden [SOH13]. Edge [Ano23l, KSSF11, Ano19a, Ano19b, Ano20b, Ano20c, Ano22d, Ano22e, Ano23a, Ano23b, Ano23f]. editable [Hin18d]. edited [Lov04]. Edition [Dub04, Seg99]. Editor [Ano07, Cyb99b, Sul99a, NLV99, Ale13, Asr04, BLMR21, Ber99, Cho03, Cyb99a, Dub07e, Dun09, Eth01, Got06, Kar02, Kax01, Kup03, MS99, Pos04a, Pos04b, Pos07, Qua18, Run00, Run05, Sul09a, Win06]. Editor-in-Chief [Cyb99a]. Editorial [GS13a, HP14b]. Editors [Ano22-68, Cyb00a, Cyb99c, Cyb00b, Kal99, Kil99, MBS^{+}00, MHDM99, Sul99b, Sul00c, Sul00b, Sul00a, BC05a, AM15, Ano04b, Ano22-59, Ano23-50, BC99, BS06a, BC05b, CE14, CN03, CZ07, CLZ13, Cho05c, Cho05d, CF13, Cyb00c, CS01b, DS00, FC09, FF03, For16b, HP14a, HS03, HG02, HP04, KS02, MF16, MR06, Mem15, NC03, NL99, PT14, PV00, PS02, RC01, SS02, Sul01a, Sul01b, Sul03b, Sul03c, Sul03d, Sul03e, Sul04a, Sul04f, Sul04b, Sul04c, Sul04d, Sul04e, TA05, TX07, Thi05, TB99, TP04, TM00, VP04, VN99, Wai16, WR00, ZW19]. Education [Adl20, Ano16c, Ano16d, Ano17b, Ano18h, Ass00, Bäc07a, Bei12d, BB06, BERT09, But99, Day06b, DdS23a, DdS23b, Don99, DL23, FGP99, Gor13, GL20, HL00, Hu07, JHJ01, JPMG08, KMB^{+}08, Lan06, LM08, Mar17, Mas06, MS07, NMCM22, PMK^{+}08, Peo20, PP20, Roo06, SDA20, SDD^{+}08, Tan21, Thi12a, TX08, TMC^{+}13, Tre99, YRT^{+}00, YMLJ06]. Educational [Chr99, MHDM99]. Effect [Chr99, Don02, KS06]. Effective [KC19, PTML11, SNCM16, SGRK^{+}18, TS02]. Effectively [Luo13]. Effects [PQQ20, ZMM03]. Efficiency [BMCC21, MM16]. Efficient [CLC03, Hoe10, RK05, SH10, SSP06, SES^{+}11, Yav06, ZJW08, ZLTX19, dKCAY00, vdWCV11]. Efficiently [CPdlF^{+}12]. Effort [Fom15, Got02b]. Efforts [MB20a]. EICs [SCBT18]. Eigenpairs [GBDW04]. Eigenstates [Nob02b]. Eigenvalues [O’L05a]. Einstein [KF03, STTV05]. EJB [Lau05, Läu06, LTG07]. Elastic [MJAK09]. Elastoplastic [O’L04a]. Electrical [CB02]. Electricity [Ass00]. Electro [Roh10]. Electro-Mechanical [Roh10]. Electrocatalytic [VWP12].

Electrodynamics [Hei20]. electrograms [SOV^{+}13]. Electrohydraulic [YMHQ19]. Electromagnetic [JLYL19, LFN^{+}11, XZL^{+}19]. Electromagnetics [KLS01]. Electron [KHC^{+}07]. Electronic [BJ02, GBDW04, GS03, Kyr08, Lew99a, NG20]. Electronic-Textiles [NG20]. Electrons [SDA^{+}14]. Electrostatic [CLC03]. Element [BKB20, Bas14, BS21, BFH21, GGJD23, IHL^{+}02, JXY^{+}19, LFC01, WWJH20]. Elementary [Ono01]. Elements [Ara99, BGHR06, Bri23, JC02, O’L05g, Pos16]. Eliciting [Rao16]. Elliptic [Don10]. ELM [LPCY19]. Elusive [Lew01a]. ELVIS [Tou03]. Email [Day15c]. Embedded [JR10, Lan19, NW15]. Embraced [RTSS14b]. Emergencies [Par16]. Emerging [Dec15, VB22]. eMinerals [BDCT05]. Empirical [SCW^{+}17]. Empowering [NMCM22]. Emptor [DC04]. Emscripten [Zak18]. Enable [DAKM16, FL21, HC99]. Enabled [BDCT05, ESO08, GK18, GWA^{+}07, JON^{+}21, PARD13, SKL10, TGP13, Thi10, VSMD^{+}09, VCGS11]. Enabling [Can99, Thi02]. Enacting [Smi16]. encapsulator [PLH^{+}18]. Encoding [LPY18]. End [FF03, LAY04, Liu11, Mes15, RP20, TW17, TFS17, YBD10, YLZ^{+}19]. End-to-End [RP20, YBD10]. Endeavor [BLT^{+}23]. Endless [Day15b]. Energy [AMS14, DGJ^{+}08, Eis17, ECK^{+}15, MB99, PL02, PSA14, PMW20, SW10, SM17, SdPV21, SPJ^{+}14, SSK02, WG15, YYL^{+}18, dKCAY00]. Energy-to-Solution [Eis17]. Engage [WMB20]. Engagement [Den16, PBSS14]. Engine [SP18, GHK^{+}08, SKP^{+}10]. Engineering [Adl20, Ama06, Ano14y, Ano14-49, Ano18z, Ano19y, Ano20-36, Ano21-48, BT17, Bar20a, Bar20b, BPW^{+}20, Bar21c, BCH^{+}09, BSD07, Bot16, BERT09, Car09a, Car09b, Car12, CHHB13, CE14, Car16, CHC17, Cas16, CC03, Col18, CW20, Cos22, DSPY05, Day06a, DMXR^{+}14, Dun09, DST^{+}09, EJ09, ESO08, FLV^{+}09, FM13, GPL09, Gor07c, GK18, HEB^{+}11, HLS^{+}16, Her22, HGA23, JRP^{+}17, JH18, KSM11, KMM^{+}11, KRH^{+}99, MTBG^{+}22, McK11, MK10, MV20, Mes15, MCGA22, Mil10, PARD13, PA21, PZJS10, PV00, Pos07, Pos09, Pos10, Pos13, PK15, PAN^{+}16a, PNL^{+}16, PAN^{+}16b, PG17, Sch23a, STWK15, SSK13, SKL10, SPB^{+}20, TB11, Thi15c, TP04, VP04, WL09, YMK11, Ano03, Ano05a, Ano12b, Car09a, CHC17, RHC^{+}23, Ano22f]. Engineers [Bar20b, CHHB13, CENT22a, CENT22b, Cho12, CMK22, GAB^{+}22, Guo23, MA11, Tho12, Wri10]. Engines [Gor07a, HSP^{+}23, MGCBI17, PA12]. Enhance [BBN03, BBD^{+}13, DD23]. Enhanced [HEH^{+}10, MSR^{+}16, Tou01, YKD^{+}03]. Enhancement [YLZ^{+}19]. Enhancing [BNNM04, Den16, EPHY18, PK15, PG17]. Enrollment [MMTD^{+}17]. Ensemble [GVB15, PSS20, SCW^{+}17]. Ensembles [ALH^{+}20, BH02]. Enthusiasm [Sul08a]. Entrepreneurs [Lew00c]. Entrepreneurship [WMB20]. Entropy [Jq19, PAF08]. Entry [CMK22]. Enumerations [Gut01]. Envelopes [GARS^{+}20]. Environment [Asr04, Bar21b, BLT^{+}23, CAP^{+}10, CCP23, CAS^{+}07, CS15, GWA^{+}07, GHKZ17, Har04a, Har23, Hun07, Ong02, PTML11, Par16, PL02, RPBE12, Shi01a, SdPV21, WQT^{+}16]. Environmental [GLS07, MKJ07]. Environments [DGR^{+}05, Fox03b, LRRK00, MHK^{+}06, PAN^{+}16a, PNL^{+}16, VSMD^{+}09, Weg00]. EPCC [WP22]. Ephemera [Bei10b]. Epidemic [O’L04d]. Epidemiology [Gor10]. Epileptic [SSG16]. Epiphany [PK18]. ePub [Ano13e, Ano14i]. Equation [JWEK06, Moy06, Osk07, RK05, Sch18, SFSK01, STTV05, XKG05]. Equation-Free [XKG05]. Equations [Bog03, GWW09, HT99, JHJ01, MSL^{+}07, O’L05f, Sil00, Ter11, Lud13]. Equipment [Tof08, Tof09a, Tof09b]. Equitable [Deb18]. Equity [BPW^{+}20, GPMSC20, HGA23, Les16, SDA20]. Equity-Based [GPMSC20]. Equivalence [Rob04, ZLW^{+}19]. Era [Bak21, BHF^{+}08, DST^{+}09, LL13, Mei03, NdSS17, Thi13b]. Erlang [Vin12]. Erratum [Ano18i, Ano18j, Rei21a]. Error [Iac21, NSR10]. Errors [O’L05e]. eScience [MOBD^{+}22]. Essay [HG00, LSV^{+}07]. Essays [Cre99, KRH^{+}99, Lan99]. Essence [Yas17a]. Essential [HLT09]. Estimating [AB03, DKCL14, Rus01b]. Estimation [FOdLVF^{+}11, KLMS99]. eTextBook [LPB13, LPB15]. Ethanol [MPR18]. ETSF [MTG^{+}12]. Eureka [Smi99d]. Europe [Ale18]. European [EAF^{+}23, Els21, FM02, Kal19]. Evaluate [GYF^{+}10]. Evaluating [yFZDY13, MPP14, PD02, Ron14]. Evaluation [ACQ^{+}20, NA07, Nan11]. Evenly [LM07a]. Event [Cas16, Gra07, Lew01b, NW15, PRCL^{+}22]. Every [Sul04d]. Everybody [Day18b]. Everyone [VGM^{+}09]. Everything [Yas17b]. Everywhere [O’L05e]. Evidence [Bar21a, Fox18, SGRK^{+}18, TMMB18, Tec18, TKM^{+}18, vGDS18]. Evidence-Based [SGRK^{+}18, TMMB18, Tec18, TKM^{+}18, vGDS18]. Evolution [BBM^{+}15, BHV21, Che16, Cho06c, CBS14, DVR^{+}19, FL99, For01, New00, VP04]. Evolutionary [RRN20]. Evolvable [For01]. Evolving [Ano23o, Cho05c]. EWD [Sul02a]. Exa [BAD^{+}21]. Exact [RLRML04a]. Exaggerated [Day07b]. Examining [BZL^{+}07, PQQ20]. Example [Lud13, Pos16, Wei21]. Examples [TNR^{+}23]. Exascale [Ano15-37, BRS22, DGK19, DS23, DMT^{+}21, GS13a, GS13b, JR10, JRP^{+}17, Kal19, KES^{+}22, KS13, KLQ19, LMPV13, LA18, NdSS17, SBW^{+}19, WSY^{+}22, WP22, MB20a, Mes17]. ExCALIBUR [BBWW22]. Excellence [Kad04, NRG^{+}17, NdS17]. Exceptionally [Thi11b]. Excited [Mor15]. Exciting [Day07c]. Execution [BHL99, MSD10, SB00]. Executions [GPC08]. Exemplifying [PQQ20]. Exercise [KWBB22]. Exercises [DM04]. Exoskeleton [YMHQ19]. Expand [WCP17]. Expanding [Her22, Pap16, Thi07]. Expect [Day09b]. Experience [CCP23]. Experiences [BB06, Bot16, LTD11, LUMM14, Lof22, MS22, Taj10]. Experiment [Dav12, Gor07d, Sch01, Ste02, Var08]. Experimental [CBF^{+}22, GHT^{+}10, Guo12, YYG^{+}19]. Experimentalists [DR05c, DR05b, Don06a, Don06b, RD05a, RD05b, DR05a]. Experimentation [PGH^{+}05, SGRK^{+}18]. Experimenter [Ben04]. Experiments [Ben04, LBS14, Liu11, MAFM21, MTG^{+}12, MGD^{+}08, Rei13, SZM^{+}13]. Expert [Sma12]. Expertise [MBH14, Mes15]. Experts [Ano16-47, Ano16-48, Bur18]. Explaining [GLS07, Jer13]. Explicit [BPMKC21]. Exploitation [PTH13, RLHGA^{+}13]. Exploiting [BBD^{+}13, DSK15, EGFL12, Muz19, XZL^{+}19]. Exploration [BHF^{+}08, BBW^{+}20, CO15, CHM^{+}20b, CHM^{+}20a, DVP^{+}17, DH12, DMT^{+}21, EPHY18, JMEL08, TDT^{+}22, Thi11b, TNV^{+}02]. Exploratory [GM02, KHG^{+}23, LSN20, WCP17]. Explore [Ano23o, LPV00]. Exploring [ALH^{+}20, BHC^{+}15, Don02, Luo12, MKH^{+}23, MB99, Ree16]. Explosions [BPH^{+}13, HMS^{+}00, Ott16]. Exponentials [O’L04b, Rus02]. Export [EHG01]. Exposing [PMM^{+}08]. Exposition [Hah04]. Expression [Kir03, OMKdSB11, YCK03]. Extend [GDDR16]. Extended [KL10]. Extending [Bas02, DY99, Dub00].

Extensibility [BKS15]. Extensible [ASM^{+}14, KSM^{+}17]. Extension [Ron14]. Extensions [EGFL12]. External [CFCD04]. Extinction [New00]. Extraction [GMPR11, KS20]. Extractors [JG12]. Extremal [Boe00]. Extreme [Che19, Cus14, Dar18, DTL^{+}17, LWG19, NWP19, PT14, Sza11, TWE14, VM15, WHW18]. Extreme-Scale [VM15]. Extremes [TM15]. Eye [CT00, KBPW15].

Face [FOdLVF^{+}11, OMKdSB11, Pos13, SMM^{+}11, SYP08]. Facebook [Day14f, FAFX20]. Faces [Got02b]. Facial [OMKdSB11, WLCJ12]. Facial-Expression [OMKdSB11]. Facilitate [MVUSK14]. Facilities [Pap16, QRP22]. Facility [AGM^{+}00, SW10, LBS14]. Factor [Els21]. Factorizations [O’L06e, O’L06g]. Factors [DCWH07, DKWL17]. Failure [GB20, MPRU23]. Faint [Bar12]. Family [Gor05a, Roc00]. Fanfic [Day18f]. Fans [Got02a]. Fantasy [Smi99c]. Far [UGV11]. Far-Infrared [UGV11]. Faraday [Day16c]. Farewell [Bei12b]. Farms [Day17d]. Fashion [Day17a]. Fast [Ara99, BB07, BS00d, KLMS99, MBS^{+}00, O’L05f, RcK99, RRAB06, SMF^{+}23, Tho99c, Zak18, Clo15, DR05c, DR05b, DR05a, Don06a, Don06b, RD05a, RD05b]. Fate.Com [Smi99c]. Fault [Jq19, JRP^{+}17, LCC^{+}19, ZZC^{+}19, GGD^{+}05]. Fault-Tolerant [LCC^{+}19]. Faults [PSR^{+}00, WCC^{+}19]. FEA [JLYL19]. Feasibility [FPRK16]. Feature [BAD^{+}21, JG12, TNV^{+}02]. Featured [Ano20-72, BPW^{+}20]. Features [Bur99, DTA21, GPMSC20, GvdWT07, PMFM14, Shi01a, Shi03]. Federated [DMXR^{+}14]. Federating [BDS13]. Feedback [YMHQ19]. Feeling [MPRU23]. Fellowship [BLMR21, BHV21, GAB^{+}22]. Female [JH16]. Fermi [HKB12]. Fermions [Ale15]. Few [MW14]. Few-Body [MW14].

FFT [Roc00]. Fiber [JXY^{+}19]. fibrillation [SOV^{+}13]. Fiction [Day16f]. Fictional [Day17d]. Fidelity [HEH^{+}10]. Field [HHZK01b, HHZK01a, Jef22, PMW20, ZMM03]. Fields [Ma16, SÜP^{+}11, SV14]. Fifteen [BKK15]. Fight [HP20, Wes21]. File [Ben09, IKMK13]. Filesystems [BBM^{+}15]. Filled [GGJD23]. Filtering [FL05]. Filters [Don06a]. Fin [Thi16]. Final [Sul05a]. Finally [Aya14]. Finance [Far99, NL99, Sha99, Vir16, Wep15]. Find [Day17c, Don10]. Finding [MGFRL^{+}12, SW10, Tou02a]. Findings [JH16]. Finite [Ara99, BKB20, Bas14, BGHR06, BFH21, Bri23, GGJD23, JXY^{+}19, LFC01, NA07, O’L05g, UZC^{+}12, WWJH20, YLR02]. Finite-Difference [UZC^{+}12]. Finite-Element [BKB20, LFC01]. Finite-Element-Based [GGJD23]. Finite-Volume [YLR02]. FiPy [GWW09]. Fire [Tha08b]. Firedrake [BFH21]. Fires [HMS^{+}00]. Fireworks [Don02]. First [BB06, BVT^{+}21, CHJC05, Day08a, Day14c, HWPS16, PRCL^{+}22, Slo16, Car09a]. First-Generation [Slo16]. First-Principles [HWPS16]. Fisher [FOdLVF^{+}11]. Fit [Ano17x, Ano17y, Ano17z, Ano17-27, Ano16y, Ano17-28]. Fitting [Don10, LWG19, O’L04b, Rus01a, Rus01b, Rus02, Rus03, SM17, TR08, WS99]. Five [KHS09, Poi10, Shi99]. Fixed [MM16]. Fixed-Wing [MM16]. Fixing [Day10b, Day15c]. Flagship [Sor19]. Flair [Wep08]. Flame [Tow09]. Flames [MRKK17]. Flash [RCD^{+}00]. Flashes [RCD^{+}00]. Fleet [MSB^{+}14]. Flexible [DKCL14, GHKR11, KB07, LVLA14]. Flight [ACS15, Sim13]. Floating [Bai05, GGJD23, PPE00, TM14, YLZ^{+}19]. Floating-Point [Bai05, TM14]. Flow [CCPS12, EPHY18, FPRK16, Ged16a, GvdWT07, GIF^{+}12, HF04, JMFJ01, Jon19, KSW^{+}12, Ma16, NW13, RSC^{+}14, TGEA09, VCGS11, WT12, YWMM04, ZSM^{+}22]. Flowers [Sul05b]. Flowfield [HWPS16]. Flowfields [MM04]. Flows [DJS13, FMKS08, GF04, LUMM14, MP09, NTW07]. Fluid [Ben04, CFCD04, Jon19, JCPS14, KSP12, KSM^{+}17, LUMM14, LWSK07, LCY^{+}04, Lui06, MB17, MB20b, Ork09, Sch20, Sch23c, SFSK01, SKC05, VB22]. Fluid-Structure [LCY^{+}04]. Fluids [Bry99]. FluoRender [WHH22]. Focus [Ano02b, Ano14l, Ano14j, Ano14k, Ano14m, Ano15j, Ano15k, Ano15l, Ano17d, EVL^{+}17, For00, For01, Lew00a, Lew00b, Lew00c, Lew01a, Luo12, YCD^{+}21]. Folding [Han03, Jav12, SK01]. Following [O’L06c]. Force [JLYL19]. Forcing [MRHP23]. Forecast [Gor07b, SS09]. Forecasting [Lum07, STHR12]. Forecasts [DWC^{+}11, Gan02, KILZ13, ZQY^{+}11]. Forefront [GLS11]. Forests [DBCN03]. Forever [Smi99f]. Forgeries [Gor05b]. Form [Yas17b]. Formal [FGG^{+}22, KNKP14]. Formalism [GW15]. Formalism-Based [GW15]. Format [Ben09, Poi10]. Formation [CDKF15, SETK05, SNTL13]. Formats [CY00]. Former [SCBT18]. Forth [Nob00b]. Fortran [Mol12, CRDO16, DNG07, DY99, Fos17, GRE99, JPE20, KAČ^{+}22, Pad00, PMM^{+}08, Rei03, RMX12]. Fortranning [Mol12]. Forty [WG15]. Forum [KFMG20]. Forums [LSPN21]. Forward [Cho05f, Her21]. Foster [For99]. Fostering [Lan19, Tur14a]. Four [GSK^{+}23, HGA23]. Fourier [DR05c, DR05b, DR05a, Don06a, Don06b, RD05a, RD05b, CS18, Cor07, Mus20, Tre99]. FP [DPP^{+}01]. FP-LAPW [DPP^{+}01]. FP7 [Ale18]. FPGA [BMP^{+}06, BCC^{+}09, HGV^{+}08, NLGNJ13, SDCV10, SG10, TJ14]. FPGA-Based [BCC^{+}09, HGV^{+}08, NLGNJ13, SDCV10]. FPGAs [AAAH^{+}16, LHZ21, VMK20]. Fractal [ARO^{+}11]. Fracture [BP99, Han05, RcK99, VN99, VKN99, XHL^{+}13]. Frame [Wil01]. Framework [ALH^{+}20, APC^{+}19, CBB06, DL00, EP10, GBDW04, GYL^{+}17, GRS08, GSB^{+}12, HC17, Ikk16, KRR^{+}12, KFS18, KSM^{+}17, LVLA14, McK11, PGC21, QRP22, Ram18, RPBE12, RPEB14, SBB^{+}15, SNCT13, SMSC22, Sto09, VCGS11, ZFS12, HDB^{+}04]. Frameworks [HMB^{+}14]. Franca [WHH22]. Frankenstein [Sul03c]. Free [Shi00c, THGS07, XKG05, dKCAY00]. Free-Energy [dKCAY00]. Freedom [Ano15-44, Lew00b]. Freestyle [TR08]. Frequency [CPdlF^{+}12, PAN^{+}16b]. Frequency-Domain [CPdlF^{+}12]. Fresnel [Tre99]. Friendship [Sim13]. Frisch [Sch23c]. Front [Ano13f, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano15r, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano17i, Ano17e, Ano17f, Ano17g, Ano17h, Ano18o, Ano18k, Ano18l, Ano18m, Ano18n, Ano19h, Ano19i, Ano19k, Ano19l, Ano19m, Ano20-32, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano23p, Ano23q, Ano23r, Ano23s, SG10, Ano19j]. Front-Side [SG10]. FRONTERA [GARS^{+}20, CED^{+}21, WNDV21]. Frontier [Gue18, Kus07, SP18]. Frontiers [HP14b, HJLH03, Pos04a, Pos04b]. Fuel [SP18]. Fuel-Engine [SP18]. Full [HLYQ19, Nob02a]. Fullerenes [SMC01]. Fully [ZZC^{+}19]. Fun [Day11e]. Function [BHC^{+}08, Don10, Iac21, KEF07, LWG19, NSR10, Rus01a]. Functional [GW15, Hin09, Kar99, LT09, MB07]. Functions [LTD11, MAC08, Rei21a, Rei21b, Rus01b, Rus02, Rus03, Tho99b, Tho00]. Fund [Ano23-30, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano23-29]. Further [FIG^{+}23]. Fusion [ECK^{+}15, HSJ^{+}19, Ma16, TWE14, SMM^{+}11]. Future [AHL^{+}11, Bec15, CENT22a, CENT22b, Cho07c, Cho08b, CW20, Dau99, Day06c, Day12a, Day19a, DSSS05, Dub07f, Dub07d, EKLY07, GB20, Got15, Got16, Hin13b, JH18, LVWK02, LL18, Mar17, MDW^{+}22, MCGA22, MBB^{+}22, Pos11, Rei03, SPS15, SACR21, Smi00d, Sul04f, Thi11a, Thi13a, Thi15a, Thi15b, TW03, WP22, YCD^{+}21, Zha11]. FutureGrid [JRD^{+}13]. Fuzzy [CS01a, DKCL14, Fra07, GYF^{+}10, LLQ18, ZLW^{+}19]. Fuzzy-Neural [Fra07]. FVDM [HLYQ19]. FVTD [SWPB00].

G [GLS11]. Gaining [Mes15]. Galaxies [Jon15, KCPFT02]. Galaxy [ASLK22, PRCL^{+}22]. Gamble [PPE00]. Game [BB20, Les16, TMC^{+}13]. Game-Based [Les16]. Gaming [Day12b]. Gap [Jef21]. Gaps [CMK22]. Garmin [Tou02a]. Gas [Par16, Sch20, Sch23c]. Gateway [GDH^{+}23, HBD^{+}23, MKH^{+}23]. Gateways [DdS23a, DdS23b, DL23, GSK^{+}23, KHG^{+}23, OHK23, WDC18]. Gauging [Day08a]. Gauss [Cor07, Nob00a, Ome06]. Gaussian [Bal99, Nan11, Wan23]. Gay [Wri16]. Gel [dA03]. Gen [Smi00c]. Gender [Les16]. Gendered [Bot16]. Gene [YCK03]. General [MW11a, SSC18, Wat21, YLR02, Zei17]. General-Purpose [SSC18]. General-Relativistic [MW11a]. Generation [BBWW22, BFH21, CHJC05, Ged16a, HP15, MAC08, McM09, MRHP23, NMCM22, RGD13, Slo16, Smi00c, SPJ^{+}14, Tha14, Thi04, Thi15c]. Generic [Kyr08, MM18, SL99, IK05]. Genes [CAS^{+}07]. Genetic [Bea00, FL99]. Genome [Mye99, STM99]. Genomes [Mis02]. Gentoo [Thi04]. Geodesic [RRH^{+}02, WLCJ12]. geodynamic [Mat05]. Geodynamics [HPC20]. Geodynamo [GC00]. Geographic [Che10, WCC^{+}02].

Geometric [Lui06, ZZPC06]. Geometries [FGR^{+}07]. Geometry [Chi21, DAKM16, MPP14, O’L12, RLRML04a, SWDR22]. Geometry-Centric [DAKM16]. Geophysical [IHL^{+}02]. Georeferenced [AAB^{+}13]. George [O’L05b]. Geoscience [Har04a, HEH^{+}10]. Geosciences [LWF10]. Geoscientists [Tha14]. Geospatial [MPP14]. Geostatistical [SMF^{+}23]. Germany [Day12c]. Gestures [YHWY05]. Get [Ano14t, Ano15s, Ano15t, Ano15u, Ano15v, Ano16s, Ano20-33, Ano20-34, Ano20-35, Ano21-42, Ano21-43, Ano21-44, Ano21-40, Ano21-41, Ano21-45, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano23t]. Gets [Shi99]. Getting [Gra09, KSB07, MAFM21, O’L05g, Wil06]. Giant [LWT^{+}13]. Gide [CO15]. Ginzburg [Osk07]. Girls [Sca16]. GIS [GYF^{+}10, JJZC10, Shi07, WPM^{+}12]. Gismo [BCA^{+}00]. Giving [GAB^{+}22]. Glacier [BZL^{+}07, Bri23]. Glaciers [TAHP23]. Glass [YZZ04]. Glasses [BHKW03, DCC10]. Glast [BCA^{+}00]. Glimpse [Hin13b]. Glimpses [Sin18, Ano18j]. Global [BB07, BBM^{+}15, CNO99, ECK^{+}15, KE05, LAY04, MRNT17, SH10, TR08, TLD02, YCD^{+}21]. Globally [KFMG20, XBK10]. go [Ano14-55, Ano14-56, Ano15-48, Ano15-49]. Goal [SBW^{+}19]. Going [Sul02c]. Golomb [Mem16]. Gone [FL21]. Good [Bei10c, Dub22, Dub99, MBB^{+}22, Poi10, Sul04b]. Goodbye [Dub08b]. Google [Cha08]. Googol [Cha08]. Got [Dub15b]. GPGPU [HKB12]. GPS [Tou02a]. GPU [EH22, BFF12, CJL^{+}18, CG09, DJS13, EKCS12, FKB^{+}13, FPRK16, Gra09, Kel10, MMG08, RR21, SES^{+}11, TGP13, VGD^{+}11, WQLZ18, Wei11, WCH12]. GPU-Accelerated [FPRK16]. GPU-Based [WCH12]. GPU-Enabled [TGP13]. gpu.js [SSC18]. GPULib [MMG08]. GPUs [Ale15, BKB20, Kin12, SSC18, STH22, Wei21]. Grabbers [Wil01]. Grace [EH22]. Graders [Ste14]. Gradients [Tof09b]. Grading [KC09a, KC09b, KC09c]. Graduate [BLMR21, BHV21, GN08, LC09]. Grain [Bil00]. Grained [BMCC21]. Grande [Fox03d, WSC^{+}04]. Granular [Saa09]. GRAPE [Mak06]. Graph [CS18, Coh09, Har23, HB08, Wan18]. Graph-Based [Wan18]. Graphene [Ste12]. Graphical [ATRA00, CO15, UM08, WOAEAG10]. Graphicality [Clo15]. Graphics [Ano20a, Ano20z, Ano20f, Ano21a, Ano21b, Ano21l, Ano22k, Ano22l, Ano22m, Ano22n, Ano23d, Ano23e, BLT^{+}23, BkCP22, CF99a, CY00, Che03, HRWS06, Hun07, KSP12, SDS00, SJDV09, Wat21, Zhu16]. Graphs [Dru20, JG13, QSEQJFH20]. Gravitational [ACF18, BVT^{+}21, LL13, SV14, SR12]. Gravity [KNKP14]. Gravity-Induced [KNKP14]. Gray [Jq19]. Great [Lov04, Smi99f]. Green [Hem10, MSD10]. Grid [BB07, CCJ04, FG01, Gor07d, GWA^{+}07, HF04, SYP08, YLCZ05, BC05a, BC05b, BDCT05, CHJC05, Fox02b, Fox03a, Fox03b, Gor07b, JRD^{+}13, Zhu02]. Grid-Enabled [GWA^{+}07, BDCT05]. Gridlock [FLV^{+}09]. Grids [BW06, CB02, FG01, Fox04a, RRH^{+}02, SAK^{+}13, WB03, CB02, Fox03c, WCGB05]. Gröbner [RLRML04a, RLRML04b]. Gross [STTV05]. Ground [Cho05e, For02, GL20, JRP^{+}17, LGW^{+}17, MR13]. Ground-Range [MR13]. Group [BB06, Cos22, Day11c]. Grow [Yas17b]. Growing [Cho04, PL07]. Growth [Amo18, GSK^{+}23, PBD^{+}11]. Guest [Ale13, AM15, Asr04, BC99, BS06a, Ber99, BC05a, BC05b, CE14, CN03, CZ07, CLZ13, Cho03, CF13, CS01b, DS00, Dub07e, Dun09, Eth01, FC09, FF03, For16b, Got06, GS13a, HP14a, HP14b, HS03, HG02, HP04, Kar02, Kax01, KS02, Kup03, MF16, MR06, Mem15, MS99, NC03, NL99, PT14, PV00, PS02, Pos04a, Pos04b, Pos07, Qua18, RC01, Run00, Run05, SS02, Sul09a, TA05, TX07, Thi05, TB99, TP04, TM00, VP04, VN99, Wai16, WR00, Win06, ZW19]. GUI [OASFLAB09]. Guide [KL15, Lew10, Pek04, Tsa14]. Guided [BT10b, DiP18b, DTA21, JT01, VWL^{+}11, ZFS12]. Guides [Sch01]. Guiding [Bur18]. GV [LGW^{+}17]. GW150914 [BVT^{+}21]. Gyrokinetic [ECK^{+}15].

H5N1 [Rao16]. Hackathons [CJL^{+}18]. Hadron [Mor15]. Hail [Bei12b]. Hair [YZC^{+}13]. Halos [Jon15]. Hand [Ano15i, YHWY05]. Hands [CBF^{+}22]. Hands-On [CBF^{+}22]. Hans [Ano21-46]. Happens [HGA23]. Haptics [EPHY18]. Hard [Sul03b]. Harder [MB17]. Hardest [CS01b]. Hardware [AAAH^{+}16, Day08c, EP10, For01, Hoe10, Hsu06, Naj08, NGGS22, SWDR22]. Hardware-Accelerated [NGGS22, SWDR22]. Hardware-Independent [EP10]. Harlan [Ano14y, Ano18p, Ano20-36, Ano21-47, Ano21-48]. Harmonic [DMR^{+}09]. Hash [MPP14]. Hashing [OOB17]. Haskell [CL12]. Hasselmann [Hei22]. Hasslacher [Sch23c]. Having [Ben04]. Hazard [KMM^{+}11]. Hazards [ASLK22, MT00]. HC [Bak10]. HC-1 [Bak10]. Headroom [Cho05a]. Healing [CJ16]. Health [Ano19f, Day13e, Jef22]. Healthcare [TGU21]. Hearing [Gio02]. Heart [ZB04]. Heart-Rate [ZB04]. Heartbeat [HMA00]. Hearted [Bar12]. Heat [Sch18]. Heider [Kul07]. Heights [Ano16-47, Ano16-48]. Helios [HWPS16]. Hell [Day13b]. Help [Ano15-37, FLV^{+}09]. Here [MM13]. Hesitant [LLQ18, ZLW^{+}19]. Heterogeneity [BBD^{+}13, DBB^{+}21]. Heterogeneous [Bak10, BKRT21, CR15, MSB^{+}14, Sch15, SGS10, TFF05, VGD^{+}11, YEC^{+}19, YB12, GGD^{+}05]. Heuristics [Boe00]. Hewlett [Got02a]. Hidden [Bal17]. Hierarchical [BZL^{+}07, DAEJ18, Muz19, QSEQJFH20, YEC^{+}19, YB12, Poi10]. High [AE22, AAGH17a, AAGH17b, Amo18, AMS14, Bai05, Bak10, BKRT21, BUS21, Bar21c, BCC^{+}09, BS21, BBW^{+}20, BCH^{+}22, CDL^{+}23, CN03, DD05, DCWH07, DWC^{+}11, DLW^{+}19, DGJ^{+}08, DSSS05, Dub22, EDJ^{+}10, EHG01, FHM99, GPMSC20, Got02b, GYJL20, HP20, HC99, HB08, HKW03, HG02, HP04, HEH^{+}10, JPMG08, JPK01, KWBB22, KTG08, KWB^{+}10, KT11, KBLD08, Lan02, LM08, Lat16, LQZL19, LAY04, Liu11, MS22, Men18, Mes15, MMG08, MWC^{+}16, MB99, MBB^{+}09, NC03, PGF^{+}15, Pap16, Par22, PSA14, PMP21, PMW20, QL19, Rag06, Rao16, RR21, SW10, SBZ^{+}08, SDA^{+}14, SBB^{+}15, SKC02, Sch18, STG11, Shi99, Shn06, SL99, SOH13, SR12, SR13, SJDV09, SS09, Str10, TAM^{+}14, TBVP^{+}21, VMK20, VM15, Vil08, WBS^{+}22, YBBP15, YCD^{+}21, YYG^{+}19, ZMM03, ZQY^{+}11, dSRT16]. High-Density [ZMM03]. High-Dimensional [BBW^{+}20, CN03, HKW03, MB99, NC03, Shn06, SR13, dSRT16]. High-Dimensionality [Vil08]. High-End [LAY04, Liu11, Mes15]. High-Energy [AMS14, DGJ^{+}08, PMW20]. High-Fidelity [HEH^{+}10]. High-Impact [STG11]. High-Level [MMG08, Rag06, VMK20]. High-Performance [AE22, Amo18, AMS14, Bak10, BKRT21, BCC^{+}09, BS21, BCH^{+}22, CDL^{+}23, DD05, DCWH07, DSSS05, EDJ^{+}10, EHG01, FHM99, HP20, HC99, HB08, HG02, HP04, JPMG08, JPK01, KWBB22, KTG08, KWB^{+}10, KT11, KBLD08, LM08, Lat16, MS22, Men18, PGF^{+}15, Pap16, Par22, PSA14, PMP21, RR21, SW10, SBZ^{+}08, SDA^{+}14, SBB^{+}15, SKC02, Sch18, SOH13, SJDV09, SS09, Str10, VM15, WBS^{+}22, YBBP15, SL99]. High-Precision [Bai05]. High-Productivity [BUS21, MBB^{+}09].

High-Quality [Dub22]. High-Resolution [DWC^{+}11, Lan02, TAM^{+}14, YCD^{+}21, ZQY^{+}11]. High-Risk [Rao16]. High-Speed [GYJL20, LQZL19, SR12, YYG^{+}19]. High-Temperature [DLW^{+}19]. High-Throughput [MWC^{+}16]. Higher [Pie04, SÜP^{+}11, SSG16]. Higher-Order [Pie04, SÜP^{+}11, SSG16]. Highlights [EH22]. Highly [HF04]. Hike [Lau05, Läu06, LTG07]. Hilbert [NA07]. Hip [XHL^{+}13]. HiQ [Shi99]. Hire [Day14f]. History [Ano21i, Ano21j, Ano21k, Ano22i, Ano22j, Ano23c, Mar17, Wil08]. Hit [Tou03]. Hitting [GL20]. Hodge [JG13]. Hole [BkCP22, BVT^{+}21, PRCL^{+}22]. Holograms [HRWS06]. Holographic [BDF^{+}20, Toh08]. Holography [Sic09]. HoloPy [BDF^{+}20]. Home [Kra03, Kra15, Thi06]. Homework [Ben00, O’L06e, O’L07b]. Hood [O’L06b]. Hopes [Got02b]. Hopkins [KBLE15]. Hopper [EH22]. Horizon [Cho05b, PRCL^{+}22]. Horizons [BMC99]. Hose [Tha08b]. HOST [Ano20-37]. Hosting [Thi07]. Hot [Ano14t, Ano17-40]. House [Ano14d, Ano14g, Ano14h, Ano14i, Ano14l, Ano14m, Ano14t, Ano14v, Ano14z, Ano14-29, Ano14-27, Ano14-28, Ano14-31, Ano14-40, Ano14-41, Ano14-42, Ano14-46, Ano14-47, Ano14-48, Ano14-49, Ano15c, Ano15e, Ano15f, Ano15i, Ano15g, Ano15k, Ano15l, Ano15u, Ano15v, Ano15w, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-35, Ano15-36, Ano15-43, Ano15-40, Ano15-41, Ano15-42, Ano15-39, Ano15-45, Ano16c, Ano16d, Ano16e, Ano16l, Ano16t, Ano16q, Ano16u, Ano16r, Ano16z, Ano16v, Ano16n, Ano16s, Ano16-27, Ano16w, Ano16o, Ano16x, Ano16p, Ano16-28, Ano16-29, Ano16-30, Ano16-38, Ano16-39, Ano16-40, Ano16-47, Ano16-48, Ano18-35, Ano18-37, Ano18-36, Ano18-40, Ano18-44, Day11a, Ano15s, Ano15t, Ano15-49]. How-to [Lew10]. HPC [CHC17, Ano18q, Bak21, BBWW22, Bor02, GK22, GDDR16, Har18, Hem10, Hoe10, HMB^{+}14, KVP^{+}16, KVP^{+}17, KFMG20, LGG^{+}22, NC18, SRKS22, SACR21, SSK13, SP18]. HPCMP [HLS^{+}16, HWPS16, KPA^{+}16, LGW^{+}17, WQT^{+}16]. HPJava [CF03]. HS [WMB20]. Hubble [Chr15]. HUBzero [MK10, GSK^{+}23]. Hulls [Muc09]. Human [Ano15-44, DL00, Gor08b, LCY08, NWP19, PBSS14, TGP^{+}06, YCL05, YCZ07, ZB04]. Humanitarian [MSB^{+}14]. Humanities [KHG^{+}23, Tan21]. Humphrey [Ano22-75]. Hundred [Sul05b]. Hurdles [Got02b]. Hurricane [DWC^{+}11, WZZ11, Zha11, ZQY^{+}11]. HWRFx [ZQY^{+}11]. Hybrid [CL14, CXC^{+}20, FGR^{+}07, FGRS17, Gaa03, Gan02, Gia22, LVLA14, LHGX18, XZL^{+}19]. Hydraulic [FGR^{+}07]. Hydrodynamic [WQT^{+}16]. Hydrodynamics [BRS22, KSM^{+}17, MSS09, Owe01, WHW18]. Hydrogen [SSK02]. Hyperbolic [LeV09]. Hyperspectral [DM12]. Hypotheses [GP15, Rus01b]. HYSPLIT [War18]. Hysteresis [KPD^{+}99, LCC^{+}19].

I-V [Azo06]. Ianus [BMP^{+}06]. Ice

[GGJD23, MRHP23, TAHP23, THGS07]. Ice-Free [THGS07]. ICTVdB [BO03]. Ideal [Smi00a]. Identification [CPdlF^{+}12]. Identifiers [HBG^{+}20]. Identifying [BH02, EUD15]. IDL [Gal11]. IEEE [Ano16b, Ano17-31, Ano18b, Ano22w, Ano22a, Ano22b, Ano13g, Ano13i, Ano13h, Ano13j, Ano14v, Ano14u, Ano14z, Ano14-29, Ano14w, Ano14-27, Ano14y, Ano14x, Ano14-28, Ano14-31, Ano14-30, Ano15w, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-35, Ano15-36, Ano16l, Ano16t, Ano16q, Ano16u, Ano16r, Ano16z, Ano16v, Ano16n, Ano16s, Ano16-27, Ano16w, Ano16o, Ano16x, Ano16p, Ano16m, Ano16y, Ano16-28, Ano16-29, Ano16-30, Ano17o, Ano17j, Ano17k, Ano17l, Ano17m, Ano17p, Ano17n, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano18s, Ano18r, Ano18y, Ano18-27, Ano18z, Ano18-28, Ano19t, Ano19u, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19v, Ano19s, Ano19w, Ano19x, Ano19y, Ano19z, Ano20-46, Ano20-38, Ano20-39, Ano20-47, Ano20-51, Ano20-40, Ano20-48]. IEEE [Ano20-52, Ano20-41, Ano20-49, Ano20-44, Ano20-42, Ano20-43, Ano20-45, Ano20-53, Ano20-50, Ano20-54, Ano20-56, Ano20-55, Ano21-49, Ano21-64, Ano21-53, Ano21-59, Ano21-54, Ano21-65, Ano21-60, Ano21-55, Ano21-56, Ano21-61, Ano21-50, Ano21-57, Ano21-62, Ano21-51, Ano21-63, Ano21-66, Ano21-52, Ano21-58, Ano21-67, Ano22x, Ano22y, Ano22z, Ano22-39, Ano22-44, Ano22-48, Ano22-55, Ano22-41, Ano22-45, Ano22-40, Ano22-56, Ano22-46, Ano22-52, Ano22-57, Ano22-49, Ano22-53, Ano22-59, Ano22-58, Ano22-47, Ano22-50, Ano22-42, Ano22-54, Ano22-51, Ano22-43, Ano22-60, Ano23-31, Ano23u, Ano23w, Ano23-37, Ano23-29, Ano23v, Ano23-34, Ano23-38, Ano23z, Ano23-30, Ano23-32, Ano23-35, Ano23-39, Ano23x, Ano23-27, Ano23-33, Ano23-36, Ano23y, Ano23-28, Ano23-40, Ano23-41, Ano23-42, Ano19c, Ano20q, Ano20r, Ano20u, Ano20v, Ano20-33]. IEEE [Ano20-34, Ano20-35, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20m, Ano20k, Ano20l, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-42, Ano21-43, Ano21-44, Ano21-40, Ano21-41, Ano21-45, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano22v, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano23t, Ano23c, Ano23f, Ano23d, Ano23e, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k]. IEEE-CS [Ano16b, Ano18b]. Ignition [Ged16b, LBS14]. II [BC05b, CENT22b, DdS23b, DR05a, Läu06, LT09, Mal07b, Pos04b, RTSS14b, Rus01b, SA08b, WHG21]. III [LTG07, Rus02, RD05a]. IK [WLL^{+}14]. IK-SVD [WLL^{+}14]. Illuminating [DRR^{+}04, Ged16b]. Illumination [BB07, Dac16]. Illustrator [Tof08, Tof09a, Tof09b]. Image [CO15, CS11, Eng15, FSED10, FKB^{+}13, KSSF11, LFC01, LPY18, LPCY19, LHGX18, MP09, NO03, PL07, PRCL^{+}22, Pey11c, TDT^{+}22, Van12, VWL^{+}11, ZFS12]. Image-Based [LFC01]. Image-Guided [VWL^{+}11, ZFS12]. Imagenation [Wil01]. Images [BBM^{+}21, Day15a, Gig00, GMPR11, JG12, MFD^{+}09, YWC02]. Imaging [JT01, MJAK09, MB07, QPCJ07, Rei21b, UGV11, Rei21a]. Imbalanced [RRN20]. Immune [KS00]. Immunology [Cyb99b]. Impact [CJL^{+}18, CR15, Day08a, ERF21, ECK^{+}15, GWMG04, Her22, HPC20, Ree16, Sch15, STG11, Van12]. Imperfect [Day16d]. Impingement [MPR18]. Implement [SDA20]. Implementation [MVUSK14, SCW^{+}17, TLD02]. Implemented [Naj08]. Implementing [GEH^{+}99, GDDR16, Joh06, ZFS12]. Implicit [RSC^{+}14]. Importance [BS99b, Chu21, Fox02c, PCY14, Thy20]. Importance-Driven [PCY14]. Important [Day07c]. Impressive [Nai15]. Improve [DL23, Eis17, STG11, WJ04]. Improvement [MBH14]. Improves [Jon19]. Improving [BMCC21, CJTH^{+}13, FAFX20, Gan02, GDH^{+}23, KL07, LPCY19, MM16, PSS20, SMM^{+}11, SNCT13, YLZ17, ZQY^{+}11]. In-Silico [KFS18]. In-Situ [BAD^{+}21]. In-Socket [SG10]. Inaugural [WDC18]. Incentivization [MMR22]. Included [Dub07e]. Inclusion [Akl18, Ano23m, Ano23n, HGA23, Leu17, Leu20, SL18]. Inclusive [Deb18]. Inclusivity [Mil21]. Income [Slo16]. Incompressible [DJS13, DGK16]. Incorporating [JLNR19]. Increase [MMTD^{+}17]. Increasing [DBB^{+}21]. Incremental [Roo06, WLL^{+}14]. Independent [EP10]. Index [Ano99, Ano00a, Ano01, Ano02a, Ano03, Ano04a, Ano05a, Ano06, Ano08, Ano09a, Ano11c]. Indifferent [Sul04b]. Individual [Biz16]. Individuals [Oli13]. Indoor [KLMS99]. Induced [FCT^{+}10, KNKP14]. Industrial [Das00, Mil10, RHC^{+}23]. Industry [Ano14t, Bea00, Bet99, Fei05, Gyu99, Haa99, TNR^{+}23, WLCD01, YAA^{+}00]. Inevitable [DS23]. Infected [Sch21]. Infection [O’L04c, O’L04d, Sch21]. Inflection [Gar17]. Influence [AM05, DKWL17, HH22]. Influencers [CHM^{+}20a, OHK23, CHM^{+}20b]. Influences [JH16, Sca16]. Inform [DBCN03]. Informatics [DBCN03, GKG^{+}15, Liu15, MSM13]. Information [Ano13i, Ano14-29, Ano18c, Ano18d, Ano18e, Ano18f, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano22f, Ano22-27, Ano22-52, Ano22-53, Ano22-54, Ano23-34, Ano23-35, Ano23-36, BBN03, BCL03, Cha08, Che10, FL99, Fox02a, Fox03c, Ged16a, HGA23, Kal99, KNV03, Lew00b, LLQ18, MO03, MSL02, Pok04, Qua18, Qua19, TDT^{+}22, TW17, ZS07, ZYKG04, SMM^{+}11]. Information-Theoretic [KNV03, TDT^{+}22, ZS07]. Infrared [ML02, UGV11]. Infrared-Scene [ML02]. Infrastructure [ABM^{+}22, Ale18, Cyb99b, Got02b, MTG^{+}12, Got15, Got16]. infrastructures [SPS15]. ing [NG20]. Initial [LTD11]. Initialization [WZZ11]. Initiative [SDA20, Ano18-41, ES18]. Inner [YZC^{+}13]. Innovating [Goo17]. Innovation [FL21, McM09, Men18, Mil17, OHK23, PG17, SSK13, WMB20]. Innovative [For99]. Inpainting [FSED10]. Inria [ACG^{+}20]. Insights [Ano14t, Dub21, Mal07a, Sch23a]. Inspection [DLLZ19, DLLZ20]. Inspired [For00, Gor08b]. Instabilities [BW01, ZB04]. Institute [Ano18d, Ano18e, Ano18f, CGK^{+}18, ABC^{+}14, CHH^{+}13, MMG^{+}05, Par12, WDC18]. Institutes [WDC18]. Institutional [Biz16]. Instruction [CL01, CW05b, CW05c]. Instructional [Ree16]. Instrument [BCE^{+}22, CBF^{+}22, DC04]. Instruments [Tou03]. Integer [AB03, Bai00]. Integral [Ama00, Ron14]. Integrals [Ron14]. Integrated [CAP^{+}10, DDV^{+}08, GF04, HKB12, HEH^{+}10, JMEL08, NCB^{+}05, PGC21, RF00, She07, WQT^{+}16, WWJH20]. Integrating [CHB19, CF99b, Fox03c, Hu07, Lof22, MCE^{+}03, YWC02]. Integration [BKB20, GK22, Nan11, Nob00a, O’L04e, Tur15, ZZYNH06]. Integrative [CL01]. Integrity [Bak21]. Intel [BHC^{+}15, DADY15, HKB12, Rob13]. Intelect [Ano14-32, Ano15-28]. Intellectual [Cyb99c, SLK^{+}20]. Intelligence [BB20, Gor06b, Har23, SGRK^{+}18, TKM^{+}18, Ano22w, Ano22a]. Intelligent [BO04, Har23, QRP22, Qua18, Qua19]. Intensive [AHL^{+}11, Bry11, HPMJ12, NCB^{+}05, SYM^{+}21, Sza11]. Interacting [CHM^{+}20b, CHM^{+}20a, DGR^{+}05, HL00]. Interaction [Kus06b, LCY^{+}04, MRNT17, MPR18]. Interactions [EUD15, LBS14, SNTL13]. Interactive [BBW^{+}20, DH12, HLYQ19, Lan02, Mar02, PG07, PFS21, PTH13, RR21, SDS00, Sic09, Sil02, SJF^{+}23, TC21, VGM^{+}09, Vor01b, WAS^{+}12]. Interactively [LM07a]. Interbeat [TGP^{+}06]. Intercalation [PAF08]. Interconnect [Mei03]. Interdisciplinary [GN08, Tan21]. Interest [JH16, O’L04b, SKA^{+}02]. Interface [BJ02, GRE99, OW01, VGM^{+}09]. Interfaces [Ben04, PV00, Shi01b, Smi99d, Smi99a, Smi99b, Smi99c, Smi99e, Smi99f, Smi00a, Smi00c, Smi00b, Smi00d, Smi01b, Smi01a]. Interferometric [Zeb00]. Interinstitutional [GCV08]. Interior [DRR^{+}04]. Intermediate [SSK02]. Internal [CFCD04]. International [Ano15b, Car09a, Car09b, CHC17, Run05]. Internet [Ano19x, Com99, CNO99, Day10b, Day11d, Day15a, Day16a, FGRS17, GHKZ17, Kil99, MHDM99, NLV99, Rob06, Shi07, Shi01a, Smi99b]. Internet-of-Things [FGRS17]. Interns [Lof22]. Interplay [BHKW03]. Interpolation [WLCJ12]. Interpretive [CNC10]. Intersection [Sha99]. Intersectional [Wri16]. Interval [LLQ18]. Intervals [TGP^{+}06]. Interview [Mar99a, Sha99, WM00]. intracardiac [SOV^{+}13]. Intracranial [WNZ^{+}17]. Introducing [KNS18, Thi09b]. Introduction [Asr04, BLMR21, BC99, BS06a, Ber99, Bli02, BC05a, BC05b, CN03, CZ07, Cho03, CS01b, DS00, Dub07e, Dun09, Eth01, Fal06, FC09, FF03, Got06, HS03, HG02, HP04, Kar02, Kax01, KS02, Kup03, LFK^{+}19, MR06, MS99, NC03, NL99, PV00, PS02, Pos04a, Pos04b, Pos07, RC01, Run00, Run05, SS02, ST99, Sul09a, TA05, TX07, Tes15, Thi05, TB99, TP04, TM00, VP04, VN99, War18, WR00, Win06, ZW19, Ale13, AM15, CE14, CLZ13, CF13, For16b, HP14a, MF16, Mem15, PT14, Qua18, Wai16]. Introductory [Ass00, Aya07, MMP^{+}22]. Intuitive [Wan23]. Invasion [Ebr10]. Inverse [Chu21, XBK10]. Inversion [CL14]. Inverter [LCC^{+}19, WCC^{+}19]. Investigations [CHB19]. Involve [DM04]. Iodine [MM04]. Ion [TLR10]. Ionic [BHKW03]. Ions [Seg99]. IPython [PG07]. IRI [Ano20-54]. Iron [BKK15]. Irony [Day08b]. Irregularities [KS06]. iSERVO [MMG^{+}05, Run05, YLCZ05]. Isn’t [RMX12]. Isosurface [PCY14]. Isosurfaces [BB07]. Isotopes [Har18]. Issue [Ano15e, Ano15f, CENT22a, CENT22b, FT08, JD03, O’L06e, O’L07b]. Issues [HBB08, TLD02]. Istanbul [DSPY05]. Iteration [vdV00]. Iterative [O’L06c]. IV [Rus03, RD05b]. Ivory [Far99].

J [Sny13, Rei13]. J2EE

[Lau05, Läu06, LTG07]. Jack

[HH22, Par22, Vud22]. Jackets [Day10a]. Janus [BCC^{+}09, Smi00a]. Japan [Sor19]. Japanese [For99]. Java [Esq11, Fox03d, Has12, HRAB05, PTML09, PTML11, Thi02, Vil08, XYC05]. Java3D [Vor01b]. JavaScript [DiP18a, DiP18b, TAF^{+}18, Zak18, dJM18]. JHelioviewer [MFD^{+}09]. jLab [PTML09]. Job [Ano14l, Ano14j, Ano14k, Ano14m, Ano15j, Ano15k, Ano15l, Ano17d, Ano19-31, Ano19-32]. Jobs [Ano13k, Ano18-29, Ano19v, Ano20-50, Ano21-64, Ano21-65, Ano21-66, Ano21-67, Ano22-55, Ano22-56, Ano22-57, Ano22-58, SOH13]. Johnny [DEK03]. Johns [KBLE15]. Joint [JXY^{+}19]. JOSS [KNS18]. Journal [Ano20-33, Ano20-34, Ano20-35, Ano21-42, Ano21-43, Ano21-44, Ano21-40, Ano21-41, Ano21-45, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano23t, Gaa03, TS10, KNS18]. Journal-Database [Gaa03]. Journals [Day08a]. Journey [WP22]. Joy [Sul00c]. JPEG [MFD^{+}09]. Jr. [Sul03c]. Julian [Slo16]. Jump [Che99]. Jungle [Ben09]. Jupyter [Bar21b, BTK^{+}21, FKD21, GP21, JON^{+}21, Moo21, PFS21, Tan21, TC21, WBB^{+}20]. Jupyter-Enabled [JON^{+}21]. Just [Bei11b, Cho05b, Cho06c, Day15d, ERS^{+}03, Gor06d].

K12 [SDA20]. Kaleao [Goo17]. Karlsson [Ano21-46]. Keeneland [VGD^{+}11]. Keep [Ano15h, Ano19-28, Ano19-27, Ano20-57]. Keeping [Cus14, Lew01b]. Kernel [Ama00, BKB20]. Kernels [DADY15]. Key [LL18, Sch99]. Kind [BC02]. Kinetics [MDK16]. Kink [LH06]. Kiosks [Day18c]. Klaus [Hei22]. KMAX [Goo17]. Knee [ZCXM99]. Knights [DADY15]. Know [DC04, Hin20b, MAFM21, O’L05g, SHPL12, Sul03a]. Knowing [Day13c]. Knowledge [GSB^{+}12, KB04, KS01, QSEQJFH20, Ano17-29, Ano17-30, Ano18-42]. Knowledge-Management [GSB^{+}12]. Kokkos [TBVP^{+}21]. Kriging [XZL^{+}19]. Krylov [Saa22, vdV00]. KSS [KSSF11].

Lab [CC99, DDV^{+}08, MVUSK14]. Lab-Data [CC99]. Label [GYL^{+}17]. Laboratories [AM15, MBB^{+}22, Sil02]. Laboratory [DZW^{+}05, KBLE15, KT08, Shi01b, MTBG^{+}22]. LabView [Shi01a]. Lake [VVNV18]. Land [WSY^{+}22]. Landau [Osk07]. Landscape [Cho06b]. Landscapes [MB99]. Lane [LL19]. Language [Cho12, Gor06b, HBG^{+}20, Mil21, SS06, Shi00c, Shi00b, Taj10]. Languages [BFS04, DS12, Hin18c, JWLG14, LSBC22, MMG08, PMM^{+}08, Wol21]. LAPW [DPP^{+}01]. Large [Ama00, BLO^{+}22, CS14, CS15, DJS13, DMXR^{+}14, Eis17, GL99, Gob05, HHF^{+}14, HMB^{+}14, JPE20, Jon15, KMSH10, LWF10, LCY^{+}04, Luo12, Ma16, MWE08, MFD^{+}09, PCY14, Sah03, SBH^{+}00, SKNV03, Ste99, TB11, TNV^{+}02, VKN99, Wes21, WNDV21, YWMM04, ZJW08]. Large-Eddy [DJS13, YWMM04]. Large-Scale [Ama00, CS14, CS15, DMXR^{+}14, Eis17, GL99, Gob05, HHF^{+}14, HMB^{+}14, JPE20, Jon15, KMSH10, LWF10, LCY^{+}04, Ma16, MWE08, Sah03, Ste99, TB11, VKN99, Wes21]. Larger [GAB^{+}22]. Laser [LBS14, MM04, TDB09]. Laser-Plasma [LBS14]. Last [Day12e, Day13d, O’L06e, O’L07b]. Later [BS99a]. Latin [MHC^{+}18]. Latina [DVS22]. Lattice [Ale15, Bog03, Mor15, Pek04, Sch20, Sch23c, SFSK01, WNZ^{+}17]. Lattices [Rei02]. Law [Fla17, FF03, Gar17, Mei03, TW17, TFS17, Wes03]. Lazy [LT09]. Leaders [Ano14t, OHK23]. Leadership [Ano17-31, Got14a, HP14a, HP14b, LUMM14, Leu20, WG15]. Leading [Ano16-47, Ano16-48]. Leaf [DB07]. Learn [Guo23, MPRU23, Ste14, Wil08]. Learned [DADY15, HLS^{+}16, KKO^{+}20, KB09, Tha08a, Wit21]. Learning [Ale13, Bar20b, Che19, Cho05f, Cho08e, Cyb01, Gia22, GVB15, GPMSC20, Les16, LWG19, MSM13, Muz19, NG20, NWP19, PTH13, Rob04, SR13, Sul17, Thi12a, VB22, WLL^{+}14, WCP17, YMHQ19, ZWYL20, Sny13]. Learns [Gor06b]. Least [Dub08a, Rus01a, Rus02]. Legacy [BJ02, Lov04, O’L06a]. Legal [Sto09]. Legate [BLP^{+}21]. Legendre [Nob00a]. Lego [Day10a]. Length [Gyu99]. Lens [SL18]. Lesbian [Wri16]. Lesions [Gig00]. Less [Ano15s, Sul03d, Ano15t, Ano15u, Ano15v]. Lessons [DPBS16, Day09c, DADY15, Fom15, HLS^{+}16, KKO^{+}20, KB09, Tha08a]. Let [Sul05b]. Lethality [SKC02]. Letters [Ano04b, Ano07, Ano18y, Ano18-30, Ano19-29, Ano19-30, Kil99, MBS^{+}00, MHDM99, NLV99]. Level [DAEJ18, MMG08, Rag06, Rob06, RR21, SDA20, VMK20]. Levels [HC17]. Leverage [MPRU23]. Leveraging [AAAH^{+}16, KFS18, MBH14]. Levitation [JLYL19]. Lexicon [TAF^{+}18]. LHCb [TJCC20]. Liberal [SL18]. libflame [VCvdG^{+}09]. Libraries [BWC01, Gar06, Mas06]. Librarization [BKS15]. Library [BS21, Cho08c, FKB^{+}13, Moo21, SL99, TAF^{+}18, Thi10, VCvdG^{+}09, dJM18, Ano14-55, Ano14-56, Ano15-48, Ano15-49]. Licensing [Sto09]. Licensure [Tho12]. Life [Esq11, Gor05a, HE05, KB09, LGJ^{+}19, Lov04, Ome06, Shi00d, Smi99c]. Lifecycle [CHB19]. Light [Fei05, JMFJ01, OR12, SYP08, TM15]. Lighting [NG20]. Like [BST^{+}13, CJ16, MPRU23, OS03, Sul06c, Wat21]. Limb [XHL^{+}13, YMHQ19]. Limits [Die12, Fra02, PS02]. LIMS [Shi03]. Line [BW14]. Linear [GS03, HH22, Nas00, O’L05c, O’L06c, Rus01a, ZZPC06]. Lines [CT00, WT12]. Lingua [WHH22]. Linkage [Dal99]. Linked [WCC^{+}02]. Linking [Ale18]. LinkSCEEM [Ale18]. Linux [Thi04]. Lip [GMPR11]. Lip-Contour [GMPR11]. Liquid [BW01]. Listening [CW06]. Literacy [Tur14b]. Literate [LCG^{+}20]. Literature [KS20, Sch01]. Litter [JJ15]. Little [O’L06f, Shi00b, Day09c]. Live [Gar17]. Living [Rad21]. LLNL [NdS17]. Load [CS01a]. Load-Balanced [CS01a]. Loading [BWC01, STG08]. Local [Thi10, TLD02, YLR02]. Localization [SS11]. Loci [ZL09]. Locomotive [WCC^{+}19]. Logical [Har23]. Logs [MSR15]. Lonely [Bau08, Sul08b]. Long [AISR^{+}21, Gar17, HS12, PSR^{+}00, RBC^{+}19, SBH^{+}00, Smi00c, TNR^{+}23]. Long-Range [PSR^{+}00]. Long-Term [AISR^{+}21, HS12, RBC^{+}19, TNR^{+}23]. Look [Bog03, Kir03, Lew02b, O’L06b, Ste99]. Looking [Ano19-31, Ano19-32, Her21, Sch17b]. Lookup [KEF07]. Loop [LM07a]. Loss [KNKP14]. Loss-of-Consciousness [KNKP14]. Lossy [SSP06]. Lot [O’L06f, Sch14]. Lots [Cho05a]. Lottery [Cyb99c]. Love [Ano18-40, Wit21, Day12f]. Low [ABNZ09, CJTH^{+}13, CS18, JPK01, RR21, Slo16]. Low-Cost [CJTH^{+}13, JPK01]. Low-Dimensional [CS18]. Low-Income [Slo16]. Low-Level [RR21]. Lower [YMHQ19]. LQCD [BST^{+}06].

M [Wep15]. M87 [PRCL^{+}22]. Mach [ABNZ09]. Machine [Ale13, Cyb01, Gia22, Gor06c, MSM13, NWP19, PTH13, SR13, VB22]. Machine-Based [NWP19]. Machines [Che19, DPP^{+}01, HP15, Hin17b, KM12, LWG19, WC17]. Macromolecular [TGP13]. Macroscopic [HT99]. Madagascar [Fom15]. Magazine [SCBT18, Thi15b]. Magazines [Ano13d, Ano14e, Ano14f, Ano14g, Ano14h]. Magic [Har18, Hin20a]. Maglev [DLW^{+}19]. Magnetic [MB07, SV14, TL08b]. Magnetism [Ass00]. Magnetohydrodynamic [MRNT17]. Magnetohydrodynamics [GPZ^{+}04]. Magnetron [FMB^{+}07]. Mainstream [DD05]. Maintaining [Dub05b]. Major [Ano16n, Ano16o, Ano16p, Ano17p, Ano20n, Ano21y, Ano21z]. Majors [Slo16]. Mak [NG20]. Makes [Bei10c, Day08c]. Making [Bar11, CE18, FS12, Kra03, LLQ18, Nob07, PF04, Rag06, RSZ^{+}20, RSZ^{+}21, SKC00, Sul02a, Wil06, YaL10]. Male [Deb18]. Manage [KB04]. Managed [NCB^{+}05]. Management [Ano18z, Ano19y, CDL^{+}23, CF13, DVR^{+}19, DBH^{+}02, FWGB07, GB20, GSB^{+}12, HL01, KPA^{+}16, MWE08, MSL02, NKV99, NSLD99, PGF^{+}15, SOH13, STM99, SBZB13, SF11, TSFG08, TL04b]. Managing [BDCT05, DD23, FB04, GP15, Hin12b, KB09, O’L06d, RMX12]. Manifold [ZWYL20]. Mantle [MGZ00]. Manufacturing [FM13, SGW02]. Many [BFL^{+}22, BOS07, CWOL11, GHKR11, HKB12, KSP12]. Many-Core [BFL^{+}22, CWOL11, GHKR11, HKB12]. Map [ZDW^{+}07]. Maple [CW05b, CW05c, CW05d, CW05e, Dub00, Ron14]. Mapping [JJ15]. MapReduce [Coh09]. Maps [KSSF11]. March [Day18c]. Marching [DSC^{+}09]. Marine [Cor07, JJ15]. Marker [ACQ^{+}20]. Marketing [Ano00b, Smi99b]. Markov [DJ02, Ran06]. Mars [MRNT17, VPL18]. MARTE [RGD13]. Mass [BZL^{+}07]. Masses [DMXR^{+}14]. Massive [Ber99, Gor06d, O’L06d, Ott16]. Massively [BGHR06, BRS22, CWOL11, KKO^{+}20]. Master [Peo20]. Mastering [Hin17a]. Masthead [Ano13l, Ano13m, Ano14-33, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano15-34, Ano15-29, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano16-31, Ano16-32, Ano16-33, Ano16-34, Ano16-35, Ano16-36, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano18-31, Ano18-32, Ano18-33, Ano18-34, Ano19-37, Ano19-33, Ano19-34, Ano19-35, Ano19-36, Ano20-63, Ano20-58, Ano20-59, Ano20-60, Ano20-61, Ano20-62, Ano21-68, Ano21-69, Ano21-70, Ano21-71, Ano21-72, Ano21-73, Ano22-61, Ano22-62, Ano22-63, Ano22-64, Ano22-65, Ano23-43, Ano23-44, Ano23-45, Ano23-46]. Matching [XXK^{+}02]. Material [ASLK22, BCB07, BMS99, CBB06, Eis17, LFC01, NKV99, SGW02, SÖS^{+}00]. Materials [AGM^{+}00, BNNM04, EKCS12, GL99, KNG10, Kax01, KES^{+}22, KS01, LHN^{+}12, Lew10, MJAK09, Osk07, PZJS10, Saa09, SDA^{+}14, SKNV03, STH22, ZS07]. Math [Bur99, Can99, Mar99a, Rei13, TM14]. Math-Style [Rei13]. Math.Js [dJM18]. Mathcad [Bur99]. Mathematica [CW05b, CW05c, CW05d, CW05e, Com99, MAC08, Shi02a]. Mathematical [BB01, DM04, HAB17, HHR02, McK00, OW01, Shi99]. Mathematicians [Sch23a]. Mathematics [Coo14, Gra08b, RHC^{+}23, dJM18]. MathML [LVWK02]. MathScript [Com99]. Matlab [CW05b, CW05c, CW05d, CW05e, Coo14, Edd09, Has12, MAC08, Ono01]. Matplotlib [Hun07]. Matrices [O’L14]. Matrix [Kir03, LGG^{+}22, O’L06e, O’L06g, SL99, Ste00, VCvdG^{+}09]. Matter [DCWH07, Deb18, O’L05d]. Matters [Ano14-39, Ano14-40, Ano14-41, Ano14-42, PKST08a, PKST08b, PKST08c]. Maximizing [CF13]. Maximum [HSJ^{+}19, PA12, Sch16]. MaxLike [SMM^{+}11]. May [Dub08a, Smi99a]. Mayavi [RV11]. MCALab [FSED10]. MDBN [GYJL20]. MDE [RGD13]. Me [BS99a, Day13c, Dub06b, Dub07c]. Meandering [O’L06c]. Means [O’L06f, Pie04]. Measure [CRDO16].

Measurement [Tou01, VSG^{+}02]. Measurements [DMA^{+}21, O’L13]. Measures [YLZ^{+}19]. Measuring [COS^{+}15, DYY^{+}17, SBZ^{+}08]. Meatspace [Day12b]. Mechanical [JS99, Roh10]. Mechanics [Ara99, BC03, BCB07, Ben00, Can99, Com99, CFCD04, Gut01, JCPS14, LWSK07, Mal00, MSR^{+}16, Ono01, Reb99, Sch16, Sil00, Tho99c, XYC^{+}09]. Mechanism [Cho08g, XDK^{+}20]. Mechanisms [WBB^{+}20]. Mechatronics [Cra03]. Media [CHM^{+}20b, CHM^{+}20a, Day13f, KM99, Sah03]. Mediated [WLCD01]. Medical [CLZ13, Eng15, Gig00, Gor08c, Luo12, PL07, QPCJ07, TMC^{+}13]. Medicine [WR00]. Mediterranean [Ale18]. Meep [LFN^{+}11]. Meets [Bei09c, CT00, Fox02a, Haa99]. Membership [Ano13n, Ano13o, Ano14-39, Ano14-40, Ano14-41, Ano14-42, Ano17x, Ano17y, Ano17z, Ano17-27, Ano18-35, Ano18-37, Ano18-36, Ano18-46, Ano17-28, Ano17-29, Ano17-30, Ano18-42]. Membrane [FGP99, Wol16]. Membranes [TLR10]. Memory [AAAH^{+}16, DAEJ18, DPP^{+}01, O’L06b, PKST08b, VM15]. Mental [Jef22]. Mentoring [Bar11]. Mercer [Mar02]. Mercury [MW14]. Merger [BVT^{+}21, DMA^{+}21, Smi99e]. Merging [WC17]. Merit [EVL^{+}17]. Merwin [Ano14-43, Ano17-31, Ano18-38, Ano21-74]. Mesh [Bry99, DBB^{+}21, LJWC06, MCAA05, NSLD99]. Meshes [O’L06d]. Mesoscale [DGR^{+}05]. Message [BBG^{+}01, Fox02b, Vin12]. Messages [Bau08]. Metadata [Fox03a]. Metal [KLS01, WM00]. Metamorphic [KC19, LSN20, LSPN21]. Metaphysics [Cho07a]. Metaprogramming [MM18]. Method [ASLK22, Ama06, Bas14, BS06b, BGHR06, CL14, CLC03, DAEJ18, Fra07, GH00, IHL^{+}02, LGJ^{+}19, Nas00, RK05, Rei02, WNZ^{+}17, WCC^{+}19]. Methodology [APC^{+}19, TKM^{+}18]. Methods [ATG05, BW06, BS06a, Coo14, Fel00, GS03, GPC08, HMB^{+}14, JSNR11, JHJ01, KS20, MS07, O’L06c, O’L06d, OL06h, Ork09, Oug03, PP20, Saa22, SÖS^{+}00, STB03, TK06, TLD02, Wep08, XBK10, Yav06, YLR02, Seg99]. Metropolis [BS00c]. MFiX [BAD^{+}21]. MFiX-Exa [BAD^{+}21]. Michelson [Ste02]. Microanatomy [DZW^{+}05]. Microarray [RRN20]. Microbiology [Nai15]. Micromagnetic [Zhu16]. Micromap [WCC^{+}02]. Microprocessor [WJLY08]. Microscopy [BDF^{+}20, SRM^{+}07]. Microsoft [Smi99e]. Microstructural [BP99]. Microstructure [CBS14]. Microstructures [LFC01]. Microwave [BCJK99, BKK15, CPdlF^{+}12]. Middle [Les16, Sca16, Thi02]. Middleware [TBM^{+}19]. Might [PK18]. Migrating [TSKG03]. Millenium [ZAF^{+}01]. Millennium [Cyb00c]. Millisecond [Fox04b]. Mills [Ano14y, Ano18p, Ano20-36, Ano21-47, Ano21-48]. Mind [Day12d]. Mind-Reading [Day12d]. Mine [Ged16a]. Mines [Jon15]. Minimal [JCC^{+}10]. Minimization [BOS07]. Mining [FAFX20, GM02, Kar02, KNV03, KJ04, RBK02, TNV^{+}02]. Minor [GCV08]. Minority [HG00]. Mirrors [Day16d]. Misinformation [Day19a]. Mission [Cho05c]. Missions [EVL^{+}17, MSB^{+}14]. Mistakes [MPRU23]. Mitigation [KMM^{+}11]. Mixed [Hin18a, Tou01]. Mixed-Signal [Tou01]. Mixing [Ran06]. Mobile [ACQ^{+}20, Ano14-47, Ano14-48, BB20, DYY^{+}17, GHKZ17, SdPV21]. Mochi [CDL^{+}23]. Modality [KEF07]. Mode [SCW^{+}17, SD11]. Model [ASM^{+}14, ACF18, Bak21, CPdlF^{+}12, Cos22, DJ02, GGJD23, GB20, GVB15, GWA^{+}07, Gyu99, HLYQ19, HSJ^{+}19, JCC^{+}10, JLYL19, Lan06, ML02, NWP19, O’L07a, O’L07b, PHW^{+}21, QSEQJFH20, Rad21, Ree16, Sch21, Sch23c, War18, YWC02, YLR02, dOMdO^{+}04]. Model-Order [CPdlF^{+}12]. Modelica [SU22]. Modeling [ASM^{+}14, ABK^{+}02, AMS14, Ano15e, BMCC21, BPH^{+}13, BPMKC21, BZL^{+}07, Bri23, CAP^{+}10, Cas16, CF99b, CYW01, Cho07a, CS01a, CNO99, FSD02, FGR^{+}07, FGG^{+}22, FGRS17, GSB^{+}12, HAB17, Har04a, HDB^{+}04, HPKS04, Hus22, Ikk16, IK05, IHL^{+}02, IBPV03, JLP^{+}10, KM99, KES^{+}22, Kel10, KL10, KHC^{+}07, KB07, LWF10, Lum07, MRNT17, MM18, Muz19, Nai15, NW15, Orf21, Owe01, Par16, PHW^{+}21, Put16, RRH^{+}02, RC01, SU22, Sch21, Seg99, STG11, SNCT13, SV14, STH22, SS02, SKA^{+}02, STB03, Str10, TM15, TA05, TX07, TAHP23, Tie16, VSE01, Wai16, WCHRM21, WHM^{+}02, WCGB05, WPW11, WSY^{+}22, WPZ00, YM14, Yas17b, ZS07, ZQY^{+}11, ZCXM99, ZS23, dA03, Beh05]. Models [ATRA00, Ben04, CS18, DLLM04, GEH^{+}99, Gia22, GW15, HHR02, HGV^{+}08, Hin23, KKO^{+}20, Liu15, Luo12, MT00, McK00, MSD10, MRHP23, New00, O’L04c, O’L04d, Pek04, Pie04, PD02, PSR^{+}00, SETK05, She07, WWJH20, XHL^{+}13]. Modern [Alt10, Cho09a, Cho09b, Day15e, DS12, Hu07, Mol12, RMX12, Sch14, SBB^{+}15, WC17]. Modernization [NdSS17, NdS17]. Modify [KRR^{+}12]. Modifying [Tof09a]. Modular [SZM^{+}13]. Module [KB07, TGEA09]. Modules [Bar20b]. Molding [JMFJ01]. Molecular [BST^{+}13, BH02, CJTH^{+}13, Cho08g, Col18, EP10, HBD^{+}23, HE05, IBPV03, KES^{+}22, KFS18, LNI^{+}19, Mal07a, Mar99b, NKV99, NMCM22, PZJS10, Rei02, TDB09, TX08, Vor01b, WDC18]. Molecular-Dynamics [NKV99]. Molecular-Scale [Rei02]. Molecule [Gor06a, NLGNJ13]. Molecules [McC21]. Moler [Mar99a]. MolSSI [NMCM22]. Money [SW10]. Monitoring [KNKP14, VVNV18]. Monotonically [GF04]. Monte [Ama06, BS06a, BS06b, BOS07, DCC10, Day13e, Eng09, EKCS12, Fel00, GH00, LSDP^{+}04, Lui06, Nob02a, OASFLAB09, Ork09, PPE00, SWDR22, SG00, Sul17, VWP12]. MOOC [Kra15, PQQ20]. Moon [Lan02, MM14]. Moore [Fla17, FF03, Gar17, Mei03, TW17, TFS17, Wes03]. Morley [Ste02]. Morphogenesis [MHK^{+}06]. Morphologies [BMCC21]. mOSAIC [JCPS14]. Most [Orf21]. Motion [JRP^{+}17, VWL^{+}11, YZC^{+}13]. Motions [BHKW03]. Motivation [MM12]. Motors [Cho08g]. Mountains [JJZC10]. Mouse [ERS^{+}03, SRM^{+}07]. Mouth [GMPR11]. Mouth-Structure [GMPR11]. Movable [And11]. Movement [DL00]. Movie [Kra03]. Moving [LTD11, MB11, WJLY08]. MPI [CNC10, Ong02]. mpi4py [DF21]. MPPs [DSSS05]. Mr [Day09b]. Mr. [Smi00b]. MRI [HMA00, MWC^{+}16]. MRICloud [MWC^{+}16]. Mrs [Smi00b]. Multi [Ano16-47, Ano16-48, RRN20]. Multi-Core [Ano16-47, Ano16-48]. Multi-Objective [RRN20]. Multiagent [YYL^{+}18]. Multiagent-Based [YYL^{+}18]. Multiagents [Naj08]. Multicanonical [GH00]. Multicellular [CAS^{+}07]. Multicharacterization [ALH15]. Multicomponent [Dub21]. Multicomputers [SWPB00]. Multicore [Gor07c, KM12, WJLY08]. Multicriteria [LLQ18]. Multics [MKM^{+}14]. Multidimensional [HW15, NSP12, O’L04e, SKNV03, YKD^{+}03]. Multigrid [BW06, BGHR06, Fal06, MR06, O’L06d, OL06h, Yav06]. Multilevel [PLW17]. Multimedia [EGFL12, Shi00c]. Multimillion [BW14]. Multimillion-Line [BW14]. Multimodal [LPB13]. Multimodel [SPJ^{+}14]. Multipart [Ben00]. Multiperspective [LHGX18]. Multiphase [BAD^{+}21, FPRK16]. Multiphase-Flow [FPRK16]. Multiphysics [Dub21, Gra08a, GZC14, KVP^{+}16, KVP^{+}17, TA05]. Multiple [Bea00, DKWL17, FM19, Kus06a, LPY18, Smi03]. Multiple-Choice [Bea00]. Multiple-Precision [Smi03]. Multiple-Scales [Kus06a]. Multipole [BS00d, MBS^{+}00]. Multiprocessing [SMF^{+}23]. Multiscale [BPH^{+}13, FMKS08, GZC14, GNB^{+}09, Hym05, IHL^{+}02, LVLA14, MM18, MCAA05, NBK^{+}01, PAF08, Pey11b, SNCT13, SFSK01, SKC05, XKG05]. Multisensor [HSJ^{+}19]. Multisensory [Har04a, Lof03, Rob04]. Multitask [GVB15]. Multiterabyte [TSKG03]. Multithreaded [MAG21, SZM^{+}13]. Multitier [PSS20]. Multivariable [XZL^{+}19]. Multivariate [AMCH07, DH12, HH99, JMEL08, Liu15, SETK05, TDT^{+}22]. Multiyear [JH16, Wol16]. Muscle [YCZ07]. Muscles [Roh10]. Musings [Cho07a]. Mutation [HK09]. My [Bei12c, Ben09, Day06b, Day10c, Day10d, Day14c, MM14, TL08a, Toh08]. myComputer [Ano13p]. myCS [Ano17w, Ano18-39, Ano18-40]. MyDB [LT08].

N [Mol12]. Naked [Läu08]. Nallatech [SG10]. Nanobiological [FMKS08]. nanoHUB.org [KMB^{+}08]. Nanoparticles [KLS01]. Nanophotonics [BVB^{+}07]. Nanoscale [SKL10]. Nanoscience [RC01]. Nanoscopic [Kyr08]. Nanostructured [KNG10, PZJS10, VWP12]. Nanosystems [NBK^{+}01]. Nanotechnology [KMB^{+}08, RC01, SMC01]. Nanotubes [SMC01]. NASA [DM12, LAY04, MB11, MR13, MM14, Mem16, SM17, SNCT13, Sim13, SV14, YLR02]. Nascent [Tan21]. National [MTBG^{+}22, MBB^{+}22, Tow18, Ano18-41, ES18, LBS14, Tou03, WG15]. Native [AAB^{+}21]. Natural [ASLK22, Asr04, FWGB07, HBG^{+}20, KB04, MT00, TL04a, YHWY05]. Nature [Hin23, NTW07, Run03, Rus01a, Rus01b, Rus02, Rus03]. Naval [PAN^{+}16b]. Navigating [PSJ^{+}21]. Navy [MSR^{+}16].

NavyFOAM [KSM^{+}17]. Near [CW05e, Cho08c, JRP^{+}17, MM13, ZMM03, Zei17]. Near-Fault [JRP^{+}17]. Near-Field [ZMM03]. Nebulae [WAS^{+}12]. Necessary [Toh07]. Necessity [Hem10]. Need [Day14e, Lew00a, WBP^{+}19]. Needed [Ano15-37, Mes15]. Needs [HG02, Jef22, Thi13d]. NEEShub [HEB^{+}11]. Negative [Bot16]. Negotiating [SDA20]. Negotiations [BFS04]. Neither [DS23]. NEMO [KL10]. Nene [KPM10]. NERSC [BKK15, ECK^{+}15, YBBP15]. Nerve [Has08]. NESM [MSR^{+}16]. Nest [Dub05c]. Netherlands [MOBD^{+}22]. Network [Cas16, NCB^{+}05, Put16, ZZS^{+}19, ZGR^{+}17, ZZC^{+}19, ZZYNH06]. Networking [ALH15, Ano18-43, Ano18-44, BMC99, Hoe10, Thi06]. Networks [ABK^{+}02, Day13c, Fox01, Gor06c, JLP^{+}10, LTD11, LWG19, PI16, Seg99, ZYKG04]. Neural [CWOL11, Fra07, Gor06c, JLP^{+}10, ZZS^{+}19, ZZYNH06]. Neurocognitive [Muz19]. Neuroimaging [SL03]. Neuroinformatics [MWC^{+}16]. Neuromorph [Boa17]. Neuronal [Seg99]. Neuroscience [BFF12]. Neutrino [Cho07b]. Neutron [ZS23]. Never [Dub07d]. News [BCC^{+}99, Bur99, CC99, GJ03a, GJ03b, Gor03, Gor04a, GH04, Gor04b, Gor04c, Gor05c, Gor05d, Har04b, JG03, Jac03, LG03, Nob00b, Shi99, Shi00c, Shi00b, Shi00d, Shi00a, Shi01a, Tou00, Tou01]. Next [BBWW22, DS23, Ged16a, HP15, McM09, NMCM22, Sul02c, Tha14, Thi04, Thi15c, Wil17]. Next-Generation [Ged16a, HP15, McM09]. Nice [Hem10]. NMR [EWN^{+}13, XXK^{+}02]. No [Day09b, Ome06]. NOAA [War18]. Nobel [Day12e]. node [YLCZ05]. Noise [ATG05, GLS07, KPD^{+}99, Kus06b]. Nominate [Ano19-38]. Nominations [Ano14y, Ano15g, Ano16c, Ano16d, Ano16n, Ano16o, Ano16p, Ano17a, Ano17b, Ano17p, Ano18b, Ano20n, Ano20s, Ano20-36, Ano21y, Ano21z, Ano21-48, Ano22c]. Nominees [Ano15c, Ano16c, Ano16d, Ano17b]. Nonconventional [ZAF^{+}01]. Nonequilibrium [MCAA05, dKCAY00]. Nonlinear [Bea00, FL05, JCC^{+}10, LWG19, MS07, Rus02, YCKK03, ZB04]. Nonstationarity [ZB04]. Nonuniformly [HH06]. Nonvolatile [VM15]. Norm [O’L05d]. Normal [KS13]. Nose [LQZL19]. Note [ACF18, NSR10]. Notebooks [PFS21, Tan21, WBB^{+}20]. Novel [CXC^{+}20, FMB^{+}07, KTG08, Kin09, LCC^{+}19]. Novelty [Cho06e]. Novice [Sma12]. Novo [GLS11]. Novo-G [GLS11]. NP [Sul04c]. NSAP [LZZ17]. NSF [Dub15b, Got15, Got16, Got17, Mor15, WNDV21]. NSF’s [WDC18]. Nuclear [Liu11]. Numba [BS21, KWBB22]. Number [ABNZ09, Ano05a, KM12, Peg12]. Number-Crunching [Peg12]. Numbers [Bau08, Hil15]. Numerical [BKB20, CBS14, Die12, Ful06, GRS08, HH22, HT99, Hu07, KL07, Lud13, LL11, MGZ00, Moy06, MSS09, Nob02b, Pes03, Pey11a, Pey11b, Pey11c, Ram18, RK05, SA08a, SA08b, SS11, STTV05, Sul99b, Sul06b, Tur14b, WWJH20, XBK10, ZZYNH06, ZS23, vdWCV11]. NumPy [PSSP15, vdWCV11]. Nvidia [EH22, HKB12].

Oak [MTBG^{+}22]. Object

[BJ02, Fox02c, GRE99, TSKG03, YaL10]. Object-Oriented [BJ02]. Objective [RRN20]. Objects [And11, Läu08, RMX12, Tof09b, IK05, Tob05]. Obscure [Shi00a]. Observation [BVT^{+}21]. Observations [The03]. Observatories [BHF^{+}08]. Observatory [Run05]. Observer [Shi02b]. Obtaining [Azo06]. Occasion [Pre09]. Occupation [HSJ^{+}19]. Ocean [BHF^{+}08, MRHP23, WHM^{+}02]. oceanographic [IK05]. off [NLV99]. Office [MWE08]. offs [PKST08c]. Oh [Sul02d].

OK [Day11f]. Olfactory [WJ04]. Olio [Shi00a]. Olive [GYF^{+}10]. OMEN [KL10]. Once [Smi01a]. One [Ano17-29, Ano17-30, Ano18-42, Bar11, BOS07, Day11b, Dub15b, Wil16]. Online [Ano15e, COS^{+}15, GPMSC20, GDDR16, Mar02, WCC^{+}19]. Only [Smi99b]. Onward [Sul01a]. OOF [LFC01]. Open [ABC^{+}14, AM15, Ano19-28, Ano19-27, Ano20-33, Ano20-34, Ano20-35, Ano20-56, Ano20-57, Ano21-42, Ano21-43, Ano21-44, Ano21-40, Ano21-41, Ano21-45, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano23t, Bar20b, BCB07, CC03, CBB06, EAF^{+}23, GSK^{+}23, KBLE15, KNS18, MMP^{+}22, MCGA22, Owe01, PGC21, Thi12a, Tow18, WCHRM21, WBB^{+}20, WCC^{+}19, JRD^{+}13, LFN^{+}11, PGF^{+}15]. Open-Circuit [WCC^{+}19]. Open-Source [CC03, CBB06, MMP^{+}22, MCGA22, Owe01, PGC21, WCHRM21]. OpenACC [STH22, SMSC22]. OpenCL [BS21, Di 14, PCY14, RGD13, SGS10]. OpenGL [XYC05]. Opening [DiD03]. OpenMM [EP10]. OpenMP [BST^{+}13]. OpenOffice.org [DiD03]. OpenPNM [Put16]. Opens [SP18]. OpenSees [McK11]. Operation [BDCT05]. Operational [Cos22]. Opinion [OHK23]. Opportunities [Ano23o, BLO^{+}22, DPG^{+}12, EUD15, GPL09, Guo23, Hin18a, LM07b, SL18, TGP13, VM15, WPW11, WCP17, THGS07]. Optical [GBPR11, HBD^{+}23, HOÖ99, HJLH03, SGA03, ZMM03]. Optical-Disk [ZMM03]. Optics [Kup03, Oug03]. Optimal [ARO^{+}11, BKB20, BHL99, Bet99, MGFRL^{+}12, WQLZ18, Zei17]. Optimization [Boe00, EVL^{+}17, PSA14, SH10, SR12, XZL^{+}19, YAA^{+}00, ZLTX19]. Optimize [For01]. Optimizing [DADY15]. Option [GEH^{+}99]. Option-Pricing [GEH^{+}99]. Options [Ano17x, Ano17y, Ano17z, Ano17-27, Ano19-28, Ano19-27, Ano20-57, Ano17-28, Ano23o]. Orbit [NLV99]. Order [ASM^{+}14, CPdlF^{+}12, Pie04, SÜP^{+}11, SSG16]. Oregon [Lan04]. Org [SD11]. Org-Mode [SD11]. Organ [XYC^{+}09]. Organic [BMCC21]. Organisms [CAS^{+}07, CFCD04]. Organizational [Cos22]. Organizations [Deb18, KMB^{+}19]. Organizing [PBD^{+}11]. Oriented [BJ02, DGR^{+}05, FGRS17, LZZ17]. Origin [Saa22]. Orthogonal [Bal99, GYL^{+}17, Rei21a, Rei21b]. Orthogonality [Nob02b]. Oscilloscopes [Tou01]. Other [BS06b, HBG^{+}20, PMM^{+}08, Shi03]. Our [Ano20-71, Bec15, Cho05f, CW06, CMN00, Day16d, Dub08a, WM00]. Outlook [ACG^{+}20]. Output [WCC^{+}19, YMHQ19]. Outreach [HPC20, JPMG08]. Outward [Cho04]. Over-the-Horizon [Cho05b]. Overcome [Gar17]. Overcoming [GK22]. Overlap [Ale15]. Overload [Thi13b]. Overview [FB04, Han03, Ork09]. Oxygen [MM04]. Oxygen-Iodine [MM04].

P [Sul04c]. Pace [Cus14]. Package [FM19, PMFM14, Put16]. Packages [BUS21, Guo12]. Packard [Got02a]. Page [Cho08e]. Pages [Mal00]. Pain [ACQ^{+}20]. Paleontology [Lew02b]. Palm [Ano15i]. Pandemic [Jef21, PHW^{+}21, PQQ20, WCHRM21]. Pandemics [RSZ^{+}20, RSZ^{+}21]. Panel [DMT^{+}21]. Papers [Ano14d, Ano14a, Ano14u, Ano15e, Ano15f, Ano15w, Ano16l, Ano18g, Ano20t, Ano20u, Ano20v, Ano20-44, Ano20-45, Ano20-54, Ano21-32, Ano21-33, Ano21-53, Ano21-54, Ano21-55, Ano21-56, Ano21-57, Ano21-58, Ano22x, Ano22y, Ano22w, Ano22z, Ano22-41, Ano22-42, Ano23w, Ano23v, Ano23x, Ano23y, Ano13a, Ano14b, Ano14c, Ano15d]. Parachute [STB03]. Paradigm [AAGH17a, AAGH17b, TX08, ZGR^{+}17]. Paradigms [PARD13, ZAF^{+}01]. Parallel [BGHR06, BBG^{+}01, BRS22, CF03, CWOL11, CNC10, CL12, CS14, CS15, DBH^{+}02, Di 14, DPP^{+}01, Esq11, Fal09, Fox02b, FPRK16, GHK^{+}08, HW15, Hil15, Hin07, Jer13, KKO^{+}20, MAG21, Ong02, QPCJ07, Raf16, Rag06, Rag07, Rob13, SA08b, SCW^{+}17, SÖS^{+}00, SGS10, TLD02, VGM^{+}09, VCGS11, Vor01a, WCAL14, Wol21, Zei17]. Parallelism [BUS21]. Parallelizing [BST^{+}13]. Parameters [Par00b]. Parametric [But99]. Parent [RMX12]. PaRSEC [BBD^{+}13]. Part [AAGH17a, AAGH17b, CENT22b, DdS23b, For16a, HHZK01b, HHZK01a, KC09a, KC09b, KC09c, Pey11b, Pey11c, PNL^{+}16, Tof08, Tof09a, Tof09b, BC05b, CENT22a, CW05b, DdS23a, DR05c, DR05a, Don06a, Don06b, Läu06, LTG07, LT09, Mal07a, Mal07b, Pos04b, Rus01a, Rus01b, Rus02, Rus03, RD05a, RD05b, SA08a, SA08b, WHG21]. Partial [GWW09, JWEK06, JHJ01, MSL^{+}07, O’L06e, O’L07b]. Participation [Ano23-41, BPW^{+}20, HG00, HC17, Slo16]. Particle [BRS22, Cre99, Dec15, JMEL08, Liu11]. Particle-in-Cell [Dec15]. Particles [HL00, MB11]. Partition [O’L04e]. Partitioning [AB03]. Partly [SS09]. Partnership [KHG^{+}23]. Partnerships [TNR^{+}23]. Party [Sul07a]. Passing [Bau08, BBG^{+}01, Fox02b, Vin12]. Past [Cho05f, Cho07c, CW20, Dau99, JH18, SACR21, WP22]. Patch [Hym05]. Pathological [Sul04a]. Paths [ARO^{+}11]. Pathway [Asr04, MPR18]. Patients [yFZDY13, SSG16]. Patron [Smi00a]. Pattern [Ano15f, SSP06]. Patterning [Rei02]. Patterns [AAB^{+}13, Bli02, QSEQJFH20, Rob13]. Pay [BS99a]. PCell [CFA04]. PCi [Tou00]. PDACS [MRU^{+}15]. PDE [Roa04, SBB^{+}15]. PDE-Solving [SBB^{+}15]. PDEs [Bal99, Wei21]. PE [BT10a]. Peaks [XXK^{+}02]. Pedagogical [NG20, Tan21]. Pedagogy [YM14]. Peer [Ano09b, Ano10c, Ano11d, Cho06f, Fox01, RBC^{+}19, ZYKG04]. Peer-to-Peer [Fox01, ZYKG04]. Pegasus [DVR^{+}19]. Pendulum [Mar02, OS03]. Pentagon [HPKS04]. People [SLK^{+}20]. Percent [Smi01b]. Perception [Lof03, Rob04]. Perceptions [CHHB13, MM12]. Percolation [Ebr10]. Perfect [Pre09, Ano16e]. Perfectly [Sch17a]. Performance [AE22, Amo18, AMS14, Bak10, BKRT21, Bar21c, BCC^{+}09, BS21, BCH^{+}22, CDL^{+}23, CJTH^{+}13, Cou22, DD05, DCWH07, DMA^{+}21, DGG^{+}21, DSSS05, DMT^{+}21, DBB^{+}21, EDJ^{+}10, EHG01, FHM99, Gob05, HP20, HC99, HB08, HG02, HP04, JPMG08, JPK01, KWBB22, KTG08, KWB^{+}10, KT11, KBLD08, LM08, Lat16, MS22, Men18, PGF^{+}15, Pap16, Par22, PA12, PSJ^{+}21, PSA14, PMP21, Pos11, QL19, RR21, SMM^{+}11, SW10, SBZ^{+}08, SDA^{+}14, SBB^{+}15, SKC02, Sch18, SOH13, SR12, SJDV09, SS09, Str10, TBVP^{+}21, VM15, WBS^{+}22, YBBP15, Zei17, SL99]. Performance-Optimization [SR12]. Performant [HLRW17, Wol21]. Periodic [Pos16]. Permanent [Gor07d]. Permeation [TLR10]. Persistence [DKWL17, JH16, TL04a]. Persistent [TMMB18]. Person [Bar11, O’L04c]. Persona [MMR22]. Persona-Based [MMR22]. Personal [Che09, The03]. Perspective [Ben04, BERT09, CSS00, CMN00, Deb18, DVS22, Fla17, HSP^{+}23, JaJ00, MDK16, Men18, PA22, Ric99, SPW^{+}13, YRT^{+}00, Wep15]. Perspectives [QL19, SACR21]. Pervasive [Ano16-39, Ano20q, Ano20r, Ano20g, Ano21-29, Ano21-30, Ano21-31, Ano22v, Ano23g, Day17d]. Petaflop [GIF^{+}12, LSV^{+}07]. Petaflops [Day10d]. Petascale [CDKF15, DW01, Dun09, DST^{+}09, GHT^{+}10, GPL09, Gro09, Kog09, Wit21, YMK11]. PETSc [ABM^{+}22]. pF3D [LBS14].

Phantom [O’L13, WNZ^{+}17]. Phase [BMS99, Hus22, IBPV03, Lan99, Lun01, WWJH20]. Phenomena [Ebr10, Lan99, Run03]. Phenotype [SRM^{+}07]. Phi [BHC^{+}15]. Photon [SG00]. Photon-Beam [SG00]. Photonic [PGC21]. Photorealistic [LCY08]. Photosynthetic [Hin17a]. Photovoltaic [KNG10]. Phylodynamic [Rao16]. Phylogenetic [HE05]. Phylogeography [Rao16]. Physical [EUD15, Fra02, Lau02, Tie16]. Physicists [Far99, Mer02, Sch23a]. Physics [AMS14, Aya07, Aya14, Bäc07a, BCB07, BT10b, BBW^{+}20, BkCP22, CF99b, CYW01, Cho06d, Cho09a, CSS00, DLB^{+}07, DAKM16, DGJ^{+}08, Ful06, GS03, Gor07d, Got17, GL20, Han05, HWPS16, KMSH10, Lan04, Lan06, LG10, LPB13, Liu11, MAFM21, MMP^{+}22, MCAA05, Mar17, Mas06, Par12, Pat02, PMW20, Raf16, Ran06, Roo06, Sch15, Sch17b, TK06, TNV^{+}02, Tow09, Tur14b, Win06, Zak18, Wep15, Ano18d, Ano18e, Ano18f]. Physics-Based [CF99b, CYW01, DAKM16, HWPS16, KMSH10, Raf16, TNV^{+}02]. Physlets [BC03]. Pi [MAFM21]. PIC [ECK^{+}15]. Picture [Rob06, Sku04]. Pierro [Wep15]. Pipeline [Che03, EWN^{+}13, STG08]. Pipelines [VSB^{+}21]. Pitaevskii [STTV05]. Pitch [OS04]. Pits [LQZL19]. Place [BSD07, Dub04]. Placenta [SRM^{+}07]. Placing [LM07a]. Plan [MKM^{+}14]. Planet [VPL18]. Planetarium [Shi02b]. Planetary [SV14]. Planning [Lew99b]. Plans [Che17a, O’L06e, O’L06g]. Plasma [CFA04, LBS14, SJDV09, TWE14]. Plasmas [GPZ^{+}04]. Plate [MGZ00]. Platform [CWOL11, DAKM16, DL23, FPRK16, GHKR11, Liu11, MK10, PTML09, SAK^{+}13, WCHRM21, WGJ16, XYC^{+}09]. Platforms [ACQ^{+}20, Has12, JON^{+}21]. Play [Bal15, DD05]. PlayStation [KBLD08]. Plone [TL04b]. Plots [WCC^{+}02]. Plotting [CCSS08]. Plug [DD05]. Plug-and-Play [DD05]. Plus [BUS21, Rob13]. pMatlab [MBB^{+}09]. Podcast [Ano22-66, Ano22-67, Ano23-47, Ano23-48, Ano23-49]. Poetry [Day07a]. Point [ASLK22, Bai05, PPE00, TM14, ZSM^{+}22]. Points [CMK22, Gar17]. Poisson [Tho01]. Polar [LM07b]. Polarization [BNNM04]. Polarized [KHC^{+}07]. Policies [LMC20]. Policy [Bei10c, SPJ^{+}14]. Polling [Dub04]. Polymerization [WPZ00]. Polynomials [Bal99, Rei21a, Rei21b, Rus01a]. Pomeau [Sch23c]. Pop [Smi01a]. Popand [Smi01a]. Popes [Day19b]. Popular [Bur99, Has12]. Popular-but-Seemingly-Dissimilar [Has12]. Population [Fen06, GLS07]. Pore [Put16, TAM^{+}14]. Pore-Scale [TAM^{+}14]. Porous [KM99, MJAK09, Sah03]. Portability [DGG^{+}21, DMT^{+}21, DBB^{+}21, PSJ^{+}21, TBVP^{+}21]. Portable [BFH21, Di 14, EGFL12, Guo12, HLRW17, STH22, Wol21]. Portal [MRU^{+}15]. Portfolio [HL01]. Portfolios [BHL99]. Position [HAB17]. Post [Lau05, Läu06, LTG07, Sor19]. Post-EJB [Lau05, Läu06, LTG07]. Post-K [Sor19]. Postal [Smi99e]. Postdocs [Ano15-37]. Posterior [HSJ^{+}19]. Posterity [PLH^{+}18]. Postprocessing [KC09a, KC09b, KC09c]. Postsecondary [Biz16]. Potential [yFZDY13]. Potentials [Hus22]. Power [Day14d, GKG^{+}15, Har18, Hin16, Hoe10, SES^{+}11, WM00]. Power-Efficient [Hoe10, SES^{+}11]. Powered [Col18]. Practical [VMH05]. Practically [TM14]. Practice [BBM^{+}21, KPA^{+}16, KJ04, MS22, MMR22, PMP21, RBC^{+}19]. Practice-Centered [KPA^{+}16]. Practices [ACG^{+}20, CJL^{+}18, CHH^{+}13, Dub22, Dub99, KHS09, KVP^{+}16, NRG^{+}17, PARD13, SHPL12, WBB^{+}20]. Praxis [Bar19]. Precession [MW14]. Precise [Nob02b]. Precision [Bai05, GLTZ10, Hin18a, Smi03]. Predator [Pek04]. predictability [Mat05]. Predicting [Hei20, Lew10, MSR^{+}16, SÖS^{+}00]. Prediction [DJ02, Jq19, LMPV13, LGJ^{+}19, LAY04, MKJ07, STG11, VWL^{+}11, WZZ11, WC15, ZZS^{+}19, Zha11]. Predictions [Jon19]. Predictive [Ano16-39, GP15, KS01, WLCD01]. Prefetching [XLLJ04]. Preliminary [JH16]. Preparation [BBWW22]. Prepare [GL20, Lat16]. Preparing [Bor02, CED^{+}21, GPL09, GN08, LCG^{+}20, MDW^{+}22]. Preprocessing [RRN20]. Prescriptions [Bal99, BS99b, BS99a, BS00a, BS00b, BT01, CT00, Nob00a, ST99, Tho99a, Tho99b, Tho00, Tho01]. Present [Cho07c, JH18]. Presented [BTL19]. Presents [Tou02b]. Preserving [Bak21, PLH^{+}18]. Press [Ano15-35, Ano15-36]. Prevails [Smi99f]. Prey [Pek04]. Pricing [GEH^{+}99, SPJ^{+}14]. Primitives [Che03]. Prince [Sny13]. Princeton [Cos22]. Principal [Nob00a, OMKdSB11]. Principles [Day08a, HWPS16, O’L05a]. Printing [Ano15-38, Ano15-39]. Priorities [SDA20]. Privacy [Ano22-66, Ano22-67, Ano23-47, Ano23-48, Ano23-49, Ano19c, Ano20h, Ano20i, Ano21w, Ano22o, Ano22p, Ano23h, Ano23i]. prizes [Day12e]. Probability [Hah04, HSJ^{+}19]. Problem [ATRA00, Bar21b, Bea00, Bre17, CAS^{+}07, FAFX20, FGP99, GPC08, Kul07, MO03, MHK^{+}06, OM03, Pes03, Smi99a]. Problem-Solving [Bar21b, CAS^{+}07, GPC08, MHK^{+}06]. Problems [Ama00, Bei09b, Ben00, BT10b, Bet99, Chu21, CLC03, CG09, CS01b, Das00, DAEJ18, DV99, DMXR^{+}14, Hym05, JCPS14, LeV09, Naj08, SH10, SAC15, SFSK01, Sul02c, WB03, XBK10]. Process [Che18, GPC08, Gyu99, MBH14, RPEB14, Wan23, WLCD01]. Processes [CBS14, JLNR19, KM99, Muz19, RcK99, TAM^{+}14, dKCAY00, Mat05]. Processing [APS10, AAAH^{+}16, BMCC21, CWOL11, CS18, CS11, DSK15, DM12, Eng15, HBG^{+}20, MP09, MR13, Pey11a, Pey11b, Pey11c, Qua18, Qua19, TL04a, UM08, Van12, WOAEAG10, Zhu16]. Processings [Pey11b]. Processor [SdPV21]. Processors [Gor07c, KSP12, SJDV09]. Product [Mil17, Pos14, PG17]. Production [GKG^{+}15]. Productive [AGC^{+}16, BFH21, HLRW17, Wil06, Wol21]. Productivity [BUS21, FLV^{+}09, Moo21, MBB^{+}09, PSJ^{+}21, PK15, Thi13b]. Products [Shi00d, Shi00a]. Professional [GPMSC20, Peo20, Tho12, Ano21v, Ano22u]. Professionally [ARAG19]. Program [BB20, Bur99, CFA04, CMN00, GCV08, Har23, Lan04, OASFLAB09, PAN^{+}16a, Vla12, BHV21, PNL^{+}16]. Programmer [Shi00a, Thi07]. Programmers [Esq11, Sma12]. Programming [AAGH17a, AAGH17b, BB20, BBG^{+}01, CF03, CL12, DS12, Dra00, DY99, Dub99, Dub00, Fal09, GRE99, Gra09, GS13b, Guo23, HC99, HHZK01b, HHZK01a, Hin09, Hin13a, Hin13b, Kar99, LT09, LPV00, LC12, MM18, MAG21, Nas00, PTML11, Rag06, SDS00, SL99, SB00, SGS10, Taj10, Wol21, XYC05, Wep15]. Programs [BCC^{+}99, CRDO16, Di 14, Dub05b, Dub12, Fos17, LDAS19]. Progress [GF04]. Project [Ale18, DMT^{+}21, KMSH10, KPA^{+}16, Mak06, NCB^{+}05, Owe01, PSSP15, PQQ20, Thi07, Fom15, KPM10, MB20a, Mes17, SGRK^{+}18, SJF^{+}23]. Project-Based [PQQ20]. Projection [MR13, NSP12, Rus03, YCKK03]. Projections [HKW03, dSRT16]. Projector [ML02]. Projects [BB06, COS^{+}15, Har23, HPMJ12, KL07, LWF10, PBSS14]. Prolog [BT10b]. Prologue [Dau99]. Promise [Gor06c, Pos09, Pos10]. Promises [Hin09, LT09]. Promising [Mar17, ZGR^{+}17]. Propagation [BPMKC21, GGJD23, LPV00, SA08a, SA08b]. Propellant [HD00]. Properties [JXY^{+}19, Lew10, MJAK09, Osk07, PI16, SÖS^{+}00]. Property [Cyb99c]. Prospectus [Boa17]. Protection [Lew00a]. Protein [Han03, Jav12, Mal07a, Mal07b, SK01, Wol16, WCH12, XDK^{+}20]. Protein-DNA [WCH12]. Proteins [PP20]. Protocol [Gal11, LZZ17, Zei17]. Prototypes [Mil17, Pos14]. Prototyping [FMB^{+}07, HD00, LRRK00, PL02, PS17]. Prove [Sul99b]. Provenance [AAH^{+}08, DD23, DGJ^{+}08, FKSS08, MGD^{+}08, SFC07, ST08, TJCC20]. Provide [PK18, Tou01]. Provides [CC99, Rob06, Tou00, Wol16]. Province [GYF^{+}10]. Provision [GHKZ17]. Proximate [JON^{+}21]. Proximity [MPP14]. Proxy [MDW^{+}22]. Pseudopotential [SAC15]. Public [Day13e, MB20b, MBB^{+}22, RTSS14b, Ree16, SKA^{+}02]. Public-Interest [SKA^{+}02]. Publication [Ano22f, Che18, CW05e]. Publications [Ano22-59, Ano22-68, Ano23-50, FS12]. Publish [KNS18, Thi14]. Published [Ano20-33, Ano20-34, Ano20-35, Ano21-42, Ano21-43, Ano21-44, Ano21-40, Ano21-41, Ano21-45, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano23t]. Publishing [Ano13c, Thi15b]. Pulsar [MKH^{+}23]. Pulsars [MKH^{+}23]. Pulses [TDB09]. Pulverized [MRKK17]. Pure [LT09]. Purpose [FHM99, Got06, SSC18]. Purposes [Wat21]. Pursuit [Sic09]. Putting [LTNME09, Mar99a, WC15]. PyCOMPSs [BCE^{+}22]. PyExaFMM [KWBB22]. PyFR [Wit21]. PyMOL [Moo21]. PyNSol [Tob05]. PyOMP [MAG21]. PySPLIT [War18]. Python [Wep15, APS10, Aya14, Bäc07a, Bar21b, BKRT21, BDF^{+}20, BUS21, BLP^{+}21, BVB^{+}07, Cot03, Day14d, Di 14, DY99, Dub07e, Gre07, Gue18, GWW09, Hin07, KWBB22, KB07, LFN^{+}11, LeV09, MAFM21, MMP^{+}22, MSL^{+}07, MAG21, MB07, MA11, MS07, MGS07, Oli07, PGH11, Ram18, Sch17a, SMF^{+}23, Shi07, Sma12, Tie16, TGEA09, VB08, Vir16, War18, Wit21]. Python-Based [Aya14, Ram18]. Python/Jupyter [Bar21b]. Pythons [Dub05c]. Pythran [Gue18].

QCD [Mor15]. QMD [KFS18]. QPACE [GHK^{+}08]. QR [ACF18]. QR-Based [ACF18]. Quadratures [Bal99]. Qualitative [DKWL17]. Quality [Ano22-75, Dub22, HLS^{+}16, MMR22, RPEB14, Ste99]. Quantification [XKG05]. Quantify [JJZC10]. Quantity [Fos17]. QuantLib [Vir16]. QuantLib-Python [Vir16]. Quantum [AE22, Ale15, Ano20-56, Ano20-55, Ano22z, Ano22-60, Ano23-40, Ano23-41, Ano23-42, Bäc07b, BC03, BCB07, Bro06, Cyb02, DCC10, Day07c, EKCS12, GHK^{+}08, KFS18, McC21, Mus20, OASFLAB09, PMW20, SDA^{+}14, SRKS22, UM08, VMH05, WOAEAG10, WB03]. Quarks [Cre04]. QUBES [DL23]. Qubit [Key05, Lew01a]. Queer [Wri16]. Queries [MPP14, O’L14, ZSM^{+}22]. Query [ALH^{+}20, AAAH^{+}16, LT08]. Query-Based [ALH^{+}20]. Quest [DW01, Mat05]. Questions [Sul03b]. Quick [Vor01b]. Quickhull [Muc09]. Quickly [Muc09]. Quicksort [JaJ00].

R [FM19, Lud13]. R&D [PS17, Sch99]. Race [DGK19]. Racial [Bot16]. Rad [Lew99b]. Radar [Cho05b, Zeb00]. Radial [LWG19]. Radiance [UGV11]. Radiation [Lew99b, MSS09, SWDR22, SG00, VWL^{+}11, ZFS12]. Radio [KLMS99, PAN^{+}16b]. Radiometers [CPdlF^{+}12]. Radiotherapy [LSDP^{+}04]. Rafters [KMM^{+}11]. Railway [ZZS^{+}19]. Railways [MGFRL^{+}12, RLHGA^{+}13]. Rainbow [Ano22-66, Ano22-67, Ano23-47, Ano23-48, Ano23-49]. Rainfall [Gan02]. Raising [SSW21]. Ramakrishna [Ano17a]. Random [ALM19, Hil15]. Range [MR13, PSR^{+}00]. Rapid [FCT^{+}10, LRRK00]. Rapidly [MPRU23, Ran06]. Rare [KHG^{+}23]. Raspberry [MAFM21]. Rate [ZB04]. Rates [O’L04b]. Ratio [GYJL20]. Rau [Ano17a]. Ray [SWDR22, ZSM^{+}22]. Ray-Tracing [ZSM^{+}22]. RBF [WLCJ12]. RBPCA [SMM^{+}11]. RC [MAFM21]. Re [Hin18d, MCE^{+}03]. Re-editable [Hin18d]. Re-Integrating [MCE^{+}03]. Reach [WC15]. Reaching [Gre07]. Reactive [JC02, LNI^{+}19, TAM^{+}14]. Reactor [BAD^{+}21]. Reader [CW06, Day14c]. Readiness [CHM^{+}20b, CHM^{+}20a]. Reading [Day12c, Day12d]. Real [CYW01, Clo15, CC99, EKCS12, HE05, LCY08, PSA14, SBZ^{+}08, VWL^{+}11, VSG^{+}02, WHM^{+}02, YZZ04, YHWY05]. Real-Life [HE05]. Real-Time [CYW01, CC99, LCY08, PSA14, VWL^{+}11, WHM^{+}02, YZZ04, YHWY05, VSG^{+}02]. Realistic [RSC^{+}14, WLCJ12]. Realistically [KL10]. Reality [ACQ^{+}20, DBJ^{+}20, MV20, SLK^{+}20, SPB^{+}20, TW03, YWC02]. Realizations [LPB15]. Realizing [CCD^{+}22]. Really [Day07c, Day16e]. Reason [Day14e]. Reasoning [SGRK^{+}18, Tec18, TKM^{+}18, vGDS18]. Reasons [Poi10]. Reckless [LGG^{+}22]. Reclusive [BC02]. Recognition [Ano15f, Ano16s, FOdLVF^{+}11, Mal07b, NWP19, OMKdSB11, SMM^{+}11, ZZC^{+}19]. Recognizing [MBB^{+}22]. Recommended [KHS09]. Recommender [FAFX20]. Recommending [Cho07f]. Reconfigurable [EVA^{+}21, GLS11, KSB07, Liu11]. Reconstruct [HMA00]. Reconstruction [FKB^{+}13, yFZDY13, Lan02, LWT^{+}13, ZDW^{+}07]. Record [MCE^{+}03]. Records [Tou03]. Recovered [Sch21]. Recurses [Nob03]. Recursive [GYL^{+}17]. Recycling [Sul99a]. Reduce [JG12]. Reduced [ASM^{+}14]. Reduced-Order [ASM^{+}14]. Reducing [DPBS16]. Reduction [ACF18, BAD^{+}21, GYL^{+}17, LQZL19, SL03, Vil08, Vor01a]. Redux [MHDM99]. Referencing [ACG^{+}20, CGZ20]. Refined [BW06]. Refinement [ATG05, Bry99, DBB^{+}21, MCAA05, NSLD99]. Reflect [SCBT18]. Reflecting [BLT^{+}23, SBW^{+}19]. Reflection [Che09, Chu21]. Reforms [For99]. Region [KSSF11, PL07]. Region-Growing [PL07]. Regional [ASLK22, PSR^{+}20, WC15]. Regional-Scale [ASLK22]. Regions [LM07b, Rao16]. Register [Ano20-37, Ano23-42]. Regression [CT00, FM19, GYL^{+}17, RPBE12, SL03, Wan23]. Regression-Test [RPBE12]. Regularization [LHGX18]. Rehabilitation [YMHQ19]. Reinforced [JXY^{+}19]. Reinventing [Har23]. Reinvigorate [TK06]. Related [HEH^{+}10]. Relation [Bai00, ZLW^{+}19]. Relational [TSKG03]. Relations [LSPN21]. Relationships [Rei21b, Rei21a]. Relativistic [FMB^{+}07, MW11a, MW11b]. Relevant [BCL03, HC17, MM13]. Reliable [GB20]. RELM [STHR12]. Remaining [LGJ^{+}19]. Remembrance [O’L05b]. Remote [DDV^{+}08, LM07b, VSMD^{+}09]. Rendered [KEF07]. Rendering [BHC^{+}08, CCSS08, Che03, CZ07, DLLZ19, DLLZ20, LYC07, NGGS22, QPCJ07, SRM^{+}07, SGA^{+}22, YWC02]. Repealed [Fla17]. Repetitive [YMHQ19]. Replicable [MB17]. Report [Car09b, CSW17, MJM^{+}06, OW01, SSW11]. Repositories [AAB^{+}21, IKMK13]. Repository [LHN^{+}12]. Representation [Biz16, QSEQJFH20]. Reproduced [Sul03e]. Reproducibility [AISR^{+}21, Bar21a, Cou22, Dav12, DD23, Die12, PRCL^{+}22, PMP21, SFC07, SMS15]. Reproducible [Ano10b, BT17, Bar19, BTK^{+}21, DMR^{+}09, FSED10, FC09, Fom15, FS12, Hil15, Hin13c, How12, LeV09, MB17, MB20b, Muh23, PE09, SKC00, Sto09, SLM12, Sto12, Sul03e].

Reproducing [BVT^{+}21]. Research [ACG^{+}20, Ano10b, BT17, BLMR21, BPW^{+}20, BLT^{+}23, BVB^{+}07, CCP23, CGK^{+}18, CENT22a, CENT22b, CHP^{+}18, Che18, Cho12, Chu21, Cos22, CMK22, CHH^{+}13, Dar18, Dav12, DD23, DdS23a, DdS23b, DMR^{+}09, Dru20, Dub05a, ERF21, EI11, ECK^{+}15, FSED10, FC09, Fom15, GAB^{+}22, Gor08a, GL20, GHKR11, Har23, Her22, Hil15, Hin13c, How12, HBG^{+}20, Hu07, JLNR19, KBLE15, KL15, KMB^{+}19, KMB^{+}08, KJ04, LC09, LeV09, Lof22, MS22, MTBG^{+}22, MOBD^{+}22, McC21, McM09, Mes15, MJM^{+}06, MBB^{+}22, MCE^{+}03, PSR^{+}20, PE09, Peo20, PAN^{+}16a, PNL^{+}16, RBC^{+}19, Rob06, Run05, ST05, SL18, SW10, SYM^{+}21, Sch23a, SP18, SKA^{+}02, Sto09, SLM12, Sto12, SJF^{+}23, Van12, WZS^{+}10, WPM^{+}12, WG15, WCP17, Wol16, ZFS12]. Research-Based [Peo20]. ResearchCompendia.org [SMS15]. Researcher [Moo21]. Researchers [CE18, Ged16a]. Residual [DKCL14]. Resistance [Pon16]. Resolution [DWC^{+}11, Lan02, TAM^{+}14, WSY^{+}22, YCD^{+}21, ZQY^{+}11]. Resolved [Smi99a]. Resolving [WZZ11]. Resonance [But99, Kus06a, MB07]. Resonances [Mor15]. Resource [FWGB07, HBG^{+}20, NCB^{+}05, TFF05]. Resource-Aware [TFF05]. Resources [Asr04, DMXR^{+}14, KB04, MHC^{+}18, Mor15, PGF^{+}15, Ano16y]. RESPECT [DE17, For16a, WBP^{+}19]. Response [AGM^{+}00, Bor02, CW05a, TDB09, WPM^{+}12, ZZS^{+}19]. Responses [ZLTX19]. Responsibly [LGG^{+}22]. Responsive [NG20]. Restoration [Pey11c]. Results [CW06, DG12, PMP21, Sul03e]. Resume [Ano19-28, Ano23o]. Retain [DBB^{+}21]. Retaining [MBB^{+}22]. Rethinking [MDW^{+}22]. Retouching [Tof09b]. Retrieval [KHE13, Pok04, ZYKG04]. Retrospective [McC21, WP22]. Reusable [ARAG19, Bar20b, Hin18d, LFK^{+}19]. Reuse [Lan19]. Reveal [Gor05b, ASM^{+}14]. Revealed [Thi13a]. Revealing [XDK^{+}20]. Review [BO04, ERF21, Gra07, KS20, Lud13, QL19, RBC^{+}19, Sha14, Sny13, Sul05a, Bas14, Kwa17, Mol12, Nai15, Tur14b, Vog13, Wep15]. reviewed [TCCC13]. Reviewer [Ano15a, Ano16a, Ano18a]. Reviewers [Ano05b, Ano09b, Ano10c, Ano11d, Ano12b, Cho06f]. Reviewing [Sha14]. Reviews [Ara99, BC02, Bil00, BCC^{+}99, Bur99, CW05c, CW05e, CC99, Cyb01, Fel00, Lov04, McK00, Nob00b, Seg99, Shi99, Shi00c, Shi00b, Shi00d, Shi00a, Shi01a, Tou00, Tou01]. Revisited [Got16]. Revisiting [SR13]. Revitalizing [Bei10d]. Revolution [Bei10e, Wei11]. Ribonucleic [Maj03]. Rica [Men18]. Richard [Ano17-31, BTL19, Ano14-43, Ano17-31, Ano21-74]. Ridge [MTBG^{+}22]. Right [BS02, Nob02b, Sul03f]. Rights [Ano15-44]. Rigid [NLGNJ13]. Rigid-Molecule [NLGNJ13]. Ring [AK04]. Rio [WSC^{+}04]. RISC [TJ14]. Rise [DiP18a]. Risk [Ano16-40, Hei20, KPA^{+}16, KVP^{+}16, Rao16]. Risk-Based [Ano16-40, KPA^{+}16, KVP^{+}16]. Risks [Tur16]. River [FGR^{+}07]. Road [Bei09c, Haa99]. Roadmap [SBW^{+}19]. Roadrunner [HW09]. Robot [Day18a, Day18d, OS03]. Robotic [DiP18b]. Robots [KB07]. Robust [GNB^{+}09, LM07a, Zei17]. Robustness [YAA^{+}00]. Rock [Ano14-44, Ano14-45, Ano14-46, Ano14-47, Ano14-48, Ano15-43, Ano15-40, Ano15-41, Ano15-42, Ano15-38, Ano15-39, Ano16z, Ano16-27, Ano16-38, Ano16-37, Ano16-39, Ano16-40, Jon19]. Rockets [HD00]. Role [BSD07, Cra03, Das00, JC02, LSBC22]. Roles [Bor02, Hin17c]. Roll [OS04]. Rolling [KS06, MGFRL^{+}12, ZZC^{+}19]. Rolling-Stock [MGFRL^{+}12]. Rotations [ZJW08]. Rotorcraft [HWPS16, Str10].

Rough [ZLW^{+}19]. Round [LQZL19]. Roundtable [ZAF^{+}01]. Route [Gor06b, Kus06a, MGFRL^{+}12]. Router [Rob06]. Router-Level [Rob06]. Routine [Ron14]. Routines [Dub00]. Rs [Day11b]. RSEs [MBB^{+}22]. Rubber [Bei09c, Haa99]. Ruby [Ong02]. Rule [Day11f, Rei21a, Rei21b]. Rulers [Mem16]. Rules [DTL^{+}17]. Run [BKS15]. Run-Time [BKS15]. Running [CJL^{+}18, GL20, RRAB06]. Runtime [DMA^{+}21]. Rupture [KMM^{+}11]. Rupture-to-Rafters [KMM^{+}11].

S [Ano22-75, Mol12]. Safari [Ben09]. Sage [Gra08b]. SambaNova [EVA^{+}21]. Sampled [HH06]. Sampling [ALM19, BS99b, Bez08, RBK02]. Samsung [Tou02b]. San [LC09]. Sand [Bil00]. SAR [MR13]. SARS [FGG^{+}22, XDK^{+}20]. SARS-CoV-2 [FGG^{+}22, XDK^{+}20]. Satellite [BCA^{+}00, HAB17, Lo99]. Satellite-Constellation [Lo99]. Savings [SW10]. Say [Bei10e, Got02a, Sul04d]. Says [Dub07b]. SC98 [Dau99]. scaffolding [Tob05]. Scala [PTML11]. Scala-Based [PTML11]. Scalability [BBD^{+}13, YLZ17]. Scalable [BFL^{+}22, BLT^{+}23, BFH21, Bry11, GLS11, GARS^{+}20, KSM^{+}17, LNI^{+}19, Liu15, MSS09, NKV99, Par16, RSC^{+}14, SJF^{+}23]. ScalaLab [PTML11]. Scale [ASLK22, Ama00, BBM^{+}15, BP99, CS14, CS15, Day14a, DMXR^{+}14, Eis17, Far99, GL99, Gob05, HHF^{+}14, HMB^{+}14, JPE20, Jon15, KMSH10, LWF10, LCY^{+}04, Ma16, MWE08, Rei02, Sah03, SNTL13, Ste99, TB11, TWE14, TAM^{+}14, VKN99, VM15, Wes21, WNDV21, WPZ00, WHW18]. Scales [Gyu99, Kus06a]. Scaling [GS03, Run03]. Scarce [RBK02]. SCC [GHKR11]. Scenario [AAGH17a, AAGH17b]. Scenario-Based [AAGH17a, AAGH17b]. Scenarios [FGG^{+}22, PQQ20, UGV11, WCHRM21].

Scene [ML02]. Scheduling [WQLZ18]. Scheme [PLW17]. Scholars [Day13a, Slo16]. Scholarship [Ano17-31, Ano21-74, Ano17-31]. Scholes [Hig04]. School [AAGH17a, AAGH17b, GPMSC20, Les16, Sca16]. Schools [Ree16, Sul05b]. Schrödinger [Moy06, RK05]. SciDB [SBZB13, YBBP15]. Science [AM18, Adl20, ABC^{+}14, AHL^{+}11, AAGH17a, AAGH17b, APC^{+}19, Ama06, AMS14, Ano03, Ano05a, Ano12b, Ano15-37, Ano22f, BCE^{+}22, BC02, BT17, Bar19, Bar20a, Bar21b, BLMR21, Bar21c, Bei10c, BSD07, BHV21, BERT09, CDL^{+}23, Car09a, Car12, CE14, Car16, CHC17, CBF^{+}22, CC03, Chr15, Chu21, CW20, COS^{+}15, CF13, DSPY05, Day06a, Day12a, Day14a, Day18d, DADY15, DdS23a, DdS23b, DPG^{+}12, DL23, Dun09, DST^{+}09, EAF^{+}23, ESO08, FKD21, FM13, FGHW20, FL99, FM02, Fox02a, FHM99, FWGB07, Gaa03, GSK^{+}23, GDH^{+}23, GPL09, GPMSC20, Gor07c, GCV08, Gor13, GK18, GHM^{+}16, HP15, HBD^{+}23, Hei22, Her22, Hin18a, Hin15a, Hin15b, Hin17c, HG00, HGA23, HPMJ12, HHR^{+}13, JH18, JRD^{+}13, KE05, KL15, Kax01, KHG^{+}23, KB04, KMM^{+}11, KRH^{+}99, LC09, Les16, LM07b, MF16, MM12, Mas06, MMTD^{+}17]. Science [MKH^{+}23, MK10, Mem15, MV20, Mes15, MCGA22, MSM13, MSR15, NC18, NCB^{+}05, OW01, OHK23, PARD13, PA21, PV00, Pos07, Pos09, Ran06, Ree16, RRAB06, Ric99, RP20, RHC^{+}23, Run00, RF12, SDA20, Sca16, SDD^{+}08, SKNV03, SCW^{+}17, Slo16, SMS15, SKL10, SPB^{+}20, Sul03a, Sul04a, Tan21, Tes15, Thi02, TB04, Thi13a, Thi15c, TX08, TNR^{+}23, THGS07, TP04, Tur14a, VGD^{+}11, WHG21, WHH22, WMB20, WNDV21, WCP17, Wil16, Wil18, WL09, WBB^{+}20, WC17, YBBP15, YRT^{+}00, Day12e, Peo20, RTSS14a, RTSS14b, WDC18]. Science-Based [Pos09]. Sciences [BBM^{+}21, Eis17, KB09, OW01, Sul09a, TS10, WDC18]. Scientific [AAB^{+}21, ARAG19, ABC^{+}14, Ale18, Ano15-44, Bai05, Bak21, BHF^{+}08, BKRT21, Bas02, BCH^{+}09, BMC99, BTK^{+}21, BCC^{+}09, Bli02, BBM^{+}15, BC05a, BHC^{+}15, Bry11, CCSS08, CHP^{+}18, CHB19, CJL^{+}18, CHJC05, Cho03, CS11, Day08a, DC04, Di 11, DSK15, DDC04, Dra00, Dub21, Dub22, DY99, Dub99, Dub00, Dub02, Dub05b, Dub12, EVA^{+}21, EVL^{+}17, ERS^{+}03, Esq11, FLV^{+}09, Gob05, GAB^{+}22, GP15, GRE99, GNB^{+}09, HLRW17, HC99, HHZK01b, HHZK01a, Her21, HCK22, Hig04, Hin13a, Hin13b, Hin15a, Hin17b, Hin18c, HXMC05, IKMK13, JLNR19, JPE20, JPK01, JRD^{+}13, KC19, KRR^{+}12, Kar99, KSB07, Kin12, KBLD08, LDAS19, LVWK02, Lan19, LFK^{+}19, LSBC22, LSN20, LSPN21, LPV00, LGG^{+}22, MWE08, MBH14, NGGS22, Oli07, PTML09, PTML11, PSSP15, PG07, PGH11, PF04, Pos13, RV11, RVG^{+}10, RPBE12, RPEB14]. Scientific [RR21, SNCM16, Sch01, SDS00, Sch23a, Sch18, SKC00, SES^{+}11, SL99, SHPL12, Sma12, SdPV21, SSW21, SB00, Sto09, SLM12, Sto12, SJF^{+}23, Sza11, TWE14, TAF^{+}18, TFF05, Ter11, TTT15, Thi07, Toh07, TCD^{+}14, VB08, VSMD^{+}09, VMK20, Wan18, WG15, WD06, Wes21, WJLY08, YKD^{+}03, YBD10, Beh05, Wep15, BC05b]. Scientist [Got14b, Men18, Tsa14]. Scientists [Bet17, CHHB13, Cho12, DPBS16, Guo23, KHS09, KSM11, MM13, MA11, NMCM22, PMK^{+}08, THLK10, Wil06, Wri10]. Scinco [SJF^{+}23]. SciPipe [LDAS19]. Scoiety [Ano21-40, Ano21-41]. SCons [Kni05]. Scope [Cho06b, Pap16]. Score [O’L14]. Screening [NLGNJ13]. Script [WHH22]. Scripting [BJ02, Bas02, Hin07, Ong02, PTML09]. SDSS [Tha08a]. Sea [KS06, WM00]. Seamless [SKC05]. Search [Ano14l, Ano14j, Ano14k, Ano14m, Ano15j, Ano15k, Ano15l, Ano17d, Gor07a, KHE13, MGCBI17, VSB^{+}21, VMH05]. Searching [ALH^{+}20, Lew01a, Pok04, Sch01]. Seas [Lew99a]. Second [Car09b]. Secondary [FCT^{+}10]. Secure [AGC^{+}16, MCGA22]. Security [Ano16-40, Ano19c, Ano20h, Ano20i, Ano22o, Ano22p, Ano22-66, Ano22-67, Ano23h, Ano23i, Ano23-47, Ano23-48, Ano23-49, PLW17, Ano21w]. Security&Privacy [Ano14-30]. See [Bil00, NO03, Sul03a]. Seeing [Bro06, Gor08c]. Seek [Ano22-59, Ano22-68, Ano23-50]. Seemingly [Has12]. Seems [Yas17b]. Segmentation [GMPR11, PL07, WNZ^{+}17]. Seismic [AMKL04, CL14, HH06]. Seismology [TB11]. Seizure [SSG16]. Selection [CPdlF^{+}12, MGFRL^{+}12, Tof08]. Self [CHHB13, CJ16, KL07]. Self-Assessment [KL07]. Self-Healing [CJ16]. Self-Perceptions [CHHB13]. Semantic [Fox03a, Zhu02]. Semiconductor [Gyu99, KL10]. Semiconductors [KHC^{+}07]. Semisupervised [GYL^{+}17]. Sense [Wri16]. Sensing [LM07b]. Sensitive [AH07]. Sensitivity [DRA11, HK09, O’L06f]. Sensor [AMKL04, JLYL19]. Sensors [NWP19]. Separate [CGZ20]. Separated [HF04]. Sepkoski [Ome06]. September [HPKS04]. Sequel [BT10a]. Sequencing [Mye99]. Sequential [SA08a]. Sequestration [TAM^{+}14]. Series [DMT^{+}21, JSNR11, Liu15]. Server [Goo17, Nei08, TSFG08, Tha08a, YKD^{+}03]. Serverless [VPL18]. Serves [TS10]. Service [Ano14-43, Ano20-51, Ano20-52, Ano20-53, Ano23-37, Ano23-38, Ano23-39, DGR^{+}05, Dub07b, GHM^{+}16, GHKZ17, LZZ17, MF16, MWC^{+}16, WPM^{+}12, DBH^{+}02, Smi99e]. Service-Based [WPM^{+}12]. Service-Oriented [DGR^{+}05, LZZ17]. Services [Ano13c, Ano19z, Fox04a, MRU^{+}15, PF04, PGH^{+}05, GGD^{+}05]. services-based [GGD^{+}05]. Serving [Dra00]. Sets [BCG^{+}99, Don02, Gor06d, GYF^{+}10, LLQ18, MFD^{+}09, Ste12, Wan18, ZLW^{+}19]. Seventh [Ste14]. Several [Tou02b]. Sez [Sul04e]. SH [WQT^{+}16]. Shader [Wat21]. Shader-Like [Wat21]. Shadow [GRE99]. Shadow-Object [GRE99]. Shake [Bet17]. Shape [Gio02, Sul10b]. Shapes [Tof09a, VSG^{+}02]. Shared [BWC01, GCV08]. Sharing [CHP^{+}18, Cho07d, DDC04, MHC^{+}18, PLH^{+}18, Van12]. Sharp [LL19]. Shear [NTW07]. Shed [TM15]. Sheet [MRHP23]. Sheets [PAF08, TAHP23]. Ship [KSM^{+}17, KS06]. Ships [Lew99a]. Shock [Ben04, BBW^{+}20, Cyb99a, MSR^{+}16, ST05]. Shock-Accelerated [Ben04]. Shock-Wave [ST05]. Shoes [MMR22]. Short [KILZ13, Pek04]. Short-Term [KILZ13]. Should [Day14f]. Show [Gor06c]. Showcase [BLMR21]. Shrinkage [Tas00]. Shunt [DKCL14]. SIAM [BERT09]. Sichuan [GYF^{+}10]. Side [SG10]. Sides [O’L05f]. Sierra [MSR^{+}16, NdS17]. Signal [APS10, CWOL11, FSED10, Kus06b, Pey11a, Pey11b, Pey11c, Tou01]. Signals [Kwa17]. Signature [JWLG14]. Silico [Han03, KFS18]. Silicon [The03, WLCD01]. SimEvents [Gra07]. Similarities [AE22]. Similarity [GYJL20, LGJ^{+}19]. Simon [Sny13]. Simphony [PGC21]. Simple [Fox04a, KC19, MAFM21, MMP^{+}22, New00, Nob07, Sul02a]. Simplex [Nas00]. Simplicity [Hin17b, Sul00a]. Simplified [SAC15, SOH13]. Simplify [Shi00d]. Simulate [Day06c, Kul07]. Simulated [Tre99]. Simulating [AGM^{+}00, BCA^{+}00, CPdlF^{+}12, Cre04, DL00, DLLM04, EKLY07, FGRS17, HMS^{+}00, KS00, MGCBI17, Par00b, Roh10, SBH^{+}00, UGV11, VWP12, WSC^{+}04, WHW18, ZMM03]. Simulation [ASM^{+}14, AM15, BFH21, BAD^{+}21, BKS15, CDF^{+}04, Cas16, CR15, CYW01, CHC^{+}11, CJTH^{+}13, CSS00, CBS14, CFCD04, DLW^{+}19, DAKM16, Die12, DH12, Don03, FGG^{+}22, FMKS08, GW15, GSB^{+}12, Got02b, Gra07, HAB17, HBD^{+}23, HWPS16, HLYQ19, HS12, HT99, Hil15, HPKS04, HXMC05, Ikk16, IBPV03, Jon15, Kad04, KNG10, KBLE15, LSDP^{+}04, LWT^{+}13, LMC20, Lui06, LL11, MJAK09, MM04, Mal07b, MW14, McK11, Mil10, MGZ00, MM16, Muz19, NBK^{+}01, NGGS22, NW15, Ork09, Par16, PV00, PCY14, PI16, PGC21, Pos04a, Pos04b, Saa09, SA08a, SA08b, SDA^{+}14, Sch23b, SS11, SPJ^{+}14, SMSC22, Str10, SKC05, TB11, TGP13, TLR10, TAM^{+}14, TL08b, UZC^{+}12, Wai16, WCHRM21, WAS^{+}12, WWJH20, WLCD01, Wol16, YZZ04, YZC^{+}13, YM14, Yas17b, YWMM04, YYG^{+}19, Zei17, ZJW08, ZCXM99, dKCAY00, Mat05]. Simulation-Based [Mil10]. Simulation-Driven [Cas16]. Simulations [Adl03, AMCH07, Ama00, AM20, AMKL04, BMS99, Boe00, CLZ13, CSS00, CDKF15, CS14, CS15, DLB^{+}07, DJS13, DDC04, EP10, EKCS12, FPRK16, Ged16a, GL99, Gia22, GWMG04, GC00, Gob05, GPZ^{+}04, GIF^{+}12, GH00, HHR02, HL00, HSG03, HHF^{+}14, HT99, HEH^{+}10, Jav12, JRP^{+}17, KM99, KSP12, KS00, KMM^{+}11, KPD^{+}99, LBS14, LUMM14, LL13, LVLA14, LNI^{+}19, LCY^{+}04, Ma16, MRU^{+}15, Mal07a, MPR18, MTG^{+}12, MB99, Muh23, NKV99, New00, NSLD99, PZJS10, PR01, PMW20, Raf16, Rao16, RSC^{+}14, SMF^{+}23, SBW^{+}19, SS06, Sta03, Ste99, SKL10, TWE14, Tow09, VKN99, YAA^{+}00, YCD^{+}21, YB12, Zhu16, dKCAY00, RF00]. Simulator [LFN^{+}11]. Single [JXY^{+}19, OR12]. Single-Strap [JXY^{+}19]. Singular [Los03]. Singularities [Ano18j, Sin18]. Sinusoids [Rus02]. Sisyphus [Chr99]. Site [DKK05]. Situ [ASLK22, BAD^{+}21, DB21]. Six [Guo23]. Skeletal [Roh10]. Skeleton [Dec15].

Skeletons [Fal09]. Skills [GDDR16]. Skinny [DSC^{+}09]. Skunqoil [Smi99b]. Sky [HHF^{+}14, Nei08, Sza99, Tha08b]. SkyQuery [BDS13]. SkyServer [RTSS14a, RTSS14b]. Slab [YLZ^{+}19]. Slices [QPCJ07]. Slide [Sul04d]. Sloan [Nei08, Sza99, Tha08b]. Slug [LTNME09]. Small [EVL^{+}17, SL18]. Smart [LMC20, SAK^{+}13]. Smarter [Bei12c]. Smelly [Dub05a]. Smoother [YaL10]. Snake [Wit21]. Snapshot [HHP19]. Snaring [Cho07d]. Snippet [Moo21]. Sociable [Day13f]. Social [Ano18-43, Ano18-44, AM05, CHM^{+}20b, CHM^{+}20a, Day13f, LTD11, PI16, Pon16, Tan21]. Society [Ano14y, Ano17-31, Ano18z, Ano19v, Ano19y, Ano20-51, Ano20-52, Ano20-53, Ano20-50, Ano21-64, Ano21-65, Ano21-66, Ano21-67, Ano22-44, Ano22-55, Ano22-45, Ano22-56, Ano22-46, Ano22-57, Ano22-58, Ano22-47, Ano23-29, RHC^{+}23, Ano13i, Ano13h, Ano14z, Ano14-29, Ano14w, Ano14-27, Ano14x, Ano14-28, Ano15x, Ano15y, Ano15z, Ano15-27, Ano16t, Ano16q, Ano16u, Ano16r, Ano16z, Ano16v, Ano16n, Ano16s, Ano16-27, Ano16w, Ano16o, Ano16x, Ano16p, Ano16m, Ano16y, Ano17o, Ano17j, Ano17k, Ano17l, Ano17m, Ano17p, Ano17n, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano18s, Ano18r, Ano18y, Ano19t, Ano19u, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19-29, Ano19-30, Ano20-46, Ano20-38, Ano20-39, Ano20-47, Ano20-40, Ano20-48, Ano20-41, Ano20-49, Ano20-44]. Society [Ano20-42, Ano20-43, Ano20-45, Ano21-49, Ano21-53, Ano21-59, Ano21-54, Ano21-60, Ano21-55, Ano21-56, Ano21-61, Ano21-50, Ano21-57, Ano21-62, Ano21-51, Ano21-63, Ano21-52, Ano21-58, Ano22x, Ano22y, Ano22-27, Ano22-39, Ano22-48, Ano22-41, Ano22-40, Ano22-52, Ano22-49, Ano22-53, Ano22-59, Ano22-50, Ano22-42, Ano22-54, Ano22-51, Ano22-43, Ano23-31, Ano23u, Ano23w, Ano23-37, Ano23v, Ano23-34, Ano23-38, Ano23z, Ano23-30, Ano23-32, Ano23-35, Ano23-39, Ano23x, Ano23-27, Ano23-33, Ano23-36, Ano23y, Ano23-28, Ano20-33, Ano20-34, Ano20-35, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano23t]. Sociophysics [Sta03]. Socket [SG10]. Soetaert [Lud13]. Soft [Day11d, Zhu02]. Soft-Devices [Zhu02]. Software [ARAG19, ABC^{+}14, ACG^{+}20, Ano14y, Ano14-49, Ano15b, Ano20-36, Ano21-48, Ano22-75, ACF18, BLO^{+}22, Bli02, BBM^{+}15, BKS15, Car09a, Car09b, Car12, CHHB13, CE14, Car16, CHC17, CGK^{+}18, CENT22a, CENT22b, CF99a, CC03, Chr99, Cos22, CMK22, CGZ20, CHH^{+}13, Day08c, DD23, Don99, Dru20, Dub21, EJ09, Edd09, EWN^{+}13, ERS^{+}03, FLV^{+}09, FIG^{+}23, Fox04b, GEH^{+}99, GAB^{+}22, Gor06b, Gor08c, Gra08b, Gro09, Guo12, HAB17, HLS^{+}16, HHP19, HCK22, Her22, Hin13c, Hin15a, Hin15c, Hin19, Hoe10, HBG^{+}20, HPC20, JH18, KWBB22, KC19, KMB^{+}19, KHC^{+}20, KHS09, KSM11, KVP^{+}16, KVP^{+}17, KB09, KSM^{+}17, Kni05, KS20, LFK^{+}19, LSN20, LSPN21, Lof22, MTBG^{+}22, MMP^{+}22, MOBD^{+}22, MBH14, MWC^{+}16, MHDM99, MBB^{+}09, MMR22, O’L06a, PSSP15, Peo20, PAN^{+}16b, PK18, PBD^{+}11, PMW20, RBC^{+}19, Reb99, Sch23a, SRKS22, STWK15]. Software [Shi02b, SAK^{+}13, SZM^{+}13, SHPL12, SSW21, STH22, SGRK^{+}18, TTT15, Tho12, TJCC20, WDC18, WW17, Wil06, WL09, KNS18, CHH^{+}13, DVR^{+}19, Her21]. Sol [dA03]. Sol-Gel [dA03]. Solar [BMCC21, Col18, MRNT17, MFD^{+}09]. Solar-Powered [Col18]. Solid [Ara99, HD00, JCPS14, Run05, WWJH20]. Solids [RcK99]. Solution [Eis17, GPZ^{+}04, Moy06, O’L06e, O’L07b, STTV05, TS02, WQLZ18]. Solution-Adaptive [GPZ^{+}04]. Solutions [BT10b, Bur18, JWEK06, Lud13]. Solve [DAEJ18, MSL^{+}07, WB03]. Solved [Sul10c]. Solver [DGK16, RSC^{+}14]. Solvers [Ara99, BFL^{+}22, O’L05f, Wei21]. Solving [ATRA00, Bar21b, Bet99, CLC03, CAS^{+}07, CG09, DM04, GPC08, JCPS14, MHK^{+}06, Naj08, O’L05c, Pes03, RK05, RLRML04a, SH10, SBB^{+}15]. Some [Kul07, RLRML04b, XBK10]. Something [Cho08e, GM06]. Sonification [KWT99]. Soon [CW05e, Gor08a]. Sophisticated [Bas14]. Sort [OOB17]. Sorting [ALH^{+}20]. Soulmate [Day17c]. Sound [Azo06, KWT99, LPV00]. Sound-Wave [LPV00]. Source [ABC^{+}14, BCB07, CC03, CBB06, CGZ20, KNS18, LFN^{+}11, MMP^{+}22, MCGA22, Owe01, PGC21, WCHRM21]. Space [AAB^{+}13, Ano18j, Chr15, EAF^{+}23, GPZ^{+}04, LMPV13, Par12, Sin18, WCHRM21]. Space-Based [WCHRM21]. Space-Time [AAB^{+}13, Ano18j, Sin18]. Spaced [LM07a]. Spaces [DAEJ18, JS99]. SPACSSIM [HAB17]. Spanish [MGFRL^{+}12, RLHGA^{+}13]. Sparse [O’L05c]. Spatial [GW15, Luo13, Sch21, WLL^{+}14]. Spatiotemporal [MRHP23, SL03]. Spawn [Gor07d]. Speaking [Sul07a]. Special [Ano05b, Ano09b, Ano10c, Ano11d, Ano15e, Ano15f, CENT22a, CENT22b, Cho06f, FHM99, Got06, MW11b, Ron14]. Special-Purpose [FHM99, Got06]. Special-Relativistic [MW11b]. Speciation [dOMdO^{+}04]. Specific [Hin18c, JWLG14]. Specification [BHC^{+}08]. Specifications [Hin15c]. Spector [Mol12]. Spectra [Gaa03]. Spectral [Cor07, IHL^{+}02, Ome06, RK05, Rei21a, Rei21b, RD05a, RD05b, SM17, SOV^{+}13]. Spectrum [Cho06g, EWN^{+}13]. Speed [GYJL20, LQZL19, SR12, YYG^{+}19]. Speeds [Che03]. Speedup [Zhu16]. Sphere [AK04]. Spherical [LPV00, RRH^{+}02]. Spheroidal [Tho99b]. Spice [Vla12]. Spike [XDK^{+}20]. Spiking [JLP^{+}10]. Spin [Adl03, DCC10, KHC^{+}07, TL08b]. Spin-Polarized [KHC^{+}07]. SpiNNaker [JLP^{+}10]. Spins [Shi02a]. Spintronics [KHC^{+}07]. Splines [FGR^{+}07]. Splitting [KKO^{+}20]. SPMD [BST^{+}13]. SPMD-Like [BST^{+}13]. Sponsors [Ano20-71]. Spotlight [FL99, For99]. Spray [MPR18]. Spray-Wall [MPR18]. Spread [Cho06g]. Spring [JCC^{+}10]. Sprinting [Dub08c]. SQL [HHR^{+}13, Joh09, RTSS14a]. sqlLoader [STG08]. Squares [Rus01a, Rus02]. Stability [YYG^{+}19]. Stable [ZJW08]. Stack [AE22, HHP19]. Stacked [PSS20]. Stage [GAB^{+}22]. Staged [Hin20b]. Stages [HC17]. Staking [Cho05e]. Stampede [Dub15b]. Standard [SGS10, SGA^{+}22, Wol21]. Standards [Ano15g, Ano21-46, Day11f]. Stanford [RF00]. Stars [Ano14-44, Ano14-45, Ano14-46, Ano14-47, Ano14-48, Ano15-43, Ano15-40, Ano15-41, Ano15-42, Ano15-38, Ano15-39, Ano16z, Ano16-27, Ano16-38, Ano16-37, Ano16-39, Ano16-40, Gor08c, Gre07, ZS23, Den16]. Start [FAFX20]. Started [Gra09, KSB07]. Starving [Alt10]. State [Bal15, CCJ04, Hin12b, KAČ^{+}22, Lan04, LC09, Moy06, SSP06, YYL^{+}18, ZZC^{+}19]. State-of-Charge [YYL^{+}18]. Statements [KS20]. States [CENT22a, CENT22b, KLQ19]. Stationary [Moy06]. Statistical [CSS00, Gut01, MRHP23]. Statistics [BCC^{+}99, Hah04, Pie04, RT12, SSG16, WCC^{+}02]. Status [DF21]. Stay [Ano15-45, Ano18-45, Ano19-39, Ano19-40, Ano20-64]. STC [Ano13j, Ano14-31]. Steered [SPJ^{+}14]. Stella [Cho04]. Stellar [MSS09, WHW18]. STEM [DVS22, DL23, Jef21]. Stencil [Wei21]. Stenosis [DKCL14]. Stepping [Hin18b]. Stereocilia [YZC^{+}13]. Stieltjes [Rei21a, Rei21b]. Stiffness [YLZ^{+}19]. Still [WBP^{+}19]. Stitching [LHGX18].

Stochastic [LWSK07, PSA14, SETK05, She07, TX07, ZS07, ZZS^{+}19, ZLTX19]. Stochastics [LH06]. Stock [Bea00, MGFRL^{+}12]. Stokes [OL06h]. Stop [Wit21]. Storage [Hin20a, SACR21, YYL^{+}18, ZMM03]. Stories [Day10c]. Story [BWC01]. Storytelling [GP21, SNCM16]. Strain [PZJS10]. Strand [DGK16]. Strap [JXY^{+}19]. Strategic [Ano18-41, ES18]. Strategies [Ano16-47, Ano16-48, NSLD99, PMW20, RBK02, SH10, SWPB00, Smi16, SLM12, Sto12, WWJH20]. Strategy [CXC^{+}20, Hin18a]. Streak [WT12]. Stream [BCH^{+}09]. Streaming [VCGS11]. Streaming-Enabled [VCGS11]. Streamlines [LM07a]. Streamlining [BW14]. Stress [GKG^{+}15, O’L04a, PSR^{+}00, WLCD01]. Stress-Mediated [WLCD01]. Strikes [Cho07b]. String [AK04, Gio02]. Stroll [Läu08]. Structural [STWK15]. Structure [BJ02, BHKW03, CDKF15, GBDW04, GS03, GMPR11, Kyr08, LCY^{+}04, LHGX18, Luo13, YCZ07, vdWCV11]. Structured [DBB^{+}21, MOBD^{+}22, TKM^{+}18, YBBP15]. Structures [FL05, Maj03]. Student [Ano17-31, Ano18-46, Ano21-74, HPMJ12, HC17, KL07, Wri16]. Students [BW10, Bei12a, CHM^{+}20b, CHM^{+}20a, Den16, GPL09, GN08, GL20, Hig04, SDS00, WCP17]. Studies [AHL^{+}11, Joh12, LC09, PHW^{+}21, SK01]. Studio [Kra03]. Study [AAH^{+}08, BBW^{+}20, BDCT05, CDL^{+}23, COS^{+}15, DPBS16, DM12, DL23, DDV^{+}08, FAFX20, HF04, JH16, KHG^{+}23, KMSH10, KPM10, MB11, MR13, MM14, Mem16, MSS09, NCM^{+}14, NC18, PSSP15, RPEB14, SDA20, SM17, Sch18, Sch20, Sch21, Sch23b, SLK^{+}20, SNCT13, Sim13, SV14, Ste02, WHH22, Wol16, YWMM04, YYG^{+}19, ZLTX19]. Studying [Ma16, MAFM21, OR12, RCD^{+}00, Zeb00].

Style [Mol12, Rei13]. Stylized [LYC07]. Subarctic [EKLY07]. Subglacial [SMF^{+}23]. Subjects [TGP^{+}06]. Submissions [Ano20-56]. Submodels [MPR18]. Subspace [Saa22, vdV00]. Subsurface [Ged16a]. Subtract [Tho99a]. Success [COS^{+}15, ERF21]. Successfully [Gar17]. Such [Guo23]. Sue [Dub06b]. Suitability [GYF^{+}10, HRRS09]. Summary [KL15, MJM^{+}06]. Summer [Day17c, TL08a]. Summit [Hin18b]. Sun [GPZ^{+}04, Tou00]. Sun-to-Earth [GPZ^{+}04]. SunRay [TS02]. Sunshine [Thi15b]. Supercomputer [Fei05, Rag06, WS99, Dub15b]. Supercomputers [Ano18j, Day12f, GIF^{+}12, Sin18]. Supercomputing [ACKW01, AGC^{+}16, BLP^{+}21, CE18, GLS11, GK18, Jon19, Mil17, TC21, VSB^{+}21, WG15]. Supercomputing-Enabled [GK18]. Superconducting [DLW^{+}19]. Supermodels [Rad21]. Supernova [Ott16]. Supernovae [OR12, Tow09]. Supply [She07]. Supply-Chain [She07]. Supplying [EDJ^{+}10]. Support [CHB19, GP15, GPMSC20, Mas06, MBH14, MSD10, SPB^{+}20, TGU21, WPM^{+}12]. Supporting [GSK^{+}23, HLS^{+}16, KHE13, LZZ17, LGW^{+}17]. Supports [SYM^{+}21]. Sure [Sul08c]. Surface [CS01a, Gaa03, GGJD23, KS01, MPR18, Pey11c, QPCJ07]. Surfaces [JCC^{+}10, LJWC06, YaL10]. Surgery [DBJ^{+}20, JT01]. Surprises [Sul01b]. Surveillance [Day13d]. Survey [Ano00b, CHHB13, CW06, FKSS08, GZC14, PMP21, VWL^{+}11, Nei08, Sza99, Tha08b]. Surveys [HHF^{+}14]. Survivability [SKC02]. Survivability-Lethality [SKC02]. Survival [AT06]. Susceptible [Sch21]. Susceptible-Infected-Recovered [Sch21]. Suspension [Sul01d]. Sustainability [CGK^{+}18, CHH^{+}13, FIG^{+}23, RBC^{+}19]. Sustainable [Ano16-30, Ano18-28, Ano20m, Ano20k, Ano20l, Ano21s, Ano21t, Ano21u, Ano22r, Ano22s, Ano22t, Ano23k, Tow18]. Sustained [BPW^{+}20, CJL^{+}18, KMB^{+}19]. Sustaining [Wes03]. SV [HLYQ19]. SV-FVDM [HLYQ19]. SVD [WLL^{+}14]. SVG [LVWK02]. SVP [Tou02b]. SVP-6000 [Tou02b]. SWARM [vGDS18]. Swatch [Kil99, NLV99]. Swiftly [Dub15a]. SWIG [Cot03]. Swimming [CFCD04]. Swinging [OS03]. Switch [Kil99, NLV99]. Sylvester [O’L05f]. Symbolic [DM04, RT12, VGM^{+}09]. SymPy [RT12]. Synchronous [Can99]. Synthesis [AMKL04, GEH^{+}99, LNI^{+}19, VMK20]. Synthesis-Based [VMK20]. Synthetic [HLYQ19, UGV11]. System [BCC^{+}09, DLW^{+}19, DAKM16, DC04, DDV^{+}08, GHT^{+}10, GVB15, Gra08b, HDB^{+}04, Hin17a, Ikk16, JLYL19, KNKP14, KB04, KS00, LWF10, Los03, MS22, MRKK17, MKM^{+}14, O’L07a, O’L07b, OL06h, PG07, RPBE12, Run00, SSG16, SLK^{+}20, Sor19, STM99, SBZB13, TKM^{+}18, TJ14, UGV11, WNDV21, ZLTX19, TSFG08]. Systematic [HMB^{+}14]. Systems [AMKL04, BKRT21, BFS04, Che10, CJTH^{+}13, DMA^{+}21, FAFX20, FGRS17, Gro09, HAB17, HLT09, JR10, KC09a, KC09b, KC09c, Kwa17, Kyr08, MW14, Muz19, MS07, MGS07, NW15, O’L05c, O’L06c, OSM^{+}19, Owe01, Par00b, PSA14, RLRML04a, SDA20, SU22, STB03, SGS10, TB11, TX07, TL08b, TBM^{+}19, VM15, VWP12, WHM^{+}02, YB12, Wil01].

Table [Ano13q, Ano13r, Ano14-50, Ano14-51, Ano14-52, Ano14-53, Ano14-54, Ano15-46, Ano15-47, Ano16-41, Ano16-42, Ano16-43, Ano16-44, Ano16-45, Ano16-46, Ano17-37, Ano17-32, Ano17-33, Ano17-34, Ano17-35, Ano17-36, Ano18-47, Ano18-48, Ano18-49, Ano18-50, Ano18-51, Ano18-52, Ano19-41, Ano19-42, Ano19-43, Ano19-44, Ano19-45, Ano19-46, Ano20-70, Ano20-65, Ano20-66, Ano20-67, Ano20-68, Ano20-69, Ano21-75, Ano21-76, Ano21-77, Ano21-78, Ano21-79, Ano21-80, Ano22-69, Ano22-70, Ano22-71, Ano22-72, Ano22-73, Ano22-74, Ano23-51, Ano23-52, Ano23-53, Ano23-54, KEF07, Pos16]. Tables [Bez08, MPP14]. Tablet [Day16c]. Tablets [Thi13d]. Tackle [DMXR^{+}14]. Tackling [Akl18]. Tagged [HMA00]. Take [Ano14-55, Ano14-56, Ano15-48, Ano15-49, Ano16-47, Ano16-48, Bei09b, Jef22]. Takes [FGP99, Jon15]. Taking [O’L05c]. Tale [Ste12]. Talk [SACR21, WBP^{+}19]. TAM [DYY^{+}17]. Tangled [Sul02d]. Tank [DDV^{+}08]. Tape [CW05d]. Tapia [BTL19]. Target [HEH^{+}10]. Targeted [SNCM16]. Task [DMA^{+}21]. Task-Based [DMA^{+}21]. Tasks [BBN03, FKSS08]. Taste [MMTD^{+}17]. TAU [MS22]. Taxonomy [GB20]. Teacher [NG20]. Teachers [GPMSC20]. Teaching [AAGH17a, AAGH17b, Bas14, GL08, JHJ01, Kin16, Lew00c, Pat02]. Team [APC^{+}19, BLO^{+}22, BLT^{+}23]. Teams [HCK22, Lof22, MOBD^{+}22]. Tears [PCBVS19]. Tech [Ano19-31, Ano19-32, WMB20]. Techies [Lew00c]. TechIngnite [Ano17-38, Ano17-39]. Technical [DSPY05, Hin15b, HSP^{+}23, PMK^{+}08, Wri10]. Technique [CCSS08, DTA21, Hin20b, LCC^{+}19]. Techniques [FAFX20, Hoe10, KLMS99, LWT^{+}13, Ma16, Muh23, NSP12, STWK15, Wan18, ZYKG04]. Technologies [Gra08a, SKC02, VP04]. Technology [Ano18z, Ano19y, Ano20-72, BPW^{+}20, BCH^{+}09, BCC^{+}99, Bur99, CC99, For99, Fox02a, Goo17, IKMK13, Kal99, LL18, Nob00b, Sha99, Shi99, Shi00c, Shi00b, Shi00d, Shi00a, Shi01a, TW17, TMC^{+}13, Tou00, Tou01, YMLJ06, YLZ17]. Tectonics [MGZ00]. Teleology [BC02]. Telescope [Chr15, PRCL^{+}22, Shn06]. Telescopes [HJLH03]. Tell [Sul00b]. Temperature [DCC10, DLW^{+}19]. Temperatures [TR08]. Template [SL99]. Templates [HC99, Kir03]. Temporal [DZW^{+}05, DB21]. TEMS [Ano18z, Ano19y]. Ten [Cho08d, Dub99, RTSS14a, RTSS14b]. Tennessee [Par12]. Tensile [JXY^{+}19]. TeraGrid [DKK05]. Terahertz [GBPR11]. Term [AISR^{+}21, HS12, KILZ13, RBC^{+}19, TNR^{+}23]. Terms [Nan11, Smi99c]. Terra [Got01]. Terrain [HEH^{+}10]. Terrain-Related [HEH^{+}10]. Terrific [Got01]. Terrorism [Bor02]. Test [AGM^{+}00, NCM^{+}14, NC18, RPBE12, STHR12]. Test-Driven [NCM^{+}14, NC18]. Testing [Clo15, DLLZ19, DLLZ20, Dub12, Edd09, Hin15a, HK09, KC19, LSN20, PD02, RPBE12, Rus01b, TLG06, WM00]. Teuscher [Lov04]. Texas [PHW^{+}21]. Texascale [CED^{+}21]. Text [Aya07, Aya14, BCG^{+}99, KHE13, KNV03]. Textbook [GL08]. Textbooks [BP10]. Textiles [NG20]. Texture [FOdLVF^{+}11, NW13]. Texture-Based [FOdLVF^{+}11, NW13]. Thank [Ano20-71]. Thanks [Ano05b, Ano09b, Ano10c, Ano11d, Ano15a, Ano16a, Ano18a, Cho06f]. Their [KHG^{+}23, KLS01, Mem16, MMR22, PI16, RLRML04a, ZLW^{+}19]. Them [Wil06]. Thematic [AAB^{+}13]. Theme [Sny13]. Theme-Based [Sny13]. Theoretic [KNV03, TDT^{+}22, ZS07]. Theorist [Cre99]. Theory [Ara99, HHZK01b, HHZK01a, Jq19, JG13, KNG10, KS01, PMW20, Sch01, ZZS^{+}19]. Therapeutics [VSB^{+}21]. Therapy [Lew99b, SG00, ZFS12]. There [Esq11]. Thermal [JC02, PZJS10]. Thermath [MAC08]. Thermodynamic [Hus22]. Thermodynamics [SR13]. Thermomechanical [CBS14]. Thermonuclear [RCD^{+}00, Tow09]. Theta [Wil18]. Things [Bet17, Dub07b, FGRS17, HGA23, Smi00d, Sul10b, CBF^{+}22]. Think [Kus07, MB17]. Thinker [Lov04]. Thinking [Day11b, GP21, PKST08c, PQQ20, Thi09a, Yas17a]. Third [Smi00c, dSRT16]. Those [Wil08]. Thought [Sul05b]. Thousand [Sku04]. Threads [Sul01b]. Threat [Bor02]. Threats [TMMB18]. Three [BFF12, Day11b, DG12, DDV^{+}08, GWMG04, Maj03, PHW^{+}21, Sil02]. Three-Dimensional [GWMG04, Maj03]. Three-Tank [DDV^{+}08]. Throughput [MWC^{+}16]. Thunderstorms [Orf21]. Tightly [GIF^{+}12]. Time [AM20, AAB^{+}13, Ano18j, BPMKC21, BKS15, CYW01, CC99, Day15b, DLLZ19, DLLZ20, Don02, DTA21, Eis17, Gor06a, Gyu99, JSNR11, Kil99, LCY08, Liu15, Ma03, MHDM99, NLV99, Oug03, PSA14, SACR21, Sin18, Sul08c, TDT^{+}22, VWL^{+}11, WHM^{+}02, WT12, WWJH20, YZZ04, YHWY05, VSG^{+}02]. Time- [Eis17]. Time-Domain [Oug03]. Time-Integrated [WWJH20]. Time-Series [Liu15]. Time-Varying [DLLZ19, DLLZ20, DTA21, Ma03, TDT^{+}22]. Times [Leu20, SBH^{+}00]. Tip [YWMM04]. Tip-Clearance [YWMM04]. Titan [Bet17, Jon15]. Title [Ano15-35, Ano15-36]. TMSCS [Ano18-53]. Today [Ano19-28, Ano19-38, Ano23-42, Bar21b, Bei12a, FM02, HG02, Ano23o]. Tolerant [LCC^{+}19]. Tomography [FL05, HBB08, Sch09, SGA03]. Tomorrow [CS01b, HG02]. Tongues [GGJD23]. Too [BOS07, SRKS22]. Tool [BBW^{+}20, Guo12, LMC20, MM16, PGF^{+}15, Tan21, TL08b]. Toolbox [CC99, MSR^{+}16, Thi07]. Toolkit [SZM^{+}13, War18, VCGS11]. Tools [BLT^{+}23, CF99a, Don03, Guo23, HLT09, LeV09, MS22, PAN^{+}16a, PNL^{+}16, Rag06, Sil02, SÖS^{+}00, SLM12, Sto12, Tof08, Wri10]. Top [Ano20-72, Bei11c, DS00, MBS^{+}00]. Top500 [Fei05]. Topic [Ano14t, TM15]. Topography [SMF^{+}23]. Topologies [AK04]. Topology [GNB^{+}09]. Topology-Based [GNB^{+}09]. TOPSIS [Jq19]. Torque [CXC^{+}20]. Torsion [O’L04a]. Total [Sch09]. Totally [Lew99b, Smi00d]. Touching [EPHY18]. Tour [SLK^{+}20]. Tours [Pey11a, Pey11b, Pey11c]. Tower [Hin15a, Cho09b]. Towers [Far99]. Toy [DiP18b]. TRACE [SGRK^{+}18]. Tracing [SWDR22, ZSM^{+}22]. Track [HSJ^{+}19, Lew01b, YLZ^{+}19]. Tracked [AMKL04]. Tracking [BAD^{+}21, DTA21, JD03, KBPW15, RTSS14a, SSCN11, TJCC20]. Trade [PKST08c]. Trade-offs [PKST08c]. Traditional [Mus20]. Traffic [CS18, CSS00, HLYQ19, HT99, LMC20, SSCN11, SU22]. Trailblazing [HW09]. Train [GYJL20, LQZL19, MPR18]. Trainer [Don10]. Training [Chu21, CMK22, JPMG08, MB20a, Tha14]. Trains [RLHGA^{+}13, YYG^{+}19]. Trajectories [BH02, WB03]. Trajectory [Don02, EVL^{+}17]. Transactions [Ano16-28, Ano16-29, Ano16-30, Ano18-27, Ano18-28, Ano20u, Ano20v, Ano20j, Ano20m, Ano20k, Ano20l, Ano21-32, Ano21-33, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano22q, Ano22r, Ano22s, Ano22t, Ano23j, Ano23k, Ano21x]. Transcendental [Tho00]. Transductive [ZWYL20]. Transfer [BHC^{+}08, KEF07, PSR^{+}00, ZWYL20]. Transfer-Function [BHC^{+}08]. Transform [CS18, Cor07, Mus20, Tre99, DR05c, DR05b, DR05a, Don06a, Don06b, RD05a, RD05b]. Transformation [Par22]. Transformations [Sah03, ZZPC06]. Transformer [Jq19]. Transforming [PMK^{+}08, SDD^{+}08]. Transforms [NA07]. Transgender [Wri16]. Transient [WCC^{+}19]. Transistor [Key05, The03]. Transition [CXC^{+}20, dA03]. Transitions [BMS99, Lan99]. Translating [MS22]. Translation [EWN^{+}13]. Translational [BCE^{+}22, BLT^{+}23, CDL^{+}23, PA21, TNR^{+}23, WHH22]. Translationists [PA22]. Translator [UZC^{+}12]. Transonic [GvdWT07]. Transparency [Bar21a]. Transparent [Che17b, DYY^{+}17, GHKZ17, LZZ17, Moy06, PLW17, YLZ17, ZGR^{+}17]. Transparently [KRR^{+}12]. Transport [KHC^{+}07, Mal07a, Mal07b, PZJS10, SWDR22, TAM^{+}14]. Travel [Day18e]. Treading [For02]. Treating [Tow09]. Treatment [VWL^{+}11]. Treatments [LSDP^{+}04]. Treats [Dau99]. Tree [Gor05a, GYF^{+}10]. Trees [DBCN03]. Trenches [PA22, PMK^{+}08]. Trends [Ano20-72, KT11, KS13, Lew99a, Lew99b, Sch99, VB22]. Triangles [DSC^{+}09]. Triangulation [Bei02]. Tridiagonal [BFL^{+}22]. Trinity [NRG^{+}17]. Tropical [DRA11, STG11, SNTL13, SNCT13]. Truels [AT06]. Trusses [O’L12]. Trust [BFS04, Sul02e]. Trustworthy [Bar21a]. Try [Bei11b]. Trying [DC04]. Tsunami [EI11, LL11]. TSUSC [Ano18-54]. Tumor [LWT^{+}13, VWL^{+}11]. Turbomachines [YWMM04]. Turbulence [KBLE15, NTW07, PR01, RF00, SJDV09, TWE14]. Turbulent [AMCH07, CCPS12, DJS13, HF04, LUMM14, MP09]. Turing [Lov04, Hin17b, Lov04]. Turn [Day06c]. Turned [Sch23a]. Turning [Cho08e, DB07, LDAS19]. Tutorial [JG13, Wan23]. Tweets [AAB^{+}13]. Twice [RRAB06]. Twiddling [Coh09]. Twins [CSR^{+}23]. Twist [O’L04a]. Twitter [FL21]. Two [BOS07, BW01, DAEJ18, GWMG04, Has12, MVUSK14, Pat02, Ron14, Ste12, Tou00, TCCC13]. Two- [GWMG04]. Two-Body [Ron14]. Two-Level [DAEJ18]. Typed [LT09]. Types [PMM^{+}08].

U.K. [BBWW22]. U.S. [HP20, MBB^{+}22]. Ubiquitous [ARO^{+}11]. UK [MWE08]. Ultimate [Smi99e, Smi00d]. Ultrafast [Oug03]. Ultrahigh [WSY^{+}22]. Ultrahigh-Resolution [WSY^{+}22].

UltraLight [NCB^{+}05]. Ultrashort [TDB09]. Ultrasonic [DLLZ19, DLLZ20]. UML [RGD13]. UML/MARTE [RGD13]. Uncertainties [Rus01b]. Uncertainty [AH07, XKG05, ZLW^{+}19]. Uncle [Day09a]. Uncovering [Har18]. Undergoing [ZJW08]. Undergraduate [Don03, DL23, Ful06, GCV08, Mar17, Pes03, TK06]. Undergraduates [Lan04, Tur14b]. Underrepresented [Den16]. Understanding [BNNM04, Che18, Com20, Cyb02, EJ09, GPC08, GvdWT07, HS12, MRNT17, Ott16, WMB20]. Unearthing [Mor15]. Uneven [Mar17]. Unexpected [PK18]. Unified [GRS08, HRRS09]. Unifying [Has12, KMM^{+}11]. Union [FM02]. Unique [GL08]. Unit [TLG06, YYL^{+}18]. United [CENT22a, CENT22b, KLQ19]. Units [CRDO16, UM08, WOAEAG10, Zhu16]. Units-of-Measure [CRDO16]. universal [GGD^{+}05, BO03]. Universe [ACS15, Bry99, Jon15, MKM^{+}14, PHL^{+}10, WS99, Ged16b]. University [Cos22, DSPY05, Lan04, LC09, Par12]. Unleashed [MGS07]. Unlimited [Ano17-29, Ano17-30, Ano18-42]. Unresolvable [Tow09]. Unsteadiness [MM04]. Unsteady [GvdWT07]. Unstructured [MOBD^{+}22]. Until [Sul10c]. Up-Winded [WWJH20]. Update [DF21, SEPC10, Tho12, WLL^{+}14]. Updating [O’L06e, O’L06g]. Upgrades [PCBVS19]. Upload [Ano19-28, Ano23o]. Upward [Cho04, Sul01a]. Urban [CBF^{+}22]. URSSI [CGK^{+}18]. USA [PHW^{+}21]. Usage [KS20, MSR15, RTSS14a, SdPV21, Mol12]. Use [BAD^{+}21, Don03, GLTZ10, HBG^{+}20, Mil21, Poi10, Roc00]. Used [Sul04f]. Useful [Cho08e, LGJ^{+}19, TLG06, Thi11b]. Useless [Smi00d]. User [LSPN21, NSP12, SPW^{+}13, SPJ^{+}14, SGRK^{+}18]. User-Centered [NSP12, SGRK^{+}18]. User-Steered [SPJ^{+}14]. Uses [SSG16]. Usher [LL13]. Using [APS10, Ano18j, ALM19, Azo06, BST^{+}13, BBN03, BCH^{+}09, BTK^{+}21, BT10b, BBM^{+}21, CS18, CS01a, Cot03, DL23, DKCL14, Eis17, FOdLVF^{+}11, FAFX20, FKB^{+}13, Gal11, GEH^{+}99, GW15, GYF^{+}10, Guo23, HMA00, HRWS06, HHR02, HH99, HWPS16, HKW03, Hin20b, HSJ^{+}19, IHL^{+}02, JHJ01, JCC^{+}10, JJZC10, JG12, JS99, JON^{+}21, KFS18, KSP12, KEF07, KSSF11, LVLA14, LWG19, LPV00, MAC08, MPP14, MSL^{+}07, Mor15, Moy06, MFD^{+}09, Naj08, Ono01, PE09, Rao16, RPBE12, RLRML04a, SMM^{+}11, SU22, SSC18, SYP08, Sin18, Smi03, SJDV09, Taj10, TK06, Tre99, Vir16, Vla12, Vor01b, WCHRM21, WLCJ12, WPM^{+}12, WCAL14, WNZ^{+}17, WD06, WOAEAG10, WT12, WWJH20, WB03, XHL^{+}13, XXK^{+}02, YCKK03, YWMM04, Zak18, Zei17, ZZPC06, ZL09, SOV^{+}13]. Utility [Cho06e]. UV [SPW^{+}13]. UV-CDAT [SPW^{+}13].

V [Azo06, Don06a, Tou02a]. Vacation [TL08a]. Vaccines [WZS^{+}10]. Vale [Cho08a]. Validatable [Roa04]. Validating [Ben04, CDF^{+}04, HLS^{+}16, Pie04]. Validation [KVP^{+}17, Ste02, TP04]. valuable [O’L05a]. Value [Nob00a]. Vanícek [Ome06, Cor07]. Variability [MRHP23, ZB04]. Variable [O’L12, Rus03]. Variable-Geometry [O’L12]. Variations [GLS07]. Varying [DLLZ19, DLLZ20, DTA21, Ma03, TDT^{+}22]. Vaults [IKMK13]. Vector [ALM19, DSK15]. Vegetation [EKLY07]. Vehicle [AMKL04, DLW^{+}19, LGW^{+}17, SSCN11, ZZS^{+}19, ZLTX19]. Vehicle/Bridge [DLW^{+}19]. Vehicles [MSB^{+}14, PAN^{+}16b]. Vehicular [CSS00, YAA^{+}00]. Velo [GSB^{+}12]. Velocity [HLYQ19]. Ventilator [KKO^{+}20]. Verifiable [DG12, Roa04]. Verification [KVP^{+}17, KNKP14, TP04]. Verify [Sul02e]. Versatile [Aya07, Shi01b].

Version [HLT09, PCBVS19]. VersionClimber [PCBVS19]. Versus [Hin18d, TLD02]. Vertical [Tur15]. Very [DSK15]. Very-Wide [DSK15]. Vessels [Luo13, PAN^{+}16b]. Vestibular [ZDW^{+}07]. Vets [Day09d]. VI [Don06b]. Via [CAS^{+}07, GDH^{+}23, Boe00, Hus22, LHGX18, NCM^{+}14, RR21, WLL^{+}14, YLZ17]. Vibrating [FGP99, Gio02]. Video [BB20, Mal00, SSCN11]. Video-Based [SSCN11]. Videos [Day19b]. View [Fox18, GVB15, PMK^{+}08]. Viewing [YCKK03]. Viewpoint [OS04]. Vintage [Day10a, Lew02a]. Violent [Orf21]. Viral [GARS^{+}20]. Virtual [AGM^{+}00, CCP23, DZW^{+}05, Day18e, DBJ^{+}20, DDV^{+}08, FMB^{+}07, Har04a, HD00, How12, KT08, LRRK00, LCY08, MVUSK14, MV20, NLGNJ13, PL02, Pos14, PS17, Run05, SLK^{+}20, SPB^{+}20, TW03, WSC^{+}04, XYC^{+}09, YWC02, YCL05, PSR^{+}20]. Virtualization [HSP^{+}23, THLK10]. VirtualLab [ERS^{+}03]. Virus [BO03]. VIS [CSW17]. VisDict [GDH^{+}23]. Vision [CCD^{+}22, DiP18b, Gor06c, Gor08b, HLYQ19, Oli13, Sny13, WP22, Wil01]. Vision-Guided [DiP18b]. Visions [LPB15]. VisTrails [FS12, TGEA09]. Visual [AAB^{+}13, DVP^{+}17, DH12, ERF21, EPHY18, GDH^{+}23, Joh12, KHE13, Lof03, MSL02, RSZ^{+}20, Rob04, Wan18, RSZ^{+}21]. Visualization [ASLK22, Adl03, APS10, Ber99, CF99a, CF99b, Che99, CY00, CYW01, CN03, CZ07, Com20, CW20, CSR^{+}23, DPBS16, DB21, GHT^{+}10, GWA^{+}07, HW15, HPKS04, JWEK06, JPK01, KWT99, KSW^{+}12, KBPW15, LRRK00, Lan02, LYC07, Lof03, Ma16, MW14, MRKK17, MJM^{+}06, MSR15, MW11a, MW11b, NKV99, NC03, NW13, NSLD99, PCY14, PFS21, PP20, RV11, Rob04, RR21, SÜP^{+}11, SKNV03, STG11, SYP08, SJDV09, TAF^{+}18, Toh07, Vor01a, WCC^{+}02, WY12, WJ04, Weg00, WT12, WAS^{+}12, YCK03, ZSM^{+}22, ZCXM99, VCGS11]. Visualizations [Oli13, SNTL13, SFC07]. Visualize [Ben09]. Visualizing [AMCH07, AK04, BB07, BH02, BCG^{+}99, CFA04, CHC^{+}11, JMEL08, KF03, KEF07, Luo13, Ma03, MFD^{+}09, SOV^{+}13, SGW02, Tho00, TS10, TW03, WGJ16, dSRT16]. Visually [HKW03]. VisWeek [SSW11, SEPC10]. VLAD [LPY18]. Vlog [CHM^{+}20b, CHM^{+}20a]. Volcanology [ST05]. Voltage [WCC^{+}19]. Volume [AMCH07, Ano03, Ano05a, CCSS08, Dac16, DLLZ19, DLLZ20, Ma03, SRM^{+}07, YLR02]. Volumes [BHC^{+}08, KEF07]. Volumetric [Luo12]. Volunteer [Ano12b, Ano20-51, Ano20-52, Ano20-53, Ano23-37, Ano23-38, Ano23-39, MSR15, PBSS14]. Voronoi [Raf16]. VPython [SDS00]. VR [Ano17-40]. Vulkan [RR21]. Vulnerability [Day11d].

W [Mol12]. W2CWM2C [PMFM14]. Waiting [Bei11d]. Wall [MPR18]. Wandered [Sul08b]. Want [Bei10e, Rag06]. Wanted [Day16e]. Wants [Cyb00a]. War [Ano17c]. Warehouse [SDCV10]. Warfare [EDJ^{+}10]. Warping [LHGX18]. Warriors [Day18a]. Was [Cho05d, Cho07e, DS23, Dub07d, Sul09b]. Watch [Ano16-47, Ano16-48, Ano22c]. Watchful [Bei11d]. Water [AM05, GKG^{+}15, GGJD23, LJWC06, LWF10, WSC^{+}04, YZZ04, ABC^{+}14]. Water-Diversion [LWF10]. Water-Filled [GGJD23]. Watersheds [WSC^{+}04]. Wattage [Don10]. Watts [Ano22-75]. Wave [ACF18, LL13, LPV00, LJWC06, SA08a, SA08b, ST05, Tho99b]. Wavefield [AMKL04]. WaveformECG [WGJ16]. Wavefunctions [AK04]. Wavelet [Ama00, DVP^{+}17, FM19, Sah03, Tas00]. Wavelet-Based [Ama00, DVP^{+}17]. Wavelets [HOÖ99]. Wavemulcor [FM19]. Waves [BVT^{+}21, SNTL13]. Way [Cho05f, GM06, O’L06c, Smi00c, ST99, Tou02a, Vog13]. Weakest [AT06]. Weapons [Day10a]. Wearable [NWP19, YLZ17]. Weather [DGR^{+}05, KILZ13, LMPV13, LAY04, SBW^{+}19, STG11, WBS^{+}22]. Weave [Sul02d]. Weaving [CB02]. Web [Lau05, ACKW01, Ara99, AGC^{+}16, BC02, Ben00, Bil00, BO04, Can99, Cho07d, Com99, Cyb01, Day08a, DJ02, Dra00, Fel00, Fox01, GGD^{+}05, GM02, JHJ01, KJ04, Läu06, LTG07, Mal00, McK00, PF04, Pok04, RTSS14a, Reb99, Seg99, Shi02a, Sil00, Sil02, Smi00d, Sul02d, Tho99c, VP04, VSMD^{+}09, WCC^{+}02, XLLJ04, Zak18]. Web-Based [AGC^{+}16, WCC^{+}02]. Web-Enabled [VSMD^{+}09]. web2py [Di 11]. WebGL [Wat21]. Webgraph [DLLM04]. Week [Ano20-56, Ano22z, Ano23-40, Ano23-41, Ano20-55, Ano22-60, Ano23-42]. Weight [Jq19, Rei21a, Rei21b]. Weighting [FOdLVF^{+}11]. Weights [Rei21a, Rei21b]. We’ll [Bei09b, Sul03a]. Wenchuan [FCT^{+}10]. We’re [CW06, Day10e, Sul02c, Hin20b]. Where [Bei09c, Cho08f, Haa99, Jon15, MM14, Toh08, Ano16y]. wherever [Ano14-55, Ano14-56, Ano15-48, Ano15-49]. Which [KMB^{+}19]. Whip [Sul10c]. White [Deb18, DMA^{+}21]. Whither [Day18f, Got15]. Who [Cyb00a, KHS09, Sul04e, Thi13d, Wil08]. Whole [Mye99, Roc00]. Whole-Genome [Mye99]. Whom [Lew00a, Sul07a]. Wide [DSK15]. Widening [Jef21]. Wildfire [Sch23b]. Will [Got01, Wil08]. Willing [Sul01d]. Win [PPE00]. Wind [MRNT17]. Winded [WWJH20]. Windows [Col18, YZZ04]. Wing [MM16, RSC^{+}14]. Winners [Don99]. Wireless [VVNV18]. Within [BBM^{+}21, DMA^{+}21, SDA20, GWA^{+}07, TGEA09]. Without [BUS21, MPRU23, PCBVS19, CW05d, DSC^{+}09, Par00b]. Wizardry [Shi02a].

Women [Biz16, DKWL17, Jef21, Smi16]. Wonderful [GM06]. Word [Day12e, Day13d, Sul04d]. Words [Day15d, Day19b, Sku04]. Work [AB03, Bei10d, LTNME09, Mar99a, Wei11]. Workbench [LT08, TX08]. Workbook [MMP^{+}22]. Workflow [APC^{+}19, DVR^{+}19, MB20b, OSM^{+}19, TBM^{+}19, YBD10, YEC^{+}19]. Workflow-Driven [APC^{+}19]. Workflows [BTK^{+}21, CHB19, CR15, DGJ^{+}08, HHR^{+}13, JWLG14, JLNR19, JRD^{+}13, LDAS19, WCAL14]. Workforce [Lat16, LCG^{+}20, SYM^{+}21]. Workload [WQLZ18]. Workshop [Car09a, Car09b, CHC17, HCK22]. Workshops [Her21]. World [And11, Ano16-47, Ano16-48, Ano19z, Bil00, Bro06, CC03, Cho08f, Coh09, Orf21, RRAB06, Ano17c]. Worlds [BBC^{+}11, Tou00]. Worries [Dub08a]. Worrying [Wit21]. Worthy [Ano19-38]. Would [Day10b]. Wouldn’t [Shi00b]. Write [KHS09, Wil06]. Writers [Day18d]. Writing [Bar12, Day16f, Hin15c, O’L06a, Wri10]. Wrong [FL21, Sul07b].

X10 [Taj10]. XDMoD [PGF^{+}15]. Xeon [BHC^{+}15]. XML [CBB06, Fox02c, FB04, IK05, LVWK02, SF11, TL04a, TB04, VB08]. XSEDE [Akl18, Mor15, TCD^{+}14, Gor13]. XtremeData [SDCV10].

Y1K [Smi99a]. Y3K [Smi99f]. Yaw [OS04]. Year [BLT^{+}23, Cho05d, Cho07e, Dub15b, Smi99a, Sul02c]. Years [BKK15, BHV21, Cho08d, DF21, RTSS14a, RTSS14b, Sch17b, WG15, WP22]. Yes [WBP^{+}19]. Yield [CF13]. Yields [CJ16, Gor07b]. You’re [Cho07f, NLV99].

Zero [DCC10]. Zero-Temperature [DCC10]. Zeros [Rei21a, Rei21b]. Zipping [BCL03]. Zoltan [DBH^{+}02]. Zone [GGJD23]. Zooniverse [COS^{+}15]. Zori [OASFLAB09].

References

Arcas-Abella:2016:HAQ

[AAAH^{+}16] O. Arcas-Abella, A. Armejach, T. Hayes, G. A. Malazgirt, O. Palomar, B. Salami, and N. Sonmez. Hardware acceleration for query processing: Leveraging FPGAs, CPUs, and memory. Computing in Science and Engineering, 18(1):80–87, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Andrienko:2013:TPG

[AAB^{+}13] Gennady Andrienko, Natalia Andrienko, Harald Bosch, Thomas Ertl, Georg Fuchs, Piotr Jankowski, and Dennis Thom. Thematic patterns in georeferenced tweets through space-time visual analytics. Computing in Science and Engineering, 15(3):72–82, May/ June 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Abernathey:2021:CNR

[AAB^{+}21] R. P. Abernathey, T. Augspurger, A. Banihirwe, C. C. Blackmon-Luca, T. J. Crone, C. L. Gentemann, J. J. Hamman, N. Henderson, C. Lepore, T. A. McCaie, N. H. Robinson, and R. P. Signell. Cloudnative repositories for big scientific data. Computing in Science and Engineering, 23 (2):26–35, March/April 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Alexandron:2017:TSBa

[AAGH17a] Giora Alexandron, Michal Armoni, Michal Gordon, and David Harel. Teaching scenario-based programming: An additional paradigm for the high school computer science curriculum, part 1. Computing in Science and Engineering, 19(5):58–67, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/05/mcs2017050058-
abs.html.

Alexandron:2017:TSBb

[AAGH17b] Giora Alexandron, Michal Armoni, Michal Gordon, and David Harel. Teaching scenario-based programming: An additional paradigm for the high school computer science curriculum, part 2. Computing in Science and Engineering, 19(6):64–71, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060064-
abs.html.

Anderson:2008:PCA

[AAH^{+}08] Erik W. Anderson, James P.

Ahrens, Katrin Heitmann, Salman Habib, and Cláudio T. Silva. Provenance in comparative analysis: a study in cosmology. Computing in Science and Engineering, 10(3): 30–37, May/June 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Andreas:2003:EWI

[AB03] April K. Andreas and Isabel Beichl. Estimating the work in integer partitioning. Computing in Science and Engineering, 5(1):48–56, January/February 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/01/c1048abs.htm;
http://csdl.computer.org/
dl/mags/cs/2003/01/c1048.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/01/
c1048.pdf.

Ahalt:2014:WSS

[ABC^{+}14] Stan Ahalt, Larry Band, Laura Christopherson, Ray Idaszak, Chris Lenhardt, Barbara Minsker, Margaret Palmer, Mary Shelley, Michael Tiemann, and Ann Zimmerman. Water Science Software Institute: Agile and open source scientific software development. Computing in Science and Engineering, 16(3):

18–26, May/June 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Alur:2002:MAB

[ABK^{+}02] Rajeev Alur, Calin Belta, Vijay Kumar, Max Mintz, George J. Pappas, Harvey Rubin, and Jonathan Schug. Modeling and analyzing biomolecular networks. Computing in Science and Engineering, 4(1):20–31, January/February 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c1020abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c1020.pdf.

Adams:2022:PCI

[ABM^{+}22] Mark Adams, Satish Balay, Oana Marin, Lois Curfman McInnes, Richard Tran Mills, Todd Munson, Hong Zhang, Junchao Zhang, Jed Brown, Victor Eijkhout, Jacob Faibussowitsch, Matthew Knepley, Fande Kong, Scott Kruger, Patrick Sanan, Barry F. Smith, and Hong Zhang. The PETSc community as infrastructure. Computing in Science and Engineering, 24(3): 6–15, May/June 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Almgren:2009:NLM

[ABNZ09] Ann S. Almgren, John B. Bell, Andy Nonaka, and Michael Zingale. A new low Mach number approach in astrophysics. Computing in Science and Engineering, 11(2): 24–33, March/April 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Antil:2018:NQB

[ACF18] Harbir Antil, Dangxing Chen, and Scott Field. A note on QR-based model reduction: Algorithm, software, and gravitational wave applications. Computing in Science and Engineering, 20(4): 10–25, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040010-
abs.html.

Alliez:2020:ARR

[ACG^{+}20] P. Alliez, R. D. Cosmo, B. Guedj, A. Girault, M. Hacid, A. Legrand, and N. Rougier. Attributing and referencing (research) software: Best practices and outlook from Inria. Computing in Science and Engineering, 22(1):39–52, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Aloisio:2001:WAS

[ACKW01] Giovanni Aloisio, Massimo Cafaro, Carl Kesselman, and Roy Williams. Web access to supercomputing. Computing in Science and Engineering, 3(6):66–72, November/December 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c6066abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2001/pdf/c6066.pdf.

Ali:2020:PME

[ACQ^{+}20] R. Ali, J. Chen, G. Qu, M. Pekarek, Y. Cai, F. Zhou, and M. Huang. Pain marker evaluation application in augmented reality and mobile platforms. Computing in Science and Engineering, 22(3): 40–50, May/June 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Aragon-Calvo:2015:FTU

[ACS15] Miguel A. Aragon-Calvo and Mark SubbaRao. A flight through the universe. Computing in Science and Engineering, 17(6):96–102, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Adler:2003:VAS

[Adl03]  Joan Adler.  Visualiza-

tion in atomistic and spin simulations. Computing in Science and Engineering, 5(5):61–65, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/05/c5061abs.htm;
http://csdl.computer.org/
dl/mags/cs/2003/05/c5061.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/05/
c5061.pdf.

Adler:2020:CSE

[Adl20] J. Adler. Computational science and engineering education. Computing in Science and Engineering, 22(4): 4–6, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Alarcon:2022:QCH

[AE22] Sonia Lopez Alarcon and Anne C. Elster. Quantum computing and highperformance computing: Compilation stack similarities. Computing in Science and Engineering, 24(6):66–71, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Atwood:2016:SWB

[AGC^{+}16] C. A. Atwood, R. C. Goebbert, J. A. Calahan, T. V. Hromadka, T. M. Proue, W. Monceaux, and J. Hirata. Secure Web-based access for productive supercomputing. Computing in Science and Engineering, 18(1):63–72, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Aivazis:2000:VTF

[AGM^{+}00] Michael Aivazis, William A. Goddard, Dan Meiron, Michael Ortiz, James Pool, and Joseph Shepherd. A virtual test facility for simulating the dynamic response of materials. Computing in Science and Engineering, 2(2): 42–53, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c2042.
pdf;   http://www.computer.
org/cse/cs1999/c2042abs.
htm.

Arriola:2007:BSU

[AH07] Leon M. Arriola and James M. Hyman. Being sensitive to uncertainty. Computing in Science and Engineering, 9(2): 10–20, March/April 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ahrens:2011:DIS

[AHL^{+}11] James Ahrens, Bruce Hendrickson, Gabrielle Long, Steve Miller, Rob Ross, and Dean Williams. Dataintensive science in the US DOE: Case studies and future challenges. Computing in Science and Engineering, 13(6): 14–24, November/December 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Alexander:2011:BD

[AHS11] Francis J. Alexander, Adolfy Hoisie, and Alexander Szalay. Big data. Computing in Science and Engineering, 13(6):10–13, November/December 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Akhlaghi:2021:TLT

[AISR^{+}21] Mohammad Akhlaghi, Raúl Infante-Sainz, Boudewijn F. Roukema, Mohammadreza Khellat, David Valls-Gabaud, and Roberto Baena-Gallé. Toward long-term and archivable reproducibility. Computing in Science and Engineering, 23(3):82–91, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ashkenazi:2004:SRS

[AK04] Guy Ashkenazi and Ronnie Kosloff. String, ring, sphere: Visualizing wavefunctions on different topologies. Computing in Science and Engineering, 6(3): 82–86, May/June 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/03/c3082abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2004/03/c3082.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/03/
                                        c3082.pdf.

Akli:2018:XTD

[Akl18] Linda Akli. XSEDE: Tackling diversity and inclusion in advanced computing. Computing in Science and Engineering, 20(3):71–72, May/ June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Alexander:2013:MLG

[Ale13] Francis J. Alexander. Machine learning [Guest Editor’s introduction]. Computing in Science and Engineering, 15 (5):9–11, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Alexandru:2015:LQC

[Ale15]  Andrei Alexandru.  Lat-

tice quantum chromodynamics with overlap fermions on GPUs. Computing in Science and Engineering, 17(2): 14–22, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/02/mcs2015020014-
                                        abs.html.

Alexandrou:2018:LFI

[Ale18] Constantia Alexandrou. The LinkSCEEM FP7 infrastructure project: Linking scientific computing in Europe and the Eastern Mediterranean. Computing in Science and Engineering, 20(3):13–20, May/ June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8358028/.

Albarado:2015:AAN

[ALH15] Kevin Albarado, Timothy Ledlow, and Roy Hartfield. Alternative analysis networking: A multicharacterization algorithm. Computing in Science and Engineering, 17(1):54–63, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/01/mcs2015010054-
abs.html.

Aldrich:2020:QBF

[ALH^{+}20] G. Aldrich, J. Lukasczyk, J. D. Hyman, G. Srinivasan, H. Viswanathan, C. Garth, H. Leitte, J. P. Ahrens, and B. Hamann. A query-based framework for searching, sorting, and exploring data ensembles. Computing in Science and Engineering, 22(2): 64–76, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Arnas:2019:RSU

[ALM19] D. Arnas, C. Leake, and D. Mortari. Random sampling using k-vector. Computing in Science and Engineering, 21(1):94–107, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Alted:2010:WMC

[Alt10] Francesc Alted. Why modern CPUs are starving and what can be done about it. Computing in Science and Engineering, 12(2):68–71, March/April 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Athanasiadis:2005:SIW

[AM05] Ioannis N. Athanasiadis and Pericles A. Mitkas. Social influence and water conservation: An agent-based approach. Computing in Science and Engineering, 7 (1):65–70, January/February 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/01/c1065.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2005/01/
                                        c1065.pdf.

Alexander:2015:OSL

[AM15] Francis J. Alexander and Charles Meneveau. Open simulation laboratories [Guest Editors’ introduction]. Computing in Science and Engineering, 17(5):7–9, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs2015050007.
html.

Adler:2018:CSD

[AM18] Joan Adler and Esteban Mocskos. Computational science in developing countries. Computing in Science and Engineering, 20(3):10–12, May/ June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8358038/.

Amaro:2020:BST

[AM20] R. E. Amaro and A. J. Mulholland. Biomolecular simulations in the time of COVID-19, and after. Computing in Science and Engineering, 22(6):30–36, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Amaratunga:2000:WBA

[Ama00]  Kevin Amaratunga.  A

wavelet-based approach for compressing kernel data in large-scale simulations of 3D integral problems. Computing in Science and Engineering, 2(4):34–45, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c4034.
pdf;   http://www.computer.
org/cse/cs1999/c4034abs.
htm.

Amar:2006:MCM

[Ama06] Jacques G. Amar. The Monte Carlo method in science and engineering. Computing in Science and Engineering, 8(2): 9–19, March/April 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Akiba:2007:VMV

[AMCH07] Hiroshi Akiba, Kwan-Liu Ma, Jacqueline H. Chen, and Evatt R. Hawkes. Visualizing multivariate volume data from turbulent combustion simulations. Computing in Science and Engineering, 9(2): 76–83, March/April 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anderson:2004:TVS

[AMKL04] Thomas S. Anderson, Mark L. Moran, Stephen A. Ketcham, and James Lacombe. Tracked vehicle simulations and seismic wavefield synthesis in seismic sensor systems. Computing in Science and Engineering, 6(6):22–28, November/December 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/06/c6022.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/06/
c6022.pdf.

Amolo:2018:GHP

[Amo18]  George O. Amolo.  The

growth of high-performance computing in Africa. Computing in Science and Engineering, 20(3):21–24, May/ June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8358041/.

Amundson:2014:HPC

[AMS14] James Amundson, Alexandru Macridin, and Panagiotis Spentzouris. Highperformance computing modeling advances accelerator science for high-energy physics. Computing in Science and Engineering, 16(6):32–41, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs2014060032-
abs.html.

Andreyev:2011:WMO

[And11]  Sergey Andreyev.  Into

the world of movable objects. Computing in Science and Engineering, 13(4):79– 84, July/August 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:1999:AI

[Ano99] Anonymous. 1999 annual index. Computing in Science and Engineering, 1(6): 88–91, November/December 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c6088.
                                        pdf.

Anonymous:2000:AI

[Ano00a] Anonymous. 2000 annual index. Computing in Science and Engineering, 2(6): 94–??, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c6094.
                                        pdf.

Anonymous:2000:MS

[Ano00b] Anonymous. Marketing survey. Computing in Science and Engineering, 2 (2):4–5, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c2004.
                                        pdf.

Anonymous:2001:AI

[Ano01] Anonymous. 2001 annual index. Computing in Science and Engineering, 3(6):100, c3, November/December 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c6100abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c6100.pdf.

Anonymous:2002:AI

[Ano02a]  Anonymous.  2002 an-

nual index. Computing in Science and Engineering, 4(6):83, C3, November/December 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/06/c6083.pdf;
http://csdl.computer.org/
dl/mags/cs/2002/06/c6083.
htm.

Anonymous:2002:F

[Ano02b] Anonymous. Focus. Computing in Science and Engineering, 4(4):6–11, July/August 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2002/04/c4006.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/04/
c4006.pdf.

Anonymous:2003:CSE

[Ano03] Anonymous. Computing in Science & Engineering 2003 annual index, volume 5. Computing in Science and Engineering, 5(6):88–91, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/06/c6088.pdf.

Anonymous:2004:AI

[Ano04a]  Anonymous.  2004 an-

nual index. Computing in Science and Engineering, 6(6):91–96, November/ December 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/06/c6091.pdf;
http://csdl.computer.org/
dl/mags/cs/2004/06/c6091.
htm.

Anonymous:2004:LE

[Ano04b] Anonymous. Letters to the Editors. Computing in Science and Engineering, 6(4): 3–5, July/August 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/04/c4003.pdf;
http://csdl.computer.org/
dl/mags/cs/2004/04/c4003.
htm.

Anonymous:2005:AIC

[Ano05a] Anonymous. 2005 annual index, Computing in Science and Engineering, volume 7, number 6. Computing in Science and Engineering, 7(6):91–95, November/December 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/06/c6091.pdf.

Anonymous:2005:STC

[Ano05b] Anonymous. Special thanks to CiSE’s reviewers. Computing in Science and Engineering, 7(1):21–??, January/February 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/01/c1021.pdf;
http://csdl.computer.org/
dl/mags/cs/2005/01/c1021.
htm.

Anonymous:2006:AI

[Ano06]  Anonymous.  2006 an-

nual index. Computing in Science and Engineering, 8(6):98–103, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2006/06/c6098.pdf.

Anonymous:2007:LE

[Ano07] Anonymous. Letters to the editor. Computing in Science and Engineering, 9(4): 4–6, July/August 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2007/04/c4004.pdf.

Anonymous:2008:AI

[Ano08] Anonymous. Annual index. Computing in Science and Engineering, 10(6):??, November/December 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2009:AI

[Ano09a] Anonymous. Annual index. Computing in Science and Engineering, 11(6):0, November/December 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2009:STC

[Ano09b] Anonymous. Special thanks to CiSE’s peer reviewers. Computing in Science and Engineering, 11(2):5, March/April 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2010:B

[Ano10a] Anonymous. Books. Computing in Science and Engineering, 12(4):7–11, July/August 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2010:RR

[Ano10b] Anonymous. Reproducible research. Computing in Science and Engineering, 12 (5):8–13, September/October 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2010:STC

[Ano10c] Anonymous. Special thanks to CiSE’s peer reviewers. Computing in Science and Engineering, 12(2):94–95, March/ April 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2011:Ba

[Ano11a] Anonymous. Books. Computing in Science and Engineering, 13(2):5–8, March/April 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2011:Bb

[Ano11b] Anonymous. Books. Computing in Science and Engineering, 13(6):6–9, November/December 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2011:CAI

[Ano11c] Anonymous. CiSE 2011 annual index. Computing in Science and Engineering, 13(6): i, November/December 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2011:STC

[Ano11d] Anonymous. Special thanks to CiSE’s peer reviewers. Computing in Science and Engineering, 13(1):87, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2012:B

[Ano12a] Anonymous. Books. Computing in Science and Engineering, 14(1):5–8, January/February 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2012:CSE

[Ano12b] Anonymous. Computing in Science & Engineering commends volunteer reviewers. Computing in Science and Engineering, 14(1):87, January/February 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:CP

[Ano13a] Anonymous. Call for papers. Computing in Science and Engineering, 15(5): c2, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:CNA

[Ano13b] Anonymous. Computing now [advertisement]. Computing in Science and Engineering, 15(6):7, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:CPS

[Ano13c] Anonymous. Conference publishing services [advertisement]. Computing in Science and Engineering, 15 (6):55, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:DMA

[Ano13d] Anonymous. Digital magazines [advertisement]. Computing in Science and Engineering, 15(6):63, November/December 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:EA

[Ano13e] Anonymous. ePub [advertisement]. Computing in Science and Engineering, 15 (6):c2, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:FC

[Ano13f] Anonymous. [front cover].

Computing in Science and Engineering, 15(5):c1, September/October 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:ICC

[Ano13g] Anonymous. IEEE Cloud Computing [advertisement]. Computing in Science and Engineering, 15(6):c4, November/December 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:ICSb

[Ano13h] Anonymous. IEEE Computer Society [advertisement]. Computing in Science and Engineering, 15(6):103, November/December 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:ICSa

[Ano13i] Anonymous. IEEE Computer Society information. Computing in Science and Engineering, 15(5):21, September/October 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:IS

[Ano13j] Anonymous. IEEE STC 2014. Computing in Science and Engineering, 15(6):73, November/December 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:JBA

[Ano13k] Anonymous. Jobs board [advertisement]. Computing in Science and Engineering, 15 (6):81, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:Ma

[Ano13l]  Anonymous.  [masthead].

Computing in Science and Engineering, 15(5):1, September/October 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:Mb

[Ano13m]  Anonymous.  [masthead].

Computing in Science and Engineering, 15(6):1, November/December 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:MAb

[Ano13n]  Anonymous.  Membership

[advertisement]. Computing in Science and Engineering, 15(5):c3, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:MAc

[Ano13o] Anonymous. Membership [advertisement]. Computing in Science and Engineering, 15 (6):c3, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:MAa

[Ano13p]  Anonymous.  myComputer

[advertisement]. Computing in Science and Engineering, 15(5):41, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:TCa

[Ano13q] Anonymous. Table of contents. Computing in Science and Engineering, 15 (5):2–3, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2013:TCb

[Ano13r] Anonymous. Table of contents. Computing in Science and Engineering, 15 (6):2–3, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:CPa

[Ano14a] Anonymous. Call for papers. Computing in Science and Engineering, 16(1): 87, January/February 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:CPb

[Ano14b] Anonymous. Call for papers. Computing in Science and Engineering, 16(5):c2, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs20140500c2.
pdf.

Anonymous:2014:CPc

[Ano14c] Anonymous. Call for papers.

Computing in Science and Engineering, 16(5):c4, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs20140500c4.
pdf.

Anonymous:2014:CPH

[Ano14d] Anonymous. Call for papers house advertisement. Computing in Science and Engineering, 16(6):c3, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs20140600c3.
pdf.

Anonymous:2014:DMAa

[Ano14e] Anonymous. Digital magazines [advertisement]. Computing in Science and Engineering, 16(3):67, May/June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:DMAb

[Ano14f] Anonymous. Digital magazines [advertisement]. Computing in Science and Engineering, 16(5):61, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs2014050061.
pdf.

Anonymous:2014:DMHa

[Ano14g] Anonymous. Digital magazines [house advertisement]. Computing in Science and Engineering, 16(1):51, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:DMHb

[Ano14h] Anonymous. Digital magazines house advertisement. Computing in Science and Engineering, 16(6):79, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs2014060079.
pdf.

Anonymous:2014:EHA

[Ano14i] Anonymous. ePub [house advertisement]. Computing in Science and Engineering, 16(1):69, January/February 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:FYJb

[Ano14j] Anonymous. Focus on your job search [advertisement]. Computing in Science and Engineering, 16(3):89, May/ June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:FYJc

[Ano14k] Anonymous. Focus on your job search [advertisement]. Computing in Science and Engineering, 16(5):7, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs2014050007.
pdf.

Anonymous:2014:FYJa

[Ano14l] Anonymous. Focus on your job search [house advertisement]. Computing in Science and Engineering, 16(2): 7, March/April 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:FYJd

[Ano14m]  Anonymous.  Focus on

your job search house advertisement. Computing in Science and Engineering, 16(6):c2, November/ December 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.

                                        computer.org/csdl/mags/
                                        cs/2014/06/mcs20140600c2.
                                        pdf.

Anonymous:2014:FCa

[Ano14n] Anonymous. Front cover. Computing in Science and Engineering, 16(1):c1, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:FCb

[Ano14o] Anonymous. Front cover. Computing in Science and Engineering, 16(2):c1, March/ April 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:FCc

[Ano14p] Anonymous. [front cover]. Computing in Science and Engineering, 16(3):c1, May/June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:FCd

[Ano14q] Anonymous. Front cover. Computing in Science and Engineering, 16(4):c1, July/ August 2014. CODEN CSENFA. ISSN 1521-9615.

Anonymous:2014:FCe

[Ano14r] Anonymous. [front cover]. Computing in Science and Engineering, 16(5):c1, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs20140500c1.
pdf.

Anonymous:2014:FCf

[Ano14s] Anonymous. Front cover.

Computing in Science and Engineering, 16(6):c1, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs20140600c1.
pdf.

Anonymous:2014:GHT

[Ano14t] Anonymous. Get hot topic insights from industry leaders house advertisement. Computing in Science and Engineering, 16(1):41, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:ICCb

[Ano14u] Anonymous. IEEE CG&A calls for papers. Computing in Science and Engineering, 16 (3):57, May/June 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:ICCa

[Ano14v] Anonymous. IEEE Cloud Computing [house advertisement]. Computing in Science and Engineering, 16(1): c3, January/February 2014.

CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:ICSc

[Ano14w] Anonymous. IEEE Computer Society [advertisement]. Computing in Science and Engineering, 16(3):27, May/June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:ICSf

[Ano14x] Anonymous. IEEE Computer Society [advertisement]. Computing in Science and Engineering, 16(5):53, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2014/05/mcs2014050053.
                                        pdf.

Anonymous:2014:ICSe

[Ano14y]  Anonymous.  IEEE Com-

puter Society Harlan D. Mills Award: Call for software engineering award nominations. Computing in Science and Engineering, 16(5):13, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2014/05/mcs2014050013.
                                        pdf.

Anonymous:2014:ICSa

[Ano14z]  Anonymous.  IEEE Com-

puter Society [house advertisement]. Computing in Science and Engineering, 16(1): 33, January/February 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:ICSd

[Ano14-27]  Anonymous.  IEEE Com-

puter Society [house advertisement]. Computing in Science and Engineering, 16(4): c2, July/August 2014. CO-DEN CSENFA. ISSN 1521-9615.

Anonymous:2014:ICSg

[Ano14-28] Anonymous. IEEE Computer Society house advertisement. Computing in Science and Engineering, 16(6):51, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs2014060051.
pdf.

Anonymous:2014:ICSb

[Ano14-29] Anonymous. IEEE Computer Society information [house advertisement]. Computing in Science and Engineering, 16(2):21, March/April 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:ISP

[Ano14-30]  Anonymous.  IEEE Secu-

rity&Privacy [advertisement]. Computing in Science and Engineering, 16(3):47, May/June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:ISH

[Ano14-31]  Anonymous.  IEEE STC

2014 [house advertisement]. Computing in Science and Engineering, 16(1):65, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:IA

[Ano14-32]  Anonymous.  Intelect ad-

vertisement. Computing in Science and Engineering, 16(6):c4, November/ December 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2014/06/mcs20140600c4.
                                        pdf.

Anonymous:2014:Ma

[Ano14-33]  Anonymous.  Masthead.

Computing in Science and Engineering, 16(1):1, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:Mb

[Ano14-34] Anonymous. Masthead. Computing in Science and Engineering, 16(2):1, March/April 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:Mc

[Ano14-35]  Anonymous.  [masthead].

Computing in Science and Engineering, 16(3):1, May/June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:Md

[Ano14-36] Anonymous. Masthead. Computing in Science and Engineering, 16(4):1, July/August 2014. CODEN CSENFA. ISSN 1521-9615.

Anonymous:2014:Me

[Ano14-37]  Anonymous.  [masthead].

Computing in Science and Engineering, 16(5):1, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs2014050001.
pdf.

Anonymous:2014:Mf

[Ano14-38] Anonymous. Masthead. Computing in Science and Engineering, 16(6):1, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs2014060001.
pdf.

Anonymous:2014:MMA

[Ano14-39]  Anonymous.  Membership

matters [advertisement]. Computing in Science and Engineering, 16(3):c2, May/June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:MMHa

[Ano14-40]  Anonymous.  Membership

matters [house advertisement]. Computing in Science and Engineering, 16(1): c2, January/February 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:MMHb

[Ano14-41]  Anonymous.  Membership

matters [house advertisement]. Computing in Science and Engineering, 16(2): c2, March/April 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:MMHc

[Ano14-42]  Anonymous.  Membership

matters [house advertisement]. Computing in Science and Engineering, 16(4): c3, July/August 2014. CO-DEN CSENFA. ISSN 1521-

Anonymous:2014:RMD

[Ano14-43]  Anonymous.  Richard E.

Merwin Distinguished Service Award. Computing in Science and Engineering, 16(5):43, September/ October 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs2014050043.
pdf.

Anonymous:2014:RSB

[Ano14-44]  Anonymous.  Rock stars

of big data analytics [advertisement]. Computing in Science and Engineering, 16(5):c3, September/ October 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs20140500c3.
pdf.

Anonymous:2014:RSCa

[Ano14-45] Anonymous. Rock stars of cybersecurity [advertisement]. Computing in Science and Engineering, 16(3):c3, May/June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:RSCb

[Ano14-46] Anonymous. Rock stars of cybersecurity [house advertisement]. Computing in Science and Engineering, 16(4): c4, July/August 2014. CO-DEN CSENFA. ISSN 1521-9615.

Anonymous:2014:RSMa

[Ano14-47] Anonymous. Rock stars of mobile cloud [house advertisement]. Computing in Science and Engineering, 16(1): 77, January/February 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:RSMb

[Ano14-48] Anonymous. Rock stars of mobile cloud [house advertisement]. Computing in Science and Engineering, 16(2): c4, March/April 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:SEC

[Ano14-49] Anonymous. Software engineering for the 21st Century [house advertisement]. Computing in Science and Engineering, 16(2):c3, March/ April 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:TCa

[Ano14-50] Anonymous. Table of contents. Computing in Science and Engineering, 16(1):2–3, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:TCb

[Ano14-51] Anonymous. Table of contents. Computing in Science and Engineering, 16(2): 2–3, March/April 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:TCc

[Ano14-52] Anonymous. Table of contents. Computing in Science and Engineering, 16(3): 2–3, May/June 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:TCd

[Ano14-53] Anonymous. Table of contents. Computing in Science and Engineering, 16(4): 2–3, July/August 2014. CO-DEN CSENFA. ISSN 1521-9615.

Anonymous:2014:TCe

[Ano14-54] Anonymous. Table of contents. Computing in Science and Engineering, 16 (5):2–3, September/October 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/csdl/
mags/cs/2014/05/mcs2014050002.
pdf.

Anonymous:2014:TCLa

[Ano14-55] Anonymous. Take the CS Library wherever you go! Computing in Science and Engineering, 16(3):35, May/June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2014:TCLb

[Ano14-56] Anonymous. Take the CS Library wherever you go! [advertisement]. Computing in Science and Engineering, 16(5):75, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs2014050075.
pdf.

Anonymous:2015:RT

[Ano15a] Anonymous. 2014 reviewer thanks. Computing in Science and Engineering, 17(1):68–70, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/01/mcs2015010068.
pdf.

Anonymous:2015:AIC

[Ano15b]  Anonymous.  39th An-

nual International Computers, Software & Applications Conference advertisement. Computing in Science and Engineering, 17(3): 43, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/

cs/2015/03/mcs2015030043.
pdf.

Anonymous:2015:CNH

[Ano15c] Anonymous. Call for nominees house advertisement. Computing in Science and Engineering, 17(6):c4, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2015:CP

[Ano15d] Anonymous. Call for papers.

Computing in Science and Engineering, 17(1):c3, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/01/mcs20150100c3.
pdf.

Anonymous:2015:CPSa

[Ano15e] Anonymous. Call for papers: Special issue on online behavioral analysis and modeling house advertisement. Computing in Science and Engineering, 17(3): 85, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/03/mcs2015030085.
pdf.

Anonymous:2015:CPSb

[Ano15f] Anonymous. Call for papers: Special issue on pattern recognition house advertisement. Computing in Science and Engineering, 17(3): c4, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/03/mcs20150300c4.
pdf.

Anonymous:2015:CSA

[Ano15g] Anonymous. Call for standards award nominations house advertisement. Computing in Science and Engineering, 17(6):c3, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2015:CWK

[Ano15h]  Anonymous.  The com-

pany we keep [advertisement]. Computing in Science and Engineering, 17(1):27, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/01/mcs2015010027.
pdf.

Anonymous:2015:CPY

[Ano15i] Anonymous. Conferences in the palm of your hand house advertisement. Computing in Science and Engineering, 17 (4):c3, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs20150400c3.
pdf.

Anonymous:2015:FYJa

[Ano15j] Anonymous. Focus on your job search [advertisement]. Computing in Science and Engineering, 17(1):71, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/01/mcs2015010071.
pdf.

Anonymous:2015:FYJb

[Ano15k] Anonymous. Focus on your job search house advertisement. Computing in Science and Engineering, 17(3): 77, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/03/mcs2015030077.
pdf.

Anonymous:2015:FYJc

[Ano15l] Anonymous. Focus on your job search house advertisement. Computing in Science and Engineering, 17(4): 51, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040051.
pdf.

Anonymous:2015:FCa

[Ano15m] Anonymous. [front cover]. Computing in Science and Engineering, 17(1):c1, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/01/mcs20150100c1.
                                        pdf.

Anonymous:2015:FCb

[Ano15n] Anonymous. Front cover. Computing in Science and Engineering, 17(2):c1, March/ April 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //csdl.computer.org/csdl/
                                        mags/cs/2015/02/mcs20150200c1.
                                        pdf.

Anonymous:2015:FCc

[Ano15o] Anonymous. Front cover. Computing in Science and Engineering, 17(3):c1, May/ June 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //csdl.computer.org/csdl/
                                        mags/cs/2015/03/mcs20150300c1.
                                        pdf.

Anonymous:2015:FCd

[Ano15p] Anonymous. Front cover. Computing in Science and Engineering, 17(4):c1, July/ August 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.

computer.org/csdl/mags/
cs/2015/04/mcs20150400c1.
pdf.

Anonymous:2015:FCe

[Ano15q] Anonymous. Front cover.

Computing in Science and Engineering, 17(5):c1, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs20150500c1.
pdf.

Anonymous:2015:FC

[Ano15r] Anonymous. Front cover.

Computing in Science and Engineering, 17(6):c1, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2015:GMLa

[Ano15s]  Anonymous.  Get more,

for less! House advertisement. Computing in Science and Engineering, 17(3): 53, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/03/mcs2015030053.
pdf.

Anonymous:2015:GMLb

[Ano15t]  Anonymous.  Get more,

for less! House advertisement. Computing in Science and Engineering, 17(4):

11, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040011.
pdf.

Anonymous:2015:GMLc

[Ano15u]  Anonymous.  Get more,

for less house advertisement. Computing in Science and Engineering, 17(4): c4, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs20150400c4.
pdf.

Anonymous:2015:GMLd

[Ano15v] Anonymous. Get more, for less house advertisement. Computing in Science and Engineering, 17(5):c3, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs20150500c3.
pdf.

Anonymous:2015:ICC

[Ano15w] Anonymous. IEEE Cloud Computing call for papers house advertisement. Computing in Science and Engineering, 17(2):c3, March/April 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/csdl/
mags/cs/2015/02/mcs20150200c3.
pdf.

Anonymous:2015:ICSa

[Ano15x] Anonymous. IEEE Computer Society house advertisement. Computing in Science and Engineering, 17(1):11, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/01/mcs2015010011.
pdf.

Anonymous:2015:ICSb

[Ano15y]  Anonymous.  IEEE Com-

puter Society house advertisement. Computing in Science and Engineering, 17(3): 9, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/03/mcs2015030009.
pdf.

Anonymous:2015:ICSc

[Ano15z]  Anonymous.  IEEE Com-

puter Society house advertisement. Computing in Science and Engineering, 17(4): 7, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040007.
pdf.

Anonymous:2015:ICSd

[Ano15-27] Anonymous. IEEE Computer Society house advertisement. Computing in Science and Engineering, 17(5):c4, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/05/mcs20150500c4.
                                        pdf.

Anonymous:2015:IA

[Ano15-28] Anonymous. Intelect [advertisement]. Computing in Science and Engineering, 17(1): c2, January/February 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:/

                                        /csdl.computer.org/csdl/
                                        mags/cs/2015/01/mcs20150100c2.
                                        pdf.

Anonymous:2015:Ma

[Ano15-29]  Anonymous.  Masthead.

Computing in Science and Engineering, 17(1):1, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/01/mcs2015010001.
                                        pdf.

Anonymous:2015:Mb

[Ano15-30] Anonymous. Masthead. Computing in Science and Engineering, 17(2):1, March/April 2015. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/csdl/
mags/cs/2015/02/mcs2015020001.
pdf.

Anonymous:2015:Mc

[Ano15-31] Anonymous. Masthead. Computing in Science and Engineering, 17(3):1, May/June 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/csdl/
mags/cs/2015/03/mcs2015030001.
pdf.

Anonymous:2015:Md

[Ano15-32]  Anonymous.  Masthead.

Computing in Science and Engineering, 17(4):1, July/ August 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040001.
pdf.

Anonymous:2015:Me

[Ano15-33] Anonymous. Masthead. Computing in Science and Engineering, 17(5):1, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs2015050001.
pdf.

Anonymous:2015:M

[Ano15-34] Anonymous. Masthead. Computing in Science and Engineering, 17(6):1, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2015:NTIa

[Ano15-35] Anonymous. New title from IEEE CS Press house advertisement. Computing in Science and Engineering, 17(2): c2, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/02/mcs20150200c2.
                                        pdf.

Anonymous:2015:NTIb

[Ano15-36] Anonymous. New title from IEEE CS Press house advertisement. Computing in Science and Engineering, 17(3): 86, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/03/mcs2015030086.
                                        pdf.

Anonymous:2015:PNH

[Ano15-37] Anonymous. Postdocs needed to help advance science at the exascale advertisement. Computing in Science and Engineering, 17(3): c2, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/03/mcs20150300c2.
pdf.

Anonymous:2015:RSPa

[Ano15-38] Anonymous. Rock stars of 3D printing [advertisement]. Computing in Science and Engineering, 17(1):c4, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/01/mcs20150100c4.
pdf.

Anonymous:2015:RSPb

[Ano15-39] Anonymous. Rock stars of 3D printing house advertisement. Computing in Science and Engineering, 17(2): c4, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/02/mcs20150200c4.
pdf.

Anonymous:2015:RSCa

[Ano15-40] Anonymous. Rock stars of cybersecurity house advertisement. Computing in Science and Engineering, 17(3): c3, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/

                                        cs/2015/03/mcs20150300c3.
                                        pdf.

Anonymous:2015:RSCb

[Ano15-41] Anonymous. Rock stars of cybersecurity house advertisement. Computing in Science and Engineering, 17(4): c2, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/04/mcs20150400c2.
                                        pdf.

Anonymous:2015:RSCc

[Ano15-42]  Anonymous.  Rock stars

of cybersecurity house advertisement. Computing in Science and Engineering, 17(5):c2, September/ October 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/05/mcs20150500c2.
                                        pdf.

Anonymous:2015:RSC

[Ano15-43] Anonymous. Rock stars of cybersecurity house advertisement. Computing in Science and Engineering, 17 (6):c2, November/December 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2015:SFH

[Ano15-44] Anonymous. Scientific freedom & human rights [advertisement]. Computing in Science and Engineering, 17(1): 67, January/February 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:/

/csdl.computer.org/csdl/
mags/cs/2015/01/mcs2015010067.
pdf.

Anonymous:2015:SCH

[Ano15-45] Anonymous. Stay connected house advertisement. Computing in Science and Engineering, 17(4):27, July/ August 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040027.
pdf.

Anonymous:2015:TCa

[Ano15-46] Anonymous. Table of contents. Computing in Science and Engineering, 17(1):2–3, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/01/mcs2015010002.
pdf.

Anonymous:2015:TCb

[Ano15-47] Anonymous. Table of contents. Computing in Science and Engineering, 17 (6):2–3, November/December 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2015:TCLa

[Ano15-48] Anonymous. Take the CS Library wherever you go! [advertisement]. Computing in Science and Engineering, 17(1):37, January/ February 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/01/mcs2015010037.
                                        pdf.

Anonymous:2015:TCLb

[Ano15-49] Anonymous. Take the CS Library wherever you go! House advertisement. Computing in Science and Engineering, 17 (3):87, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/03/mcs2015030087.
                                        pdf.

Anonymous:2016:RT

[Ano16a] Anonymous. 2015 reviewer thanks. Computing in Science and Engineering, 18(1):6–8, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICCb

[Ano16b]  Anonymous.  2016 IEEE-

CS Charles Babbage Award. Computing in Science and Engineering, 18(6):c3, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs20160600c3.
pdf.

Anonymous:2016:CNEa

[Ano16c] Anonymous. Call for nominees: Education awards nominations house advertisement. Computing in Science and Engineering, 18(3):69, May/June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:CNEb

[Ano16d] Anonymous. Call for nominees education awards nominations house advertisement. Computing in Science and Engineering, 18(5):c3, September/October 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:CPC

[Ano16e] Anonymous. CiSE is the perfect combination house advertisement. Computing in Science and Engineering, 18(3): c4, May/June 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:FCa

[Ano16f] Anonymous. Front cover.

Computing in Science and Engineering, 18(1):c1, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:FCb

[Ano16g] Anonymous. Front cover. Computing in Science and Engineering, 18(2):c1, March/ April 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:FCc

[Ano16h] Anonymous. Front cover. Computing in Science and Engineering, 18(3):c1, May/June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:FCd

[Ano16i] Anonymous. Front cover. Computing in Science and Engineering, 18(4):c1, July/ August 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:FCe

[Ano16j] Anonymous. Front cover. Computing in Science and Engineering, 18(5):c1, September/October 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:FCf

[Ano16k] Anonymous. Front cover. Computing in Science and Engineering, 18(6):c1, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs20160600c1.
pdf.

Anonymous:2016:ICCa

[Ano16l] Anonymous. IEEE Cloud Computing call for papers house advertisement. Computing in Science and Engineering, 18(2):c4, March/April 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSo

[Ano16m]  Anonymous.  IEEE Com-

puter Society. Computing in Science and Engineering, 18(6):57, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs2016060057.
pdf.

Anonymous:2016:ICSg

[Ano16n] Anonymous. IEEE Computer Society 2016 call for major award nominations house advertisement. Computing in Science and Engineering, 18 (3):95, May/June 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSk

[Ano16o] Anonymous. IEEE Computer Society 2016 call for major award nominations house advertisement. Computing in Science and Engineering, 18 (4):c4, July/August 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSm

[Ano16p]  Anonymous.  IEEE Com-

puter Society 2016 call for major award nominations house advertisement. Computing in Science and Engineering, 18(5):93, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSb

[Ano16q]  Anonymous.  IEEE Com-

puter Society: Be at the center of it all house advertisement. Computing in Science and Engineering, 18(1): c3, January/February 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSd

[Ano16r]  Anonymous.  IEEE Com-

puter Society: Be at the center of it all house advertisement. Computing in Science and Engineering, 18(2): 57, March/April 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSh

[Ano16s] Anonymous. IEEE Computer Society: Get the recognition you deserve house advertisement. Computing in Science and Engineering, 18(3): c3, May/June 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSa

[Ano16t]  Anonymous.  IEEE Com-

puter Society house advertisement. Computing in Science and Engineering, 18(1):9, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSc

[Ano16u]  Anonymous.  IEEE Com-

puter Society house advertisement. Computing in Science and Engineering, 18(2): 9, March/April 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSf

[Ano16v] Anonymous. IEEE Computer Society house advertisement. Computing in Science and Engineering, 18(3):31, May/June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSj

[Ano16w]  Anonymous.  IEEE Com-

puter Society house advertisement. Computing in Science and Engineering, 18(4): c3, July/August 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSl

[Ano16x]  Anonymous.  IEEE Com-

puter Society house advertisement. Computing in Science and Engineering, 18(5): 47, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSp

[Ano16y]  Anonymous.  IEEE Com-

puter Society is where you choose the resources that fit your career. Computing in Science and Engineering, 18(6):c2, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs20160600c2.
pdf.

Anonymous:2016:ICSe

[Ano16z] Anonymous. IEEE Computer Society: Rock stars of big data house advertisement. Computing in Science and Engineering, 18(3):c2, May/June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ICSi

[Ano16-27]  Anonymous.  IEEE Com-

puter Society: Rock stars of big data house advertisement. Computing in Science and Engineering, 18(4): c2, July/August 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ITBa

[Ano16-28] Anonymous. IEEE Transactions on Big Data house advertisement. Computing in Science and Engineering, 18(1): c4, January/February 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ITBb

[Ano16-29] Anonymous. IEEE Transactions on Big Data house advertisement. Computing in Science and Engineering, 18(2): 89, March/April 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:ITS

[Ano16-30] Anonymous. IEEE Transactions on Sustainable Computing house advertisement. Computing in Science and Engineering, 18(1):c2, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:Ma

[Ano16-31]  Anonymous.  Masthead.

Computing in Science and Engineering, 18(1):1, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:Mb

[Ano16-32] Anonymous. Masthead. Computing in Science and Engineering, 18(2):1, March/April 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:Mc

[Ano16-33] Anonymous. Masthead. Computing in Science and Engineering, 18(3):1, May/June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:Md

[Ano16-34] Anonymous. Masthead. Computing in Science and Engineering, 18(4):1, July/August 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:Me

[Ano16-35] Anonymous. Masthead. Computing in Science and Engineering, 18(5):1, September/October 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:Mf

[Ano16-36] Anonymous. Masthead. Computing in Science and Engineering, 18(6):1, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs2016060001.
pdf.

Anonymous:2016:RSBb

[Ano16-37]  Anonymous.  Rock stars

of big data. Computing in Science and Engineering, 18(6):c4, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs20160600c4.
pdf.

Anonymous:2016:RSBa

[Ano16-38] Anonymous. Rock stars of big data house advertisement. Computing in Science and Engineering, 18(5): c2, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:RSP

[Ano16-39] Anonymous. Rock stars of pervasive, predictive analytics house advertisement. Computing in Science and Engineering, 18(5):c4, September/October 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:RSR

[Ano16-40] Anonymous. Rock stars of risk-based security house advertisement. Computing in Science and Engineering, 18 (2):c2, March/April 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:TCa

[Ano16-41] Anonymous. Table of contents. Computing in Science and Engineering, 18(1):2–3, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:TCb

[Ano16-42] Anonymous. Table of contents. Computing in Science and Engineering, 18(2): 2–3, March/April 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:TCc

[Ano16-43] Anonymous. Table of contents. Computing in Science and Engineering, 18(3): 2–3, May/June 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:TCd

[Ano16-44] Anonymous. Table of contents. Computing in Science and Engineering, 18(4): 2–3, July/August 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:TCe

[Ano16-45] Anonymous. Table of contents. Computing in Science and Engineering, 18 (5):2–3, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:TCf

[Ano16-46] Anonymous. Table of contents. Computing in Science and Engineering, 18(6):2– 3, November/December 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2016/06/mcs2016060002.
pdf.

Anonymous:2016:WWLa

[Ano16-47]  Anonymous.  Watch the

world’s leading experts take multi-core strategies to new heights house advertisement. Computing in Science and Engineering, 18(1):101, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2016:WWLb

[Ano16-48]  Anonymous.  Watch the

world’s leading experts take multi-core strategies to new heights house advertisement. Computing in Science and Engineering, 18(2):c3, March/ April 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2017:BRR

[Ano17a] Anonymous. 2017 B. Ramakrishna Rau Award call for nominations. Computing in Science and Engineering, 19(3): 51, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs2017030051.
pdf.

Anonymous:2017:CNE

[Ano17b] Anonymous. Call for nominees education awards nominations. Computing in Science and Engineering, 19(5): c3, September/October 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2017/05/mcs20170500c3.
pdf.

Anonymous:2017:CWW

[Ano17c]  Anonymous.  Computing

in World War I. Computing in Science and Engineering, 19(4):88, July/ August 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/04/mcs2017040088.
html.

Anonymous:2017:FYJ

[Ano17d] Anonymous. Focus on your job search. Computing in Science and Engineering, 19 (3):c4, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs20170300c4.
pdf.

Anonymous:2017:FCa

[Ano17e] Anonymous. Front cover.

Computing in Science and Engineering, 19(1):c1, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs20170100c1.
pdf.

Anonymous:2017:FCb

[Ano17f] Anonymous. Front cover.

Computing in Science and Engineering, 19(2):c1, March/ April 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://www.computer.org/
csdl/mags/cs/2017/02/mcs20170200c1.
pdf.

Anonymous:2017:FCc

[Ano17g] Anonymous. Front cover. Computing in Science and Engineering, 19(3):c1, May/ June 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cs/2017/03/mcs20170300c1.
                                        pdf.

Anonymous:2017:FCd

[Ano17h] Anonymous. Front cover. Computing in Science and Engineering, 19(4):c1, July/ August 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/04/mcs20170400c1.
                                        pdf.

Anonymous:2017:FC

[Ano17i] Anonymous. Front cover. Computing in Science and Engineering, 19(6):c1, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/06/mcs20170600c1.
                                        pdf.

Anonymous:2017:ICSa

[Ano17j] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 19(1):c3, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs20170100c3.
pdf.

Anonymous:2017:ICSb

[Ano17k] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 19(2): c3, March/April 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/02/mcs20170200c3.
pdf.

Anonymous:2017:ICSc

[Ano17l] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 19(3): 19, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs2017030019.
pdf.

Anonymous:2017:ICSd

[Ano17m] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 19(4): 29, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/04/mcs2017040029.
pdf.

Anonymous:2017:ICSf

[Ano17n]  Anonymous.  IEEE Com-

puter Society. Computing in Science and Engineering, 19(5):c2, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/05/mcs20170500c2.
pdf.

Anonymous:2017:ICS

[Ano17o]  Anonymous.  IEEE Com-

puter Society. Computing in Science and Engineering, 19(6):c2, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs20170600c2.
pdf.

Anonymous:2017:ICSe

[Ano17p]  Anonymous.  IEEE Com-

puter Society 2017 call for major award nominations. Computing in Science and Engineering, 19(4):c3, July/ August 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/04/mcs20170400c3.
pdf.

Anonymous:2017:Ma

[Ano17q]  Anonymous.  Masthead.

Computing in Science and Engineering, 19(1):1, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs2017010001.
pdf.

Anonymous:2017:Mb

[Ano17r] Anonymous. Masthead. Computing in Science and Engineering, 19(2):1, March/April 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://www.computer.org/
csdl/mags/cs/2017/02/mcs2017020001.
pdf.

Anonymous:2017:Mc

[Ano17s] Anonymous. Masthead. Computing in Science and Engineering, 19(3):1, May/June 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://www.computer.org/
csdl/mags/cs/2017/03/mcs2017030001.
pdf.

Anonymous:2017:Md

[Ano17t]  Anonymous.  Masthead.

Computing in Science and Engineering, 19(4):1, July/ August 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/04/mcs2017040001.
pdf.

Anonymous:2017:Me

[Ano17u] Anonymous. Masthead. Computing in Science and Engineering, 19(5):1, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/05/mcs2017050001.
                                        pdf.

Anonymous:2017:Mf

[Ano17v] Anonymous. Masthead. Computing in Science and Engineering, 19(6):1, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/06/mcs2017060001.
                                        pdf.

Anonymous:2017:Mg

[Ano17w] Anonymous. myCS. Computing in Science and Engineering, 19(6):c4, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/06/mcs20170600c4.
                                        pdf.

Anonymous:2017:NMOa

[Ano17x] Anonymous. New membership options for a better fit. Computing in Science and Engineering, 19(1):c2, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs20170100c2.
pdf.

Anonymous:2017:NMOb

[Ano17y]  Anonymous.  New mem-

bership options for a better fit. Computing in Science and Engineering, 19(2): c4, March/April 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/02/mcs20170200c4.
pdf.

Anonymous:2017:NMOc

[Ano17z]  Anonymous.  New mem-

bership options for a better fit. Computing in Science and Engineering, 19(3): c2, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs20170300c2.
pdf.

Anonymous:2017:NMOe

[Ano17-27] Anonymous. New membership options for a better fit. Computing in Science and Engineering, 19(5):c4, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        cs/2017/05/mcs20170500c4.
                                        pdf.

Anonymous:2017:NMOd

[Ano17-28] Anonymous. New membership options for a better fit [advertisement]. Computing in Science and Engineering, 19(4): c4, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/04/mcs20170400c4.
                                        pdf.

Anonymous:2017:OMUa

[Ano17-29] Anonymous. One membership. Unlimited knowledge [advertisement]. Computing in Science and Engineering, 19 (4):c2, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/04/mcs20170400c2.
                                        pdf.

Anonymous:2017:OMUb

[Ano17-30]  Anonymous.  One mem-

bership. Unlimited knowledge [advertisement]. Computing in Science and Engineering, 19(6):c3, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/06/mcs20170600c3.
                                        pdf.

Anonymous:2017:RMS

[Ano17-31] Anonymous. Richard E. Merwin Scholarship IEEE Computer Society. Richard E. Merwin Student Leadership Scholarship. Computing in Science and Engineering, 19(3): c3, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs20170300c3.
pdf.

Anonymous:2017:TCa

[Ano17-32] Anonymous. Table of contents. Computing in Science and Engineering, 19(1):2–3, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs2017010002.
pdf.

Anonymous:2017:TCb

[Ano17-33] Anonymous. Table of contents. Computing in Science and Engineering, 19(2): 2–3, March/April 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/02/mcs2017020002.
pdf.

Anonymous:2017:TCc

[Ano17-34] Anonymous. Table of contents. Computing in Science and Engineering, 19(3):

2–3, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs2017030002.
pdf.

Anonymous:2017:TCd

[Ano17-35] Anonymous. Table of contents. Computing in Science and Engineering, 19(4): 2–3, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/04/mcs2017040002.
pdf.

Anonymous:2017:TCe

[Ano17-36] Anonymous. Table of contents. Computing in Science and Engineering, 19(5):2– 3, September/October 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2017/05/mcs2017050002.
pdf.

Anonymous:2017:TC

[Ano17-37] Anonymous. Table of contents. Computing in Science and Engineering, 19(6):2– 3, November/December 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2017/06/mcs2017060002.
pdf.

Anonymous:2017:Ta

[Ano17-38]  Anonymous.  TechIngnite.

Computing in Science and Engineering, 19(1):c4, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs20170100c4.
pdf.

Anonymous:2017:Tb

[Ano17-39]  Anonymous.  TechIngnite.

Computing in Science and Engineering, 19(2):c2, March/ April 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://www.computer.org/
csdl/mags/cs/2017/02/mcs20170200c2.
pdf.

Anonymous:2017:VHW

[Ano17-40] Anonymous. VR is hot, but why? Computing in Science and Engineering, 19(4): 4–5, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/04/mcs2017040004.
html.

Anonymous:2018:RT

[Ano18a] Anonymous. 2017 reviewer thanks. Computing in Science and Engineering, 20(2): 4–5, March/April 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:ICC

[Ano18b] Anonymous. 2019 IEEE-CS Charles Babbage Award: Call for nominations ad. Computing in Science and Engineering, 20(3):c4, May/June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:AI

[Ano18c] Anonymous. AIP information. Computing in Science and Engineering, 20 (2):110, March/April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8317972/.

Anonymous:2018:AIPa

[Ano18d] Anonymous. American Institute of Physics information. Computing in Science and Engineering, 20(3):102, May/ June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8358042/.

Anonymous:2018:AIPb

[Ano18e] Anonymous. American Institute of Physics information. Computing in Science and Engineering, 20(4): 126, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

cs/2018/04/mcs2018040126.
pdf.

Anonymous:2018:AIPc

[Ano18f] Anonymous. American Institute of Physics information. Computing in Science and Engineering, 20(5):129, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/05/mcs2018050129.
pdf.

Anonymous:2018:CP

[Ano18g] Anonymous. Call for papers. Computing in Science and Engineering, 20(6):95, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2018/06/08627355.
pdf.

Anonymous:2018:EA

[Ano18h]  Anonymous.  Education

awards. Computing in Science and Engineering, 20(4): 128, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040128.
pdf.

Anonymous:2018:E

[Ano18i] Anonymous. Erratum. Computing in Science and Engineering, 20(6):4, November/December 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/06/08625901.pdf.

Anonymous:2018:EGS

[Ano18j] Anonymous. Erratum: Glimpses of space-time beyond the singularities using supercomputers. Computing in Science and Engineering, 20(5):4, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/05/mcs2018050004.

html. See [Sin18].

Anonymous:2018:FCa

[Ano18k] Anonymous. Front cover. Computing in Science and Engineering, 20(1):c1, January/February 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:FCb

[Ano18l] Anonymous. Front cover. Computing in Science and Engineering, 20(2):c1, March/ April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:FCc

[Ano18m] Anonymous. Front cover. Computing in Science and Engineering, 20(3):c1, May/June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:FCd

[Ano18n] Anonymous. Front cover.

Computing in Science and Engineering, 20(4):c1, July/ August 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs20180400c1.
pdf.

Anonymous:2018:FC

[Ano18o] Anonymous. Front cover.

Computing in Science and Engineering, 20(6):C1, November/December 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/06/08627362.pdf.

Anonymous:2018:HMA

[Ano18p] Anonymous. Harlan Mills Award. Computing in Science and Engineering, 20(4): 89, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040089.
pdf.

Anonymous:2018:HAA

[Ano18q] Anonymous. HPC awards ad. Computing in Science and Engineering, 20(3):c2, May/ June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:ICSg

[Ano18r] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 20(6):C3, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/cs/2018/06/08627356.
                                        pdf.

Anonymous:2018:ICSf

[Ano18s] Anonymous. IEEE Computer Society: Be at the center of it all. Computing in Science and Engineering, 20(6):C2, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/cs/2018/06/08627359.
                                        pdf.

Anonymous:2018:ICSa

[Ano18t]  Anonymous.  IEEE Com-

puter Society information. Computing in Science and Engineering, 20(1):c3, January/February 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:ICSb

[Ano18u] Anonymous. IEEE Computer Society information. Computing in Science and Engineering, 20(2):c3, March/April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8317989/.

Anonymous:2018:ICSc

[Ano18v] Anonymous. IEEE Computer Society information. Computing in Science and Engineering, 20(3):c3, May/June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8358024/.

Anonymous:2018:ICSd

[Ano18w]  Anonymous.  IEEE Com-

puter Society information. Computing in Science and Engineering, 20(4):4, July/ August 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040004.
pdf.

Anonymous:2018:ICSe

[Ano18x] Anonymous. IEEE Computer Society information. Computing in Science and Engineering, 20(5):c3, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/05/mcs20180500c3.
pdf.

Anonymous:2018:ILC

[Ano18y]  Anonymous.  IEEE let-

ters of the Computer Society. Computing in Science and Engineering, 20(6):80, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/cs/2018/06/08627360.
                                        pdf.

Anonymous:2018:ITE

[Ano18z] Anonymous. IEEE Technology and Engineering Management Society TEMS Administrative Committee. Computing in Science and Engineering, 20(6):3, November/December 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/06/08627365.pdf.

Anonymous:2018:ITB

[Ano18-27] Anonymous. IEEE Transactions on Big Data. Computing in Science and Engineering, 20(6):5, November/December 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/06/08627358.pdf.

Anonymous:2018:ITS

[Ano18-28] Anonymous. IEEE Transactions on Sustainable Computing. Computing in Science and Engineering, 20(6):66, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2018/06/08627357.
pdf.

Anonymous:2018:JB

[Ano18-29]  Anonymous.  Jobs board.

Computing in Science and Engineering, 20(4):74, July/ August 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040074.
pdf.

Anonymous:2018:LC

[Ano18-30] Anonymous. Letters of the CS. Computing in Science and Engineering, 20(4): c2, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs20180400c2.
pdf.

Anonymous:2018:Ma

[Ano18-31] Anonymous. Masthead. Computing in Science and Engineering, 20(2):3, March/April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8317993/.

Anonymous:2018:Mb

[Ano18-32] Anonymous. Masthead. Computing in Science and Engineering, 20(3):3, May/June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8357985/.

Anonymous:2018:Mc

[Ano18-33]  Anonymous.  Masthead.

Computing in Science and Engineering, 20(4):3, July/ August 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/04/mcs2018040003.
                                        pdf.

Anonymous:2018:Md

[Ano18-34] Anonymous. Masthead. Computing in Science and Engineering, 20(5):3, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/05/mcs2018050003.
                                        pdf.

Anonymous:2018:MHAa

[Ano18-35]  Anonymous.  Membership

house ad. Computing in Science and Engineering, 20(2): c2, March/April 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:MHAc

[Ano18-36]  Anonymous.  Membership

house ad. Computing in Science and Engineering, 20(5): c4, September/October 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2018/05/mcs20180500c4.
pdf.

Anonymous:2018:MHAb

[Ano18-37]  Anonymous.  Membership

house advertisement. Computing in Science and Engineering, 20(4):127, July/ August 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040127.
pdf.

Anonymous:2018:MAb

[Ano18-38] Anonymous. Merwin Award.

Computing in Science and Engineering, 20(4):c3, July/ August 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs20180400c3.
pdf.

Anonymous:2018:MAa

[Ano18-39] Anonymous. myCS ad. Computing in Science and Engineering, 20(3):104, May/June 2018. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:MML

[Ano18-40] Anonymous. myCS more to love house ad. Computing in Science and Engineering, 20 (2):c4, March/April 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:NSC

[Ano18-41] Anonymous. National Strategic Computing Initiative. Computing in Science and Engineering, 20(5):c1, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/05/mcs20180500c1.
                                        pdf.

Anonymous:2018:OMU

[Ano18-42]  Anonymous.  One mem-

bership. Unlimited knowledge. Computing in Science and Engineering, 20(6):72, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/cs/2018/06/08627364.
                                        pdf.

Anonymous:2018:SN

[Ano18-43]  Anonymous.  Social net-

working. Computing in Science and Engineering, 20(4): c4, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs20180400c4.
pdf.

Anonymous:2018:SNH

[Ano18-44]  Anonymous.  Social net-

working house ad. Computing in Science and Engineering, 20(5):c2, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/05/mcs20180500c2.
pdf.

Anonymous:2018:SC

[Ano18-45]  Anonymous.  Stay con-

nected. Computing in Science and Engineering, 20(6):C4, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2018/06/08627361.
pdf.

Anonymous:2018:SMA

[Ano18-46] Anonymous. Student membership ad. Computing in Science and Engineering, 20(3): 103, May/June 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:TCa

[Ano18-47] Anonymous. Table of contents. Computing in Science and Engineering, 20(1):1–2, January/February 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:TCb

[Ano18-48] Anonymous. Table of contents. Computing in Science and Engineering, 20(2): 1–2, March/April 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:TCc

[Ano18-49] Anonymous. Table of contents. Computing in Science and Engineering, 20(3): 1–2, May/June 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:TCd

[Ano18-50] Anonymous. Table of contents. Computing in Science and Engineering, 20(4): 1–2, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/04/mcs2018040001.
                                        pdf.

Anonymous:2018:TCe

[Ano18-51] Anonymous. Table of contents. Computing in Science and Engineering, 20(5):1– 2, September/October 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2018/05/mcs2018050001.
pdf.

Anonymous:2018:TCf

[Ano18-52] Anonymous. Table of contents. Computing in Science and Engineering, 20(6):1– 2, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2018/06/08627363.
pdf.

Anonymous:2018:TAa

[Ano18-53]  Anonymous.  TMSCS ad.

Computing in Science and Engineering, 20(1):107, January/February 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2018:TAb

[Ano18-54]  Anonymous.  TSUSC ad.

Computing in Science and Engineering, 20(1):108, January/February 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:CEa

[Ano19a] Anonymous. Computing Edge.

Computing in Science and Engineering, 21(5):C4, September/October 2019. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:CEb

[Ano19b] Anonymous. Computing Edge. Computing in Science and Engineering, 21(6):C4, November/December 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:ISP

[Ano19c] Anonymous. IEEE Security & Privacy. Computing in Science and Engineering, 21(5): 94, September/October 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:Ca

[Ano19d]  Anonymous.  COMPSAC

2019. Computing in Science and Engineering, 21(3): C2, May/June 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:Cb

[Ano19e]  Anonymous.  COMPSAC

2019. Computing in Science and Engineering, 21(4): C2, July/August 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:CH

[Ano19f]  Anonymous.  Connected

health. Computing in Science and Engineering, 21(2): C3, March/April 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:Cc

[Ano19g] Anonymous. Cover 4. Computing in Science and Engineering, 21(4):C4, July/ August 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:FCa

[Ano19h] Anonymous. Front cover.

Computing in Science and Engineering, 21(1):C1, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:FCb

[Ano19i] Anonymous. Front cover.

Computing in Science and Engineering, 21(2):C1, March/ April 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:FCc

[Ano19j] Anonymous. [Front cover].

Computing in Science and Engineering, 21(3):C1, May/ June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:FCd

[Ano19k] Anonymous. Front cover.

Computing in Science and Engineering, 21(4):C1, July/ August 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:FCe

[Ano19l] Anonymous. Front cover. Computing in Science and Engineering, 21(5):C1, September/October 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:FCf

[Ano19m] Anonymous. Front cover. Computing in Science and Engineering, 21(6):C1, November/December 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:ICSb

[Ano19n] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 21(1): C3, January/February 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:ICSc

[Ano19o] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 21(2): 2, March/April 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:ICSd

[Ano19p] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 21(2): 96, March/April 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:ICSe

[Ano19q] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 21(3): C3, May/June 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:ICSf

[Ano19r] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 21(4): C3, July/August 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:ICSh

[Ano19s] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 21(5): C3, September/October 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:ICS

[Ano19t]  Anonymous.  IEEE Com-

puter Society. Computing in Science and Engineering, 21 (6):C3, November/December 2019. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:ICSa

[Ano19u] Anonymous. IEEE Computer Society: Be at the center of it all. Computing in Science and Engineering, 21(1):C2, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:ICSg

[Ano19v] Anonymous. IEEE Computer Society Jobs Board. Computing in Science and Engineering, 21(5):C2, September/October 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:ID

[Ano19w] Anonymous. IEEE Data-Port. Computing in Science and Engineering, 21(5): 95, September/October 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:IIC

[Ano19x] Anonymous. IEEE Internet computing. Computing in Science and Engineering, 21(2): 1, March/April 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:ITE

[Ano19y] Anonymous. IEEE Technology and Engineering Management Society TEMS Administrative Committee. Computing in Science and Engineering, 21(1):65, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:IWC

[Ano19z] Anonymous. IEEE World Congress on Services 2019. Computing in Science and Engineering, 21(3):C4, May/ June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:KYCb

[Ano19-27] Anonymous. Keep your career options open. Computing in Science and Engineering, 21 (6):C2, November/December 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:KYCa

[Ano19-28] Anonymous. Keep your career options open: Upload your resume today! Computing in Science and Engineering, 21(4):87, July/August 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:LCSa

[Ano19-29] Anonymous. Letters of the Computer Society. Computing in Science and Engineering, 21(2):C2, March/April 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:LCSb

[Ano19-30] Anonymous. Letters of the Computer Society. Computing in Science and Engineering, 21 (3):103, May/June 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:LBTa

[Ano19-31]  Anonymous.  Looking for

the BEST tech job for you? Computing in Science and Engineering, 21(2): C4, March/April 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:LBTb

[Ano19-32] Anonymous. Looking for the BEST tech job for you? Computing in Science and Engineering, 21(3):120, May/June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:Ma

[Ano19-33]  Anonymous.  Masthead.

Computing in Science and Engineering, 21(1):1, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:Mb

[Ano19-34] Anonymous. Masthead. Computing in Science and Engineering, 21(3):1, May/June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8701546/.

Anonymous:2019:Mc

[Ano19-35] Anonymous. Masthead. Computing in Science and Engineering, 21(4):1, July/August 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:Md

[Ano19-36] Anonymous. Masthead. Computing in Science and Engineering, 21(5):1, September/October 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:M

[Ano19-37] Anonymous. Masthead. Computing in Science and Engineering, 21(6):1, November/December 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:NWC

[Ano19-38] Anonymous. Nominate a worthy colleague today! Computing in Science and Engineering, 21(3):119, May/June 2019. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:SCa

[Ano19-39] Anonymous. Stay connected. Computing in Science and Engineering, 21(1):C4, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:SCb

[Ano19-40] Anonymous. Stay connected. Computing in Science and Engineering, 21(5):96, September/October 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:TCa

[Ano19-41] Anonymous. Table of contents. Computing in Science and Engineering, 21(1):2–3, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:TCb

[Ano19-42] Anonymous. Table of contents. Computing in Science and Engineering, 21(2): 3–4, March/April 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:TCc

[Ano19-43] Anonymous. Table of contents. Computing in Science and Engineering, 21(3):

2–3, May/June 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2019:TCd

[Ano19-44] Anonymous. Table of contents. Computing in Science and Engineering, 21(4): 2–3, July/August 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:TCe

[Ano19-45] Anonymous. Table of contents. Computing in Science and Engineering, 21 (5):2–3, September/October 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anonymous:2019:TCf

[Ano19-46] Anonymous. Table of contents. Computing in Science and Engineering, 21 (6):2–3, November/December 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CGAa

[Ano20a]  Anonymous.  Computer

Graphics and Applications. Computing in Science and Engineering, 22(4):103, July/ August 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CEa

[Ano20b] Anonymous. Computing Edge. Computing in Science and Engineering, 22(1):C4, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CEb

[Ano20c]  Anonymous.  Computing

Edge. Computing in Science and Engineering, 22(2): C4, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:Cd

[Ano20d] Anonymous. ComputingEdge. Computing in Science and Engineering, 22(4):C4, July/ August 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:Ce

[Ano20e] Anonymous. ComputingEdge. Computing in Science and Engineering, 22(6):C4, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICG

[Ano20f] Anonymous. IEEE Computer Graphics and Applications. Computing in Science and Engineering, 22(6):59, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:IPC

[Ano20g] Anonymous. IEEE Pervasive Computing. Computing in Science and Engineering, 22(4): 25, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ISPa

[Ano20h] Anonymous. IEEE Security & Privacy. Computing in Science and Engineering, 22(3): 119, May/June 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ISPb

[Ano20i] Anonymous. IEEE Security & Privacy. Computing in Science and Engineering, 22 (6):20, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ITB

[Ano20j] Anonymous. IEEE Transactions on Big Data. Computing in Science and Engineering, 22 (6):111, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ITSa

[Ano20k] Anonymous. IEEE Transactions on Sustainable Computing. Computing in Science and Engineering, 22(3):39, May/ June 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ITSb

[Ano20l] Anonymous. IEEE Transactions on Sustainable Computing. Computing in Science and Engineering, 22(4): 60, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ITS

[Ano20m] Anonymous. IEEE Transactions on Sustainable Computing. Computing in Science and Engineering, 22(6):86, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CMA

[Ano20n] Anonymous. Call for 2020 major awards nominations. Computing in Science and Engineering, 22(3):120, May/June 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CAb

[Ano20o] Anonymous. Call for articles. Computing in Science and Engineering, 22(4): 109, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CAc

[Ano20p] Anonymous. Call for articles. Computing in Science and Engineering, 22(5): 94, September/October 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CAIa

[Ano20q] Anonymous. Call for articles IEEE Pervasive Computing. Computing in Science and Engineering, 22(2): 7, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CAIb

[Ano20r] Anonymous. Call for articles:

IEEE Pervasive Computing. Computing in Science and Engineering, 22(6):102, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CN

[Ano20s] Anonymous. Call for nominations. Computing in Science and Engineering, 22(3): 51, May/June 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CP

[Ano20t] Anonymous. Call for papers.

Computing in Science and Engineering, 22(3):63, May/June 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CPIa

[Ano20u] Anonymous. Call for papers: IEEE Transactions on Computers. Computing in Science and Engineering, 22(5): 82, September/October 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CPIb

[Ano20v] Anonymous. Call for papers: IEEE Transactions on Computers. Computing in Science and Engineering, 22 (6):114, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CAa

[Ano20w] Anonymous. CGA ad. Computing in Science and Engineering, 22(2):87, March/ April 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:Cb

[Ano20x]  Anonymous.  COMPSAC

2020. Computing in Science and Engineering, 22(1):101, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:Cc

[Ano20y]  Anonymous.  COMPSAC

2020. Computing in Science and Engineering, 22(2): C2, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:CGAb

[Ano20z] Anonymous. Computer graphics and applications. Computing in Science and Engineering, 22(5):71, September/October 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:FCa

[Ano20-27] Anonymous. [front cover].

Computing in Science and Engineering, 22(1):C1, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:FCb

[Ano20-28] Anonymous. Front cover.

Computing in Science and Engineering, 22(2):C1, March/ April 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:FCc

[Ano20-29] Anonymous. Front cover.

Computing in Science and Engineering, 22(3):C1, May/ June 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:FCd

[Ano20-30] Anonymous. Front cover. Computing in Science and Engineering, 22(4):C1, July/ August 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:FCe

[Ano20-31] Anonymous. Front cover. Computing in Science and Engineering, 22(5):C1, September/October 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:FC

[Ano20-32] Anonymous. Front cover. Computing in Science and Engineering, 22(6):C1, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:GPNa

[Ano20-33] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 22(3):C4, May/ June 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:GPNb

[Ano20-34] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 22(5):

95, September/October 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:GPNc

[Ano20-35] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 22(6):C3, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:HDM

[Ano20-36] Anonymous. Harlan D. Mills Award: Call for software engineering award nominations. Computing in Science and Engineering, 22(5): 83, September/October 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:HRN

[Ano20-37] Anonymous. HOST 2020 register now. Computing in Science and Engineering, 22(2): 96, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSb

[Ano20-38] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 22(1): C3, January/February 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSc

[Ano20-39] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 22(2): C3, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSf

[Ano20-40] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 22(3): C3, May/June 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSi

[Ano20-41] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 22(4): C3, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSl

[Ano20-42] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 22(5): C3, September/October 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSm

[Ano20-43] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 22(6):

4, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSk

[Ano20-44]  Anonymous.  IEEE Com-

puter Society: Call for papers. Computing in Science and Engineering, 22(5): 28, September/October 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSn

[Ano20-45] Anonymous. IEEE Computer Society: Call for papers. Computing in Science and Engineering, 22(6):29, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSa

[Ano20-46]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 22(1): 52, January/February 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSd

[Ano20-47]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 22(3): C2, May/June 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSg

[Ano20-48]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 22(4): C2, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSj

[Ano20-49]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 22(5): C2, September/October 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSp

[Ano20-50] Anonymous. IEEE Computer Society Jobs Board. Computing in Science and Engineering, 22(6):113, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSe

[Ano20-51]  Anonymous.  IEEE Com-

puter Society Volunteer Service Awards. Computing in Science and Engineering, 22 (3):93, May/June 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSh

[Ano20-52]  Anonymous.  IEEE Com-

puter Society Volunteer Service Awards. Computing in Science and Engineering, 22(4):61, July/August 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:ICSo

[Ano20-53]  Anonymous.  IEEE Com-

puter Society Volunteer Service Awards. Computing in Science and Engineering, 22 (6):60, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:IIC

[Ano20-54] Anonymous. IEEE IRI 2020 call for papers. Computing in Science and Engineering, 22(2):95, March/April 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:IQWb

[Ano20-55] Anonymous. IEEE Quantum Week. Computing in Science and Engineering, 22(5): C4, September/October 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:IQWa

[Ano20-56] Anonymous. IEEE Quantum Week 2020 is open for submissions. Computing in Science and Engineering, 22(2):

77, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:KYC

[Ano20-57] Anonymous. Keep your career options open. Computing in Science and Engineering, 22(1):C2, January/February 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:Ma

[Ano20-58]  Anonymous.  Masthead.

Computing in Science and Engineering, 22(1):1, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:Mb

[Ano20-59] Anonymous. Masthead. Computing in Science and Engineering, 22(2):1, March/April 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:Mc

[Ano20-60] Anonymous. Masthead. Computing in Science and Engineering, 22(3):1, May/June 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:Md

[Ano20-61] Anonymous. Masthead. Computing in Science and Engineering, 22(4):1, July/August 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:Me

[Ano20-62] Anonymous. Masthead. Computing in Science and Engineering, 22(5):1, September/October 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:M

[Ano20-63] Anonymous. Masthead. Computing in Science and Engineering, 22(6):1, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:SC

[Ano20-64] Anonymous. Stay connected.

Computing in Science and Engineering, 22(6):8, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:TCa

[Ano20-65] Anonymous. Table of contents. Computing in Science and Engineering, 22(1):2–3, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:TCb

[Ano20-66] Anonymous. Table of contents. Computing in Science and Engineering, 22(2): 2–3, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:TCc

[Ano20-67] Anonymous. Table of contents. Computing in Science and Engineering, 22(3): 2–3, May/June 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:TCd

[Ano20-68] Anonymous. Table of contents. Computing in Science and Engineering, 22(4): 2–3, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:TCe

[Ano20-69] Anonymous. Table of contents. Computing in Science and Engineering, 22 (5):2–3, September/October 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:TC

[Ano20-70] Anonymous. Table of contents. Computing in Science and Engineering, 22 (6):2–3, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:TYO

[Ano20-71] Anonymous. Thank you to our sponsors! Computing in Science and Engineering, 22 (6):C2, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2020:TTT

[Ano20-72] Anonymous. Top technology trends for 2020 featured in Computer. Computing in Science and Engineering, 22(2): 54, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:CGAa

[Ano21a]  Anonymous.  Computer

Graphics and Applications. Computing in Science and Engineering, 23(1):74, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:CGAb

[Ano21b]  Anonymous.  Computer

Graphics and Applications. Computing in Science and Engineering, 23(3):41, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:Ca

[Ano21c] Anonymous. ComputingEdge.

Computing in Science and Engineering, 23(1):C2, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:Cb

[Ano21d] Anonymous. ComputingEdge. Computing in Science and Engineering, 23(2):C2, March/ April 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:Cd

[Ano21e] Anonymous. ComputingEdge. Computing in Science and Engineering, 23(3):C2, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:Ce

[Ano21f] Anonymous. ComputingEdge. Computing in Science and Engineering, 23(4):C2, July/ August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:Cf

[Ano21g] Anonymous. ComputingEdge. Computing in Science and Engineering, 23(5):C2, September/October 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:Cg

[Ano21h] Anonymous. ComputingEdge. Computing in Science and Engineering, 23(6):C2, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:IAHa

[Ano21i] Anonymous. IEEE Annals of the History of Computing. Computing in Science and Engineering, 23(1):88, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:IAHb

[Ano21j] Anonymous. IEEE Annals of the History of Computing. Computing in Science and Engineering, 23(2): 92, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:IAHc

[Ano21k] Anonymous. IEEE Annals of the History of Computing. Computing in Science and Engineering, 23(6):51, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICG

[Ano21l] Anonymous. IEEE Computer Graphics and Applications. Computing in Science and Engineering, 23(6):15, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ITBa

[Ano21m] Anonymous. IEEE Transactions on Big Data. Computing in Science and Engineering, 23(1):101, January/February 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ITBb

[Ano21n] Anonymous. IEEE Transactions on Big Data. Computing in Science and Engineering, 23 (3):13, May/June 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ITBc

[Ano21o] Anonymous. IEEE Transactions on Big Data. Computing in Science and Engineering, 23(4):28, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ITBd

[Ano21p] Anonymous. IEEE Transactions on Big Data. Computing in Science and Engineering, 23(5):38, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ITBe

[Ano21q] Anonymous. IEEE Transactions on Big Data. Computing in Science and Engineering, 23(6):41, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ITC

[Ano21r] Anonymous. IEEE Transactions on Computers. Computing in Science and Engineering, 23(2):46, March/April 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ITSa

[Ano21s] Anonymous. IEEE Transactions on Sustainable Computing. Computing in Science and Engineering, 23(1):107, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ITSb

[Ano21t] Anonymous. IEEE Transactions on Sustainable Computing. Computing in Science and Engineering, 23(4): 54, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ITSc

[Ano21u] Anonymous. IEEE Transactions on Sustainable Computing. Computing in Science and Engineering, 23(6):33, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:IPC

[Ano21v] Anonymous. IT Professional: Call for articles. Computing in Science and Engineering, 23(4):17, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:SPa

[Ano21w] Anonymous. Security & Privacy. Computing in Science and Engineering, 23(1): 92, January/February 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:TBD

[Ano21x] Anonymous. Transactions on Big Data. Computing in Science and Engineering, 23(2): 98, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:CMAa

[Ano21y] Anonymous. Call for 2021 major awards nominations. Computing in Science and Engineering, 23(1):108, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:CMAb

[Ano21z] Anonymous. Call for 2021 major awards nominations. Computing in Science and Engineering, 23(5):72, September/October 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:CAa

[Ano21-27] Anonymous. Call for articles. Computing in Science and Engineering, 23(1): 34, January/February 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:CAb

[Ano21-28] Anonymous. Call for articles. Computing in Science and Engineering, 23(5): 9, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:CAIa

[Ano21-29] Anonymous. Call for articles: IEEE Pervasive Computing. Computing in Science and Engineering, 23(1):106, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:CAIb

[Ano21-30] Anonymous. Call for articles: IEEE Pervasive Computing. Computing in Science and Engineering, 23(3):72, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:CAIc

[Ano21-31] Anonymous. Call for articles: IEEE Pervasive Computing. Computing in Science and Engineering, 23(4): 46, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:CPIa

[Ano21-32] Anonymous. Call for papers: IEEE Transactions on Computers. Computing in Science and Engineering, 23(4): 64, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:CPIb

[Ano21-33] Anonymous. Call for papers: IEEE Transactions on Computers. Computing in Science and Engineering, 23 (6):24, November/December 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:FCa

[Ano21-34] Anonymous. Front cover. Computing in Science and Engineering, 23(1):C1, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:FCb

[Ano21-35] Anonymous. Front cover. Computing in Science and Engineering, 23(2):C1, March/ April 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:FCc

[Ano21-36] Anonymous. [front cover].

Computing in Science and Engineering, 23(3):C1, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:FCd

[Ano21-37] Anonymous. Front cover.

Computing in Science and Engineering, 23(4):C1, July/ August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:FCe

[Ano21-38] Anonymous. [front cover].

Computing in Science and Engineering, 23(5):C1, September/October 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:FCf

[Ano21-39] Anonymous. [front cover].

Computing in Science and Engineering, 23(6):C1, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:GPNd

[Ano21-40] Anonymous. Get published in the new IEEE Open Journal of the Computer Scoiety. Computing in Science and Engineering, 23(4): C3, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:GPNe

[Ano21-41] Anonymous. Get published in the new IEEE Open Journal of the Computer Scoiety. Computing in Science and Engineering, 23(5):C3, September/October 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:GPNa

[Ano21-42] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 23(1):C3, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:GPNb

[Ano21-43] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 23(2): C3, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:GPNc

[Ano21-44] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 23(3):C3, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:GPNf

[Ano21-45] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 23(6):C3, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:HKS

[Ano21-46] Anonymous. Hans Karlsson Standards Award. Computing in Science and Engineering, 23(2):83, March/April 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:HDMa

[Ano21-47] Anonymous. Harlan D. Mills Award. Computing in Science and Engineering, 23(2): 120, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:HDMb

[Ano21-48] Anonymous. Harlan D. Mills Award: Call for software engineering award nominations. Computing in Science and Engineering, 23(4):

92, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSa

[Ano21-49] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 23(1):1, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSk

[Ano21-50] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 23(4): 1, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSn

[Ano21-51] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 23(5): 1, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSq

[Ano21-52] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 23(6): 1, November/December 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSc

[Ano21-53]  Anonymous.  IEEE Com-

puter Society: Call for papers. Computing in Science and Engineering, 23(1): 82, January/February 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSe

[Ano21-54] Anonymous. IEEE Computer Society: Call for papers. Computing in Science and Engineering, 23(2):1, March/April 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSh

[Ano21-55] Anonymous. IEEE Computer Society: Call for papers. Computing in Science and Engineering, 23(3):1, May/June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSi

[Ano21-56] Anonymous. IEEE Computer Society: Call for papers. Computing in Science and Engineering, 23(3):108, May/June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSl

[Ano21-57]  Anonymous.  IEEE Com-

puter Society: Call for papers. Computing in Science and Engineering, 23(4):

90, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSr

[Ano21-58] Anonymous. IEEE Computer Society: Call for papers. Computing in Science and Engineering, 23(6):57, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSd

[Ano21-59]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 23(1): C4, January/February 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSg

[Ano21-60]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 23(2): C4, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSj

[Ano21-61]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 23(3): C4, May/June 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSm

[Ano21-62]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 23(4): C4, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSo

[Ano21-63]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 23(5): 61, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSb

[Ano21-64]  Anonymous.  IEEE Com-

puter Society Jobs Board. Computing in Science and Engineering, 23(1):46, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSf

[Ano21-65] Anonymous. IEEE Computer Society Jobs Board. Computing in Science and Engineering, 23(2):64, March/April 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSp

[Ano21-66] Anonymous. IEEE Computer Society Jobs Board. Computing in Science and Engineering, 23(5):C4, September/October 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:ICSs

[Ano21-67] Anonymous. IEEE Computer Society Jobs Board. Computing in Science and Engineering, 23(6):C4, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:Ma

[Ano21-68]  Anonymous.  Masthead.

Computing in Science and Engineering, 23(1):2, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:Mb

[Ano21-69] Anonymous. Masthead. Computing in Science and Engineering, 23(2):2, March/April 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:Mc

[Ano21-70]  Anonymous.  [masthead].

Computing in Science and Engineering, 23(3):2, May/June 2021. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:Md

[Ano21-71] Anonymous. Masthead. Computing in Science and Engineering, 23(4):2, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:Me

[Ano21-72] Anonymous. Masthead. Computing in Science and Engineering, 23(5):2, September/October 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:Mf

[Ano21-73] Anonymous. Masthead. Computing in Science and Engineering, 23(6):2, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:RMS

[Ano21-74]  Anonymous.  Richard E.

Merwin student scholarship. Computing in Science and Engineering, 23(4):68, July/ August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:TCa

[Ano21-75] Anonymous. Table of contents. Computing in Science and Engineering, 23(1):3–4, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:TCb

[Ano21-76] Anonymous. Table of contents. Computing in Science and Engineering, 23(2): 3–4, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:TCc

[Ano21-77] Anonymous. Table of contents. Computing in Science and Engineering, 23(3): 3–4, May/June 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:TCd

[Ano21-78] Anonymous. Table of contents. Computing in Science and Engineering, 23(4): 3–4, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:TCe

[Ano21-79] Anonymous. Table of contents. Computing in Science and Engineering, 23 (5):3–4, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2021:TCf

[Ano21-80] Anonymous. Table of contents. Computing in Science and Engineering, 23 (6):3–4, November/December 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICA

[Ano22a]  Anonymous.  2023 IEEE

Conference on Artificial Intelligence. Computing in Science and Engineering, 24 (6):C2, November/December 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICC

[Ano22b]  Anonymous.  2023 IEEE

CS Charles Babbage Award. Computing in Science and Engineering, 24(1):87, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:AWC

[Ano22c]  Anonymous.  AI’s 10 to

watch: Call for nominations. Computing in Science and Engineering, 24(1):88, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:CEa

[Ano22d] Anonymous. Computing Edge.

Computing in Science and Engineering, 24(5):C4, September/October 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:CEb

[Ano22e] Anonymous. Computing Edge. Computing in Science and Engineering, 24(6):C4, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:CSE

[Ano22f] Anonymous. Computing in Science & Engineering publication information. Computing in Science and Engineering, 24(3):1, May/June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:Ca

[Ano22g] Anonymous. ComputingEdge. Computing in Science and Engineering, 24(1):C2, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:Cb

[Ano22h] Anonymous. ComputingEdge. Computing in Science and Engineering, 24(2):C2, March/ April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:IAHa

[Ano22i] Anonymous. IEEE Annals of the History of Computing. Computing in Science and Engineering, 24(1):7, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:IAHb

[Ano22j] Anonymous. IEEE Annals of the History of Computing. Computing in Science and Engineering, 24(2): 30, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICGa

[Ano22k]  Anonymous.  IEEE Com-

puter Graphics and Applications. Computing in Science and Engineering, 24(1): 45, January/February 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICGb

[Ano22l]  Anonymous.  IEEE Com-

puter Graphics and Applications. Computing in Science and Engineering, 24(2): 61, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICGc

[Ano22m]  Anonymous.  IEEE Com-

puter Graphics and Applications. Computing in Science and Engineering, 24(3): 85, May/June 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICGd

[Ano22n]  Anonymous.  IEEE Com-

puter Graphics and Applications. Computing in Science and Engineering, 24(4): 64, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ISPa

[Ano22o] Anonymous. IEEE Security & Privacy. Computing in Science and Engineering, 24(1): 77, January/February 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ISPb

[Ano22p] Anonymous. IEEE Security & Privacy. Computing in Science and Engineering, 24(2): 94, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ITBa

[Ano22q] Anonymous. IEEE Transactions on Big Data. Computing in Science and Engineering, 24(1):55, January/February 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ITSa

[Ano22r] Anonymous. IEEE Transactions on Sustainable Computing. Computing in Science and Engineering, 24(1):25, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ITSb

[Ano22s] Anonymous. IEEE Transactions on Sustainable Computing. Computing in Science and Engineering, 24(3):76, May/ June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ITSc

[Ano22t] Anonymous. IEEE Transactions on Sustainable Computing. Computing in Science and Engineering, 24(4): 53, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:IPCa

[Ano22u] Anonymous. IT Professional:

Call for articles. Computing in Science and Engineering, 24(1):13, January/February 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:CAI

[Ano22v] Anonymous. Call for articles: IEEE Pervasive Computing. Computing in Science and Engineering, 24(2):

6, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:CPIc

[Ano22w] Anonymous. Call for papers: 2023 IEEE Conference on Artificial Intelligence. Computing in Science and Engineering, 24(5):C2, September/October 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:CPIa

[Ano22x] Anonymous. Call for papers: IEEE Computer Society. Computing in Science and Engineering, 24(3):24, May/ June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:CPIb

[Ano22y] Anonymous. Call for papers: IEEE Computer Society. Computing in Science and Engineering, 24(4): 11, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:CPId

[Ano22z] Anonymous. Call for papers: IEEE quantum week. Computing in Science and Engineering, 24(5):32, September/October 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:CSI

[Ano22-27] Anonymous. Computer Society information. Computing in Science and Engineering, 24(5):64, September/October 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:FCa

[Ano22-28] Anonymous. [front cover].

Computing in Science and Engineering, 24(1):C1, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:FCb

[Ano22-29] Anonymous. Front cover.

Computing in Science and Engineering, 24(2):C1, March/ April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:FCc

[Ano22-30] Anonymous. Front cover.

Computing in Science and Engineering, 24(3):C1, May/ June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:FCd

[Ano22-31] Anonymous. Front cover.

Computing in Science and Engineering, 24(4):C1, July/ August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:FCe

[Ano22-32] Anonymous. Front cover. Computing in Science and Engineering, 24(5):C1, September/October 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:FCf

[Ano22-33] Anonymous. Front cover. Computing in Science and Engineering, 24(6):C1, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:GPNa

[Ano22-34] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 24(1):C3, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:GPNb

[Ano22-35] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 24(2): C3, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:GPNc

[Ano22-36] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 24(3):C3, May/ June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:GPNd

[Ano22-37] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 24(4): 54, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:GPNe

[Ano22-38] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 24(5): C3, September/October 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSa

[Ano22-39] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 24(1):4, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSg

[Ano22-40] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 24(2): 73, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSe

[Ano22-41]  Anonymous.  IEEE Com-

puter Society: Call for papers. Computing in Science and Engineering, 24(2): 18, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSr

[Ano22-42]  Anonymous.  IEEE Com-

puter Society call for papers. Computing in Science and Engineering, 24(6):5, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSu

[Ano22-43] Anonymous. IEEE Computer Society career center. Computing in Science and Engineering, 24(6):C3, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSb

[Ano22-44] Anonymous. IEEE Computer Society D&I Fund. Computing in Science and Engineering, 24(1):56, January/February 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSf

[Ano22-45] Anonymous. IEEE Computer Society D&I Fund. Computing in Science and Engineering, 24(2):62, March/April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSi

[Ano22-46] Anonymous. IEEE Computer Society D&I Fund. Computing in Science and Engineering, 24 (3):51, May/June 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSp

[Ano22-47] Anonymous. IEEE Computer Society D&I Fund. Computing in Science and Engineering, 24(4):C4, July/August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSc

[Ano22-48]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 24(1): 71, January/February 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSl

[Ano22-49]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 24(4): C2, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSq

[Ano22-50]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 24(5): 31, September/October 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSt

[Ano22-51]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 24(6):43, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSj

[Ano22-52] Anonymous. IEEE Computer Society information. Computing in Science and Engineering, 24(3):71, May/June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSm

[Ano22-53] Anonymous. IEEE Computer Society information. Computing in Science and Engineering, 24(4):23, July/August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSs

[Ano22-54] Anonymous. IEEE Computer Society information. Computing in Science and Engineering, 24(6):28, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSd

[Ano22-55]  Anonymous.  IEEE Com-

puter Society Jobs Board. Computing in Science and Engineering, 24(1):C4, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSh

[Ano22-56] Anonymous. IEEE Computer Society Jobs Board. Computing in Science and Engineering, 24(2):C4, March/April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSk

[Ano22-57] Anonymous. IEEE Computer Society Jobs Board. Computing in Science and Engineering, 24(3):C4, May/June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSo

[Ano22-58] Anonymous. IEEE Computer Society Jobs Board. Computing in Science and Engineering, 24(4):C3, July/August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:ICSn

[Ano22-59]  Anonymous.  IEEE Com-

puter Society: Publications seek 2024 Editors in Chief. Computing in Science and Engineering, 24(4): 76, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:IQW

[Ano22-60] Anonymous. IEEE Quantum Week. Computing in Science and Engineering, 24(3): C2, May/June 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:Ma

[Ano22-61]  Anonymous.  Masthead.

Computing in Science and Engineering, 24(1):3, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:Mb

[Ano22-62] Anonymous. Masthead. Computing in Science and Engineering, 24(2):1, March/April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:Mc

[Ano22-63] Anonymous. Masthead. Computing in Science and Engineering, 24(4):3, July/August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:Md

[Ano22-64] Anonymous. Masthead. Computing in Science and Engineering, 24(5):3, September/October 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:Me

[Ano22-65] Anonymous. Masthead. Computing in Science and Engineering, 24(6):3, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:RCSa

[Ano22-66] Anonymous. Over the rainbow: 21st Century security & privacy podcast. Computing in Science and Engineering, 24(2):90, March/April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:RCSb

[Ano22-67] Anonymous. Over the rainbow: 21st century security & privacy podcast. Computing in Science and Engineering, 24(6):37, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:PSE

[Ano22-68]  Anonymous.  Publications

seek 2023 Editors in Chief. Computing in Science and Engineering, 24(1):86, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:TCa

[Ano22-69] Anonymous. Table of contents. Computing in Science and Engineering, 24(1):1–2, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:TCb

[Ano22-70] Anonymous. Table of contents. Computing in Science and Engineering, 24(2): 2–3, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:TCc

[Ano22-71] Anonymous. Table of contents. Computing in Science and Engineering, 24(3): 2–3, May/June 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:TCd

[Ano22-72] Anonymous. Table of contents. Computing in Science and Engineering, 24(4): 1–2, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:TCe

[Ano22-73] Anonymous. Table of contents. Computing in Science and Engineering, 24 (5):1–2, September/October 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:TCf

[Ano22-74] Anonymous. Table of contents. Computing in Science and Engineering, 24 (6):1–2, November/December 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2022:WHS

[Ano22-75]  Anonymous.  Watts S.

Humphrey Software Quality Award. Computing in Science and Engineering, 24 (6):76, November/December 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:CEa

[Ano23a] Anonymous. Computing Edge.

Computing in Science and Engineering, 25(1):C4, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:CEb

[Ano23b] Anonymous. Computing Edge. Computing in Science and Engineering, 25(2):C4, February 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:IAH

[Ano23c] Anonymous. IEEE Annals of the History of Computing. Computing in Science and Engineering, 25(2): 68, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICG

[Ano23d]  Anonymous.  IEEE Com-

puter Graphics and Applications. Computing in Science and Engineering, 25(2): 56, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICGb

[Ano23e] Anonymous. IEEE Computer Graphics and Applications. Computing in Science and Engineering, 25(3):49, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICE

[Ano23f]  Anonymous.  IEEE Com-

puting Edge. Computing in Science and Engineering, 25 (3):C4, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:IPC

[Ano23g] Anonymous. IEEE Pervasive Computing. Computing in Science and Engineering, 25(2): 39, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ISP

[Ano23h] Anonymous. IEEE Security & Privacy. Computing in Science and Engineering, 25(2): 29, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ISPb

[Ano23i] Anonymous. IEEE Security & Privacy. Computing in Science and Engineering, 25 (3):78, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ITC

[Ano23j] Anonymous. IEEE Transactions on Computers. Computing in Science and Engineering, 25(3):83, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ITS

[Ano23k] Anonymous. IEEE Transactions on Sustainable Computing. Computing in Science and Engineering, 25(3):28, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:CEc

[Ano23l] Anonymous. Computing edge. Computing in Science and Engineering, 25(4):C4, April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:DDI

[Ano23m] Anonymous. Drive diversity & inclusion in computing. Computing in Science and Engineering, 25(1):4, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:DDIb

[Ano23n] Anonymous. Drive diversity & inclusion in computing. Computing in Science and Engineering, 25(4):C2, April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ECO

[Ano23o] Anonymous. Evolving career opportunities: Explore new options upload your resume today. Computing in Science and Engineering, 25(1): C3, January/February 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:FCa

[Ano23p] Anonymous. Front cover.

Computing in Science and Engineering, 25(1):C1, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:FCb

[Ano23q] Anonymous. Front cover.

Computing in Science and Engineering, 25(2):C1, March/ April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:FCc

[Ano23r] Anonymous. Front cover.

Computing in Science and Engineering, 25(3):C1, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:FCd

[Ano23s] Anonymous. Front cover.

Computing in Science and Engineering, 25(4):C1, April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:GPN

[Ano23t] Anonymous. Get published in the new IEEE Open Journal of the Computer Society. Computing in Science and Engineering, 25(3):67, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSb

[Ano23u] Anonymous. IEEE Computer Society. Computing in Science and Engineering, 25(1): 16, January/February 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSf

[Ano23v]  Anonymous.  IEEE Com-

puter Society — call for papers. Computing in Science and Engineering, 25(2): 5, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSc

[Ano23w]  Anonymous.  IEEE Com-

puter Society call for papers. Computing in Science and Engineering, 25(1):56, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSn

[Ano23x] Anonymous. IEEE Computer Society call for papers. Computing in Science and Engineering, 25(3):41, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSr

[Ano23y] Anonymous. IEEE Computer Society call for papers. Computing in Science and Engineering, 25(4):46, April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSi

[Ano23z]  Anonymous.  IEEE Com-

puter Society career center. Computing in Science and Engineering, 25(2): C3, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSo

[Ano23-27] Anonymous. IEEE Computer Society career center. Computing in Science and Engineering, 25(3):C3, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSs

[Ano23-28] Anonymous. IEEE Computer Society career center. Computing in Science and Engineering, 25(4):C3, April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSe

[Ano23-29] Anonymous. IEEE Computer Society D&I Fund. Computing in Science and Engineering, 25(2):C2, March/April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSj

[Ano23-30] Anonymous. IEEE Computer Society DI fund. Computing in Science and Engineering, 25(3):C2, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSa

[Ano23-31]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 25(1): 15, January/February 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSk

[Ano23-32]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 25(3):7, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSp

[Ano23-33]  Anonymous.  IEEE Com-

puter Society has you covered! Computing in Science and Engineering, 25(4):24, April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSg

[Ano23-34] Anonymous. IEEE Computer Society information. Computing in Science and Engineering, 25(2):40, March/April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSl

[Ano23-35] Anonymous. IEEE Computer Society information. Computing in Science and Engineering, 25(3):17, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSq

[Ano23-36] Anonymous. IEEE Computer Society information. Computing in Science and Engineering, 25(4):34, April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSd

[Ano23-37]  Anonymous.  IEEE Com-

puter Society volunteer service awards. Computing in Science and Engineering, 25(1): 72, January/February 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSh

[Ano23-38]  Anonymous.  IEEE Com-

puter Society volunteer service awards. Computing in Science and Engineering, 25(2):

57, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:ICSm

[Ano23-39]  Anonymous.  IEEE Com-

puter Society volunteer service awards. Computing in Science and Engineering, 25 (3):29, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:IQWa

[Ano23-40] Anonymous. IEEE quantum week. Computing in Science and Engineering, 25(1): C2, January/February 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:IQWb

[Ano23-41]  Anonymous.  IEEE quan-

tum week: Call for participation. Computing in Science and Engineering, 25(1): 51, January/February 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:IQWc

[Ano23-42] Anonymous. IEEE Quantum Week: Register today! Computing in Science and Engineering, 25(2):6, March/April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:Ma

[Ano23-43]  Anonymous.  Masthead.

Computing in Science and Engineering, 25(1):1, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:Mb

[Ano23-44] Anonymous. Masthead. Computing in Science and Engineering, 25(2):1, March/April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:Mc

[Ano23-45] Anonymous. Masthead. Computing in Science and Engineering, 25(3):1, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:Md

[Ano23-46] Anonymous. Masthead. Computing in Science and Engineering, 25(4):1, April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:RCSa

[Ano23-47] Anonymous. Over the rainbow: 21st century security & privacy podcast. Computing in Science and Engineering, 25(1):71, January/February 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:RCSb

[Ano23-48] Anonymous. Over the rainbow: 21st century security & privacy podcast. Computing in Science and Engineering, 25(2):49, March/April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:RCSc

[Ano23-49] Anonymous. Over the rainbow: 21st century security & privacy podcast. Computing in Science and Engineering, 25(3):6, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:PSE

[Ano23-50]  Anonymous.  Publications

seek: 2025 Editors in Chief. Computing in Science and Engineering, 25(3):66, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:TCa

[Ano23-51] Anonymous. Table of contents. Computing in Science and Engineering, 25(1):2–3, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:TCb

[Ano23-52] Anonymous. Table of contents. Computing in Science and Engineering, 25(2):

2–3, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:TCc

[Ano23-53] Anonymous. Table of contents. Computing in Science and Engineering, 25(3): 2–3, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Anonymous:2023:TCd

[Ano23-54] Anonymous. Table of contents. Computing in Science and Engineering, 25 (4):2–3, April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Altintas:2019:TMF

[APC^{+}19] I. Altintas, S. Purawat, D. Crawl, A. Singh, and K. Marcus. Toward a methodology and framework for workflow-driven team science. Computing in Science and Engineering, 21(4):37– 48, July/August 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Anderson:2010:UPS

[APS10] Erik W. Anderson, Gilbert A.

Preston, and Claudio T. Silva. Using Python for signal processing and visualization. Computing in Science and Engineering, 12(4):90– 95, July/August 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Aravas:1999:BWR

[Ara99] Nikolaos Aravas. Book & Web reviews: Finite elements: Theory, fast solvers, and applications in solid mechanics. Computing in Science and Engineering, 1(2): 81–82, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2081.
pdf.

Adorf:2019:HPD

[ARAG19] C. S. Adorf, V. Ramasubramani, J. A. Anderson, and S. C. Glotzer. How to professionally develop reusable scientific software and when not to. Computing in Science and Engineering, 21(2): 66–79, March/April 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Andrade:2011:UFD

[ARO^{+}11] Jose S. Andrade, Jr., Saulo D. S. Reis, Erneson A. Oliveira, Eric Fehr, and Hans J. Herrmann. Ubiquitous fractal dimension of optimal paths. Computing in Science and Engineering, 13 (1):74–81, January/February 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Abram:2022:SVR

[ASLK22] Gregory Abram, Andrew Solis, Yong Liang, and Krishna Kumar. In situ visualization of regional-scale natural hazards with galaxy and material point method. Computing in Science and Engineering, 24(2):31–39, March/April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Agarwal:2014:RER

[ASM^{+}14] Khushbu Agarwal, Poorva Sharma, Jinliang Ma, Chaomei Lo, Ian Gorton, and Yan Liu. Reveal: An extensible reduced-order model builder for simulation and modeling. Computing in Science and Engineering, 16(2):44– 53, March/April 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Asrar:2004:GEI

[Asr04] Ghassem R. Asrar. Guest Editor’s introduction: a pathway to decisions on Earth’s environment and natural resources. Computing in Science and Engineering, 6(1):13–16, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/

                                        cs/2004/01/c1013.pdf;
                                        http://csdl.computer.org/
                                        comp/mags/cs/2004/01/c1013abs.
                                        htm.

Assimakopoulos:2000:ECA

[Ass00] Panayotis A. Assimakopoulos. Education: a computer-aided introductory course in electricity and magnetism. Computing in Science and Engineering, 2(6):88–93, November/December 2000. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c6088.
                                        pdf.

Amengual:2006:TSW

[AT06] Pau Amengual and Raúl Toral. Truels, or survival of the weakest. Computing in Science and Engineering, 8(5): 88–95, September/October 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Alexander:2005:NAR

[ATG05] Francis J. Alexander, Daniel M. Tartakovsky, and Alejandro L. Garcia. Noise in algorithm refinement methods. Computing in Science and Engineering, 7(3): 32–38, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2005/03/c3032abs.htm;

 http://csdl.computer.org/
dl/mags/cs/2005/03/c3032.
pdf.

Alberola:2000:GMP

[ATRA00] Carlos Alberola, Lorenzo Tardón, and Juan Ruiz-Alzola. Graphical models for problem solving. Computing in Science and Engineering, 2(4):46–57, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c4046.
pdf;   http://www.computer.
org/cse/cs1999/c4046abs.
htm.

Ayars:2007:VTI

[Aya07] Eric Ayars. A versatile text for the introductory computational physics course. Computing in Science and Engineering, 9(6):60–61, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/06/mcs2007060060.
pdf.

Ayars:2014:FPB

[Aya14] Eric Ayars. Finally, a Pythonbased computational physics text. Computing in Science and Engineering, 16(1):6–7, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Azooz:2006:ADA

[Azo06] Aasim Azooz. Analog data acquisition for obtaining I-V characteristics using sound cards. Computing in Science and Engineering, 8(4): 10–15, July/August 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Backer:2007:CPE

[Bäc07a]  Arnd Bäcker.  Computa-

tional physics education with Python. Computing in Science and Engineering, 9(3): 30–33, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Backer:2007:QCB

[Bäc07b] Arnd Bäcker. Quantum chaos in billiards. Computing in Science and Engineering, 9(3): 60–64, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bacon:2010:B

[Bac10] Dave Bacon. Books. Computing in Science and Engineering, 12(5):5–7, September/October 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Biswas:2021:FAT

[BAD^{+}21] A. Biswas, J. P. Ahrens, S. Dutta, J. M. Musser, A. S.

Almgren, and T. L. Turton. Feature analysis, tracking, and data reduction: An application to multiphase reactor simulation MFiX-Exa for in-situ use case. Computing in Science and Engineering, 23(1):75–82, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bailey:2000:IRD

[Bai00] David H. Bailey. Integer relation detection. Computing in Science and Engineering, 2 (1):24–28, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1024.
pdf;   http://www.computer.
org/cse/cs1999/c1024abs.
htm.

Bailey:2005:HPF

[Bai05]  David H. Bailey.  High-

precision floating-point arithmetic in scientific computation. Computing in Science and Engineering, 7(3): 54–61, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/03/c3054abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2005/03/c3054.
pdf.

Bakos:2010:HPH

[Bak10]  Jason D. Bakos.  High-

performance heterogeneous computing with the Convey HC-1. Computing in Science and Engineering, 12(6): 80–87, November/December 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Baker:2021:PSI

[Bak21] Allison H. Baker. On preserving scientific integrity for climate model data in the HPC era. Computing in Science and Engineering, 23(6): 16–24, November/December 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ball:1999:CPO

[Bal99] James S. Ball. Computing prescriptions: Orthogonal polynomials, Gaussian quadratures, and PDEs. Computing in Science and Engineering, 1(6):92–95, November/December 1999. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c6092.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c6092abs.
                                        htm.

Balaji:2015:CCS

[Bal15] V. Balaji. Climate computing: The state of play. Computing in Science and Engineering, 17(6):9–13, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Baldwin:2017:HC

[Bal17] Melinda Baldwin. Hidden computers. Computing in Science and Engineering, 19(2): 96, March/April 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/02/mcs2017020096.
html.

Barrett:2011:MMD

[Bar11] Steven F. Barrett. Mentoring and making a difference: What can one person do? Computing in Science and Engineering, 13(1):70–73, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Barrett:2012:BWF

[Bar12] Steven F. Barrett. Book writing: Not for the faint hearted. Computing in Science and Engineering, 14(4): 90–91, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Barba:2019:PRC

[Bar19] L. A. Barba. Praxis of reproducible computational science. Computing in Science and Engineering, 21(1):73–78, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Barba:2020:CSE

[Bar20a]  L. A. Barba.  Computa-

tional science and engineering in 2020. Computing in Science and Engineering, 22 (6):5–7, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Barba:2020:ECR

[Bar20b] L. A. Barba. Engineers code: Reusable open learning modules for engineering computations. Computing in Science and Engineering, 22(4): 26–35, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Barba:2021:TCE

[Bar21a]  L. A. Barba.  Trustwor-

thy computational evidence through transparency and reproducibility. Computing in Science and Engineering, 23 (1):58–64, January/February 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Barba:2021:PJE

[Bar21b]  Lorena A. Barba.  The

Python/Jupyter ecosystem: Today’s problem-solving environment for computational science. Computing in Science and Engineering, 23(3): 5–9, May/June 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Barton:2021:HPC

[Bar21c] Joseph Michael Barton. High performance computing for science and engineering in the Department of Defense. Computing in Science and Engineering, 23(6):58–62, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Basegmez:2002:ESA

[Bas02] Fahri Basegmez. Extending scientific application with scripting capabilities. Computing in Science and Engineering, 4(6):52–59, November/December 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2002/06/c6052.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/06/
c6052.pdf.

Basu:2014:TFE

[Bas14] Prodyot K. Basu. Teaching the finite element method: A sophisticated approach [book review]. Computing in Science and Engineering, 16(5):8–9, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2014/05/mcs2014050008.
                                        html.

Bauke:2008:PML

[Bau08]  Heiko Bauke.  Passing

messages to lonely numbers. Computing in Science and Engineering, 10(2):32– 40, March/April 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Borwein:2001:CMC

[BB01] Jonathan M. Borwein and Peter B. Borwein. Challenges in mathematical computing. Computing in Science and Engineering, 3(3):48–53, May/ June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2001/pdf/c3048.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999c3048abs.
                                        htm.

Bernreuther:2006:FEG

[BB06] Martin Bernreuther and Hans-Joachim Bungartz. First experiences with group projects in CSE education. Computing in Science and Engineering, 8(4): 16–25, July/August 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Banks:2007:FGI

[BB07] David C. Banks and Kevin Beason. Fast global illumination for visualizing isosurfaces with a 3D illumination grid. Computing in Science and Engineering, 9(1):48–54, January/February 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Barriga:2020:AIM

[BB20] N. A. Barriga and F. Besoaín.

Artificial intelligence and mobile programming courses for a video game development program in Chile. Computing in Science and Engineering, 22 (4):17–25, July/August 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Behnel:2011:CBB

[BBC^{+}11] Stefan Behnel, Robert Bradshaw, Craig Citro, Lisandro Dalcin, Dag Sverre Seljebotn, and Kurt Smith. Cython: The best of both worlds. Computing in Science and Engineering, 13(2):31–39, March/April 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bosilca:2013:PEH

[BBD^{+}13] George Bosilca, Aurelien Bouteiller, Anthony Danalis, Mathieu Faverge, Thomas Herault, and Jack J. Dongarra. PaRSEC: Exploiting heterogeneity to enhance scalability. Computing in Science and Engineering, 15(6): 36–45, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bova:2001:PPM

[BBG^{+}01] Steve W. Bova, Clay P. Breshears, Henry Gabb, Bob Kuhn, Bill Magro, Rudolf Eigenmann, Greg Gaertner, Stefano Salvini, and Howard Scott. Parallel programming with message passing and directives. Computing in Science and Engineering, 3(5):22–37, September/October 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c5022abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c5022.pdf.

Blomer:2015:EGS

[BBM^{+}15] Jakob Blomer, Predrag Buncic, Rene Meusel, Gerardo Ganis, Igor Sfiligoi, and Douglas Thain. The evolution of global scale filesystems for scientific software distribution. Computing in Science and Engineering, 17(6):61–71, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Brooks:2021:BCC

[BBM^{+}21] Tessa Durham Brooks, Raychelle Burks, Mark Meysenburg, Erin Doyle, and Chris Huber. Building a culture of computing in the sciences using images as data within a community of practice. Computing in Science and Engineering, 23(5):67– 71, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ballas:2003:UAE

[BBN03] James A. Ballas, Derek Brock, and Justin Nevitt. Using audio to enhance information tasks. Computing in Science and Engineering, 5(5):66–71, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/05/c5066abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/05/c5066.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/05/
c5066.pdf.

Biswas:2020:IET

[BBW^{+}20] A. Biswas, C. M. Biwer, D. J.

Walters, J. Ahrens, D. Francom, E. Lawrence, R. L. Sandberg, D. A. Fredenburg, and C. Bolme. An interactive exploration tool for highdimensional datasets: A shock physics case study. Computing in Science and Engineering, 22(2):44–54, March/April 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Betcke:2022:EUK

[BBWW22] Timo Betcke, Hans-Joachim Bungartz, Garth Wells, and Scott Woodley. ExCALIBUR: U.K.’s preparation for the arrival of the next generation of HPC. Computing in Science and Engineering, 24(1):5–7, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beckmann:1999:GEI

[BC99] Carl J. Beckmann and George Cybenko. Guest Editors’ introduction: Computation in communication. Computing in Science and Engineering, 1(1): 22, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c1022.
                                        pdf.

Bailey:2002:BWR

[BC02] David H. Bailey and George Cybenko. Book and Web reviews: a reclusive kind of science; a new kind of teleology. Computing in Science and Engineering, 4(5): 79–83, September/October 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2002/05/c5079.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/05/
c5079.pdf.

Belloni:2003:PQM

[BC03] Mario Belloni and Wolfgang Christian. Physlets for quantum mechanics. Computing in Science and Engineering, 5(1): 90–96, c3, January/February 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/01/c1090.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/01/
c1090.pdf.

Boghosian:2005:GEIa

[BC05a] B. Boghosian and P. Coveney.

Guest Editors’ introduction: Scientific applications of Grid computing. Computing in Science and Engineering, 7(5):10–13, September/October 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501734.pdf?
isnumber=32219&prod=JNL&
arnumber=1501734&arSt=+
10&ared=+13&arAuthor=+Boghosian%
2C+B.%3B++Coveney%2C+P.;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
32219&arnumber=1501734&
count=14&index=2.

Boghosian:2005:GEIb

[BC05b] Bruce M. Boghosian and Peter V. Coveney. Guest Editors’ introduction: Scientific applications of Grid computing, Part II. Computing in Science and Engineering, 7(6):10–11, November/December 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2005/06/c6010.pdf.

Burnett:2000:SGS

[BCA^{+}00] Toby Burnett, Chris Chaput, Heather Arrighi, Jay Norris, and Daniel J. Suson. Simulating the glast satellite with gismo. Computing in Science and Engineering, 2(4):9–18, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c4009.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c4009abs.
                                        htm.

Belloni:2007:OSP

[BCB07] Mario Belloni, Wolfgang Christian, and Douglas Brown. Open source physics curricular material for quantum mechanics. Computing in Science and Engineering, 9(4): 24–31, July/August 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Burr:1999:TNRa

[BCC^{+}99] Alex F. Burr, Bruce A. Craig, Choong-Geun Chung, Tak-Shing Harry So, Terry J. Deveau, and Donald L. Shirer. Technology news and reviews: Programs for statistics and data analysis. Computing in Science and Engineering, 1 (1):17–21, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c1017.
pdf.

Belletti:2009:JFB

[BCC^{+}09] Francesco Belletti, Maria Cotallo, Andres Cruz, Luis Antonio Fernandez, Antonio Gordillo-Guerrero, Marco Guidetti, Andrea Maiorano, Filippo Mantovani, Enzo Marinari, Victor Martin-Mayor, Antonio Munoz-Sudupe, Denis Navarro, Giorgio Parisi, Sergio Perez-Gaviro, Mauro Rossi, Juan Jesus Ruiz-Lorenzo, Sebastiano Fabio Schifano, Daniele Sciretti, Alfonso Tarancon, Raffaele (lele) Tripiccione, Jose Luis Velasco, David Yllanes, and Gianpaolo Zanier. Janus: An FPGA-based system for high-performance scientific computing. Computing in Science and Engineering, 11 (1):48–58, January/February 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Badia:2022:PIT

[BCE^{+}22] Rosa M. Badia, Javier Conejero, Jorge Ejarque, Daniele Lezzi, and Francesc Lordan. PyCOMPSs as an instrument for translational computer science. Computing in Science and Engineering, 24(2): 79–84, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Booker:1999:VTD

[BCG^{+}99] Andrew Booker, Michelle Condliff, Mark Greaves, Frederick B. Holt, Anne Kao, Daniel J. Pierce, Stephen Poteet, and Yuan-Jye Jason Wu. Visualizing text data sets. Computing in Science and Engineering, 1 (4):26–35, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c4026.
                                        pdf.

Bayoumi:2009:SEC

[BCH^{+}09] Amr Bayoumi, Michael Chu, Yasser Hanafy, Patricia Harrell, and Gamal Refai-Ahmed. Scientific and engineering computing using ATI stream technology. Computing in Science and Engineering, 11(6): 92–97, November/December 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Brase:2022:CHP

[BCH^{+}22] Jim Brase, Nancy Campbell, Barbara Helland, Thuc Hoang, Manish Parashar, Michael Rosenfield, James Sexton, and John Towns. The COVID-19 High-Performance Computing Consortium. Computing in Science and Engineering, 24(1):78–85, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bond:1999:CCC

[BCJK99] J. Richard Bond, Robert G.

Crittenden, Andrew H. Jaffe, and Lloyd Knox. Computing challenges of the cosmic microwave background. Computing in Science and Engineering, 1(2):21–35, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2021.
pdf;   http://www.computer.
org/cse/cs1999/c2021abs.
htm.

Benedetto:2003:ZRI

[BCL03] Dario Benedetto, Emanuele Caglioti, and Vittorio Loreto. Zipping out relevant information. Computing in Science and Engineering, 5 (1):80–85, January/February 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/

                                        dl/mags/cs/2003/01/c1080.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/01/
                                        c1080.pdf.

Bruin:2005:BME

[BDCT05] Richard P. Bruin, Martin T. Dove, Mark Calleja, and Matthew G. Tucker. Building and managing the eMinerals clusters: a case study in Grid-enabled cluster operation. Computing in Science and Engineering, 7(6): 30–37, November/December 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Barkley:2020:HMP

[BDF^{+}20] S. Barkley, T. G. Dimiduk, J. Fung, D. M. Kaz, V. N. Manoharan, R. McGorty, R. W. Perry, and A. Wang. Holographic microscopy with Python and HoloPy. Computing in Science and Engineering, 22(5):72–82, September/October 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Budavari:2013:SFA

[BDS13] Tamas Budavari, Laszlo Dobos, and Alexander S. Szalay. SkyQuery: Federating astronomy archives. Computing in Science and Engineering, 15(3):12–20, May/ June 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bean:2000:CIM

[Bea00] James C. Bean. CSE in industry: a multiple-choice genetic algorithm for a nonlinear cutting stock problem. Computing in Science and Engineering, 2(2):80–83, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c2080.
pdf;   http://www.computer.
org/cse/cs1999/c2080abs.
htm.

Beckman:2015:ODF

[Bec15] Pete Beckman. Our dynamic future. Computing in Science and Engineering, 17(6): 88–90, November/December 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Behrens:2005:AAM

[Beh05] J. Behrens. Adaptive atmospheric modeling: scientific computing at its best. Computing in Science and Engineering, 7(4): 76–83, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463140.pdf?
isnumber=31456&prod=JNL&
arnumber=1463140&arSt=+
76&ared=+83&arAuthor=Behrens%
2C+J.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=31456&arnumber=

                                        1463140&count=14&index=
                                        10.

Beichl:2002:DDT

[Bei02]  Isabel Beichl.  Dealing

with degeneracy in triangulation. Computing in Science and Engineering, 4(6): 70–74, November/December 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2002/06/c6070.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/06/
                                        c6070.pdf.

Beichl:2009:BB

[Bei09a] Isabel Beichl. Begin at the beginning. Computing in Science and Engineering, 11(1):3, January/February 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2009:WTC

[Bei09b] Isabel Beichl. We’ll take care of all your problems. Computing in Science and Engineering, 11(2):3–4, March/April 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2009:WRM

[Bei09c] Isabel Beichl. Where the rubber meets the road. Computing in Science and Engineering, 11(5):3, September/October 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2010:D

[Bei10a] Isabel Beichl. Dystopia. Computing in Science and Engineering, 12(6):4, November/December 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2010:E

[Bei10b]  Isabel Beichl.  Ephemera.

Computing in Science and Engineering, 12(1):3, January/February 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2010:GPM

[Bei10c] Isabel Beichl. Good policy makes good science. Computing in Science and Engineering, 12(3):5, May/June 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2010:RWC

[Bei10d] Isabel Beichl. Revitalizing work in CiSE. Computing in Science and Engineering, 12(5):4, September/October 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2010:YSY

[Bei10e] Isabel Beichl. You say you want a revolution. Computing in Science and Engineering, 12 (2):5, March/April 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2011:C

[Bei11a]  Isabel Beichl.  Changes.

Computing in Science and Engineering, 13(1):4, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2011:JT

[Bei11b] Isabel Beichl. Just try. Computing in Science and Engineering, 13(6):4–5, November/December 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2011:TC

[Bei11c] Isabel Beichl. Top of the charts. Computing in Science and Engineering, 13(5): 4, September/October 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2011:WW

[Bei11d] Isabel Beichl. Watchful waiting. Computing in Science and Engineering, 13(2): 3–4, March/April 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2012:CTS

[Bei12a] Isabel Beichl. Counting on today’s students. Computing in Science and Engineering, 14(5):4, September/October 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2012:HF

[Bei12b]  Isabel Beichl.  Hail and

farewell. Computing in Science and Engineering, 14(6): 4, November/December 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2012:MCS

[Bei12c] Isabel Beichl. Is my car smarter than my cat? Computing in Science and Engineering, 14(3):4–5, May/June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2012:ACE

[Bei12d] Isabel Beichl. Of art and CS education. Computing in Science and Engineering, 14(1):4, January/February 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bennett:2000:WMM

[Ben00] Chuck Bennett. Web mechanics: Multipart homework problems over the Web. Computing in Science and Engineering, 2(1):87–91, 95– 95, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1087.
pdf;   http://www.computer.
org/cse/cs1999/c1087abs.
htm.

Benjamin:2004:EPV

[Ben04] Robert F. Benjamin. An experimenter’s perspective on validating codes and models with experiments having shock-accelerated fluid interfaces. Computing in Science and Engineering, 6(5): 40–49, September/October 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/05/c5040.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/05/
c5040.pdf.

Benger:2009:SFF

[Ben09] Werner Benger. On safari in the file format jungle — why can’t you visualize my data? Computing in Science and Engineering, 11(6): 98–102, November/December 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Berry:1999:GEI

[Ber99] Michael W. Berry. Guest Editor’s introduction: Massive data visualization. Computing in Science and Engineering, 1(4):16–17, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c4016.
pdf.

Bungartz:2009:CSE

[BERT09] Hans-Joachim Bungartz, Donald Estep, Ulrich Rude, and Peter Turner. Computational science and engineering education: SIAM’s perspective. Computing in Science and Engineering, 11(6):5–11, November/December 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Betts:1999:CID

[Bet99] John T. Betts. CSE in industry: a direct approach to solving optimal control problems. Computing in Science and Engineering, 1(3):73–75, May/ June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3073.
pdf;   http://www.computer.
org/cse/cs1999/c3073abs.
htm.

Bethea:2017:ESS

[Bet17] Katie Bethea. Earth scientists shake things up on Titan. Computing in Science and Engineering, 19(4):68– 71, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/04/mcs2017040068.
                                        html.

Bezakova:2008:SBC

[Bez08]  Ivona Bezáková.  Sam-

pling binary contingency tables. Computing in Science and Engineering, 10(2):26– 31, March/April 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Baladron:2012:TAG

[BFF12] Javier Baladron, Diego Fasoli, and Olivier Faugeras. Three applications of GPU computing in neuroscience. Computing in Science and Engineering, 14(3):40–47, May/ June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Betteridge:2021:CGP

[BFH21] Jack D. Betteridge, Patrick E. Farrell, and David A. Ham. Code generation for productive, portable, and scalable finite element simulation in Firedrake. Computing in Science and Engineering, 23(4): 8–17, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Balogh:2022:SMC

[BFL^{+}22] Gábor D. Balogh, Tobias S.

Flynn, Sylvain Laizet, Gihan R. Mudalige, and István Z. Reguly. Scalable manycore algorithms for tridiagonal solvers. Computing in Science and Engineering, 24(1):26–35, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bertino:2004:TNC

[BFS04] Elisa Bertino, Elena Ferrari, and Anna Squicciarini. Trust negotiations: Concepts, systems, and languages. Computing in Science and Engineering, 6(4):27–34, July/August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/04/c4027.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/04/
c4027.pdf.

Bergen:2006:MPM

[BGHR06] Benjamin Bergen, Tobias Gradl, Frank Hülsemann, and Ulrich Rüde. A massively parallel multigrid method for finite elements. Computing in Science and Engineering, 8(6): 56–62, November/December 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Best:2002:VIC

[BH02] Christoph Best and Hans-Christian Hege. Visualizing and identifying conformational ensembles in molecular dynamics trajectories. Computing in Science and Engineering, 4(3): 68–75, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/03/c3068abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/03/c3068.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/03/
c3068.pdf.

Bernardon:2008:TFS

[BHC^{+}08] Fabio F. Bernardon, Linh K. Ha, Steven P. Callahan, Joao L. D. Comba, and Claudio T. Silva. Transfer-function specification for rendering disparate volumes. Computing in Science and Engineering, 10(6):82–89, November/December 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Brook:2015:BED

[BHC^{+}15] R. Glenn Brook, Alexander Heinecke, Anthony B. Costa, Paul Peltz, Vincent C. Betro, Troy Baer, Michael Bader, and Pradeep Dubey. Beacon: Exploring the deployment and application of Intel Xeon Phi coprocessors for scientific computing. Computing in Science and Engineering, 17(2): 1, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/02/mcs2015020065-
abs.html.

Baptista:2008:SEE

[BHF^{+}08] António Baptista, Bill Howe, Juliana Freire, David Maier, and Cláudio T. Silva. Scientific exploration in the era of ocean observatories. Computing in Science and Engineering, 10(3):53–58, May/ June 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Binder:2003:IBS

[BHKW03] Kurt Binder, Jürgen Horbach, Walter Kob, and Anke Winkler. The interplay between structure and ionic motions in glasses. Computing in Science and Engineering, 5(2): 60–66, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/02/c2060abs.htm;
http://csdl.computer.org/
dl/mags/cs/2003/02/c2060.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/02/
c2060.pdf.

Bertsimas:1999:OCE

[BHL99] Dimitris Bertsimas, Paul Hummell, and Andrew W. Lo. Optimal control of execution costs for portfolios. Computing in Science and Engineering, 1(6): 40–53, November/December 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c6040.
pdf;   http://www.computer.
org/cse/cs1999/c6040abs.
htm.

Brown:2021:EYE

[BHV21] David Brown, James Hack, and Robert Voigt. The early years and evolution of the DOE Computational Science Graduate Fellowship Program. Computing in Science and Engineering, 23(6): 9–15, November/December 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bilar:2000:BWR

[Bil00] Daniel Bilar. Book and Web reviews: To see the world in a grain of sand. Computing in Science and Engineering, 2(2):104, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c2104.
pdf;   http://www.computer.
org/cse/cs1999/c2104abs.
htm.

Bizot:2016:RWP

[Biz16] Elizabeth B. Bizot. Representation of women in postsecondary computing: Disciplinary, institutional, and individual characteristics. Computing in Science and Engineering, 18(2):40– 56, March/April 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bahn:2002:OOS

[BJ02] Sune R. Bahn and Karsten W.

Jacobsen. An object-oriented scripting interface to a legacy electronic structure code. Computing in Science and Engineering, 4(3):56–66, May/ June 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2002/03/c3056abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/03/
c3056.htm;   http://csdl.
computer.org/dl/mags/cs/
2002/03/c3056.pdf.

Bana:2020:OKD

[BKB20] K. Bana , F. Kru el, and J. Biela ski. Optimal kernel design for finiteelement numerical integration on GPUs. Computing in Science and Engineering, 22(6): 61–74, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bozzola:2022:BHP

[BkCP22] Gabriele Bozzola, Chi kwan Chan, and Vasileios Paschalidis. Black hole physics and computer graphics. Computing in Science and Engineering, 24(2):19–30, March/April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Borrill:2015:BBB

[BKK15] Julian Borrill, Reijo Keskitalo, and Theodore Kisner. Big bang, big data, big iron: Fifteen years of cosmic microwave background data analysis at NERSC. Computing in Science and Engineering, 17(3):22–29, May/ June 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //csdl.computer.org/csdl/
                                        mags/cs/2015/03/mcs2015030022-
                                        abs.html.

Barba:2021:SCP

[BKRT21] Lorena A. Barba, Andreas Klöckner, Prabhu Ramachandran, and Rollin Thomas. Scientific computing with Python on high-performance heterogeneous systems. Computing in Science and Engineering, 23(4):5–7, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Brown:2015:RTE

[BKS15] Jed Brown, Matthew G. Knepley, and Barry F. Smith.

Run-time extensibility and librarization of simulation software. Computing in Science and Engineering, 17(1):38–45, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/01/mcs2015010038-
abs.html.

Blilie:2002:PSS

[Bli02] Charles Blilie. Patterns in scientific software: An introduction. Computing in Science and Engineering, 4(3): 48–53, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/03/c3048abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/03/c3048.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/03/
c3048.pdf.

Bardhan:2021:EID

[BLMR21] Jaydeep Bardhan, Mary Ann Leung, Eileen Martin, and Amanda Randles. 2Guest Editor’s introduction: DOE Computational Science Graduate Fellowship Research Showcase. Computing in Science and Engineering, 23(6):5– 8, November/December 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Balos:2022:COL

[BLO^{+}22] Cody J. Balos, Piotr Luszczek, Sarah Osborn, James Willenbring, and Ulrike M. Yang. Challenges of and opportunities for a large diverse software team. Computing in Science and Engineering, 24(3): 16–24, May/June 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bauer:2021:SPL

[BLP^{+}21] Michael Bauer, Wonchan Lee, Manolis Papadakis, Marcin Zalewski, and Michael Garland. Supercomputing in Python with Legate. Computing in Science and Engineering, 23(4):73–79, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Belcaid:2023:RSA

[BLT^{+}23] Mahdi Belcaid, Jason Leigh, Ryan Theriot, Nurit Kirshenbaum, Roderick Tabalba, Michael Rogers, Andrew Johnson, Maxine Brown, Luc Renambot, Lance Long, Arthur Nishimoto, Chris North, and Jesse Harden. Reflecting on the scalable adaptive graphics environment team’s 20-year translational research endeavor in digital collaboration tools. Computing in Science and Engineering, 25(2):50–56, March/April 2023. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Beckmann:1999:HSN

[BMC99] Carl J. Beckmann, Donald D.

McManus, and George Cybenko. Horizons in scientific networking and computing. Computing in Science and Engineering, 1(1):23– 30, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c1023.
pdf;   http://www.computer.
org/cse/cs1999/c1023abs.
htm.

Balasubramanian:2021:TIE

[BMCC21]  Ganesh  Balasubramanian,

Joydeep Munshi, Wei Chen, and TeYu Chien. Towards improving the efficiency of organic solar cells by coarsegrained atomistic modeling of processing dependent morphologies. Computing in Science and Engineering, 23(3): 48–55, May/June 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Belletti:2006:IAF

[BMP^{+}06] Francesco Belletti, Filippo Mantovani, Giorgio Poli, Sebastiano Fabio Schifano, R. Tripiccione, Isabel Campos, Andres Cruz Flor, Denis Navarro, Sergio Perez Gaviro, Daniele Sciretti, Alfonso Tarancon, José Luis Velasco, Pedro Tellez, Luis Antonio Fernandez, Victor Martin-Mayor, Antonio Munoz Sudupe, Sergio Jimenez, Andrea Maiorano, Enzo Marinari, and Juan Jesus Ruiz-Lorenzo. Ianus: An adaptive FPGA computer. Computing in Science and Engineering, 8 (1):41–49, January/February 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Binder:1999:HSC

[BMS99] Kurt Binder, Marcus Muller, and Friedericke Schmid. How simulations clarify complex material phase transitions. Computing in Science and Engineering, 1(3):10–12, May/ June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3010.
pdf;   http://www.computer.
org/cse/cs1999/c3010abs.
htm.

Bernholc:2004:UEP

[BNNM04] Jerry Bernholc, Serge M. Nakhmanson, Marco Buongiorno Nardelli, and Vincent Meunier. Understanding and enhancing polarization in complex materials. Computing in Science and Engineering, 6(6): 12–21, November/December 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/06/c6012.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/06/
c6012.pdf.

Buchen-Osmond:2003:UVD

[BO03]  Cornelia  Büchen-Osmond.

The Universal Virus Database ICTVdB. Computing in Science and Engineering, 5(3): 16–25, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/03/c3016abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/03/c3016.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/03/
c3016.pdf.

Boudriga:2004:IAW

[BO04] Noureddine Boudriga and Mohammad S. Obaidat. Intelligent agents on the Web: a review. Computing in Science and Engineering, 6 (4):35–42, July/August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/04/c4035.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/04/
c4035.pdf.

Boahen:2017:NP

[Boa17] Kwabena Boahen. A neuromorph’s prospectus. Computing in Science and Engineering, 19(2):14–28, March/April 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cs/2017/02/mcs2017020014-
                                        abs.html.

Boettcher:2000:CSE

[Boe00] Stefan Boettcher. Computer simulations: Extremal optimization:heuristics via coevolutionary avalanches. Computing in Science and Engineering, 2(6):75–82, November/December 2000. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c6075.
                                        pdf;   pdf/c6075.pdf.

Boghosian:2003:LLB

[Bog03] Bruce M. Boghosian. A look at lattice Boltzmann equations. Computing in Science and Engineering, 5(2): 86–87, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/02/c2086abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/02/c2086.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/02/
                                        c2086.pdf.

Boghosian:2005:CCC

[Bog05] Bruce Boghosian. A crash course on computing. Computing in Science and Engineering, 7(1):17–20, January/February 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/01/c1017.pdf;
http://csdl.computer.org/
dl/mags/cs/2005/01/c1017.
htm.

Boris:2002:TCB

[Bor02] Jay Boris. The threat of chemical and biological terrorism: Roles for HPC in preparing a response. Computing in Science and Engineering, 4 (2):22–32, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c2022abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c2022.pdf.

Beichl:2007:MCM

[BOS07] Isabel Beichl, Dianne P.

O’Leary, and Francis Sullivan. Monte Carlo minimization and counting: One, two, …, too many. Computing in Science and Engineering, 9 (1):72–80, January/February 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Botchway:2016:ANR

[Bot16] Portia K. Botchway. Addressing negative racial and gendered experiences that discourage academic careers in engineering. Computing in Science and Engineering, 18 (2):29–39, March/April 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bouchaud:1999:FDM

[BP99] Elisabeth Bouchaud and Florin Paun. Fracture and damage at a microstructural scale. Computing in Science and Engineering, 1(5): 32–38, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c5032.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c5032abs.
                                        htm.

Barrett:2010:T

[BP10] Steven Barrett and Daniel Pack. Textbooks 101. Computing in Science and Engineering, 12(2):72–77, March/ April 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beckvermit:2013:MMA

[BPH^{+}13] Jacqueline Beckvermit, Joseph Peterson, Todd Harman, Scott Bardenhagen, Charles Wight, Qingyu Meng, and Martin Berzins. Multiscale modeling of accidental explosions and detonations. Computing in Science and Engineering, 15 (4):76–86, July/August 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bergonzi:2021:DTM

[BPMKC21] M. Bergonzi, E. Pecker-Marcosig, E. Kofman, and R. Castro. Discrete-time modeling of COVID-19 propagation in Argentina with explicit delays. Computing in Science and Engineering, 23(1):35–45, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Barnes:2020:FRE

[BPW^{+}20] T. Barnes, J. Payton, N. Washington, F. Stukes, A. Peterfreund, and S. Dunton. Featured research on equity and sustained participation in engineering, computing, and technology. Computing in Science and Engineering, 22 (5):4–6, September/October 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Breaux:2017:CBD

[Bre17] Justin H. S. Breaux. Cancer’s big data problem. Computing in Science and Engineering, 19 (2):79–81, March/April 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2017/02/mcs2017020079.
html.

Brinkerhoff:2023:CFE

[Bri23] Douglas J. Brinkerhoff. Compatible finite elements for glacier modeling. Computing in Science and Engineering, 25 (3):18–28, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Brown:2006:SQW

[Bro06] James Brown. Seeing into the quantum world. Computing in Science and Engineering, 8(3): 44–45, May/June 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2006/03/c3044.pdf.

Bower:2022:MPP

[BRS22] Richard Bower, Benedict D. Rogers, and Matthieu Schaller. Massively parallel particle hydrodynamics at exascale. Computing in Science and Engineering, 24(1):14–25, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bryan:1999:FUA

[Bry99]  Greg L. Bryan.  Fluids

in the universe: Adaptive mesh refinement in cosmology. Computing in Science and Engineering, 1(2): 46–53, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2046.
pdf;   http://www.computer.
org/cse/cs1999/c2046abs.
htm.

Bryant:2011:DIS

[Bry11] Randal E. Bryant. Dataintensive scalable computing for scientific applications. Computing in Science and Engineering, 13(6): 25–33, November/December 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:1999:CPP

[BS99a] Isabel Beichl and Francis Sullivan. Computing prescriptions: Pay me now or pay me later. Computing in Science and Engineering, 1 (4):59–62, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c4059.
pdf.

Beichl:1999:CPI

[BS99b] Isabel Beichl and Francis Sullivan. Computing prescriptions: The importance of importance sampling. Computing in Science and Engineering, 1(2):71–73, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/

                                        books/cs1999/pdf/c2071.
                                        pdf.

Beichl:2000:CP

[BS00a] Isabel Beichl and Francis Sullivan. Computing prescriptions: A = B? Computing in Science and Engineering, 2(3):84–87, May/ June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c3084.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c3084abs.
                                        htm.

Beichl:2000:CPD

[BS00b] Isabel Beichl and Francis Sullivan. Computing prescriptions: Determining the derminant. Computing in Science and Engineering, 2(5): 63–66, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c5063.
                                        pdf.

Beichl:2000:MA

[BS00c] Isabel Beichl and Francis Sullivan. The Metropolis algorithm. Computing in Science and Engineering, 2 (1):65–69, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c1065.

pdf;   http://www.computer.
org/cse/cs1999/c1065abs.
htm.

Board:2000:FMA

[BS00d] John Board and Klaus Schulten. The fast multipole algorithm. Computing in Science and Engineering, 2 (1):76–79, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1076.
pdf;   http://www.computer.
org/cse/cs1999/c1076abs.

htm. See correspondence [MBS^{+}00].

Beichl:2002:BR

[BS02] Isabel Beichl and Francis Sullivan. It’s bound to be right. Computing in Science and Engineering, 4(2): 86–89, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c2086abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c2086.pdf.

Beichl:2006:GEI

[BS06a] Isabel Beichl and Francis Sullivan. Guest Editors’ introduction: Monte Carlo methods. Computing in Science and Engineering, 8(2): 7–8, March/April 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2006/02/c2007.pdf.

Beichl:2006:OMC

[BS06b] Isabel Beichl and Francis Sullivan. The other Monte Carlo method. Computing in Science and Engineering, 8(2): 42–47, March/April 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beichl:2008:CC

[BS08] Isabel Beichl and Francis Sullivan. Combinatorics in computing. Computing in Science and Engineering, 10(2): 12–13, March/April 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2008/02/mcs2008020012.
                                        pdf.

Beichl:2009:CI

[BS09] Isabel Beichl and Francis Sullivan. Cut it out! Computing in Science and Engineering, 11(3):74–79, May/ June 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Betcke:2021:DHP

[BS21] Timo Betcke and Matthew W. Scroggs. Designing a highperformance boundary element library with OpenCL and Numba. Computing in Science and Engineering, 23 (4):18–28, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Berzins:2007:RPC

[BSD07] Martin Berzins, Bradley A.

Shadwick, and Denis Donnelly. On the role and place of computation in science and engineering. Computing in Science and Engineering, 9 (1):98–103, January/February 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Belletti:2006:CLA

[BST^{+}06] Francesco Belletti, Sebastiano Fabio Schifano, Raffaele Tripiccione, François Bodin, Philippe Boucaud, Jacques Micheli, Olivier Pène, France Nicola Cabibbo, Sergio de Luca, Alessandro Lonardo, Davide Rossetti, Piero Vicini, Maxim Lukyanov, Laurent Morin, Norbert Paschedag, Hubert Simma, Vincent Morenas, Dirk Pleiter, and Federico Rapuano. Computing for LQCD: apeNEXT. Computing in Science and Engineering, 8 (1):18–29, January/February 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bai:2013:SLA

[BST^{+}13] Mingze Bai, Shixin Sun, Hong Tang, Yusheng Dou, and Glenn V. Lo. An SPMDlike algorithm for parallelizing molecular dynamics using OpenMP. Computing in Science and Engineering, 15(4): 48–56, July/August 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Black:2001:CPB

[BT01] Timothy C. Black and William J. Thompson. Computing prescriptions: Bayesian data analysis. Computing in Science and Engineering, 3(4): 86–91, July/August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2001/pdf/c4086.
                                        pdf.

Barrett:2010:PPS

[BT10a] Steven F. Barrett and Mitchell A. Thornton. To PE or not to PE: the sequel. Computing in Science and Engineering, 12 (4):62–65, July/August 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bensky:2010:CGS

[BT10b] Thomas J. Bensky and Catherine A. Taff. Computerguided solutions to physics problems using Prolog. Computing in Science and Engineering, 12(1):88–95, January/February 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Barba:2017:RRC

[BT17] Lorena A. Barba and George K.

Thiruvathukal. Reproducible research for computing in science & engineering. Computing in Science and Engineering, 19(6):85–87, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060085.
html.

Beg:2021:UJR

[BTK^{+}21] M. Beg, J. Taka, T. Kluyver, A. Konovalov, M. Ragan-Kelley, N. M. Thiéry, and H. Fangohr. Using Jupyter for reproducible scientific workflows. Computing in Science and Engineering, 23(2): 36–46, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Barrett:2019:ART

[BTL19] J. Barrett, V. E. Taylor, and M. A. Leung. The ACM Richard Tapia Celebration of Diversity in Computing Conferences, Presented by CMD-IT. Computing in Science and Engineering, 21(4): 88–92, July/August 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Burr:1999:TNRb

[Bur99] Alex F. Burr. Technology news & reviews: Mathcad 8: New features for a popular math program. Computing in Science and Engineering, 1(4):6–9, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c4006.
pdf.

Burt:2018:EGD

[Bur18] Timothy C. Burt. From experts guiding data to data guiding solutions. Computing in Science and Engineering, 20 (1):76–78, January/February 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bartlett:2021:HPP

[BUS21] John Bartlett, Chris Uchytil, and Duane Storti. Highproductivity parallelism with Python plus packages (but without a cluster). Computing in Science and Engineering, 23(4):38–46, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Butikov:1999:EPR

[But99] Eugene I. Butikov. Education: Parametric resonance. Computing in Science and Engineering, 1(3):76–83, May/ June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/

books/cs1999/pdf/c3076.
pdf;   http://www.computer.
org/cse/cs1999/c3076abs.
htm.

Bienstman:2007:PNR

[BVB^{+}07] Peter Bienstman, Lieven Vanholme, Wim Bogaerts, Pieter Dumon, and Peter Vandersteegen. Python in nanophotonics research. Computing in Science and Engineering, 9(3): 46–47, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Brown:2021:RGF

[BVT^{+}21] D. A. Brown, K. Vahi, M. Taufer, V. Welch, and E. Deelman. Reproducing GW150914: The first observation of gravitational waves from a binary black hole merger. Computing in Science and Engineering, 23(2): 73–82, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bradley:2001:ITL

[BW01] Geoff Bradley and Denis Weaire. Instabilities of two liquid drops in contact. Computing in Science and Engineering, 3(5):16–21, September/October 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c5016abs.htm;   http://

                                        dlib.computer.org/cs/books/
                                        cs2001/pdf/c5016.pdf.

Bastian:2006:MMA

[BW06] Peter Bastian and Christian Wieners. Multigrid methods on adaptively refined grids. Computing in Science and Engineering, 8(6):44–54, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Barrett:2010:SS

[BW10] Steven Barrett and Cameron H. G. Wright. For students by students. Computing in Science and Engineering, 12(6):70–73, November/December 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Betz:2014:SDM

[BW14] Robin M. Betz and Ross C. Walker. Streamlining development of a multimillionline computational chemistry code. Computing in Science and Engineering, 16(3): 10–17, May/June 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beazley:2001:ISS

[BWC01] David M. Beazley, Brian D. Ward, and Ian R. Cooke. The inside story on shared libraries and dynamic loading. Computing in Science and Engineering, 3(5):90– 97, September/October 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c5090abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c5090.pdf.

Bhatt:2007:EGM

[BZL^{+}07] Uma S. Bhatt, Jing Zhang, Craig S. Lingle, Lisa M. Phillips, and Wendell V. Tangborn. Examining glacier mass balances with a hierarchical modeling approach. Computing in Science and Engineering, 9(2):60–67, March/ April 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Canessa:1999:WME

[Can99]  Enrique Canessa.  Web

mechanics: Enabling synchronous math discussions on the Web. Computing in Science and Engineering, 1(5): 74–76, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5074.
pdf;   http://www.computer.
org/cse/cs1999/c5074abs.
htm.

Campbell:2010:IMB

[CAP^{+}10] Tim Campbell, Richard Allard, Ruth Preller, Lucy Smedstad, Alan Wallcraft, Sue Chen, Hao Jin, Sasa Gabersek, Richard Hodur, and Joseph Reich. Integrated modeling of the battlespace environment. Computing in Science and Engineering, 12(5):36–45, September/October 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Carver:2009:FIW

[Car09a] Jeffrey C. Carver. First International Workshop on Software Engineering for Computational Science & Engineering. Computing in Science and Engineering, 11(2): 7–11, March/April 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Carver:2009:RSI

[Car09b]  Jeffrey C. Carver.  Re-

port: The Second International Workshop on Software Engineering for CSE. Computing in Science and Engineering, 11(6):14–19, November/December 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Carver:2012:SEC

[Car12]  Jeffrey C. Carver.  Soft-

ware engineering for computational science and engineering. Computing in Science and Engineering, 14(2): 8–11, March/April 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Carver:2016:SES

[Car16] Jeffrey C. Carver. Software engineering for science. Computing in Science and Engineering, 18(2):4–5, March/April 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cickovski:2007:GOC

[CAS^{+}07] Trevor Cickovski, Kedar Aras, Maciej Swat, Roeland M. H. Merks, Tilmann Glimm, H. George E. Hentschel, Mark S. Alber, James A. Glazier, Stuart A. Newman, and Jesus A. Izaguirre. From genes to organisms via the cell: a problem-solving environment for multicellular development. Computing in Science and Engineering, 9(4): 50–60, July/August 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Castro:2016:DEM

[Cas16]  Rodrigo Castro.  Discrete

event modeling and simulationdriven engineering for the AT-LAS data acquisition network. Computing in Science and Engineering, 18(3):70–83, May/ June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chetty:2002:WCG

[CB02] Madhu Chetty and Rajkumar Buyya. Weaving computational Grids: How analogous are they with electrical grids? Computing in Science and Engineering, 4(4): 61–71, July/August 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/04/c4061abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/04/c4061.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/04/
c4061.pdf.

Christian:2006:OSX

[CBB06] Wolfgang Christian, Mario Belloni, and Douglas Brown. An open-source XML framework for authoring curricular material. Computing in Science and Engineering, 8(5): 51–58, September/October 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Catlett:2022:HCS

[CBF^{+}22] Charlie Catlett, Pete Beckman, Nicola Ferrier, Michael E. Papka, Rajesh Sankaran, Jeff Solin, Valerie Taylor, Douglas Pancoast, and Daniel Reed. Hands-on computer science: The Array of Things experimental urban instrument. Computing in Science and Engineering, 24(1):57–63, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Colombo:2014:NST

[CBS14] Tiago C. A. Colombo, Alberto M. G. Brito, and Lirio Schaeffer. Numerical simulation of thermomechanical processes coupled with microstructure evolution. Computing in Science and Engineering, 16(2): 10–15, March/April 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Coisson:1999:TNR

[CC99] Robert Coïsson and Luigi Cristofolini. Technology news and reviews: Real-time toolbox provides convenient labdata acquisition. Computing in Science and Engineering, 1(6):10–13, November/December 1999. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c6010.
pdf;   http://www.computer.
org/cse/cs1999/c6010abs.
htm.

Chonacky:2003:SED

[CC03] Norman Chonacky and Dante Choi. Science and engineering databases in an open-source software world. Computing in Science and Engineering, 5(3): 10–13, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/03/c3010abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/03/c3010.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/03/
                                        c3010.pdf.

Cary:2022:RVC

[CCD^{+}22] Andrew Cary, John Chawner, Earl Duque, William Gropp, Bil Kleb, Ray Kolonay, Eric Nielsen, and Brian Smith. Realizing the vision of CFD in 2030. Computing in Science and Engineering, 24(1):64–70, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chin:2004:CSG

[CCJ04] Jonathan Chin, Peter V. Coveney, and Shantenu Jha. The current state of the grid. Computing in Science and Engineering, 6(5):75–77, September/October 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/05/c5075.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/cs/2004/05/c5075.
                                        htm.

Candela:2023:DEV

[CCP23] Leonardo Candela, Donatella Castelli, and Pasquale Pagano. The D4Science experience on virtual research environment development. Computing in Science and Engineering, 25 (2):12–19, March/April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Choi:2012:DFT

[CCPS12] Woosong Choi, Yong S. Chen, Stefanos Papanikolaou, and James P. Sethna. Is dislocation flow turbulent in deformed crystals? Computing in Science and Engineering, 14 (1):33–39, January/February 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Callahan:2008:DVR

[CCSS08] Steven P. Callahan, Jason H.

Callahan, Carlos E. Scheidegger, and Claudio T. Silva. Direct volume rendering: a 3D plotting technique for scientific data. Computing in Science and Engineering, 10 (1):88–92, January/February 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Calder:2004:VAS

[CDF^{+}04] Alan Calder, Jonathan Dursi, Bruce Fryxell, Tomek Plewa, Greg Weirs, Todd Dupont, Harry Robey, Jave Kane, Bruce Remington, Frank Timmes, Guy Dimonte, John Hayes, Mike Zingale, Paul Drake, Paul Ricker, Jim Stone, and Kevin Olson. Validating astrophysical simulation codes. Computing in Science and Engineering, 6(5): 10–20, September/October 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/05/c5010.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/05/
c5010.pdf.

Croft:2015:PCS

[CDKF15] Rupert Croft, Tiziana Di Matteo, Nishikanta Khandai, and Yu Feng. Petascale cosmology: Simulations of structure formation. Computing in Science and Engineering, 17(2): 40–46, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/02/mcs2015020040-
abs.html.

Carns:2023:MCS

[CDL^{+}23] Philip Carns, Matthieu Dorier, Rob Latham, Robert B. Ross, Shane Snyder, and Jerome Soumagne. Mochi: a case study in translational computer science for highperformance computing data management. Computing in Science and Engineering, 25 (4):35–41, April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Carver:2014:SEC

[CE14] Jeffrey C. Carver and Tom Epperly. Software engineering for computational science and engineering [Guest Editors’ introduction]. Computing in Science and Engineering, 16 (3):6–9, May/June 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cruz-Enriquez:2018:MSA

[CE18] Hector Cruz-Enriquez. Making supercomputing available to all Cuban researchers. Computing in Science and Engineering, 20(3):25–30, May/ June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8358046/.

Cazes:2021:PFT

[CED^{+}21] John Cazes, Richard Todd Evans, Aaron Dubrow, Lei Huang, Si Liu, and Robert McLay. Preparing Frontera for Texascale Days. Computing in Science and Engineering, 23(3):42–47, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Carver:2022:SIFa

[CENT22a] Jeffrey C. Carver, Nasir Eisty, Hai Ah Nam, and Irina Tezaur. Special issue on the future of research software engineers in the United States. Part I. Computing in Science and Engineering, 24 (5):4–5, September/October 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Carver:2022:SIFb

[CENT22b] Jeffrey C. Carver, Nasir Eisty, Hai Ah Nam, and Irina Tezaur. Special issue on the future of research software engineers in the United States part II. Computing in Science and Engineering, 24 (6):4–5, November/December 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chen:1999:VCA

[CF99a] Jim X. Chen and Ophir Frieder. Visualization corner: Applications of computer graphics software tools. Computing in Science and Engineering, 1(6):82–87, November/December 1999. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c6082.
pdf;   http://www.computer.
org/cse/cs1999/c6082abs.
htm.

Chen:1999:VCI

[CF99b] Jim X. Chen and Xiaodong Fu. Visualization corner: Integrating physics-based computing & visualization: Modeling dust behavior. Computing in Science and Engineering, 1 (1):12–16, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c1012.
pdf.

Carpenter:2003:HDP

[CF03] Bryan Carpenter and Geoffrey Fox. HPJava: a data parallel programming alternative. Computing in Science and Engineering, 5(3): 60–64, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/03/c3060abs.htm;
http://csdl.computer.org/
dl/mags/cs/2003/03/c3060.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/03/
c3060.pdf.

Cushing:2013:SDM

[CF13] Judith Bayard Cushing and James French. Science data management: Maximizing the yield [Guest Editors’ introduction]. Computing in Science and Engineering, 15(3): 8–10, May/June 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Carboni:2004:PPV

[CFA04] Rodrigo Carboni and Francisco Frutos-Alfaro. PCell: a 2D program for visualizing convective plasma cells. Computing in Science and Engineering, 6(4): 101–104, July/August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/04/c4101.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/04/
c4101.pdf.

Cortez:2004:SSO

[CFCD04] Ricardo Cortez, Lisa Fauci, Nathaniel Cowen, and Robert Dillon. Simulation of swimming organisms: Coupling internal mechanics with external fluid dynamics. Computing in Science and Engineering, 6(3): 38–45, May/June 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/03/c3038abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2004/03/c3038.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/03/
c3038.pdf.

Clay:2008:B

[CG08] R. Torsten Clay and Richard Gass. Books. Computing in Science and Engineering, 10 (2):85–87, March/April 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:/

/csdl.computer.org/comp/

mags/cs/2008/02/mcs2008020085.
pdf.

Cohen:2009:SCP

[CG09] Jonathan Cohen and Michael Garland. Solving computational problems with GPU computing. Computing in Science and Engineering, 11(5): 58–63, September/October 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Carver:2018:CUR

[CGK^{+}18] Jeffrey C. Carver, Sandra Gesing, Daniel S. Katz, Karthik Ram, and Nicholas Weber. Conceptualization of a US Research Software Sustainability Institute (URSSI). Computing in Science and Engineering, 20(3):4–9, May/ June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8358036/.

Cosmo:2020:RSC

[CGZ20] R. D. Cosmo, M. Gruenpeter, and S. Zacchiroli. Referencing source code artifacts: A separate concern in software citation. Computing in Science and Engineering, 22(2): 33–43, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chartier:2008:GIA

[Cha08]  Timothy P. Chartier.  A

Googol of information about Google. Computing in Science and Engineering, 10(6): 11–12, November/December 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Catlin:2019:LSS

[CHB19] Ann Christine Catlin, Chandima HewaNadungodage, and Andres Bejarano. Lifecycle support for scientific investigations: Integrating data, computing, and workflows. Computing in Science and Engineering, 21(4):49– 61, July/August 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Chen:2011:VES

[CHC^{+}11] Cheng-Kai Chen, Chris Ho, Carlos Correa, Kwan-Liu Ma, and Ahmed Elgamal. Visualizing 3D earthquake simulation data. Computing in Science and Engineering, 13(6): 52–63, November/December 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Carver:2017:IWS

[CHC17] Jeffrey C. Carver, Neil Chue Hong, and Selim Ciraci. The 4th International Workshop on Software Engineering for HPC in Computational Science and Engineering. Computing in Science and Engineering, 19(2):91– 95, March/April 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/02/mcs2017020091-
abs.html.

Chen:1999:VCJ

[Che99] Jim X. Chen. Visualization corner: a jump on visualization: The bottomsup approach. Computing in Science and Engineering, 1 (2):83–87, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2083.
pdf.

Chen:2003:NGP

[Che03] Jim X. Chen. New graphics pipeline approach speeds up atomic primitives rendering. Computing in Science and Engineering, 5(3):86–88, c3, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/03/c3086abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/03/c3086.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/03/
c3086.pdf.

Chen:2009:PR

[Che09] Jim X. Chen. A personal reflection. Computing in Science and Engineering, 11(4): 3, July/August 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chen:2010:GIS

[Che10] Jim X. Chen. Geographic information systems. Computing in Science and Engineering, 12(1):8–9, January/February 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chen:2015:AC

[Che15] Jim X. Chen. The advancement of computing. Computing in Science and Engineering, 17(5):4, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/05/mcs2015050004.
                                        html.

Chen:2016:ECA

[Che16] Jim X. Chen. The evolution of computing: AlphaGo. Computing in Science and Engineering, 18(4): 4–7, July/August 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chen:2017:PC

[Che17a] Jim X. Chen. Plans for CiSE. Computing in Science and Engineering, 19(1):4, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs2017010004.
html.

Chen:2017:TC

[Che17b] Jim X. Chen. Transparent computing. Computing in Science and Engineering, 19(1):5–6, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs2017010005.
html.

Chen:2018:URP

[Che18]  Jim X. Chen.  Under-

standing the research publication process. Computing in Science and Engineering, 20(2):6–7, March/April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8317988/.

Chen:2019:AEL

[Che19] J. X. Chen. Applications of extreme learning machines. Computing in Science and Engineering, 21(5):4–5, September/October 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Crouch:2013:SSI

[CHH^{+}13] Stephen Crouch, Neil Chue Hong, Simon Hettrick, Mike Jackson, Aleksandra Pawlik, Shoaib Sufi, Les Carr, David De Roure, Carole Goble, and Mark Parsons. The Software Sustainability Institute: Changing research software attitudes and practices. Computing in Science and Engineering, 15(6): 74–80, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Carver:2013:SPA

[CHHB13] Jeffrey Carver, Dustin Heaton, Lorin Hochstein, and Roscoe Bartlett. Self-perceptions about software engineering: A survey of scientists and engineers. Computing in Science and Engineering, 15(1):7–11, January/February 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chi:2021:DGD

[Chi21] Eric C. Chi. Discovering geometry in data arrays. Computing in Science and Engineering, 23(6):42–51, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chin:2005:SGC

[CHJC05] J. Chin, M. J. Harvey, S. Jha, and P. V. Coveney. Scientific Grid computing: The first generation. Computing in Science and Engineering, 7(5):24–32, September/October 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501736.pdf?
isnumber=32219&prod=JNL&
arnumber=1501736&arSt=+
24&ared=+32&arAuthor=Chin%
2C+J.%3B+Harvey%2C+M.J.%
3B+Jha%2C+S.%3B+Coveney%
2C+P.V;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=32219&arnumber=
1501736&count=14&index=4.

Cummings:2020:EBS

[CHM^{+}20a] R. T. Cummings, E. H.

Huff, N. A. Mack, K. Womack, A. Reid, B. Ghoram, K. Gosha, and J. E. Gilbert. An exploration of black students interacting with computing college and career readiness Vlog commentary social media influencers. Computing in Science and Engineering, 22(5):29–40, September/October 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cummings:2020:CEB

[CHM^{+}20b] R. T. Cummings, E. W.

Huff, N. A. Mack, K. Womack, A. Reid, B. Ghoram, K. Gosha, and J. E. Gilbert. Corrections to An Exploration of Black Students Interacting With Computing College and Career Readiness Vlog Commentary Social Media Influencers. Computing in Science and Engineering, 22 (6):112, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chonacky:2003:GEI

[Cho03] Norman Chonacky. Guest Editor’s introduction: Scientific databases. Computing in Science and Engineering, 5(3): 14–15, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/03/c3014abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/03/c3014.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/03/
                                        c3014.pdf.

Chonacky:2004:SGU

[Cho04] Norman Chonacky. Stella: Growing upward, downward, and outward. Computing in Science and Engineering, 6(3): 8–15, May/June 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/03/c3008abs.htm;
                                         http://csdl.computer.org/

dl/mags/cs/2004/03/c3008.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/03/
c3008.pdf.

Chonacky:2005:LH

[Cho05a] N. Chonacky. Lots of headroom. Computing in Science and Engineering, 7(5): 2, September/October 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501731.pdf?
isnumber=32219&prod=JNL&
arnumber=1501731&arSt=+2&
ared=+2&arAuthor=+Chonacky%
2C+N.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=32219&arnumber=
1501731&count=14&index=0.

Chonacky:2005:HJR

[Cho05b]  N. Chonacky.  Over-the-

horizon: Not just for radar anymore. Computing in Science and Engineering, 7 (4):4–5, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463128.pdf?
isnumber=31456&prod=JNL&
arnumber=1463128&arSt=+4&
ared=+5&arAuthor=Chonacky%
2C+N.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=31456&arnumber=
1463128&count=14&index=0.

Chonacky:2005:EEM

[Cho05c] Norman Chonacky. From the Editors: An evolving mission. Computing in Science and Engineering, 7(1):3–4, January/February 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2005/01/c1003.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/01/c1003.
                                        htm.

Chonacky:2005:EYW

[Cho05d] Norman Chonacky. From the Editors: This is the year that was. Computing in Science and Engineering, 7(6):4, November/December 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2005/06/c6004.pdf.

Chonacky:2005:SNG

[Cho05e] Norman Chonacky. Staking new ground. Computing in Science and Engineering, 7(2): 3–4, March/April 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2005/02/c2003.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/02/c2003.
                                        htm.

Chonacky:2005:WFL

[Cho05f] Norman Chonacky. The way forward: Learning from our past. Computing in Science and Engineering, 7(3): 2, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/03/c3002.pdf;
http://csdl.computer.org/
comp/mags/cs/2005/03/c3002abs.
htm.

Chonacky:2006:N

[Cho06a]  Norman Chonacky.  And

now. . .. Computing in Science and Engineering, 8(2): 2, March/April 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chonacky:2006:CLB

[Cho06b] Norman Chonacky. A change of landscape and a broadened scope. Computing in Science and Engineering, 8(4): 4–5, July/August 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2006/04/c4004.pdf.

Chonacky:2006:EJB

[Cho06c] Norman Chonacky. Evolution: It’s not just for biology. Computing in Science and Engineering, 8(3):2, May/ June 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chonacky:2006:CCP

[Cho06d]  Norman Chonacky.  Has

computing changed physics courses? Computing in Science and Engineering, 8(5):4–5, September/ October 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2006/05/c5004.pdf.

Chonacky:2006:NUL

[Cho06e]  Norman Chonacky.  Nov-

elty, utility, … and more. Computing in Science and Engineering, 8(6):3, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2006/06/c6003.pdf.

Chonacky:2006:STC

[Cho06f] Norman Chonacky. Special thanks to CiSE’s peer reviewers. Computing in Science and Engineering, 8(1):14, January/February 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2006/01/c1014.pdf.

Chonacky:2006:SSC

[Cho06g]  Norman Chonacky.  The

spread spectrum in computing. Computing in Science and Engineering, 8(1):2–3, January/February 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chonacky:2007:MMM

[Cho07a] Norman Chonacky. Musings on a metaphysics of modeling. Computing in Science and Engineering, 9(5):2–3, September/October 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/05/c5002.pdf.

Chonacky:2007:NSA

[Cho07b] Norman Chonacky. The neutrino strikes again. Computing in Science and Engineering, 9 (2):2, March/April 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/02/c2002.pdf.

Chonacky:2007:PPF

[Cho07c] Norman Chonacky. The past, present, and future. Computing in Science and Engineering, 9(4):2–3, July/ August 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/04/c4002.pdf.

Chonacky:2007:WSS

[Cho07d] Norman Chonacky. The Web:

For sharing or snaring? Computing in Science and Engineering, 9(6):3–4, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/06/mcs2007060003.
pdf.

Chonacky:2007:YWL

[Cho07e] Norman Chonacky. The year that was … and yet to be. Computing in Science and Engineering, 9(1):3–5, January/February 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/01/c1003.pdf.

Chonacky:2007:YRW

[Cho07f] Norman Chonacky. You’re recommending what?! Computing in Science and Engineering, 9(3):2, May/June 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2007/03/c3002.pdf.

Chonacky:2008:AAV

[Cho08a]  Norman Chonacky.  Ave

atque vale. Computing in Science and Engineering, 10 (6):3–5, November/December 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chonacky:2008:BF

[Cho08b] Norman Chonacky. Back to the future? Computing in Science and Engineering, 10 (3):3–4, May/June 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2008/03/mcs2008030003.
pdf.

Chonacky:2008:CDL

[Cho08c] Norman Chonacky. Coming to a digital library near you. Computing in Science and Engineering, 10(5):3–4, September/October 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chonacky:2008:TYC

[Cho08d] Norman Chonacky. Ten years and counting. . .. Computing in Science and Engineering, 10 (4):3, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2008/04/mcs2008040003.
pdf.

Chonacky:2008:TPL

[Cho08e] Norman Chonacky. Turning a page and learning something useful. Computing in Science and Engineering, 10(2): 4–5, March/April 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2008/02/mcs2008020004.
                                        pdf.

Chonacky:2008:WWY

[Cho08f] Norman Chonacky. Where in this world are you? Computing in Science and Engineering, 10(1):3–6, January/February 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2008/01/mcs2008010003.
                                        pdf.

Chowdhury:2008:MMD

[Cho08g] Debashish Chowdhury. Molecular motors: Design, mechanism, and control. Computing in Science and Engineering, 10(2):70–77, March/April 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chonacky:2009:CMP

[Cho09a] Norman Chonacky. Computation in modern physics. Computing in Science and Engineering, 11(4):6–7, July/ August 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chonacky:2009:MTB

[Cho09b] Norman Chonacky. A modern Tower of Babel. Computing in Science and Engineering, 11 (3):80, May/June 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chonacky:2012:CCL

[Cho12] Norman Chonacky. Crosstalk:

Computation as a language for research conversations among scientists and engineers. Computing in Science and Engineering, 14(1):9–11, January/February 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Catlin:2018:CSS

[CHP^{+}18] Ann Christine Catlin, Chandima HewaNadungodage, Santiago Pujol, Lucas Laughery, Chungwook Sim, Aishwarya Puranam, and Andres Bejarano. A cyberplatform for sharing scientific research data at DataCenterHub. Computing in Science and Engineering, 20(3):49–70, May/ June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8358033/.

Christian:1999:ESS

[Chr99] Wolfgang Christian. Educational software and the Sisyphus effect. Computing in Science and Engineering, 1 (3):13–15, May/June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3013.
pdf;   http://www.computer.
org/cse/cs1999/c3013abs.
htm.

Christian:2015:CSH

[Chr15] Carol Christian. Citizen science with Hubble space telescope data. Computing in Science and Engineering, 17(4): 12–19, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040012-
abs.html.

Chung:2021:RIP

[Chu21] Julianne Chung. Research in inverse problems and training in computational science: a reflection on the importance of community. Computing in Science and Engineering, 23(6):25–33, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cunningham:2016:ADY

[CJ16] Greg Cunningham and Katie Elyce Jones. Argonne discovery yields self-healing diamondlike carbon. Computing in Science and Engineering, 18(6):77–79, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/

cs/2016/06/mcs2016060077.
html.

Chandrasekaran:2018:BPR

[CJL^{+}18] Sunita Chandrasekaran, Guido Juckeland, Meifeng Lin, Matthew Otten, Dirk Pleiter, John E. Stone, Juan Lucio-Vega, Michael Zingale, and Fernanda Foertter. Best practices in running collaborative GPU hackathons: Advancing scientific applications with a sustained impact. Computing in Science and Engineering, 20(4):95– 106, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040095-
abs.html.

Chew:2013:IMD

[CJTH^{+}13] Teong Han Chew, Kwee Hong Joyce-Tan, Zeti Azura Mohamed Hussein, Pek Iee Elizabeth-Chia, and Mohd Shahir Shamsir. Improving molecular dynamics simulation performance on low-cost systems. Computing in Science and Engineering, 15(3):64–70, May/ June 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Calder:2009:ACA

[CK09] Alan C. Calder and Richard T.

Kouzes. Advances in computational astrophysics. Computing in Science and Engineering, 11(2):15–17, March/ April 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chonacky:2001:CII

[CL01] Norman Chonacky and Mitchell Lit. Computers for integrative instruction in bioengineering. Computing in Science and Engineering, 3(5):73– 83, September/October 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c5073abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2001/pdf/c5073.pdf.

Coutts:2012:DPP

[CL12] Duncan Coutts and Andres Loh. Deterministic parallel programming with Haskell. Computing in Science and Engineering, 14(6): 36–43, November/December 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cao:2014:ABH

[CL14] Danping Cao and Wenyuan Liao. An adjoint-based hybrid computational method for crosswell seismic inversion. Computing in Science and Engineering, 16(6): 60–67, November/December 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //csdl.computer.org/csdl/

mags/cs/2014/06/mcs2014060060-
abs.html.

Chyuan:2003:EMS

[CLC03] Shiang-Woei Chyuan, Yunn-Shiuan Liao, and Jeng-Tzong Chen. An efficient method for solving electrostatic problems. Computing in Science and Engineering, 5(3): 52–58, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/03/c3052abs.htm;
http://csdl.computer.org/
dl/mags/cs/2003/03/c3052.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/03/
c3052.pdf.

Cloteaux:2015:RFG

[Clo15]  Brian Cloteaux.  Is this

for real? Fast graphicality testing. Computing in Science and Engineering, 17(6): 91–95, November/December 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chen:2013:MSG

[CLZ13] Jim X. Chen, Yanlin Li, and Tianyu Zhang. Medical simulations [Guest Editors’ introduction]. Computing in Science and Engineering, 15(2): 8–9, March/April 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cosden:2022:RSE

[CMK22] Ian A. Cosden, Kenton McHenry, and Daniel S. Katz. Research software engineers: Career entry points and training gaps. Computing in Science and Engineering, 24(6): 14–21, November/December 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Crawford:2000:OPA

[CMN00] Dona Crawford, Donald Mc-Coy, and David Nowak. Our perspective on the alliance program’s benefits. Computing in Science and Engineering, 2(2):77–79, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c2077.
pdf;   http://www.computer.
org/cse/cs1999/c2077abs.
htm.

Chen:2003:GEI

[CN03] Jim X. Chen and Aiichiro Nakano. Guest Editors’ introduction: High-dimensional data acquisition, computing, and visualization. Computing in Science and Engineering, 5(2):12–13, March/April 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2003/02/c2012abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/02/

c2012.htm;   http://csdl.
computer.org/dl/mags/cs/
2003/02/c2012.pdf.

Chou:2010:CMI

[CNC10] Yu-Cheng Chou, Stephen S.

Nestinger, and Harry H. Cheng. Ch MPI: Interpretive parallel computing in C. Computing in Science and Engineering, 12(2):54– 67, March/April 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cowie:1999:MGI

[CNO99] James H. Cowie, David M.

Nicol, and Andy T. Ogielski. Modeling the global Internet. Computing in Science and Engineering, 1(1):42– 50, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c1042.
pdf;   http://www.computer.
org/cse/cs1999/c1042abs.
htm.

Cash:2015:GGI

[CO15] Brianna R. Cash and Dianne P. O’Leary. Gide: Graphical image deblurring exploration. Computing in Science and Engineering, 17 (3):62–67, May/June 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:/

/csdl.computer.org/csdl/

                                        mags/cs/2015/03/mcs2015030062-
                                        abs.html.

Cohen:2009:GTM

[Coh09]  Jonathan Cohen.  Graph

twiddling in a MapReduce world. Computing in Science and Engineering, 11(4): 29–41, July/August 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cole:2018:DDM

[Col18] Jacqueline M. Cole. Datadriven molecular engineering of solar-powered windows. Computing in Science and Engineering, 20 (1):84–87, January/February 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8291776/.

Compton:1999:WMM

[Com99] Tom Compton. Web mechanics: MathScript, Internet connection to Mathematica. Computing in Science and Engineering, 1(1):92– 95, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c1092.
                                        pdf.

Comba:2020:DVU

[Com20] J. L. D. Comba. Data visualization for the understanding of COVID-19. Computing in Science and Engineering, 22(6):81–86, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cook:2014:AMM

[Coo14] Brandon G. Cook. Applied mathematics: Methods and Matlab. Computing in Science and Engineering, 16(4): 6–7, July/August 2014. CO-DEN CSENFA. ISSN 1521-9615.

Cornette:2007:GVC

[Cor07] James L. Cornette. Gauss– Vaníček and Fourier transform spectral analyses of marine diversity. Computing in Science and Engineering, 9 (4):61–63, July/August 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:/

/csdl.computer.org/comp/
mags/cs/2007/04/WebExtra.
pdf.

Cox:2015:DMS

[COS^{+}15] Joe Cox, Eun Young Oh, Brooke Simmons, Chris Lintott, Karen Masters, Anita Greenhill, Gary Graham, and Kate Holmes. Defining and measuring success in online citizen science: A case study of zooniverse projects. Computing in Science and Engineering, 17(4): 28–41, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/04/mcs2015040028-
                                        abs.html.

Cosden:2022:PUR

[Cos22] Ian A. Cosden. The Princeton University Research Software Engineering Group model: Operational and organizational approaches. Computing in Science and Engineering, 24(5):24–31, September/October 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cottom:2003:USB

[Cot03]  Teresa L. Cottom.  Us-

ing SWIG to bind C++ to Python. Computing in Science and Engineering, 5(2):88–96, c3, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/02/c2088abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/02/c2088.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/02/
                                        c2088.pdf.

Courtes:2022:RPW

[Cou22]  Ludovic Courtès.  Repro-

ducibility and performance: Why choose? Computing in Science and Engineering, 24(3):77–80, May/June 2022.

CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Casas:2012:FDI

[CPdlF^{+}12] Francisco Javier Casas, Juan Pablo Pascual, Luisa de la Fuente, Eduardo Artal, and Joaquin Portilla. Frequency-domain identification and model-order selection for efficiently simulating microwave radiometers. Computing in Science and Engineering, 14(3):68–77, May/ June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cheatham:2015:IHC

[CR15] Thomas E. Cheatham and Daniel R. Roe. The impact of heterogeneous computing on workflows for biomolecular simulation and analysis. Computing in Science and Engineering, 17(2):30– 39, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/02/mcs2015020030-
abs.html.

Craig:2003:RCM

[Cra03]  Kevin Craig.  The role

of computers in mechatronics. Computing in Science and Engineering, 5(2):80– 85, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.

                                        computer.org/comp/mags/
                                        cs/2003/02/c2080abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/02/c2080.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/02/
                                        c2080.pdf.

Contrastin:2016:UMC

[CRDO16] M. Contrastin, A. Rice, M. Danish, and D. Orchard. Units-of-measure correctness in Fortran programs. Computing in Science and Engineering, 18(1):102–107, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Creutz:1999:ECP

[Cre99]  Michael Creutz.  Essays:

Computers and the particle theorist. Computing in Science and Engineering, 1 (4):14–15, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c4014.
                                        pdf.

Creutz:2004:SQ

[Cre04] Michael Creutz. Simulating quarks. Computing in Science and Engineering, 6(2): 80–85, March/April 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/02/c2080abs.htm;

http://csdl.computer.org/
dl/mags/cs/2004/02/c2080.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/02/
c2080.pdf.

Clematis:2001:AFS

[CS01a] Andrea Clematis and Michela Spagnuolo. Analyzing fuzzy surface modeling using loadbalanced computation. Computing in Science and Engineering, 3(6):74–81, November/December 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c6074abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c6074.pdf.

Cybenko:2001:GEI

[CS01b] George Cybenko and Francis Sullivan. Guest Editors’ introduction: Tomorrow’s hardest problems. Computing in Science and Engineering, 3(3):40–50, May/ June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3040.
pdf;   http://www.computer.
org/cse/cs1999c3040abs.
htm.

Conci:2011:SIP

[CS11] Aura Conci and Angel Sanchez.

Scientific image processing. Computing in Science and Engineering, 13(3):6–8, May/ June 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cytowski:2014:LSP

[CS14] Maciej Cytowski and Zuzanna Szymanska. Large-scale parallel simulations of 3D Cell colony dynamics. Computing in Science and Engineering, 16(5):86–95, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs2014050086-
abs.html.

Cytowski:2015:LSP

[CS15] Maciej Cytowski and Zuzanna Szymanska. Large-scale parallel simulations of 3D cell colony dynamics: The cellular environment. Computing in Science and Engineering, 17(5):44–48, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs2015050044-
abs.html.

Chindanur:2018:LDM

[CS18] Narendra Babu Chindanur and Pallaviram Sure. Lowdimensional models for traffic data processing using graph Fourier transform. Computing in Science and Engineering, 20(2):24–37, March/April 2018. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8254302/.

Comba:2023:DVD

[CSR^{+}23] João L. D. Comba, Nicolau O.

Santos, Jonathan C. Rivera, Regis K. Romeu, and Mara Abel. Data visualization for digital twins. Computing in Science and Engineering, 25 (2):58–63, March/April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chowdhury:2000:CSS

[CSS00] Debashish Chowdhury, Ludger Santen, and Andres Schadschneider. Computer simulations: Simulation of vehicular traffic: a statistical physics perspective. Computing in Science and Engineering, 2(5): 80–87, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c5080.
pdf.

Comba:2017:RV

[CSW17] Joao Comba, Filip Sadlo, and Daniel Weiskopf. A report from VIS 2016. Computing in Science and Engineering, 19 (2):82–90, March/April 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/

                                        mags/cs/2017/02/mcs2017020082-
                                        abs.html.

Concannon:2000:CPR

[CT00] Kristi D. Concannon and William J. Thompson. Computing prescriptions: Regression lines: More than meets the eye. Computing in Science and Engineering, 2(4):78–81, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c4078.
                                        pdf.

Cushing:2013:BBD

[Cus13] Judith Bayard Cushing. Beyond big data? Computing in Science and Engineering, 15 (5):4–5, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Cushing:2014:KPE

[Cus14] Judith Bayard Cushing. Keeping pace with extreme data. Computing in Science and Engineering, 16(4):4–5, July/ August 2014. CODEN CSENFA. ISSN 1521-9615.

Chonacky:2005:R

[CW05a] N. Chonacky and D. Winch. 3Ms: a response. Computing in Science and Engineering, 7(5):7–9, September/October 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501733.pdf?
isnumber=32219&prod=JNL&
arnumber=1501733&arSt=+7&
ared=+9&arAuthor=+Chonacky%
2C+N.%3B++Winch%2C+D.;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
32219&arnumber=1501733&
count=14&index=1.

Chonacky:2005:IPM

[CW05b] N. Chonacky and D. Winch.

3Ms for instruction, Part 2: Maple, Mathematica, and Matlab. Computing in Science and Engineering, 7(4): 14–23, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463131.pdf?
isnumber=31456&prod=JNL&
arnumber=1463131&arSt=+
14&ared=+23&arAuthor=Chonacky%
2C+N.%3B+Winch%2C+D.;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31456&arnumber=1463131&
count=14&index=1.

Chonacky:2005:IRM

[CW05c] Norman Chonacky and David Winch. 3Ms for instruction: Reviews of Maple, Mathematica, and Matlab. Computing in Science and Engineering, 7(3): 7–13, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/

                                        cs/2005/03/c3007.pdf;
                                        http://csdl.computer.org/
                                        comp/mags/cs/2005/03/c3007abs.
                                        htm.

Chonacky:2005:MMM

[CW05d] Norman Chonacky and David Winch. Maple, Mathematica, and Matlab: The 3M’s without the tape. Computing in Science and Engineering, 7(1):8–16, January/February 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2005/01/c1008.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/01/c1008.
                                        htm.

Chonacky:2005:RMM

[CW05e] Norman Chonacky and David Winch. Reviews of Maple, Mathematica, and Matlab: Coming soon to a publication near you. Computing in Science and Engineering, 7(2):9– 10, March/April 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2005/02/c2009.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/02/c2009.
                                        htm.

Chonacky:2006:WLY

[CW06] Norman Chonacky and David Winch. We’re listening to you! results from our recent reader survey. Computing in Science and Engineering, 8(3): 46–53, May/June 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Comba:2020:PFV

[CW20] J. Comba and D. Weiskopf.

The past and future of visualization for computing in science and engineering. Computing in Science and Engineering, 22(3):94–99, May/ June 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chen:2011:MPN

[CWOL11] Dan Chen, Lizhe Wang, Gaoxiang Ouyang, and Xiaoli Li. Massively parallel neural signal processing on a many-core platform. Computing in Science and Engineering, 13(6): 42–51, November/December 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chen:2020:NTS

[CXC^{+}20] C. Chen, C. Xiang, S. Cheng, L. Zhi, K. Li, Y. Chen, and X. Li. A novel transition strategy of hybrid direct torque control. Computing in Science and Engineering, 22(1):53–63, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Chen:2000:VCG

[CY00] Jim X. Chen and Yonggao Yang. Visualization corner: 3D graphics formats and conversions. Computing in Science and Engineering, 2(5): 67–73, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c5067.
pdf.

Cybenko:1999:ECC

[Cyb99a] George Cybenko. From the Editor-in-Chief: Computational culture shock. Computing in Science and Engineering, 1(1):1–3, January/February 1999. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c1001.
pdf.

Cybenko:1999:EII

[Cyb99b] George Cybenko. From the editor: Infrastructure immunology. Computing in Science and Engineering, 1 (6):1–3, November/December 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c6001.
pdf;   http://www.computer.
org/cse/cs1999/c6001abs.
htm.

Cybenko:1999:EIP

[Cyb99c]  George Cybenko.  From

the editors: The intellectual property lottery. Computing in Science and Engineering, 1(4):4, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c4004.
pdf.

Cybenk:2000:EWW

[Cyb00a] George Cybenk. From the editors: Who wants more? Computing in Science and Engineering, 2(4):3, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c4003.
pdf.

Cybenko:2000:EDD

[Cyb00b]  George Cybenko.  From

the editors: The death of disciplines. Computing in Science and Engineering, 2 (2):2–3, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c2002.
pdf;   http://www.computer.
org/cse/cs1999/c2002abs.
htm.

Cybenko:2000:ENM

[Cyb00c] George Cybenko. From the Editors: The New Millennium challenge. Computing in Science and Engineering, 2(6): 4–??, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c6004.
pdf;   pdf/c6004.pdf.

Cybenko:2001:BWR

[Cyb01] George Cybenko. Book and Web reviews: Machine learning. Computing in Science and Engineering, 3 (3):95–96, May/June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3095.
pdf.

Cybenko:2002:UQC

[Cyb02]  George Cybenko.  Under-

standing quantum computing. Computing in Science and Engineering, 4(4): 92–93, July/August 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2002/04/c4092.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/04/
c4092.pdf.

Chen:2001:VCP

[CYW01] Jim X. Chen, Yonggao Yang, and Xusheng Wang. Visualization corner: Physicsbased modeling and real-time simulation. Computing in Science and Engineering, 3 (3):98–102, May/June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3098.
pdf;   http://www.computer.
org/cse/cs1999c3098abs.
htm.

Chen:2007:GEI

[CZ07] Jim X. Chen and Tianyu Zhang. Guest Editors’ introduction: Anatomic rendering and visualization. Computing in Science and Engineering, 9(1):11–12, January/February 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/01/c1011.pdf.

deArcangelis:2003:MSG

[dA03] Lucilla de Arcangelis. Modeling the sol-gel transition. Computing in Science and Engineering, 5(6):78–87, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/dl/mags/cs/
2003/06/c6078.pdf.

Dachsbacher:2016:AVI

[Dac16] Carsten Dachsbacher. Ambient volume illumination. Computing in Science and Engineering, 18(2):90–97, March/ April 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Deslippe:2015:LLO

[DADY15] Jack Deslippe, Brian Austin, Chris Daley, and Woo-Sun Yang. Lessons learned from optimizing science kernels for Intel’s “Knights Corner” architecture. Computing in Science and Engineering, 17(3): 30–42, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/03/mcs2015030030-
                                        abs.html.

Datta:2018:ATL

[DAEJ18] Debanjan Datta, David Appelhans, Constantinos Evangelinos, and Kirk Jordan. An asynchronous two-level checkpointing method to solve adjoint problems on hierarchical memory spaces. Computing in Science and Engineering, 20(4):39–55, July/ August 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/04/mcs2018040039-
                                        abs.html.

Dey:2016:CGC

[DAKM16] S. Dey, R. M. Aubry, B. K. Karamete, and E. L. Mestreau. Capstone: A geometry-centric platform to enable physics-based simulation and system design. Computing in Science and Engineering, 18(1):32–39, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Daly:1999:CCL

[Dal99] Mark J. Daly. The computational challenge of linkage analysis: What causes diseases? Computing in Science and Engineering, 1 (3):18–25, May/June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3018.
pdf;   http://www.computer.
org/cse/cs1999/c3018abs.
htm.

Darooneh:2018:DRU

[Dar18] Amir Hossein Darooneh.

Doing research under extreme conditions. Computing in Science and Engineering, 20(3):31–35, May/June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8357993/.

Dasgupta:2000:RCC

[Das00] Siddharth Dasgupta. The role and challenges of computational chemistry in industrial problems. Computing in Science and Engineering, 2(6):

52–60, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c6052.
pdf;   http://www.computer.
org/cse/cs1999c6052abs.
htm.

Daukantas:1999:CST

[Dau99] Patricia Daukantas. Conferences: SC98 treats the past as prologue to computing’s future. Computing in Science and Engineering, 1(2): 94–95, 97–97, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2094.
pdf.

Davison:2012:ACE

[Dav12]  Andrew Davison.  Auto-

mated capture of experiment context for easier reproducibility in computational research. Computing in Science and Engineering, 14(4): 48–56, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2006:ECS

[Day06a] Charles Day. Engineering in computing and science. Computing in Science and Engineering, 8(2):88, March/April 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2006:MCE

[Day06b] Charles Day. My computational education. Computing in Science and Engineering, 8(5):104, September/October 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cs/2006/05/
c5104.pdf.

Day:2006:TYC

[Day06c] Charles Day. Turn on your computer and simulate the future! Computing in Science and Engineering, 8(3): 88, May/June 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Day:2007:CP

[Day07a]  Charles Day.  The com-

putation of poetry. Computing in Science and Engineering, 9(6):96, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/06/mcs2007060096.
pdf.

Day:2007:DDB

[Day07b] Charles Day. The death of distance has been exaggerated. Computing in Science and Engineering, 9(1):104, January/February 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cs/2007/01/
c1104.pdf.

Day:2007:QCE

[Day07c]  Charles Day.  Quantum

computing is exciting and important–really! Computing in Science and Engineering, 9(2):104, March/April 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://bell.computer.org/
dlcomments/;   http://csdl.
computer.org/comp/mags/
cs/2007/02/c2104.pdf.

Day:2007:WBB

[Day07d] Charles Day. When bits bite. Computing in Science and Engineering, 9(4): 96, July/August 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cs/2007/04/
c4096.pdf.

Day:2008:FPG

[Day08a]  Charles Day.  From first

principles—gauging the Web’s impact on scientific journals. Computing in Science and Engineering, 10(4):

96, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://bell.
computer.org/dlcomments/.

Day:2008:IC

[Day08b] Charles Day. The irony of craigslist. Computing in Science and Engineering, 10(2): 88, March/April 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cs/2008/02/
mcs2008020088.pdf.

Day:2008:SMH

[Day08c] Charles Day. Software makes hardware. Computing in Science and Engineering, 10 (6):104, November/December 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2009:AUC

[Day09a] Charles Day. Advice from Uncle Charles. Computing in Science and Engineering, 11(5): 96, September/October 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2009:CNM

[Day09b] Charles Day. “Compute? No, Mr. Bond, I Expect You to Die!”. Computing in Science and Engineering, 11(2):

88, March/April 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2009:LLD

[Day09c] Charles Day. Lessons from Little Dorrit. Computing in Science and Engineering, 11 (6):104, November/December 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2009:VD

[Day09d] Charles Day. Vets 1, docs 1. Computing in Science and Engineering, 11(4):96, July/August 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2010:CJL

[Day10a] Charles Day. Chihuahua jackets, Lego weapons, and vintage bags. Computing in Science and Engineering, 12(4): 96, July/August 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2010:FIA

[Day10b] Charles Day. Fixing Internet advertising: What would Donald Draper do? Computing in Science and Engineering, 12(6):88, November/December 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2010:MCS

[Day10c] Charles Day. My CiSEiest stories. Computing in Science and Engineering, 12(1): 96, January/February 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2010:MPC

[Day10d] Charles Day. My petaflops challenge. Computing in Science and Engineering, 12(3): 82, May/June 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2010:WAC

[Day10e] Charles Day. We’re all celebrities now. Computing in Science and Engineering, 12 (5):104, September/October 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2011:BHT

[Day11a] Charles Day. Buying a house, then and now. Computing in Science and Engineering, 13(2):88, March/April 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2011:CTB

[Day11b] Charles Day. Computational thinking is becoming one of the three Rs. Computing in Science and Engineering, 13(1):88, January/February 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2011:DBG

[Day11c] Charles Day. The demise of Borders Group and the challenge of browsing. Computing in Science and Engineering, 13(3):96, 95, May/ June 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2011:ESB

[Day11d] Charles Day. Earthquakes, soft bombs, and Internet vulnerability. Computing in Science and Engineering, 13(4): 104, July/August 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2011:MFC

[Day11e] Charles Day. It’s more fun to compose. Computing in Science and Engineering, 13 (6):96, November/December 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2011:SRO

[Day11f] Charles Day. Standards rule OK. Computing in Science and Engineering, 13(5): 96, September/October 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2012:FCS

[Day12a] Charles Day. The future of computational science — in 1977. Computing in Science and Engineering, 14(3): 104, May/June 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2012:GM

[Day12b] Charles Day. Gaming in meatspace. Computing in Science and Engineering, 14(5): 88, September/October 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2012:GCR

[Day12c]  Charles Day.  Germany’s

CiSE-reading Chancellor. Computing in Science and Engineering, 14(2):104, March/ April 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2012:MRC

[Day12d] Charles Day. Mind-reading computers. Computing in Science and Engineering, 14(4): 104, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2012:NPC

[Day12e] Charles Day. Nobel prizes for computational science [the last word]. Computing in Science and Engineering, 14 (6):88, November/December 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2012:SAW

[Day12f] Charles Day. “supercomputers are awesome and why I love what I DO!!!”. Computing in Science and Engineering, 14(1):88, January/February 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2013:BS

[Day13a] Charles Day. Blogging scholars. Computing in Science and Engineering, 15(6):104, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2013:CH

[Day13b]  Charles Day.  Computing

hell. Computing in Science and Engineering, 15(1):104, January/February 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2013:KMK

[Day13c] Charles Day. Knowing me, knowing you: Networks of cooperation. Computing in Science and Engineering, 15(3): 104, May/June 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2013:LWB

[Day13d] Charles Day. The last word:

Big surveillance. Computing in Science and Engineering, 15(4):88, July/August 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2013:MCC

[Day13e] Charles Day. Monte Carlo, colloids, and public health. Computing in Science and Engineering, 15(2):88, March/ April 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2013:SSM

[Day13f] Charles Day. Sociable social media. Computing in Science and Engineering, 15(5): 88, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2014:ASS

[Day14a] Charles Day. App-scale science. Computing in Science and Engineering, 16(2): 72, March/April 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2014:CC

[Day14b] Charles Day. Control and creativity. Computing in Science and Engineering, 16(4): 96, July/August 2014. CO-DEN CSENFA. ISSN 1521-9615.

Day:2014:MFR

[Day14c] Charles Day. My first Ereader. Computing in Science and Engineering, 16(3): 96, May/June 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2014:PP

[Day14d] Charles Day. Python power. Computing in Science and Engineering, 16(1):88, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2014:RN

[Day14e] Charles Day. Reason the need! Computing in Science and Engineering, 16(6):104, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs2014060104.
html.

Day:2014:WFS

[Day14f] Charles Day. Why Facebook should hire astronomers. Computing in Science and Engineering, 16(5):104, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs2014050104.
html.

Day:2015:AIB

[Day15a] Charles Day. Astronomical images before the Internet. Computing in Science and Engineering, 17(6):108, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2015:DET

[Day15b] Charles Day. Days of endless time. Computing in Science and Engineering, 17(4): 80, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040080.
html.

Day:2015:FE

[Day15c] Charles Day. Fixing email.

Computing in Science and Engineering, 17(3):88, May/ June 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/csdl/
mags/cs/2015/03/mcs2015030088.
html.

Day:2015:JW

[Day15d]  Charles Day.  Just 10

words. Computing in Science and Engineering, 17(2): 96, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/

                                        cs/2015/02/mcs2015020096.
                                        html.

Day:2015:MC

[Day15e] Charles Day. Modern computing — in 1949. Computing in Science and Engineering, 17(1):72, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/01/mcs2015010072.
                                        pdf.

Day:2016:AI

[Day16a] Charles Day. Archaeology by Internet. Computing in Science and Engineering, 18(2): 108, March/April 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2016:CC

[Day16b] Charles Day. Computers in cars. Computing in Science and Engineering, 18 (5):108, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2016:FT

[Day16c]  Charles Day.  Faraday’s

tablet. Computing in Science and Engineering, 18(4): 84, July/August 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2016:OBI

[Day16d] Charles Day. Our black, imperfect mirrors. Computing in Science and Engineering, 18(1):108, January/February 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2016:WYR

[Day16e] Charles Day. What you really wanted. Computing in Science and Engineering, 18 (3):96, May/June 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2016:WF

[Day16f] Charles Day. Writing fiction.

Computing in Science and Engineering, 18(6):84, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs2016060084.
html.

Day:2017:CAF

[Day17a] Charles Day. Computer-aided fashion. Computing in Science and Engineering, 19(3): 88, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs2017030088.
html.

Day:2017:CWC

[Day17b] Charles Day. Crisis? What crisis? Computing in Science and Engineering, 19(6):88, November/December 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2017/06/mcs2017060088.
html.

Day:2017:FYS

[Day17c] Charles Day. Find your summer soulmate with blur! Computing in Science and Engineering, 19(5):96, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/05/mcs2017050096.
html.

Day:2017:PCF

[Day17d] Charles Day. Pervasive computing on fictional farms. Computing in Science and Engineering, 19(1):88, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs2017010088.
html.

Day:2018:CRW

[Day18a] Charles Day. The consciences of robot warriors. Computing in Science and Engineering, 20(1):105–106, January/February 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Day:2018:DEC

[Day18b] Charles Day. Did everybody come? Computing in Science and Engineering, 20(2):111–112, March/April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8317997/.

Day:2018:MK

[Day18c] Charles Day. The march of kiosks. Computing in Science and Engineering, 20(6):104, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2018/06/08625917-
abs.html.

Day:2018:RSW

[Day18d]  Charles Day.  Robot sci-

ence writers. Computing in Science and Engineering, 20(3):101, May/June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8357997/.

Day:2018:VT

[Day18e] Charles Day. Virtual travel.

Computing in Science and Engineering, 20(4):125, July/ August 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/04/mcs2018040125.
                                        html.

Day:2018:WF

[Day18f] Charles Day. Whither Fanfic? Computing in Science and Engineering, 20(5):128, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/05/mcs2018050128.
                                        html.

Day:2019:FM

[Day19a] C. Day. The future of misinformation. Computing in Science and Engineering, 21(1):108, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Day:2019:WVP

[Day19b] Charles Day. Words, videos, and popes. Computing in Science and Engineering, 21(3):117–118, May/June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8701537/.

Dennin:2007:TNL

[DB07] Michael Dennin and Steven Barrett. Turning over a new leaf. Computing in Science and Engineering, 9(6): 67–69, November/December 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

DeMarle:2021:SVT

[DB21] David E. DeMarle and Andrew C. Bauer. In situ visualization with temporal caching. Computing in Science and Engineering, 23(3):25–33, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubey:2021:SAM

[DBB^{+}21] Anshu Dubey, Martin Berzins, Carsten Burstedde, Michael L. Norman, Didem Unat, and Mohammed Wahib. Structured adaptive mesh refinement adaptations to retain performance portability with increasing heterogeneity. Computing in Science and Engineering, 23(5):62– 66, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dial:2003:HTF

[DBCN03] Roman Dial, Barbara Bond, Judith Bayard Cushing, and Nalini Nadkarni. How trees and forests inform biodiversity and ecosystem informatics. Computing in Science and Engineering, 5(3): 32–43, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/03/c3032abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/03/c3032.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/03/
                                        c3032.pdf.

Devine:2002:ZDM

[DBH^{+}02] Karen Devine, Erik Boman, Robert Heapby, Bruce Hendrickson, and Courtenay Vaughan. Zoltan Data Management Service for parallel dynamic applications. Computing in Science and Engineering, 4(2):90–97, March/ April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://computer.org/cise/
                                        cs2001/c2090abs.htm;   http:
                                        //dlib.computer.org/cs/
                                        books/cs2002/pdf/c2090.
                                        pdf.

Desselle:2020:AVR

[DBJ^{+}20] M. R. Desselle, R. A. Brown, A. R. James, M. J. Midwinter, S. K. Powell, and M. A. Woodruff. Augmented and virtual reality in surgery. Computing in Science and Engineering, 22(3):18–26, May/ June 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

DiDio:2004:CEW

[DC04] Joseph DiDio III and Norman Chonacky. Caveat emptor: What to know before trying to beat a consumer system into a scientific instrument. Computing in Science and Engineering, 6(1):5–11, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/01/c1005abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2004/01/c1005.
pdf.

Das:2010:ZTQ

[DCC10] Arnab Das, Anjan K. Chandra, and Bikas K. Chakrabarti. A zero-temperature quantum Monte Carlo algorithm and quantum spin glasses. Computing in Science and Engineering, 12(1):64–72, January/February 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Davis:2007:HPC

[DCWH07] Larry P. Davis, Roy L. Campbell, Jr., William A. Ward, Jr., and Cray J. Henry. Highperformance computing acquisitions based on the factors that matter. Computing in Science and Engineering, 9(6):35–44, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dauger:2005:PPC

[DD05] Dean E. Dauger and Viktor K. Decyk. Plug-and-play cluster computing: Highperformance computing for the mainstream. Computing in Science and Engineering, 7(2):27–33, March/April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/02/c2027.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2005/02/
                                        c2027.pdf.

Dieter:2007:DCY

[DD07] William R. Dieter and Henry G. Dietz. Designing a cluster for your application. Computing in Science and Engineering, 9 (4):72–79, July/August 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dhruv:2023:MSP

[DD23] Akash Dhruv and Anshu Dubey. Managing software provenance to enhance reproducibility in computational research. Computing in Science and Engineering, 25(3): 60–65, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Downing:2004:DSS

[DDC04] Glenn Downing, Paul F. Dubois, and Teresa Cottom.

Data sharing in scientific simulations. Computing in Science and Engineering, 6(3): 87–96, May/June 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/03/c3087abs.htm;
http://csdl.computer.org/
dl/mags/cs/2004/03/c3087.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/03/
c3087.pdf.

Diehl:2023:SGAa

[DdS23a] Patrick Diehl and Rafael Ferreira da Silva. Science gateways: Accelerating research and education: Part I. Computing in Science and Engineering, 25(1):5–6, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Diehl:2023:SGAb

[DdS23b] Patrick Diehl and Rafael Ferreira da Silva. Science gateways: Accelerating research and education part II. Computing in Science and Engineering, 25(2):4– 5, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Duro:2008:IVR

[DDV^{+}08] Natividad Duro, Raquel Dormido, Héctor Vargas, Sebastián Dormido-Canto, José Sánchez, Gonzalo Farias, Francisco Esquembre, and Sebastián Dormido. An integrated virtual and remote control lab: The three-tank system as a case study. Computing in Science and Engineering, 10(4):50–59, July/August 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Decker:2017:BR

[DE17] Adrienne Decker and Kurt Eiselt. Best of RESPECT 2016. Computing in Science and Engineering, 19(3): 6–7, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs2017030006.
html.

Debusschere:2018:HDW

[Deb18] Bert Debusschere. How do we create more equitable, diverse, and inclusive organizations, and why does it matter? A white male’s perspective. Computing in Science and Engineering, 20(1):79–83, January/February 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Decyk:2015:SPC

[Dec15]  Viktor K. Decyk.  Skele-

ton particle-in-cell codes on emerging computer architectures. Computing in Science and Engineering, 17(2):

47–52, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/02/mcs2015020047-
abs.html.

Dubois:2003:WJC

[DEK03] Paul F. Dubois, Thomas Epperly, and Gary Kumfert. Why Johnny can’t build. Computing in Science and Engineering, 5(5):83–88, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/05/c5083abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/05/c5083.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/05/
c5083.pdf.

Dennis:2016:SCC

[Den16] Lawrence Dennis. STARS Computing Corps: Enhancing engagement of underrepresented students and building community in computing. Computing in Science and Engineering, 18(3):44–57, May/ June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dalcin:2021:MSU

[DF21] Lisandro Dalcin and Yao-Lung L. Fang. mpi4py: Status update after 12 years of development. Computing in Science and Engineering, 23(4):

47–54, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Donoho:2012:TDA

[DG12] David Donoho and Matan Gavish. Three dream applications of verifiable computational results. Computing in Science and Engineering, 14 (4):26–31, July/August 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Doerfler:2021:PPA

[DGG^{+}21] Douglas Doerfler, Steven Gottlieb, William Gropp, Barry I. Schneider, and Alan Sussman. Performance portability for advanced architectures. Computing in Science and Engineering, 23(5):7–9, September/October 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dolgert:2008:PHE

[DGJ^{+}08] Andrew Dolgert, Lawrence Gibbons, Christopher D. Jones, Valentin Kuznetsov, Mirek Riedewald, Daniel Riley, Gregory J. Sharp, and Peter Wittich. Provenance in high-energy physics workflows. Computing in Science and Engineering, 10(3): 22–29, May/June 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Delaney:2016:DAI

[DGK16] Keegan Delaney, Joseph Gorski, and Aaron Katz. Development and application of an incompressible Strand solver. Computing in Science and Engineering, 18(6):27–34, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs2016060027-
abs.html.

Dongarra:2019:RE

[DGK19] J. Dongarra, S. Gottlieb, and W. T. C. Kramer. Race to exascale. Computing in Science and Engineering, 21(1):4–5, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Droegemeier:2005:SOE

[DGR^{+}05] Kelvin K. Droegemeier, Dennis Gannon, Daniel Reed, Beth Plale, Jay Alameda, Tom Baltzer, Keith Brewster, Richard Clark, Ben Domenico, Sara Graves, Everette Joseph, Donald Murray, Rahul Ramachandran, Mohan Ramamurthy, Lavanya Ramakrishnan, John A. Rushing, Daniel Weber, Robert Wilhelmson, Anne Wilson, Ming Xue, and Sepideh Yalda. Service-oriented environments for dynamically interacting with mesoscale weather. Computing in Science and Engineering, 7(6):12–29, November/December 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Doleisch:2012:IVE

[DH12] Helmut Doleisch and Helwig Hauser. Interactive visual exploration and analysis of multivariate simulation data. Computing in Science and Engineering, 14(2): 70–77, March/April 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

DiPierro:2011:WSA

[Di 11] Massimo Di Pierro. web2py for scientific applications. Computing in Science and Engineering, 13(2):64–69, March/ April 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

DiPierro:2014:PPP

[Di 14] Massimo Di Pierro. Portable parallel programs with Python and OpenCL. Computing in Science and Engineering, 16 (1):34–40, January/February 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

DiPierro:2017:WB

[Di 17] Massimo Di Pierro. What is the blockchain? Computing in Science and Engineering, 19(5):92–95, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/05/mcs2017050092-
abs.html.

DiDio:2003:ODO

[DiD03] Joseph DiDio III. Opening doors with OpenOffice.org. Computing in Science and Engineering, 5(5):10–13, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/05/c5010abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/05/c5010.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/05/
c5010.pdf.

Diethelm:2012:LRN

[Die12] Kai Diethelm. The limits of reproducibility in numerical simulation. Computing in Science and Engineering, 14 (1):64–72, January/February 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

DiPierro:2018:RJ

[DiP18a] Massimo DiPierro. The rise of JavaScript. Computing in Science and Engineering, 20(1): 9–10, January/February 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

DiPierro:2018:TVG

[DiP18b] Massimo DiPierro. Toy visionguided 3D robotic arm in JavaScript. Computing in Science and Engineering, 20 (1):43–49, January/February 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8291773/.

Dongshan:2002:NMM

[DJ02] Xing Dongshan and Shen Junyi. A new Markov model for Web access prediction. Computing in Science and Engineering, 4(6):34–39, November/December 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/06/c6034abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/06/c6034.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/06/
                                        c6034.pdf.

deJong:2018:MJA

[dJM18] Jos de Jong and Eric Mansfield. Math.Js: An advanced mathematics library for JavaScript. Computing in Science and Engineering, 20 (1):20–32, January/February 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8291769/.

DeLeon:2013:LES

[DJS13] Rey DeLeon, Dana Jacobsen, and Inanc Senocak. Largeeddy simulations of turbulent incompressible flows on GPU clusters. Computing in Science and Engineering, 15 (1):26–33, January/February 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

deKoning:2000:CSE

[dKCAY00] Maurice de Koning, Wei Cai, Alex Antonelli, and Sidney Yip. Computer simulations: Efficient free-energy calculations by the simulation of nonequilibrium processes. Computing in Science and Engineering, 2(3):88–96, May/ June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3088.
pdf;   http://www.computer.
org/cse/cs1999/c3088abs.
htm.

Du:2014:ERS

[DKCL14] Yi-Chun Du, Chung-Dann Kan, Wei-Ling Chen, and Chia-Hung Lin. Estimating residual stenosis for an arteriovenous shunt using a flexible fuzzy classifier. Computing in Science and Engineering, 16(6):80–91, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.

                                        computer.org/csdl/mags/
                                        cs/2014/06/mcs2014060080-
                                        abs.html.

Dong:2005:CSC

[DKK05] Suchuan Dong, G. E. Karniadakis, and N. T. Karonis. Cross-site computations on the TeraGrid. Computing in Science and Engineering, 7(5):14–23, September/October 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        5992/32219/01501735.pdf?
                                        isnumber=32219&prod=JNL&
                                        arnumber=1501735&arSt=+
                                        14&ared=+23&arAuthor=+Suchuan+
                                        Dong%3B++Karniadakis%2C+
                                        G.E.%3B++Karonis%2C+N.T.;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        32219&arnumber=1501735&
                                        count=14&index=3.

DuBow:2017:MFC

[DKWL17] Wendy DuBow, Alexis Kaminsky, and Joanna Weidler-Lewis. Multiple factors converge to influence women’s persistence in computing: A qualitative analysis. Computing in Science and Engineering, 19(3):30–39, May/ June 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cs/2017/03/mcs2017030030.
                                        html.

Delp:2000:CFS

[DL00] Scott L. Delp and J. Peter Loan. A computational framework for simulating and analyzing human and animal movement. Computing in Science and Engineering, 2(5): 46–55, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c5046.
pdf;   http://www.computer.
org/cse/cs1999/c5046abs.
htm.

Donovan:2023:USE

[DL23] Sam Donovan and M. Drew LaMar. Using science education gateways to improve undergraduate STEM education: The QUBES platform as a case study. Computing in Science and Engineering, 25(2): 20–29, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Danese:2007:APS

[DLB^{+}07] Giovanni Danese, Francesco Leporati, Marco Bera, Mauro Giachero, Nelson Nazzicari, and Alvaro Spelgatti. An accelerator for physics simulations. Computing in Science and Engineering, 9(5): 16–25, September/October 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Donato:2004:SWC

[DLLM04] Debora Donato, Luigi Laura, Stefano Leonardi, and Stefano Millozzi. Simulating the Webgraph: a comparative analysis of models. Computing in Science and Engineering, 6(6): 84–89, November/December 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/06/c6084.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/06/
c6084.pdf.

Deng:2019:CBV

[DLLZ19] Y. Deng, T. Li, Y. Luo, and X. Zhao. CUDA-based volume rendering and inspection for time-varying ultrasonic testing datasets. Computing in Science and Engineering, 21(5):76–86, September/October 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic). See corrections [DLLZ20].

Deng:2020:CCB

[DLLZ20] Y. Deng, T. Li, Y. Luo, and X. Zhao. Corrections to “CUDA-Based Volume Rendering and Inspection for Time-Varying Ultrasonic Testing Datasets”. Computing in Science and Engineering, 22 (1):4, January/February 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). See [DLLZ19].

Deng:2019:DSV

[DLW^{+}19] Zigang Deng, Jipeng Li, Hongdi Wang, Yanxing Li, and Jun Zheng. Dynamic simulation of the vehicle/ bridge coupled system in high-temperature superconducting Maglev. Computing in Science and Engineering, 21(3):60–71, May/June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8667020/.

Diez:2004:SME

[DM04] Fernando Díez and Roberto Moriyón. Solving mathematical exercises that involve symbolic computations. Computing in Science and Engineering, 6(1):81–84, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/01/c1081abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2004/01/c1081.
pdf.

Doggett:2012:NCC

[DM12] Thomas Doggett and Nargess Memarsadeghi. NASA computational case study, hyperspectral data processing: Cryospheric change detection. Computing in Science and Engineering, 14(4):92– 97, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Diehl:2021:PMW

[DMA^{+}21] Patrick Diehl, Dominic Marcello, Parsa Amini, Hartmut Kaiser, Sagiv Shiber, Geoffrey C. Clayton, Juhan Frank, Gregor Daiß, Dirk Pflüger, David Eder, Alice Koniges, and Kevin Huck. Performance measurements within asynchronous task-based runtime systems: a double white dwarf merger as an application. Computing in Science and Engineering, 23(3): 73–81, May/June 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Donoho:2009:RRC

[DMR^{+}09] David L. Donoho, Arian Maleki, Inam Ur Rahman, Morteza Shahram, and Victoria Stodden. Reproducible research in computational harmonic analysis. Computing in Science and Engineering, 11(1):8–18, January/February 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubey:2021:PPE

[DMT^{+}21] Anshu Dubey, Lois Curfman McInnes, Rajeev Thakur, Erik W. Draeger, Thomas Evans, Timothy C. Germann, and William E. Hart. Performance portability in the exascale computing project: Exploration through a panel series. Computing in Science and Engineering, 23(5):46– 54, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Diaz-Montes:2014:FCM

[DMXR^{+}14] Javier Diaz-Montes, Yu Xie, Ivan Rodero, Jaroslaw Zola, Baskar Ganapathysubramanian, and Manish Parashar. Federated computing for the masses — aggregating resources to tackle large-scale engineering problems. Computing in Science and Engineering, 16(4):62–72, July/ August 2014. CODEN CSENFA. ISSN 1521-9615.

Decyk:2007:WF

[DNG07] Viktor K. Decyk, Charles D.

Norton, and Henry J. Gardner. Why Fortran? Computing in Science and Engineering, 9(4):68–71, July/August 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

deOliveira:2004:PMB

[dOMdO^{+}04] Suzana Moss de Oliveira, Jorge S. Sá Martins, Paulo Murilo C. de Oliveira, Karen Luz-Burgoa, Armando Ticona, and Thadeu J. P. Penna. The penna model for biological aging and speciation. Computing in Science and Engineering, 6(3): 74–81, May/June 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/03/c3074abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2004/03/c3074.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/03/
                                        c3074.pdf.

Donnelly:1999:EWC

[Don99] Denis Donnelly. Education: Winners of the CiSE software contest. Computing in Science and Engineering, 1(6): 65–69, November/December 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c6065.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c6065abs.
                                        htm.

Donnelly:2002:FEE

[Don02] Denis Donnelly. The fireworks effect: Exploring trajectory sets in time. Computing in Science and Engineering, 4(1):92–97, January/February 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c1092abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2002/pdf/c1092.pdf.

Donohue:2003:UUC

[Don03] George L. Donohue. Undergraduate use of complex simulation tools for airspace design. Computing in Science and Engineering, 5(5):72–75, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/05/c5072abs.htm;
http://csdl.computer.org/
dl/mags/cs/2003/05/c5072.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/05/
c5072.pdf.

Donnelly:2006:FFTa

[Don06a] Denis Donnelly. The Fast Fourier Transform for experimentalists, Part V: Filters. Computing in Science and Engineering, 8(1):92–95, January/February 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Donnelly:2006:FFTb

[Don06b] Denis Donnelly. The Fast Fourier Transform for experimentalists, Part VI: Chirp of a bat. Computing in Science and Engineering, 8(2): 72–78, March/April 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Donnelly:2010:ETW

[Don10]  Denis Donnelly.  Elliptic

trainer wattage: Fitting to find a function. Computing in Science and Engineering, 12(3):74–77, May/June 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dasgupta:2016:RAB

[DPBS16] A. Dasgupta, J. Poco, E. Bertini, and C. T. Silva. Reducing the analytical bottleneck for domain scientists: Lessons from a climate data visualization case study. Computing in Science and Engineering, 18 (1):92–100, January/February 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Donnellan:2012:DAC

[DPG^{+}12] Andrea Donnellan, Jay Parker, Margaret Glasscoe, Eric De Jong, Marlon Pierce, Geoffrey Fox, Dennis McLeod, John Rundle, and Lisa Grant Ludwig. A distributed approach to computational earthquake science: Opportunities and challenges. Computing in Science and Engineering, 14(5):31– 42, September/October 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dohmen:2001:PFL

[DPP^{+}01] Renate Dohmen, Jakob Pichlmeier, Max Petersen, Frank Wagner, and Matthias Scheffler. Parallel FP-LAPW for distributedmemory machines. Computing in Science and Engineering, 3(4):18–29, July/August 2001. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c4018.
pdf;   http://www.computer.
org/cse/cs1999c4018abs.
htm.

Donnelly:2005:FFTc

[DR05a] D. Donnelly and B. Rust.

The Fast Fourier Transform for experimentalists. Part II. Convolutions. Computing in Science and Engineering, 7 (4):92–95, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463142.pdf?
isnumber=31456&prod=JNL&
arnumber=1463142&arSt=+
92&ared=+95&arAuthor=Donnelly%
2C+D.%3B+Rust%2C+B.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=31456&
arnumber=1463142&count=
14&index=12.

Donnelly:2005:FFTb

[DR05b] Denis Donnelly and Bert Rust. The Fast Fourier Transform for experimentalists. Computing in Science and Engineering, 7(3): 71, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/03/c3071abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2005/03/c3071.
pdf.

Donnelly:2005:FFTa

[DR05c] Denis Donnelly and Bert Rust. The Fast Fourier Transform for experimentalists, Part I: Concepts. Computing in Science and Engineering, 7(2):80–88, March/April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/02/c2080.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2005/02/
                                        c2080.pdf.

Drach:2000:SPS

[Dra00] Robert Drach. Scientific programming: Serving scientific data over the Web. Computing in Science and Engineering, 2(6):14–18, November/December 2000. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c6014.
                                        pdf.

Doyle:2011:DTC

[DRA11] James Doyle, Carolyn Reynolds, and Clark Amerault. Diagnosing tropical cyclone sensitivity. Computing in Science and Engineering, 13(1):31–39, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Donnellan:2004:IEI

[DRR^{+}04] Andrea Donnellan, John Rundle, John Ries, Geoffrey Fox, Marlon Pierce, Jay Parker, Robert Crippen, Eric DeJong, Ben Chao, Weijia Kuang, Dennis McLeod, Mitsuhiro Matu’ura, and Jeremy Bloxham. Illuminating Earth’s interior through advanced computing. Computing in Science and Engineering, 6(1):36–44, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/01/c1036abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2004/01/c1036.
pdf.

Druskat:2020:SDR

[Dru20] S. Druskat. Software and dependencies in research citation graphs. Computing in Science and Engineering, 22(2): 8–21, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dongarra:2000:GEI

[DS00] Jack Dongarra and Francis Sullivan. Guest Editors’ introduction: The top 10 algorithms. Computing in Science and Engineering, 2 (1):22–23, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/

                                        books/cs2000/pdf/c1022.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c1022abs.

htm. See correspondence [MBS^{+}00].

DiPierro:2012:CMP

[DS12] Massimo Di Pierro and David Skinner. Concurrency in modern programming languages. Computing in Science and Engineering, 14(6):8–10, November/December 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Draeger:2023:EWI

[DS23] Erik W. Draeger and Andrew Siegel. Exascale was not inevitable; neither is what comes next. Computing in Science and Engineering, 25 (3):79–83, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dietrich:2009:MCS

[DSC^{+}09] Carlos A. Dietrich, Carlos E. Scheidegger, Joao L. D. Comba, Luciana P. Nedel, and Claudio T. Silva. Marching cubes without skinny triangles. Computing in Science and Engineering, 11(2): 82–87, March/April 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Diavastos:2015:EVW

[DSK15] Andreas Diavastos, Giannos Stylianou, and Giannis Koutsou. Exploiting very-wide vector processing for scientific applications. Computing in Science and Engineering, 17(6):83–87, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dag:2005:CSE

[DSPY05] Hasan Dag, Gürkan Soykan, Senol Piskin, and Osman Yasar. Computational science and engineering at Istanbul Technical University. Computing in Science and Engineering, 7 (1):72–77, January/February 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2005/01/c1072.
htm;   http://csdl.computer.
org/dl/mags/cs/2005/01/
c1072.pdf.

daSilva:2016:BTD

[dSRT16] Renato R. O. da Silva, Paulo E. Rauber, and Alexandru C. Telea. Beyond the third dimension: Visualizing highdimensional data with projections. Computing in Science and Engineering, 18(5): 98–107, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dongarra:2005:HPC

[DSSS05] Jack Dongarra, Thomas Sterling, Horst Simon, and Erich Strohmaier. Highperformance computing: Clusters, constellations, MPPs, and future directions. Computing in Science and Engineering, 7(2):51–59, March/ April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/02/c2051.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2005/02/
                                        c2051.pdf.

Dunning:2009:SEP

[DST^{+}09] Thom H. Dunning, Klaus Schulten, Jeroen Tromp, Jeremiah P. Ostriker, Kelvin Droegemeier, Ming Xue, and Paul Fussell. Science and engineering in the petascale era. Computing in Science and Engineering, 11(5):28– 37, September/October 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dutta:2021:CGT

[DTA21] S. Dutta, T. L. Turton, and J. P. Ahrens. A confidence-guided technique for tracking time-varying features. Computing in Science and Engineering, 23(2): 84–92, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dongarra:2017:ECR

[DTL^{+}17] Jack Dongarra, Stanimire Tomov, Piotr Luszczek, Jakub Kurzak, Mark Gates, Ichitaro Yamazaki, Hartwig Anzt, Azzam Haidar, and Ahmad Abdelfattah. With extreme computing, the rules have changed. Computing in Science and Engineering, 19(3): 52–62, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs2017030052-
abs.html.

Dubois:1999:SPT

[Dub99] Paul F. Dubois. Scientific programming: Ten good practices in scientific programming. Computing in Science and Engineering, 1(1): 7–11, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c1007.
pdf.

Dubois:2000:SPE

[Dub00]  Paul Dubois.  Scientific

programming: Extending Maple with compiled routines. Computing in Science and Engineering, 2(4): 82–86, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/

                                        books/cs2000/pdf/c4082.
                                        pdf.

Dubois:2002:DSC

[Dub02]  Paul F. Dubois.  De-

signing scientific components. Computing in Science and Engineering, 4(5): 84–90, September/October 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2002/05/c5084.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/05/
                                        c5084.pdf.

Dubois:2004:CDP

[Dub04] Paul Dubois. Cafe Dubois, polling place edition. Computing in Science and Engineering, 6(6):67–71, November/December 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2004/06/c6067.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/06/
                                        c6067.pdf.

Dubois:2005:BCR

[Dub05a] Paul F. Dubois. Bean-counted research is smelly. Computing in Science and Engineering, 7(6):96, November/December 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2005:MCS

[Dub05b] Paul F. Dubois. Maintaining correctness in scientific programs. Computing in Science and Engineering, 7(3): 80–85, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/03/c3080abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2005/03/c3080.
pdf.

Dubois:2005:NP

[Dub05c] Paul F. Dubois. A nest of Pythons. Computing in Science and Engineering, 7(6): 81–84, November/December 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2006:HB

[Dub06a] Paul Dubois. He’s baaack!

Computing in Science and Engineering, 8(5):68–69, September/October 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2006:SM

[Dub06b] Paul Dubois. So sue me. Computing in Science and Engineering, 8(6):64–65, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2007:CC

[Dub07a] Paul F. Dubois. Career contradictions. Computing in Science and Engineering, 9(6): 53–54, November/December 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2007:CSS

[Dub07b] Paul F. Dubois. Customer service says the darnedest things. Computing in Science and Engineering, 9(4): 66–67, July/August 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2007:DM

[Dub07c]  Paul F. Dubois.  Djange

me. Computing in Science and Engineering, 9(1):96–97, January/February 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2007:FNW

[Dub07d] Paul F. Dubois. The future that never was. Computing in Science and Engineering, 9(2): 68–69, March/April 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2007:GEI

[Dub07e] Paul F. Dubois. Guest Editor’s introduction: Python: Batteries included. Computing in Science and Engineering, 9 (3):7–9, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/03/c3007.pdf.

Dubois:2007:F

[Dub07f] Paul F. Dubois. Into the future. Computing in Science and Engineering, 9(5): 68–69, September/October 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2008:BCM

[Dub08a] Paul F. Dubois. Brain cancer may be the least of our worries. Computing in Science and Engineering, 10(3): 7–8, May/June 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2008:GA

[Dub08b] Paul F. Dubois. Goodbye to all that. Computing in Science and Engineering, 10(4): 93–95, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2008:SAE

[Dub08c] Paul F. Dubois. Sprinting ain’t easy. Computing in Science and Engineering, 10 (1):70–71, January/February 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2012:TSP

[Dub12] Paul F. Dubois. Testing scientific programs. Computing in Science and Engineering, 14 (4):69–73, July/August 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2015:AS

[Dub15a] Paul F. Dubois. Aging swiftly. Computing in Science and Engineering, 17(5):80, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs2015050080.
html.

Dubrow:2015:WGD

[Dub15b]  Aaron Dubrow.  What

got done in one year at NSF’s Stampede Supercomputer. Computing in Science and Engineering, 17(2): 83–88, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/02/mcs2015020083-
abs.html.

Dubey:2021:ISD

[Dub21] Anshu Dubey. Insights from the software design of a multiphysics multicomponent scientific code. Computing in Science and Engineering, 23(3): 92–95, May/June 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubey:2022:GPH

[Dub22] Anshu Dubey. Good practices for high-quality scientific computing. Computing in Science and Engineering, 24(6):72–76, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dunning:2009:GEI

[Dun09] Thom H. Dunning. Guest Editor’s introduction: Science and engineering on petascale computers. Computing in Science and Engineering, 11 (5):7–9, September/October 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

DeLisi:1999:CPC

[DV99] Charles DeLisi and Sandor Vajda. Computational problems in cell biology. Computing in Science and Engineering, 1(3):26–32, May/ June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3026.
pdf;   http://www.computer.
org/cse/cs1999/c3026abs.
htm.

DalCol:2017:WBV

[DVP^{+}17] Alcebiades Dal Col, Paola Valdivia, Fabiano Petronetto, Fabio Dias, Claudio T. Silva, and L. Gustavo Nonato. Wavelet-based visual analysis for data exploration. Computing in Science and Engineering, 19(5):85–91, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/05/mcs2017050085-
abs.html.

Deelman:2019:EPW

[DVR^{+}19] E. Deelman, K. Vahi, M. Rynge, R. Mayani, R. F. da Silva, G. Papadimitriou, and M. Livny. The evolution of the Pegasus Workflow Management Software. Computing in Science and Engineering, 21(4):22– 36, July/August 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Delgado:2022:CSL

[DVS22] Andrea Delgado, Veronica G. Melesse Vergara, and Andrea Schneibel. Careers in STEM: a Latina perspective. Computing in Science and Engineering, 24(3):81–85, May/ June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dongarra:2001:QPC

[DW01] Jack J. Dongarra and David W. Walker. The quest for petascale computing. Computing in Science and Engineering, 3(3):32–39, May/June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3032.
pdf;   http://ieeexplore.
ieee.org/iel5/5992/19880/
00919263.pdf;   http://
www.computer.org/cse/cs1999c3032abs.
htm;   http://www.netlib.
org/utk/people/JackDongarra/
PAPERS/petascale.pdf.

Davis:2011:HRH

[DWC^{+}11] Christopher Davis, Wei Wang, Steven Cavallo, James Done, Jimy Dudhia, Sherrie Fredrick, John Michalakes, Ginger Caldwell, Thomas Engel, and Ryan Torn. High-resolution hurricane forecasts. Computing in Science and Engineering, 13(1):22–30, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:1999:SPE

[DY99] Paul F. Dubois and T.-Y.

Yang. Scientific programming: Extending Python with Fortran. Computing in Science and Engineering, 1(5): 66–73, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5066.
pdf;   http://www.computer.
org/cse/cs1999/c5066abs.
htm.

Dong:2017:TTA

[DYY^{+}17] Jiaqing Dong, Hao Yin, Lyu Yongqiang, Hao Li, and Wei Wang. TAM: A transparent agent architecture for measuring mobile applications. Computing in Science and Engineering, 19(1):54–61, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/01/mcs2017010054-
                                        abs.html.

Dai:2005:VLT

[DZW^{+}05] Pei-Dong Dai, Tian-Yu Zhang, Zheng-Min Wang, Jim X. Chen, and Ke-Qiang Wang. A virtual laboratory for temporal bone microanatomy. Computing in Science and Engineering, 7(2):75–79, March/ April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/02/c2075.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2005/02/
                                        c2075.pdf.

Elia:2023:DSC

[EAF^{+}23] Donatello Elia, Fabrizio Antonio, Sandro Fiore, Paola Nassisi, and Giovanni Aloisio. A data space for climate science in the European Open Science Cloud. Computing in Science and Engineering, 25(1):7–15, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ebrahimi:2010:IPC

[Ebr10] Fatemeh Ebrahimi. Invasion percolation: a computational algorithm for complex phenomena. Computing in Science and Engineering, 12(2): 84–93, March/April 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ethier:2015:NIA

[ECK^{+}15] Stephane Ethier, Choon-Seock Chang, Seung-Hoe Ku, Wei li Lee, Weixing Wang, Zhihong Lin, and William Tang. NERSC’s impact on advances of global gyrokinetic PIC codes for fusion energy research. Computing in Science and Engineering, 17(3): 10–21, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/03/mcs2015030010-
abs.html.

Eddins:2009:AST

[Edd09] Steven L. Eddins. Automated software testing for Matlab. Computing in Science and Engineering, 11(6): 48–55, November/December 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ellison:2010:SAW

[EDJ^{+}10] Steven Ellison, John Dean, Michael Johnson, Cindy Prebola, Charles Fabozzi, and Alexis Cenko. Supplying air warfare capability through high-performance computing. Computing in Science and Engineering, 12(5):18– 26, September/October 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Esterie:2012:EME

[EGFL12] Pierre Esterie, Mathias Gaunard, Joel Falcou, and Jean-Thierry Lapreste. Exploiting multimedia extensions in C++: a portable approach. Computing in Science and Engineering, 14(5):72– 77, September/October 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Elster:2022:NHG

[EH22] Anne C. Elster and Tor A. Haugdahl. Nvidia Hopper GPU and Grace CPU highlights. Computing in Science and Engineering, 24(2):95– 100, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Etter:2001:ECH

[EHG01] Delores M. Etter, Charles J. Holland, and John Grosh. Export control of highperformance computing: Analysis and alternatives. Computing in Science and Engineering, 3(3):24–31, May/ June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3024.
pdf;   http://www.computer.
org/cse/cs1999c3024abs.
htm; pdf/c3024.pdf.

Eigenmann:2011:CET

[EI11] Rudolf Eigenmann and Ayhan Irfanoglu. Computational earthquake and tsunami research. Computing in Science and Engineering, 13(4): 11–13, July/August 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Eisenbach:2017:LSC

[Eis17] Markus Eisenbach. Largescale calculations for material sciences using accelerators to improve time- and energy-to-solution. Computing in Science and Engineering, 19(1):83–85, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs2017010083.
html.

Easterbrook:2009:ESU

[EJ09] Steve M. Easterbrook and Timothy C. Johns. Engineering the software for understanding climate change. Computing in Science and Engineering, 11(6):65–74, November/December 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Esler:2012:AQM

[EKCS12] Kenneth P. Esler, Jeongnim Kim, David M. Ceperley, and Luke Shulenburger. Accelerating quantum Monte Carlo simulations of real materials on GPU clusters. Computing in Science and Engineering, 14 (1):40–51, January/February 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Epstein:2007:SFC

[EKLY07] Howard E. Epstein, Jed O. Kaplan, Heike Lischke, and Qin Yu. Simulating future changes in Arctic and subarctic vegetation. Computing in Science and Engineering, 9(4): 12–23, July/August 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Elster:2021:EFA

[Els21] A. C. Elster. The European factor: From ARM to Atos. Computing in Science and Engineering, 23(1):102–105, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Engelhardt:2009:DCM

[Eng09]  Larry Engelhardt.  De-

scriptions and comparisons of Monte Carlo algorithms. Computing in Science and Engineering, 11(6):103, November/December 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Engdahl:2015:MIP

[Eng15] John C. Engdahl. Medical image processing you can apply. Computing in Science and Engineering, 17(4):78– 79, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040078.
html.

Eastman:2010:OHI

[EP10] Peter Eastman and Vijay Pande. OpenMM: a hardware-independent framework for molecular simulations. Computing in Science and Engineering, 12(4):34– 39, July/August 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Englund:2018:TDE

[EPHY18] Rickard Englund, Karljohan Lundin Palmerius, Ingrid Hotz, and Anders Ynnerman. Touching data: Enhancing visual exploration of flow data with haptics. Computing in Science and Engineering, 20(3):89–100, May/ June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8358030/.

Ebert:2021:VAR

[ERF21] David S. Ebert, Audrey Reinert, and Brian Fisher. Visual analytics review: an early and continuing success of convergent research with impact. Computing in Science and Engineering, 23(3):99–108, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ernst:2003:DVS

[ERS^{+}03] Thilo Ernst, Tom Rother, Franz Schreier, Jochen Wauer, and Wolfgang Balzer. DLR’s VirtualLab: Scientific software just a mouse click away. Computing in Science and Engineering, 5 (1):70–79, January/February 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/01/c1070.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/01/
c1070.pdf.

Eigenmann:2018:NSC

[ES18] Rudi Eigenmann and Barry I. Schneider. National Strategic Computing Initiative. Computing in Science and Engineering, 20(5):5–7, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/05/mcs2018050005.
html.

Elmagarmid:2008:CCE

[ESO08] Ahmed K. Elmagarmid, Arjmand Samuel, and Mourad Ouzzani. Community-cyberinfrastructure-enabled discovery in science and engineering. Computing in Science and Engineering, 10(5):46–53, September/October 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Esquembre:2011:TPL

[Esq11] Francisco Esquembre. There is parallel life for Java scientific programmers! Computing in Science and Engineering, 13(4):6–10, July/August 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ethier:2001:GEI

[Eth01] C. Ross Ethier. Guest Editor’s introduction: Bioengineering and biophysics. Computing in Science and Engineering, 3(5):38–39, September/October 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c5038abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2001/pdf/c5038.pdf.

Ebert-Uphoff:2015:IPI

[EUD15] Imme Ebert-Uphoff and Yi Deng. Identifying physical interactions from climate data: Challenges and opportunities. Computing in Science and Engineering, 17(6): 27–34, November/December 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Emani:2021:ASA

[EVA^{+}21] M. Emani, V. Vishwanath, C. Adams, M. E. Papka, R. Stevens, L. Florescu, S. Jairath, W. Liu, T. Nama, and . Sujeeth. Accelerating scientific applications with SambaNova reconfigurable dataflow architecture. Computing in Science and Engineering, 23(2):114–119, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Englander:2017:TOM

[EVL^{+}17] Jacob A. Englander, Matthew A. Vavrina, Lucy F. Lim, Lucy A. McFadden, Alyssa R. Rhoden, and Keith S. Noll. Trajectory optimization for missions to small bodies with a focus on scientific merit. Computing in Science and Engineering, 19(4):

18–28, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/04/mcs2017040018-
abs.html.

Ellis:2013:PSA

[EWN^{+}13] Heidi J. C. Ellis, Gerard Weatherby, Ronald J. Nowling, Jay Vyas, Matt Fenwick, and Michael Gryk. A pipeline software architecture for NMR spectrum data translation. Computing in Science and Engineering, 15(1):76–83, January/February 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Forouzandeh:2020:ACS

[FAFX20] S. Forouzandeh, A. R. Aghdam, S. Forouzandeh, and S. Xu. Addressing the coldstart problem using data mining techniques and improving recommender systems by cuckoo algorithm: A case study of Facebook. Computing in Science and Engineering, 22(4):62–73, July/August 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Falgout:2006:IAM

[Fal06] Robert D. Falgout. An introduction to algebraic multigrid. Computing in Science and Engineering, 8(6):24–33, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Falcou:2009:PPS

[Fal09] Joel Falcou. Parallel programming with skeletons. Computing in Science and Engineering, 11(3):58–63, May/ June 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Farmer:1999:PAS

[Far99] J. Doyne Farmer. Physicists attempt to scale the ivory towers of finance. Computing in Science and Engineering, 1(6): 26–39, November/December 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c6026.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c6026abs.
                                        htm.

Freire:2004:MXD

[FB04] Juliana Freire and Michael Benedikt. Managing XML data: An abridged overview. Computing in Science and Engineering, 6(4):12–19, July/ August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/cs/
                                        2004/04/c4012.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/cs/2004/04/c4012.pdf.

Fomel:2009:GEI

[FC09] Sergey Fomel and Jon F.

Claerbout. Guest Editors’ introduction: Reproducible research. Computing in Science and Engineering, 11(1):5–7, January/February 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Fan:2010:RAS

[FCT^{+}10] Jianrong Fan, Jim X. Chen, Bingwei Tian, Dong Yan, Genwei Cheng, Peng Cui, and Wen Zhang. Rapid assessment of secondary disasters induced by the Wenchuan earthquake. Computing in Science and Engineering, 12(1):10–19, January/February 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Feitelson:2005:SIL

[Fei05] Dror G. Feitelson. The supercomputer industry in light of the Top500 data. Computing in Science and Engineering, 7 (1):42–47, January/February 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2005/01/c1042.
htm;   http://csdl.computer.
org/dl/mags/cs/2005/01/
c1042.pdf.

Feldman:2000:BWR

[Fel00] David P. Feldman. Book and Web reviews: Monte Carlo methods. Computing in Science and Engineering, 2(6): 73–??, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c6073.
pdf.

Fent:2006:CAP

[Fen06] Thomas Fent. Counting (on) an aging population. Computing in Science and Engineering, 8(6):88–96, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Forbes:2003:GEI

[FF03] Nancy Forbes and Mike Foster. Guest Editors’ introduction: The end of Moore’s Law? Computing in Science and Engineering, 5 (1):18–19, January/February 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/01/c1018.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/01/
c1018.pdf.

Fox:2001:GCC

[FG01] Geoffrey Fox and Dennis Gannon. Grid computing: Computational grids. Computing in Science and Engineering, 3(4):74–77, July/August 2001. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c4074.
pdf.

FonsecaICasas:2022:FMS

[FGG^{+}22] P. Fonseca I Casas, J. Garcia I Subirana, V. Garcia I Carrasco, X. Pi I Palomés, and Gabriel Wainer. Formal modeling and simulation for SARS-CoV-2 containment scenarios in Catalonia. Computing in Science and Engineering, 24(3):86–90, May/ June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Foertter:2020:CSB

[FGHW20] F. Foertter, K. Gaither, K. Hinsen, and J. West. Computational science in the battle against COVID-19. Computing in Science and Engineering, 22(6):9–10, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Frenkel:1999:ECA

[FGP99] Daniel Frenkel, Laura Golebiowski, and Renato Portugal. Education: Computer algebra takes on the vibrating membrane problem. Computing in Science and Engineering, 1(2):88–93, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c2088.
                                        pdf.

Flanagin:2007:HSH

[FGR^{+}07] Maik Flanagin, Aurélien Grenotton, Jay Ratcliff, Kevin B. Shaw, John Sample, and Mahdi Abdelguerfi. Hydraulic splines: a hybrid approach to modeling river channel geometries. Computing in Science and Engineering, 9(5):4– 15, September/October 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Fortino:2017:MSI

[FGRS17] Giancarlo Fortino, Raffaele Gravina, Wilma Russo, and Claudio Savaglio. Modeling and simulating Internetof-things systems: A hybrid agent-oriented approach. Computing in Science and Engineering, 19(5):68–76, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/05/mcs2017050068-
                                        abs.html.

Fukushige:1999:HPS

[FHM99] Toshiyuki Fukushige, Piet Hut, and Junichiro Makino. High-performance specialpurpose computers in science. Computing in Science and Engineering, 1(2): 12–13, 16, March/April 1999.

CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2012.
pdf.

Fouilloux:2023:BCF

[FIG^{+}23] Anne Fouilloux, Jean Iaquinta, Alok Kumar Gupta, Hamish Struthers, Oskar Landgren, Prashanth Dwarakanath, Tommi Bergman, and Yanchun He. Building on communities to further software sustainability. Computing in Science and Engineering, 25(3):84–88, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Freiberger:2013:ALB

[FKB^{+}13] Manuel Freiberger, Florian Knoll, Kristian Bredies, Hermann Scharfetter, and Rudolf Stollberger. The Agile library for biomedical image reconstruction using GPU acceleration. Computing in Science and Engineering, 15(1):34–44, January/February 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Fangohr:2021:JCS

[FKD21] H. Fangohr, T. Kluyver, and M. DiPierro. Jupyter in computational science. Computing in Science and Engineering, 23(2):5–6, March/April 2021. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Faghmous:2015:CC

[FKS15] James H. Faghmous, Vipin Kumar, and Shashi Shekhar. Computing and climate. Computing in Science and Engineering, 17(6):6–8, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Freire:2008:PCT

[FKSS08] Juliana Freire, David Koop, Emanuele Santos, and Cláudio T. Silva. Provenance for computational tasks: a survey. Computing in Science and Engineering, 10(3):11–21, May/ June 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Forbes:1999:SCS

[FL99] Nancy A. Forbes and Laura F. Landweber. Spotlight: Computer science and the evolution of genetic information. Computing in Science and Engineering, 1(5): 12–15, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c5012.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c5012abs.
                                        htm.

Fernandez:2005:ANF

[FL05] J. Fernandez and S. Li.

Anisotropic nonlinear filtering of cellular structures in cryoelectron tomography. Computing in Science and Engineering, 7(5):54–61, September/October 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501741.pdf?
isnumber=32219&prod=JNL&
arnumber=1501741&arSt=+
54&ared=+61&arAuthor=+Fernandez%
2C+J.%3B++Li%2C+S.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=32219&
arnumber=1501741&count=
14&index=7.

Frehill:2021:TGW

[FL21] L. M. Frehill and M. A. Leung.

Twitter gone wrong: How constructive dialog and collaboration enable innovation. Computing in Science and Engineering, 23(1):97–101, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Flamm:2017:MLB

[Fla17]  Kenneth Flamm.  Has

Moore’s Law been repealed? an Economist’s perspective. Computing in Science and Engineering, 19(2):29–40, March/ April 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cs/2017/02/mcs2017020029-
                                        abs.html.

Faulk:2009:SCP

[FLV^{+}09] Stuart Faulk, Eugene Loh, Michael L. Van De Vanter, Susan Squires, and Lawrence G. Votta. Scientific computing’s productivity gridlock: How software engineering can help. Computing in Science and Engineering, 11(6): 30–39, November/December 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Forbes:2002:CST

[FM02] Nancy Forbes and Paul Messina. Computer science today in the European Union. Computing in Science and Engineering, 4(1):10– 14, January/February 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2002/
                                        c1010abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2002/pdf/c1010.pdf.

Feiereisen:2013:CSE

[FM13] William J. Feiereisen and James D. Myers. Computing in science and engineering — in manufacturing. Computing in Science and Engineering, 15 (1):12–15, January/February 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Fernandez-Macho:2019:PPW

[FM19] J. Fernández-Macho. Package Wavemulcor: Wavelet multiple regression and correlation in R. Computing in Science and Engineering, 21(6):63–73, November/December 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Fleming:2007:VPN

[FMB^{+}07] Timothy P. Fleming, Peter J. Mardahl, Lester Bowers, Keith L. Cartwright, Matthew T. Bettencourt, and Michael D. Haworth. Virtual prototyping of novel cathode designs for the relativistic magnetron. Computing in Science and Engineering, 9(6): 18–28, November/December 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Fyta:2008:MSN

[FMKS08] Maria Fyta, Simone Melchionna, Efthimios Kaxiras, and Sauro Succi. Multiscale simulation of nanobiological flows. Computing in Science and Engineering, 10(4): 10–19, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Feitosa:2011:WET

[FOdLVF^{+}11] Raul Queiroz Feitosa, Dario Augusto Borges Oliveira, Alvaro de Lima Veiga-Filho, Raphael Pithan Brito, Jose Luiz Buonomo de Pinho, and Antonio Carlos Censi. Weighting estimation for texturebased face recognition using the Fisher discriminant. Computing in Science and Engineering, 13(3):31–37, May/ June 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Fomel:2015:RRC

[Fom15] Sergey Fomel. Reproducible research as a community effort: Lessons from the Madagascar Project. Computing in Science and Engineering, 17(1):20–26, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/01/mcs2015010020-
abs.html.

Forbes:1999:SJR

[For99] Nancy A. Forbes. Spotlight: Japanese reforms foster innovative technology. Computing in Science and Engineering, 1 (6):4–9, November/December 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c6004.
pdf;   http://www.computer.
org/cse/cs1999/c6004abs.
htm.

Forbes:2000:FBI

[For00] Nancy Forbes. Focus: Biologically inspired computing. Computing in Science and Engineering, 2(6): 83–87, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c6083.
pdf.

Forbes:2001:FEC

[For01] Nancy Forbes. Focus: Evolution on a chip: Evolvable hardware aims to optimize circuit design. Computing in Science and Engineering, 3(3):6–10, May/June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3006.
pdf;   http://www.computer.
org/cse/cs1999c3006abs.
htm; pdf/c3006.pdf.

Forbes:2002:TNG

[For02]  Nancy Forbes.  Tread-

ing new ground. Computing in Science and Engineering, 4(5):12–13, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/05/c5012abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/05/c5012.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/05/
c5012.pdf.

Forbes:2016:BRP

[For16a] Jeff Forbes. Best of RE-SPECT, part 2. Computing in Science and Engineering, 18(3):11–13, May/June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Forbes:2016:GEI

[For16b] Jeff Forbes. Guest Editors’ introduction. Computing in Science and Engineering, 18(2): 6–8, March/April 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Foster:2017:QCF

[Fos17] Marcus P. Foster. Quantity correctness in Fortran programs. Computing in Science and Engineering, 19(4): 83–87, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/04/mcs2017040083-
                                        abs.html.

Fox:2001:WCP

[Fox01] Geoffrey Fox. Web computing: Peer-to-peer networks. Computing in Science and Engineering, 3(3):75–77, May/ June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2001/pdf/c3075.
                                        pdf;   http://www.computer.

org/cse/cs1999c3075abs.
htm.

Fox:2002:SMC

[Fox02a]  Geoffrey Fox.  E-science

meets computational science and information technology. Computing in Science and Engineering, 4(4): 84–85, July/August 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/04/c4084abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/04/c4084.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/04/
c4084.pdf.

Fox:2002:MPP

[Fox02b] Geoffrey Fox. Message passing: From parallel computing to the Grid. Computing in Science and Engineering, 4(5):70–73, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/05/c5070abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/05/c5070.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/05/
c5070.pdf.

Fox:2002:XIB

[Fox02c]  Geoffrey Fox.  XML and

the importance of being an object. Computing in Science and Engineering, 4(3):

96–98, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/03/c3096abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/03/c3096.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/03/
                                        c3096.pdf.

Fox:2003:DMS

[Fox03a] Geoffrey Fox. Data and metadata on the Semantic Grid. Computing in Science and Engineering, 5(5):76–78, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/05/c5076abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/05/c5076.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/05/
                                        c5076.pdf.

Fox:2003:GCE

[Fox03b] Geoffrey Fox. Grid computing environments. Computing in Science and Engineering, 5(2): 68–72, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/02/c2068abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/02/c2068.
                                        htm;   http://csdl.computer.

org/dl/mags/cs/2003/02/
c2068.pdf.

Fox:2003:ICI

[Fox03c]  Geoffrey Fox.  Integrat-

ing computing and information on Grids. Computing in Science and Engineering, 5(4):94–96, July/August 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/04/c4094.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/04/
c4094.pdf.

Fox:2003:JGA

[Fox03d]  Geoffrey Fox.  Java and

Grande applications. Computing in Science and Engineering, 5(1):60–62, January/February 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/01/c1060.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/01/
c1060.pdf.

Fox:2004:GGS

[Fox04a] Geoffrey Fox. Grids of grids of simple services. Computing in Science and Engineering, 6(4):84–87, July/August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/04/c4084.

                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/04/
                                        c4084.pdf.

Fox:2004:SDA

[Fox04b] Geoffrey Fox. Software development around a millisecond. Computing in Science and Engineering, 6(2): 93–96, March/April 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/02/c2093abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2004/02/c2093.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/02/
                                        c2093.pdf.

Fox:2018:EAC

[Fox18] John Fox. Evidence and argument: A cognitive view. Computing in Science and Engineering, 20(6):43–53, November/December 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/06/08565979-abs.
                                        html.

Freniere:2016:FAC

[FPRK16] Cole Freniere, Ashish Pathak, Mehdi Raessi, and Gaurav Khanna. The feasibility of Amazon’s cloud computing platform for parallel, GPU-accelerated, multiphaseflow simulations. Computing in Science and Engineering, 18(5):68–77, September/October 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Frank:2002:PLC

[Fra02]  Michael P. Frank.  The

physical limits of computing. Computing in Science and Engineering, 4(3): 16–26, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/03/c3016abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/03/c3016.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/03/
c3016.pdf.

Franklin:2007:ADF

[Fra07] William Franklin. Algorithmic design of a fuzzy-neural method. Computing in Science and Engineering, 9(5): 26–31, September/October 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
comp/mags/cs/2007/extras/
c5026_x1_Franklin.pdf.

Freire:2012:MCP

[FS12] Juliana Freire and Claudio T.

Silva. Making computations and publications reproducible with VisTrails. Computing in Science and Engineering, 14 (4):18–25, July/August 2012.

CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Feichter:2002:MCC

[FSD02] Johann Feichter, Martin Schultz, and Thomas Diehl. Modeling chemical constituents of the atmosphere. Computing in Science and Engineering, 4(5):56–63, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/05/c5056abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/05/c5056.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/05/
                                        c5056.pdf.

Fadili:2010:MRR

[FSED10] Jalal M. Fadili, Jean-Luc Starck, Michael Elad, and David L. Donoho. MCALab: Reproducible research in signal and image decomposition and inpainting. Computing in Science and Engineering, 12 (1):44–63, January/February 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Feiereisen:2008:I

[FT08] Bill Feiereisen and George K. Thiruvathukal. At issue. Computing in Science and Engineering, 10(2):60–64, March/ April 2008. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Fuller:2006:NCU

[Ful06] Robert G. Fuller. Numerical computations in US undergraduate physics courses. Computing in Science and Engineering, 8(5):16–21, September/October 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2006/extras/c5016x1.
doc;   http://csdl.computer.
org/comp/mags/cs/2006/extras/
c5016x2.doc;   http://csdl.
computer.org/comp/mags/
cs/2006/extras/c5016x3.
doc.

Fuller:2007:CSN

[FWGB07] Michael M. Fuller, Dali Wang, Louis J. Gross, and Michael W. Berry. Computational science for natural resource management. Computing in Science and Engineering, 9(4):40–48, July/August 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gaarenstroom:2003:SSS

[Gaa03] Stephen W. Gaarenstroom.

Surface science spectra: a hybrid journal-database. Computing in Science and Engineering, 5(3):26–30, May/ June 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/cs/2003/03/c3026abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/03/
                                        c3026.htm;   http://csdl.
                                        computer.org/dl/mags/cs/
                                        2003/03/c3026.pdf.

Godoy:2022:GRS

[GAB^{+}22] William F. Godoy, Ritu Arora, Keith Beattie, David E. Bernholdt, Sarah E. Bratt, Daniel S. Katz, Ignacio Laguna, Amiya K. Maji, Addi Malviya-Thakur, Rafael M. Mudafort, Nitin Sukhija, Damian Rouson, Cindy Rubio-González, and Karan Vahi. Giving research software engineers a larger stage through the better scientific software fellowship. Computing in Science and Engineering, 24 (5):6–13, September/October 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Galloy:2011:UDA

[Gal11]  Michael D. Galloy.  Us-

ing the data access protocol with IDL. Computing in Science and Engineering, 13(6):90–95, November/December 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ganguly:2002:HAI

[Gan02] Auroop R. Ganguly. A hybrid approach to improving rainfall forecasts. Computing in Science and Engineering, 4(4): 14–21, July/August 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/04/c4014abs.htm;
http://csdl.computer.org/
dl/mags/cs/2002/04/c4014.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/04/
c4014.pdf.

Gardner:2006:DLC

[Gar06] Pam Frost Gardner. Digital libraries come of age. Computing in Science and Engineering, 8(5):6–10, September/October 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2006/05/c5006.pdf.

Gargini:2017:HSO

[Gar17] Paolo A. Gargini. How to successfully overcome inflection points, or long live Moore’s Law. Computing in Science and Engineering, 19(2):51– 62, March/April 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/02/mcs2017020051-
abs.html.

Gonzlez-Arias:2020:SAA

[GARS^{+}20] F. González-Arias, T. Reddy, J. E. Stone, J. A. Hadden-Perilla, and J. R. Perilla.

Scalable analysis of authentic viral envelopes on FRON-TERA. Computing in Science and Engineering, 22(6): 11–20, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gill:2020:FMR

[GB20] S. S. Gill and R. Buyya. Failure management for reliable cloud computing: a taxonomy, model, and future directions. Computing in Science and Engineering, 22(3):52–63, May/ June 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gansterer:2004:FAE

[GBDW04] Wilfried N. Gansterer, Yihua Bai, Robert M. Day, and Robert C. Ward. A framework for approximating eigenpairs in electronic structure computations. Computing in Science and Engineering, 6(5): 50–59, September/October 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/05/c5050.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/05/
c5050.pdf.

Gayen:2011:AOB

[GBPR11] Dilip Gayen, Arunava Bhattacharyya, Rajat Pal, and Jitendra Nath Roy. All-optical binary-coded decimal adder with a terahertz optical asymmetric demultiplexer. Computing in Science and Engineering, 13(1):50–57, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Glatzmaier:2000:CAG

[GC00] Gary A. Glatzmaier and Thomas Clune. Computational aspects of geodynamo simulations. Computing in Science and Engineering, 2 (3):61–67, May/June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3061.
pdf;   http://www.computer.
org/cse/cs1999/c3061abs.
htm.

Gordon:2008:SIU

[GCV08] Steven I. Gordon, Kate Carey, and Ignatios Vakalis. A shared, interinstitutional undergraduate minor program in computational science. Computing in Science and Engineering, 10(5):12–16, September/October 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gordon:2016:ICO

[GDDR16] S. I. Gordon, J. Demmel, L. Destefano, and L. Rivera. Implementing a collaborative online course to extend access to HPC skills. Computing in Science and Engineering, 18 (1):73–79, January/February 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gesing:2023:VIC

[GDH^{+}23] Sandra Gesing, Ewa Deelman, Michael Hildreth, Ramandeep Makhija, Mary Ann McDowell, Natalie K. Meyers, and Douglas Thain. VisDict: Improving communication via a visual dictionary in a science gateway. Computing in Science and Engineering, 25(2): 7–11, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gedenk:2016:RMI

[Ged16a]  E. Gedenk.  Researchers

mine information from nextgeneration subsurface flow simulations. Computing in Science and Engineering, 18 (1):88–91, January/February 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gedenk:2016:IUI

[Ged16b] Eric Gedenk. Illuminating the Universe’s ignition. Computing in Science and Engineering, 18(4):80–83, July/August 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gatheral:1999:IOP

[GEH^{+}99] James Gatheral, Yonathan Epelbaum, Jining Han, Kishor Laud, Olga Lubovitsky, Elaine Kant, and Curt Randall. Implementing option-pricing models using software synthesis. Computing in Science and Engineering, 1(6): 54–64, November/December 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c6054.
pdf;   http://www.computer.
org/cse/cs1999/c6054abs.
htm.

Grinstein:2004:CCF

[GF04] Fernando F. Grinstein and Christer Fureby. From canonical to complex flows: Recent progress on monotonically integrated LES. Computing in Science and Engineering, 6(2): 36–49, March/April 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/02/c2036abs.htm;
http://csdl.computer.org/
dl/mags/cs/2004/02/c2036.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/02/
c2036.pdf.

Grant:2005:WSB

[GGD^{+}05] L. B. Grant, M. M. Gould, A. Donnellan, D. McLeod, A. Yun-An Chen, Sang-Soo Sung, M. Pierce, G. C. Fox, and P. Rundle. A Web services-based universal approach to heterogeneous fault databases. Computing in Science and Engineering, 7 (4):51–57, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463136.pdf?
isnumber=31456&prod=JNL&
arnumber=1463136&arSt=+
51&ared=+57&arAuthor=Grant%
2C+L.B.%3B+Gould%2C+M.M.
%3B+Donnellan%2C+A.%3B+
McLeod%2C+D.%3B+Yun-An+
Chen%2C+A.%3B+Sang-Soo+
Sung%3B+Pierce%2C+M.%3B+
Fox%2C+G.C.%3B+Rundle%2C+
P.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=31456&arnumber=
1463136&count=14&index=6.

Gao:2023:FEB

[GGJD23] Yuxiang Gao, Gourab Ghosh, Stephen Jiménez, and Ravindra Duddu. A finite-element-based cohesive zone model of water-filled surface crevasse propagation in floating ice tongues. Computing in Science and Engineering, 25(3): 8–16, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gubernatis:2000:CSM

[GH00] Jim Gubernatis and Naomichi Hatano. Computer simulations: The multicanonical Monte Carlo method. Computing in Science and Engineering, 2(2):95–103, March/ April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c2095.
pdf;   http://www.computer.
org/cse/cs1999/c2095abs.
htm.

Gorder:2004:Nb

[GH04] Pam Frost Gorder and Lissa E.

Harris. News. Computing in Science and Engineering, 6(3): 4–7, May/June 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/03/c3004abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2004/03/c3004.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/03/
c3004.pdf.

Goldrian:2008:QQC

[GHK^{+}08] Gottfried Goldrian, Thomas Huth, Benjamin Krill, Jack Lauritsen, Heiko Schick, Ibrahim Ouda, Simon Heybrock, Dieter Hierl, Thilo Maurer, Nils Meyer, Andreas Schafer, Stefan Solbrig, Thomas Streuer, Tilo Wettig, Dirk Pleiter, Karl-Heinz Sulanke, Frank Winter, Hubert Simma, Sebastiano Fabio Schifano, Raffaele Tripiccione, Andrea Nobile, Matthias Drochner, Thomas Lippert, and Zoltan Fodor. QPACE: Quantum chromodynamics parallel computing on the Cell Broadband Engine. Computing in Science and Engineering, 10(6): 46–54, November/December 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gries:2011:SFA

[GHKR11] Matthias Gries, Ulrich Hoffmann, Michael Konow, and Michael Riepen. SCC: a flexible architecture for manycore platform research. Computing in Science and Engineering, 13(6):79–83, November/December 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Guo:2017:CCA

[GHKZ17] Kehua Guo, Yujian Huang, Li Kuang, and Yaoxue Zhang. CASP: A context-aware transparent active service provision architecture in a mobile Internet environment. Computing in Science and Engineering, 19(1):38–45, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/01/mcs2017010038-
                                        abs.html.

Grossman:2016:CDC

[GHM^{+}16] Robert L. Grossman, Allison Heath, Mark Murphy, Maria Patterson, and Walt Wells. A case for data commons: Toward data science as a service. Computing in Science and Engineering, 18(5):10– 20, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ge:2010:EDV

[GHT^{+}10] Jinghua Ge, Andrei Hutanu, Cornelius Toole, Robert Kooima, Imtiaz Hossain, and Gabrielle Allen. An experimental distributed visualization system for petascale computing. Computing in Science and Engineering, 12(5):78– 83, September/October 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Giabbanelli:2022:HMC

[Gia22] Philippe J. Giabbanelli. Hybrid models that combine machine learning and simulations. Computing in Science and Engineering, 24(5):72– 76, September/October 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Grinberg:2012:TCA

[GIF^{+}12] Leopold Grinberg, Joseph A.

Insley, Dmitry A. Fedosov, Vitali Morozov, Michael E. Papka, and George Em Karniadakis. Tightly coupled atomistic-continuum simulations of brain blood flow on petaflop supercomputers. Computing in Science and Engineering, 14(6):58–67, November/December 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Giger:2000:CAD

[Gig00] Maryellen L. Giger. Computeraided diagnosis of breast lesions in medical images. Computing in Science and Engineering, 2(5):39–45, September/October 2000. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c5039.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c5039abs.
                                        htm.

Giordano:2002:HSV

[Gio02] Nicholas Giordano. On hearing the “Shape” of a vibrating string. Computing in Science and Engineering, 4(3):100, C3, May/June 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2002/03/c3100.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/03/
                                        c3100.pdf.

Gorder:2003:Na

[GJ03a] Pam Frost Gorder and Anne Jacobson. News. Computing in Science and Engineering, 5(2):4–8, March/April 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2003/02/c2004abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/02/
c2004.htm;   http://csdl.
computer.org/dl/mags/cs/
2003/02/c2004.pdf.

Gorder:2003:Nb

[GJ03b] Pam Frost Gorder and Anne Jacobson. News. Computing in Science and Engineering, 5 (3):4–8, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/03/c3004abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/03/c3004.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/03/
c3004.pdf.

Gottlieb:2018:SEA

[GK18] Sigal Gottlieb and Gaurav Khanna. Supercomputingenabled advances in science and engineering. Computing in Science and Engineering, 20(4):8–9, July/ August 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040008.
html.

Gamblin:2022:OCC

[GK22] Todd Gamblin and Daniel S.

Katz. Overcoming challenges to continuous integration in HPC. Computing in Science and Engineering, 24(6): 54–59, November/December 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ganguly:2015:CAI

[GKG^{+}15] Auroop R. Ganguly, Devashish Kumar, Poulomi Ganguli, Geoffrey Short, and James Klausner. Climate adaptation informatics: Water stress on power production. Computing in Science and Engineering, 17(6):53–60, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Germann:1999:RAL

[GL99] Timothy C. Germann and Peter S. Lomdahl. Recent advances in large-scale atomistic materials simulations. Computing in Science and Engineering, 1(2): 10–11, 16, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c2010.
                                        pdf.

Gregerson:2008:UTT

[GL08] Robert G. Gregerson and Tim H. Lindblom. A unique textbook for teaching courses in bioinformatics. Computing in Science and Engineering, 10(1):7–8, January/February 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2008/01/mcs2008010007.
pdf.

Graves:2020:HGR

[GL20] A. L. Graves and A. D.

Light. Hitting the ground running: Computational physics education to prepare students for computational physics research. Computing in Science and Engineering, 22(4): 50–60, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ginn:2007:ENE

[GLS07] Timothy R. Ginn, Frank J.

Loge, and Timothy D. Scheibe. Explaining ‘noise’ as environmental variations in population dynamics. Computing in Science and Engineering, 9(2): 40–49, March/April 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

George:2011:NGF

[GLS11] Alan George, Herman Lam, and Greg Stitt. Novo-G: At the forefront of scalable reconfigurable supercomputing. Computing in Science and Engineering, 13(1):82–86, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ghazi:2010:WHU

[GLTZ10] Kaveh R. Ghazi, Vincent Lefevre, Philippe Theveny, and Paul Zimmermann. Why and how to use arbitrary precision. Computing in Science and Engineering, 12(3): 5, May/June 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Grossman:2002:DDW

[GM02] Robert Grossman and Marco Mazzucco. DataSpace: a data Web for the exploratory analysis and mining of data. Computing in Science and Engineering, 4(4):44–51, July/ August 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/04/c4044abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/04/c4044.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/04/
                                        c4044.pdf.

Gray:2006:SWW

[GM06] Paul Gray and Thomas Murphy. Something wonderful this way comes. Computing in Science and Engineering, 8(3): 82–87, May/June 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gomez-Mendoza:2011:ALC

[GMPR11] Juan-Bernardo Gomez-Mendoza, Flavio Prieto, and Tanneguy Redarce. Automatic lip-contour extraction and mouth-structure segmentation in images. Computing in Science and Engineering, 13(3): 22–30, May/June 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gobbert:2008:PGS

[GN08] Matthias K. Gobbert and Nagaraj K. Neerchal. Preparing graduate students for interdisciplinary careers. Computing in Science and Engineering, 10(1):93–95, January/February 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gyulassy:2009:RTB

[GNB^{+}09] Attila Gyulassy, Luis Gustavo Nonato, Peer-Timo Bremer, Claudio Silva, and Valerio Pascucci. Robust topologybased multiscale analysis of scientific data. Computing in Science and Engineering, 11(5):88–95, September/October 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gobbert:2005:CPB

[Gob05] Matthias K. Gobbert. Configuration and performance of a Beowulf cluster for large-scale scientific simulations. Computing in Science and Engineering, 7(2): 14–26, March/April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2005/02/c2014.
htm;   http://csdl.computer.
org/dl/mags/cs/2005/02/
c2014.pdf.

Goodacre:2017:IDS

[Goo17] John Goodacre. Innovating the delivery of server technology with Kaleao KMAX. Computing in Science and Engineering, 19(5):77–81, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/05/mcs2017050077.
html.

Gorder:2003:Nc

[Gor03] Pam Frost Gorder. News. Computing in Science and Engineering, 5(6):6–9, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/06/c6006abs.htm;
 http://csdl.computer.org/

dl/mags/cs/2003/06/c6006.
pdf.

Gorder:2004:Na

[Gor04a] Pam Frost Gorder. News.

Computing in Science and Engineering, 6(2):6–11, March/ April 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2004/02/c2006abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/02/
c2006.htm;   http://csdl.
computer.org/dl/mags/cs/
2004/02/c2006.pdf.

Gorder:2004:Nc

[Gor04b] Pam Frost Gorder. News.

Computing in Science and Engineering, 6(4):6–9, July/ August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/04/c4006.pdf;
http://csdl.computer.org/
dl/mags/cs/2004/04/c4006.
htm.

Gorder:2004:Nd

[Gor04c] Pam Frost Gorder. News.

Computing in Science and Engineering, 6(6):4–7, November/December 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/06/c6004.pdf;
http://csdl.computer.org/

                                        dl/mags/cs/2004/06/c6004.
                                        htm.

Gorder:2005:CLF

[Gor05a] Pam Frost Gorder. Computing life’s family tree. Computing in Science and Engineering, 7(3):3–6, May/June 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/cs/2005/03/c3003.pdf;
                                         http://csdl.computer.org/
                                        comp/mags/cs/2005/03/c3003abs.
                                        htm.

Gorder:2005:DDR

[Gor05b] Pam Frost Gorder. Digital detectives reveal art forgeries. Computing in Science and Engineering, 7(2): 5–8, March/April 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2005/02/c2005.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/02/c2005.
                                        htm.

Gorder:2005:Na

[Gor05c] Pam Frost Gorder. News. Computing in Science and Engineering, 7(1):5–7, January/February 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2005/01/c1005.pdf;
                                        http://csdl.computer.org/

dl/mags/cs/2005/01/c1005.
htm.

Gorder:2005:Nb

[Gor05d] Pam Frost Gorder. News.

Computing in Science and Engineering, 7(6):6–9, November/December 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/06/c6006.pdf.

Gorder:2006:CBT

[Gor06a] Pam Frost Gorder. Computing in biological time: The design of an anticocaine molecule. Computing in Science and Engineering, 8(4): 6–9, July/August 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2006/04/c4006.pdf.

Gorder:2006:RAI

[Gor06b]  Pam Frost Gorder.  En

route to artificial intelligence, software learns language. Computing in Science and Engineering, 8(1):4–7, January/February 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2006/01/c1004.pdf.

Gorder:2006:NNS

[Gor06c] Pam Frost Gorder. Neural networks show new promise for machine vision. Computing in Science and Engineering, 8(6):4–8, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2006/06/c6004.pdf.

Gorder:2006:JBA

[Gor06d] Pam Frost Gorder. Not just for the birds: Archiving massive data sets. Computing in Science and Engineering, 8(3): 3–7, May/June 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2006/03/c3003.pdf.

Gorder:2007:BBS

[Gor07a] Pam Frost Gorder. Building better search engines. Computing in Science and Engineering, 9(4):7–11, July/ August 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gorder:2007:GCY

[Gor07b] Pam Frost Gorder. Grid computing yields earthquake forecast. Computing in Science and Engineering, 9(1):6–10, January/February 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/01/c1006.pdf.

Gorder:2007:MPS

[Gor07c] Pam Frost Gorder. Multicore processors for science and engineering. Computing in Science and Engineering, 9(2): 3–7, March/April 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/02/c2003.pdf.

Gorder:2007:PEC

[Gor07d] Pam Frost Gorder. Physics experiment could spawn permanent computing grid. Computing in Science and Engineering, 9(6):5–9, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/06/mcs2007060005.
pdf.

Gorder:2008:CSR

[Gor08a] Pam Frost Gorder. Coming soon: Research in a cloud. Computing in Science and Engineering, 10(6):6–10, November/December 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gorder:2008:CVI

[Gor08b] Pam Frost Gorder. Computer vision, inspired by the human brain. Computing in Science and Engineering, 10(2): 6–11, March/April 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gorder:2008:MSA

[Gor08c] Pam Frost Gorder. Medical software has astronomers seeing stars. Computing in Science and Engineering, 10(4): 4–9, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gorder:2010:CE

[Gor10] Pam Frost Gorder. Computational epidemiology. Computing in Science and Engineering, 12(1):4–6, January/February 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gordon:2013:ACS

[Gor13] Steven I. Gordon. Advancing computational science education through Xsede. Computing in Science and Engineering, 15(1):90–92, January/February 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Goth:2001:WTT

[Got01] Greg Goth. Will Terra be terrific? Computing in Science and Engineering, 3(6):4– 8, November/December 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c6004abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c6004.pdf.

Goth:2002:FHP

[Got02a] Greg Goth. Fans of Hewlett– Packard calculators say “it all adds up”. Computing in Science and Engineering, 4 (2):5–8, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c2005abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c2005.pdf.

Goth:2002:ISE

[Got02b] Greg Goth. Infrastructure simulation effort has high hopes, faces high hurdles. Computing in Science and Engineering, 4(5):4–11, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/05/c5004abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/05/c5004.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/05/
c5004.pdf.

Gottlieb:2006:GEI

[Got06] Steven Gottlieb. Guest Editor’s introduction: Specialpurpose computing. Computing in Science and Engineering, 8(1):15–17, January/February 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gottlieb:2014:ALC

[Got14a] Steven Gottlieb. Anecdotes on leadership computing. Computing in Science and Engineering, 16(5):4–6, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2014/05/mcs2014050004.
                                        html.

Gottlieb:2014:BCS

[Got14b]  Steven Gottlieb.  Becom-

ing a computational scientist. Computing in Science and Engineering, 16(1):4–5, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gottlieb:2015:WFN

[Got15] Steven Gottlieb. Whither the future of NSF Advanced Computing Infrastructure? Computing in Science and Engineering, 17(2): 4–6, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/02/mcs2015020004.
                                        html.

Gottlieb:2016:FNA

[Got16] Steven Gottlieb. The future of NSF Advanced Computing Infrastructure revisited. Computing in Science and Engineering, 18(5):4–7, September/October 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gottlieb:2017:CNC

[Got17] Steven Gottlieb. Changes at the NSF and computing in the physics curriculum. Computing in Science and Engineering, 19(3): 4–5, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs2017030004.
html.

Goncalves:2015:MSH

[GP15] Bernardo Goncalves and Fabio Porto. Managing scientific hypotheses as data with support for predictive analytics. Computing in Science and Engineering, 17(5):35–43, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs2015050035-
abs.html.

Granger:2021:JTS

[GP21] B. E. Granger and F. Pérez.

Jupyter: Thinking and storytelling with code and data. Computing in Science and Engineering, 23(2): 7–14, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gomez-Perez:2008:PSM

[GPC08] Jose Manuel Gómez-Pérez and Oscar Corcho. Problemsolving methods for understanding process executions. Computing in Science and Engineering, 10(3):47–52, May/ June 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Glotzer:2009:COP

[GPL09] Sharon C. Glotzer, Bob Panoff, and Scott Lathrop. Challenges and opportunities in preparing students for petascale computational science and engineering. Computing in Science and Engineering, 11(5):22–27, September/October 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Goode:2020:OPD

[GPMSC20] J. Goode, K. Peterson, J. Malyn-Smith, and G. Chapman. Online professional development for high school computer science teachers: Features that support an equity-based professional learning community. Computing in Science and Engineering, 22(5):51–59, September/October 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gombosi:2004:SAM

[GPZ^{+}04] Tamas I. Gombosi, Kenneth G. Powell, Darren L. De Zeeuw, C. Robert Clauer, Kenneth C. Hansen, Ward B. Manchester, Aaron J. Ridley, Ilia I. Roussev, Igor V. Sokolov, Quentin F. Stout, and Gábor Tóth. Solution-adaptive magnetohydrodynamics for space plasmas: Sun-to-Earth simulations. Computing in Science and Engineering, 6(2): 14–35, March/April 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/02/c2014abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2004/02/c2014.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/02/
c2014.pdf.

Gray:2007:DES

[Gra07] Michael A. Gray. Discrete event simulation: a review of SimEvents. Computing in Science and Engineering, 9(6): 62–66, November/December 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gray:2008:MT

[Gra08a] Michael A. Gray. Multiphysics technologies. Computing in Science and Engineering, 10 (2):82–84, March/April 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gray:2008:SNM

[Gra08b] Michael A. Gray. Sage: a new mathematics software system. Computing in Science and Engineering, 10(6): 72–75, November/December 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gray:2009:GSG

[Gra09] Michael A. Gray. Getting started with GPU programming. Computing in Science and Engineering, 11(4): 61–64, July/August 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gray:1999:SPS

[GRE99] Mark G. Gray, Randy M. Roberts, and Tom M. Evans. Scientific programming: Shadowobject interface between Fortran 95 and C++. Computing in Science and Engineering, 1(2):63–70, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c2063.
                                        pdf.

Greenfield:2007:RSP

[Gre07] Perry Greenfield. Reaching for the stars with Python. Computing in Science and Engineering, 9(3):38–40, May/ June 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gropp:2009:SPC

[Gro09] William D. Gropp. Software for petascale computing systems. Computing in Science and Engineering, 11(5):17– 21, September/October 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gilbert:2008:UFN

[GRS08] John R. Gilbert, Steve Reinhardt, and Viral B. Shah. A unified framework for numerical and combinatorial computing. Computing in Science and Engineering, 10(2): 20–25, March/April 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Goedecker:2003:LSE

[GS03] Stefan Goedecker and Gustavo E. Scuseria. Linear scaling electronic structure methods in chemistry and physics. Computing in Science and Engineering, 5(4): 14–21, July/August 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.

computer.org/comp/mags/
cs/2003/04/c4014abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/04/c4014.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/04/
c4014.pdf.

Gottlieb:2013:ECG

[GS13a] Steven Gottlieb and Thomas Sterling. Exascale computing [Guest editorial]. Computing in Science and Engineering, 15(6):12–15, November/December 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gropp:2013:PEC

[GS13b] William Gropp and Marc Snir. Programming for exascale computers. Computing in Science and Engineering, 15(6): 27–35, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gorton:2012:VKM

[GSB^{+}12] Ian Gorton, Chandrika Sivaramakrishnan, Gary Black, Signe White, Sumit Purohit, Carina Lansing, Michael Madison, Karen Schuchardt, and Yan Liu. Velo: a knowledge-management framework for modeling and simulation. Computing in Science and Engineering, 14(2): 12–23, March/April 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gesing:2023:HCG

[GSK^{+}23] Sandra Gesing, Claire Stirm, Gerhard Klimeck, Lynn Zentner, Su Wang, Braulio M. Villegas-Martinez, Hector M. Moya-Cessa, Carrie Diaz Eaton, Sam Donovan, Carol Song, Lan Zhao, I Luk Kim, Alejandro Strachan, Michael Zentner, and Rajesh Kalyanam. Hubzero: Community growth for four science gateways supporting open science. Computing in Science and Engineering, 25(1):34–42, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Guelton:2018:PCP

[Gue18]  Serge Guelton.  Pythran:

Crossing the Python frontier. Computing in Science and Engineering, 20 (2):83–89, March/April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8317992/.

Guo:2012:CTC

[Guo12] Philip Guo. CDE: a tool for creating portable experimental software packages. Computing in Science and Engineering, 14(4):32– 35, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Guo:2023:SOS

[Guo23] Philip J. Guo. Six opportunities for scientists and engineers to learn programming using AI tools such as Chat-GPT. Computing in Science and Engineering, 25(3): 73–78, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Guttmann:2001:ESM

[Gut01] Anthony Guttmann. Enumerations in statistical mechanics and combinatorics. Computing in Science and Engineering, 3(3):42–47, May/ June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2001/pdf/c3042.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999c3042abs.
                                        htm.

Goncalves:2015:MLV

[GVB15] Andre R. Goncalves, Fernando J. Von Zuben, and Arindam Banerjee. A multitask learning view on the earth system model ensemble. Computing in Science and Engineering, 17(6):35–42, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gorrell:2007:UUF

[GvdWT07] Steven E. Gorrell, Allan van de Wall, and Fu-Lin Tsung. Understanding unsteady flow features in transonic compressors. Computing in Science and Engineering, 9(6):12–17, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Goldstein:2015:DBS

[GW15] Rhys Goldstein and Gabriel A.

Wainer. Designing biological simulation models using formalism-based functional and spatial decompositions. Computing in Science and Engineering, 17(6): 72–82, November/December 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Grimstead:2007:AMV

[GWA^{+}07] Ian J. Grimstead, David W.

Walker, Nick J. Avis, Frederic Kleinermann, and John Mc-Clure. 3D anatomical model visualization within a gridenabled environment. Computing in Science and Engineering, 9(5):32–38, September/October 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gisler:2004:TTD

[GWMG04] Galen R. Gisler, Robert P.

Weaver, Charles L. Mader, and Michael L. Gittings. Twoand three-dimensional asteroid impact simulations. Computing in Science and Engineering, 6(3):46–55, May/ June 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2004/03/c3046abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/03/
c3046.htm;   http://csdl.
computer.org/dl/mags/cs/
2004/03/c3046.pdf.

Guyer:2009:FPD

[GWW09] Jonathan E. Guyer, Daniel Wheeler, and James A. Warren. FiPy: Partial differential equations with Python. Computing in Science and Engineering, 11(3):6–15, May/ June 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Guo:2010:UGF

[GYF^{+}10] Xiang Guo, Dong Yan, Jianrong Fan, Wanze Zhu, and Mai-He Li. Using GIS and fuzzy sets to evaluate the olive tree’s ecological suitability in Sichuan Province. Computing in Science and Engineering, 12 (1):20–27, January/February 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Guo:2020:CAH

[GYJL20] C. Guo, Y. Yang, Y. Jiang, and T. Li. Condition analysis of a high-speed train based on similarity ratio and MDBN. Computing in Science and Engineering, 22(1):64–76, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Gao:2017:ROL

[GYL^{+}17] Shangbing Gao, Qiaolin Ye, Limin Luo, Zhigeng Pan, Yunyang Yan, and Hao Zheng. Recursive orthogonal label regression: A framework for semisupervised dimension reduction. Computing in Science and Engineering, 19(4): 30–43, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/04/mcs2017040030-
abs.html.

Gyure:1999:CIB

[Gyu99] Mark F. Gyure. CiSE in industry: Bridging time and length scales in semiconductor process model development. Computing in Science and Engineering, 1(5): 100–103, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5100.
pdf;   http://www.computer.
org/cse/cs1999/c5100abs.
htm.

Groen:2014:SMM

[GZC14] Derek Groen, Stefan J. Zasada, and Peter V. Coveney. Survey of multiscale and multiphysics applications and communities. Computing in Science and Engineering, 16 (2):34–43, March/April 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Haas:1999:CIW

[Haa99] Ken Haas. CSE in industry: Where the rubber meets the road. Computing in Science and Engineering, 1(1): 4–6, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c1004.
                                        pdf.

Habibkhah:2017:SSA

[HAB17] Shahnaz Habibkhah, Jalal Arasi, and Hossein Bolandi. SPACSSIM: Simulation and analysis software for mathematical modeling of satellite position and attitude control systems. Computing in Science and Engineering, 19(5):38–48, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/05/mcs2017050038-
                                        abs.html.

Hahn:2004:NEP

[Hah04]  Marjorie G. Hahn.  New

exposition on probability and statistics. Computing in Science and Engineering, 6(1):85–88, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/01/c1085abs.htm;
http://csdl.computer.org/
dl/mags/cs/2004/01/c1085.
pdf.

Hailperin:2010:B

[Hai10] Max Hailperin. Books. Computing in Science and Engineering, 12(2):8–11, March/ April 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hansmann:2003:PFS

[Han03] Ulrich H. E. Hansmann. Protein folding in silico: An overview. Computing in Science and Engineering, 5 (1):64–69, January/February 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/01/c1064.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/01/
c1064.pdf.

Hansen:2005:PF

[Han05] A. Hansen. Physics and fracture. Computing in Science and Engineering, 7(5):90– 95, September/October 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501748.pdf?
isnumber=32219&prod=JNL&
arnumber=1501748&arSt=+
90&ared=+95&arAuthor=+Hansen%
2C+A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=32219&arnumber=
1501748&count=14&index=
12.

Harding:2004:MGD

[Har04a] Chris Harding. Modeling geoscience data in a multisensory virtual environment. Computing in Science and Engineering, 6(1):89–92, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/01/c1089abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2004/01/c1089.
pdf.

Harris:2004:N

[Har04b] Lissa E. Harris. News. Computing in Science and Engineering, 6(5):4–7, September/October 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/05/c5004.pdf;
http://csdl.computer.org/

dl/mags/cs/2004/05/c5004.
htm.

Harken:2018:UMI

[Har18] Rachel Harken. Uncovering magic isotopes with the power of HPC. Computing in Science and Engineering, 20(4):107– 110, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040107.
html.

Harrod:2023:IAR

[Har23] William J. Harrod. The intelligence advanced research projects activity advanced graph intelligent logical computing environment program: Reinventing computing. Computing in Science and Engineering, 25(4):12–24, April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hasbun:2008:HDN

[Has08] Javier E. Hasbun. How do nerve cells compute? Computing in Science and Engineering, 10(3):64–65, May/ June 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hasbun:2012:UTP

[Has12] Javier E. Hasbun. Unifying two popular-but-seemingly-dissimilar platforms: Matlab and Java. Computing in Science and Engineering, 14(3):

6–7, May/June 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hendrickson:2008:GAH

[HB08] Bruce Hendrickson and Jonathan W. Berry. Graph analysis with high-performance computing. Computing in Science and Engineering, 10(2):14– 19, March/April 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ham:2008:BIT

[HBB08] Kyungmin Ham, Heath A. Barnett, and Leslie G. Butler. Burning issues in tomography analysis. Computing in Science and Engineering, 10(2): 78–81, March/April 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hamilton:2023:SAA

[HBD^{+}23] Kathryn R. Hamilton, Klaus Bartschat, Nicolas Douguet, Sudhakar V. Pamidighantam, and Barry I. Schneider. Simulation for all: The atomic, molecular, and optical science gateway. Computing in Science and Engineering, 25(3): 68–72, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hsu:2020:CUR

[HBG^{+}20] C. Hsu, A. E. Bandrowski, T. H. Gillespie, J. Udell, K. Lin, I. B. Ozyurt, J. S. Grethe, and M. E. Martone. Comparing the use of research resource identifiers and natural language processing for citation of databases, software, and other digital artifacts. Computing in Science and Engineering, 22(2): 22–32, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Haney:1999:SPH

[HC99] Scott Haney and James Crotinger. Scientific programming: How templates enable high-performance scientific computing in C++. Computing in Science and Engineering, 1(4):66–72, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c4066.
pdf.

Hu:2017:FLS

[HC17] Helen H. Hu and Patricia B.

Campbell. A framework for levels of student participation and stages of relevant curriculum. Computing in Science and Engineering, 19(3): 20–29, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.

                                        computer.org/csdl/mags/
                                        cs/2017/03/mcs2017030020-
                                        abs.html.

Heroux:2022:CWS

[HCK22] Michael A. Heroux, Jeffrey C. Carver, and Sarah Knepper. Collegeville Workshop 2021: Scientific software teams. Computing in Science and Engineering, 24(3): 4–5, May/June 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Heath:2000:VPS

[HD00] Michael T. Heath and William A. Dick. Virtual prototyping of solid propellant rockets. Computing in Science and Engineering, 2(2): 21–32, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c2021.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c2021abs.
                                        htm.

Hill:2004:AES

[HDB^{+}04] Chris Hill, Cecelia DeLuca, Balaji, Max Suarez, and Arlindo da Silva. The architecture of the Earth System Modeling Framework. Computing in Science and Engineering, 6(1):18–28, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.

computer.org/comp/mags/
cs/2004/01/c1018abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2004/01/c1018.
pdf.

Hoef-Emden:2005:MPA

[HE05] Kerstin Hoef-Emden. Molecular phylogenetic analyses and real-life data. Computing in Science and Engineering, 7(3): 86–91, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/03/c3086abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2005/03/c3086.
pdf.

Hacker:2011:NCE

[HEB^{+}11] Thomas J. Hacker, Rudi Eigenmann, Saurabh Bagchi, Ayhan Irfanoglu, Santiago Pujol, Ann Catlin, and Ellen Rathje. The NEEShub cyberinfrastructure for earthquake engineering. Computing in Science and Engineering, 13 (4):67–78, July/August 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Horner:2010:IHF

[HEH^{+}10] David Horner, Owen Eslinger, Stacy Howington, Steve Ketcham, John Peters, and Jerry Ballard. Integrated high-fidelity geoscience simulations for enhanced terrainrelated target detection. Computing in Science and Engineering, 12(5):56–63, September/October 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Heinonen:2020:PRC

[Hei20] N. Heinonen. Predicting the risk of cancer with computational electrodynamics. Computing in Science and Engineering, 22(1):117–120, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Heimbach:2022:CSK

[Hei22] Patrick Heimbach. The computational science of Klaus Hasselmann. Computing in Science and Engineering, 24 (4):40–53, July/August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hemmert:2010:GHN

[Hem10] Scott Hemmert. Green HPC: From nice to necessity. Computing in Science and Engineering, 12(6):8–10, November/December 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Heroux:2021:CWS

[Her21] M. A. Heroux. The Collegeville Workshops on Scientific Software: Looking back and forward. Computing in Science and Engineering, 23 (4):69–72, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Heroux:2022:RSS

[Her22]  Michael A. Heroux.  Re-

search software science: Expanding the impact of research software engineering. Computing in Science and Engineering, 24(6):22–27, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hansen:2004:GCS

[HF04] Robert Hansen and James Forsythe. A grid convergence study of a highly separated turbulent flow. Computing in Science and Engineering, 6(6): 30–37, November/December 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/06/c6030.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/06/
c6030.pdf.

Holmes:2000:EMP

[HG00] Raquell Holmes and Roscoe Giles. Essay: Minority participation in computational science. Computing in Science and Engineering, 2(2): 11–13, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c2011.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c2011abs.
                                        htm.

Holland:2002:GEI

[HG02] Charles J. Holland and John Grosh. Guest Editors’ introduction: High-performance computing: Addressing defense needs of today and tomorrow. Computing in Science and Engineering, 4 (2):11–14, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c2011abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2002/pdf/c2011.pdf.

Holmes:2023:DEI

[HGA23] Raquell Holmes, Roscoe Giles, and Dorian Arnold. Diversity, equity, and inclusion for computer and information science and engineering conferences: How change happens and four things you can do now. Computing in Science and Engineering, 25(1):57–60, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Herbordt:2008:CMF

[HGV^{+}08] Martin C. Herbordt, Yongfeng Gu, Tom VanCourt, Josh Model, Bharat Sukhwani, and Matt Chiu. Computing models for FPGA-based accelerators. Computing in Science and Engineering, 10(6): 35–45, November/December 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hargrove:1999:UMC

[HH99] William W. Hargrove and Forrest M. Hoffman. Using multivariate clustering to define ecoregion borders. Computing in Science and Engineering, 1(4):18–25, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c4018.
pdf.

Hennenfent:2006:SDN

[HH06] Gilles Hennenfent and Felix J.

Herrmann. Seismic denoising with nonuniformly sampled curvelets. Computing in Science and Engineering, 8(3): 16–25, May/June 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hammarling:2022:ICJ

[HH22] Sven Hammarling and Nicholas J.

Higham. The influence and contribution of Jack Dongarra to numerical linear algebra. Computing in Science and Engineering, 24(4): 6–11, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Heitmann:2014:LSS

[HHF^{+}14] Katrin Heitmann, Salman Habib, Hal Finkel, Nicholas Frontiere, Adrian Pope, Vitali Morozov, Steve Rangel, Eve Kovacs, Juliana Kwan, Nan Li, Silvio Rizzi, Joe Insley, Venkatram Vishwanath, Tom Peterka, David Daniel, Patricia Fasel, and George Zagaris. Large-scale simulations of sky surveys. Computing in Science and Engineering, 16(5):14–23, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2014/05/mcs2014050014-
                                        abs.html.

Heinonen:2019:SSS

[HHP19] Nils Heinonen, James J. Hack, and Michael E. Papka. Software stack in a snapshot. Computing in Science and Engineering, 21 (3):114–116, May/June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8701539/.

Hamalainen:2002:UMM

[HHR02] Juhani Hämäläinen, Kari Hirvi, and Hannu Rajaniemi. Using mathematical models to cope with complex computer simulations. Computing in Science and Engineering, 4(1):64–72, January/February 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c1064abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c1064.pdf.

Howe:2013:CSW

[HHR^{+}13] Bill Howe, Daniel Halperin, François Ribalet, Sagar Chitnis, and E. Virginia Armbrust. Collaborative science workflows in SQL. Computing in Science and Engineering, 15(3):22–31, May/ June 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hannemann:2001:SPSb

[HHZK01a] Jens Hannemann, Regina Hannemann, Michael Zellerhoff, and Ludger Klinkenbusch. Scientific programming: Scientific programming in field theory,part 2. Computing in Science and Engineering, 3(4):78–85, July/August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c4078.
pdf.

Hannemann:2001:SPSa

[HHZK01b] Regina Hannemann, Jens Hannemann, Michael Zellerhoff, and Ludger Klinkenbusch. Scientific programming: Scientific programming in field theory, part 1. Computing in Science and Engineering, 3(3):66–74, May/ June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3066.
pdf;   http://www.computer.
org/cse/cs1999c3066abs.
htm.

Higham:2004:BSS

[Hig04] Desmond J. Higham. Black– Scholes for scientific computing students. Computing in Science and Engineering, 6(6): 72–79, November/December 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/06/c6072.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/06/
c6072.pdf.

Hill:2015:PRN

[Hil15] David R. C. Hill. Parallel random numbers, simulation, and reproducible research. Computing in Science and Engineering, 17(4): 66–71, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040066-
abs.html.

Hinsen:2007:PSP

[Hin07]  Konrad Hinsen.  Parallel

scripting with Python. Computing in Science and Engineering, 9(6):82–89, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinsen:2009:PFP

[Hin09] Konrad Hinsen. The promises of functional programming. Computing in Science and Engineering, 11(4):86–90, July/ August 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinsen:2012:CYD

[Hin12a] Konrad Hinsen. Caring for your data. Computing in Science and Engineering, 14(6): 70–74, November/December 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinsen:2012:MS

[Hin12b] Konrad Hinsen. Managing state. Computing in Science and Engineering, 14(1):80–86, January/February 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinsen:2013:DAS

[Hin13a] Konrad Hinsen. Daydreaming about scientific programming. Computing in Science and Engineering, 15(5):77– 79, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinsen:2013:GFS

[Hin13b] Konrad Hinsen. A glimpse of the future of scientific programming. Computing in Science and Engineering, 15 (1):84–88, January/February 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinsen:2013:SDR

[Hin13c] Konrad Hinsen. Software development for reproducible research. Computing in Science and Engineering, 15(4): 60–63, July/August 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinsen:2015:ATC

[Hin15a] Konrad Hinsen. The approximation tower in computational science: Why testing scientific software is difficult. Computing in Science and Engineering, 17(4): 72–77, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040072-
abs.html.

Hinsen:2015:TDC

[Hin15b] Konrad Hinsen. Technical debt in computational science.

Computing in Science and Engineering, 17(6):103–107, November/December 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinsen:2015:WSS

[Hin15c] Konrad Hinsen. Writing software specifications. Computing in Science and Engineering, 17(3):54–61, May/ June 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/csdl/
mags/cs/2015/03/mcs2015030054-
abs.html.

Hinsen:2016:PCC

[Hin16] Konrad Hinsen. The power to create chaos. Computing in Science and Engineering, 18 (4):75–79, July/August 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hines:2017:MBP

[Hin17a]  Jonathan Hines.  Master-

ing a bacterial photosynthetic system. Computing in Science and Engineering, 19(6):82–84, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060082.
html.

Hinsen:2017:DSS

[Hin17b] Konrad Hinsen. A dream of simplicity: Scientific computing on Turing machines. Computing in Science and Engineering, 19(3):78–85, May/ June 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://www.computer.org/
                                        csdl/mags/cs/2017/03/mcs2017030078-
                                        abs.html.

Hinsen:2017:RCC

[Hin17c] Konrad Hinsen. The roles of code in computational science. Computing in Science and Engineering, 19(1):78–82, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/01/mcs2017010078-
                                        abs.html.

Hines:2018:MPS

[Hin18a] Jonathan Hines. Mixed precision: A strategy for new science opportunities. Computing in Science and Engineering, 20(6):67–71, November/December 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/06/08625902-abs.
                                        html.

Hines:2018:SS

[Hin18b] Jonathan Hines. Stepping up to Summit. Computing in Science and Engineering, 20(2): 78–82, March/April 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinsen:2018:DSL

[Hin18c] Konrad Hinsen. Domainspecific languages in scientific computing. Computing in Science and Engineering, 20 (1):88–92, January/February 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://ieeexplore.ieee.
org/document/8291790/.

Hinsen:2018:RVR

[Hin18d] Konrad Hinsen. Reusable versus re-editable code. Computing in Science and Engineering, 20(3):78–83, May/ June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8358032/.

Hinsen:2019:DSC

[Hin19] Konrad Hinsen. Dealing with software collapse. Computing in Science and Engineering, 21(3):104–108, May/June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8701540/.

Hinsen:2020:MCA

[Hin20a]  K. Hinsen.  The magic

of content-addressable storage. Computing in Science and Engineering, 22(3):113– 119, May/June 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinsen:2020:SCT

[Hin20b] K. Hinsen. Staged computation: The technique you did not know you were using. Computing in Science and Engineering, 22(4):99– 103, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinsen:2023:NCM

[Hin23] Konrad Hinsen. The nature of computational models. Computing in Science and Engineering, 25(1):61–66, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hege:2003:CTF

[HJLH03] E. Keith Hege, Stuart M. Jefferies, and Michael Lloyd-Hart. Computing and telescopes at the frontiers of optical astronomy. Computing in Science and Engineering, 5(6):42–51, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/06/c6042abs.htm;
 http://csdl.computer.org/

dl/mags/cs/2003/06/c6042.
pdf.

Hook:2009:MST

[HK09] Daniel Hook and Diane Kelly.

Mutation sensitivity testing. Computing in Science and Engineering, 11(6): 40–47, November/December 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Heinecke:2012:GMC

[HKB12] Alexander Heinecke, Michael Klemm, and Hans-Joachim Bungartz. From GPGPU to many-core: Nvidia Fermi and Intel Many Integrated Core architecture. Computing in Science and Engineering, 14(2): 78–83, March/April 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinneburg:2003:UPV

[HKW03] Alexander Hinneburg, Daniel Keim, and Markus Wawryniuk. Using projections to visually cluster high-dimensional data. Computing in Science and Engineering, 5(2): 14–25, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/02/c2014abs.htm;
http://csdl.computer.org/
dl/mags/cs/2003/02/c2014.
htm;   http://csdl.computer.

org/dl/mags/cs/2003/02/
c2014.pdf.

Hartel:2000:ESI

[HL00] Hermann Härtel and Michael Lüdke. Education: 3D simulations of interacting particles. Computing in Science and Engineering, 2(4): 87–90, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c4087.
pdf.

Haugh:2001:CCP

[HL01] Martin B. Haugh and Andrew W. Lo. Computational challenges in portfolio management. Computing in Science and Engineering, 3(3):54–59, May/June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3054.
pdf;   http://www.computer.
org/cse/cs1999c3054abs.
htm.

Hale:2017:CPP

[HLRW17] Jack S. Hale, Lizao Li, Christopher N. Richardson, and Garth N. Wells. Containers for portable, productive, and performant scientific computing. Computing in Science and Engineering, 19(6):40–50, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060040-
abs.html.

Hallissy:2016:HCA

[HLS^{+}16] B. P. Hallissy, J. P. Laiosa, T. C. Shafer, D. H. Hine, J. R. Forsythe, J. Abras, N. S. Hariharan, and C. Dahl. HPCMP CREATE-AV quality assurance: Lessons learned by validating and supporting computation-based engineering software. Computing in Science and Engineering, 18 (1):52–62, January/February 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hinsen:2009:ETV

[HLT09] Konrad Hinsen, Konstantin Laufer, and George K. Thiruvathukal. Essential tools: Version control systems. Computing in Science and Engineering, 11(6):84–91, November/December 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

He:2019:SFS

[HLYQ19] W. He, Z. Li, S. Yang, and W. Quan. SV-fvdm: A synthetic vision based full velocity difference model for interactive traffic simulation. Computing in Science and Engineering, 21(5):35–45, September/October 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Haber:2000:UTM

[HMA00] Idith Haber, Dimitris N. Metaxas, and Leon Axel. Using tagged MRI to reconstruct a 3D heartbeat. Computing in Science and Engineering, 2(5): 18–30, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c5018.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c5018abs.
                                        htm.

Humphrey:2014:SDM

[HMB^{+}14] Alan Humphrey, Qingyu Meng, Martin Berzins, Diego Caminha B. de Oliveira, Zvonimir Rakamaric, and Ganesh Gopalakrishnan. Systematic debugging methods for large-scale HPC computational frameworks. Computing in Science and Engineering, 16(3):48–56, May/June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Henderson:2000:SAF

[HMS^{+}00] Thomas C. Henderson, Patrick A. McMurtry, Philip J. Smith, Gregory A. Voth, Charles A. Wight, and David F. Pershing. Simulating accidental fires and explosions. Computing in Science and Engineering, 2(2):64–76, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c2064.
pdf;   http://www.computer.
org/cse/cs1999/c2064abs.
htm.

Hoefler:2010:SHT

[Hoe10] Torsten Hoefler. Software and hardware techniques for power-efficient HPC networking. Computing in Science and Engineering, 12(6): 30–37, November/December 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Healey:1999:WOC

[HOÖ99] Dennis Healey, Tim Olson, and Ulf Österberg. Wavelets in optical communications. Computing in Science and Engineering, 1 (1):51–57, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c1051.
pdf;   http://www.computer.
org/cse/cs1999/c1051abs.
htm.

Howe:2012:VAC

[How12] Bill Howe. Virtual appliances, cloud computing, and reproducible research. Computing in Science and Engineering, 14 (4):36–41, July/August 2012.

CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Holland:2004:GEI

[HP04] Charles J. Holland and Robert E. Peterkin, Jr. Guest Editors’ introduction: High-performance computing. Computing in Science and Engineering, 6(6):8–10, November/December 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/06/c6008.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/cs/2004/06/c6008.
                                        htm.

Hack:2014:ALC

[HP14a] James J. Hack and Michael E. Papka. Advances in leadership computing [Guest Editors’ introduction]. Computing in Science and Engineering, 16(5):10–12, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2014/05/mcs2014050010.
                                        html.

Hack:2014:NFL

[HP14b] James J. Hack and Michael E. Papka. New frontiers in leadership computing [Guest editorial]. Computing in Science and Engineering, 16(6): 10–12, November/December 2014. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/csdl/
mags/cs/2014/06/mcs2014060010.
html.

Hack:2015:BDN

[HP15] James J. Hack and Michael E.

Papka. Big data: Nextgeneration machines for big science. Computing in Science and Engineering, 17(4): 63–65, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040063-
abs.html.

Hack:2020:UHP

[HP20] J. J. Hack and M. E. Papka.

The U.S. High-Performance Computing Consortium in the fight against COVID-

19. Computing in Science and Engineering, 22(6):75–80, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hwang:2020:AIO

[HPC20] L. J. Hwang, R. A. Pauloo, and J. Carlen. Assessing the impact of outreach through software citation for community software in geodynamics. Computing in Science and Engineering, 22(1):16–25, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hoffmann:2004:MSV

[HPKS04] Chris Hoffmann, Voicu Popescu, Sami Kilic, and Mete Sozen. Modeling, simulation, and visualization: The Pentagon on September 11th. Computing in Science and Engineering, 6(1):52–60, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/01/c1052abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2004/01/c1052.
                                        pdf.

Howard:2012:ACS

[HPMJ12] Jessica Howard, Omar Padron, Patricia Morreale, and David Joiner. Applications of computational science: Dataintensive computing for student projects. Computing in Science and Engineering, 14 (2):84–89, March/April 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hatcher:2005:CCJ

[HRAB05] Philip Hatcher, Mathew Reno, Gabriel Antoniu, and Luc Bougé. Cluster computing with Java. Computing in Science and Engineering, 7(2):34–39, March/April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/02/c2034.

htm;   http://csdl.computer.
org/dl/mags/cs/2005/02/
c2034.pdf.

Hwu:2009:CUD

[HRRS09] Wen-Mei Hwu, Christopher Rodrigues, Shane Ryoo, and John Stratton. Compute unified device architecture application suitability. Computing in Science and Engineering, 11(3):16–26, May/ June 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Haist:2006:UGB

[HRWS06] Tobias Haist, Marcus Reicherter, Min Wu, and Lars Seifert. Using graphics boards to compute holograms. Computing in Science and Engineering, 8(1):8–13, January/February 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hase:2003:GEI

[HS03] William L. Hase and Gustavo E. Scuseria. Guest Editors’ introduction: Computational chemistry. Computing in Science and Engineering, 5(4):12–13, July/August 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/04/c4012.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/04/
c4012.pdf.

Heien:2012:ULT

[HS12] Eric M. Heien and Michael Sachs. Understanding longterm earthquake behavior through simulation. Computing in Science and Engineering, 14(5):10–20, September/October 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hase:2003:DDS

[HSG03] William L. Hase, Kihyung Song, and Mark S. Gordon. Direct dynamics simulations. Computing in Science and Engineering, 5(4): 36–44, July/August 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/04/c4036abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/04/c4036.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/04/
                                        c4036.pdf.

Hua:2019:TOD

[HSJ^{+}19] Z. Hua, Z. Sun, H. Jia, Z. Zhao, and L. Hou. Track occupation detection based on a maximum posterior probability model using multisensor data fusion. Computing in Science and Engineering, 21(6):40–54, November/December 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hoefler:2023:EVE

[HSP^{+}23] Torsten Hoefler, Bjorn Stevens, Andreas F. Prein, Johanna Baehr, Thomas Schulthess, Thomas F. Stocker, John Taylor, Daniel Klocke, Pekka Manninen, Piers M. Forster, Tobias Kölling, Nicolas Gruber, Hartwig Anzt, Claudia Frauen, Florian Ziemen, Milan Klöwer, Karthik Kashinath, Christoph Schär, Oliver Fuhrer, and Bryan N. Lawrence. Earth virtualization engines: a technical perspective. Computing in Science and Engineering, 25 (3):50–59, March 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hsu:2006:CHD

[Hsu06] Feng-Hsiung Hsu. Chess hardware in Deep Blue. Computing in Science and Engineering, 8(1):50–60, January/February 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Helbling:1999:CSN

[HT99] Dirk Helbling and Martin Treiber. Computer simulations: Numerical simulation of macroscopic traffic equations. Computing in Science and Engineering, 1(5): 89–99, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5089.

                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c5089abs.
                                        htm.

Hu:2007:IMR

[Hu07]  Chenglie Hu.  Integrating

modern research into numerical computation education. Computing in Science and Engineering, 9(5):78–81, September/October 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hunter:2007:MGE

[Hun07] John D. Hunter. Matplotlib: a 2D graphics environment. Computing in Science and Engineering, 9(3):90–95, May/ June 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Husowitz:2022:MPC

[Hus22] Barry C. Husowitz. Modeling of phase coexistence via thermodynamic potentials. Computing in Science and Engineering, 24(1):72–77, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Henning:2009:TR

[HW09] Paul Henning and Andrew B. White, Jr. Trailblazing with Roadrunner. Computing in Science and Engineering, 11 (4):91–95, July/August 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Heinrich:2015:PCM

[HW15] Julian Heinrich and Daniel Weiskopf. Parallel coordinates for multidimensional data visualization: Basic concepts. Computing in Science and Engineering, 17(3): 70–76, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/03/mcs2015030070-
abs.html.

Hariharan:2016:FPP

[HWPS16] Nathan Hariharan, Andrew Wissink, Mark Potsdam, and Roger Strawn. First-principles physics-based rotorcraft flowfield simulation using HPCMP CREATE-AV Helios. Computing in Science and Engineering, 18(6):19–26, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs2016060019-
abs.html.

Huang:2005:ABS

[HXMC05] Yingping Huang, Xiaorong Xiang, Gregory Madey, and Steve E. Cabaniss. Agentbased scientific simulation. Computing in Science and Engineering, 7(1):22–29, January/February 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/01/c1022.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2005/01/
                                        c1022.pdf.

Hyman:2005:PDM

[Hym05] James M. Hyman. Patch dynamics for multiscale problems. Computing in Science and Engineering, 7(3): 47–53, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2005/03/c3047abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2005/03/c3047.
                                        pdf.

Iacono:2021:BEF

[Iac21] Roberto Iacono. Bounding the error function. Computing in Science and Engineering, 23(4):65–68, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Iyengar:2003:MCP

[IBPV03] Srinivasan S. Iyengar, Christian J. Burnham, Matt K. Petersen, and Gregory A. Voth. Modeling condensedphase chemistry through molecular dynamics simulation. Computing in Science and Engineering, 5(4): 31–35, July/August 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.

computer.org/comp/mags/
cs/2003/04/c4031abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/04/c4031.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/04/
c4031.pdf.

Iskandarani:2002:MGM

[IHL^{+}02] Mohamed Iskandarani, Dale B.

Haidvogel, Julia C. Levin, Enrique Curchitser, and Christopher A. Edwards. Multiscale geophysical modeling using the spectral element method. Computing in Science and Engineering, 4(5):42–48, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/05/c5042abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/05/c5042.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/05/
c5042.pdf.

Isenor:2005:MGO

[IK05] A. W. Isenor and J. R.

Keeley. Modeling generic oceanographic data objects in XML. Computing in Science and Engineering, 7 (4):58–66, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463137.pdf?
isnumber=31456&prod=JNL&
arnumber=1463137&arSt=+

                                        58&ared=+66&arAuthor=Isenor%
                                        2C+A.W.%3B+Keeley%2C+J.
                                        R.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=31456&arnumber=
                                        1463137&count=14&index=7.

Ikkala:2016:CMS

[Ikk16] Kimmo Ikkala. A conceptual modeling and simulation framework for system design. Computing in Science and Engineering, 18(4): 42–52, July/August 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ivanova:2013:DVD

[IKMK13] Milena Ivanova, Martin Kersten, Stefan Manegold, and Yagiz Kargin. Data vaults: Database technology for scientific file repositories. Computing in Science and Engineering, 15(3):32–42, May/ June 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jacobson:2003:Nb

[Jac03] Anne Jacobson. News. Computing in Science and Engineering, 5(5):6–9, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/05/c5006abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/05/c5006.
                                        htm;   http://csdl.computer.

org/dl/mags/cs/2003/05/
c5006.pdf.

JaJa:2000:PQ

[JaJ00] Joseph JaJa. A perspective on quicksort. Computing in Science and Engineering, 2 (1):43–49, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1043.
pdf;   http://www.computer.
org/cse/cs1999/c1043abs.
htm.

Javidpour:2012:CSP

[Jav12] Leili Javidpour. Computer simulations of protein folding. Computing in Science and Engineering, 14(2):97– 103, March/April 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jarvis:2002:RRE

[JC02] Emily A. Jarvis and Emily A.

Carter. The role of reactive elements in the bond coat for thermal barrier coatings. Computing in Science and Engineering, 4(2): 33–41, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c2033abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c2033.pdf.

Jiang:2010:CDM

[JCC^{+}10] Yongquan Jiang, Li Chen, Qishu Chen, Qiang Peng, and Jim X. Chen. Computing discrete minimal surfaces using a nonlinear spring model. Computing in Science and Engineering, 12(6):74–79, November/December 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Juzna:2014:SSF

[JCPS14] Jernej Juzna, Peter Cesarek, Dana Petcu, and Vlado Stankovski. Solving solid and fluid mechanics problems in the cloud with mOSAIC. Computing in Science and Engineering, 16(3):68–77, May/ June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Johnson:2003:IT

[JD03] Jeffrey N. Johnson and Paul F. Dubois. Issue tracking. Computing in Science and Engineering, 5(6):71–77, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/cs/
                                        2003/06/c6071.pdf.

Jeffers:2021:CPW

[Jef21] Ann E. Jeffers. The COVID-19 pandemic is widening the gap for women in STEM. Computing in Science and Engineering, 23(3):96–98, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jeffers:2022:FCN

[Jef22] Ann E. Jeffers. The field of computing needs to take care of its mental health. Computing in Science and Engineering, 24(2):91–94, March/April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jerger:2013:EPA

[Jer13] Natalie Enright Jerger. Explaining parallel architecture design. Computing in Science and Engineering, 15(4): 6–7, July/August 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jacobson:2003:Na

[JG03] Anne Jacobson and Pam Frost Gorder. News. Computing in Science and Engineering, 5(4):3–7, July/August 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/04/c4003.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/04/
c4003.pdf.

Joliveau:2012:UBF

[JG12] Marc Joliveau and Michel Gendreau. Using bilevel feature extractors to reduce dimensionality in images. Computing in Science and Engineering, 14(3):60–67, May/ June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Johnson:2013:DHT

[JG13] James L. Johnson and Tom Goldring. Discrete Hodge theory on graphs: A tutorial. Computing in Science and Engineering, 15(5):42– 55, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

James-Hawkins:2016:WIF

[JH16] Laurie James-Hawkins. What influences female interest and persistence in computing?: Preliminary findings from a multiyear study. Computing in Science and Engineering, 18 (2):58–67, March/April 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Johanson:2018:SEC

[JH18] Arne Johanson and Wilhelm Hasselbring. Software engineering for computational science: Past, present, future. Computing in Science and Engineering, 20(2):90–109, March/April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8317991/.

Jaun:2001:ETC

[JHJ01] André Jaun, Johan Hedin, and Thomas Johnson. Education: Teaching computational methods for partial differential equations using the Web. Computing in Science and Engineering, 3(3):83–85, May/ June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3083.
pdf;   http://www.computer.
org/cse/cs1999c3083abs.
htm.

Jambeck:2015:CBL

[JJ15] Jenna R. Jambeck and Kyle Johnsen. Citizen-based litter and marine debris data collection and mapping. Computing in Science and Engineering, 17(4):20–26, July/ August 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040020-
abs.html.

Jiang:2010:UGQ

[JJZC10] Xiaobo Jiang, Wei Ji, Hongcheng Zeng, and Leiting Chen. Using GIS to quantify mountains in China. Computing in Science and Engineering, 12(1):36–43, January/February 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jha:2019:ISW

[JLNR19] S. Jha, S. Lathrop, J. Nabrzyski, and L. Ramakrishnan. Incorporating scientific workflows in computing research processes. Computing in Science and Engineering, 21(4): 4–6, July/August 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Jin:2010:MSN

[JLP^{+}10] Xin Jin, Mikel Lujan, Luis A. Plana, Sergio Davies, Steve Temple, and Steve B. Furber. Modeling spiking neural networks on SpiNNaker. Computing in Science and Engineering, 12(5):91–97, September/October 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jiang:2019:FSM

[JLYL19] Q. Jiang, D. Liang, F. Yan, and D. Liu. Force sensor model based on FEA for the electromagnetic levitation system. Computing in Science and Engineering, 21(6):20–25, November/December 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jones:2008:IEA

[JMEL08] Chad Jones, Kwan-Liu Ma, Stéphane Ethier, and Wei-Li Lee. An integrated exploration approach to visualizing multivariate particle data. Computing in Science and Engineering, 10(4):20–29, July/August 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Johnson:2001:MFL

[JMFJ01] Steven G. Johnson, Attila Mekis, Shanhui Fan, and John D. Joannopoulos. Molding the flow of light. Computing in Science and Engineering, 3(6):38–47, November/December 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c6038abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c6038.pdf.

Johnston:2006:ICC

[Joh06] Marty Johnston. Implementing curricular change. Computing in Science and Engineering, 8(5):32–37, September/October 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Johnson:2009:SC

[Joh09] James L. Johnson. SQL in the clouds. Computing in Science and Engineering, 11(4): 12–28, July/August 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Johnson:2012:BVC

[Joh12] Christopher Johnson. Biomedical visual computing: Case studies and challenges. Computing in Science and Engineering, 14(1):12–21, January/February 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jones:2015:TTU

[Jon15]  Katie Elyce Jones.  Ti-

tan takes on the universe: Large-scale cosmology simulation of universe mines for halos where galaxies are born. Computing in Science and Engineering, 17(3): 68–69, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/03/mcs2015030068-
                                        abs.html.

Jones:2019:SIP

[Jon19] K. E. Jones. Supercomputing improves predictions of fluid flow in rock. Computing in Science and Engineering, 21(6):74–76, November/December 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Juneau:2021:JEA

[JON^{+}21] S. Juneau, K. Olsen, R. Nikutta, A. Jacques, and S. Bailey. Jupyter-enabled astrophysical analysis using dataproximate computing platforms. Computing in Science and Engineering, 23(2): 15–25, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Johnson:2020:AFC

[JPE20] S. R. Johnson, A. Prokopenko, and K. J. Evans. Automated Fortran C++ bindings for large-scale scientific applications. Computing in Science and Engineering, 22(5):84– 94, September/October 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jones:2001:LCH

[JPK01] Samuel T. Jones, Scott E.

Parker, and Charlson C. Kim. Low-cost high-performance scientific visualization. Computing in Science and Engineering, 3(4):12–17, July/ August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://dlib.
computer.org/cs/books/cs2001/
pdf/c4012.pdf;   http://
www.computer.org/cse/cs1999c4012abs.
htm.

Joiner:2008:EOT

[JPMG08] David Joiner, Charles Peck, Thomas Murphy, and Paul Gray. Education, outreach, and training for highperformance computing. Computing in Science and Engineering, 10(5):40–45, September/October 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jin-qiang:2019:FPT

[Jq19] C. Jin-qiang. Fault prediction of a transformer bushing based on entropy weight TOP-SIS and gray theory. Computing in Science and Engineering, 21(6):55–62, November/December 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jensen:2010:ESE

[JR10] David Jensen and Arun Rodrigues. Embedded systems and exascale computing. Computing in Science and Engineering, 12(6):20–29, November/December 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Juve:2013:CFA

[JRD^{+}13] Gideon Juve, Mats Rynge, Ewa Deelman, Jens-S. Vockler, and G. Bruce Berriman. Comparing FutureGrid, Amazon EC2, and Open Science Grid for scientific workflows. Computing in Science and Engineering, 15(4): 20–29, July/August 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Johansen:2017:TEE

[JRP^{+}17] Hans Johansen, Arthur Rodgers, N. Anders Petersson, David McCallen, Bjorn Sjogreen, and Mamun Miah. Toward exascale earthquake ground motion simulations for near-fault engineering analysis. Computing in Science and Engineering, 19(5):27–37, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/05/mcs2017050027-
abs.html.

Joskowicz:1999:CAM

[JS99] Leo Joskowicz and Elisha Sacks. Computer-aided mechanical design using configuration spaces. Computing in Science and Engineering, 1(6): 14–21, November/December 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c6014.
pdf;   http://www.computer.
org/cse/cs1999/c6014abs.
htm.

Jafari:2011:CTS

[JSNR11] Gholamreza Jafari, Amir Hossein Shirazi, Ali Namaki, and Reza Raei. Coupled time series analysis: Methods and applications. Computing in Science and Engineering, 13(6): 84–89, November/December 2011. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Joskowicz:2001:CIG

[JT01] Leo Joskowicz and Russell H. Taylor. Computers in imaging and guided surgery. Computing in Science and Engineering, 3(5):65– 72, September/October 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c5065abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2001/pdf/c5065.pdf.

Jacquot:2006:VPD

[JWEK06] Raymond G. Jacquot, Cameron H. G. Wright, Thomas V. Edgar, and Robert F. Kubichek. Visualization of partial differential equation solutions. Computing in Science and Engineering, 8(1):73–77, January/February 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jacob:2014:DSL

[JWLG14] Ferosh Jacob, Adam Wynne, Yan Liu, and Jeff Gray. Domain-specific languages for developing and deploying signature discovery workflows. Computing in Science and Engineering, 16(1):52–64, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jiang:2019:FEA

[JXY^{+}19] Lanxin Jiang, Shoune Xiao, Bing Yang, Guangwu Yang, Tao Zhu, Dawei Dong, and Jingke Zhang. Finite element analysis of tensile properties for single-strap butt joint of carbon fiber reinforced composite. Computing in Science and Engineering, 21(3):42–50, May/June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8543178/.

Kedward:2022:SF

[KAČ^{+}22] Laurence J. Kedward, Bálint Aradi, Ondřej Čertík, Milan Curcic, Sebastian Ehlert, Philipp Engel, Rohit Goswami, Michael Hirsch, Asdrubal Lozada-Blanco, Vincent Magnin, Arjen Markus, Emanuele Pagone, Ivan Pribec, Brad Richardson, Harris Snyder, John Urban, and Jérémie Vandenplas. The state of Fortran. Computing in Science and Engineering, 24(2):63– 72, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kadanoff:2004:ECS

[Kad04]  Leo P. Kadanoff.  Excel-

lence in computer simulation. Computing in Science and Engineering, 6(2): 57–67, March/April 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/02/c2057abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2004/02/c2057.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/02/
                                        c2057.pdf.

Kalia:1999:ECT

[Kal99] Rajiv Kalia. From the editors: Consilience through information technology. Computing in Science and Engineering, 1(3):2–4, May/June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c3002.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c3002abs.
                                        htm.

Kalbe:2019:EAE

[Kal19] G. Kalbe. The European approach to the exascale challenge. Computing in Science and Engineering, 21(1):42–47, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Karczmarczuk:1999:SPS

[Kar99] Jerzy Karczmarczuk. Scientific programming: Scientific computation and functional programming. Computing in Science and Engineering, 1(3):64–72, May/ June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3064.
pdf;   http://www.computer.
org/cse/cs1999/c3064abs.
htm.

Karypis:2002:GEI

[Kar02] George Karypis. Guest Editor’s introduction: Data mining. Computing in Science and Engineering, 4(4): 12–13, July/August 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/04/c4012abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/04/c4012.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/04/
c4012.pdf.

Kaxiras:2001:GEI

[Kax01] Efthimios Kaxiras. Guest Editor’s introduction: Materials science. Computing in Science and Engineering, 3(6):14–15, November/December 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c6014abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c6014.pdf.

King:2004:DES

[KB04] Roger L. King and Ronald J.

Birk. Developing Earth system science knowledge to manage Earth’s natural resources. Computing in Science and Engineering, 6(1):45–51, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/01/c1045abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2004/01/c1045.
                                        pdf.

Krauss:2007:PMM

[KB07] Ryan W. Krauss and Wayne J. Book. A Python module for modeling and control design of flexible robots. Computing in Science and Engineering, 9(3): 41–45, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Killcoyne:2009:MCL

[KB09] Sarah Killcoyne and John Boyle. Managing chaos: Lessons learned developing software in the life sciences. Computing in Science and Engineering, 11(6): 20–29, November/December 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kurzak:2008:PHP

[KBLD08] Jakub Kurzak, Alfredo Buttari, Piotr Luszczek, and Jack Dongarra. The PlayStation 3 for high-performance scientific computing. Computing in Science and Engineering, 10(3):

84–87, May/June 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kanov:2015:JHT

[KBLE15] Kalin Kanov, Randal Burns, Cristian Lalescu, and Gregory Eyink. The Johns Hopkins Turbulence Databases: An open simulation laboratory for turbulence research. Computing in Science and Engineering, 17(5):10–17, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs2015050010-
abs.html.

Kurzhals:2015:ETC

[KBPW15] Kuno Kurzhals, Michael Burch, Thies Pfeiffer, and Daniel Weiskopf. Eye tracking in computer-based visualization. Computing in Science and Engineering, 17(5):64–71, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs2015050064-
abs.html.

Knipp:2009:PAGa

[KC09a] Peter A. Knipp and S. Raj Chaudhury. Postprocessing in automated grading systems, part 1. Computing in Science and Engineering, 11(3):

64–67, May/June 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Knipp:2009:PAGb

[KC09b] Peter A. Knipp and S. Raj Chaudhury. Postprocessing in automated grading systems, part 2. Computing in Science and Engineering, 11(4): 82–85, July/August 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Knipp:2009:PAGc

[KC09c] Peter A. Knipp and S. Raj Chaudhury. Postprocessing in automated grading systems, part 3. Computing in Science and Engineering, 11(5):64– 67, September/October 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kanewala:2019:MTS

[KC19] U. Kanewala and T. Yueh Chen. Metamorphic testing: A simple yet effective approach for testing scientific software. Computing in Science and Engineering, 21 (1):66–72, January/February 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Kamath:2002:CBD

[KCPFT02] Chandrika Kamath, Erick Cantú-Paz, Imola K. Fodor, and Nu Ai Tang. Classifying of bent-double galaxies. Computing in Science and Engineering, 4(4):52– 60, July/August 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/04/c4052abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/04/c4052.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/04/
c4052.pdf.

Karmous-Edwards:2005:GSC

[KE05] Gigi Karmous-Edwards. Global E-science collaboration. Computing in Science and Engineering, 7(2):67–74, March/ April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2005/02/c2067.
htm;   http://csdl.computer.
org/dl/mags/cs/2005/02/
c2067.pdf.

Kim:2007:VDM

[KEF07] Jinman Kim, Stefan Eberl, and Dagan Feng. Visualizing dual-modality rendered volumes using a dual-lookup table transfer function. Computing in Science and Engineering, 9(1):20–25, January/February 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kelly:2010:GCA

[Kel10]  Rory Kelly.  GPU com-

puting for atmospheric modeling. Computing in Science and Engineering, 12(4): 26–33, July/August 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Keal:2022:MMM

[KES^{+}22] Thomas W. Keal, Alin-Marin Elena, Alexey A. Sokol, Karen Stoneham, Matt I. J. Probert, Clotilde S. Cucinotta, David J. Willock, Andrew J. Logsdail, Andrea Zen, Phil J. Hasnip, Ian J. Bush, Matthew Watkins, Dario Alfè, Chris-Kriton Skylaris, Basile F. E. Curchod, Qiong Cai, and Scott M. Woodley. Materials and molecular modeling at the exascale. Computing in Science and Engineering, 24 (1):36–45, January/February 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Keyes:2005:ATQ

[Key05] Robert W. Keyes. After the transistor, the qubit? Computing in Science and Engineering, 7(1):36–41, January/February 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/01/c1036.
                                        htm;   http://csdl.computer.

org/dl/mags/cs/2005/01/
c1036.pdf.

Ketcham:2003:VBE

[KF03] Peter M. Ketcham and David L. Feder. Visualizing Bose–Einstein condensates. Computing in Science and Engineering, 5 (1):86–89, January/February 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/01/c1086.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/01/
c1086.pdf.

Kunkel:2020:HCF

[KFMG20] J. Kunkel, W. Filinger, C. Meesters, and A. Gerbes. The HPC certification forum: Toward a globally acknowledged HPC certification. Computing in Science and Engineering, 22(4):110– 114, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Keinan:2018:LCC

[KFS18] Shahar Keinan, Elizabeth Hatcher Frush, and William J. Shipman. Leveraging cloud computing for in-silico drug design using the quantum molecular design (QMD) framework. Computing in Science and Engineering, 20(4):66– 73, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/04/mcs2018040066-
                                        abs.html.

Kos:2007:MSP

[KHC^{+}07] Šimon Kos, Marina Hruška, Scott A. Crooker, Avadh Saxena, and Darryl L. Smith. Modeling spin-polarized electron transport in semiconductors for spintronics applications. Computing in Science and Engineering, 9(5): 46–52, September/October 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Katz:2020:SDC

[KHC^{+}20] D. S. Katz, N. P. C. Hong, T. Clark, M. Fenner, and M. E. Martone. Software and data citation. Computing in Science and Engineering, 22(2):4–7, March/April 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Koch:2013:VDR

[KHE13] Steffen Koch, Florian Heimerl, and Thomas Ertl. Visual document retrieval: Supporting text search and analysis with visual analytics. Computing in Science and Engineering, 15 (4):66–74, July/August 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kee:2023:SGH

[KHG^{+}23] Kerk F. Kee, Cassandra Hayes, Sandra Gesing, Annelie Rugg, Shannon Bradley, Steven R. Brandt, Natalie K. Meyers, Richard P. Johnson, and Quinn Dombrowski. Science gateways and the humanities: an exploratory study of their rare partnership. Computing in Science and Engineering, 25(1):25–33, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kelly:2009:FRP

[KHS09] Diane Kelly, Daniel Hook, and Rebecca Sanders. Five recommended practices for computational scientists who write software. Computing in Science and Engineering, 11(5): 48–53, September/October 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kilpatrick:1999:LEI

[Kil99] Alex Kilpatrick. Letters to the editors: Internet time: Switch to swatch? Computing in Science and Engineering, 1(4):5, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c4005.
pdf.

Krishnappa:2013:CCC

[KILZ13] D. K. Krishnappa, D. Irwin, E. Lyons, and M. Zink. CloudCast: Cloud computing for short-term weather forecasts. Computing in Science and Engineering, 15(4): 30–37, July/August 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kindratenko:2009:NCA

[Kin09] Volodymyr Kindratenko. Novel computing architectures. Computing in Science and Engineering, 11(3):54–57, May/ June 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kindratenko:2012:SCG

[Kin12] Volodymyr Kindratenko. Scientific computing with GPUs. Computing in Science and Engineering, 14(3):8–9, May/ June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kingsford:2016:TCB

[Kin16] Carl Kingsford. Teaching computation to biologists. Computing in Science and Engineering, 18(1):4–5, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kirby:2003:NLE

[Kir03] Robert C. Kirby. A new look at expression templates for matrix computation. Computing in Science and Engineering, 5(3): 66–70, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/03/c3066abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/03/c3066.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/03/
c3066.pdf.

Kolari:2004:WMR

[KJ04] Pranam Kolari and Anupam Joshi. Web mining: Research and practice. Computing in Science and Engineering, 6(4):49–53, July/August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/04/c4049.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/04/
c4049.pdf.

Kaplan:2020:CCC

[KKO^{+}20] M. Kaplan, C. Kneifel, V. Orlikowski, J. Dorff, M. Newton, A. Howard, D. Shinn, M. Bishawi, S. Chidyagwai, P. Balogh, and A. Randles. Cloud computing for COVID-19: Lessons learned from massively parallel models of ventilator splitting. Computing in Science and Engineering, 22(6):37–47, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kung:2007:ISS

[KL07] Rebecca Lippmann Kung and Cedric Linder. Improving students’ self-assessment of numerical analysis projects. Computing in Science and Engineering, 9(4):92–95, July/ August 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Klimeck:2010:AMR

[KL10] Gerhard Klimeck and Mathieu Luisier. Atomistic modeling of realistically extended semiconductor devices with NEMO and OMEN. Computing in Science and Engineering, 12 (2):28–35, March/April 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Karpatne:2015:GES

[KL15] Anuj Karpatne and Stefan Liess. A guide to earth science data: Summary and research challenges. Computing in Science and Engineering, 17(6):14–18, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kimpe:1999:FCT

[KLMS99] Marc Kimpe, Harry Leib, Olivier Maquelin, and Ted H.

Szymanski. Fast computational techniques for indoor radio channel estimation. Computing in Science and Engineering, 1(1):31– 41, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c1031.
pdf;   http://www.computer.
org/cse/cs1999/c1031abs.
htm.

Kothe:2019:ECU

[KLQ19] D. Kothe, S. Lee, and I. Qualters. Exascale computing in the United States. Computing in Science and Engineering, 21 (1):17–29, January/February 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Kelly:2001:CEM

[KLS01] K. Lance Kelly, Anne A.

Lazarides, and George C. Schatz. Computational electromagnetics of metal nanoparticles and their aggregates. Computing in Science and Engineering, 3(4): 67–73, July/August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c4067.
pdf;   http://www.computer.
org/cse/cs1999c4067abs.
htm.

Kapoustin:1999:CSM

[KM99] Grigori V. Kapoustin and Jian Ma. Computing simulations: Modeling adsorptiondesorption processes in porous media. Computing in Science and Engineering, 1(1):84– 91, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c1084.
                                        pdf.

Kalin:2012:CNC

[KM12] Martin Kalin and David Miller. Clojure for number crunching on multicore machines. Computing in Science and Engineering, 14(6): 12–23, November/December 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Klimeck:2008:NOA

[KMB^{+}08] Gerhard Klimeck, Michael McLennan, Sean P. Brophy, George B. Adams III, and Mark S. Lundstrom. nanoHUB.org: Advancing education and research in nanotechnology. Computing in Science and Engineering, 10(5):17–23, September/October 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Katz:2019:COC

[KMB^{+}19] D. S. Katz, L. C. McInnes, D. E. Bernholdt, A. C. Mayes, N. P. C. Hong, J. Duckles, S. Gesing, M. A. Heroux, S. Hettrick, R. C. Jimenez, M. Pierce, B. Weaver, and N. Wilkins-Diehr. Community organizations: Changing the culture in which research software is developed and sustained. Computing in Science and Engineering, 21(2): 8–24, March/April 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Krishnan:2011:RRS

[KMM^{+}11] Swaminathan Krishnan, Matthew Muto, Ramses Mourhatch, Arnar Bjorn Bjornsson, and Hemanth Siriki. Ruptureto-rafters simulations: Unifying science and engineering for earthquake hazard mitigation. Computing in Science and Engineering, 13(4): 28–43, July/August 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kendall:2010:CCS

[KMSH10] Richard P. Kendall, Andrew Mark, Susan E. Squires, and Christine A. Halverson. Condor: Case study of a largescale, physics-based code development project. Computing in Science and Engineering, 12(3):22–27, May/June 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kanai:2010:TSN

[KNG10] Yosuke Kanai, Jeffrey B. Neaton, and Jeffrey C. Grossman. Theory and simulation of nanostructured materials for photovoltaic applications. Computing in Science and Engineering, 12(2): 18–27, March/April 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Knight:2005:BSS

[Kni05] Steven Knight. Building software with SCons. Computing in Science and Engineering, 7 (1):79–88, January/February 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/01/c1079.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2005/01/
                                        c1079.pdf.

Kim:2014:FVG

[KNKP14] Seonmo Kim, Wonhong Nam, Hyunyoung Kil, and Myunghwan Park. Formal verification of a gravity-induced loss-of-consciousness monitoring system for aircraft. Computing in Science and Engineering, 16(5):96–103, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.

computer.org/csdl/mags/
cs/2014/05/mcs2014050096-
abs.html.

Katz:2018:PYS

[KNS18] Daniel S. Katz, Kyle E.

Niemeyer, and Arfon M. Smith. Publish your software: Introducing the Journal of Open Source Software (JOSS). Computing in Science and Engineering, 20(3): 84–88, May/June 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kogan:2003:TMI

[KNV03] Jacob Kogan, Charles Nicholas, and Vladimir Volkovich. Text mining with informationtheoretic clustering. Computing in Science and Engineering, 5(6):52–59, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/06/c6052abs.htm;
http://csdl.computer.org/
dl/mags/cs/2003/06/c6052.
pdf.

Kogge:2009:CPA

[Kog09] Peter M. Kogge. The challenges of petascale architectures. Computing in Science and Engineering, 11(5):10– 16, September/October 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kendall:2016:RBP

[KPA^{+}16] R. P. Kendall, D. E. Post, C. A. Atwood, K. P. Newmeyer, L. G. Votta, P. A. Gibson, D. L. Borovitcky, L. K. Miller, R. L. Meakin, M. M. Hurwitz, S. Dey, J. N. D’Angelo, R. L. Vogelsong, O. A. Goldfarb, and S. B. Allwerdt. A risk-based, practice-centered approach to project management for HPCMP CREATE. Computing in Science and Engineering, 18(1):40–51, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kuntz:1999:CSH

[KPD^{+}99] Matthew C. Kuntz, Olga Perkovic, Karin A. Dahmen, Bruce W. Roberts, and James P. Sethna. Computer simulations: Hysteresis, avalanches, and noise. Computing in Science and Engineering, 1(4):73–81, July/ August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://dlib.
computer.org/cs/books/cs1999/
pdf/c4073.pdf.

Kendall:2010:CSN

[KPM10] Richard P. Kendall, Douglass E. Post, and Andrew Mark. Case study of the Nene Code Project. Computing in Science and Engineering, 12(3):28–33, May/ June 2010. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Kraft:2003:MHD

[Kra03] L. Alan Kraft. Making home DVDs: Dell’s Movie Studio bundle. Computing in Science and Engineering, 5(2):9– 10, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/02/c2009abs.htm;
http://csdl.computer.org/
dl/mags/cs/2003/02/c2009.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/02/
c2009.pdf.

Krauth:2015:CHM

[Kra15] Werner Krauth. Coming home from a MOOC. Computing in Science and Engineering, 17(2):91–95, March/April 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/csdl/
mags/cs/2015/02/mcs2015020091.
html.

Kuck:1999:HCA

[KRH^{+}99] David J. Kuck, D. C. Rapaport, Hans J. Herrmann, H. Eugene Stanley, and Dietrich Stauffer. How computation affects science and engineering: 5 essays. Computing in Science and Engineering, 1 (1):68–80, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c1068.
pdf.

Kang:2012:ACC

[KRR^{+}12] Pilsung Kang, Naren Ramakrishnan, Calvin J. Ribbens, Srinidhi Varadarajan, and Michael Heffner. Adaptive code collage: a framework to transparently modify scientific codes. Computing in Science and Engineering, 14(1):52–63, January/February 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kleinstein:2000:CSS

[KS00] Steven H. Kleinstein and Philip E. Seiden. Computer simulations: Simulating the immune system. Computing in Science and Engineering, 2(4):69–77, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c4069.
pdf.

Kratzer:2001:SKT

[KS01] Peter Kratzer and Matthias Scheffler. Surface knowledge: Toward a predictive theory of materials. Computing in Science and Engineering, 3(6):16–25, November/December 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c6016abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c6016.pdf.

Kumar:2002:GEI

[KS02] Srikanta Kumar and Shankar Sastry. Guest Editors’ introduction: Biocomputation. Computing in Science and Engineering, 4(1):18–19, January/February 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c1018abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c1018.pdf.

Kreuzer:2006:ESI

[KS06] Edwin Kreuzer and Wolfgang Sichermann. The effect of sea irregularities on ship rolling. Computing in Science and Engineering, 8(3):26–34, May/ June 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kogge:2013:ECT

[KS13] Peter Kogge and John Shalf.

Exascale computing trends: Adjusting to the “new normal” for computer architecture. Computing in Science and Engineering, 15(6): 16–26, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kruger:2020:LRM

[KS20] F. Krüger and D. Schindler. A literature review on methods for the extraction of usage statements of software and data. Computing in Science and Engineering, 22(1):26–38, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kindratenko:2007:ASA

[KSB07] Volodymyr V. Kindratenko, Craig P. Steffen, and Robert J. Brunner. Accelerating scientific applications with reconfigurable computing: Getting started. Computing in Science and Engineering, 9(5): 70–77, September/October 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kelly:2011:SES

[KSM11] Diane Kelly, Spencer Smith, and Nicholas Meng. Software engineering for scientists. Computing in Science and Engineering, 13(5):7–11, September/October 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kim:2017:SEC

[KSM^{+}17] Sung-Eun Kim, Hua Shan, Ronald Miller, Bong Rhee, Abel Vargas, Shawn Aram, and Joseph Gorski. A scalable and extensible computational fluid dynamics software framework for ship hydrodynamics applications: NavyFOAM. Computing in Science and Engineering, 19(6):33–39, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060033-
abs.html.

Khajeh-Saeed:2012:CFD

[KSP12] Ali Khajeh-Saeed and J. Blair Perot. Computational fluid dynamics simulations using many graphics processors. Computing in Science and Engineering, 14(3):10–19, May/ June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Komati:2011:KUR

[KSSF11] Karin S. Komati, Evandro O. T. Salles, and Mario Sarcinelli-Filho. KSS: Using region and edge maps to detect image boundaries. Computing in Science and Engineering, 13(3):46–52, May/ June 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Karch:2012:DBF

[KSW^{+}12] Grzegorz K. Karch, Filip Sadlo, Daniel Weiskopf, Charles D. Hansen, Guo-Shi Li, and Thomas Ertl. Dye-based flow visualization. Computing in Science and Engineering, 14(6):80–86, November/December 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kaylor:2008:VCL

[KT08] Joe Kaylor and George K. Thiruvathukal. A virtual computing laboratory. Computing in Science and Engineering, 10(2):65–69, March/April 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kindratenko:2011:THP

[KT11] Volodymyr Kindratenko and Pedro Trancoso. Trends in high-performance computing. Computing in Science and Engineering, 13(3):92–95, May/ June 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kindratenko:2008:HPC

[KTG08] Volodymyr Kindratenko, George K. Thiruvathukal, and Steven Gottlieb. High-performance computing applications on novel architectures. Computing in Science and Engineering, 10(6):13–15, November/December 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kulakowski:2007:SRA

[Kul07] Krzysztof Kulakowski. Some recent attempts to simulate the Heider balance problem. Computing in Science and Engineering, 9(4): 80–85, July/August 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kupinski:2003:GEI

[Kup03] Matthew A. Kupinski. Guest Editor’s introduction: Computing in optics. Computing in Science and Engineering, 5(6):13–14, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/06/c6013.pdf;
http://csdl.computer.org/
comp/mags/cs/2003/06/c6013abs.
htm.

Kuske:2006:MSA

[Kus06a] Rachel Kuske. Multiple-scales approximation of a coherence resonance route to chatter. Computing in Science and Engineering, 8(3):35–43, May/ June 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kuske:2006:NSI

[Kus06b] Rachel Kuske. Noise and signal interaction. Computing in Science and Engineering, 8 (3):8–9, May/June 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2006/03/c3008.pdf.

Kusnezov:2007:HBC

[Kus07] Dimitri Kusnezov. How big can you think? challenges at the frontier. Computing in Science and Engineering, 9(5): 62–67, September/October 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kendall:2016:RBS

[KVP^{+}16] Richard P. Kendall, Lawrence G. Votta, Douglass E. Post, Chris A. Atwood, Nathan Hariharan, Scott A. Morton, Michael Gilbert, E. Thomas Moyer, Robert P. McNally, and Anthony J. Wilson. Risk-based software development practices for CRE-ATE multiphysics HPC software applications. Computing in Science and Engineering, 18(6):35–46, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2016/06/mcs2016060035-
                                        abs.html.

Kendall:2017:VVC

[KVP^{+}17] Richard P. Kendall, Lawrence G. Votta, Douglass E. Post, E. Thomas Moyer, and Scott A. Morton. Verification and validation in CRE-ATE multiphysics HPC software applications. Computing in Science and Engineering, 19(6):18–26, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060018-
abs.html.

Kwasinski:2017:SSD

[Kwa17] Andres Kwasinski. Signals, systems, and design [book review]. Computing in Science and Engineering, 19(4): 72–73, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/04/mcs2017040072.
html.

Kindratenko:2010:HPC

[KWB^{+}10] Volodymyr Kindratenko, Robert Wilhelmson, Robert Brunner, Todd J. Martinez, and Wen mei Hwu. High-performance computing with accelerators. Computing in Science and Engineering, 12(4):12– 16, July/August 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kailasa:2022:PED

[KWBB22] Srinath Kailasa, Tingyu Wang, Lorena A. Barba, and Timo Betcke. PyExaFMM: an exercise in designing highperformance software with Python and Numba. Computing in Science and Engineering, 24(5):77–84, September/October 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kaper:1999:DSS

[KWT99] Hans G. Kaper, Elizabeth Wiebel, and Sever Tipei. Data sonification and sound visualization. Computing in Science and Engineering, 1 (4):48–58, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c4048.
                                        pdf.

Kyriakidis:2008:GAE

[Kyr08] Jordan Kyriakidis. A generic approach to electronic structure calculations in nanoscopic systems. Computing in Science and Engineering, 10 (1):60–69, January/February 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lee:2018:ECN

[LA18] Christopher T. Lee and Rommie E. Amaro. Exascale computing: A new dawn for computational biology. Computing in Science and Engineering, 20(5):18–25, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/05/mcs2018050018-
                                        abs.html.

Landau:1999:EPT

[Lan99] David P. Landau. Essays:

Phase transitions and critical phenomena. Computing in Science and Engineering, 1(5): 10–11, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5010.
pdf;   http://www.computer.
org/cse/cs1999/c5010abs.
htm.

Lanzagorta:2002:IVH

[Lan02] Marco Lanzagorta. Interactive visualization of a highresolution reconstruction of the moon. Computing in Science and Engineering, 4(6): 78–82, November/December 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2002/06/c6078.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/06/
c6078.pdf.

Landau:2004:CPU

[Lan04] Rubin H. Landau. Computational physics for undergraduates: The CPUG degree program at Oregon State University. Computing in Science and Engineering, 6(2): 68–75, March/April 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/

                                        cs/2004/02/c2068abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2004/02/c2068.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/02/
                                        c2068.pdf.

Landau:2006:CPB

[Lan06] Rubin Landau. Computational physics: a better model for physics education? Computing in Science and Engineering, 8(5):22–30, September/October 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lanore:2019:FRS

[Lan19] V. Lanore. Fostering reuse in scientific computing with embedded components. Computing in Science and Engineering, 21(2):36–47, March/April 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Lathrop:2016:CAP

[Lat16]  Scott Lathrop.  A call

to action to prepare the high-performance computing workforce. Computing in Science and Engineering, 18(6):80–83, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2016/06/mcs2016060080-
                                        abs.html.

Laughlin:2002:PBC

[Lau02] Robert B. Laughlin. The physical basis of computability. Computing in Science and Engineering, 4(3):27– 30, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/03/c3027abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/03/c3027.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/03/
c3027.pdf.

Laufer:2005:HTP

[Lau05] K. Laufer. A hike through post-EJB J2EE Web Application Architecture. Computing in Science and Engineering, 7(5):80–88, September/October 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501746.pdf?
isnumber=32219&prod=JNL&
arnumber=1501746&arSt=+
80&ared=+88&arAuthor=+Laufer%
2C+K.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=32219&arnumber=
1501746&count=14&index=
11.

Laufer:2006:HTP

[Läu06] Konstantin Läufer. A hike through post-EJB J2EE Web application architecture, Part II. Computing in Science and Engineering, 8(2):79– 87, March/April 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Laufer:2008:STD

[Läu08] Konstantin Läufer. A stroll through domain-driven development with naked objects. Computing in Science and Engineering, 10(3):76–83, May/ June 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lin:2004:GWP

[LAY04] Shian-Jiann Lin, Robert Atlas, and Kao-San Yeh. Global weather prediction and highend computing at NASA. Computing in Science and Engineering, 6(1):29–35, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/01/c1029abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2004/01/c1029.
                                        pdf.

Langer:2014:PSL

[LBS14] Steven H. Langer, Abhinav Bhatele, and Charles H. Still. pF3D simulations of laser-plasma interactions in National Ignition Facility experiments. Computing in Science and Engineering, 16(6):42–50, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs2014060042-
abs.html.

Landau:2009:CSR

[LC09] Rubin Landau and Jose E.

Castillo. Computational science research and graduate studies at San Diego State University. Computing in Science and Engineering, 11(4): 5, July/August 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Luger:2012:PBB

[LC12] George F. Luger and Norman Chonacky. Programming from the bottom up [books]. Computing in Science and Engineering, 14(4): 6–10, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Li:2019:NFT

[LCC^{+}19] Kaidi Li, Chunyang Chen, Shu Cheng, Tianjian Yu, Chaoqun Xiang, and Xun Wu. A novel fault-tolerant control technique for an inverter with hysteresis current. Computing in Science and Engineering, 21(3):92–102, May/ June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8573828/.

Lathrop:2020:PCL

[LCG^{+}20] S. A. Lathrop, K. Cahill, S. I. Gordon, J. Houchins, R. M. Panoff, and A. Weeden. Preparing a computationally literate workforce. Computing in Science and Engineering, 22(4):7–16, July/August 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lohner:2004:LSF

[LCY^{+}04] Rainald Löhner, Juan Cebral, Chi Yang, Joseph D. Baum, Eric Mestreau, Charles Charman, and Daniele Pelessone. Large-scale fluidstructure interaction simulations. Computing in Science and Engineering, 6(3): 27–37, May/June 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/03/c3027abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2004/03/c3027.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/03/
                                        c3027.pdf.

Liu:2008:RTP

[LCY08] Yanling Liu, Jim X. Chen, and Lin Yang. Real-time photorealistic virtual human anatomy. Computing in Science and Engineering, 10(2): 41–47, March/April 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lampa:2019:SSW

[LDAS19] Samuel Lampa, Martin Dahlö, Jonathan Alvarsson, and Ola Spjuth. SciPipe — turning scientific workflows into computer programs. Computing in Science and Engineering, 21(3):109–113, May/ June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8701541/.

Lester:2016:CGE

[Les16] James C. Lester. Collaboration and gender equity in game-based learning for middle school computer science. Computing in Science and Engineering, 18(2): 18–28, March/April 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Leung:2017:AND

[Leu17] Mary Ann Leung. Announcing the new diversity and inclusion department. Computing in Science and Engineering, 19(6):79–81, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060079.
html.

Leung:2020:DIT

[Leu20]  M. A. Leung.  Diversity

and inclusion through leadership during challenging times. Computing in Science and Engineering, 22(6): 92–96, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

LeVeque:2009:PTR

[LeV09] Randall J. LeVeque. Python tools for reproducible research on hyperbolic problems. Computing in Science and Engineering, 11(1):19–27, January/February 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lewin:1999:TCD

[Lew99a]  David I. Lewin.  Trends:

Cruising digital ships on electronic seas. Computing in Science and Engineering, 1(5): 16–19, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5016.
pdf;   http://www.computer.
org/cse/cs1999/c5016abs.
htm.

Lewin:1999:TTR

[Lew99b] David L. Lewin. Trends: Totally rad: Planning radiation therapy in 3D. Computing in Science and Engineering, 1(2):6–9, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2006.
pdf.

Lewin:2000:FDD

[Lew00a] David I. Lewin. Focus: Do databases need protection? from whom? Computing in Science and Engineering, 2(5): 11–13, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c5011.
pdf.

Lewin:2000:FFI

[Lew00b] David I. Lewin. Focus: Freedom of information. Computing in Science and Engineering, 2(1):6–9, January/February 2000. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1006.
pdf;   http://www.computer.
org/cse/cs1999/c1006abs.
htm.

Lewin:2000:FTT

[Lew00c]  David I. Lewin.  Focus:

Teaching techies to become entrepreneurs. Computing in Science and Engineering, 2(3):6–9, May/June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3006.

pdf;   http://www.computer.
org/cse/cs1999/c3006abs.
htm.

Lewin:2001:FSE

[Lew01a]  David I. Lewin.  Focus:

Searching for the elusive qubit. Computing in Science and Engineering, 3 (4):4–7, July/August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c4004.
pdf.

Lewin:2001:KTB

[Lew01b] David I. Lewin. Keeping track of the big event. Computing in Science and Engineering, 3(5):8–11, September/October 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c5008abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c5008.pdf.

Lewin:2002:VC

[Lew02a] David Lewin. Vintage computing. Computing in Science and Engineering, 4(6): 4–10, November/December 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2002/06/c6004.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/06/
c6004.pdf.

Lewin:2002:CAP

[Lew02b] David I. Lewin. Computeraided paleontology: a new look for dinosaurs. Computing in Science and Engineering, 4(1):5–9, January/February 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2002/
c1005abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c1005.pdf.

Lewin:2002:DC

[Lew02c] David I. Lewin. DNA computing. Computing in Science and Engineering, 4(3): 5–8, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/03/c3005abs.htm;
http://csdl.computer.org/
dl/mags/cs/2002/03/c3005.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/03/
c3005.pdf.

Lewis:2010:HGP

[Lew10] Steven P. Lewis. A how-to guide for predicting properties of materials with DFT. Computing in Science and Engineering, 12(6):5–7, November/December 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Langer:2001:OIB

[LFC01] Stephen A. Langer, Edwin R. Fuller, Jr., and W. Craig Carter. OOF: An image-based finite-element analysis of material microstructures. Computing in Science and Engineering, 3(3):15–23, May/ June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3015.
pdf;   http://www.computer.
org/cse/cs1999c3015abs.
htm.

Lathrop:2019:IAS

[LFK^{+}19] S. Lathrop, M. Folk, D. S. Katz, L. Curfman McInnes, and A. Terrel. Introduction to accelerating scientific discovery with reusable software. Computing in Science and Engineering, 21(2): 5–7, March/April 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Lambert:2011:PBO

[LFN^{+}11] Emmanuel Lambert, Martin Fiers, Shavkat Nizamov, Martijn Tassaert, Steven G. Johnson, Peter Bienstman, and Wim Bogaerts. Python bindings for the Open Source Electromagnetic Simulator Meep. Computing in Science and Engineering, 13(3):53–65, May/ June 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lewin:2003:N

[LG03] David I. Lewin and Greg Goth. News. Computing in Science and Engineering, 5(1):4–13, January/February 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/01/c1004.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/01/
c1004.pdf.

Landau:2010:ACC

[LG10] Rubin Landau and Steven Gottlieb. Advancing the cause of computation in the physics curriculum. Computing in Science and Engineering, 12(2): 6–7, March/April 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ltaief:2022:RRM

[LGG^{+}22] Hatem Ltaief, Marc G.

Genton, Damien Gratadour, David E. Keyes, and Matteo Ravasi. Responsibly reckless matrix algorithms for HPC scientific applications. Computing in Science and Engineering, 24(4):12– 22, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Liang:2019:DDC

[LGJ^{+}19] Z. Liang, J. Gao, H. Jiang, X. Gao, Z. Gao, and R. Wang.

A degradation degree considered method for remaining useful life prediction based on similarity. Computing in Science and Engineering, 21 (1):50–64, January/February 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Lynch:2017:HCG

[LGW^{+}17] Larry N. Lynch, Christopher Goodin, Kevin Walker, Jody D. Priddy, and Michael Puhr. HPCMP CREATE-GV: Supporting ground vehicle acquisition. Computing in Science and Engineering, 19(6):27–32, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060027-
abs.html.

Lythe:2006:KS

[LH06] Grant Lythe and Salman Habib. Kink stochastics. Computing in Science and Engineering, 8(3):10–15, May/ June 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lu:2018:MIS

[LHGX18] Yan Lu, Zizheng Hua, Kun Gao, and Tingfa Xu. Multiperspective image stitching and regularization via hybrid structure warping. Computing in Science and Engineering, 20(2):10–23, March/April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8254328/.

Landis:2012:CMR

[LHN^{+}12] David D. Landis, Jens S.

Hummelshøj, Svetlozar Nestorov, Jeff Greeley, Marcin Dułak, Thomas Bligaard, Jens K. Nørskov, and Karsten W. Jacobsen. The computational materials repository. Computing in Science and Engineering, 14(6):51–57, November/December 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Leeser:2021:FC

[LHZ21] Miriam Leeser, Suranga Handagala, and Michael Zink. FP-GAs in the cloud. Computing in Science and Engineering, 23(6):72–76, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Liu:2011:HER

[Liu11] Ming Liu. A high-end reconfigurable computation platform for nuclear and particle physics experiments. Computing in Science and Engineering, 13(2):52–63, March/April 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Liu:2015:SMT

[Liu15] Yan Liu. Scalable multivariate time-series models for climate informatics. Computing in Science and Engineering, 17(6):19–26, November/December 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Liu:2006:WWA

[LJWC06] Shengjun Liu, Xiaogang Jin, Charlie C. L. Wang, and Jim X. Chen. Water wave animation on mesh surfaces. Computing in Science and Engineering, 8(5):81–87, September/October 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lynett:2011:NSC

[LL11] Patrick Lynett and Philip Lui. Numerical simulation of complex tsunami behavior. Computing in Science and Engineering, 13(4):50– 57, July/August 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lehner:2013:SUE

[LL13] Luis Lehner and Steven L. Liebling. Simulations to usher in the era of gravitational wave astronomy. Computing in Science and Engineering, 15(2): 60–65, March/April 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Li:2018:BDK

[LL18] Xuewei Li and Xueyan Li. Big data and its key technology in the future. Computing in Science and Engineering, 20(4): 75–88, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040075-
abs.html.

Liu:2019:SCL

[LL19] H. Liu and X. Li. Sharp curve lane detection for autonomous driving. Computing in Science and Engineering, 21(2): 80–95, March/April 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Liang:2018:IAH

[LLQ18] Decui Liang, Dun Liu, and Wei Quan. Information aggregation of hesitant fuzzy interval sets for multicriteria decision-making. Computing in Science and Engineering, 20(2):38–51, March/April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8254317/.

Liu:2007:RLD

[LM07a] Zhanping Liu and Robert J.

Moorhead, II. Robust loop detection for interactively placing evenly spaced streamlines. Computing in Science and Engineering, 9(4): 86–91, July/August 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lubin:2007:RSE

[LM07b] Dan Lubin and Robert Massom. Remote sensing of Earth’s polar regions: Opportunities for computational science. Computing in Science and Engineering, 9(1):58–71, January/February 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lathrop:2008:HPC

[LM08] Scott Lathrop and Thomas Murphy. High-performance computing education. Computing in Science and Engineering, 10(5):9–11, September/October 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Linares:2020:TST

[LMC20] M. P. Linares, L. Montero, and J. Casanovas. A traffic simulation tool for assessing smart city policies (CitScale). Computing in Science and Engineering, 22(3): 100–112, May/June 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lapenta:2013:SWP

[LMPV13] Giovanni Lapenta, Stefano Markidis, Stefaan Poedts, and Dean Vucinic. Space weather prediction and exascale computing. Computing in Science and Engineering, 15(5):68– 76, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Li:2019:SRM

[LNI^{+}19] Y. Li, K. Nomura, J. A. Insley, V. Morozov, K. Kumaran, N. A. Romero, W. A. Goddard, R. K. Kalia, A. Nakano, and P. Vashishta. Scalable reactive molecular dynamics simulations for computational synthesis. Computing in Science and Engineering, 21(5): 64–75, September/October 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Lo:1999:SCD

[Lo99] Martin W. Lo. Satelliteconstellation design. Computing in Science and Engineering, 1(1):58–67, January/February 1999. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c1058.
pdf;   http://www.computer.
org/cse/cs1999/c1058abs.
htm.

Loftin:2003:MPB

[Lof03] R. Bowen Loftin. Multisensory perception: Beyond the visual in visualization. Computing in Science and Engineering, 5(4):56–58, July/August 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2003/04/c4056.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/04/
                                        c4056.pdf.

Lofstead:2022:EII

[Lof22] Jay Lofstead. Experiences integrating interns into research software teams. Computing in Science and Engineering, 24(3):33–41, May/June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lossen:2003:SCA

[Los03] Christoph Lossen. Singular: a computer algebra system. Computing in Science and Engineering, 5(4): 45–55, July/August 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/04/c4045abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/04/c4045.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/04/
                                        c4045.pdf.

Love:2004:BRL

[Lov04] Peter Love. Book reviews:

The legacy of Alan Turing: Alan Turing: Life and Legacy of a Great Thinker, edited by Christof Teuscher. Computing in Science and Engineering, 6(4):97–99, July/ August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/04/c4097.pdf;
http://csdl.computer.org/
dl/mags/cs/2004/04/c4097.
htm.

Landau:2013:BMA

[LPB13] Rubin H. Landau, Manuel J.

Paez, and Cristian C. Bordeianu. A blended, multimodal access eTextBook in computational physics. Computing in Science and Engineering, 15(3):96–103, May/ June 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Landau:2015:VRC

[LPB15] Rubin H. Landau, Manuel J.

Paez, and Cristian C. Bordeianu. Visions and realizations of a computational eTextbook. Computing in Science and Engineering, 17(5):72–79, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/

                                        cs/2015/05/mcs2015050072-
                                        abs.html.

Li:2019:IIC

[LPCY19] Q. Li, Q. Peng, J. Chen, and C. Yan. Improving image classification accuracy with ELM and CSIFT. Computing in Science and Engineering, 21(5):26–34, September/October 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Lipshitz:2000:SPU

[LPV00] Stanley P. Lipshitz, Renato Portugal, and John Vanderkooy. Scientific programming: Using computer algebra to explore sound-wave propagation in spherical cavities. Computing in Science and Engineering, 2(2): 84–94, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c2084.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c2084abs.
                                        htm.

Li:2018:MVE

[LPY18] Qing Li, Qiang Peng, and Chuan Yan. Multiple VLAD encoding of CNNs for image classification. Computing in Science and Engineering, 20(2):52–63, March/April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8254309/.

Li:2019:ADR

[LQZL19] Tian Li, Deng Qin, Jiye Zhang, and Ming Li. Aerodynamic drag reduction of high-speed train nose with bionic round pits. Computing in Science and Engineering, 21(3):31–41, May/June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8656573/.

Lanzagorta:2000:VCR

[LRRK00] Marco Lanzagorta, Robert Rosenberg, Lawrence J. Rosenblum, and Eddy Y. Kuo. Visualization corner: Rapid prototyping of virtual environments. Computing in Science and Engineering, 2 (3):68–73, May/June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3068.
pdf;   http://www.computer.
org/cse/cs1999/c3068abs.
htm.

Leroy:2022:RCL

[LSBC22] Dorian Leroy, June Sallou, Johann Bourcier, and Benoit Combemale. On the role of computer languages in scientific computing. Computing in Science and Engineering, 24 (4):55–59, July/August 2022.

CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Leal:2004:MCS

[LSDP^{+}04] Antonio Leal, Franciso Sánchez-Doblado, María Perucha, Ester Carrasco, Magnolia Rincón, Rafael Arrans, and Carlos Bernal. Monte Carlo simulation of complex radiotherapy treatments. Computing in Science and Engineering, 6(4):60–68, July/August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2004/04/c4060.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/04/
                                        c4060.pdf.

Lin:2020:EMT

[LSN20] X. Lin, M. Simon, and N. Niu. Exploratory metamorphic testing for scientific software. Computing in Science and Engineering, 22(2): 78–87, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lin:2021:DMR

[LSPN21] X. Lin, M. Simon, Z. Peng, and N. Niu. Discovering metamorphic relations for scientific software from user forums. Computing in Science and Engineering, 23(2): 65–72, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Luu:2007:PEC

[LSV^{+}07] Thomas Luu, Ron Soltz, Pavlos Vranas, Christopher M. Frenz, and Francis Sullivan. Petaflop essay contest. Computing in Science and Engineering, 9(6):55–59, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Li:2008:CMB

[LT08] Nolan Li and Ani R. Thakar.

CasJobs and MyDB: a batch query workbench. Computing in Science and Engineering, 10 (1):18–29, January/February 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Laufer:2009:PTP

[LT09] Konstantin Laufer and George K.

Thiruvathukal. The promises of typed, pure, and lazy functional programming: Part II. Computing in Science and Engineering, 11(5):68– 75, September/October 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Laufer:2011:MAD

[LTD11] Konstantin Laufer, George K.

Thiruvathukal, and David Dennis. Moving academic department functions to social networks and clouds: Initial experiences. Computing in Science and Engineering, 13(5):84–89, September/October 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Laufer:2007:HTP

[LTG07] Konstantin Läufer, George K. Thiruvathukal, and Benjamín González. A hike through a post-EJB J2EE Web application architecture, Part III. Computing in Science and Engineering, 9(1):82–95, January/February 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Laufer:2009:PSW

[LTNME09] Konstantin Laufer, George K. Thiruvathukal, Ryohei Nishimura, and Carlos Ramirez Martinez-Eiroa. Putting a slug to work. Computing in Science and Engineering, 11(2):62– 68, March/April 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ludlam:2013:BEN

[Lud13] John Ludlam. R by Example: Numerical Solutions to Differential Equations (Soetaert, K., et al; 2012) [book review]. Computing in Science and Engineering, 15(2): 6–7, March/April 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Luijten:2006:FSG

[Lui06] Erik Luijten. Fluid simulation with the geometric cluster Monte Carlo algorithm. Computing in Science and Engineering, 8(2): 20–29, March/April 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lummerzheim:2007:MFA

[Lum07] Dirk Lummerzheim. Modeling and forecasting aurora. Computing in Science and Engineering, 9(5):53–61, September/October 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lee:2014:ELC

[LUMM14] Myoungkyu Lee, Rhys Ulerich, Nicholas Malaya, and Robert D. Moser. Experiences from leadership computing in simulations of turbulent fluid flows. Computing in Science and Engineering, 16(5):24– 31, September/October 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:/

/csdl.computer.org/csdl/
mags/cs/2014/05/mcs2014050024-
abs.html.

Lunney:2001:CDD

[Lun01] Elizabeth A. Lunney. Computing in drug discovery:

The design phase. Computing in Science and Engineering, 3(5):105–108, September/October 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c5105abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c5105.pdf.

Luo:2012:DBF

[Luo12] Yanlin Luo. Distance-based focus + context models for exploring large volumetric medical datasets. Computing in Science and Engineering, 14(5):63–71, September/October 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Luo:2013:EVS

[Luo13] Yan-Lin Luo. Effectively visualizing the spatial structure of cerebral blood vessels. Computing in Science and Engineering, 15(2): 41–46, March/April 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Leukkunen:2014:MCF

[LVLA14] Lauri Leukkunen, Tuukka Verho, and Olga Lopez-Acevedo. A multiscale code for flexible hybrid simulations using ASE framework. Computing in Science and Engineering, 16(2):54–62, March/April 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Landau:2002:FSD

[LVWK02] Rubin H. Landau, David Vediner, Pornrat Wattanakasiwich, and Kevin R. Kyle. Future scientific digital documents with MathML, XML, and SVG. Computing in Science and Engineering, 4 (2):77–85, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c2077abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c2077.pdf.

Liu:2010:GMS

[LWF10] Siyuan Liu, Gaojin Wen, and Jianping Fan. A 3D geosciences modeling system for large-scale water-diversion projects. Computing in Science and Engineering, 12 (1):28–35, January/February 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Li:2019:NCF

[LWG19] M. Li, S. Wibowo, and W. Guo. Nonlinear curve fitting using extreme learning machines and radial basis function networks. Computing in Science and Engineering, 21 (5):6–15, September/October 2019. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366x (electronic).

Lin:2007:SCF

[LWSK07] Guang Lin, Xiaoliang Wan, Chau-Hsing Su, and George Em Karniadakis. Stochastic computational fluid mechanics. Computing in Science and Engineering, 9(2):21– 29, March/April 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Li:2013:CST

[LWT^{+}13] Jihui Li, Felasfa Wodajo, Mark Theiss, Michelle Kew, and Alison Jarmas. Computer simulation techniques in giant cell tumor curettage and defect reconstruction. Computing in Science and Engineering, 15(2):21–26, March/April 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lee:2007:SRA

[LYC07] Tong-Yee Lee, Chung-Ren Yan, and Ming-Te Chi. Stylized rendering for anatomic visualization. Computing in Science and Engineering, 9 (1):13–19, January/February 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Li:2017:NST

[LZZ17] Shuang Li, Yuezhi Zhou, and Yaoxue Zhang. NSAP+:

Supporting transparent computing applications with a service-oriented protocol. Computing in Science and Engineering, 19(1):21–28, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs2017010021-
abs.html.

Ma:2003:VTV

[Ma03]  Kwan-Liu Ma.  Visual-

izing time-varying volume data. Computing in Science and Engineering, 5(2): 34–42, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/02/c2034abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/02/c2034.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/02/
c2034.pdf.

Millman:2011:PSE

[MA11] K. Jarrod Millman and Michael Aivazis. Python for scientists and engineers. Computing in Science and Engineering, 13(2):9–12, March/ April 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ma:2016:VTS

[Ma16] Kwan-Liu Ma. Visualization techniques for studying largescale flow fields from fusion simulations. Computing in Science and Engineering, 18 (2):68–77, March/April 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Macedo:2008:AGM

[MAC08] Raquel S. Macedo, Marcelo F. Alfradique, and Marcelo Castier. Automatic generation of Matlab functions using Mathematica and Thermath. Computing in Science and Engineering, 10(4): 41–49, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mandanici:2021:SPG

[MAFM21] A. Mandanici, S. Alessandro Sarà, G. Fiumara, and G. Mandaglio. Studying physics, getting to know Python: RC circuit, simple experiments, coding, and data analysis with Raspberry Pi. Computing in Science and Engineering, 23(1):93–96, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mattson:2021:PPM

[MAG21] Timothy G. Mattson, Todd A. Anderson, and Giorgis Georgakoudis. PyOMP: Multithreaded parallel programming in Python. Computing in Science and Engineering, 23(6):77–80, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Major:2003:BTD

[Maj03]  François Major.  Building

three-dimensional ribonucleic acid structures. Computing in Science and Engineering, 5(5):44–53, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/05/c5044abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/05/c5044.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/05/
c5044.pdf.

Makino:2006:GP

[Mak06]  Junichiro Makino.  The

GRAPE project. Computing in Science and Engineering, 8 (1):30–40, January/February 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Malak:2000:WMA

[Mal00] Michael Malak. Web mechanics: Adding video to your Web pages. Computing in Science and Engineering, 2(3):74–77, May/June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3074.

                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c3074abs.
                                        htm.

Malek:2007:TPCa

[Mal07a] Kourosh Malek. Transport in protein crystals, Part I: Insights from molecular simulations. Computing in Science and Engineering, 9(5): 90–95, September/October 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Malek:2007:TPCb

[Mal07b] Kourosh Malek. Transport in protein crystals, Part II: Diffusion simulation and chiral recognition. Computing in Science and Engineering, 9(6):70–75, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Marchioro:1999:ICM

[Mar99a] Tom Marchioro. Interview: With Cleve Moler putting math to work. Computing in Science and Engineering, 1(4):10–13, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c4010.
                                        pdf.

Marder:1999:MDC

[Mar99b] Michael Marder. Molecular dynamic of cracks. Computing in Science and Engineering, 1(5):48–55, September/October 1999. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5048.
pdf;   http://www.computer.
org/cse/cs1999/c5048abs.
htm.

Marone:2002:MOI

[Mar02] Matt Marone. The Mercer online interactive chaotic pendulum. Computing in Science and Engineering, 4(4):94, C3, July/August 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/04/c4094abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/04/c4094.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/04/
c4094.pdf.

Martin:2017:UCP

[Mar17]  Richard F. Martin.  Un-

dergraduate computational physics education: Uneven history and promising future. Computing in Science and Engineering, 19(2):70– 78, March/April 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/02/mcs2017020070-
abs.html.

Mason:2006:DLS

[Mas06] Bruce Mason. Digital libraries in support of science education: a case for computational physics. Computing in Science and Engineering, 8(4): 62–65, July/August 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Matsuura:2005:QPG

[Mat05] M. Matsu’ura. Quest for predictability of geodynamic processes through computer simulation. Computing in Science and Engineering, 7 (4):43–50, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        5992/31456/01463135.pdf?
                                        isnumber=31456&prod=JNL&
                                        arnumber=1463135&arSt=+
                                        43&ared=+50&arAuthor=Matsu%
                                        27ura%2C+M.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=31456&arnumber=
                                        1463135&count=14&index=5.

Mousseau:1999:CSE

[MB99] Normand Mousseau and Gerard T. Barkema. Computer simulations: Exploring high-dimensional energy landscapes. Computing in Science and Engineering, 1 (2):74–80, 82, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/

books/cs1999/pdf/c2074.
pdf.

Millman:2007:AFM

[MB07] K. Jarrod Millman and Matthew Brett. Analysis of functional magnetic resonance imaging in Python. Computing in Science and Engineering, 9(3):52–55, May/ June 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Memarsadeghi:2011:NCC

[MB11] Nargess Memarsadeghi and Brent J. Bos. NASA computational case study characterizing moving particles. Computing in Science and Engineering, 13(6):76–78, November/December 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mesnard:2017:RRC

[MB17] Olivier Mesnard and Lorena A.

Barba. Reproducible and replicable computational fluid dynamics: It’s harder than you think. Computing in Science and Engineering, 19(4): 44–55, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/04/mcs2017040044-
abs.html.

Marques:2020:TEE

[MB20a] O. Marques and A. Barker.

Training efforts in the Exascale Computing Project. Computing in Science and Engineering, 22(5):103–107, September/October 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mesnard:2020:RWP

[MB20b] O. Mesnard and L. A. Barba. Reproducible workflow on a public cloud for computational fluid dynamics. Computing in Science and Engineering, 22(1):102–116, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mullen:2009:HPS

[MBB^{+}09] Julie Mullen, Nadya Bliss, Robert Bond, Jeremy Kepner, Hahn Kim, and Albert Reuther. High-productivity software development with pMatlab. Computing in Science and Engineering, 11 (1):75–79, January/February 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mundt:2022:PGC

[MBB^{+}22] Miranda R. Mundt, Keith Beattie, Jonathan Bisila, Charles R. Ferenbaugh, William F. Godoy, Rinku Gupta, Jonathan E. Guyer, Miriam Kiran, Addi Malviya-Thakur, Reed Milewicz, Benjamin H. Sims, Vanessa Sochat, and Joshua B. Teves. For the public good: Connecting, retaining, and recognizing current and future RSEs at U.S. National Research Laboratories and Agencies. Computing in Science and Engineering, 24(6):6–13, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mesh:2014:LES

[MBH14] Erika S. Mesh, Gabbie Burns, and J. Scott Hawker. Leveraging expertise to support scientific software process improvement decisions. Computing in Science and Engineering, 16(3):28–34, May/ June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Makino:2000:LEF

[MBS^{+}00] Jun Makino, John Board, Klaus Schulten, Peter Borcherds, and Rubin D. Orduz Z. Letters to the editors: “The Fast Multipole Algorithm” and “The Top 10 Algorithms”. Computing in Science and Engineering, 2(3):4–5, May/ June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3004.

pdf. See [DS00, BS00d].

Martin:2005:AMR

[MCAA05] Daniel F. Martin, Phillip Colella, Marian Anghel, and Francis J. Alexander. Adaptive mesh refinement for multiscale nonequilibrium physics. Computing in Science and Engineering, 7(3):24–31, May/ June 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2005/03/c3024abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2005/03/
c3024.pdf.

McClean:2021:MQC

[McC21] Jarrod R. McClean. From molecules to quantum computers: a research retrospective. Computing in Science and Engineering, 23(6): 52–57, November/December 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Myers:2003:RIR

[MCE^{+}03] James D. Myers, Alan Chappell, Matthew Elder, Al Geist, and Jens Schwidder. Reintegrating the research record. Computing in Science and Engineering, 5(3):44–50, May/ June 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2003/03/c3044abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/03/
c3044.htm;   http://csdl.
computer.org/dl/mags/cs/
2003/03/c3044.pdf.

Milewicz:2022:SFO

[MCGA22] Reed Milewicz, Jeffrey Carver, Samuel Grayson, and Travis Atkison. A secure future for open-source computational science and engineering. Computing in Science and Engineering, 24(4):65– 69, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

McKay:2000:BWR

[McK00]  Susan R. McKay.  Book

and Web reviews: Mathematical models. Computing in Science and Engineering, 2(4):96, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c4096.
pdf.

McKenna:2011:OFE

[McK11] Frank McKenna. OpenSees: a framework for earthquake engineering simulation. Computing in Science and Engineering, 13(4):58–66, July/August 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

McMail:2009:NGR

[McM09] Thomas C. McMail. Nextgeneration research and breakthrough innovation. Computing in Science and Engineering, 11(6):76–84, November/December 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mehta:2016:CKC

[MDK16] Dinesh P. Mehta, Anthony M. Dean, and Tina M. Kouri. Chemical kinetics: A CS perspective. Computing in Science and Engineering, 18(5): 48–55, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Matsuoka:2022:PFR

[MDW^{+}22] Satoshi Matsuoka, Jens Domke, Mohamed Wahib, Aleksandr Drozd, Andrew A. Chien, Raymond Bair, Jeffrey S. Vetter, and John Shalf. Preparing for the future rethinking proxy applications. Computing in Science and Engineering, 24(2):85–90, March/April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Meindl:2003:BML

[Mei03]  James D. Meindl.  Be-

yond Moore’s Law: The interconnect era. Computing in Science and Engineering, 5(1):20–24, January/February 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/01/c1020abs.htm;
 http://csdl.computer.org/

dl/mags/cs/2003/01/c1020.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/01/
c1020.pdf.

Meisel:2010:BCC

[Mei10] David D. Meisel. Bringing computation to the classroom. Computing in Science and Engineering, 12(3):6–11, May/ June 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Memarsadeghi:2015:CSG

[Mem15] Nargess Memarsadeghi. Citizen science [Guest Editors’ introduction]. Computing in Science and Engineering, 17(4):8–10, July/ August 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040008.
html.

Memarsadeghi:2016:NCC

[Mem16] Nargess Memarsadeghi. NASA computational case study: Golomb rulers and their applications. Computing in Science and Engineering, 18(6):58–62, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs2016060058-
abs.html.

Meneses:2018:ADI

[Men18] Esteban Meneses. Accelerating discovery and innovation through advanced computing: Perspective of a highperformance computing scientist in Costa Rica. Computing in Science and Engineering, 20(3):36–38, May/ June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8357991/.

Mertens:2002:CCP

[Mer02] Stephan Mertens. Computational complexity for physicists. Computing in Science and Engineering, 4(3): 31–47, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/03/c3031abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/03/c3031.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/03/
                                        c3031.pdf.

Messina:2015:GBE

[Mes15] Paul Messina. Gaining the broad expertise needed for high-end computational science and engineering research. Computing in Science and Engineering, 17(2): 89–90, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.

computer.org/csdl/mags/
cs/2015/02/mcs2015020089-
abs.html.

Messina:2017:ECP

[Mes17] Paul Messina. The Exascale Computing Project. Computing in Science and Engineering, 19(3):63–67, May/ June 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://www.computer.org/
csdl/mags/cs/2017/03/mcs2017030063.
html.

Madduri:2016:SSG

[MF16] Ravi Madduri and Ian Foster. Science as a service [Guest Editors’ introduction]. Computing in Science and Engineering, 18(5):8–9, September/October 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Muller:2009:JVL

[MFD^{+}09] Daniel Muller, Bernhard Fleck, George Dimitoglou, Benjamin W. Caplins, Desmond E. Amadigwe, Juan Pablo Garcia Ortiz, Benjamin Wamsler, Alen Alexanderian, V. Keith Hughitt, and Jack Ireland. JHelioviewer: Visualizing large sets of solar images using JPEG 2000. Computing in Science and Engineering, 11(5):38–47, September/October 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Marin:2017:SSE

[MGCBI17] Mauricio Marin, Veronica Gil-Costa, Carolina Bonacic, and Alonso Inostrosa. Simulating search engines. Computing in Science and Engineering, 19(1):62–73, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/01/mcs2017010062-
                                        abs.html.

Miles:2008:PBB

[MGD^{+}08] Simon Miles, Paul Groth, Ewa Deelman, Karan Vahi, Gaurang Mehta, and Luc Moreau. Provenance: The bridge between experiments and data. Computing in Science and Engineering, 10(3):38–46, May/ June 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mesa:2012:ORF

[MGFRL^{+}12] Luis Eduardo Mesa, Ignacio Gonzalez-Franco, Eugenio Roanes-Lozano, Alberto Garcia-Alvarez, and Antonio Hernando. Optimal route finding and rolling-stock selection for the Spanish railways. Computing in Science and Engineering, 14(4): 82–89, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Myers:2007:PUS

[MGS07] Christopher R. Myers, Ryan N.

Gutenkunst, and James P. Sethna. Python unleashed on systems biology. Computing in Science and Engineering, 9(3): 34–37, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Moresi:2000:PTC

[MGZ00] Louis Moresi, Michael Gurnis, and Shijie Zhong. Plate tectonics and convection in the Earth’s mantle: Toward a numerical simulation. Computing in Science and Engineering, 2(3):22–33, May/ June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3022.
pdf;   http://www.computer.
org/cse/cs1999/c3022abs.
htm.

Mocskos:2018:BAC

[MHC^{+}18] Esteban Mocskos, Carlos J. Barrios H., Harold Castro, Dennis Cazar Ramírez, Sergio Nesmachnow, and Rafael Mayo-García. Boosting advanced computational applications and resources in Latin America through collaboration and sharing. Computing in Science and Engineering, 20(3):39–48, May/June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8358001/.

Muir:1999:LEI

[MHDM99] Douglas W. Muir, William G. Harter, Denis Donnelly, and Thomas L. Marchioro II. Letters to the editors: Internet time redux; educational software contest. Computing in Science and Engineering, 1 (5):4–9, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c5004.
                                        pdf.

Merks:2006:PSE

[MHK^{+}06] Roeland M. H. Merks, Alfons G. Hoekstra, Jaap A. Kaandorp, Peter M. A. Sloot, and Paulien Hogeweg. Problem-solving environments for biological morphogenesis. Computing in Science and Engineering, 8(1):61–72, January/February 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Miller:2010:SBE

[Mil10] Loren K. Miller. Simulationbased engineering for industrial competitive advantage. Computing in Science and Engineering, 12(3):14–21, May/ June 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Miller:2017:PIT

[Mil17] Loren Miller. Product innovation through computational prototypes and supercomputing. Computing in Science and Engineering, 19(6):9–17, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060009-
abs.html.

Miller:2021:IBL

[Mil21] Mark C. Miller. Inclusivity bugs and the language we use. Computing in Science and Engineering, 23(6): 69–71, November/December 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mishra:2002:CG

[Mis02]  Bud Mishra.  Compar-

ing genomes. Computing in Science and Engineering, 4(1):42–49, January/February 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c1042abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c1042.pdf.

Madadi:2009:ISE

[MJAK09] Mahyar Madadi, Anthony C.

Jones, Christoph H. Arns, and Mark A. Knackstedt. 3D imaging and simulation of elastic properties of porous materials. Computing in Science and Engineering, 11(4): 65–73, July/August 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Moorhead:2006:VRC

[MJM^{+}06] Robert Moorhead, Chris Johnson, Tamara Munzner, Hanspeter Pfister, Penny Rheingans, and Terry S. Yoo. Visualization research challenges: a report summary. Computing in Science and Engineering, 8(4): 66–73, July/August 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

McLennan:2010:HPD

[MK10] Michael McLennan and Rick Kennell. HUBzero: a platform for dissemination and collaboration in computational science and engineering. Computing in Science and Engineering, 12(2):48– 53, March/April 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

McLaughlin:2023:PSC

[MKH^{+}23] Maura McLaughlin, Nicholas Kisseberth, Sue Ann Heatherly, Harsha Blumer, Timothy Olszanski, Claire Stirm, Jack A. Smith, and Duncan Lorimer. The Pulsar Science Collaboratory: Exploring pulsars through a science gateway. Computing in Science and Engineering, 25(1):17–24, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Maslowski:2007:TPE

[MKJ07] Wieslaw Maslowski, Jaclyn Clement Kinney, and Jaromir Jakacki. Toward prediction of environmental Arctic change. Computing in Science and Engineering, 9(6):29–34, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mirtaheri:2014:MPB

[MKM^{+}14] Seyedeh Leili Mirtaheri, Ehsan Mousavi Khaneghah, Amir Saman Memaripour, Lucio Grandinetti, Mohsen Sharifi, and Zarrintaj Bornaee. Multics and Plan 9: The big bangs in the distributed computing system universe. Computing in Science and Engineering, 16(5):76–85, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs2014050076-
abs.html.

Manzardo:2002:ISP

[ML02] Mark A. Manzardo and Kenneth G. LeSueur. An infrared-scene projector digital model. Computing in Science and Engineering, 4 (2):58–65, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c2058abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2002/pdf/c2058.pdf.

Madden:2004:SUC

[MM04] Timothy J. Madden and James H. Miller. Simulation of unsteadiness in chemical oxygen-iodine laser flowfields. Computing in Science and Engineering, 6(6): 47–54, November/December 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2004/06/c6047.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/06/
                                        c6047.pdf.

Magana:2012:MAP

[MM12] Alejandra J. Magana and Jyoti I. Mathur. Motivation, awareness, and perceptions of computational science. Computing in Science and Engineering, 14(1):74–79, January/February 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Megler:2013:DNH

[MM13] V. M. Megler and David Maier. Data near here: Bringing relevant data closer to scientists. Computing in Science and Engineering, 15(3): 44–53, May/June 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Memarsadeghi:2014:NCC

[MM14] Nargess Memarsadeghi and Lucy McFadden. NASA computational case study: Where is my Moon? Computing in Science and Engineering, 16(6):92–99, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs2014060092-
abs.html.

Morton:2016:FWA

[MM16] S. A. Morton and D. R. Mc-Daniel. A fixed-wing aircraft simulation tool for improving DoD acquisition efficiency. Computing in Science and Engineering, 18(1):25–31, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Maciol:2018:AMG

[MM18] Piotr Maciol and Kazimierz Michalik. Application of metaprogramming and generic programming in multiscale modeling. Computing in Science and Engineering, 20(6):

81–94, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://www.computer.org/
csdl/mags/cs/2018/06/08500322-
abs.html.

Mora:2005:AAC

[MMG^{+}05] P. Mora, H. Muhlhaus, L. Gross, Huilin Xing, D. Weatherly, S. Abe, S. Latham, and L. Moresi. ACcESS: Australia’s contribution to the iSERVO Institute’s development. Computing in Science and Engineering, 7 (4):27–37, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463133.pdf?
isnumber=31456&prod=JNL&
arnumber=1463133&arSt=+
27&ared=+37&arAuthor=Mora%
2C+P.%3B+Muhlhaus%2C+H.%
3B+Gross%2C+L.%3B+Huilin+
Xing%3B+Weatherly%2C+D.%
3B+Abe%2C+S.%3B+Latham%
2C+S.%3B+Moresi%2C+L.;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31456&arnumber=1463133&
count=14&index=3.

Messmer:2008:GGC

[MMG08] Peter Messmer, Paul J. Mullowney, and Brian E. Granger. GPULib: GPU computing in high-level languages. Computing in Science and Engineering, 10(5):70–73, September/October 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mandanici:2022:SPP

[MMP^{+}22] Andrea Mandanici, Giuseppe Mandaglio, Giovanni Pirrotta, Valeria Conti Nibali, and Giacomo Fiumara. Simple physics with Python: a workbook on introductory physics with open-source software. Computing in Science and Engineering, 24(2):74– 78, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mundt:2022:TSP

[MMR22] Miranda R. Mundt, Reed M.

Milewicz, and Elaine M. Raybourn. In their shoes: Persona-based approaches to software quality practice incentivization. Computing in Science and Engineering, 24 (3):42–50, May/June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

McGee:2017:DTC

[MMTD^{+}17] Steven McGee, Randi McGee-Tekula, Jennifer Duck, Ronald I. Greenberg, Lucia Dettori, Dale F. Reed, Brenda Wilkerson, Don Yanek, Andrew M. Rasmussen, and Gail Chapman. Does a taste of computing increase computer science enrollment? Computing in Science and Engineering, 19(3):8–18, May/June 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2017/03/mcs2017030008-
abs.html.

Memarsadeghi:2003:CID

[MO03] Nargess Memarsadeghi and Dianne P. O’Leary. Classified information: The data clustering problem. Computing in Science and Engineering, 5(5):54–60, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/05/c5054abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/05/c5054.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/05/
c5054.pdf.

Martinez-Ortiz:2022:SUT

[MOBD^{+}22] Carlos Martinez-Ortiz, Rena Bakhshi, Yifat Dzigan, Nicolas Renaud, Faruk Diblen, Berend Weel, Maarten van Meersbergen, Niels Drost, Sven van der Burg, and Fakhereh Alidoost. Structured and unstructured teams for research software development at The Netherlands eScience Center. Computing in Science and Engineering, 24(3): 25–32, May/June 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Moldenhauer:2012:FIS

[Mol12] Jacob Moldenhauer. Fortranning it with style [review of Modern Fortran: Style and Usage (Clerman, N. S. and Spector, W.; 2012)]. Computing in Science and Engineering, 14(5):5–6, September/October 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mooers:2021:PSL

[Moo21] B. H. M. Mooers. A PyMOL snippet library for Jupyter to boost researcher productivity. Computing in Science and Engineering, 23(2): 47–53, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Morningstar:2015:UEH

[Mor15]  Colin Morningstar.  Un-

earthing excited hadron resonances in lattice QCD using NSF XSEDE resources. Computing in Science and Engineering, 17(2):23–29, March/ April 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/csdl/
mags/cs/2015/02/mcs2015020023-
abs.html.

Moyer:2006:NSS

[Moy06] Curt A. Moyer. Numerical solution of the stationary state Schrödinger equation using transparent boundary conditions. Computing in Science and Engineering, 8(4): 32–40, July/August 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ma:2009:CCI

[MP09] Jianwei Ma and Gerlind Plonka. Computing with curvelets: From image processing to turbulent flows. Computing in Science and Engineering, 11(2):72–80, March/ April 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Malensek:2014:EGG

[MPP14] Matthew Malensek, Sangmi Pallickara, and Shrideep Pallickara. Evaluating geospatial geometry and proximity queries using distributed hash tables. Computing in Science and Engineering, 16(4):53–61, July/August 2014. CODEN CSENFA. ISSN 1521-9615.

Markt:2018:ACS

[MPR18] David P. Markt, Ashish Pathak, and Mehdi Raessi. Advanced computational simulations of surface impingement of a train of ethanol drops: A pathway to developing spray-wall interaction submodels. Computing in Science and Engineering, 20(4): 56–65, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.

computer.org/csdl/mags/
cs/2018/04/mcs2018040056-
abs.html.

Margulieux:2023:LBL

[MPRU23] Lauren E. Margulieux, James Prather, Masoumeh Rahimi, and Gozde Cetin Uzun. Leverage biology to learn rapidly from mistakes without feeling like a failure. Computing in Science and Engineering, 25 (2):44–49, March/April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

McCormick:2006:GEI

[MR06] Steve McCormick and Ulrich Rüde. Guest Editors’ introduction: Multigrid computing. Computing in Science and Engineering, 8(6):10–11, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2006/06/c6010.pdf.

Memarsadeghi:2013:NCC

[MR13] Nargess Memarsadeghi and Rafael Rincon. NASA computational case study: SAR data processing — groundrange projection. Computing in Science and Engineering, 15(6):92–95, November/December 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Muruganandham:2023:SGO

[MRHP23] Shivaprakash Muruganandham, Alexander A. Robel, Matthew J. Hoffman, and Stephen F. Price. Statistical generation of ocean forcing with spatiotemporal variability for ice sheet models. Computing in Science and Engineering, 25(3):30–41, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Michalski:2017:CVS

[MRKK17] Michael Michalski, Martin Rieth, Andreas Kempf, and Jens Kruger. CoFlaVis: A visualization system for pulverized coal flames. Computing in Science and Engineering, 19(6):72–78, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/06/mcs2017060072-
                                        abs.html.

Ma:2017:USW

[MRNT17] Yingjuan Ma, C. T. Russell, Andrew Nagy, and Gabor Toth. Understanding the solar wind–Mars interaction with global magnetohydrodynamic modeling. Computing in Science and Engineering, 19(4): 6–17, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/

cs/2017/04/mcs2017040006-
abs.html.

Madduri:2015:PPD

[MRU^{+}15] Ravi Madduri, Alex Rodriguez, Thomas Uram, Katrin Heitmann, Tanu Malik, Saba Sehrish, Ryan Chard, Shreyas Cholia, Marc Paterno, Jim Kowalkowski, and Salman Habib. PDACS: A portal for data analysis services for cosmological simulations. Computing in Science and Engineering, 17(5):18–26, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs2015050018-
abs.html.

Mesirov:1999:GEI

[MS99] Jill P. Mesirov and Donna K.

Slonim. Guest Editor’s introduction: Computational biology. Computing in Science and Engineering, 1 (3):16–17, May/June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3016.
pdf;   http://www.computer.
org/cse/cs1999/c3016abs.
htm.

Myers:2007:PEC

[MS07] Christopher R. Myers and James P. Sethna. Python for education: Computational methods for nonlinear systems. Computing in Science and Engineering, 9(3): 75–79, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Malony:2022:THP

[MS22] Allen D. Malony and Sameer S. Shende. Translating highperformance computing tools from research to practice: Experiences with the TAU performance system. Computing in Science and Engineering, 24(5):65–71, September/October 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mosterman:2014:HFV

[MSB^{+}14] Pieter J. Mosterman, David Escobar Sanabria, Enes Bilgin, Kun Zhang, and Justyna Zander. A heterogeneous fleet of vehicles for automated humanitarian missions. Computing in Science and Engineering, 16(3):90–95, May/ June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Murphy:2010:AAE

[MSD10] Richard Murphy, Thomas Sterling, and Chirag Dekate. Advanced architectures and execution models to support green computing. Computing in Science and Engineering, 12(6):38–47, November/December 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Murtagh:2002:DVI

[MSL02] Fionn Murtagh, Jean-Luc Starck, and Mireille Louys. Distributed visual information management in astronomy. Computing in Science and Engineering, 4(6):14–23, November/December 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/06/c6014abs.htm;
http://csdl.computer.org/
dl/mags/cs/2002/06/c6014.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/06/
c6014.pdf.

Mardal:2007:UPS

[MSL^{+}07] Kent-Andre Mardal, Ola Skavhaug, Glenn T. Lines, Gunnar A. Staff, and Åsmund Ødegård. Using Python to solve partial differential equations. Computing in Science and Engineering, 9(3): 48–51, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Monteleoni:2013:CIA

[MSM13] Claire Monteleoni, Gavin A.

Schmidt, and Scott McQuade. Climate informatics: Accelerating discovering in climate science with machine learning. Computing in Science and Engineering, 15(5):32– 40, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Morais:2015:VCS

[MSR15] Alessandra Marli M. Morais, Rafael D. C. Santos, and M. Jordan Raddick. Visualization of citizen science volunteers’ behaviors with data from usage logs. Computing in Science and Engineering, 17(4):42–50, July/ August 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040042-
abs.html.

Moyer:2016:NES

[MSR^{+}16] Thomas Moyer, Jonathan Stergiou, Garth Reese, James Luton, and Najib Abboud. Navy Enhanced Sierra Mechanics (NESM): Toolbox for predicting Navy shock and damage. Computing in Science and Engineering, 18(6):10–18, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs2016060010-
abs.html.

Myra:2009:SCC

[MSS09] Eric S. Myra, F. Douglas Swesty, and Dennis C. Smolarski. Stellar core collapse: a case study in the design of numerical algorithms for scalable radiation hydrodynamics. Computing in Science and Engineering, 11(2): 34–44, March/April 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Malamud:2000:CAM

[MT00] Bruce D. Malamud and Donald L. Turcotte. Cellularautomata models applied to natural hazards. Computing in Science and Engineering, 2(3):42–51, May/June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3042.
pdf;   http://www.computer.
org/cse/cs1999/c3042abs.
htm.

Malviya-Thakur:2022:RSE

[MTBG^{+}22] Addi Malviya-Thakur, David E.

Bernholdt, William F. Godoy, Gregory R. Watson, Mathieu Doucet, Mark A. Coletti, David M. Rogers, Marshall McDonnell, Jay Jay Billings, and Barney Maccabe. Research software engineering at Oak Ridge National Laboratory. Computing in Science and Engineering, 24(5):14– 23, September/October 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Matsuura:2012:EIB

[MTG^{+}12] Anne Matsuura, Nicola Thrupp, Xavier Gonze, Yann Pouillon, Gaelle Bruant, and Giovanni Onida. The ETSF: An e-infrastructure that bridges simulations and experiments. Computing in Science and Engineering, 14(1):22–32, January/February 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mucke:2009:QCC

[Muc09]  Ernst Mucke.  Quickhull:

Computing convex hulls quickly. Computing in Science and Engineering, 11(5):54–57, September/October 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Muhlberger:2023:CTR

[Muh23] Curran D. Muhlberger. Challenges and techniques for reproducible simulations. Computing in Science and Engineering, 25(4):42–46, April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Musk:2020:CQT

[Mus20] D. R. Musk. A comparison of quantum and traditional Fourier transform computations. Computing in Science and Engineering, 22(6):103–110, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Muzy:2019:EAM

[Muz19] A. Muzy. Exploiting activity for the modeling and simulation of dynamics and learning processes in hierarchical (neurocognitive) systems. Computing in Science and Engineering, 21(1):84–93, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Memarsadeghi:2020:VAR

[MV20] N. Memarsadeghi and A. Varshney. Virtual and augmented reality applications in science and engineering. Computing in Science and Engineering, 22 (3):4–6, May/June 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Martin-Villalba:2014:TAF

[MVUSK14] Carla Martin-Villalba, Alfonso Urquia, Yuri Senichenkov, and Yuri Kolesov. Two approaches to facilitate virtual lab implementation. Computing in Science and Engineering, 16(1):78–86, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Muller:2011:GRV

[MW11a] Thomas Muller and Daniel Weiskopf. General-relativistic visualization. Computing in Science and Engineering, 13(6):64–71, November/December 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Muller:2011:SRV

[MW11b] Thomas Muller and Daniel Weiskopf. Special-relativistic visualization. Computing in Science and Engineering, 13 (4):85–93, July/August 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Marmaras:2014:SVF

[MW14] Brett Marmaras and J. Jay Wang. Simulation and visualization of few-body systems and the differential precession of mercury. Computing in Science and Engineering, 16 (1):42–50, January/February 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mori:2016:MDH

[MWC^{+}16] Susumu Mori, Dan Wu, Can Ceritoglu, Yue Li, Anthony Kolasny, Marc A. Vaillant, Andreia V. Faria, Kenichi Oishi, and Michael I. Miller. MRICloud: Delivering highthroughput MRI neuroinformatics as cloud-based software as a service. Computing in Science and Engineering, 18(5): 21–35, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Matthews:2008:CML

[MWE08] David Matthews, Greg Wilson, and Steve Easterbrook. Configuration management for large-scale scientific computing at the UK Met Office. Computing in Science and Engineering, 10(6): 56–64, November/December 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Myers:1999:WGD

[Mye99] Gene Myers. Whole-genome DNA sequencing. Computing in Science and Engineering, 1(3):33–43, May/June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3033.
pdf;   http://www.computer.
org/cse/cs1999/c3033abs.
htm.

Nandagopal:2007:EFH

[NA07] Mohankumar Nandagopal and Natarajan Arunajadai. On the evaluation of finite Hilbert transforms. Computing in Science and Engineering, 9(6): 90–95, November/December 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Nair:2015:IAM

[Nai15] Kalyani Nair. An impressive anthology of microbiology modeling [book review]. Computing in Science and Engineering, 17(5):5–6, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/05/mcs2015050005.
                                        html.

Naji:2008:SCC

[Naj08] Hamid Reza Naji. Solving complex computational problems using multiagents implemented in hardware. Computing in Science and Engineering, 10(5):54–63, September/October 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Nandagopal:2011:ECT

[Nan11] Mohankumar Nandagopal. On the evaluation of correction terms of Gaussian integration. Computing in Science and Engineering, 13(1):58–61, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Nash:2000:DSM

[Nas00] John C. Nash. The (Dantzig) simplex method for linear programming. Computing in Science and Engineering, 2 (1):29–31, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1029.
pdf;   http://www.computer.
org/cse/cs1999/c1029abs.
htm.

Nakano:2001:MSN

[NBK^{+}01] Aiichiro Nakano, Martina E.

Bachlechner, Rajiv K. Kalia, Elefterios Lidorikis, Priya Vashishta, George Z. Voyiadjis, Timothy J. Campbell, Shuji Ogata, and Fuyucki Shimojo. Multiscale simulation of nanosystems. Computing in Science and Engineering, 3(4):56–66, July/August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c4056.
pdf;   http://www.computer.
org/cse/cs1999c4056abs.
htm.

Nakano:2003:GEI

[NC03] Aiichiro Nakano and Jim X.

Chen. Guest Editors’ introduction: High-dimensional data acquisition, computing, and visualization. Computing in Science and Engineering, 5(5):14–15, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/05/c5014.pdf;
http://csdl.computer.org/

                                        comp/mags/cs/2003/05/c5014abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/05/
                                        c5014.htm.

Nanthaamornphong:2018:TDD

[NC18] Aziz Nanthaamornphong and Jeffrey C. Carver. Test-driven development in HPC science: A case study. Computing in Science and Engineering, 20(5):98–113, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/05/mcs2018050098-
                                        abs.html.

Newman:2005:UPN

[NCB^{+}05] Harvey Newman, Richard Cavanaugh, Julian James Bunn, Iosif Legrand, Steven H. Low, Dan Nae, Sylvain Ravot, Conrad D. Steenberg, Xun Su, Michael Thomas, Frank van Lingen, Yang Xia, and Shawn McKee. The Ultra-Light project: The network as an integrated and managed resource for data-intensive science. Computing in Science and Engineering, 7(6): 38–47, November/December 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Nanthaamornphong:2014:BCT

[NCM^{+}14] Aziz Nanthaamornphong, Jeffrey C. Carver, Karla Morris, Hope A. Michelsen, and Damian W. I. Rouson. Building CLiiME via test-driven development: a case study. Computing in Science and Engineering, 16(3):36–46, May/ June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Neely:2017:AML

[NdS17] J. Robert Neely and Bronis R.

de Supinski. Application modernization at LLNL and the Sierra Center of Excellence. Computing in Science and Engineering, 19(5):9–18, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/05/mcs2017050009-
abs.html.

Neely:2017:AME

[NdSS17] J. Robert Neely, Bronis R.

de Supinski, and Charles H. Still. Application modernization for the exascale era. Computing in Science and Engineering, 19(5):6–8, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/05/mcs2017050006.
html.

Neilsen:2008:SDS

[Nei08] Eric H. Neilsen, Jr. The Sloan Digital Sky Survey data archive server. Computing in Science and Engineering, 10 (1):13–17, January/February 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Newman:2000:CSS

[New00] Mark Newman. Computer simulations: Simple models of evolution and extinction. Computing in Science and Engineering, 2(1):80– 86, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c1080.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c1080abs.
                                        htm.

Nakajima:2020:LLT

[NG20] T. M. Nakajima and J. Goode. Lighting up learning: Teachers’ pedagogical approaches for mak(e)ing computing culturally responsive in electronictextiles classrooms. Computing in Science and Engineering, 22(5):41–50, September/October 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Navratil:2022:ARH

[NGGS22] Paul Navrátil, Christiaan Gribble, Pascal Grosset, and John E. Stone. Analytic rendering and hardwareaccelerated simulation for scientific applications. Computing in Science and Engineering, 24(2):4–6, March/April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Nakano:1999:SMD

[NKV99] Aiichiro Nakano, Rajiv K.

Kalia, and Priya Vashishta. Scalable molecular-dynamics, visualization, and data-management algorithms for material simulations. Computing in Science and Engineering, 1(5): 39–47, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5039.
pdf;   http://www.computer.
org/cse/cs1999/c5039abs.
htm.

Northover:1999:GEI

[NL99] Kevin Northover and Andrew W. Lo. Guest Editors’ introduction: Computational finance. Computing in Science and Engineering, 1(6): 22–23, November/December 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c6022.
pdf;   http://www.computer.
org/cse/cs1999/c6022abs.
htm.

Navarro:2013:FBV

[NLGNJ13] Denis Navarro, Oscar Lucia, Jose M. Gil-Narvion, and Oscar Jimenez. FPGA-based virtual screening acceleration of rigid-molecule docking. Computing in Science and Engineering, 15(6):64–72, November/December 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

North:1999:LEY

[NLV99] Joseph ‘Roy’ North, Martin Lo, and Gregory S. Vigneault. Letters to the editor: You’re off your orbit; Internet time: Why switch for swatch? Computing in Science and Engineering, 1(3):5, May/June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c3005.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c3005abs.
                                        htm.

Nash:2022:MEE

[NMCM22] Jessica A. Nash, Mohammad Mostafanejad, T. Daniel Crawford, and Ashley Ringer McDonald. MolSSI education: Empowering the next generation of computational molecular scientists. Computing in Science and Engineering, 24(3):72–76, May/ June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Nagy:2003:IDC

[NO03] James G. Nagy and Dianne P. O’Leary. Image deblurring: I can see clearly now. Computing in Science and Engineering, 5(3): 82–84, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/03/c3082abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/03/c3082.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/03/
c3082.pdf.

Noble:2000:CPG

[Nob00a] Julian V. Noble. Computing prescriptions: Gauss– Legendre principal value integration. Computing in Science and Engineering, 2 (1):92–95, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1092.
pdf;   http://www.computer.
org/cse/cs1999/c1092abs.
htm.

Noble:2000:TNR

[Nob00b] Julian V. Noble. Technology news & reviews: Adventures in the Forth dimension. Computing in Science and Engineering, 2(5):6– 10, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c5006.
pdf.

Noble:2002:FM

[Nob02a] Julian V. Noble. The full Monte. Computing in Science and Engineering, 4(3): 76–81, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/03/c3076abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/03/c3076.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/03/
                                        c3076.pdf.

Noble:2002:RAP

[Nob02b]  Julian V. Noble.  The

right angle: Precise numerical orthogonality in eigenstates. Computing in Science and Engineering, 4(5): 91–96, c3, September/October 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2002/05/c5091.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/05/
                                        c5091.pdf.

Noble:2003:R

[Nob03]  Julian V. Noble.  Re-

curses! Computing in Science and Engineering, 5(3): 76–81, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/03/c3076abs.htm;

 http://csdl.computer.org/
dl/mags/cs/2003/03/c3076.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/03/
c3076.pdf.

Noble:2007:MCS

[Nob07] Julian V. Noble. Making the complex simple. Computing in Science and Engineering, 9(3): 84–89, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Nam:2017:TCE

[NRG^{+}17] Hai Ah Nam, Gabriel Rockefeller, Mike Glass, Shawn Dawson, John Levesque, and Victor Lee. The Trinity Center of Excellence co-design best practices. Computing in Science and Engineering, 19(5):19–26, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/05/mcs2017050019-
abs.html.

Norman:1999:DDD

[NSLD99] Michael L. Norman, John Shalf, Stuart Levy, and Greg Daues. Diving deep: Data management and visualization strategies for adaptive mesh refinement simulations. Computing in Science and Engineering, 1(4): 36–47, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c4036.
                                        pdf.

Nonato:2012:UCM

[NSP12] Luis Gustavo Nonato, Claudio T. Silva, and Fernando V. Paulovich. User-centered multidimensional projection techniques. Computing in Science and Engineering, 14(4): 74–81, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Nandagopal:2010:NEF

[NSR10] Mohankumar Nandagopal, Soubhadra Sen, and Ajay Rawat. A note on the error function. Computing in Science and Engineering, 12(4): 84–88, July/August 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). See improvements [Iac21].

Newman:2007:SFT

[NTW07] David E. Newman, Paul W. Terry, and Andrew S. Ware. Shear flows and turbulence in nature. Computing in Science and Engineering, 9(6): 45–52, November/December 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Netzel:2013:TBF

[NW13] Rudolf Netzel and Daniel Weiskopf. Texture-based flow visualization. Computing in Science and Engineering, 15(6):96–102, November/December 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Niyonkuru:2015:DEM

[NW15] Daniella Niyonkuru and Gabriel A.

Wainer. Discrete-event modeling and simulation for embedded systems. Computing in Science and Engineering, 17(5):52–63, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs2015050052-
abs.html.

Niu:2019:ELM

[NWP19] X. Niu, Z. Wang, and Z. Pan.

Extreme learning machinebased deep model for human activity recognition with wearable sensors. Computing in Science and Engineering, 21(5):16–25, September/October 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Olivares-Amaya:2009:CGZ

[OASFLAB09] Roberto Olivares-Amaya, Romelia Salomon-Ferrer, William A.

Lester Jr., and Carlos Amador-Bedolla. Creating a GUI for Zori, a quantum Monte Carlo program. Computing in Science and Engineering, 11 (1):41–47, January/February 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Olshansky:2023:CIO

[OHK23] Alex Olshansky, Cassandra Hayes, and Kerk F. Kee. The characteristics of influencers and opinion leaders of science gateways and cyberinfrastructure for innovation diffusion. Computing in Science and Engineering, 25(2): 30–39, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Osburn:2010:DA

[OKS10] Jeanie Osburn, Aram Kevorkian, and Balu Sekar. Defense applications. Computing in Science and Engineering, 12(5):14– 17, September/October 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2004:ETT

[O’L04a] Dianne P. O’Leary. Elastoplastic torsion: Twist and stress. Computing in Science and Engineering, 6(4): 74–83, July/August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/04/c4074.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/04/
c4074.pdf.

OLeary:2004:FEI

[O’L04b] Dianne P. O’Leary. Fitting exponentials: An interest in rates. Computing in Science and Engineering, 6(3): 66–72, May/June 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/03/c3066abs.htm;
http://csdl.computer.org/
dl/mags/cs/2004/03/c3066.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/03/
c3066.pdf.

OLeary:2004:MIPa

[O’L04c] Dianne P. O’Leary. Models of infection: Person to person. Computing in Science and Engineering, 6(1):68–73, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/01/c1068abs.htm;
http://csdl.computer.org/
dl/mags/cs/2004/01/c1068.
pdf.

OLeary:2004:MMI

[O’L04d] Dianne P. O’Leary. More models of infection: It’s epidemic. Computing in Science and Engineering, 6(2):

50–56, March/April 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/02/c2050abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2004/02/c2050.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/02/
c2050.pdf.

OLeary:2004:MIPb

[O’L04e] Dianne P. O’Leary. Multidimensional integration: Partition and conquer. Computing in Science and Engineering, 6(6):58–66, November/December 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/06/c6058.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/06/
c6058.pdf.

OLeary:2005:EVP

[O’L05a]  D. P. O’Leary.  Eigen-

values: valuable principles. Computing in Science and Engineering, 7(4): 68–70, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463139.pdf?
isnumber=31456&prod=JNL&
arnumber=1463139&arSt=+
68&ared=+70&arAuthor=O%
27Leary%2C+D.P.;   http:/

/ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=31456&
arnumber=1463139&count=
14&index=9.

OLeary:2005:GDR

[O’L05b]  D. P. O’Leary.  George

Dantzig: a remembrance. Computing in Science and Engineering, 7(5):96, September/October 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501749.pdf?
isnumber=32219&prod=JNL&
arnumber=1501749&arSt=+
96&ared=+96&arAuthor=+O%
27Leary%2C+D.P.;   http:/
/ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=32219&
arnumber=1501749&count=
14&index=13.

OLeary:2005:SSL

[O’L05c] D. P. O’Leary. Solving sparse linear systems: Taking the direct approach. Computing in Science and Engineering, 7(5):62–70, September/October 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501742.pdf?
isnumber=32219&prod=JNL&
arnumber=1501742&arSt=+
62&ared=+70&arAuthor=+O%
27Leary%2C+D.P.;   http:/
/ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=32219&

                                        arnumber=1501742&count=
                                        14&index=8.

OLeary:2005:BDM

[O’L05d] Dianne P. O’Leary. Blind deconvolution: a matter of norm. Computing in Science and Engineering, 7(2): 60–66, March/April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/02/c2060.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2005/02/
                                        c2060.pdf.

OLeary:2005:BDE

[O’L05e] Dianne P. O’Leary. Blind deconvolution: Errors, errors everywhere. Computing in Science and Engineering, 7 (1):56–63, January/February 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2005/01/c1056.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2005/01/
                                        c1056.pdf.

OLeary:2005:FSS

[O’L05f] Dianne P. O’Leary. Fast solvers and Sylvester equations: Both sides now. Computing in Science and Engineering, 7(6):74–80, November/December 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2005:FDF

[O’L05g]  Dianne P. O’Leary.  Fi-

nite differences and finite elements: Getting to know you. Computing in Science and Engineering, 7(3): 72–79, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/03/c3072abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2005/03/c3072.
pdf.

OLeary:2006:CSW

[O’L06a] Dianne P. O’Leary. Computational software: Writing your legacy. Computing in Science and Engineering, 8(1):78–83, January/February 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2006:CMA

[O’L06b] Dianne P. O’Leary. Computer memory and arithmetic: a look under the hood. Computing in Science and Engineering, 8(3):54–57, May/ June 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2006:IML

[O’L06c] Dianne P. O’Leary. Iterative methods for linear systems: Following the meandering way. Computing in Science and Engineering, 8(4):

74–80, July/August 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2006:MMM

[O’L06d] Dianne P. O’Leary. Multigrid methods: Managing massive meshes. Computing in Science and Engineering, 8(5): 96–103, September/October 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2006:PSL

[O’L06e] Dianne P. O’Leary. Partial solution to last issue’s homework assignment: Updating and downdating matrix factorizations: a change in plans. Computing in Science and Engineering, 8(3):58–59, May/ June 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2006:SAW

[O’L06f] Dianne P. O’Leary. Sensitivity analysis: When a little means a lot. Computing in Science and Engineering, 8(6):70–75, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2006:UDM

[O’L06g] Dianne P. O’Leary. Updating and downdating matrix factorizations: a change in plans. Computing in Science and Engineering, 8(2): 66–71, March/April 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Oosterlee:2006:MMS

[OL06h] Cornelis W. Oosterlee and Francisco José Gaspar Lorenz. Multigrid methods for the Stokes system. Computing in Science and Engineering, 8(6):34–43, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2007:BCC

[O’L07a] Dianne P. O’Leary. Beetles, cannibalism, and chaos: Analyzing a dynamical system model. Computing in Science and Engineering, 9 (2):96–103, March/April 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2007:PSL

[O’L07b] Dianne P. O’Leary. A partial solution to last issue’s homework assignment: Beetles, cannibalism, and chaos– analyzing a dynamical system model. Computing in Science and Engineering, 9(3): 80–83, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2012:VGT

[O’L12] Dianne P. O’Leary. Variablegeometry trusses: What’s your angle? Computing in Science and Engineering, 14(6):76–79, November/December 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2013:PM

[O’L13] Dianne P. O’Leary. Phantom measurements. Computing in Science and Engineering, 15(4):57–59, July/August 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2014:WSM

[O’L14] Dianne P. O’Leary. What’s the score? Matrices, documents, and queries. Computing in Science and Engineering, 16(2):63–67, March/April 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Oliphant:2007:PSC

[Oli07] Travis E. Oliphant. Python for scientific computing. Computing in Science and Engineering, 9(3):10–20, May/ June 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Oliveira:2013:TMA

[Oli13] Manuel M. Oliveira. Towards more accessible visualizations for color-vision-deficient individuals. Computing in Science and Engineering, 15(5):80– 87, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2003:DAP

[OM03] Dianne P. O’Leary and Nargess Memarsadeghi. The direction-of-arrival problem: Coming at you. Computing in Science and Engineering, 5(6):60–70, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/06/c6060abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/06/c6060.
pdf.

Omerbashich:2006:GVC

[Ome06] Mensur Omerbashich. Gauss– Vanícek spectral analysis of the Sepkoski Compendium: No new life cycles. Computing in Science and Engineering, 8(4):26–30, July/August 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Oliveira:2011:PCA

[OMKdSB11] Luiz Oliveira, Marcelo Mansano, Alessandro Koerich, and Alceu de Souza Britto, Jr. 2D principal component analysis for face and facial-expression recognition. Computing in Science and Engineering, 13 (3):9–13, May/June 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ong:2002:MRS

[Ong02] Emil Ong. MPI Ruby: Scripting in a parallel environment. Computing in Science and Engineering, 4(4): 78–82, July/August 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/04/c4078abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/04/c4078.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/04/
                                        c4078.pdf.

Onofri:2001:ECM

[Ono01] Enrico Onofri. Elementary celestial mechanics using Matlab. Computing in Science and Engineering, 3(6):48–53, November/December 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c6048abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2001/pdf/c6048.pdf.

Omar:2017:DHS

[OOB17] Yasser M. K. Omar, Hoda Osama, and Amr Badr. Double hashing sort algorithm. Computing in Science and Engineering, 19(2):

63–69, March/April 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/02/mcs2017020063-
abs.html.

OLeary:2012:LSC

[OR12] Dianne P. O’Leary and Bert W. Rust. Light a single candle: Studying supernovae. Computing in Science and Engineering, 14(2): 90–96, March/April 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Orf:2021:MWM

[Orf21] Leigh Orf. Modeling the world’s most violent thunderstorms. Computing in Science and Engineering, 23(3): 14–24, May/June 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Orkoulas:2009:OMC

[Ork09] Gerassimos Orkoulas. An overview of Monte Carlo methods for fluid simulation. Computing in Science and Engineering, 11(5):76– 87, September/October 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

OLeary:2003:RCS

[OS03] Dianne O’Leary and Yalin E.

Sagduyu. Robot control:

Swinging like a pendulum. Computing in Science and Engineering, 5(4):68–74, July/ August 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/dl/mags/cs/
2003/04/c4068.htm;   http:
//csdl.computer.org/dl/
mags/cs/2003/04/c4068.pdf.

OLeary:2004:ACV

[OS04] Dianne P. O’Leary and David A. Schug. Achieving a common viewpoint: Yaw, pitch, and roll. Computing in Science and Engineering, 6(5): 60–65, September/October 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/05/c5060.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/05/
c5060.pdf.

Oskoee:2007:CPM

[Osk07] Ehsan Nedaaee Oskoee. Computing properties of materials based on the Ginzburg– Landau equation. Computing in Science and Engineering, 9(2):84–95, March/April 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Oliver:2019:WAC

[OSM^{+}19] H. Oliver, M. Shin, D. Matthews, O. Sanders, S. Bartholomew, A. Clark, B. Fitzpatrick, R. van Haren, N. Drost, and R. Hut. Workflow automation for cycling systems. Computing in Science and Engineering, 21(4):7–21, July/August 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Ott:2016:MCU

[Ott16] Christian D. Ott. Massive computation for understanding core-collapse supernova explosions. Computing in Science and Engineering, 18(5): 78–92, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Oughstun:2003:CMU

[Oug03] Kurt Edmund Oughstun.

Computational methods in ultrafast time-domain optics. Computing in Science and Engineering, 5(6):22–32, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/06/c6022abs.htm;
http://csdl.computer.org/
dl/mags/cs/2003/06/c6022.
pdf.

OLeary:2001:CRI

[OW01] Dianne P. O’Leary and Scott T. Weidman. Conference report: The interface between computer science and the mathematical sciences. Computing in Science and Engineering, 3(3):60–65, May/ June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3060.
pdf;   http://www.computer.
org/cse/cs1999c3060abs.
htm.

Owen:2001:OSP

[Owe01] J. Michael Owen. An opensource project for modeling hydrodynamics in astrophysical systems. Computing in Science and Engineering, 3(6):54–59, November/December 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c6054abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c6054.pdf.

Pell:2012:MPC

[PA12] Oliver Pell and Vitali Averbukh. Maximum performance computing with dataflow engines. Computing in Science and Engineering, 14(4):98– 103, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Parashar:2021:TCS

[PA21] Manish Parashar and David Abramson. Translational computer science for science and engineering. Computing in Science and Engineering, 23 (5):5–6, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Parashar:2022:ATP

[PA22] Manish Parashar and David Abramson. Accidental translationists: a perspective from the trenches. Computing in Science and Engineering, 24 (4):70–75, July/August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Padua:2000:FC

[Pad00] David Padua. The Fortran I compiler. Computing in Science and Engineering, 2 (1):70–75, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1070.
pdf;   http://www.computer.
org/cse/cs1999/c1070abs.
htm.

Pandey:2008:SED

[PAF08] Ras B. Pandey, Kelly L. Anderson, and Barry L. Farmer. Sheets: Entropy dissipation, multiscale dynamics, dispersion, and intercalation. Computing in Science and Engineering, 10(6):90–98, November/December 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Post:2016:CREa

[PAN^{+}16a] D. Post, C. Atwood, K. Newmeyer, S. Landsberg, and F. Shull. The Computational Research and Engineering Acquisition Tools and Environments (CREATE) program. Computing in Science and Engineering, 18(1):10–13, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Post:2016:CSE

[PAN^{+}16b] D. E. Post, C. A. Atwood, K. P. Newmeyer, R. L. Meakin, M. M. Hurwitz, S. Dey, J. N. D’Angelo, R. L. Vogelsong, N. S. Hariharan, R. P. Kendall, O. A. Goldfarb, and L. K. Miller. CREATE: Software engineering applications for the design and analysis of air vehicles, Naval vessels, and radio frequency antennas. Computing in Science and Engineering, 18(1):14–24, January/February 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Papka:2016:ESH

[Pap16] Michael E. Papka. Expanding the scope of high-performance computing facilities. Computing in Science and Engineering, 18(3):84–87, May/ June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Parlett:2000:A

[Par00a] Beresford N. Parlett. The QR algorithm. Computing in Science and Engineering, 2 (1):38–42, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1038.
pdf;   http://www.computer.
org/cse/cs1999/c1038abs.
htm.

Parrinello:2000:SCS

[Par00b] Michele Parrinello. Simulating complex systems without adjustable parameters. Computing in Science and Engineering, 2(6):22–27, November/December 2000. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c6022.
pdf;   http://www.computer.
org/cse/cs1999c6022abs.
htm.

Parigger:2012:CPA

[Par12] Christian G. Parigger. Computational physics activities at the University of Tennessee Space Institute. Computing in Science and Engineering, 14(3):94–97, May/June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Park:2016:SMS

[Par16] Daejin Park. A scalable modeling and simulation environment for chemical gas emergencies. Computing in Science and Engineering, 18(4): 25–33, July/August 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Parashar:2022:JDC

[Par22] Manish Parashar. Jack Dongarra: Catalyzing the transformation of high-performance computing. Computing in Science and Engineering, 24(4): 4–5, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Parashar:2013:CPP

[PARD13] Manish Parashar, Moustafa AbdelBaky, Ivan Rodero, and Aditya Devarakonda. Cloud paradigms and practices for computational and data-enabled science and engineering. Computing in Science and Engineering, 15(4): 10–18, July/August 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Patterson:2002:TAT

[Pat02] Charles Patterson. Two approaches to teaching computational physics. Computing in Science and Engineering, 4(6): 64–68, November/December 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2002/06/c6064.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/06/
c6064.pdf.

Pouillon:2011:OSG

[PBD^{+}11] Yann Pouillon, Jean-Michel Beuken, Thierry Deutsch, Marc Torrent, and Xavier Gonze. Organizing software growth and distributed development: The case of Abinit. Computing in Science and Engineering, 13(1):62–69, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ponciano:2014:VEH

[PBSS14] Lesandro Ponciano, Francisco Brasileiro, Robert Simpson, and Arfon Smith. Volunteers’ engagement in human computation for astronomy projects. Computing in Science and Engineering, 16(6):52–59, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs2014060052-
abs.html.

Pradal:2019:VVU

[PCBVS19] C. Pradal, S. Cohen-Boulakia, P. Valduriez, and D. Shasha. VersionClimber: Version upgrades without tears. Computing in Science and Engineering, 21(5):87–93, September/October 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Peng:2014:IDI

[PCY14] Yi Peng, Li Chen, and Jun-Hai Yong. Importance-driven isosurface decimation for visualization of large simulation data based on OpenCL. Computing in Science and Engineering, 16(1):24–32, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pope:2002:TEA

[PD02] Vicky Pope and Terry Davies. Testing and evaluating atmospheric climate models. Computing in Science and Engineering, 4(5):64–69, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/05/c5064abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/05/c5064.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/05/
                                        c5064.pdf.

Peng:2009:DRR

[PE09] Roger D. Peng and Sandrah P. Eckel. Distributed reproducible research using cached computations. Computing in Science and Engineering, 11 (1):28–34, January/February 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pegg:2012:NCC

[Peg12]  Ed Pegg, Jr.  Number-

crunching could have crunched more. Computing in Science and Engineering, 14 (6):6–7, November/December 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pekalski:2004:SGP

[Pek04] Andrzej Pekalski. A short guide to predator–prey lattice models. Computing in Science and Engineering, 6(1):62–66, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/01/c1062abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2004/01/c1062.
pdf.

Peoples:2020:RBE

[Peo20]  C. Peoples.  Research-

based education on a Master of Science degree in professional software development. Computing in Science and Engineering, 22(4): 36–44, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Peskin:2003:NPS

[Pes03] Michael E. Peskin. Numerical problem solving for undergraduate core courses. Computing in Science and Engineering, 5(6):92–96, C3, November/December 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2003/06/c6092.
                                        pdf.

Peyre:2011:NTSa

[Pey11a]  Gabriel Peyre.  The nu-

merical tours of signal processing. Computing in Science and Engineering, 13(4): 94–97, July/August 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Peyre:2011:NTSb

[Pey11b] Gabriel Peyre. The numerical tours of signal processing part 2: Multiscale processings. Computing in Science and Engineering, 13(5):68– 71, September/October 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Peyre:2011:NTSc

[Pey11c]  Gabriel Peyre.  The nu-

merical tours of signal processing part 3: Image and surface restoration. Computing in Science and Engineering, 13(6):72–75, November/December 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pierce:2004:MSA

[PF04] Marlon Pierce and Geoffrey Fox. Making scientific applications as Web services. Computing in Science and Engineering, 6(1):93–96, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/01/c1093abs.htm;
http://csdl.computer.org/
dl/mags/cs/2004/01/c1093.
pdf.

PiazentinOno:2021:IDV

[PFS21] J. Piazentin Ono, J. Freire, and C. T. Silva. Interactive data visualization in Jupyter notebooks. Computing in Science and Engineering, 23 (2):99–106, March/April 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Perez:2007:ISI

[PG07] Fernando Pérez and Brian E.

Granger. IPython: a system for interactive scientific computing. Computing in Science and Engineering, 9(3): 21–29, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Post:2017:EPI

[PG17] Douglass E. Post and Oscar Goldfarb. Enhancing product innovation with computational engineering. Computing in Science and Engineering, 19(6):4–5, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060004.
html.

Ploeg:2021:SOS

[PGC21] S. Ploeg, H. Gunther, and R. M. Camacho. Simphony: An open-source photonic integrated circuit simulation framework. Computing in Science and Engineering, 23 (1):65–74, January/February 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Palmer:2015:OXT

[PGF^{+}15] Jeffrey T. Palmer, Steven M. Gallo, Thomas R. Furlani, Matthew D. Jones, Robert L. DeLeon, Joseph P. White, Nikolay Simakov, Abani K. Patra, Jeanette Sperhac, Thomas Yearke, Ryan Rathsam, Martins Innus, Cynthia D. Cornelius, James C. Browne, William L. Barth, and Richard T. Evans. Open XDMoD: A tool for the comprehensive management of high-performance computing resources. Computing in Science and Engineering, 17(4): 52–62, July/August 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040052-
abs.html.

Plale:2005:CSD

[PGH^{+}05] B. Plale, D. Gannon, Yi Huang, G. Kandaswamy, S. Lee Pallickara, and A. Slominski. Cooperating services for data-driven computational experimentation. Computing in Science and Engineering, 7(5):34–43, September/October 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501738.pdf?
isnumber=32219&prod=JNL&
arnumber=1501738&arSt=+
34&ared=+43&arAuthor=+Plale%
2C+B.%3B++Gannon%2C+D.%
3B++Yi+Huang%3B++Kandaswamy%
2C+G.%3B++Lee+Pallickara%
2C+S.%3B++Slominski%2C+
A.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=32219&arnumber=
1501738&count=14&index=5.

Perez:2011:PES

[PGH11] Fernando Perez, Brian E.

Granger, and John D. Hunter. Python: An ecosystem for scientific computing. Computing in Science and Engineering, 13(2):13–21, March/April 2011. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Pope:2010:AU

[PHL^{+}10] Adrian Pope, Salman Habib, Zarija Lukic, David Daniel, Patricia Fasel, Nehal Desai, and Katrin Heitmann. The accelerated universe. Computing in Science and Engineering, 12(4):17–25, July/August 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pierce:2021:ECP

[PHW^{+}21] K. A. Pierce, E. Ho, X. Wang, R. Pasco, Z. Du, G. Zynda, J. Song, G. Wells, S. J. Fox, and L. Ancel Meyers. Early COVID-19 pandemic modeling: Three compartmental model case studies from Texas, USA. Computing in Science and Engineering, 23(1):25–34, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Piskor-Ignatowicz:2016:ABC

[PI16]  Cezary  Piskor-Ignatowicz.

Agent-based creation and simulation of artificial social networks and the analysis of their properties. Computing in Science and Engineering, 18(4): 34–41, July/August 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pierce:2004:BMV

[Pie04] David W. Pierce. Beyond the means: Validating climate models with higherorder statistics. Computing in Science and Engineering, 6(5): 22–29, September/October 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/05/c5022.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/05/
c5022.pdf.

Post:2015:EEP

[PK15] Douglass Post and Richard Kendall. Enhancing engineering productivity. Computing in Science and Engineering, 17(4):4–6, July/ August 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/04/mcs2015040004.
html.

Post:2018:SEU

[PK18] Douglass Post and Richard Kendall. A software epiphany by the US defense community might provide an unexpected boost for agile software development. Computing in Science and Engineering, 20(4): 5–7, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/

                                        cs/2018/04/mcs2018040005.
                                        html.

Pancratov:2008:WCAa

[PKST08a] Cosmin Pancratov, Jacob M. Kurzer, Kelly A. Shaw, and Matthew L. Trawick. Why computer architecture matters. Computing in Science and Engineering, 10(3): 59–63, May/June 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pancratov:2008:WCAb

[PKST08b] Cosmin Pancratov, Jacob M. Kurzer, Kelly A. Shaw, and Matthew L. Trawick. Why computer architecture matters: Memory access. Computing in Science and Engineering, 10(4):71–75, July/August 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pancratov:2008:WCAc

[PKST08c] Cosmin Pancratov, Jacob M. Kurzer, Kelly A. Shaw, and Matthew L. Trawick. Why computer architecture matters: Thinking through tradeoffs in your code. Computing in Science and Engineering, 10(5):74–79, September/October 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Peterkin:2002:VPE

[PL02] Robert E. Peterkin and John W. Luginsland. A virtual prototyping environment for directed-energy concepts. Computing in Science and Engineering, 4(2): 42–49, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c2042abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c2042.pdf.

Pan:2007:BBR

[PL07] Zhigeng Pan and Jianfeng Lu.

A Bayes-based region-growing algorithm for medical image segmentation. Computing in Science and Engineering, 9(4): 32–38, July/August 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pasquier:2018:SPC

[PLH^{+}18] Thomas Pasquier, Matthew K.

Lau, Xueyuan Han, Elizabeth Fong, Barbara S. Lerner, Emery R. Boose, Merce Crosas, Aaron M. Ellison, and Margo Seltzer. Sharing and preserving computational analyses for pos-

terity with encapsulator.

Computing in Science and Engineering, 20(4):111–124, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040111-
abs.html.

Peng:2017:MAC

[PLW17] Tao Peng, Qin Liu, and Guojun Wang. A multilevel access control scheme for data security in transparent computing. Computing in Science and Engineering, 19(1):46–53, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/01/mcs2017010046-
                                        abs.html.

Polanco-Martinez:2014:PWD

[PMFM14] Josue M. Polanco-Martinez and F. Javier Fernandez-Macho. Package W2CWM2C: Description, features, and applications. Computing in Science and Engineering, 16(6):68–78, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2014/06/mcs2014060068-
                                        abs.html.

Pennington:2008:TST

[PMK^{+}08] Deana D. Pennington, William K. Michener, Samantha Katz, Laura L. Downey, and Mark Schildhauer. Transforming scientists through technical education: a view from the trenches. Computing in Science and Engineering, 10(5): 28–33, September/October 2008. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Pletzer:2008:EFD

[PMM^{+}08] Alexander Pletzer, Douglas McCune, Stefan Muszala, Srinath Vadlamani, and Scott Kruger. Exposing Fortran derived types to C and other languages. Computing in Science and Engineering, 10(4): 86–92, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Plale:2021:RPH

[PMP21] Beth A. Plale, Tanu Malik, and Line C. Pouchard. Reproducibility practice in high-performance computing: Community survey results. Computing in Science and Engineering, 23(5):55– 60, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Przedzinski:2020:SDS

[PMW20] T. Przedzinski, M. Malawski, and Z. Wa s. Software development strategies for highenergy physics simulations based on quantum field theory. Computing in Science and Engineering, 22(4):86– 98, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Post:2016:CREb

[PNL^{+}16] Douglass Post, Kevin Newmeyer, Sandra Landsberg, Forrest Shull, and Scott Sundt. The Computational Research and Engineering Acquisition Tools and Environments (CREATE) Program, part 2. Computing in Science and Engineering, 18(6):7–9, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs2016060007.
html.

Poinot:2010:FGR

[Poi10]  Marc Poinot.  Five good

reasons to use the Hierarchical Data Format. Computing in Science and Engineering, 12(5):84–90, September/October 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pokorny:2004:WSI

[Pok04]  Jaroslav Pokorný.  Web

searching and information retrieval. Computing in Science and Engineering, 6(4): 43–48, July/August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/04/c4043.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/04/
c4043.pdf.

Pong:2016:SR

[Pon16] Ting Kei Pong. Social resistance. Computing in Science and Engineering, 18(2):98– 103, March/April 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Post:2004:GEIa

[Pos04a] Douglass Post. Guest Editor’s introduction: Frontiers of simulation. Computing in Science and Engineering, 6(2): 12–13, March/April 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/02/c2012.pdf;
http://csdl.computer.org/
comp/mags/cs/2004/02/c2012abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/02/
c2012.htm.

Post:2004:GEIb

[Pos04b] Douglass Post. Guest Editor’s introduction: Frontiers of simulation, Part II. Computing in Science and Engineering, 6(3): 16–17, May/June 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/03/c3016.pdf;
http://csdl.computer.org/
comp/mags/cs/2004/03/c3016abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/03/
c3016.htm.

Post:2007:GEI

[Pos07] Douglass E. Post. Guest Editor’s introduction: Computational science and engineering for the US department of defense. Computing in Science and Engineering, 9(6):10–11, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2007/06/mcs2007060010.
                                        pdf.

Post:2009:PSB

[Pos09] Douglass Post. The promise of science-based computational engineering. Computing in Science and Engineering, 11 (3):3–4, May/June 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Post:2010:CEP

[Pos10] Douglass E. Post. Computational engineering: Its promise and challenges. Computing in Science and Engineering, 12(3):12–13, May/June 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Post:2011:FCP

[Pos11]  Douglass Post.  The fu-

ture of computing performance. Computing in Science and Engineering, 13(4): 4–5, July/August 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Post:2013:CFS

[Pos13] Douglass Post. The changing face of scientific and engineering computing. Computing in Science and Engineering, 15 (6):4–6, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Post:2014:PDV

[Pos14]  Douglass Post.  Product

development with virtual prototypes. Computing in Science and Engineering, 16(6):4–7, November/ December 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs2014060004.
html.

Post:2016:PTE

[Pos16] Douglass E. Post. The periodic table of elements, an early example of “Big Data”. Computing in Science and Engineering, 18(3):4–7, May/June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pine:2020:CMC

[PP20] P. Pine and L. I. Paina. Computational methods in chemistry and biochemistry education: Visualization of proteins. Computing in Science and Engineering, 22(4): 45–49, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Parker:2000:MCA

[PPE00] D. Stott Parker, Brad Pierce, and Paul R. Eggert. Monte Carlo arithmetic: How to gamble with floating point and win. Computing in Science and Engineering, 2 (4):58–68, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c4058.
pdf;   http://www.computer.
org/cse/cs1999/c4058abs.
htm.

Pinto:2020:ECT

[PQQ20] J. D. Pinto, C. Quintana, and R. M. Quintana. Exemplifying computational thinking scenarios in the age of COVID-19: Examining the pandemic’s effects in a project-based MOOC. Computing in Science and Engineering, 22(6): 97–102, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Peterson:2001:CST

[PR01] Mark A. Peterson and Yanir Rubinstein. Computer simulations: Turbulence on a desktop. Computing in Science and Engineering, 3 (3):86–94, May/June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3086.
pdf;   http://www.computer.
org/cse/cs1999c3086abs.
htm.

Patel:2022:RFI

[PRCL^{+}22] Ria Patel, Brandan Roachell, Silvina Caíno-Lores, Ross Ketron, Jacob Leonard, Nigel Tan, Karan Vahi, Duncan A. Brown, Ewa Deelman, and Michela Taufer. Reproducibility of the first image of a black hole in the galaxy M87 from the event horizon telescope collaboration. Computing in Science and Engineering, 24(5):42–52, September/October 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Press:2009:PAO

[Pre09] William H. Press. That perfect algorithm for any occasion. Computing in Science and Engineering, 11 (5):5–6, September/October 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Post:2002:GEI

[PS02] Douglass E. Post and Francis Sullivan. Guest Editors’ introduction: Limits on computation. Computing in Science and Engineering, 4(3): 14–15, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/03/c3014abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/03/c3014.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/03/
c3014.pdf.

Post:2017:CAA

[PS17] Douglass Post and Scott Sundt. CREATE: Acceptance and adoption of virtual prototyping across the defense R&D and acquisition communities. Computing in Science and Engineering, 19(6):6–8, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060006.
html.

Petra:2014:RTS

[PSA14] Cosmin G. Petra, Olaf Schenk, and Mihai Anitescu. Real-time stochastic optimization of complex energy systems on highperformance computers. Computing in Science and Engineering, 16(5):32–42, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/

cs/2014/05/mcs2014050032-
abs.html.

Pennycook:2021:NPP

[PSJ^{+}21] S. John Pennycook, Jason D.

Sewall, Douglas W. Jacobsen, Tom Deakin, and Simon McIntosh-Smith. Navigating performance, portability, and productivity. Computing in Science and Engineering, 23(5):28–38, September/October 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Preston:2000:MEF

[PSR^{+}00] Eric F. Preston, Jorge S.

SáMartins, John B. Rundle, Marian Anghel, and William Klein. Models of earthquake faults with long-range stress transfer. Computing in Science and Engineering, 2(3):34–41, May/June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3034.
pdf;   http://www.computer.
org/cse/cs1999/c3034abs.
htm.

Parashar:2020:VDC

[PSR^{+}20] M. Parashar, A. Simonet, I. Rodero, F. Ghahramani, G. Agnew, R. Jantz, and V. Honavar. The Virtual Data Collaboratory: a regional cyberinfrastructure for collaborative data-driven research.

Computing in Science and Engineering, 22(3):79–92, May/ June 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pari:2020:MSE

[PSS20] R. Pari, M. Sandhya, and S. Sankar. A multitier stacked ensemble algorithm for improving classification accuracy. Computing in Science and Engineering, 22(4):74– 85, July/August 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pawlik:2015:CSS

[PSSP15] Aleksandra Pawlik, Judith Segal, Helen Sharp, and Marian Petre. Crowdsourcing scientific software documentation: A case study of the NumPy documentation project. Computing in Science and Engineering, 17(1):28–36, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/01/mcs2015010028-
                                        abs.html.

Parashar:2014:EDG

[PT14] Manish Parashar and George K. Thiruvathukal. Extreme data [Guest Editors’ introduction]. Computing in Science and Engineering, 16(4):8–10, July/August 2014. CODEN CSENFA. ISSN 1521-9615.

Porter:2013:IML

[PTH13] Reid Porter, James Theiler, and Don Hush. Interactive machine learning in data exploitation. Computing in Science and Engineering, 15(5): 12–20, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Papadimitriou:2009:SSJ

[PTML09] Stergios Papadimitriou, Konstantinos Terzidis, Seferina Mavroudi, and Spiridon Likothanassis. Scientific scripting for the Java platform with jLab. Computing in Science and Engineering, 11(4):50– 60, July/August 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Papadimitriou:2011:SES

[PTML11] Stergios Papadimitriou, Konstantinos Terzidis, Seferina Mavroudi, and Spiridon Likothanassis. ScalaLab: An effective Scala-based scientific programming environment for Java. Computing in Science and Engineering, 13(5):43– 55, September/October 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Putz:2016:OPN

[Put16] Andreas Putz. OpenPNM: A pore network modeling package. Computing in Science and Engineering, 18(4):

60–74, July/August 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Patrick:2000:GEI

[PV00] Merrell Patrick and Robert Voigt. Guest Editors’ introduction: Advancing simulation science and engineering at disciplinary interfaces. Computing in Science and Engineering, 2(2): 14–16, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c2014.
pdf;   http://www.computer.
org/cse/cs1999/c2014abs.
htm.

Park:2010:MDS

[PZJS10] Yumi Park, Ya Zhou, Janam Jhaveri, and Alejandro Strachan. Molecular dynamics simulations of strain engineering and thermal transport in nanostructured materials. Computing in Science and Engineering, 12(2): 36–42, March/April 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Qian:2019:HPC

[QL19] D. Qian and Z. Luan. High performance computing development in China: a brief review and perspectives. Computing in Science and Engineering, 21(1):6–16, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Qiang:2007:SRP

[QPCJ07] Wang Qiang, Zhigeng Pan, Chen Chun, and Bu Jianjun. Surface rendering for parallel slices of contours from medical imaging. Computing in Science and Engineering, 9(1):32–37, January/February 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Qin:2022:TDA

[QRP22] Yubo Qin, Ivan Rodero, and Manish Parashar. Toward democratizing access to facilities data: a framework for intelligent data discovery and delivery. Computing in Science and Engineering, 24(3): 52–60, May/June 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Quezada-Sarmiento:2020:KRM

[QSEQJFH20] P. A. Quezada-Sarmiento, L. E. Enciso-Quispe, L. A. Jumbo-Flores, and W. Hernandez. Knowledge representation model for bodies of knowledge based on design patterns and hierarchical graphs. Computing in Science and Engineering, 22(2): 55–63, March/April 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Quan:2018:IIP

[Qua18] Wei Quan. Intelligent information processing [Guest Editor’s introduction]. Computing in Science and Engineering, 20(2):8–9, March/April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8318000/.

Quan:2019:IIP

[Qua19] W. Quan. Intelligent information processing. Computing in Science and Engineering, 21 (6):4–5, November/December 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Radhakrishnan:2021:HML

[Rad21] Mala L. Radhakrishnan. How to model for a living: The CSGF as a catalyst for supermodels. Computing in Science and Engineering, 23(6): 34–41, November/December 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Raffin:2016:PVC

[Raf16] Bruno Raffin. Parallel Voronoi computation for physics-based simulations. Computing in Science and Engineering, 18 (3):88–94, May/June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Raghunathan:2006:MSD

[Rag06] Sudarshan Raghunathan. Making a supercomputer do what you want: High-level tools for parallel programming. Computing in Science and Engineering, 8(5):70–80, September/October 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Raghunathan:2007:PCA

[Rag07] Sudarshan Raghunathan. Parallel computing algorithms and applications. Computing in Science and Engineering, 9(4):64–65, July/ August 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/04/c4064.pdf.

Ramachandran:2018:APB

[Ram18] Prabhu Ramachandran. automan:

A Python-based automation framework for numerical computing. Computing in Science and Engineering, 20(5):81–97, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/05/mcs2018050081-
abs.html.

Randall:2006:RMM

[Ran06] Dana Randall. Rapidly mixing Markov chains with applications in computer science and physics. Computing in Science and Engineering, 8(2): 30–41, March/April 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Rao:2016:ECH

[Rao16] Dhananjai M. Rao. Eliciting characteristics of H5N1 in high-risk regions using phylogeography and phylodynamic simulations. Computing in Science and Engineering, 18 (4):11–24, July/August 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ram:2019:CPA

[RBC^{+}19] K. Ram, C. Boettiger, S. Chamberlain, N. Ross, M. Salmon, and S. Butland. A community of practice around peer review for long-term research software sustainability. Computing in Science and Engineering, 21(2):59–65, March/April 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Ramakrishnan:2002:SSM

[RBK02] Naren Ramakrishnan and Chris Bailey-Kellogg. Sampling strategies for mining in data-scarce domains. Computing in Science and Engineering, 4(4):31–43, July/ August 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/04/c4031abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/04/c4031.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/04/
c4031.pdf.

Ratner:2001:GEI

[RC01] Mark A. Ratner and James R.

Chelikowsky. Guest Editors’ introduction: Nanoscience, nanotechnology, and modeling. Computing in Science and Engineering, 3(4): 40–41, July/August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c4040.
pdf;   http://www.computer.
org/cse/cs1999c4040abs.
htm.

Rosner:2000:FCS

[RCD^{+}00] Robert Rosner, Alan Calder, Jonathan Dursi, Bruce Fryxell, Donald Q. Lamb, Jens C. Niemeyer, Kevin Olson, Paul Ricker, Frank X. Timmes, James W. Truran, Henry Tufo, Yuan-Nan Young, Michael Zingale, Ewing Lusk, and Rick Stevens. Flash code: Studying astrophysical thermonuclear flashes. Computing in Science and Engineering, 2(2):33–41, March/April 2000. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c2033.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c2033abs.
                                        htm.

Ravi-chandar:1999:PCF

[RcK99] Krishnaswamy Ravi-chandar and Wolfgang Knauss. Processes controlling fast fracture of brittle solids. Computing in Science and Engineering, 1(5): 24–31, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c5024.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c5024abs.
                                        htm.

Rust:2005:FFTa

[RD05a] B. Rust and D. Donnelly. The Fast Fourier Transform for experimentalists. Part III: Classical spectral analysis. Computing in Science and Engineering, 7(5):74– 78, September/October 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        5992/32219/01501744.pdf?
                                        isnumber=32219&prod=JNL&
                                        arnumber=1501744&arSt=+
                                        74&ared=+78&arAuthor=+Rust%
                                        2C+B.%3B++Donnelly%2C+D.;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=

32219&arnumber=1501744&
count=14&index=10.

Rust:2005:FFTb

[RD05b] Bert Rust and Denis Donnelly. The Fast Fourier Transform for experimentalists, Part IV: Autoregressive spectral analysis. Computing in Science and Engineering, 7(6):85–90, November/December 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Rebello:1999:WMH

[Reb99]  N. Sanjay Rebello.  Web

mechanics: How to distribute your software over the Web. Computing in Science and Engineering, 1(6): 79–81, November/December 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c6079.
pdf;   http://www.computer.
org/cse/cs1999/c6079abs.
htm.

Reed:2016:IEC

[Ree16] Dale Reed. The impact of the exploring computer science instructional model in Chicago public schools. Computing in Science and Engineering, 18 (2):10–17, March/April 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Reif:2002:DLM

[Rei02] John H. Reif. DNA lattices: a method for molecularscale patterning and computation. Computing in Science and Engineering, 4(1):32– 41, January/February 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c1032abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2002/pdf/c1032.pdf.

Reid:2003:FF

[Rei03]  John Reid.  The future

of Fortran. Computing in Science and Engineering, 5(4):59–67, July/August 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2003/04/c4059.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/04/
                                        c4059.pdf.

Reiter:2013:MSE

[Rei13] Cliff Reiter. Math-style experiments in J. Computing in Science and Engineering, 15(3): 84–88, May/June 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Reinhardt:2021:ERB

[Rei21a] William P. Reinhardt. Erratum to Relationships Between the Zeros, Weights, and Weight Functions of Orthogonal Polynomials: Derivative Rule Conjecture (the DRC) Applied to Stieltjes and Spectral Imaging. Computing in Science and Engineering, 23(4):91, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). See [Rei21b].

Reinhardt:2021:RBZ

[Rei21b] William P. Reinhardt. Relationships between the zeros, weights, and weight functions of orthogonal polynomials: Derivative rule conjecture (the DRC) applied to Stieltjes and spectral imaging. Computing in Science and Engineering, 23(3):56–64, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). See erratum [Rei21a].

Reynolds:2000:SCI

[RF00] William C. Reynolds and Massimiliano Fatica. Stanford Center for Integrated Turbulence Simulations. Computing in Science and Engineering, 2(2):54–63, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c2054.
pdf;   http://www.computer.
org/cse/cs1999/c2054abs.
htm.

Rundle:2012:CES

[RF12] John Rundle and Geoffrey Fox. Computational earthquake science. Computing in Science and Engineering, 14 (5):7–9, September/October 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Rodrigues:2013:MAA

[RGD13] A. Wendell O. Rodrigues, Frédéric Guyomarc’h, and Jean-Luc Dekeyser. An MDE approach for automatic code generation from UML/ MARTE to OpenCL. Computing in Science and Engineering, 15(1):46–55, January/February 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Rouson:2023:SIA

[RHC^{+}23] Damian Rouson, Konrad Hinsen, Jeffrey Carver, Irina Tezaur, John Shalf, Rio Yokota, and Anshu Dubey. The 2023 Society for Industrial and Applied Mathematics Conference on Computational Science and Engineering. Computing in Science and Engineering, 25(2):41– 43, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Rice:1999:PCS

[Ric99] John R. Rice. A perspective on computational science in the 21st Century. Computing in Science and Engineering, 1(2):14–16, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2014.
pdf.

Rawitscher:2005:ENS

[RK05] George H. Rawitscher and Israel Koltracht. An efficient numerical spectral method for solving the Schrödinger equation. Computing in Science and Engineering, 7(6): 58–66, November/December 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
comp/mags/cs/2005/extras/
c6058x1.htm;   http://csdl.
computer.org/comp/mags/
cs/2005/extras/c6058x1.
pdf;   http://csdl.computer.
org/comp/mags/cs/2005/extras/
c6058x2.pdf;   http://csdl.
computer.org/comp/mags/
cs/2005/extras/c6058x3.
pdf;   http://csdl.computer.
org/comp/mags/cs/2005/extras/
c6058x4.pdf.

Roanes-Lozano:2013:CEC

[RLHGA^{+}13] Eugenio Roanes-Lozano, Antonio Hernando, Alberto Garcia-Alvarez, Luis Eduardo Mesa, and Ignacio Gonzalez-Franco. Calculating the exploitation costs of trains in the Spanish railways. Computing in Science and Engineering, 15(3):89–95, May/ June 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Roanes-Lozano:2004:GAS

[RLRML04a] Eugenio Roanes-Lozano, Eugenio Roanes-Macías, and Luis M. Laita. The geometry of algebraic systems and their exact solving using Gröbner bases. Computing in Science and Engineering, 6(2): 76–79, March/April 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/02/c2076abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2004/02/c2076.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/02/
                                        c2076.pdf.

Roanes-Lozano:2004:SAG

[RLRML04b] Eugenio Roanes-Lozano, Eugenio Roanes-Macías, and Luis M. Laita. Some applications of Gröbner bases. Computing in Science and Engineering, 6(3):56–60, May/ June 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/cs/2004/03/c3056abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/03/
                                        c3056.htm;   http://csdl.
                                        computer.org/dl/mags/cs/
                                        2004/03/c3056.pdf.

Rouson:2012:IYP

[RMX12] Damian Rouson, Karla Morris, and Jim Xia. This isn’t your parents’ Fortran: Managing C++ objects with modern Fortran. Computing in Science and Engineering, 14(2): 46–54, March/April 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Roache:2004:BPC

[Roa04] Patrick J. Roache. Building PDE codes to be verifiable and validatable. Computing in Science and Engineering, 6(5): 30–38, September/October 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/05/c5030.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/05/
c5030.pdf.

Roberts:2004:VEM

[Rob04] Jonathan C. Roberts. Visualization equivalence for multisensory perception: Learning from the visual. Computing in Science and Engineering, 6(3): 61–65, May/June 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/03/c3061abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2004/03/c3061.
htm;   http://csdl.computer.

                                        org/dl/mags/cs/2004/03/
                                        c3061.pdf.

Robinson:2006:RRP

[Rob06] Sara Robinson. Recent research provides new picture of router-level Internet. Computing in Science and Engineering, 8(2): 3–6, March/April 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2006/02/c2003.pdf.

Robison:2013:CPP

[Rob13] Arch D. Robison. Composable parallel patterns with Intel Cilk Plus. Computing in Science and Engineering, 15(2): 66–71, March/April 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Rockmore:2000:FAW

[Roc00] Daniel N. Rockmore. The FFT: An algorithm the whole family can use. Computing in Science and Engineering, 2 (1):60–64, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c1060.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c1060abs.
                                        htm.

Rohrle:2010:SEM

[Roh10] Oliver Rohrle. Simulating the electro-mechanical behavior of skeletal muscles. Computing in Science and Engineering, 12(6):48–58, November/December 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ronnow:2014:ESC

[Ron14] Troels F. Ronnow. Evaluating a special class of two-body integrals through Maple’s integral routine extension. Computing in Science and Engineering, 16(1):18–23, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Roos:2006:IAC

[Roo06] Kelly R. Roos. An incremental approach to computational physics education. Computing in Science and Engineering, 8(5):44–50, September/October 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Rodero:2020:DCE

[RP20] I. Rodero and M. Parashar.

Data cyberinfrastructure for end-to-end science. Computing in Science and Engineering, 22(5):60–71, September/October 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Remmel:2012:STS

[RPBE12] Hanna Remmel, Barbara Paech, Peter Bastian, and Christian Engwer. System testing a scientific framework using a regression-test environment. Computing in Science and Engineering, 14(2): 38–45, March/April 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Remmel:2014:CSQ

[RPEB14] Hanna Remmel, Barbara Paech, Christian Engwer, and Peter Bastian. A case study on a quality assurance process for a scientific framework. Computing in Science and Engineering, 16(3):58–66, May/ June 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Rossant:2021:HPI

[RR21] Cyrille Rossant and Nicolas P. Rougier. High-performance interactive scientific visualization with Datoviz via the Vulkan low-level GPU API. Computing in Science and Engineering, 23(4):85– 90, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ribeiro:2006:SDW

[RRAB06] Leonardo C. Ribeiro, Ricardo M. Ruiz, Eduardo M. Albuquerque, and Américo T.

Bernardes. Science in the developing world: Running twice as fast? Computing in Science and Engineering, 8(4): 81–87, July/August 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Randall:2002:CMS

[RRH^{+}02] David A. Randall, Todd D.

Ringler, Ross P. Heikes, Phil Jones, and John Baumgardner. Climate modeling with spherical geodesic grids. Computing in Science and Engineering, 4(5):32–41, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/05/c5032abs.htm;
http://csdl.computer.org/
dl/mags/cs/2002/05/c5032.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/05/
c5032.pdf.

Rangasamy:2020:MOE

[RRN20] D. Rangasamy, S. Rajappan, and M. Natesan. A multiobjective evolutionary approach for preprocessing imbalanced microarray datasets. Computing in Science and Engineering, 22(1):88–100, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Rasquin:2014:SIF

[RSC^{+}14] Michel Rasquin, Cameron Smith, Kedar Chitale, E. Seegyoung Seol, Benjamin A. Matthews, Jeffrey L. Martin, Onkar Sahni, Raymond M. Loy, Mark S. Shephard, and Kenneth E. Jansen. Scalable implicit flow solver for realistic wing simulations with flow control. Computing in Science and Engineering, 16(6):13–21, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs2014060013-
abs.html.

Reinert:2020:VAD

[RSZ^{+}20] A. Reinert, L. S. Snyder, J. Zhao, A. S. Fox, D. F. Hougen, C. Nicholson, and D. S. Ebert. Visual analytics for decision-making during pandemics. Computing in Science and Engineering, 22(6): 48–59, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). See correction [RSZ^{+}21].

Reinert:2021:CVA

[RSZ^{+}21] A. Reinert, L. S. Snyder, J. Zhao, A. S. Fox, D. F. Hougen, C. Nicholson, and D. S. Ebert. Corrections to Visual Analytics for Decision-Making During Pandemics. Computing in Science and Engineering, 23(1):106, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). See [RSZ^{+}20].

Rocklin:2012:SSS

[RT12] Matthew Rocklin and Andy R.

Terrel. Symbolic statistics with SymPy. Computing in Science and Engineering, 14(3):88–93, May/June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Raddick:2014:TYSa

[RTSS14a] M. Jordan Raddick, Ani R.

Thakar, Alexander S. Szalay, and Rafael D. C. Santos. Ten years of SkyServer I: Tracking Web and SQL e-Science usage. Computing in Science and Engineering, 16(4):22–31, July/August 2014. CODEN CSENFA. ISSN 1521-9615.

Raddick:2014:TYSb

[RTSS14b] M. Jordan Raddick, Ani R.

Thakar, Alexander S. Szalay, and Rafael D. C. Santos. Ten years of SkyServer II: How astronomers and the public have embraced e-Science. Computing in Science and Engineering, 16(4):32–40, July/August 2014. CODEN CSENFA. ISSN 1521-9615.

Rundle:2000:GEI

[Run00] John B. Rundle. Guest Editor’s introduction: Computational Earth system science.

Computing in Science and Engineering, 2(3):20–21, May/ June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3020.
pdf;   http://www.computer.
org/cse/cs1999/c3020abs.
htm.

Rundle:2003:SCP

[Run03] John B. Rundle. Scaling and critical phenomena in nature. Computing in Science and Engineering, 5(5):80–82, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/05/c5080abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/05/c5080.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/05/
c5080.pdf.

Rundle:2005:GEI

[Run05] J. B. Rundle. Guest Editor’s introduction: iSERVO— The International Solid Earth Research Virtual Observatory. Computing in Science and Engineering, 7(4): 24–26, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463132.pdf?
isnumber=31456&prod=JNL&
arnumber=1463132&arSt=+
24&ared=+26&arAuthor=Rundle%

2C+J.B.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=31456&arnumber=
1463132&count=14&index=2.

Rust:2001:FNBa

[Rus01a] Bert W. Rust. Fitting nature’s basic function. Part I: Polynomials and linear least squares. Computing in Science and Engineering, 3(5):84– 89, September/October 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c5084abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c5084.pdf.

Rust:2001:FNBb

[Rus01b] Bert W. Rust. Fitting nature’s basic functions. Part II: Estimating uncertainties and testing hypotheses. Computing in Science and Engineering, 3(6):60–64, November/December 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c6060abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c6060.pdf.

Rust:2002:FNB

[Rus02] Bert W. Rust. Fitting nature’s basic functions. Part III: Exponentials, sinusoids, and nonlinear least squares. Computing in Science and Engineering, 4(4):72–77, July/ August 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/04/c4072abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/04/c4072.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/04/
                                        c4072.pdf.

Rust:2003:FNB

[Rus03] Bert W. Rust. Fitting nature’s basic functions. Part IV: The variable projection algorithm. Computing in Science and Engineering, 5(2): 74–79, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/02/c2074abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/02/c2074.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/02/
                                        c2074.pdf.

Ramachandran:2011:MVS

[RV11] Prabhu Ramachandran and Gael Varoquaux. Mayavi: 3D visualization of scientific data. Computing in Science and Engineering, 13(2): 40–51, March/April 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Rehr:2010:SCC

[RVG^{+}10] John J. Rehr, Fernando D.

Vila, Jeffrey P. Gardner, Lucas Svec, and Micah Prange. Scientific computing in the cloud. Computing in Science and Engineering, 12(3): 34–43, May/June 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sahimi:2008:NSWa

[SA08a] Muhammad Sahimi and S. Mehdi Vaez Allaei. Numerical simulation of wave propagation, Part I: Sequential computing. Computing in Science and Engineering, 10(3):66–75, May/ June 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sahimi:2008:NSWb

[SA08b] Muhammad Sahimi and S. Mehdi Vaez Allaei. Numerical simulation of wave propagation, Part II: Parallel computing. Computing in Science and Engineering, 10(4):76–83, July/August 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Saadatfar:2009:CSG

[Saa09] Mohammad Saadatfar. Computer simulation of granular materials. Computing in Science and Engineering, 11 (1):66–74, January/February 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Saad:2022:ODK

[Saa22] Yousef Saad. The origin and development of Krylov subspace methods. Computing in Science and Engineering, 24 (4):28–39, July/August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Salagaram:2015:SPP

[SAC15] Trisha Salagaram, Richard Charles Andrew, and Nithaya Chetty. Simplified pseudopotential problems for the classroom. Computing in Science and Engineering, 17(1):46–53, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/01/mcs2015010046-
                                        abs.html.

Settlemyer:2021:TTA

[SACR21] Bradley Settlemyer, George Amvrosiadis, Philip Carns, and Robert Ross. It’s time to talk about HPC storage: Perspectives on the past and future. Computing in Science and Engineering, 23(6):63–68, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sahimi:2003:LSP

[Sah03] Muhammad Sahimi. Largescale porous media and wavelet transformations. Computing in Science and Engineering, 5(4):75–87, 92– 93, July/August 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/04/c4075.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/04/
c4075.pdf.

Simmhan:2013:CBS

[SAK^{+}13] Yogesh Simmhan, Saima Aman, Alok Kumbhare, Rongyang Liu, Sam Stevens, Qunzhi Zhou, and Viktor Prasanna. Cloud-based software platform for big data analytics in smart grids. Computing in Science and Engineering, 15 (4):38–47, July/August 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Steinmacher-Burow:2000:SPD

[SB00] Burkhard D. Steinmacher-Burow. Scientific programming: Dividing the application definition from the execution. Computing in Science and Engineering, 2 (3):78–83, May/June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3078.
pdf;   http://www.computer.
org/cse/cs1999/c3078abs.
htm.

Schnetter:2015:CPS

[SBB^{+}15] Erik Schnetter, Marek Blazewicz, Steven R. Brandt, David M. Koppelman, and Frank Loffler. Chemora: A PDE-solving framework for modern highperformance computing architectures. Computing in Science and Engineering, 17(2): 53–64, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/02/mcs2015020053-
                                        abs.html.

Schlick:2000:CCS

[SBH^{+}00] Tamar Schlick, Daniel A. Beard, Jing Huang, Daniel A. Strahs, and Xiaoliang Qian. Computational challenges in simulating large DNA over long times. Computing in Science and Engineering, 2(6): 38–51, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c6038.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999c6038abs.
                                        htm.

Schulthess:2019:RGB

[SBW^{+}19] T. C. Schulthess, P. Bauer, N. Wedi, O. Fuhrer, T. Hoefler, and C. Schär. Reflecting on the goal and baseline for exascale computing: A roadmap based on weather and climate simulations. Computing in Science and Engineering, 21(1):30–41, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Sayeed:2008:MHP

[SBZ^{+}08] Mohamed Sayeed, Hansang Bae, Yili Zheng, Brian Armstrong, Rudolf Eigenmann, and Faisal Saied. Measuring high-performance computing with real applications. Computing in Science and Engineering, 10(4):60– 70, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Stonebraker:2013:SDM

[SBZB13] Michael Stonebraker, Paul Brown, Donghui Zhang, and Jacek Becla. SciDB: A database management system for applications with complex analytics. Computing in Science and Engineering, 15(3): 54–62, May/June 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Scales:2016:AAM

[Sca16]  Glenda Scales.  African-

American middle school girls: Influences on attitudes toward computer science. Computing in Science and Engineering, 18(3):14–23, May/ June 2016. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Sullivan:2018:FCE

[SCBT18] Francis Sullivan, Norman Chonacky, Isabel Beichl, and George K. Thiruvathukal. Former CiSE EICs reflect on the magazine’s 20th anniversary. Computing in Science and Engineering, 20(1):3–7, January/February 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Schreiner:1999:TCK

[Sch99] Keri Schreiner. Trends: Creativity is key to drug R&D. Computing in Science and Engineering, 1(3):6–9, May/ June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3006.
pdf;   http://www.computer.
org/cse/cs1999/c3006abs.
htm.

Schaible:2001:SSD

[Sch01]  Max Schaible.  Searching

the scientific database literature for guides to experiment and theory. Computing in Science and Engineering, 3(4):30–39, July/August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c4030.
pdf;   http://www.computer.

org/cse/cs1999c4030abs.
htm.

Schillaci:2007:CC

[Sch07] Michael Jay Schillaci. Computationally complete. Computing in Science and Engineering, 9(3):3–6, May/June 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2007/03/c3003.pdf.

Schillaci:2009:TT

[Sch09] Michael Jay Schillaci. Total tomography. Computing in Science and Engineering, 11 (2):12–13, March/April 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Schnack:2014:MCA

[Sch14]  Jurgen Schnack.  Modern

computer algebra and a lot more. Computing in Science and Engineering, 16(3): 4–5, May/June 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Schneider:2015:IHC

[Sch15] Barry I. Schneider. The impact of heterogeneous computer architectures on computational physics. Computing in Science and Engineering, 17(2):9–13, March/April 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/csdl/
mags/cs/2015/02/mcs2015020009.
html.

Schillaci:2016:MM

[Sch16] Michael Jay Schillaci. Maximum mechanics. Computing in Science and Engineering, 18(6):66–67, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs2016060066.
html.

Schillaci:2017:PP

[Sch17a] Michael Jay Schillaci. Perfectly Python. Computing in Science and Engineering, 19(6):51–53, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060051.
html.

Schneider:2017:LBY

[Sch17b] Barry I. Schneider. Looking back at 45 years of computational physics and chemistry. Computing in Science and Engineering, 19(5):4–5, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/05/mcs2017050004.
html.

Schuster:2018:HEH

[Sch18] Micah D. Schuster. The heat equation: High-performance scientific computing case study. Computing in Science and Engineering, 20(5): 114–127, September/October 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://www.computer.org/
csdl/mags/cs/2018/05/mcs2018050114-
abs.html.

Schuster:2020:LGC

[Sch20] M. D. Schuster. Lattice gas cellular automata fluid dynamics case study. Computing in Science and Engineering, 22(6):87–91, November/December 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Schuster:2021:IMC

[Sch21] M. D. Schuster. Infection modeling case study: Discrete spatial susceptible-infected-recovered model. Computing in Science and Engineering, 23 (1):83–88, January/February 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Schonborn:2023:ASE

[Sch23a] Marc Thomas Schönborn.

Adopting software engineering concepts in scientific research: Insights from physicists and mathematicians turned consultants. Computing in Science and Engineering, 25(4): 25–33, April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Schuster:2023:DWS

[Sch23b] Micah D. Schuster. Discrete wildfire simulation case study. Computing in Science and Engineering, 25(4):47–52, April 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Schuster:2023:LGC

[Sch23c] Micah D. Schuster. Lattice gas cellular automata fluid dynamics: The model of Frisch, Hasslacher, and Pomeau. Computing in Science and Engineering, 25(2): 64–68, March/April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Shen:2017:PIE

[SCW^{+}17] Bo-Wen Shen, Samson Cheung, Yu-Ling Wu, Jui-Lin F. Li, and David Kao. Parallel implementation of the ensemble empirical mode decomposition and its application for Earth science data analysis. Computing in Science and Engineering, 19(5):49–57, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/

cs/2017/05/mcs2017050049-
abs.html.

Schulte:2011:ADO

[SD11] Eric Schulte and Dan Davison. Active documents with Org-Mode. Computing in Science and Engineering, 13(3): 66–73, May/June 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Schleife:2014:QDS

[SDA^{+}14] Andre Schleife, Erik W.

Draeger, Victor M. Anisimov, Alfredo A. Correa, and Yosuke Kanai. Quantum dynamics simulation of electrons in materials on highperformance computers. Computing in Science and Engineering, 16(5):54–60, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs2014050054-
abs.html.

Santo:2020:NEP

[SDA20] R. Santo, L. A. DeLyser, and J. Ahn. Negotiating equity priorities within systems change: a case study of a district-level initiative to implement K12 computer science education. Computing in Science and Engineering, 22 (5):7–19, September/October 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Scofield:2010:XDF

[SDCV10] Todd C. Scofield, Jeffrey A. Delmerico, Vipin Chaudhary, and Geno Valente. XtremeData dbX: An FPGAbased data warehouse appliance. Computing in Science and Engineering, 12(4):66– 73, July/August 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sendlinger:2008:TCE

[SDD^{+}08] Shawn C. Sendlinger, Don J. DeCoste, Thom H. Dunning, Diana Avalos Dummitt, Eric Jakobsson, Dave R. Mattson, and Edee Norman Wiziecki. Transforming chemistry education through computational science. Computing in Science and Engineering, 10(5):34– 39, September/October 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Smit:2021:MPE

[SdPV21] Willem A. Smit, Johan A. du Preez, and Guy A. E. Vandenbosch. Mobile processor energy usage in the scientific environment. Computing in Science and Engineering, 23(3):65–72, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Scherer:2000:SPV

[SDS00] David Scherer, Paul Dubois, and Bruce Sherwood. Scientific programming: VPython: 3D interactive scientific graphics for students. Computing in Science and Engineering, 2(5): 56–62, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c5056.
pdf.

Segee:1999:BWR

[Seg99]  Bruce Segee.  Book &

Web reviews: Methods in Neuronal Modeling: From Ions to Networks, 2nd edition. Computing in Science and Engineering, 1(1): 81, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c1081.
pdf.

Silva:2010:UV

[SEPC10] Claudio Silva, David Ebert, Hanspeter Pfister, and Sheelagh Carpendale. An update from VisWeek 2009. Computing in Science and Engineering, 12(1):82–87, January/February 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Showerman:2011:EPE

[SES^{+}11] Mike Showerman, Jeremy Enos, Craig Steffen, Sean Treichler, William Gropp, and Wen mei W. Hwu. EcoG: a power-efficient GPU cluster architecture for scientific computing. Computing in Science and Engineering, 13(2): 83–87, March/April 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

SanMiguel:2005:BMS

[SETK05] Maxi San Miguel, Victor M. Eguíluz, Raul Toral, and Konstantin Klemm. Binary and multivariate stochastic models of consensus formation. Computing in Science and Engineering, 7(6):67–73, November/December 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Stromback:2011:XMB

[SF11] Lena Stromback and Juliana Freire. XML management for bioinformatics applications. Computing in Science and Engineering, 13(5):12– 23, September/October 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Silva:2007:PVR

[SFC07] Claudio T. Silva, Juliana Freire, and Steven P. Callahan. Provenance for visualizations: Reproducibility and beyond. Computing in Science and Engineering, 9(5): 82–89, September/October 2007. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Succi:2001:ALB

[SFSK01] Sauro Succi, Olga Filippova, Greg Smith, and Efthimios Kaxiras. Applying the lattice Boltzmann equation to multiscale fluid problems. Computing in Science and Engineering, 3(6):26–37, November/December 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c6026abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c6026.pdf.

Siantar:2000:PBR

[SG00] Christine L. Hartmann Siantar and Dewey Garrett. Photonbeam radiation therapy and Monte Carlo. Computing in Science and Engineering, 2(5): 31–38, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c5031.
pdf;   http://www.computer.
org/cse/cs1999/c5031abs.
htm.

Steffen:2010:NSF

[SG10] Craig Steffen and Gildas Genest. Nallatech in-socket FPGA front-side bus accelerator. Computing in Science and Engineering, 12(2): 78–83, March/April 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Schweiger:2003:CAD

[SGA03] Martin Schweiger, Adam Gibson, and Simon R. Arridge. Computational aspects of diffuse optical tomography. Computing in Science and Engineering, 5(6):33–41, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/06/c6033abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/06/c6033.
                                        pdf.

Stone:2022:ADR

[SGA^{+}22] John E. Stone, Kevin S. Griffin, Jefferson Amstutz, David E. DeMarle, William R. Sherman, and Johannes Günther. ANARI: a 3-D rendering API standard. Computing in Science and Engineering, 24(2): 7–18, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Stromer-Galley:2018:UCD

[SGRK^{+}18] Jennifer Stromer-Galley, Patricia G. C. Rossini, Kate Kenski, James Folkestad, Brian McKernan, Rosa Mikeal Martey, Benjamin Clegg, Carsten Osterlund, and Lael Schooler. User-centered design and experimentation to develop effective software for evidence-based reasoning in the intelligence community: The TRACE Project. Computing in Science and Engineering, 20(6):35–42, November/December 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/06/08532281-abs.
html.

Stone:2010:OPP

[SGS10] John E. Stone, David Gohara, and Guochun Shi. OpenCL: a parallel programming standard for heterogeneous computing systems. Computing in Science and Engineering, 12(3):66–73, May/June 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Shires:2002:DCV

[SGW02] Dale R. Shires, William Green, and Shawn Walsh. Designing, controlling, and visualizing composite material manufacturing. Computing in Science and Engineering, 4(4): 86–91, July/August 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/04/c4086abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/04/c4086.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/04/
c4086.pdf.

Sahimi:2010:ECS

[SH10] Muhammad Sahimi and Hossein Hamzehpour. Efficient computational strategies for solving global optimization problems. Computing in Science and Engineering, 12(4): 74–83, July/August 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Shaw:1999:IIT

[Sha99] David Shaw. Interview: At the intersection of technology and finance. Computing in Science and Engineering, 1(6): 24–25, November/December 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c6024.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c6024abs.
                                        htm.

Shallit:2014:RED

[Sha14] Jeffrey Shallit. Reviewing the early days of computing [book review]. Computing in Science and Engineering, 16(6):8–9, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2014/06/mcs2014060008.
                                        html.

Shen:2007:ISS

[She07] Zuo-Jun (Max) Shen. Integrated stochastic supply-chain design models. Computing in Science and Engineering, 9(2): 50–59, March/April 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Shirer:1999:TNR

[Shi99] Donald L. Shirer. Technology news & reviews: HiQ gets a high five for mathematical analysis. Computing in Science and Engineering, 1(2):4–5, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2004.
pdf.

Shirer:2000:TNRd

[Shi00a] Donald L. Shirer. Technology news & reviews: An olio of almost obscure programmers’ products. Computing in Science and Engineering, 2(6): 6–??, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c6006.
pdf;   pdf/c6006.pdf.

Shirer:2000:TNRb

[Shi00b] Donald L. Shirer. Technology news & reviews: Basic: The little language that wouldn’t die. Computing in Science and Engineering, 2 (2):6–10, March/April 2000.

CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c2006.
pdf;   http://www.computer.
org/cse/cs1999/c2006abs.
htm.

Shirer:2000:TNRa

[Shi00c] Donald L. Shirer. Technology news & reviews: Multimedia cT language is now free. Computing in Science and Engineering, 2(1): 3–5, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1003.
pdf;   http://www.computer.
org/cse/cs1999/c1003abs.
htm.

Shirer:2000:TNRc

[Shi00d] Donald L. Shirer. Technology news & reviews: Products to simplify your life. Computing in Science and Engineering, 2(3):10–11, May/ June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3010.
pdf;   http://www.computer.
org/cse/cs1999/c3010abs.
htm.

Shirer:2001:TNR

[Shi01a] Donald L. Shirer. Technology news & reviews: Lab-View 6i adds Internet features to data acquisition environment. Computing in Science and Engineering, 3 (4):8–11, July/August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c4008.
pdf.

Shirer:2001:VLC

[Shi01b] Donald L. Shirer. Versatile laboratory computer interfaces. Computing in Science and Engineering, 3(6):9–13, November/December 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c6009abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c6009.pdf.

Shirer:2002:MSW

[Shi02a] Donald L. Shirer. Mathematica spins Web wizardry. Computing in Science and Engineering, 4 (2):9–10, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c2009abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c2009.pdf.

Shirer:2002:POS

[Shi02b] Donald L. Shirer. Planetarium and observer software. Computing in Science and Engineering, 4(5):14–23, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/05/c5014abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/05/c5014.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/05/
                                        c5014.pdf.

Shirer:2003:LOA

[Shi03]  Don Shirer.  Out on a

LIMS (and other arboreal features). Computing in Science and Engineering, 5 (1):14–16, January/February 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2003/01/c1014.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/01/
                                        c1014.pdf.

Shi:2007:PIG

[Shi07] Xuan Shi. Python for Internet GIS applications. Computing in Science and Engineering, 9(3):56–59, May/ June 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Shneiderman:2006:THD

[Shn06] Ben Shneiderman. A telescope for high-dimensional data. Computing in Science and Engineering, 8(2): 48–53, March/April 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sletholt:2012:WDW

[SHPL12] Magnus Thorstein Sletholt, Jo Erskine Hannay, Dietmar Pfahl, and Hans Petter Langtangen. What do we know about scientific software development’s agile practices? Computing in Science and Engineering, 14(2): 24–37, March/April 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sich:2009:IHP

[Sic09] Mark Sich. Interactive holography: Pursuit of a dream. Computing in Science and Engineering, 11(1):62–65, January/February 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Silbar:2000:WMA

[Sil00] Richard R. Silbar. Web mechanics: Animating equations on the Web. Computing in Science and Engineering, 2(4):91–95, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c4091.
pdf.

Silbar:2002:WDI

[Sil02]  Richard R. Silbar.  Web

delivery of interactive laboratories: Comparing three authoring tools. Computing in Science and Engineering, 4(5):74–78, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/05/c5074abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/05/c5074.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/05/
c5074.pdf.

Simpson:2013:NCC

[Sim13] David G. Simpson. NASA computational case study: The flight of Friendship 7. Computing in Science and Engineering, 15(2):72– 78, March/April 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Singh:2018:GST

[Sin18] Parampreet Singh. Glimpses of space-time beyond the singularities using supercomputers. Computing in Science and Engineering, 20(4):26– 38, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040026.

html. See erratum [Ano18j].

Stantchev:2009:UGP

[SJDV09] George Stantchev, Derek Juba, William Dorland, and Amitabh Varshney. Using graphics processors for highperformance computation and visualization of plasma turbulence. Computing in Science and Engineering, 11(2): 52–59, March/April 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Stubbs:2023:SSI

[SJF^{+}23] Joe Stubbs, Anagha Jamthe, Nathan Freeman, Mike Packard, Gilbert Curbelo, and Cody Hammock. Scalable scientific interactive research computing with Project Scinco. Computing in Science and Engineering, 25(1):43–50, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Skolnick:2001:CSP

[SK01] Jeffrey Skolnick and Andrzej Kolinski. Computational studies of protein folding. Computing in Science and Engineering, 3(5):40– 50, September/October 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c5040abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c5040.pdf.

Stainforth:2002:DCP

[SKA^{+}02] Dave Stainforth, Jamie Kettleborough, Myles Allen, Mat Collins, Andy Heaps, and James Murphy. Distributed computing for publicinterest climate modeling research. Computing in Science and Engineering, 4(3): 82–89, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/03/c3082abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/03/c3082.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/03/
                                        c3082.pdf.

Schwab:2000:MSC

[SKC00] Matthias Schwab, Martin Karrenbach, and Jon Claerbout. Making scientific computations reproducible. Computing in Science and Engineering, 2(6):61–67, November/December 2000. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c6061.
                                        pdf.

Schraml:2002:HPC

[SKC02] Stephen J. Schraml, Kent D. Kimsey, and Jerry A. Clarke. High-performance computing applications for survivabilitylethality technologies. Computing in Science and Engineering, 4(2):16–21, March/ April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://computer.org/cise/
cs2001/c2016abs.htm;   http:
//dlib.computer.org/cs/
books/cs2002/pdf/c2016.
pdf.

Symeonidis:2005:SAM

[SKC05] Vasileios Symeonidis, George Em Karniadakis, and Bruce Caswell. A seamless approach to multiscale complex fluid simulation. Computing in Science and Engineering, 7(3): 39–46, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/03/c3039abs.htm;
http://csdl.computer.org/
dl/mags/cs/2005/03/c3039.
pdf.

Strachan:2010:CES

[SKL10] Alejandro Strachan, Gerhard Klimeck, and Mark Lundstrom. Cyber-enabled simulations in nanoscale science and engineering. Computing in Science and Engineering, 12 (2):12–17, March/April 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sharma:2003:LMD

[SKNV03] Ashish Sharma, Rajiv K.

Kalia, Aiichiro Nakano, and Priya Vashishta. Large multidimensional data visualization for materials science. Computing in Science and Engineering, 5(2): 26–33, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/02/c2026abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/02/c2026.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/02/
c2026.pdf.

Shi:2010:AAC

[SKP^{+}10] Guochun Shi, Volodymyr Kindratenko, Frederico Pratas, Pedro Trancoso, and Michael Gschwind. Application acceleration with the Cell Broadband Engine. Computing in Science and Engineering, 12 (1):76–81, January/February 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Skupin:2004:PTW

[Sku04] André Skupin. A picture from a thousand words. Computing in Science and Engineering, 6(5):84–88, September/October 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/05/c5084.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/05/
c5084.pdf.

Siek:1999:SPM

[SL99] Jeremy G. Siek and Andrew Lumsdaine. Scientific programming: The matrix template library: Generic components for high-performance scientific computing. Computing in Science and Engineering, 1(6):70–78, November/December 1999. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c6070.
pdf;   http://www.computer.
org/cse/cs1999/c6070abs.
htm.

Shedden:2003:DRS

[SL03] Kerby Shedden and Ker-Chau Li. Dimension reduction and spatiotemporal regression: Applications to neuroimaging. Computing in Science and Engineering, 5(5):30–36, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/05/c5030abs.htm;
http://csdl.computer.org/
dl/mags/cs/2003/05/c5030.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/05/
c5030.pdf.

Sapp:2018:DLD

[SL18] Wendi K. Sapp and Mary Ann Leung. A different lens on diversity and inclusion: Creating research opportunities for small liberal arts colleges. Computing in Science and Engineering, 20(4):90– 94, July/August 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040090-
abs.html.

Shaker:2020:DVR

[SLK^{+}20] A. Shaker, X. Lin, D. Y. Kim, J. Kim, G. Sharma, and M. A. Devine. Design of a virtual reality tour system for people with intellectual and developmental disabilities: a case study. Computing in Science and Engineering, 22(3): 7–17, May/June 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Stodden:2012:RRS

[SLM12] Victoria Stodden, Randall LeVeque, and Ian Mitchell. Reproducible research for scientific computing: Tools and strategies for changing the culture. Computing in Science and Engineering, 14(4): 13–17, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sloan:2016:JSB

[Slo16] Kay Sloan. Julian scholars: Broadening participation of low-income, firstgeneration computer science majors. Computing in Science and Engineering, 18(3): 32–43, May/June 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Schuster:2017:NCC

[SM17] Micah D. Schuster and Nargess Memarsadeghi. NASA computational case study: Spectral energy distribution fitting. Computing in Science and Engineering, 19(6):54–63, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060054-
abs.html.

Small:2012:SPB

[Sma12] Alex Small. Scientific Python for both expert and novice programmers. Computing in Science and Engineering, 14 (2):6–7, March/April 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Srivastava:2001:CNC

[SMC01] Deepak Srivastava, Madhu Menon, and Kyeongjae Cho. Computational nanotechnology with carbon nanotubes and fullerenes. Computing in Science and Engineering, 3 (4):42–55, July/August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2001/pdf/c4042.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999c4042abs.
                                        htm.

Schoedl:2023:PMA

[SMF^{+}23] Nathan W. Schoedl, Emma J. MacKie, Michael J. Field, Eric A. Stubbs, Allan Zhang, Matthew Hibbs, and Mathieu Gravey. A Python multiprocessing approach for fast geostatistical simulations of subglacial topography. Computing in Science and Engineering, 25(3):42–49, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Smith:1999:IHY

[Smi99a] Norris Parker Smith. Interfaces: How the Y1K problem may have been resolved in the year 999. Computing in Science and Engineering, 1(2):96–97, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c2096.
                                        pdf.

Smith:1999:IME

[Smi99b] Norris Parker Smith. Interfaces: Marketing Eau De Skunqoil: Only on the Internet. Computing in Science and Engineering, 1 (3):88–89, May/June 1999.

CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3088.
pdf;   http://www.computer.
org/cse/cs1999/c3088abs.
htm.

Smith:1999:ITL

[Smi99c] Norris Parker Smith. Interfaces: Terms of life defined in agreement with fate.com — a fantasy. Computing in Science and Engineering, 1(5): 104–105, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5104.
pdf;   http://www.computer.
org/cse/cs1999/c5104abs.
htm.

Smith:1999:IEA

[Smi99d] Norris Parker Smith. Interfaces: The Eureka anomaly. Computing in Science and Engineering, 1(6):96–96, 95–95, November/December 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c6096.
pdf;   http://www.computer.
org/cse/cs1999/c6096abs.
htm.

Smith:1999:IUM

[Smi99e] Norris Parker Smith. Interfaces: The ultimate merger: Microsoft and the Postal Service. Computing in Science and Engineering, 1(1):96– 97, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c1096.
                                        pdf.

Smith:1999:IYP

[Smi99f] Norris Parker Smith. Interfaces: Y3K prevails: Is the great database deleted forever? Computing in Science and Engineering, 1 (4):88–89, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c4088.
                                        pdf.

Smith:2000:IJI

[Smi00a] Norris Parker Smith. Interfaces: Janus: The ideal patron for computing. Computing in Science and Engineering, 2(6): 104–??, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c6104.
                                        pdf.

Smith:2000:IMM

[Smi00b] Norris Parker Smith. Interfaces: Mr. & Mrs. Average America. Computing in Science and Engineering, 2 (1):96–97, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1096.
pdf;   http://www.computer.
org/cse/cs1999/c1096abs.
htm.

Smith:2000:ILW

[Smi00c] Norris Parker Smith. Interfaces: The long way from 2.5 gen to the third generation. Computing in Science and Engineering, 2(5): 88–89, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c5088.
pdf.

Smith:2000:ITU

[Smi00d] Norris Parker Smith. Interfaces: The totally useless things dot.com: Is it the ultimate future for the Web? Computing in Science and Engineering, 2(3): 97–97, 87–87, May/June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c30c3.
pdf;   http://www.computer.
org/cse/cs1999/c30c3abs.
htm.

Smith:2001:IHB

[Smi01a] Norris Parker Smith. Interfaces: How bubbles pop… and could pop once more. Computing in Science and Engineering, 3(4):96–97, July/August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2001/pdf/c4096.
                                        pdf.

Smith:2001:ICP

[Smi01b] Norris Parker Smith. Interfaces: The crucial 0.1 percent difference. Computing in Science and Engineering, 3(3):104–??, May/June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2001/pdf/c3104.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999c3104abs.
                                        htm.

Smith:2003:UMP

[Smi03]  David M. Smith.  Us-

ing multiple-precision arithmetic. Computing in Science and Engineering, 5(4): 88–93, July/August 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2003/04/c4088.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/04/
                                        c4088.pdf.

Smith:2016:EAS

[Smi16]  Janet M. Smith.  Enact-

ing agency: The strategies of women of color in computing. Computing in Science and Engineering, 18(3):58–68, May/ June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Salvadeo:2011:IFR

[SMM^{+}11] Denis Salvadeo, Nelson Mascarenhas, Jander Moreira, Alexandre Levada, and Debora C. Correa. Improving face recognition performance using RBPCA MaxLike and Information Fusion. Computing in Science and Engineering, 13(3):14–21, May/ June 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Stodden:2015:ROC

[SMS15] Victoria Stodden, Sheila Miguez, and Jennifer Seiler. ResearchCompendia.org: Cyberinfrastructure for reproducibility and collaboration in computational science. Computing in Science and Engineering, 17(1):12–19, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/01/mcs2015010012-
abs.html.

Stack:2022:OAA

[SMSC22] Matt Stack, Paul Macklin, Robert Searles, and Sunita Chandrasekaran. OpenACC acceleration of an agent-based biological simulation framework. Computing in Science and Engineering, 24(5):53– 63, September/October 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sauer:2016:ATD

[SNCM16] Franz Sauer, Tyson Neuroth, Jacqueline Chu, and Kwan-Liu Ma. Audiencetargeted design considerations for effective scientific storytelling. Computing in Science and Engineering, 18(6):68–76, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2016/06/mcs2016060068-
                                        abs.html.

Shen:2013:INM

[SNCT13] Bo-Wen Shen, Bron Nelson, Samson Cheung, and Wei-Kuo Tao. Improving NASA’s multiscale modeling framework for tropical cyclone climate study. Computing in Science and Engineering, 15(5):56– 67, September/October 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Shen:2013:AVS

[SNTL13] Bo-Wen Shen, Bron Nelson, Wei-Kuo Tao, and Yuh-Lang Lin. Advanced visualizations of scale interactions of tropical cyclone formation and tropical waves. Computing in Science and Engineering, 15(2): 47–59, March/April 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Snyder:2013:BRB

[Sny13] Wesley Snyder. Book review:

Computer Vision with Theme-Based Learning, by Simon J. D. Prince. Computing in Science and Engineering, 15(3): 6–7, May/June 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Simmerman:2013:ESM

[SOH13] Scott Simmerman, James Osborne, and Jian Huang. Eden: Simplified management of atypical high-performance computing jobs. Computing in Science and Engineering, 15(6):46–54, November/December 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sorensen:2019:JFP

[Sor19] B. Sorensen. Japan’s flagship 2020 Post-K system. Computing in Science and Engineering, 21(1):48–49, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Stathopoulos:2000:PMT

[SÖS^{+}00] Andreas Stathopoulos, Serdar Ögüt, Yousef Saad, James Chelikowsky, and Hanchul Kim. Parallel methods and tools for predicting material properties. Computing in Science and Engineering, 2 (4):19–33, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c4019.
pdf;   http://www.computer.
org/cse/cs1999/c4019abs.
htm.

Salinet:2013:VIA

[SOV^{+}13] J. L. Salinet, G. N. Oliveira, F. J. Vanheusden, J. L. D. Comba, G. A. Ng, and F. S. Schlindwein. Visualizing intracardiac atrial fibrillation electrograms using spectral analysis. Computing in Science and Engineering, 15(2): 79–87, March/April 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Som:2018:HON

[SP18] Sibendu Som and Yuanjiang Pei. HPC opens a new frontier in fuel-engine research. Computing in Science and Engineering, 20(5):77–80, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/

cs/2018/05/mcs2018050077.
html.

Su:2020:VAR

[SPB^{+}20] S. Su, V. Perry, L. Bravo, S. Kase, H. Roy, K. Cox, and V. R. Dasari. Virtual and augmented reality applications to support data analysis and assessment of science and engineering. Computing in Science and Engineering, 22(3): 27–39, May/June 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Smith:2014:USE

[SPJ^{+}14] Harrison B. Smith, Amy Pielow, Adithya Jayakumar, Matteo Muratori, B. J. Yurkovich, Ramteen Sioshansi, Ashok Krishnamurthy, Giorgio Rizzoni, and Matthew C. Roberts. User-steered energy generation and consumption multimodel simulation for pricing and policy development. Computing in Science and Engineering, 16(2): 22–33, March/April 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Segal:2015:FI

[SPS15] Judith Segal, Marian Petre, and Helen Sharp. The future of e-infrastructures. Computing in Science and Engineering, 17(3):78–84, May/ June 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/csdl/
mags/cs/2015/03/mcs2015030078.
html.

Santos:2013:UCA

[SPW^{+}13] Emanuele Santos, Jorge Poco, Yaxing Wei, Shishi Liu, Bob Cook, Dean N. Williams, and Claudio T. Silva. UV-CDAT: Analyzing climate datasets from a user’s perspective. Computing in Science and Engineering, 15(1):94–103, January/February 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sohrabi:2012:HSP

[SR12] Mina Sohrabi and Esmat Rashedi. A high-speed, performance-optimization algorithm based on a gravitational approach. Computing in Science and Engineering, 14(5):56–62, September/October 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Srinivasan:2013:RCT

[SR13] Srikant Srinivasan and Krishna Rajan. Revisiting computational thermodynamics through machine learning of high-dimensional data. Computing in Science and Engineering, 15(5):22–31, September/October 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Schulz:2022:AHQ

[SRKS22] Martin Schulz, Martin Ruefenacht, Dieter Kranzlmüller, and Laura Brandon Schulz. Accelerating HPC with quantum computing: It is a software challenge too. Computing in Science and Engineering, 24(4):60–64, July/August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sharp:2007:VRP

[SRM^{+}07] Richard P. Sharp, Randall Ridgway, Kishore Mosaliganti, Pamela Wenzel, Tony Pan, Alain de Bruin, Raghu Machiraju, Kun Huang, Gustavo Leone, and Joel H. Saltz. Volume rendering phenotype differences in mouse placenta microscopy data. Computing in Science and Engineering, 9(1):38–47, January/February 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Spotz:2002:GEI

[SS02] William F. Spotz and Paul Swarztrauber. Guest Editors’ introduction: Climate modeling. Computing in Science and Engineering, 4(5):24–25, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/05/c5024abs.htm;

                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/05/c5024.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/05/
                                        c5024.pdf.

Schulze:2006:RDC

[SS06] Christian Schulze and Dietrich Stauffer. Recent developments in computer simulations of language competition. Computing in Science and Engineering, 8(3):60–67, May/ June 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sterling:2009:HPC

[SS09] Thomas Sterling and Dylan Stark. A high-performance computing forecast: Partly cloudy. Computing in Science and Engineering, 11(4): 42–49, July/August 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sepehrinia:2011:NSA

[SS11] Reza Sepehrinia and Ameneh Sheikhan. Numerical simulation of Anderson localization. Computing in Science and Engineering, 13(3):74–83, May/ June 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sapuan:2018:GPC

[SSC18] Fazli Sapuan, Matthew Saw, and Eugene Cheah. Generalpurpose computation on GPUs in the browser using gpu.js.

Computing in Science and Engineering, 20(1):33–42, January/February 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://ieeexplore.ieee.
org/document/8254318/.

Sanchez:2011:VBD

[SSCN11] Angel Sanchez, Pedro D.

Suarez, Aura Conci, and Eldman Nunes. Video-based distance traffic analysis: Application to vehicle tracking and counting. Computing in Science and Engineering, 13(3): 38–45, May/June 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sareen:2016:CBS

[SSG16] Sanjay Sareen, Sandeep K.

Sood, and Sunil Kumar Gupta. A cloud-based seizure alert system for epileptic patients that uses higher-order statistics. Computing in Science and Engineering, 18(5): 56–67, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Suzuki:2002:AHA

[SSK02] Reiko Suzuki, Hiroshi Sato, and Mineo Kimura. Antiprotonhydrogen atom collision at intermediate energy. Computing in Science and Engineering, 4(6):24–33, November/December 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/06/c6024abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/06/c6024.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/06/
                                        c6024.pdf.

Shephard:2013:BHE

[SSK13] Mark S. Shephard, Cameron Smith, and John E. Kolb. Bringing HPC to engineering innovation. Computing in Science and Engineering, 15 (1):16–25, January/February 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Schmiedekamp:2006:CCS

[SSP06] Mendel Schmiedekamp, Aparna Subbu, and Shashi Phoha. The clustered causal state algorithm: Efficient pattern discovery for lossy data-compression applications. Computing in Science and Engineering, 8(5):59– 67, September/October 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Scheidegger:2011:RV

[SSW11] Carlos Scheidegger, Claudio T. Silva, and Daniel Weiskopf. A report from VisWeek 2010. Computing in Science and Engineering, 13 (2):70–77, March/April 2011.

CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Smith:2021:RBA

[SSW21] S. Smith, M. Sayari Nejad, and A. Wassyng. Raising the bar: Assurance cases for scientific software. Computing in Science and Engineering, 23 (1):47–57, January/February 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sullivan:1999:CPW

[ST99] Francis Sullivan and William J.

Thompson. Computing prescriptions: By way of introduction. Computing in Science and Engineering, 1 (1):82–83, January/February 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c1082.
pdf.

Saito:2005:ASW

[ST05] Tsutomu Saito and Kazuyoshi Takayama. Applying shockwave research to volcanology. Computing in Science and Engineering, 7 (1):30–35, January/February 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2005/01/c1030.
htm;   http://csdl.computer.

                                        org/dl/mags/cs/2005/01/
                                        c1030.pdf.

Silva:2008:CP

[ST08] Cláudio T. Silva and Joel E. Tohline. Computational provenance. Computing in Science and Engineering, 10 (3):9–10, May/June 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:/

                                        /csdl.computer.org/comp/
                                        mags/cs/2008/03/mcs2008030009.
                                        pdf.

Stauffer:2003:SS

[Sta03] Dietrich Stauffer. Sociophysics simulations. Computing in Science and Engineering, 5(3): 71–75, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/03/c3071abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/03/c3071.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/03/
                                        c3071.pdf.

Stein:2003:CMM

[STB03] Keith Stein, Tayfun Tezduyar, and Richard Benney. Computational methods for modeling parachute systems. Computing in Science and Engineering, 5(1):39–46, January/February 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.

computer.org/comp/mags/
cs/2003/01/c1039abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/01/c1039.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/01/
c1039.pdf.

Stevenson:1999:CLQ

[Ste99] D. E. Stevenson. A critical look at quality in largescale simulations. Computing in Science and Engineering, 1(3):53–63, May/June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3053.
pdf;   http://www.computer.
org/cse/cs1999/c3053abs.
htm.

Stewart:2000:DAM

[Ste00] G. W. Stewart. The decompositional approach to matrix computation. Computing in Science and Engineering, 2 (1):50–59, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1050.
pdf;   http://www.computer.
org/cse/cs1999/c1050abs.
htm.

Stevenson:2002:MME

[Ste02]  D. E. Stevenson.  The

Michelson–Morley experiment: a case study in validation. Computing in Science and Engineering, 4(6):40–51, November/December 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/06/c6040abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/06/c6040.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/06/
                                        c6040.pdf.

Stewart:2012:CTT

[Ste12] Derek Stewart. A cautionary tale of two basis sets and graphene. Computing in Science and Engineering, 14(2): 55–59, March/April 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Stevenson:2014:WCW

[Ste14] D. E. Stevenson. What can we learn from seventh graders? Computing in Science and Engineering, 16(1):66–68, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Szalay:2008:SDL

[STG08] Alex Szalay, Ani R. Thakar, and Jim Gray. The sqlLoader data-loading pipeline. Computing in Science and Engineering, 10(1):38–48, January/February 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Shen:2011:CAM

[STG11] Bo-Wen Shen, Wei-Kuo Tao, and Bryan Green. Coupling advanced modeling and visualization to improve highimpact tropical weather prediction. Computing in Science and Engineering, 13(5):56– 67, September/October 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Smith:2022:PAM

[STH22] Matthew Smith, Arjen Tamerus, and Phil Hasnip. Portable acceleration of materials modeling software: CASTEP, GPUs, and OpenACC. Computing in Science and Engineering, 24(1):46–55, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sachs:2012:FER

[STHR12] Michael Sachs, Donald L.

Turcotte, James R. Holliday, and John Rundle. Forecasting earthquakes: The RELM test. Computing in Science and Engineering, 14(5):43– 48, September/October 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Stein:1999:AGD

[STM99] Lincoln D. Stein and Jean Thierry-Mieg. AceDB: a genome database management system. Computing in Science and Engineering, 1 (3):44–52, May/June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3044.
pdf;   http://www.computer.
org/cse/cs1999/c3044abs.
htm.

Stodden:2009:LFR

[Sto09] Victoria Stodden. The legal framework for reproducible scientific research: Licensing and copyright. Computing in Science and Engineering, 11 (1):35–40, January/February 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Stodden:2012:RRT

[Sto12] Victoria Stodden. Reproducible research: Tools and strategies for scientific computing. Computing in Science and Engineering, 14(4): 11–12, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Strawn:2010:HPC

[Str10] Roger Strawn. High-performance computing for rotorcraft modeling and simulation. Computing in Science and Engineering, 12(5):27–35, September/October 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Succi:2005:BEC

[STTV05] Sauro Succi, Federico Toschi, Mario P. Tosi, and Patrizia Vignolo. Bose–Einstein condensates and the numerical solution of the Gross–Pitaevskii equation. Computing in Science and Engineering, 7(6): 48–57, November/December 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sempolinski:2015:ACS

[STWK15] Peter Sempolinski, Douglas Thain, Zhigang Wei, and Ahsan Kareem. Adapting collaborative software development techniques to structural engineering. Computing in Science and Engineering, 17(5):27–34, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs2015050027-
abs.html.

Sanz:2022:ABM

[SU22] Victorino Sanz and Alfonso Urquia. Agent-based modeling of traffic systems using Modelica. Computing in Science and Engineering, 24(6): 38–43, November/December 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sullivan:1999:ERC

[Sul99a] Francis Sullivan. From the editor: Recycling center. Computing in Science and Engineering, 1(2):1, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2001.
pdf.

Sullivan:1999:ECN

[Sul99b] Francis Sullivan. From the editors: Can a numerical computation prove anything? Computing in Science and Engineering, 1(5):2– 3, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5002.
pdf;   http://www.computer.
org/cse/cs1999/c5002abs.
htm.

Sullivan:2000:ESC

[Sul00a] Francis Sullivan. From the editors: Simplicity and complexity. Computing in Science and Engineering, 2(5):4– 5, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c5004.
pdf.

Sullivan:2000:ENT

[Sul00b] Francis Sullivan. From the editors: So now they tell us! Computing in Science and Engineering, 2(3):2–3, May/ June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3002.
pdf;   http://www.computer.
org/cse/cs1999/c3002abs.
htm.

Sullivan:2000:EJA

[Sul00c] Francis Sullivan. From the editors: The joy of algorithms. Computing in Science and Engineering, 2(1): 2, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1002.
pdf;   http://www.computer.
org/cse/cs1999/c1002abs.
htm.

Sullivan:2001:EOU

[Sul01a] Francis Sullivan. From the Editors: Onward and upward. Computing in Science and Engineering, 3(3):2–??, May/ June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3002.
pdf;   http://www.computer.
org/cse/cs1999c3002abs.
htm; pdf/c3002.pdf.

Sullivan:2001:ETS

[Sul01b] Francis Sullivan. From the Editors: Threads and surprises. Computing in Science and Engineering, 3 (4):3–??, July/August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2001/pdf/c4003.
                                        pdf.

Sullivan:2001:C

[Sul01c] Francis Sullivan. On certainty. Computing in Science and Engineering, 3(5):5– 7, September/October 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c5005abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2001/pdf/c5005.pdf.

Sullivan:2001:WSD

[Sul01d] Francis Sullivan. Willing suspension of disbelief. Computing in Science and Engineering, 3(6):3–??, November/December 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c6003abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2001/pdf/c6003.pdf.

Sullivan:2002:EMI

[Sul02a]  Francis Sullivan.  EWD:

Making it simple is not easy. Computing in Science and Engineering, 4 (6):2–3, November/December 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2002/06/c6002.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/06/
c6002.pdf.

Sullivan:2002:IAC

[Sul02b]  Francis Sullivan.  It all

comes down to B NEW ⇐ B_{0} ⊕ (B_{1} ∨ B_{2}). Computing in Science and Engineering, 4(5):2–3, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/05/c5002abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/05/c5002.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/05/
c5002.pdf.

Sullivan:2002:LNY

[Sul02c] Francis Sullivan. “… And next year, we’re going to do 3D problems”. Computing in Science and Engineering, 4(3): 3–4, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/03/c3003abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/03/c3003.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/03/
c3003.pdf.

Sullivan:2002:OWT

[Sul02d] Francis Sullivan. Oh, what a tangled Web we weave. Computing in Science and Engineering, 4(4):3–5, July/ August 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/04/c4003abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/04/c4003.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/04/
                                        c4003.pdf.

Sullivan:2002:TV

[Sul02e]  Francis Sullivan.  Trust

but verify. Computing in Science and Engineering, 4 (2):3–4, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c2003abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2002/pdf/c2003.pdf.

Sullivan:2002:WN

[Sul02f]  Francis Sullivan.  What’s

new. Computing in Science and Engineering, 4(1): 3–4, January/February 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2002/
                                        c1003abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2002/pdf/c1003.pdf.

Sullivan:2003:CSW

[Sul03a] Francis Sullivan. Computational science: We’ll know it when we see it. Computing in Science and Engineering, 5 (4):2, July/August 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/04/c4002.pdf;
http://csdl.computer.org/
dl/mags/cs/2003/04/c4002.
htm.

Sullivan:2003:EAH

[Sul03b] Francis Sullivan. From the Editors: Ask the hard questions. Computing in Science and Engineering, 5(6):3–5, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/06/c6003.pdf;
http://csdl.computer.org/
comp/mags/cs/2003/06/c6003abs.
htm.

Sullivan:2003:EFJ

[Sul03c]  Francis Sullivan.  From

the Editors: Frankenstein Jr. Computing in Science and Engineering, 5 (1):2–3, January/February 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/01/c1002.
htm;   http://csdl.computer.

                                        org/dl/mags/cs/2003/01/
                                        c1002.pdf.

Sullivan:2003:ELM

[Sul03d] Francis Sullivan. From the Editors: Less is more (is less). Computing in Science and Engineering, 5(2): 3, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/02/c2003abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/02/c2003.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/02/
                                        c2003.pdf.

Sullivan:2003:ERR

[Sul03e] Francis Sullivan. From the Editors: Reproduced and reproducible results. Computing in Science and Engineering, 5(5):3–4, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/05/c5003.pdf;
                                        http://csdl.computer.org/
                                        comp/mags/cs/2003/05/c5003abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/05/
                                        c5003.htm.

Sullivan:2003:HRY

[Sul03f] Francis Sullivan. How right you are! Computing in Science and Engineering, 5(3): 2–3, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/03/c3002abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/03/c3002.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/03/
c3002.pdf.

Sullivan:2004:ECS

[Sul04a] Francis Sullivan. From the Editors: Computational science and pathological science. Computing in Science and Engineering, 6(3): 2–3, May/June 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/03/c3002.pdf;
http://csdl.computer.org/
comp/mags/cs/2004/03/c3002abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/03/
c3002.htm.

Sullivan:2004:EGB

[Sul04b]  Francis Sullivan.  From

the Editors: Good, bad, or indifferent? Computing in Science and Engineering, 6(6):3, November/December 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/06/c6003.pdf;
http://csdl.computer.org/
dl/mags/cs/2004/06/c6003.
htm.

Sullivan:2004:EPN

[Sul04c] Francis Sullivan. From the Editors: P ≠ NP. Computing in Science and Engineering, 6(4): 2–??, July/August 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/04/c4002.pdf;
                                        http://csdl.computer.org/
                                        dl/mags/cs/2004/04/c4002.
                                        htm.

Sullivan:2004:ESE

[Sul04d] Francis Sullivan. From the Editors: Say every word on every slide. Computing in Science and Engineering, 6(1):3–4, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/01/c1003.pdf;
                                        http://csdl.computer.org/
                                        comp/mags/cs/2004/01/c1003abs.
                                        htm.

Sullivan:2004:ESW

[Sul04e] Francis Sullivan. From the Editors: Sez who? Computing in Science and Engineering, 6(2): 4–5, March/April 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2004/02/c2004.pdf;
                                        http://csdl.computer.org/
                                        comp/mags/cs/2004/02/c2004abs.
                                        htm;   http://csdl.computer.

org/dl/mags/cs/2004/02/
c2004.htm.

Sullivan:2004:EFA

[Sul04f] Francis Sullivan. From the Editors: The future ain’t what it used to be. Computing in Science and Engineering, 6(5):3, September/October 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/05/c5003.pdf;
http://csdl.computer.org/
dl/mags/cs/2004/05/c5003.
htm.

Sullivan:2005:BRY

[Sul05a]  F. Sullivan.  Book re-

view: Is that your final answer? Computing in Science and Engineering, 7 (4):67, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463138.pdf?
isnumber=31456&prod=JNL&
arnumber=1463138&arSt=+
67&ared=+67&arAuthor=Sullivan%
2C+F.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=31456&arnumber=
1463138&count=14&index=8.

Sullivan:2005:LHF

[Sul05b] F. Sullivan. Let a hundred flowers bloom and a hundred schools of thought contend! Computing in Science and Engineering, 7 (4):96, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        5992/31456/01463143.pdf?
                                        isnumber=31456&prod=JNL&
                                        arnumber=1463143&arSt=+
                                        96&ared=+96&arAuthor=Sullivan%
                                        2C+F.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=31456&arnumber=
                                        1463143&count=14&index=
                                        13.

Sullivan:2006:BC

[Sul06a] Francis Sullivan. Born to compute. Computing in Science and Engineering, 8(4): 88, July/August 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/.

Sullivan:2006:NAB

[Sul06b] Francis Sullivan. Is numerical analysis boring? Computing in Science and Engineering, 8(6):104, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/
                                        ;   http://csdl.computer.
                                        org/comp/mags/cs/2006/06/
                                        c6104.pdf.

Sullivan:2006:WIL

[Sul06c] Francis Sullivan. What is it like to be a bot? Computing in Science and Engineering, 8(1):

96, January/February 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sullivan:2007:PWS

[Sul07a] Francis Sullivan. Is this the party to whom I am speaking? Computing in Science and Engineering, 9(5):96, September/October 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cs/2007/05/
c5096.pdf.

Sullivan:2007:WA

[Sul07b]  Francis Sullivan.  Wrong

again! Computing in Science and Engineering, 9(3): 96, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.
org/comp/mags/cs/2007/03/
c3096.pdf.

Sullivan:2008:CYE

[Sul08a]  Francis Sullivan.  Curb

your enthusiasm. Computing in Science and Engineering, 10(1):96, January/ February 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://bell.
computer.org/dlcomments/
;   http://csdl.computer.

                                        org/comp/mags/cs/2008/01/
                                        mcs2008010096.pdf.

Sullivan:2008:WLC

[Sul08b] Francis Sullivan. I wandered lonely as a cloud. Computing in Science and Engineering, 10 (3):88, May/June 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://bell.
                                        computer.org/dlcomments/
                                        ;   http://csdl.computer.
                                        org/comp/mags/cs/2008/03/
                                        mcs2008030088.pdf.

Sullivan:2008:TS

[Sul08c] Francis Sullivan. This time for sure. Computing in Science and Engineering, 10(5): 80, September/October 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sullivan:2009:GEI

[Sul09a] Francis Sullivan. Guest Editor’s introduction: Cloud computing for the sciences. Computing in Science and Engineering, 11(4):10–11, July/ August 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sullivan:2009:WTN

[Sul09b] Francis Sullivan. That was then, this is now. Computing in Science and Engineering, 11(1):80, January/February 2009. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Sullivan:2010:AA

[Sul10a] Francis Sullivan. The art of approximation. Computing in Science and Engineering, 12(6):59–61, November/December 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sullivan:2010:STC

[Sul10b] Francis Sullivan. The shape of things to come. Computing in Science and Engineering, 12(2):96, March/April 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sullivan:2010:WUS

[Sul10c] Francis Sullivan. Whip until solved. Computing in Science and Engineering, 12(1):73–75, January/February 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sullivan:2017:LMC

[Sul17]  Francis Sullivan.  Learn-

ing Monte Carlo. Computing in Science and Engineering, 19(1):86–87, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs2017010086.
html.

Sadlo:2011:VCB

[SÜP^{+}11] Filip Sadlo, Markus Üffinger, Christian Pagot, Daniel Osmari, Joao Comba, Thomas Ertl, Claus-Dieter Munz, and Daniel Weiskopf. Visualization of cell-based higher-order fields. Computing in Science and Engineering, 13(3): 84–91, May/June 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Simpson:2014:NCC

[SV14] David G. Simpson and Adolfo F. Vinas. NASA computational case study: Modeling planetary magnetic and gravitational fields. Computing in Science and Engineering, 16 (4):73–79, July/August 2014. CODEN CSENFA. ISSN 1521-9615.

Sartor:2010:MRE

[SW10] Dale Sartor and Mark Wilson. Money for research, not energy bills: Finding energy and cost savings in highperformance computer facility designs. Computing in Science and Engineering, 12(6): 11–19, November/December 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Shriwise:2022:HAR

[SWDR22] Patrick C. Shriwise, Paul P. H. Wilson, Andrew Davis, and Paul K. Romano. Hardwareaccelerated ray tracing of CAD-based geometry for Monte Carlo radiation transport. Computing in Science and Engineering, 24(2):52– 61, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Shang:2000:SAF

[SWPB00] J. S. Shang, Marcus Wagner, Yi Pan, and Douglas C. Blake. Strategies for adopting FVTD on multicomputers. Computing in Science and Engineering, 2(1):10– 21, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1010.
pdf;   http://www.computer.
org/cse/cs1999/c1010abs.
htm.

Schmitz:2021:AWS

[SYM^{+}21] Patrick Schmitz, Scott Yockel, Claire Mizumoto, Thomas Cheatham, and Dana Brunson. Advancing the workforce that supports computationally and data intensive research. Computing in Science and Engineering, 23(5):19– 27, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Shi:2008:FVU

[SYP08] Lei Shi, Xin Yang, and Hailang Pan. 3D face visualization using grid light. Computing in Science and Engineering, 10(2):48–54, March/ April 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Szalay:1999:SDS

[Sza99] Alexander S. Szalay. The Sloan Digital Sky Survey. Computing in Science and Engineering, 1(2):54–62, March/ April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c2054.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c2054abs.
                                        htm.

Szalay:2011:EDI

[Sza11] Alex Szalay. Extreme dataintensive scientific computing. Computing in Science and Engineering, 13(6): 34–41, November/December 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sinenian:2013:MMS

[SZM^{+}13] Nareg Sinenian, Alex B. Zylstra, Mario J.-E. Manuel, Johan A. Frenje, Atma D. Kanojia, Joshua Stillerman, and Richard D. Petrasso. A multithreaded modular software toolkit for control of complex experiments. Computing in Science and Engineering, 15 (1):66–75, January/February 2013. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic).

Tartakovsky:2005:GEI

[TA05] Daniel M. Tartakovsky and Francis J. Alexander. Guest Editors’ introduction: Multiphysics modeling. Computing in Science and Engineering, 7(3):14–15, May/ June 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2005/03/c3014.pdf;
http://csdl.computer.org/
comp/mags/cs/2005/03/c3014abs.
htm.

Tanyalcin:2018:LVL

[TAF^{+}18] Ibrahim Tanyalcin, Carla Al Assaf, Julien Ferte, François Ancien, Taushif Khan, Guillaume Smits, Marianne Rooman, and Wim Vranken. Lexicon visualization library and JavaScript for scientific data visualization. Computing in Science and Engineering, 20 (1):50–65, January/February 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://ieeexplore.ieee.
org/document/8291800/.

Tezaur:2023:CMI

[TAHP23] Irina Tezaur, Josefin Ahlkrona, Matthew Hoffman, and Mauro Perego. Computational modeling of ice sheets and glaciers. Computing in Science and Engineering, 25(3):4–6, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tajchman:2010:PEU

[Taj10] Marc Tajchman. Programming experiences using the X10 language. Computing in Science and Engineering, 12(6):62–69, November/December 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Trebotich:2014:HRS

[TAM^{+}14] David Trebotich, Mark F. Adams, Sergi Molins, Carl I. Steefel, and Chaopeng Shen. High-resolution simulation of pore-scale reactive transport processes associated with carbon sequestration. Computing in Science and Engineering, 16(6):22–31, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs2014060022-
abs.html.

Tan:2021:NCA

[Tan21] C. R. Tan. The nascent case for adopting Jupyter notebooks as a pedagogical tool for interdisciplinary humanities, social science, and arts education. Computing in Science and Engineering, 23(2): 107–113, March/April 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Taswell:2000:WHW

[Tas00] Carl Taswell. The what, how, and why of wavelet shrinkage denoising. Computing in Science and Engineering, 2(3):12–19, May/ June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3012.
pdf;   http://www.computer.
org/cse/cs1999/c3012abs.
htm.

Tohline:1999:GEI

[TB99] Joel Tohline and Greg L.

Bryan. Guest Editors’ introduction: Cosmology and computation. Computing in Science and Engineering, 1 (2):17–20, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2017.
pdf.

Thiruvathukal:2004:XCS

[TB04] George K. Thiruvathukal and David Beazley. XML and computational science. Computing in Science and Engineering, 6(1):74–80, January/February 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/01/c1074abs.htm;

                                         http://csdl.computer.org/
                                        dl/mags/cs/2004/01/c1074.
                                        pdf.

Taborda:2011:LSE

[TB11] Ricardo Taborda and Jacobo Bielak. Large-scale earthquake simulation: Computational seismology and complex engineering systems. Computing in Science and Engineering, 13(4):14–27, July/August 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Turilli:2019:MBB

[TBM^{+}19] M. Turilli, V. Balasubramanian, A. Merzky, I. Paraskevakos, and S. Jha. Middleware building blocks for workflow systems. Computing in Science and Engineering, 21(4): 62–75, July/August 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Trott:2021:KEC

[TBVP^{+}21] Christian Trott, Luc Berger-Vergiat, David Poliakoff, Sivasankaran Rajamanickam, Damien Lebrun-Grandie, Jonathan Madsen, Nader Al Awar, Milos Gligoric, Galen Shipman, and Geoff Womeldorff. The Kokkos EcoSystem: Comprehensive performance portability for high performance computing. Computing in Science and Engineering, 23(5):10– 18, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thomas:2021:ISJ

[TC21] R. Thomas and S. Cholia. Interactive supercomputing with Jupyter. Computing in Science and Engineering, 23(2): 93–98, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Troyer:2013:BTB

[TCCC13] Matthias Troyer, Maggie Cox, Jordan Cox, and Anne J. Cox. Books [two books reviewed]. Computing in Science and Engineering, 15(6):8–11, November/December 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Towns:2014:XAS

[TCD^{+}14] John Towns, Timothy Cockerill, Maytal Dahan, Ian Foster, Kelly Gaither, Andrew Grimshaw, Victor Hazlewood, Scott Lathrop, Dave Lifka, Gregory D. Peterson, Ralph Roskies, J. Ray Scott, and Nancy Wilkins-Diehr. XSEDE: Accelerating scientific discovery. Computing in Science and Engineering, 16(5):62–74, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/

                                        cs/2014/05/mcs2014050062-
                                        abs.html.

Tang:2009:MRU

[TDB09] Hong Tang, Yusheng Dou, and Mingze Bai. Molecular response to ultrashort laser pulses. Computing in Science and Engineering, 11(3): 47–53, May/June 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tasnim:2022:ITE

[TDT^{+}22] Humayra Tasnim, Soumya Dutta, Terece L. Turton, David H. Rogers, and Melanie E. Moses. Information-theoretic exploration of multivariate time-varying image databases. Computing in Science and Engineering, 24(3):61–70, May/ June 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tecuci:2018:EBRa

[Tec18] Gheorghe Tecuci. Evidencebased reasoning and applications. Computing in Science and Engineering, 20(6):6– 8, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/cs/2018/06/08625868-
                                        abs.html.

Terrel:2011:ECA

[Ter11] Andy R. Terrel. From equations to code: Automated scientific computing. Computing in Science and Engineering, 13(2):78–82, March/April 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Teske:2015:CIC

[Tes15] Daniel Teske. A comprehensive introduction to computational science. Computing in Science and Engineering, 17(2):7–8, March/April 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/csdl/
mags/cs/2015/02/mcs2015020007.
html.

Teresco:2005:RAS

[TFF05] James D. Teresco, Jamal Faik, and Joseph E. Flaherty. Resource-aware scientific computation on a heterogeneous cluster. Computing in Science and Engineering, 7(2):40–50, March/April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2005/02/c2040.
htm;   http://csdl.computer.
org/dl/mags/cs/2005/02/
c2040.pdf.

Track:2017:EML

[TFS17] Elie Track, Nancy Forbes, and George Strawn. The end of Moore’s Law. Computing in Science and Engineering, 19 (2):4–6, March/April 2017.

CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/cs/2017/02/mcs2017020004.
                                        html.

Tohline:2009:CPM

[TGEA09] Joel E. Tohline, Jinghya Ge, Wesley Even, and Erik Anderson. A customized Python module for CFD flow analysis within VisTrails. Computing in Science and Engineering, 11(3):68–73, May/ June 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tabar:2006:NCA

[TGP^{+}06] Mohammad Reza Rahimi Tabar, Fatemeh Ghasemi, Joachim Peinke, Rudolf Friedrich, Kamran Kaviani, Fatemeh Taghavi, Sara Sadeghi, Golnoosh Bizhani, and Muhammad Sahimi. New computational approaches to the analysis of interbeat intervals in human subjects. Computing in Science and Engineering, 8(2): 54–65, March/April 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Taufer:2013:GEM

[TGP13] Michela Taufer, Narayan Ganesan, and Sandeep Patel. GPU-enabled macromolecular simulation: Challenges and opportunities. Computing in Science and Engineering, 15 (1):56–57, January/February 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tolk:2021:CDS

[TGU21] A. Tolk, C. Glazner, and J. Ungerleider. Computational decision support for the COVID-19 healthcare coalition. Computing in Science and Engineering, 23(1):17–24, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thakar:2008:LLS

[Tha08a]  Ani R. Thakar.  Lessons

learned from the SDSS catalog archive server. Computing in Science and Engineering, 10(6):65–71, November/December 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thakar:2008:SDS

[Tha08b] Ani R. Thakar. The Sloan Digital Sky Survey: Drinking from the fire hose. Computing in Science and Engineering, 10(1):9–12, January/February 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Than:2014:TNG

[Tha14]  Ker Than.  Training the

next generation of computational geoscientists. Computing in Science and Engineering, 16(6):100–103, November/December 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/06/mcs2014060100-
abs.html.

Theis:2003:BST

[The03] Thomas N. Theis. Beyond the silicon transistor: Personal observations. Computing in Science and Engineering, 5(1):25–29, January/February 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/01/c1025abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/01/c1025.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/01/
c1025.pdf.

Tremblay:2007:IFA

[THGS07] L. Bruno Tremblay, Marika M. Holland, Irina V. Gorodetskaya, and Gavin A. Schmidt. An ice-free Arctic? Opportunities for computational science. Computing in Science and Engineering, 9(3): 65–74, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2002:JMA

[Thi02] George K. Thiruvathukal. Java at middle age: Enabling Java for computational science. Computing in Science and Engineering, 4(1):74– 84, January/February 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c1074abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c1074.pdf.

Thiruvathukal:2004:GLN

[Thi04] George K. Thiruvathukal.

Gentoo Linux: The next generation of Linux. Computing in Science and Engineering, 6(5):66–74, September/October 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/05/c5066.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/05/
c5066.pdf.

Thiruvathukal:2005:GEI

[Thi05] George K. Thiruvathukal.

Guest Editors’ introduction: Cluster computing. Computing in Science and Engineering, 7(2):11–13, March/April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//csdl.computer.org/comp/
mags/cs/2005/02/c2011.pdf;
 http://csdl.computer.org/
dl/mags/cs/2005/02/c2011.
htm.

Thiruvathukal:2006:HN

[Thi06] George K. Thiruvathukal. Home networking. Computing in Science and Engineering, 8 (1):84–91, January/February 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2007:PHE

[Thi07] George K. Thiruvathukal. Project hosting: Expanding the scientific programmer’s toolbox. Computing in Science and Engineering, 9(2): 70–75, March/April 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2009:CTD

[Thi09a] George K. Thiruvathukal. Computational thinking — and doing. Computing in Science and Engineering, 11(6): 4, November/December 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2009:ICN

[Thi09b] George K. Thiruvathukal. Introducing computing now. Computing in Science and Engineering, 11(4):8–10, July/ August 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2010:YLC

[Thi10] George K. Thiruvathukal. Your local cloud-enabled library. Computing in Science and Engineering, 12(4): 5–6, July/August 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2011:BCB

[Thi11a] George K. Thiruvathukal. Beyond CiSE and back to the future. Computing in Science and Engineering, 13(3): 4–5, May/June 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2011:EUE

[Thi11b] George K. Thiruvathukal. An exceptionally useful exploration. Computing in Science and Engineering, 13(1):5–8, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2012:ALD

[Thi12a] George K. Thiruvathukal. Accelerating learning with distance education and open courseware. Computing in Science and Engineering, 14(4): 4–5, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2012:DD

[Thi12b] George K. Thiruvathukal.

Digging into data. Computing in Science and Engineering, 14(2):4–5, March/April 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2013:CSD

[Thi13a] George K. Thiruvathukal. Computational science, demystified — the future, revealed — and CiSE, 2013. Computing in Science and Engineering, 15(2):4–5, March/ April 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2013:PCO

[Thi13b] George K. Thiruvathukal. Productivity in the cognitive overload era. Computing in Science and Engineering, 15 (3):4–5, May/June 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2013:WA

[Thi13c] George K. Thiruvathukal. What’s in an algorithm? Computing in Science and Engineering, 15(4):4–5, July/ August 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2013:WNT

[Thi13d] George K. Thiruvathukal. Who needs tablets? We do. Computing in Science and Engineering, 15(1):4–6, January/February 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2014:WWP

[Thi14] George K. Thiruvathukal.

What we publish in CiSE. Computing in Science and Engineering, 16(2):4–6, March/ April 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2015:ADF

[Thi15a] George K. Thiruvathukal. The all-digital future and digital CiSE. Computing in Science and Engineering, 17(3): 4–5, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/03/mcs2015030004.
html.

Thiruvathukal:2015:CCS

[Thi15b] George K. Thiruvathukal.

Cloudy with a chance of sunshine, or the future of magazine publishing. Computing in Science and Engineering, 17 (6):4–5, November/December 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thiruvathukal:2015:NGC

[Thi15c] George K. Thiruvathukal. The next generation of computational science and engineering. Computing in Science and Engineering, 17(1):4–5, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/01/mcs2015010004.
pdf.

Thiruvathukal:2016:F

[Thi16] George K. Thiruvathukal. El fin. Computing in Science and Engineering, 18(6):4– 6, November/December 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2016/06/mcs2016060004.
html.

Thiruvathukal:2010:VCS

[THLK10] George K. Thiruvathukal, Konrad Hinsen, Konstantin Laufer, and Joe Kaylor. Virtualization for computational scientists. Computing in Science and Engineering, 12(4): 52–61, July/August 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thompson:1999:CPD

[Tho99a] William J. Thompson. Computing prescriptions: Don’t subtract the background. Computing in Science and Engineering, 1(5):84–88, September/October 1999. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/

books/cs1999/pdf/c5084.
pdf;   http://www.computer.
org/cse/cs1999/c5084abs.
htm.

Thompson:1999:CPS

[Tho99b] William J. Thompson. Computing prescriptions: Spheroidal wave functions. Computing in Science and Engineering, 1(3):84–87, May/June 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c3084.
pdf;   http://www.computer.
org/cse/cs1999/c3084abs.
htm.

Thoresen:1999:WMT

[Tho99c] Michael Thoresen. Web mechanics: TEX to Web — fast and easy. Computing in Science and Engineering, 1 (4):63–65, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c4063.
pdf.

Thompson:2000:CPV

[Tho00] William J. Thompson. Computing prescriptions: Visualizing transcendental functions. Computing in Science and Engineering, 2(6): 98–103, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/

                                        books/cs2000/pdf/c6098.
                                        pdf.

Thompson:2001:CPP

[Tho01] William J. Thompson. Computing prescriptions: Poisson distributions. Computing in Science and Engineering, 3(3):78–82, May/June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2001/pdf/c3078.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999c3078abs.
                                        htm.

Thornton:2012:PLS

[Tho12] Mitchell A. Thornton. Professional licensure for software engineers: An update. Computing in Science and Engineering, 14(5):85–87, September/October 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thyng:2020:IC

[Thy20] K. M. Thyng. The importance of colormaps. Computing in Science and Engineering, 22(5):96–102, September/October 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tien:2016:PPM

[Tie16] Vivienne Tien. Python and physical modeling. Computing in Science and Engineering, 18(3):8–10, May/June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tomar:2014:BRD

[TJ14] Amit Kumar Singh Tomar and Rita Jain. 20-bit RISC and DSP system design in an FPGA. Computing in Science and Engineering, 16(2): 16–20, March/April 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Trisovic:2020:PTL

[TJCC20] A. Trisovic, C. R. Jones, B. Couturier, and M. Clemencic. Provenance tracking in the LHCb software. Computing in Science and Engineering, 22(2):88–94, March/April 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Taylor:2006:UCM

[TK06] Jaime R. Taylor and B. Alex King, III. Using computational methods to reinvigorate an undergraduate physics curriculum. Computing in Science and Engineering, 8(5): 38–43, September/October 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tecuci:2018:EBRb

[TKM^{+}18] Gheorghe Tecuci, Louis Kaiser, Dorin Marcu, Chirag Uttamsingh, and Mihai Boicu.

Evidence-based reasoning in intelligence analysis: Structured methodology and system. Computing in Science and Engineering, 20(6):9–21, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/
mags/cs/2018/06/08625900-
abs.html.

Thiruvathukal:2004:NXD

[TL04a] George K. Thiruvathukal and Konstantin Läufer. Natural XML for data binding, processing, and persistence. Computing in Science and Engineering, 6(2): 86–92, March/April 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/02/c2086abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2004/02/c2086.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/02/
c2086.pdf.

Thiruvathukal:2004:PCM

[TL04b] George K. Thiruvathukal and Konstantin Läufer. Plone and content management. Computing in Science and Engineering, 6(4):88–95, July/ August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/dl/mags/cs/
2004/04/c4088.htm;   http:

//csdl.computer.org/dl/
mags/cs/2004/04/c4088.pdf.

Thiruvathukal:2008:WDM

[TL08a] George K. Thiruvathukal and Konstantin Laufer. What I did on my summer vacation. Computing in Science and Engineering, 10(6):76–81, November/December 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tsai:2008:SDA

[TL08b] Shan-Ho Tsai and David P.

Landau. Spin dynamics: An atomistic simulation tool for magnetic systems. Computing in Science and Engineering, 10 (1):72–79, January/February 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thomas:2002:PII

[TLD02] Stephen J. Thomas, Richard D.

Loft, and John M. Dennis. Parallel implementation issues: Global versus local methods. Computing in Science and Engineering, 4(5):26–31, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/05/c5026abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/05/c5026.
htm;   http://csdl.computer.

                                        org/dl/mags/cs/2002/05/
                                        c5026.pdf.

Thiruvathukal:2006:UTC

[TLG06] George K. Thiruvathukal, Konstantin Läufer, and Benjamin Gonzalez. Unit testing considered useful. Computing in Science and Engineering, 8(6):76–87, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Toghraee:2010:SIP

[TLR10] Reza Toghraee, Kyu-Il Lee, and Umberto Ravaioli. Simulation of ion permeation in biological membranes. Computing in Science and Engineering, 12(2):43–47, March/April 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Truhlar:2000:GEI

[TM00] Donald G. Truhlar and Vincent McKoy. Guest Editors’ introduction: Computational chemistry. Computing in Science and Engineering, 2(6): 19–21, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c6019.
                                        pdf;   pdf/c6019.pdf.

Toronto:2014:PAF

[TM14] Neil Toronto and Jay Mc-Carthy. Practically accurate floating-point math. Computing in Science and Engineering, 16(4):80–95, July/August 2014. CODEN CSENFA. ISSN 1521-9615.

Tang:2015:CTM

[TM15] Cheng Tang and Claire Monteleoni. Can topic modeling shed light on climate extremes? Computing in Science and Engineering, 17(6): 43–52, November/December 2015. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tolk:2013:AAG

[TMC^{+}13] Andreas Tolk, Geoffrey T.

Miller, Andrew E. Cross, Justin Maestri, and Benjamin Cawrse. AIMS: Applying game technology to advance medical education. Computing in Science and Engineering, 15(6):82–91, November/December 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tecuci:2018:EBD

[TMMB18] Gheorghe Tecuci, Dorin Marcu, Steven Meckl, and Mihai Boicu. Evidence-based detection of advanced persistent threats. Computing in Science and Engineering, 20(6):54–65, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/

                                        mags/cs/2018/06/08492519-
                                        abs.html.

Travnicek:2023:ELT

[TNR^{+}23] Cornelia Travnicek, Veronika Nowak, Rudolf Ramler, Lukas Fischer, and Katja Bühler. Examples of long-term science industry partnerships for translational computer science. Computing in Science and Engineering, 25(1):52–56, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thompson:2002:PBF

[TNV^{+}02] David S. Thompson, Jaya Sreevalsan Nair, Satya Sridhar Dusi Venkata, Raghu K. Machiraju, Ming Jiang, and Gheorghe Craciun. Physicsbased feature mining for large data exploration. Computing in Science and Engineering, 4(4):22–30, July/ August 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/04/c4022abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/04/c4022.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/04/
                                        c4022.pdf.

Tobis:2005:POS

[Tob05] M. Tobis. PyNSol: objects as scaffolding. Computing in Science and Engineering, 7 (4):84–91, July/August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463141.pdf?
isnumber=31456&prod=JNL&
arnumber=1463141&arSt=+
84&ared=+91&arAuthor=Tobis%
2C+M.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=31456&arnumber=
1463141&count=14&index=
11.

Tofan:2008:DCE

[Tof08] Daniel Tofan. Drawing chemical equipment with Adobe Illustrator, part 1: Basic drawing and selection tools. Computing in Science and Engineering, 10(6):100–102, November/December 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tofan:2009:DCEa

[Tof09a] Daniel Tofan. Drawing chemical equipment with Adobe Illustrator, part 2: Creating and modifying shapes. Computing in Science and Engineering, 11(1):59–61, January/February 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tofan:2009:DCEb

[Tof09b] Daniel Tofan. Drawing chemical equipment with Adobe Illustrator, part 3: Gradients, retouching, and more objects. Computing in Science and Engineering, 11(2): 60–61, March/April 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tohline:2007:SVN

[Toh07] Joel E. Tohline. Scientific visualization: a necessary chore. Computing in Science and Engineering, 9(6):76–81, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tohline:2008:WMD

[Toh08] Joel E. Tohline. Where is my digital holographic display? Computing in Science and Engineering, 10(4): 84–85, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tougaw:2000:TNR

[Tou00] Douglas Tougaw. Technology news & reviews: Sun PCi provides the best of two worlds. Computing in Science and Engineering, 2 (4):4–8, July/August 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c4004.
                                        pdf.

Tougaw:2001:TNR

[Tou01] Douglas Tougaw. Technology news & reviews: Agilent’s mixed-signal oscilloscopes provide enhanced measurement capabilities. Computing in Science and Engineering, 3(3):11–14, May/ June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c3011.
pdf;   http://www.computer.
org/cse/cs1999c3011abs.
htm.

Tougaw:2002:FYW

[Tou02a] Doug Tougaw. Finding your way with the Garmin GPS V. Computing in Science and Engineering, 4(3):10– 13, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/03/c3010abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/03/c3010.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/03/
c3010.pdf.

Tougaw:2002:SSP

[Tou02b] Douglas Tougaw. The Samsung SVP-6000 presents several benefits. Computing in Science and Engineering, 4(6): 11–13, November/December 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2002/06/c6011.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/06/
                                        c6011.pdf.

Tougaw:2003:NIR

[Tou03]  Doug Tougaw.  National

Instruments records a hit with ELVIS. Computing in Science and Engineering, 5(6):10–12, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/06/c6010abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/06/c6010.
                                        pdf.

Townsley:2009:TUF

[Tow09] Dean M. Townsley. Treating unresolvable flame physics in simulations of thermonuclear supernovae. Computing in Science and Engineering, 11(2):18–23, March/April 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Towns:2018:TOS

[Tow18] John Towns. Toward an open, sustainable national advanced computing ecosystem. Computing in Science and Engineering, 20(5):39–46, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.

computer.org/csdl/mags/
cs/2018/05/mcs2018050039-
abs.html.

Trucano:2004:GEI

[TP04] Timothy Trucano and Douglass Post. Guest Editors’ introduction: Verification and validation in computational science and engineering. Computing in Science and Engineering, 6(5):8–9, September/October 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/05/c5008.pdf;
http://csdl.computer.org/
dl/mags/cs/2004/05/c5008.
htm.

Thiruvathukal:2013:CC

[TP13] George K. Thiruvathukal and Manish Parashar. Cloud computing. Computing in Science and Engineering, 15(4): 8–9, July/August 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thijsse:2008:FDF

[TR08] Barend J. Thijsse and Bert W.

Rust. Freestyle data fitting and global temperatures. Computing in Science and Engineering, 10(1):49–59, January/February 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Trester:1999:ECS

[Tre99] Seymour Trester. Education: Computer-simulated Fresnel diffraction using the Fourier transform. Computing in Science and Engineering, 1(5): 77–83, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c5077.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c5077abs.
                                        htm.

Tougaw:2002:SCE

[TS02] Douglas Tougaw and Jon Sanders. SunRay: a costeffective desktop computer solution. Computing in Science and Engineering, 4(1):15– 17, January/February 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c1015abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2002/pdf/c1015.pdf.

Tohline:2010:VJS

[TS10] Joel E. Tohline and Emanuele Santos. Visualizing a journal that serves the computational sciences community. Computing in Science and Engineering, 12(3):78–81, May/ June 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Tsaftaris:2014:SGC

[Tsa14] Sotirios A. Tsaftaris. A scientist’s guide to cloud computing. Computing in Science and Engineering, 16(1):70–76, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thakar:2008:CAS

[TSFG08] Ani R. Thakar, Alex Szalay, George Fekete, and Jim Gray. The Catalog Archive Server Database Management System. Computing in Science and Engineering, 10(1):30–37, January/February 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thakar:2003:MMA

[TSKG03] Ani Thakar, Alex Szalay, Peter Kunszt, and Jim Gray. Migrating a multiterabyte archive from object to relational databases. Computing in Science and Engineering, 5(5):16–29, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/05/c5016abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/05/c5016.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/05/
c5016.pdf.

Terrel:2015:SSC

[TTT15] Andy Terrel, Michael Tobis, and George K. Thiruvathukal. Scientific software communities. Computing in Science and Engineering, 17(1):8–10, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/01/mcs2015010008.
                                        pdf.

Turk:2014:FCC

[Tur14a] Matthew Turk. Fostering collaborative computational science. Computing in Science and Engineering, 16(2): 68–71, March/April 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Turley:2014:NLP

[Tur14b] R. Steven Turley. Numerical literacy for physics undergraduates [book review]. Computing in Science and Engineering, 16(2):8–9, March/April 2014. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Turk:2015:VI

[Tur15] Matthew Turk. Vertical integration. Computing in Science and Engineering, 17(1):64–66, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.

computer.org/csdl/mags/
cs/2015/01/mcs2015010064-
abs.html.

Turk:2016:R

[Tur16] Matthew Turk. Risks. Computing in Science and Engineering, 18(6):63–65, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs2016060063-
abs.html.

Tougaw:2003:VFV

[TW03] Doug Tougaw and Jeff Will.

Visualizing the future of virtual reality. Computing in Science and Engineering, 5(4):8–11, July/August 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/04/c4008.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/04/
c4008.pdf.

Theis:2017:EML

[TW17] Thomas N. Theis and H.- S. Philip Wong. The end of Moore’s Law: A new beginning for information technology. Computing in Science and Engineering, 19(2): 41–50, March/April 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/

cs/2017/02/mcs2017020041-
abs.html.

Tang:2014:SDF

[TWE14] William Tang, Bei Wang, and Stephane Ethier. Scientific discovery in fusion plasma turbulence simulations at extreme scale. Computing in Science and Engineering, 16(5):44–52, September/October 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2014/05/mcs2014050044-
abs.html.

Tartakovsky:2007:GEI

[TX07] Daniel M. Tartakovsky and Dongbin Xiu. Guest Editors’ introduction: Stochastic modeling of complex systems. Computing in Science and Engineering, 9(2): 8–9, March/April 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2007/02/c2008.pdf.

Tinker:2008:ACS

[TX08] Robert F. Tinker and Qian Xie. Applying computational science to education: The molecular workbench paradigm. Computing in Science and Engineering, 10(5): 24–27, September/October 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Utz:2011:SFI

[UGV11] Alexander Utz, Lutz Gendrisch, and Holger Vogt. Simulating far-infrared scenarios with the radiance synthetic imaging system. Computing in Science and Engineering, 13 (4):98–103, July/August 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ufimtsev:2008:GPU

[UM08] Ivan S. Ufimtsev and Todd J.

Martinez. Graphical processing units for quantum chemistry. Computing in Science and Engineering, 10(6): 26–34, November/December 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Unat:2012:AFD

[UZC^{+}12] Didem Unat, Jun Zhou, Yifeng Cui, Scott B. Baden, and Xing Cai. Accelerating a 43D finite-difference earthquake simulation with a Cto-CUDA translator. Computing in Science and Engineering, 14(3):48–59, May/ June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vandewalle:2012:CSA

[Van12] Patrick Vandewalle. Code sharing is associated with research impact in image processing. Computing in Science and Engineering, 14(4):

42–47, July/August 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Varoquaux:2008:ACC

[Var08] Gaël Varoquaux. Agile computer control of a complex experiment. Computing in Science and Engineering, 10(2): 55–59, March/April 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vallisneri:2008:PXA

[VB08] Michele Vallisneri and Stanislav Babak. Python and XML for agile scientific computing. Computing in Science and Engineering, 10(1):80–87, January/February 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vinuesa:2022:ETM

[VB22] Ricardo Vinuesa and Steven L. Brunton. Emerging trends in machine learning for computational fluid dynamics. Computing in Science and Engineering, 24(5):33–41, September/October 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vo:2011:SEP

[VCGS11] Huy T. Vo, Joao L. D. Comba, Berk Geveci, and Claudio T. Silva. Streaming-enabled parallel data flow framework in the Visualization ToolKit. Computing in Science and Engineering, 13(5): 72–83, September/October 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

VanZee:2009:LLD

[VCvdG^{+}09] Field G. Van Zee, Ernie Chan, Robert A. van de Geijn, Enrique S. Quintana-Orti, and Gregorio Quintana-Orti. The libflame library for dense matrix computations. Computing in Science and Engineering, 11(6):56–63, November/December 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

vanderVorst:2000:KSI

[vdV00] Henk A. van der Vorst. Krylov subspace iteration. Computing in Science and Engineering, 2 (1):32–37, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c1032.
pdf;   http://www.computer.
org/cse/cs1999/c1032abs.
htm.

vanderWalt:2011:NAS

[vdWCV11] Stefan van der Walt, S. Chris Colbert, and Gael Varoquaux. The NumPy array: a structure for efficient numerical computation. Computing in Science and Engineering, 13(2):

22–30, March/April 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vetter:2011:KBH

[VGD^{+}11] Jeffrey S. Vetter, Richard Glassbrook, Jack Dongarra, Karsten Schwan, Bruce Loftis, Stephen McNally, Jeremy Meredith, James Rogers, Philip Roth, Kyle Spafford, and Sudhakar Yalamanchili. Keeneland: Bringing heterogeneous GPU computing to the computational science community. Computing in Science and Engineering, 13(5): 90–95, September/October 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

vanGelder:2018:SCE

[vGDS18] Tim van Gelder, Richard De Rozario, and Richard O. Sinnott. SWARM: Cultivating evidence-based reasoning. Computing in Science and Engineering, 20(6):22–34, November/December 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/cs/2018/06/08487002-
                                        abs.html.

Vesperini:2009:BAS

[VGM^{+}09] Enrico Vesperini, David M. Goldberg, Stephen L. W. McMillan, James Dura, and Douglas Jones. The Beowulf Analysis Symbolic INterface: Interactive parallel data analysis for everyone. Computing in Science and Engineering, 11(2):45–51, March/April 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Villalon:2008:HDD

[Vil08] Elena Villalon. High-dimensionality data reduction with Java. Computing in Science and Engineering, 10(5):64–69, September/October 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vinoski:2012:CMP

[Vin12] Steve Vinoski. Concurrency and message passing in Erlang. Computing in Science and Engineering, 14(6): 24–34, November/December 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Virmani:2016:CFU

[Vir16] Vineet Virmani. Computational finance using QuantLib-Python. Computing in Science and Engineering, 18(2): 78–88, March/April 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vashishta:1999:LSA

[VKN99] Priya Vashishta, Rajiv K.

Kalia, and Aiichori Nakano.

Large-scale atomistic simulations of dynamic fracture. Computing in Science and Engineering, 1(5): 56–65, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5056.
pdf;   http://www.computer.
org/cse/cs1999/c5056abs.
htm.

Vlasak:2012:AAU

[Vla12] Weldon Vlasak. Analyzing atoms using the Spice computer program. Computing in Science and Engineering, 14(3):98–103, May/June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vetter:2015:ONM

[VM15] Jeffrey S. Vetter and Sparsh Mittal. Opportunities for nonvolatile memory systems in extreme-scale high-performance computing. Computing in Science and Engineering, 17(2): 73–82, March/April 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/02/mcs2015020073-
abs.html.

Viamontes:2005:QSP

[VMH05] George F. Viamontes, Igor L. Markov, and John P. Hayes. Is quantum search practical? Computing in Science and Engineering, 7(3): 62–70, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/03/c3062abs.htm;
http://csdl.computer.org/
dl/mags/cs/2005/03/c3062.
pdf.

Venkatakrishnan:2020:HLS

[VMK20] R. Venkatakrishnan, A. Misra, and V. Kindratenko. Highlevel synthesis-based approach for accelerating scientific codes on FPGAs. Computing in Science and Engineering, 22(4):104–109, July/August 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vashishta:1999:GEI

[VN99] Priya Vashishta and Aiichiro Nakano. Guest Editors’ introduction: Dynamic fracture analysis. Computing in Science and Engineering, 1(5): 20–23, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5020.
pdf;   http://www.computer.
org/cse/cs1999/c5020abs.
htm.

Vogel:2013:BWC

[Vog13] Thomas Vogel. All the Way to CUDA [book review]. Computing in Science and Engineering, 15(5):6–8, September/October 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vormoor:2001:DVP

[Vor01a] Oliver Vormoor. Data visualization: Parallel coordinates and dimension reduction. Computing in Science and Engineering, 3(5):110–113, September/ October 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c5110abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c5110.pdf.

Vormoor:2001:QEI

[Vor01b] Oliver Vormoor. Quick and easy interactive molecular dynamics using Java3D. Computing in Science and Engineering, 3(5):98–104, September/October 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c5098abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c5098.pdf.

Vakali:2004:GEI

[VP04] Athena I. Vakali and Georgios I. Papadimitriou. Guest Editors’ introduction: Web engineering–the evolution of new technologies. Computing in Science and Engineering, 6(4):10–11, July/ August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/04/c4010.pdf;
http://csdl.computer.org/
dl/mags/cs/2004/04/c4010.
htm.

Vazquez-Poletti:2018:SCP

[VPL18] Jose Luis Vazquez-Poletti and Ignacio Martin Llorente. Serverless computing: From planet Mars to the cloud. Computing in Science and Engineering, 20(6):73–79, November/December 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/06/08625892-abs.
html.

Vermaas:2021:SPS

[VSB^{+}21] J. V. Vermaas, A. Sedova, M. B. Baker, S. Boehm, D. M. Rogers, J. Larkin, J. Glaser, M. D. Smith, O. Hernandez, and J. C. Smith. Supercomputing pipelines search for therapeutics against COVID-19. Computing in Science and Engineering, 23(1):7–16, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vorp:2001:CMA

[VSE01] David A. Vorp, David A.

Steinman, and C. Ross Ethier. Computational modeling of arterial biomechanics. Computing in Science and Engineering, 3(5):51–64, September/October 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c5051abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c5051.pdf.

Vos:2002:RTA

[VSG^{+}02] Frans M. Vos, Hans J. W. Spoelder, Desmond M. Germans, Rutger Hofman, and Henri Bal. Real-time, adaptive measurement of corneal shapes. Computing in Science and Engineering, 4(2): 66–76, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c2066abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c2066.pdf.

Vargas:2009:WER

[VSMD^{+}09] Hector Vargas, Jose Sanchez-Moreno, Sebastian Dormido, Christophe Salzmann, Denis Gillet, and Francisco Esquembre. Web-enabled remote scientific environments. Computing in Science and Engineering, 11(3):36–46, May/ June 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vuduc:2022:JA

[Vud22] Richard W. Vuduc. Jack, the autotuner. Computing in Science and Engineering, 24(4): 24–27, July/August 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vieira:2018:AWL

[VVNV18] Luiz Filipe M. Vieira, Marcos Augusto M. Vieira, José Augusto M. Nacif, and Alex Borges Vieira. Autonomous wireless lake monitoring. Computing in Science and Engineering, 20 (1):66–75, January/February 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://ieeexplore.ieee.
org/document/8291775/.

Verma:2011:SRT

[VWL^{+}11] Poonam S. Verma, Huanmei Wu, Mark P. Langer, Indra J. Das, and George Sandison. Survey: Real-time tumor motion prediction for image-guided radiation treatment. Computing in Science and Engineering, 13(5):24– 35, September/October 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Viswanathan:2012:MCB

[VWP12] Venkatasubramanian Viswanathan, Frank Wang, and Heinz Pitsch. Monte Carlo-based approach for simulating nanostructured catalytic and electrocatalytic systems. Computing in Science and Engineering, 14(2):60–69, March/April 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wainer:2016:DMS

[Wai16]  Gabriel A. Wainer.  Dis-

crete modeling and simulation [Guest Editors’ introduction]. Computing in Science and Engineering, 18(4): 8–10, July/August 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wang:2018:GBT

[Wan18] Chaoli Wang. Graph-based techniques for visual analytics of scientific data sets. Computing in Science and Engineering, 20 (1):93–103, January/February 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://ieeexplore.ieee.
org/document/8291793/.

Wang:2023:ITG

[Wan23] Jie Wang. An intuitive tutorial to Gaussian process regression. Computing in Science and Engineering, 25(4): 4–11, April 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Warner:2018:IPP

[War18] Mellissa S. C. Warner. Introduction to PySPLIT: A Python toolkit for NOAA ARL’s HYSPLIT model. Computing in Science and Engineering, 20(5):47–62, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/05/mcs2018050047-
abs.html.

Wenger:2012:IVS

[WAS^{+}12] Stephan Wenger, Marco Ament, Wolfgang Steffen, Nico Koning, Daniel Weiskopf, and Marcus Magnor. Interactive visualization and simulation of astronomical nebulae. Computing in Science and Engineering, 14(3):78–87, May/ June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Watters:2021:SLC

[Wat21] A. R. Watters. Shader-like computations in WebGL for advanced graphics and general purposes. Computing in Science and Engineering, 23(2): 54–63, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wyatt:2003:UQT

[WB03] Robert E. Wyatt and Eric R.

Bittner. Using quantum trajectories and adaptive grids to solve quantum dynamical problems. Computing in Science and Engineering, 5(4): 22–30, July/August 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/04/c4022abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2003/04/c4022.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/04/
c4022.pdf.

Wofford:2020:JND

[WBB^{+}20] M. F. Wofford, B. M. Boscoe, C. L. Borgman, I. V. Pasquetto, and M. S. Golshan. Jupyter notebooks as discovery mechanisms for open science: Citation practices in the astronomy community. Computing in Science and Engineering, 22(1):5–15, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Washington:2019:RYW

[WBP^{+}19] N. Washington, T. Barnes, J. Payton, S. Dunton, F. Stukes, and A. Peterfreund. RE-SPECT 2019: Yes, we still need to talk about diversity in computing. Computing in Science and Engineering, 21 (1):79–83, January/February 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Wedi:2022:DEH

[WBS^{+}22] Nils Wedi, Peter Bauer, Irina Sandu, Jörn Hoffmann, Sophia Sheridan, Rafael Cereceda, Tiago Quintino, Daniel Thiemert, and Thomas Geenen. Destination Earth: Highperformance computing for weather and climate. Computing in Science and Engineering, 24(6):29–37, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wolf:2015:PRC

[WC15] Laura Wolf and Jim Collins.

Putting regional climate prediction in reach. Computing in Science and Engineering, 17(5):49–51, September/October 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/05/mcs2015050049-
abs.html.

Wolf:2017:MMM

[WC17] Laura Wolf and Jim Collins.

Merging the machines of modern science. Computing in Science and Engineering, 19(5):82–84, September/October 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/05/mcs2017050082.
html.

Wang:2014:BDA

[WCAL14] Jianwu Wang, Daniel Crawl, Ilkay Altintas, and Weizhong Li. Big data applications using workflows for data parallel computing. Computing in Science and Engineering, 16(4): 11–21, July/August 2014. CO-DEN CSENFA. ISSN 1521-9615.

Wang:2002:GSV

[WCC^{+}02] Xusheng Wang, Jim X. Chen, Daniel B. Carr, B. Sue Bell, and Linda Williams Pickle. Geographic statistics visualization: Web-based linked micromap plots. Computing in Science and Engineering, 4(3): 90–94, May/June 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2002/03/c3090abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2002/03/c3090.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/03/
                                        c3090.pdf.

Wu:2019:ODM

[WCC^{+}19] Xun Wu, Tefang Chen, Yating Chen, Chaoqun Xiang, Zhi Liu, and Kaidi Li. An online diagnostic method for opencircuit faults of locomotive inverter based on output voltage transient detection. Computing in Science and Engineering, 21(3):83–91, May/ June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8550717/.

Wang:2005:TEM

[WCGB05] Dali Wang, E. Carr, L. J.

Gross, and M. W. Berry. Toward ecosystem modeling on computing Grids. Computing in Science and Engineering, 7(5):44–52, September/October 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501739.pdf?
isnumber=32219&prod=JNL&
arnumber=1501739&arSt=+
44&ared=+52&arAuthor=+Dali+
Wang%3B++Carr%2C+E.%3B++
Gross%2C+L.J.%3B++Berry%
2C+M.W.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=32219&arnumber=
1501739&count=14&index=6.

Wu:2012:GBA

[WCH12] Jiadong Wu, Chunlei Chen, and Bo Hong. A GPU-based approach to accelerate computational protein-DNA docking. Computing in Science and Engineering, 14(3):20– 29, May/June 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wainer:2021:MSS

[WCHRM21] Gabriel Wainer, Román Cárdenas, Kevin Henares, and Cristina Ruiz-Martín. Modeling and simulation of space-based pandemic scenarios using an opensource platform. Computing in Science and Engineering, 23 (4):80–84, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wille:2017:ERE

[WCP17] Sarah Wille, Jeanne Century, and Miriam Pike. Exploratory research to expand opportunities in computer science for students with learning differences. Computing in Science and Engineering, 19(3): 40–50, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/03/mcs2017030040-
                                        abs.html.

Watson:2006:DSA

[WD06] Gregory R. Watson and Nathan A. DeBardeleben. Developing scientific applications using Eclipse. Computing in Science and Engineering, 8(4): 50–61, July/August 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wilkins-Diehr:2018:NIS

[WDC18] Nancy Wilkins-Diehr and T. Daniel Crawford. NSF’s inaugural software institutes: The Science Gateways Community Institute and the Molecular Sciences Software Institute. Computing in Science and Engineering, 20(5): 26–38, September/October 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://www.computer.org/
csdl/mags/cs/2018/05/mcs2018050026-
abs.html.

Wegman:2000:VCA

[Weg00] Edward J. Wegman. Visualization corner: Affordable environments for 3D collaborative data visualization. Computing in Science and Engineering, 2(6):68– 72, 74, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c6068.
pdf.

Weigel:2011:GRW

[Wei11] Martin Weigel. The GPU revolution at work. Computing in Science and Engineering, 13 (5):5–6, September/October 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Weiner:2021:SSP

[Wei21] Zachary J. Weiner. Stencil solvers for PDEs on GPUs: an example from cosmology. Computing in Science and Engineering, 23(4):55– 64, July/August 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Weppner:2008:CMD

[Wep08] Stephen Weppner. Computational methods with depth and flair. Computing in Science and Engineering, 10 (5):5–8, September/October 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Weppner:2015:DPS

[Wep15] Stephen Weppner. A different perspective on scientific programming [review of “Annotated algorithms in Python; with applications in physics, biology, and finance’ (Di Pierro, M.; 2013)]. Computing in Science and Engineering, 17(1):6–7, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/01/mcs2015010006.
                                        pdf.

Wessner:2003:SML

[Wes03] Charles Wessner. Sustaining Moore’s Law and the US economy. Computing in Science and Engineering, 5(1):30–38, January/February 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/01/c1030abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/01/c1030.
                                        htm;   http://csdl.computer.

org/dl/mags/cs/2003/01/
c1030.pdf.

West:2021:LSSa

[Wes21] J. West. Large-scale scientific computing in the fight against COVID-19. Computing in Science and Engineering, 23 (1):89–92, January/February 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wasserman:2015:NER

[WG15] Harvey J. Wasserman and Richard A. Gerber. The National Energy Research Scientific Computing Center: Forty years of supercomputing leadership. Computing in Science and Engineering, 17(3): 6–8, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/03/mcs2015030006.
html.

Winslow:2016:WPV

[WGJ16] Raimond L. Winslow, Stephen Granite, and Christian Jurado. WaveformECG: A platform for visualizing, annotating, and analyzing ECG data. Computing in Science and Engineering, 18(5):36– 46, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wainer:2021:CSB

[WHG21] G. Wainer, K. Hinsen, and K. Gaither. Computational science in the battle against COVID-19: Part II. Computing in Science and Engineering, 23(1):5–6, January/February 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wan:2022:FSC

[WHH22] Yong Wan, Holly A. Holman, and Charles Hansen. Fluo-Render script: A case study of lingua franca in translational computer science. Computing in Science and Engineering, 24(6):60–65, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wallcraft:2002:RTO

[WHM^{+}02] Alan J. Wallcraft, Jarley E. Hurlburt, E. Joseph Metzger, Robert C. Rhodes, Jay F. Shriver, and Ole Martin Smedstad. Real-time ocean modeling systems. Computing in Science and Engineering, 4 (2):50–57, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c2050abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c2050.pdf.

Woodward:2018:SSH

[WHW18] Paul R. Woodward, Falk Herwig, and Ted Wetherbee. Simulating stellar hydrodynamics at extreme scale. Computing in Science and Engineering, 20(5):8–17, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/05/mcs2018050008-
abs.html.

Will:2001:IFG

[Wil01] Jeff Will. Imagenation frame grabbers for Computer Vision Systems. Computing in Science and Engineering, 3(5):12–15, September/October 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c5012abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c5012.pdf.

Wilson:2006:SCG

[Wil06] Greg Wilson. Software carpentry: Getting scientists to write better code by making them more productive. Computing in Science and Engineering, 8(6):66–69, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wilson:2008:WWL

[Wil08] Greg Wilson. Those who will not learn from history . . .. Computing in Science and Engineering, 10(3):5–6, May/ June 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Williams:2016:DSD

[Wil16] Timothy J. Williams. Delivering science on day one. Computing in Science and Engineering, 18(2):104–107, March/April 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Williams:2017:WN

[Wil17] R. Stanley Williams. What’s next? Computing in Science and Engineering, 19(2): 7–13, March/April 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/02/mcs2017020007-
                                        abs.html.

Williams:2018:EST

[Wil18] Timothy J. Williams. Early science on Theta. Computing in Science and Engineering, 20(3):73–77, May/June 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Winch:2006:GEI

[Win06] David Winch. Guest Editor’s introduction: Computation in physics courses. Computing in Science and Engineering, 8(5):11–15, September/October 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2006/05/c5011.pdf;
http://csdl.computer.org/
comp/mags/cs/2006/extras/
c5011_x1_Landau.ppt;   http:
//csdl.computer.org/comp/
mags/cs/2006/extras/c5011_
x2_Sullivan.pdf;   http:
//csdl.computer.org/comp/
mags/cs/2006/extras/c5011_
x3_Lagan.doc;   http://
csdl.computer.org/comp/
mags/cs/2006/extras/c5011_
x4_Chabay.ppt;   http://
csdl.computer.org/comp/
mags/cs/2006/extras/c5011_
x5_Dennin.pdf;   http://
csdl.computer.org/comp/
mags/cs/2006/extras/c5011_
x6_Cook.pdf;   http://csdl.
computer.org/comp/mags/
cs/2006/extras/c5011_x7_
Dennin.ppt;   http://csdl.
computer.org/comp/mags/
cs/2006/extras/c5011_x8_
Johnston.pdf.

Witherden:2021:PPP

[Wit21] Freddie D. Witherden. Python at petascale with PyFR or: How I learned to stop worrying and love the snake. Computing in Science and Engineering, 23(4):29–37, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Washburn:2004:COD

[WJ04] Donald A. Washburn and Lauriann M. Jones. Could olfactory displays improve data visualization? Computing in Science and Engineering, 6(6): 80–83, November/December 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2004/06/c6080.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/06/
                                        c6080.pdf.

Woodward:2008:MSC

[WJLY08] Paul R. Woodward, Jagan Jayaraj, Pei-Hung Lin, and Pen-Chung Yew. Moving scientific codes to multicore microprocessor CPUs. Computing in Science and Engineering, 10(6):16–25, November/December 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wilson:2009:SEC

[WL09] Greg Wilson and Andrew Lumsdaine. Software engineering and computational science. Computing in Science and Engineering, 11(6): 12–13, November/December 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Windl:2001:CIP

[WLCD01] Wolfgang Windl, Matthew Laudon, Neil N. Carlson, and Murray S. Daw. CSE in industry: Predictive process simulation and stress-mediated diffusion in silicon. Computing in Science and Engineering, 3(4):92–95, July/August 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2001/pdf/c4092.
pdf.

Wan:2012:GDB

[WLCJ12] Xianmei Wan, Shengjun Liu, Jim X. Chen, and Xiaogang Jin. Geodesic distancebased realistic facial animation using RBF interpolation. Computing in Science and Engineering, 14(5):49– 55, September/October 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wang:2014:ISD

[WLL^{+}14] Lizhe Wang, Ke Lu, Peng Liu, Rajiv Ranjan, and Lajiao Chen. IK-SVD: Dictionary learning for spatial big data via incremental atom update. Computing in Science and Engineering, 16(4):41–52, July/August 2014. CODEN CSENFA. ISSN 1521-9615.

Weigand:2000:ISC

[WM00] Gil Weigand and Paul Messina.

Interview: Sea changes in computational power: Testing our metal. Computing in Science and Engineering, 2(2):17–20, March/April 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c2017.
pdf;   http://www.computer.
org/cse/cs1999/c2017abs.
htm.

Washington:2020:UHE

[WMB20] G. J. Washington, M. Meijias, and L. Burge. Understanding how to engage black HS boys in computer science through tech innovation and entrepreneurship. Computing in Science and Engineering, 22(5):20–28, September/October 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

West:2021:LSSb

[WNDV21] John West, Paul A. Navrátil, Maytal Dahan, and Matthew Vaughn. Large scale science on NSF’s Frontera system. Computing in Science and Engineering, 23(3):10–13, May/ June 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wang:2017:IAP

[WNZ^{+}17] Yan Wang, Laurent Navarro, Yue Zhang, Evan Kao, Yuemin Zhu, and Guy Courbebaisse. Intracranial aneurysm phantom segmentation using a 4D lattice Boltzmann method. Computing in Science and Engineering, 19(4): 56–67, July/August 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/04/mcs2017040056-
abs.html.

Watson:2010:ACQ

[WOAEAG10] Mark Watson, Roberto Olivares-Amaya, Richard G. Edgar, and Alan Aspuru-Guzik. Accelerating correlated quantum chemistry calculations using graphical processing units. Computing in Science and Engineering, 12(4): 40–51, July/August 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wolf:2016:MSS

[Wol16] Laura Wolf. Multiyear simulation study provides breakthrough in membrane protein research. Computing in Science and Engineering, 18(5): 94–97, September/October 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wolfe:2021:PPP

[Wol21] Michael Wolfe. Performant, portable, and productive parallel programming with standard languages. Computing in Science and Engineering, 23(5):39–45, September/October 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Weiland:2022:EEJ

[WP22] Michèle Weiland and Mark Parsons. EPCC’s exascale journey: a retrospective of the past 10 years and a vision of the future. Computing in Science and Engineering, 24(1): 8–13, January/February 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wang:2012:USB

[WPM^{+}12] Jun Wang, Marlon Pierce, Yu Ma, Geoffrey Fox, Andrea Donnellan, Jay Parker, and Margaret Glasscoe. Using service-based GIS to support earthquake research and disaster response. Computing in Science and Engineering, 14(5):21–30, September/October 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wang:2011:CCM

[WPW11] Dali Wang, Wilfred M. Post, and Bruce E. Wilson. Climate change modeling: Computational opportunities and challenges. Computing in Science and Engineering, 13(5):36– 42, September/October 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Woo:2000:ASM

[WPZ00] Tom K. Woo, Serguei Patchkovskii, and Tom Ziegler. Atomic scale modeling of polymerization catalysts. Computing in Science and Engineering, 2(6): 28–37, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c6028.
pdf;   http://www.computer.
org/cse/cs1999c6028abs.
htm.

Wang:2018:AOS

[WQLZ18] Yanhua Wang, Jianzhong Qiao, Shukuan Lin, and Tinglei Zhao. An approximate optimal solution to GPU workload scheduling. Computing in Science and Engineering, 20(5):63–76, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/05/mcs2018050063-
abs.html.

Wilson:2016:HCS

[WQT^{+}16] Wesley Wilson, Tony Quezon, Vu Trinh, Cullen Sarles, Jun Li, and Joseph Gorski. HPCMP CREATE-SH integrated hydrodynamic design environment. Computing in Science and Engineering, 18(6):47–56, November/December 2016. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2016/06/mcs2016060047-
                                        abs.html.

Weinhous:2000:GEI

[WR00] Martin S. Weinhous and Joseph M. Rosen. Guest Editors’ introduction: Computing in medicine. Computing in Science and Engineering, 2(5): 14–17, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c5014.
                                        pdf.

Wright:2010:TWT

[Wri10] Cameron H. G. Wright. Technical writing tools for engineers and scientists. Computing in Science and Engineering, 12(5):98–103, September/October 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wright:2016:LGB

[Wri16] Heather M. Wright. Lesbian, gay, bisexual, transgender, and queer students’ sense of belonging in computing: An intersectional approach. Computing in Science and Engineering, 18(3):24–30, May/ June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

White:1999:FUS

[WS99] Simon D. M. White and Volker Springel. Fitting the universe on a supercomputer. Computing in Science and Engineering, 1(2): 36–45, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2036.
pdf;   http://www.computer.
org/cse/cs1999/c2036abs.
htm.

Winter:2004:VWS

[WSC^{+}04] C. L. Winter, Everett P.

Springer, Keeley Costigan, Patricia Fasel, Sue Mniewski, and George Zyvoloski. Virtual watersheds: Simulating the water balance of the Rio Grande basin. Computing in Science and Engineering, 6(3): 18–26, May/June 2004. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2004/03/c3018abs.htm;
http://csdl.computer.org/
dl/mags/cs/2004/03/c3018.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/03/
c3018.pdf.

Wang:2022:TUR

[WSY^{+}22] Dali Wang, Peter Schwartz, Fengming Yuan, Peter Thornton, and Weijian Zheng. Toward ultrahigh-resolution E3SM land modeling on exascale computers. Computing in Science and Engineering, 24(6):44–53, November/December 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Weinkauf:2012:FVA

[WT12] Tino Weinkauf and Holger Theisel. Flow visualization and analysis using streak and time lines. Computing in Science and Engineering, 14(5): 78–84, September/October 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Willhoft:2017:DSD

[WW17] Blaine Willhoft and Rob Willhoft. Decoding software design. Computing in Science and Engineering, 19(3):86– 87, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs2017030086.
html.

Wilson:2020:NSS

[WWJH20] A. B. Wilson, J. Wang, E. W. Jenkins, and S. M. Husson. Numerical simulation of solid phase adsorption models using time-integrated, up-winded finite element strategies. Computing in Science and Engineering, 22(3):64–78, May/ June 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wang:2012:ADC

[WY12] Xusheng Wang and Beifang Yi. The application of data cubes in business data visualization. Computing in Science and Engineering, 14(6): 44–50, November/December 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wallqvist:2010:ABR

[WZS^{+}10] Anders Wallqvist, Nela Zavaljevski, Ravi Vijaya Satya, Rajkumar Bondugula, Valmik Desai, Xin Hu, Kamal Kumar, Michael Lee, In-Chul Yeh, and Chenggang Yu. Accelerating biomedical research in designing diagnostic assays, drugs, and vaccines. Computing in Science and Engineering, 12(5):46–55, September/October 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Weng:2011:ADA

[WZZ11] Yonghui Weng, Meng Zhang, and Fuqing Zhang. Advanced data assimilation for cloudresolving hurricane initialization and prediction. Computing in Science and Engineering, 13(1):40–49, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Xin:2010:GCN

[XBK10] Jianguo Xin, Larisa Beilina, and Michael Klibanov. Globally convergent numerical methods for some coefficient inverse problems. Computing in Science and Engineering, 12(5):64–77, September/October 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Xie:2020:RMS

[XDK^{+}20] Y. Xie, D. Du, C. B. Karki, W. Guo, A. E. Lopez-Hernandez, S. Sun, B. Y. Juarez, H. Li, J. Wang, and L. Li. Revealing the mechanism of SARS-CoV-2 spike protein binding with ACE2. Computing in Science and Engineering, 22(6): 21–29, November/December 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Xing:2013:AAL

[XHL^{+}13] Qi Xing, Rui Han, Yanlin Li, Wenzhen Yang, and Jim X. Chen. Automatically assessing limb alignment and hip fracture using 3D models. Computing in Science and Engineering, 15(2): 10–20, March/April 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Xiu:2005:EFM

[XKG05] Dongbin Xiu, Ioannis G.

Kevrekidis, and Roger Ghanem. An equation-free, multiscale approach to uncertainty quantification. Computing in Science and Engineering, 7(3): 16–23, May/June 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2005/03/c3016abs.htm;
http://csdl.computer.org/
dl/mags/cs/2005/03/c3016.
pdf.

Xu:2004:CPW

[XLLJ04] Jianliang Xu, Jiangchuan Liu, Bo Li, and Xiaohua Jia. Caching and prefetching for Web content distribution. Computing in Science and Engineering, 6(4):54–59, July/August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/04/c4054.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/04/
c4054.pdf.

Xu:2002:AAB

[XXK^{+}02] Ying Xu, Dong Xu, Dongsup Kim, Victor Olman, Jane Razumovskaya, and Tao Jiang. Automated assignment of backbone NMR peaks using constrained bipartite matching. Computing in Science and Engineering, 4(1):50– 62, January/February 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c1050abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c1050.pdf.

Xu:2005:OPJ

[XYC05] Zhigen Xu, Yusong Yan, and Jim X. Chen. OpenGL programming in Java. Computing in Science and Engineering, 7 (1):51–55, January/February 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2005/01/c1051.
htm;   http://csdl.computer.
org/dl/mags/cs/2005/01/
c1051.pdf.

Xie:2009:VPA

[XYC^{+}09] You-Zhou Xie, Lin Yang, Li-Fen Chen, Pei-Dong Dai, Tian-Yu Zhang, Jim X. Chen, and Zheng-Min Wang. A virtual platform for auditory organ mechanics analysis. Computing in Science and Engineering, 11(4):74– 80, July/August 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Xiao:2019:MEO

[XZL^{+}19] S. Xiao, K. Zhang, G. Liu, Y. Wang, J. Duan, and L. Kong. Multivariable electromagnetic optimization design exploiting hybrid kriging. Computing in Science and Engineering, 21(5):46– 54, September/October 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Yang:2000:CIR

[YAA^{+}00] Ren-Jye Yang, Alexander Akkerman, Daniel F. Anderson, Omar M. Faruque, and Lei Gu. CSE in industry: Robustness optimization for vehicular crash simulations. Computing in Science and Engineering, 2(6): 8–13, November/December 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c6008.
pdf.

Yang:2010:MOS

[YaL10] Yonggao Yang and Jian ao Lian. Making 3D object surfaces smoother. Computing in Science and Engineering, 12(3):44–51, May/ June 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Yasar:2017:ECT

[Yas17a] Osman Yasar. The essence of computational thinking. Computing in Science and Engineering, 19(4):74–82, July/ August 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/04/mcs2017040074-
                                        abs.html.

Yasar:2017:MSH

[Yas17b] Osman Yasar. Modeling and simulation: How everything seems to form and grow. Computing in Science and Engineering, 19(1):74–77, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/01/mcs2017010074-
                                        abs.html.

Yavneh:2006:WMM

[Yav06] Irad Yavneh. Why multigrid methods are so efficient. Computing in Science and Engineering, 8(6):12–22, November/December 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Yokota:2012:HBS

[YB12] Rio Yokota and Lorena A. Barba. Hierarchical N-body simulations with autotuning for heterogeneous systems. Computing in Science and Engineering, 14(3):30–39, May/ June 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Yao:2015:SHP

[YBBP15] Yushu Yao, Benjamin P. Bowen, Dalya Baron, and Dovi Poznanski. SciDB for high-performance arraystructured science data at NERSC. Computing in Science and Engineering, 17(3): 44–52, May/June 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
cs/2015/03/mcs2015030044-
abs.html.

Yang:2010:DEE

[YBD10] Xiaoyu Yang, Richard P.

Bruin, and Martin T. Dove. Developing an end-to-end scientific workflow. Computing in Science and Engineering, 12(3):52–61, May/June 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Yeager:2021:BFF

[YCD^{+}21] Stephen G. Yeager, Ping Chang, Gokhan Danabasoglu, James Edwards, Nan Rosenbloom, Qiuying Zhang, Dan Fu, Xue Liu, and Fred Castruccio. Bringing the future into focus: Benefits and challenges of high-resolution global climate change simulations. Computing in Science and Engineering, 23(3): 34–41, May/June 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Yang:2003:GEC

[YCK03] Yonggao Yang, Jim X. Chen, and Woosung Kim. Gene expression clustering and 3D visualization. Computing in Science and Engineering, 5(5):37–43, September/October 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/05/c5037abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/05/c5037.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/05/
                                        c5037.pdf.

Yang:2003:NPU

[YCKK03] Yonggao Yang, Jim X. Chen, Woosung Kim, and Changjin Kee. Nonlinear projection: Using deformations in 3D viewing. Computing in Science and Engineering, 5(2): 54–59, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/02/c2054abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/02/c2054.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/02/
                                        c2054.pdf.

Yang:2005:VHA

[YCL05] Lin Yang, J. X. Chen, and Yanling Liu. Virtual human anatomy. Computing in Science and Engineering, 7(5):71–73, September/October 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/32219/01501743.pdf?
isnumber=32219&prod=JNL&
arnumber=1501743&arSt=+
71&ared=+73&arAuthor=+Lin+
Yang%3B++Chen%2C+J.X.%3B+
+Yanling+Liu;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=32219&
arnumber=1501743&count=
14&index=9.

Yang:2007:AHM

[YCZ07] Xiaosong Yang, Jian Chang, and Jian J. Zhang. Animating the human muscle structure. Computing in Science and Engineering, 9(5):39– 45, September/October 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Yildiz:2019:HHW

[YEC^{+}19] O. Yildiz, J. Ejarque, H. Chan, S. Sankaranarayanan, R. M. Badia, and T. Peterka. Heterogeneous hierarchical workflow composition. Computing in Science and Engineering, 21(4):76–86, July/August 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Fu:2013:EPE

[yFZDY13] Yao yao Fu, Tianyu Zhang, Peidong Dai, and Lin Yang. Evaluating potential ear canal reconstruction for congenital aural atresia patients. Computing in Science and Engineering, 15(2):34–40, March/ April 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Yi:2005:RTN

[YHWY05] Beifang Yi, Frederick C. Harris, Jr., Ling Wang, and Yusong Yan. Real-time natural hand gestures. Computing in Science and Engineering, 7(3):92–96, c3, May/June 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2005/03/c3092abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2005/03/c3092.
                                        pdf.

Yang:2003:DES

[YKD^{+}03] Ruixin Yang, Menas Kafatos, Brian Doty, James L. Kinter, III, and Long Pham. A distributed enhanced server for multidimensional scientific data. Computing in Science and Engineering, 5(2): 44–52, March/April 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/02/c2044abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/02/c2044.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/02/
                                        c2044.pdf.

Yin:2005:CAI

[YLCZ05] Xiang-Chu Yin, Zhonghua Lu, Xuebin Chi, and Huihui Zhang. The China ACESiSERVO grid node. Computing in Science and Engineering, 7(4):38–42, July/ August 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

ieeexplore.ieee.org/iel5/
5992/31456/01463134.pdf?
isnumber=31456&prod=JNL&
arnumber=1463134&arSt=+
38&ared=+42&arAuthor=Xiang-
Chu+Yin%3B+Zhonghua+Lu%
3B+Xuebin+Chi%3B+Huihui+
Zhang;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=31456&arnumber=
1463134&count=14&index=4.

Yeh:2002:ALD

[YLR02] Kao-San Yeh, Shian-Jiann Lin, and Richard B. Rood. Applying local discretization methods in the NASA finite-volume general circulation model. Computing in Science and Engineering, 4(5):49–54, September/October 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2002/05/c5049abs.htm;
 http://csdl.computer.org/
dl/mags/cs/2002/05/c5049.
htm;   http://csdl.computer.
org/dl/mags/cs/2002/05/
c5049.pdf.

Yi:2017:ISW

[YLZ17] Letian Yi, Jianbin Li, and Yaoxue Zhang. Improving the scalability of wearable devices via transparent computing technology. Computing in Science and Engineering, 19(1):29–37, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/01/mcs2017010029-
                                        abs.html.

Yang:2019:DAS

[YLZ^{+}19] Jianjin Yang, Kai Lan, Shengyang Zhu, Kaiyun Wang, Wanming Zhai, and Xuancheng Yuan. Dynamic analysis on stiffness enhancement measures of slab end for discontinuous floating slab track. Computing in Science and Engineering, 21(3):51–59, May/ June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8540034/.

Yasar:2014:CPM

[YM14] Osman Yasar and Jose Maliekal. Computational pedagogy: a modeling and simulation approach. Computing in Science and Engineering, 16(3): 78–88, May/June 2014. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Yang:2019:OFR

[YMHQ19] Y. Yang, L. Ma, D. Huang, and N. Qin. Output feedback repetitive learning control of an electrohydraulic actuator of a lower limb rehabilitation exoskeleton. Computing in Science and Engineering, 21(6):6–19, November/December 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Yamashita:2011:PCE

[YMK11] Takuzo Yamashita, Hori Muneo, and Koichi Kajiwara. Petascale computation for earthquake engineering. Computing in Science and Engineering, 13(4):44– 49, July/August 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Yasar:2006:CTA

[YMLJ06] Osman Yasar, Jose Maliekal, Leigh J. Little, and Dawn Jones. A computational technology approach to education. Computing in Science and Engineering, 8(3):76–81, May/ June 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Yasar:2000:ENP

[YRT^{+}00] Osman Yasar, Kulathur S.

Rajasethupathy, Robert E. Tuzun, R. Alan McCoy, and Joseph Harkin. Education: a new perspective on computational science education. Computing in Science and Engineering, 2(5): 74–79, September/October 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c5074.
pdf.

Yang:2002:RAV

[YWC02] Yonggao Yang, Xusheng Wang, and Jim X. Chen. Rendering avatars in virtual reality: Integrating a 3D model with 2D images. Computing in Science and Engineering, 4(1):86–91, January/February 2002. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c1086abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c1086.pdf.

You:2004:STC

[YWMM04] Donghyun You, Meng Wang, Parviz Moin, and Rajat Mittal. Study of tipclearance flow in turbomachines using large-eddy simulation. Computing in Science and Engineering, 6(6): 38–46, November/December 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/06/c6038.
htm;   http://csdl.computer.

org/dl/mags/cs/2004/06/
c6038.pdf.

Yue:2019:SES

[YYG^{+}19] Hou Yue, Song Yadong, Wu Guosong, Luo Yun, and Yao Yuan. Simulation and experimental study on active stability of high-speed trains. Computing in Science and Engineering, 21(3):72–82, May/ June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8613801/.

Yang:2018:MBC

[YYL^{+}18] Hanqing Yang, Liangzhen Yin, Qi Li, Weirong Chen, and Lijun Zhou. Multiagentbased coordination consensus algorithm for state-of-charge balance of energy storage unit. Computing in Science and Engineering, 20(2):64–77, March/April 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8013914/.

Yang:2013:MSI

[YZC^{+}13] Lin Yang, Tian-Yu Zhang, Jim X. Chen, Pei-Dong Dai, Wei Quan, Ke-Qiang Wang, and Zheng-Min Wang. Motion simulation of inner hair cell stereocilia. Computing in Science and Engineering, 15(2): 27–33, March/April 2013. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Yang:2004:RTS

[YZZ04] Yonggao Yang, Changqian Zhu, and Hua Zhang. Realtime simulation: Water droplets on glass windows. Computing in Science and Engineering, 6(4):69–73, July/ August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/cs/
                                        2004/04/c4069.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/cs/2004/04/c4069.pdf.

Zomaya:2001:NCP

[ZAF^{+}01] Albert Y. Zomaya, James A. Anderson, David B. Fogel, Gerard J. Milburn, and Grzegorz Rozenberg. Nonconventional computing paradigms in the New Millenium: a roundtable. Computing in Science and Engineering, 3(6):82–99, November/December 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c6082abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2001/pdf/c6082.pdf.

Zakai:2018:FPW

[Zak18] Alon Zakai. Fast physics on the Web using C++, JavaScript, and Emscripten. Computing in Science and Engineering, 20(1):11–19, January/February 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://ieeexplore.ieee.
org/document/8254329/.

Zebrowski:2004:NIN

[ZB04] Jan J. Zebrowski and Rafal Baranowski. Nonlinear instabilities and nonstationarity in human heart-rate variability. Computing in Science and Engineering, 6(5): 78–83, September/October 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/05/c5078.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/05/
c5078.pdf.

Zhu:1999:VCK

[ZCXM99] Ying Zhu, Jim X. Chen, Shide Xiao, and Edward B. MacMahon. Visualization corner: 3D knee modeling and biomechanical simulation. Computing in Science and Engineering, 1(4):82–87, July/August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c4082.
pdf.

Zhang:2007:CME

[ZDW^{+}07] Tian-Yu Zhang, Pei-Dong Dai, Zheng-Min Wang, Ke-Qiang Wang, Jim X. Chen, and Le Xie. A contour map of the ear’s vestibular apparatus based on 3D reconstruction. Computing in Science and Engineering, 9(1):26–31, January/February 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zebker:2000:SEI

[Zeb00] Howard A. Zebker. Studying the Earth with interferometric radar. Computing in Science and Engineering, 2(3):52–60, May/June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs2000/pdf/c3052.
pdf;   http://www.computer.
org/cse/cs1999/c3052abs.
htm.

Zeigler:2017:UPD

[Zei17] Bernard P. Zeigler. Using the parallel DEVS protocol for general robust simulation with near optimal performance. Computing in Science and Engineering, 19(3): 68–77, May/June 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs2017030068-
abs.html.

Zhang:2012:DIC

[ZFS12] Baoshe Zhang, Mirek Fatyga, and William C. Sleeman.

Designing and implementing a computing framework for image-guided radiation therapy research. Computing in Science and Engineering, 14 (4):57–68, July/August 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zhang:2017:TCP

[ZGR^{+}17] Yaoxue Zhang, Kehua Guo, Ju Ren, Yuezhi Zhou, Jianxin Wang, and Jianer Chen. Transparent computing: A promising network computing paradigm. Computing in Science and Engineering, 19(1):7–20, January/February 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/01/mcs2017010007-
abs.html.

Zhang:2011:FHP

[Zha11] Fuqing Zhang. The future of hurricane prediction. Computing in Science and Engineering, 13(1):9–12, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zhuge:2002:CSD

[Zhu02]  Hai Zhuge.  Clustering

soft-devices in the Semantic Grid. Computing in Science and Engineering, 4(6): 60–62, November/December 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2002/06/c6060.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2002/06/
                                        c6060.pdf.

Zhu:2016:SMS

[Zhu16] Ru Zhu. Speedup of micromagnetic simulations with C++ AMP on graphics processing units. Computing in Science and Engineering, 18 (4):53–59, July/August 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zhou:2008:ESS

[ZJW08] Chuan Zhou, Xiaogang Jin, and Charlie C. L. Wang. Efficient and stable simulation of cloth undergoing large rotations. Computing in Science and Engineering, 10(4): 30–40, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zhang:2009:CCU

[ZL09] Yang Zhang and Edward Luke. Concurrent composition using loci. Computing in Science and Engineering, 11(3): 27–35, May/June 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zhu:2019:EOA

[ZLTX19] Siyu Zhu, Yongle Li, Koffi Togbenou, and Tianyu Xiang. An efficient optimization algorithm to study the stochastic responses of vehicle–bridge coupling system. Computing in Science and Engineering, 21(3):6–17, May/June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8540406/.

Zhang:2019:UER

[ZLW^{+}19] H. Zhang, D. Li, T. Wang, T. Li, X. Yu, and A. Bouras. Uncertainty and equivalence relation analysis for hesitant fuzzy rough sets and their applications in classification. Computing in Science and Engineering, 21(6): 26–39, November/December 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zakharian:2003:SNF

[ZMM03] Aramais R. Zakharian, Jerome V.

Moloney, and Masud Mansuripur. Simulating near-field effects in high-density optical-disk data storage. Computing in Science and Engineering, 5(6):15–21, November/December 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
cs/2003/06/c6015abs.htm;

                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/06/c6015.
                                        pdf.

Zhang:2011:HIH

[ZQY^{+}11] Xuejin Zhang, Thiago Quirino, Kao-San Yeh, Sundararaman Gopalakrishnan, Frank Marks, Stanley Goldenberg, and Sim Aberson. HWRFx: Improving hurricane forecasts with high-resolution modeling. Computing in Science and Engineering, 13(1):13–21, January/February 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zabaras:2007:ITA

[ZS07] Nicholas Zabaras and Sethuraman Sankaran. An informationtheoretic approach to stochastic materials modeling. Computing in Science and Engineering, 9(2):30–39, March/ April 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zubairi:2023:NMN

[ZS23] Omair M. Zubairi and Micah D. Schuster. Numerical modeling of neutron stars. Computing in Science and Engineering, 25(1):67–71, January/February 2023. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zellmann:2022:PCQ

[ZSM^{+}22] Stefan Zellmann, Daniel Seifried, Nate Morrical, Ingo Wald, Will Usher, Jamie A. P. Law-Smith, Stefanie Walch-Gassner, and André Hinkenjann. Point containment queries on ray-tracing cores for AMR flow visualization. Computing in Science and Engineering, 24(2): 40–51, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zhao:2019:GEI

[ZW19] Xin Zhao and S. C. Wu. Guest Editors’ introduction. Computing in Science and Engineering, 21(3):4–5, May/June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zhao:2020:TTL

[ZWYL20] P. Zhao, G. Wu, S. Yao, and H. Liu. A transductive transfer learning approach based on manifold learning. Computing in Science and Engineering, 22(1):77–87, January/February 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zeinalipour-Yazti:2004:IRT

[ZYKG04] Demetrios Zeinalipour-Yazti, Vana Kalogeraki, and Dimitrios Gunopulos. Information retrieval techniques for peerto-peer networks. Computing in Science and Engineering, 6(4):20–26, July/August 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/04/c4020.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/04/
c4020.pdf.

Zhang:2019:FSR

[ZZC^{+}19] W. Zhang, F. Zhang, W. Chen, Y. Jiang, and D. Song. Fault state recognition of rolling bearing based fully convolutional network. Computing in Science and Engineering, 21(5):55–63, September/October 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Zhang:2006:UGA

[ZZPC06] Hua Zhang, Changqian Zhu, Qiang Peng, and Jim X. Chen. Using geometric algebra for 3D linear transformations. Computing in Science and Engineering, 8(3): 68–75, May/June 2006. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Zeng:2019:RPS

[ZZS^{+}19] Yuanchen Zeng, Weihua Zhang, Dongli Song, Zhenchen Chang, and Haifeng Zhang. Response prediction of stochastic dynamics by neural network: Theory and application on railway vehicle. Computing in Science and Engineering, 21(3):18–30, May/ June 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8601360/.

Zhe-Zhao:2006:NIB

[ZZYNH06] Zeng Zhe-Zhao, Wang Yao-Nan, and Wen Hui. Numerical integration based on a neural network algorithm. Computing in Science and Engineering, 8(4):42–48, July/August 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).